A Complete Bibliography of Publications in Statistical Papers

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

08 October 2025 Version 1.60

Title word cross-reference

(k, l) [1885]. (n − k + 1) [1753, 1623]. 0 [772]. 0 < ρ < ρ_{0} [2026]. 1 [772, 548]. $109.95 [2218]. $125.00 [2161]. $129.95 [2259]. 1 ≤ p < ∞ [562]. 2 [616, 407]. $229.00 [2019]. 2^{K} [2597, 2513]. 2^{m}4^{1} [1398]. 2^{m}4^{n} [1476]. 2^{s−k} [2039]. 3 [852, 2125]. 40m [2502]. $47.96 [2556]. $49.95 [2161]. $59.95 [2219]. $89.95 [2217]. ^{(R)} [2002, 1962]. _{1} [1276]. _{2} [1417]. _{k} [798]. A [1577, 2765, 2852, 2312, 2346]. α [858, 1840, 2242, 1465, 1271, 2418, 1945, 2327, 1926]. AR(1) [1650, 790]. AR(p) [1346]. R̄^{2} [652]. β [668, 856, 497]. c [2332]. χ^{2} [919]. D [2880, 2135, 2268, 2569, 2567, 836, 948, 1324, 2309, 2758, 2346, 2098, 2176, 2923]. δ [1959]. e [596]. E(x ⊗ xx′) [1805]. E(xx′ ⊗ xx′)

[1805]. F ; [1901, 915, 1361, 407, 681, 2599, 2543]. G ; [548]. Γ [722]. H [1664]. I(1) [845]. K ; [2308, 1551, 1223, 2725, 1837, 368, 479, 2177, ; 2968, 3069, 2410, 2502, 1322, 1198, 2224, ; 2309, 407, 1932]. k − 2 [1781]. L [2926, 439, ; 1974, 2541, 2516, 1584, 1399, 2502, 1322]. L^{p} ; [2984]. L_{1} [895, 2992]. L_{2} ; [2195, 2363, 2848, 2039]. L_{P} [562]. L_{q} [2729]. ; LM [1343]. LR [1343]. M [1054, 1107, 2465, ; 2445, 2570, 2094, 2048, 2398, 2180, 2602]. L_{p} ; [1552]. N [816, 752, 48, 1753, 1623, 548]. ; N ≡ 0(mod9) [2592]. p [1245, 3064, 2794, ; 1461, 2378, 2880, 1794, 1868, 2961, 1604, ; 1791, 1589, 2074, 2227, 3003]. P (X < Y) ; [1867, 1492, 1955]. P (Y < X) [2222, 1380]. Φ ; [1345, 3012]. P (X > Y) [2425]. ; Pr(a < x < b) [263]. q [2416, 1651, 2585]. R ; [2483, 1238, 1848, 616, 1932]. r − k

[2082, 1480]. R/S [1996]. R = Pr(X > Y) [2007]. r^{th} [1446]. ρ [3059]. ρ^{∗} [1884]. S [2151, 1105]. S^{2} [848, 790, 909]. σ [1141, 684]. SUN_{n,2} [2046]. t [2737, 1054, 1777, 327, 2924, 1533, 1722, 2944, 982, 856, 1031, 1780, 3048, 2732, 1559, 1094, 1391, 1342, 1680]. T^{2} [1377, 1967, 2966]. τ [3027]. U [2469, 2427, 2748, 2745, 2123, 2648, 2457, 2951]. φ [1116]. k [2195]. W [1343]. X [1105, 1037, 1564, 1611, 1805]. X = (x_{ij}) [1577]. X^{2} [1334]. x_{ij} = −1, 0, 1 [1577].

-asymptotic [2602]. -Boxplots [2224]. -charts [2332, 1238]. -class [1198]. -coherence [684]. -component [2195]. -content [668]. -control [1105]. -coordinate [1564]. -coordinates [1611]. -copula [1722]. -dimensional [2309]. -dimensionalen [48]. -discrepancy [2039, 2363]. -distribution [2737, 2944, 2046, 439, 836, 681]. -distributions [1054]. -disturbances [790]. -divergence [3012, 2151, 1345]. -divergences [1116]. -empirical [2123]. -estimates [1054]. -estimation [1107, 2445, 2398]. -estimator [2570, 2048]. -expectation [856]. -factor [2308]. -functionals [1399]. -GARCH [2502]. -Laplace [1465, 1271]. -level [2585]. -likelihood [1664, 2729]. -lists [2968, 3069]. -Medians [368]. -minimax [722]. -mixing [1840, 2418, 1884, 3059, 1926]. -model [982]. -moment [2543]. -negatively [2180]. -norm [2992]. -norms [562]. -observations [1037]. -optimal [2427, 2880, 2135, 1577, 2268, 2569, 2483, 2567, 1324, 2765, 2309, 2758, 2852, 2312, 2346, 2098, 2176, 2725]. -out-of- [1753, 1623, 407, 548]. -overlapping [1322]. -parameter [852]. -policy [816]. -populations [1461]. -processes [2457]. -record [1837]. -shock [1959]. -statistic [2926, 2748, 2516]. -statistics [1974, 2648, 2469, 2541, 2745, 2951].

-step-ahead [2880]. -sub-fractional [2242]. -test [919, 915]. -th [1848]. -type [2094]. -unimodality [858]. -value [2794, 1604, 1791]. -values [1245, 3064, 2961]. -version [895]. -Weibull [2416, 1651].

/1 [798].

1 [816, 2019, 798]. 10.1007/s00362 [1728]. 10.1007/s00362-007-0113-0 [1728]. 117 [787]. 13 [2757]. 17th [2105]. 1963 [93]. 1963/64 [93]. 1965 [63, 66]. 1966 [147, 155, 163]. 1967 [293]. 1969 [166]. 1972 [264]. 1975 [364].

2 [230, 2147, 2112, 2259, 2160]. 2015 [2116]. 21 [480]. 2nd [2349, 1685, 1770]. 2SHI [1094].

3 [2182, 2199, 833]. 3-component [2940]. 36 [912]. 3rd [2055, 1725].

4/SP06 [1371]. 425pp [2019]. 46 [1607]. 4th [1526].

51 [1728].

6 [2218, 2556, 799]. 64 [93].

7 [2198, 2219].

8 [2217, 2161]. 80th [1570].

9 [2675]. 978

[2675, 2019, 2198, 2147, 2219, 2112, 2218, 2259, 2182, 2217, 2199, 2556, 2160, 2161]. 978-0-12-803768-3 [2182]. 978-0-19-872646-3 [2199]. 978-0-19-872906-8 [2161]. 978-0-19-873691-2 [2112]. 978-1-138-59146-2 [2259]. 978-1-138-60079-9 [2675]. 978-1-138-63197-7 [2198].

978-1-1387-1113-6 [2556]. 978-1-4987-1523-2 [2160]. 978-1-4987-1956-8 [2217]. 978-1-4987-3422-6 [2218]. 978-1-4987-6974-7 [2219]. 978-1-4987-7505-2 [2147]. 978-1-4987-7571-7 [2197]. 978-3-642-34467-1 [2019].

A. [1472, 1544]. aberration [1305, 1347, 1398, 1420]. aberrations [2307]. Abgrenzung [131]. Abhängigkeit [539]. Abraham [18]. Abschätzung [610]. Abschätzungen [579]. absence [1147]. absolut [302]. absolute [1695, 1207, 1309, 2989]. absolutely [1407]. absorbing [1898, 2907]. Academic [2182]. Accelerated [2751, 990, 2960, 2073, 1692, 2622, 1953]. Acceptance [1609, 647, 2604]. accommodate [2915]. according [1156, 505]. account [1886, 420]. Accounting [2763, 1367, 1537]. accounts [468, 130, 468]. Accuracy [1517, 947, 1484, 3024, 2683]. Acknowledgement [1337]. across [3016]. Active [2886, 2770]. Active-set [2886]. actuarial [2949]. acyclic [2488]. Adademie [265]. adapt [2842]. adaptation [2980]. Adaptive [2807, 622, 2291, 2883, 2593, 484, 2300, 2767, 2353, 2897, 2643, 2771, 2794, 1083, 1799, 2880, 2073, 2527, 2394, 2519, 2138, 1072, 2071, 2947, 536]. adaptive-to-model [2947]. Adäquation [565]. Adäquationsproze [551]. addendum [661]. additional [630, 755, 1575, 2923, 1654]. Additionsproblem [233]. Additive [1824, 2964, 1431, 1954, 1141, 2712, 3065, 2896, 2503, 2517, 3002, 3051, 2622, 2861, 2714, 2884, 1792, 2450, 2343, 2370, 2154, 2414]. additive-multiplicative [2517]. additives [572]. Additivität [378]. additivity [742, 1587]. adequacy [459, 2583]. adequations [551]. ADF [2382]. Adjusted [1958, 2626, 711, 2519, 2858, 2996]. adjusting [2053]. adjustment [1366, 2649, 635, 818, 2708]. Admissibility [1403, 1211, 1582]. Admissible [2574, 1407, 2719, 2150, 797, 1249, 728]. adolescents [688]. advanced [1681]. Advances [2299]. advantages [1053]. Affiliation [2651]. affine [2316, 2028, 1236, 1212]. Affinitätsmodell [491]. Afshin [2367]. after [2092, 1971, 1848, 1745, 2539]. again [1903]. against [2335, 791, 3068, 2636, 1672, 897, 2766, 1225]. age [2208]. ageing [898, 2322]. Aggarwal [2089]. agglomeratives [452]. aggregate [1775]. Aggregated [2838]. Aggregaten [268]. aggregation [906, 803, 672, 411, 785, 1400, 450, 73]. Aggregationsproblem [39]. aging [1952, 1295, 1430, 1498]. Agresti [2107]. Agung [1706]. ahead [2880, 860]. air [2271]. Ajit [1771]. Akaike [835]. Akkomodation [35]. Aktionen [658]. Aktualisierung [571]. Aktuelle [224, 603]. al [1523]. Alan [2218, 1832, 2107, 1723]. Albert [293, 1661, 1683]. Alex [2144, 1771]. Alexandre [2162]. Alfredo [2162]. Algebra [2111, 2331]. Algebraische [883]. Algorithm [3061, 711, 2195, 1019, 2010, 2880, 1760, 334, 2380, 2844, 2726, 2909, 2982, 2886, 2713, 965, 1922, 763, 2709]. algorithmic [1943]. algorithms [1572, 647, 2568, 2089, 2124]. Alkire [2106]. Alleviating [2881]. Allgemeine [15]. allgemeinen [88]. allgemeines [532, 22]. allocation [964, 608]. allocations [860]. allometric [1960, 3061]. allow [1879]. Allphasen [255]. Allphasen-Umsatzsteuer [255]. Almon [589]. Almost [2043, 2364, 1132, 870, 1883, 1807, 2283]. along [1973]. alten [336]. Alternative [1838, 1253, 1653, 1053, 2694, 1794, 1156, 1581, 1361, 2882, 2893, 1563]. alternatives [915, 3045, 2772, 667, 2636, 2792, 1670, 1115]. Altersstruktur [81]. Altersverteilung [115]. Altertum [3]. Alzheimer [1902]. am [337, 192, 437, 130]. ambiguity [1140]. ambiguous [1149]. Ameet [2367]. Amerika [249]. among [643, 1678]. amplitude [2673]. Amputation [1460]. Amtlichen [272, 64, 92]. analogue [172]. Analyse [571, 678, 206, 613, 642, 480, 28, 2053, 1501, 41, 36, 508]. Analyses [271, 2730, 2155]. Analysis [864, 1644, 2299, 2218, 1915, 2889, 1706, 2217, 1586, 2270, 2160, 3074, 1822, 2186, 2903, 3054, 2693, 771, 1260, 816, 602, 620, 2406, 1574, 2155, 1764, 1496, 1034, 1927, 2776, 2816, 1073, 1184, 1664, 597, 1624, 708, 835, 2985, 2206, 2447, 881, 2389, 693, 286, 1155, 2349, 429, 814, 1458, 804, 1755, 1743, 896, 1493, 1555, 1727, 199, 1580, 793, 1630, 1714, 2215, 1789, 2900, 2732, 2998, 2959, 2802, 1681, 1962, 1248, 1009, 1857, 27, 1371, 2096, 2352, 2496, 2125, 2433, 1547, 407, 128, 393, 1199, 2110, 709, 1673, 1953, 1739, 2318, 2676, 2855, 2486, 1922, 789, 1443]. analysis [2044, 2884, 1400, 2757, 3075, 2935, 2857, 2877, 2778, 2162, 1705, 1660, 2146, 2106, 2674, 2107, 1985, 1748]. Analytic [2265, 1333, 2555]. Analytical [720]. analytics [2088]. analytisch [246]. analytisch-numerisches [246]. analyzers [2600]. Analyzing [2940, 1992, 2038, 1645, 1665]. and/or [1650]. andere [498]. Andrew [971, 2556, 1727]. Andrews [1510]. Andy [2161]. Angaben [281]. Angel [1962, 1906]. Angewandte [912]. Angrist [1708, 2104]. angular [2276]. Anirban [1662]. Anmerkung [596, 235, 555]. Anmerkungen [451, 165]. Announcement [472, 1082]. annual [522]. Anova [1155, 1980, 915, 1927, 1354, 1470, 2137, 1385, 2038, 2792, 1587, 2848, 1589, 324, 325, 326]. ANOVA-type [1980]. Ansatz [425, 576].

Ansätze [188]. Ansgar [1666].

Antedependence [1870]. antiquity [3]. Antón [1870]. Antony [1546]. Antwort [127]. Anwendbarkeit [315]. Anwendung [533, 631, 516, 377, 432, 417, 452, 476, 431, 451, 426]. anwendungen [883]. any [157]. AOQL [929, 1440]. aperture [2777]. Appendage [2576]. Application [421, 2287, 854, 2183, 2831, 1158, 2317, 1957, 2553, 1012, 2791, 3064, 2652, 2964, 2733, 1935, 1722, 2465, 2157, 1594, 2769, 3067, 1128, 2491, 334, 1475, 1332, 2119, 2718, 2837, 2339, 1375, 2966, 2899, 1841, 2457, 352, 21, 3046, 2990, 2067, 2543, 2867, 1896, 979, 2989, 2723, 2563]. Applications [2675, 445, 2243, 1568, 3074, 1401, 1736, 1728, 2272, 2293, 2975, 1837, 1139, 1104, 1842, 2776, 1593, 1019, 1185, 878, 1043, 2660, 2670, 2695, 1938, 3015, 836, 1018, 2598, 1356, 2541, 1458, 1981, 2177, 2002, 1430, 1716, 2419, 866, 3072, 2140, 2984, 3000, 1656, 1729, 2413, 2995, 2189, 2180, 2269, 1864, 1421, 1444, 1795, 2444, 2601, 3043, 1873, 2602, 2736, 2468, 2935, 2798, 2949, 2680, 2162, 2217, 2108, 1663, 2113, 2056, 2089]. Applied [3075, 226, 1073, 361, 2145, 2777, 2359, 340, 2218, 2432, 1726, 2433, 1891]. Applying [2730]. approach [924, 1728, 1552, 1573, 553, 2012, 2626, 410, 941, 1366, 2169, 2515, 2913, 2885, 2375, 2091, 2157, 1593, 1886, 2904, 2412, 2670, 2151, 828, 3020, 1332, 822, 1530, 2038, 2446, 3002, 2998, 2764, 2516, 2802, 1556, 751, 2360, 2036, 1791, 2843, 2735, 2882, 1808, 1966, 2066, 1433, 2236, 2075, 3010, 2919, 2095, 2932, 1525]. approaches [1762, 1734, 1257, 2137, 1156, 1681, 43, 1749]. appropriate [1702]. Approximate [1860, 2513, 1809, 2659, 979, 695, 1114, 947, 923, 1530, 1376, 1517, 2128, 709, 680, 869, 3010]. Approximation [356, 1007, 1854, 1235, 1484, 1712, 2479, 3011, 1196, 191]. Approximations [2263, 2783, 2965, 1223, 1798]. approximativen [506]. arbeitenden [376]. Arbeiter [93]. Arbeitsgegenstände [74]. Arbeitskräfte [83]. Arbeitslosen [509]. Arbeitslosenquoten [225]. Arbeitsmittel [74]. Arbeitsproduktivität [220]. Arbeitsteilung [164]. arbitrarily [1658]. arbitrary [1103, 993, 2796, 548, 1086]. ARCH [2520, 1716, 2536, 1285]. ARCH-type [1285]. Archimedean [1782]. area [2292, 2333, 2522, 2348, 2584, 1950]. area-specific [2292]. arguments [2828, 404]. ARIMA [905, 654]. arising [1837, 1559]. arithmetic [2724]. ARL [1836, 1064, 2332]. ARL-unbiased [2332]. ARMA [1340, 1530, 1196, 1262]. armed [1648, 1299]. arranged [730]. arrays [2592, 2307, 2624, 2657, 3050, 2979]. arrival [2275]. Art [185]. ARTICLE [2746]. artificial [1936]. Arup [2259]. ASN [1169, 1769]. ASN-minimax [1169, 1769]. aspects [767, 672, 503, 2836, 1058]. Aspekte [595, 606, 1503]. Aspekten [240]. Assemblage [69]. Assessing [2779, 2154, 241]. asset [2053, 1897, 1558, 2296]. Assistance [265]. Associate [1906]. associated [1103, 629, 2747, 3067, 1336, 1542, 2180, 2440, 2602]. Association [472, 1402, 1675, 559, 2247, 2413]. assumption [1418, 2881, 2067, 2154]. assumptions [2924, 1680]. asymmetric [2207, 2452, 2822, 1142, 2258, 2665, 1681, 2363, 956, 2860, 826, 1803, 1951]. asymmetrical [2473]. asymmetry [1852, 2573, 2710, 2192, 1826, 1880]. Asymptotic [1978, 1662, 891, 1063, 2520, 2179, 692, 704, 1690, 1369, 881, 846, 2090, 1561, 2057, 866, 1845, 1223, 2953, 1456, 2340, 2676, 3042, 2301, 2497, 390, 2643, 2382, 2035, 764, 1306, 1974, 1720, 832, 767, 2636, 790, 775, 2196, 943, 1997, 2766, 2510, 909, 1387, 2204, 2381, 2074, 2602, 3039, 1996, 2778]. Asymptotically [1454, 242, 749, 3067, 2612, 1608].

Asymptotics [2570, 1333, 2269, 805]. asymptotische [362, 181]. Ätialität [532, 565]. ATLA [2733]. atmospheric [2274]. Attempting [3]. attribute [2212]. auditing [2626]. Aufbaus [183]. Aufbereiten [199]. aufgrund [370]. Aufsatz [273]. augmentation [1673]. augmented [2654, 2494]. augmented-estimating-equations [2494]. Ausbaustufe [246]. Ausbildungsförderung [369]. Ausgabe [558, 47, 57, 68, 77, 86, 95, 102, 111, 119, 126, 134, 142, 150, 159, 169, 178, 187, 194, 202, 213, 223, 239, 248, 257, 267, 276, 285, 295, 304, 311, 322, 333, 342, 351, 358, 374, 384, 392, 409, 419, 428, 436, 447, 458, 466, 474, 482, 493, 502, 511, 520, 535, 542, 550, 567, 583, 593, 401]. ausgewählte [467]. ausgewählter [642]. Ausmaß [48]. Ausprägungen [161]. außenwirtschaftlichen [143]. Auswahl [362, 96, 104]. Auswertung [377, 432, 431]. Auswirkung [678, 217]. Auswirkungen [208, 281, 367, 211, 65]. Auszug [176]. auto [1743]. auto-multivariate [1743]. autocorrelated [359, 2082, 755, 1998, 641, 353, 382, 2098, 1387, 1689]. autocorrelation [2193, 1650, 845]. autocorrelations [925]. autocovariance [461]. autocovariances [2501, 1378]. autokorrelierten [642]. Automaten [136]. Automatic [1731]. Automobilindustrie [218, 230, 244]. autoprojektiver [345]. autoregression [2304, 1307, 3068]. autoregressions [2389]. Autoregressive [2351, 2193, 415, 2355, 3041, 2532, 2964, 2943, 749, 2993, 2682, 2985, 1414, 828, 1227, 1479, 611, 1915, 1271, 2647, 2845, 3060, 3051, 1877, 2502, 2076, 2289, 2861, 2460, 2266, 2886, 1825, 2190, 2475, 2102, 1485, 2288, 1196, 2359, 2963, 2978, 2401, 2933, 3034, 2443, 2254, 2640, 3007, 461]. auxiliary [1318, 811, 1132, 1565, 1866, 2583, 1038, 1463]. availability [548]. available [2172, 1654]. Avant [1]. Avant-propos [1]. average [2918, 390, 415, 2832, 2724, 1696, 1296, 2874, 2812, 1775, 1698, 861, 1275, 1300, 2396, 2029, 461]. averages [2037]. averaging [1829, 2493, 785, 2623, 2631, 3039, 2471]. Axiomatic [450, 612]. Axiomatik [612]. Axiomensystems [444]. axioms [810]. Aylin [2144].

B [757, 1486, 2145, 780, 789, 440, 3028, 2337, 2170, 3002, 1801, 2631]. B-spline [3028, 2337, 3002, 1801, 2631]. background [2735]. Bahadur [2419, 2374, 1926]. Bailer [1665]. Balakrishnan [1768, 1499]. balance [1347, 2098, 2060]. balanced [1447, 2303, 2576, 2592, 216, 1264, 1755, 2719, 1249, 2784, 1677, 2168]. balancing [2807, 2955, 897]. Balk [661]. ball [2309, 2758]. Ballon [2106]. Band [833, 480]. Bandbreite [385]. bandit [1299]. bands [1905, 1815, 2453]. bandwidth [1259, 846]. Banerjee [574]. Bank [632, 632]. Bankensektor [184]. Bankensektors [585]. Bartlett [1610, 2858, 2996]. Bartlett-type [1610]. base [1875, 751]. Baseball [2555]. based [2226, 1365, 1812, 1718, 2058, 2364, 2940, 2402, 2643, 1464, 2467, 895, 997, 2234, 1928, 2228, 2707, 1067, 2021, 2013, 2321, 2931, 2902, 2012, 1412, 2438, 2667, 1894, 3018, 2578, 2512, 2262, 2552, 2604, 891, 1114, 2083, 2690, 2913, 946, 1419, 3045, 2425, 2727, 1565, 1867, 1886, 991, 2738, 2790, 1853, 2233, 1383, 2412, 2879, 2993, 1970, 2220, 1690, 1096, 2568, 2451, 2135, 152, 1827, 2695, 2185, 2817, 2447, 1348, 2608, 2390, 1975, 2748, 1872, 2598, 982, 2023, 2710, 2415, 2255, 2213, 2386, 2627, 2968, 3069, 2192, 1672, 2022, 2165, 2501, 1321, 1912, 2718, 2752, 2853]. based [1065, 1789, 1358, 2732, 2422, 2315, 2827, 2164, 1983, 1756, 2122, 2740, 1375, 2123, 1844, 2479, 2946, 3011, 2034, 1857, 753, 2777, 2886, 2235, 1292, 2648, 2141, 1655, 2954, 2360, 2036, 1791, 1237, 2992, 2711, 2127, 2311, 2843, 2529, 2735, 1796, 1759, 2860, 2347, 1494, 2048, 1208, 1056, 2929, 2318, 2510, 2801, 2781, 1085, 2848, 1618, 1859, 2350, 2506, 1922, 1823, 2267, 1515, 2149, 2302, 2476, 2951, 1098, 2994, 3001, 2221, 2254, 2343, 2798, 2095, 2761, 1917, 2384, 2723, 2011, 1508, 2947, 2344]. Basic [1052, 1072, 805]. basics [440, 2111]. basing [695]. Basis [290, 2668, 1946, 1275, 1300]. Basiswissen [1666]. Basketball [2555, 2675]. Basu [2151]. bathtub [1634]. Baumgartner [2479, 1292]. Bayes [1367, 1474, 404, 2782, 1561, 1602, 1332, 2881, 2312, 794, 1515, 1486]. Bayesian [1504, 1547, 1891, 1684, 2186, 1365, 924, 2940, 3054, 1552, 1030, 2975, 3064, 1635, 359, 1764, 1435, 1246, 430, 2222, 2351, 2091, 1550, 2435, 1853, 2818, 1624, 2660, 2670, 835, 2985, 197, 1414, 1257, 1247, 860, 3015, 3012, 2795, 981, 1454, 729, 1755, 1767, 1332, 1982, 1380, 1633, 1671, 2255, 2527, 1630, 2290, 1682, 3013, 2502, 2786, 776, 1375, 1816, 3026, 1186, 1955, 1309, 1683, 2852, 2238, 1673, 1953, 2287, 2066, 2188, 1086, 970, 1433, 2587, 3034, 3033, 1097, 1400, 1591, 2640, 2605, 2334]. Bayesian-motivated [2587]. Bayeslösungen [645]. Bayes’sche [16, 659, 307]. Beale [334]. Bearbeitung [96, 104]. Bécue [2368]. Begriff [366, 2]. Begriffe [273]. begründete [262]. Begründung [614]. Behandlung [123]. behavior [1959, 2016, 859, 3042, 1996]. behaviors [2986]. behaviour [764, 2635, 767, 866]. Behrens [1158, 711, 172]. beiden [92]. beim [44, 539]. beiökonomischen [442]. Beispiel [337, 130]. Beispiele [291]. Beispielen [371]. Beitrag [514, 521, 171, 344, 515, 224]. Beiträgen [912]. Beitragsmöglichkeiten [28]. BEKK [2786]. Belastung [255]. beliefs [1147]. Bellman [18]. Belzunce [2182]. Bemerkung [471, 388]. Bemerkungen [563, 319, 298, 441, 536, 480, 437, 347, 167, 515, 273, 210]. Benchmark [1725]. Bender [211]. benefit [1796].

Benford [2056, 3072]. Benini [1904]. berechnet [258]. Berechnung [72, 9, 601, 220, 467]. Berechnungsmethodik [81]. Bereinigung [148]. Bericht [440, 6]. Berichts [165]. Berk [1747]. Berlin [1666]. Bernhard [1705]. Bernoulli [2667, 1299, 782]. Bernoullinutzens [167]. Bernstein [3046, 2997]. Berry [2550]. Bertaut [2368]. Berücksichtigung [81, 96, 104]. Beruf [165]. beschreibenden [291]. Beschreibung [456]. besonderer [96, 104]. Besonderheiten [72]. Bessel [2673]. Best [497, 540, 562, 680, 1573, 242, 729, 1212, 1931, 2256, 2060]. Bestandsaufnahme [112]. Bestimmung [658, 185, 217]. Bestrebungen [536]. Beta [1613, 1831, 1534, 2791, 2356, 1299, 2821, 1842, 2670, 1862, 653, 1676, 1344, 1841, 1598, 1361, 1182, 2729, 1449]. beta-binomial [2821]. beta-normal [1862]. beta-prime [1534]. betas [2230]. Bethlehem [2219]. Betrachtungen [365, 15]. betreffend [399]. Betriebs [203]. Betriebs- [203]. Betriebsökonometrie [349]. better [2257, 980]. betting [1147]. between [2054, 2081, 1600, 3004, 2620, 2876, 878, 686, 1941, 2033, 2644, 2833, 2710, 803, 848, 1057, 1885, 3025, 2927, 1778, 1420, 2413, 2312, 1966, 2650, 2939, 779, 2440, 1925, 1174, 2770, 1758]. Beurteilung [241]. Bevölkerungsprozeß [626]. Bevölkerungsstruktur [626]. Bewertung [545, 236, 2055, 627]. beyond [2774, 2893]. Beziehungen [372]. Bezugssysteme [261]. BGS [2735]. bi [822]. bi-variate [822]. Bias [2209, 1735, 1456, 1260, 669, 2078, 1949, 2708, 2955, 2642, 1910, 2691, 1133, 2971, 1387, 1346, 1858]. Bias-corrected [2209, 1949]. Bias-variance [1735]. biased [838, 1258, 1785, 2517, 2124, 1494, 779]. biasing [889]. big [2764, 2824, 2701]. Bilanz [315]. bilateral [2906]. Bildungsstatistik [208, 188]. bildungsstatistische [208, 173].

Bildungswesens [71]. bilinear

[2914, 286, 2475]. bimodal [2898]. binäre [479]. Binary [826, 1787, 863, 1073, 1495, 1658, 1881, 2314, 2740, 2124, 897, 2758, 2523, 2405]. binomial [838, 924, 2282, 1159, 1435, 2821, 1495, 2666, 1185, 2993, 2177, 2240, 1322, 1838, 2350, 2506, 237, 1174, 1873, 2489, 2254, 2800]. binomial-Poisson [1159]. binormal [2385]. biologists [2161]. Biometrische [96, 104]. biparametric [1088]. Birkenfelds [399]. Birnbaum [1738, 2359]. birth [1336]. Birthday [1570]. bis [293, 63, 3]. Bivariate [933, 2387, 1675, 1348, 2782, 1465, 2377, 2336, 1428, 2831, 630, 1742, 1395, 2638, 1462, 2743, 1722, 1852, 2699, 1538, 1639, 3037, 867, 2850, 1819, 1956, 1977, 1676, 934, 958, 1024, 1509, 1657, 2314, 2941, 899, 1733, 2213, 2745, 2434, 2777, 2086, 2099, 2475, 1880, 1940, 2043, 3071, 2216, 815, 666, 2967, 2334]. black [1687]. Blackjack [219]. Bleymüller [2105]. blind [1916]. Block [2261, 2396, 2303, 2576, 216, 1576, 2304, 2744, 1579, 2704, 1581, 1248, 2886, 1861, 3035, 2678, 2357]. blockcollinearity [189]. blocked [2599]. blocks [1353]. BLUE [1941, 1520, 2766]. BLUEs [2077]. BLUP [1941, 1183]. BLUPs [2590, 2408]. Bonferroni [1354]. bonus [1550]. Book [771, 1704, 1027, 381, 1749, 757, 1707, 1984, 2367, 1150, 1048, 1660, 1664, 626, 682, 746, 1486, 1643, 1525, 1663, 2331, 2675, 1189, 211, 1202, 1834, 1870, 2019, 2113, 2144, 2198, 2147, 2219, 1567, 1708, 1039, 1079, 950, 2112, 2146, 2218, 2349, 1099, 1069, 1110, 2259, 2239, 1661, 1381, 778, 1136, 1217, 1230, 2182, 2197, 960, 971, 1015, 1059, 1727, 1703, 1746, 1963, 2109, 2104, 2106, 2108, 2432, 2674, 754, 994, 349, 1526, 1645, 1682, 1685, 1771, 2002, 2072, 2107, 2555, 1545, 2111, 2145, 1705, 1706, 1985, 1242, 1267, 1289, 1302, 1315, 1327, 1392, 1546, 1642, 1665, 1681, 1770, 1890, 1983, 1962]. Book [2089, 2088, 2217, 2431, 2554, 1833, 1943, 2003, 265, 2199, 833, 1726, 2433, 2556, 1124, 1425, 1437, 1504, 1547, 1566, 1683, 1686, 1687, 1891, 1869, 1503, 2368, 2238, 2298, 1487, 464, 725, 1351, 1410, 2110, 912, 733, 1089, 1568, 1450, 1279, 799, 1725, 805, 1254, 780, 789, 1176, 787, 783, 1684, 2160, 1524, 1338, 1372, 1748, 1813, 1162, 1724, 2105, 2161, 1005, 2162, 1545]. Books [724, 750]. Bootstrap [2614, 1597, 741, 2193, 1353, 1901, 1971, 764, 2865, 1065, 2554, 690, 2063, 1346, 807, 2951]. bootstrapped [2469, 993, 3042]. Bootstrapping [2915, 2539]. Bose [2259]. both [863]. Bottle [2144]. bottom [2968, 3069]. bottom- [2968, 3069]. bound [2473, 1308, 2281, 1778, 2980, 1174]. boundaries [1253]. boundary [2263, 2565]. Bounded [3055, 359, 2403, 2542, 2728, 1601]. Bounds [1335, 2626, 2863, 1490, 679, 2798, 858, 2078, 2438, 2862, 2550, 2572, 777, 2448, 2039, 2922, 2260, 2973, 2269, 3023]. Box [1652]. Boxplots [2224]. BP [2147]. branching [955]. BRD [281, 211, 166, 258, 289, 272, 509, 312, 315]. BRD-Japan [272]. break [2609, 1720, 2642, 1783]. breakdown [1107]. breaking [2292]. breaks [1894, 1900, 1192, 1282, 2531]. breast [2831]. Bregman [2311]. Breitung [2899]. Bretz [1771]. bridge [2828, 1982, 2837]. bridges [2242]. brief [3063]. Broemeling [2238]. Brownian [2463]. Bruce [2674]. Bruno [783]. Brutto [255, 54]. Brutto- [54]. Bruttoproduktionswerte [148]. bubbles [1894]. Buchbesprechung [82, 274, 299, 132]. Buchbesprechungen [662, 670, 175, 702, 689, 696, 713, 707]. Bücherbesprechung [487]. Budapest [265, 264]. Budgetvergleich [8]. build [2096]. Bundes [131, 293]. Bundes- [131]. Bundesregierung [165]. Bundesrepublik [147, 155, 467, 480, 130, 71]. bundled [2895]. Burr [1736, 2321, 1603, 470, 1475, 1380, 1186, 1876]. Burr-type [1380]. burst [1894]. business [1834, 2108, 503, 1726]. buying [2129]. bzw [117].

C [1660, 1834, 2147, 2218, 2182, 544, 1771, 2145, 2089, 799, 1774, 2430]. C-optimale [544]. C. [2110]. CADF [2074]. calcium [2563]. calculate [334]. calculating [1312]. calculation [279, 601, 1612, 2533]. Calculations [1685]. calculus [2108, 2873]. Calendar [633, 640, 646, 663, 671, 683, 703, 714, 726, 747, 769, 792, 800, 812, 819, 831, 841, 874, 884, 893, 902, 911, 921, 930, 939, 951, 961, 972, 985, 995, 1016, 1028, 1040, 1049, 1060, 1070, 1080, 1090, 1100, 1111, 1125, 1137, 1151, 1163, 1177, 1190, 1203, 1218, 1231, 1243, 1255, 1268, 1290, 1328, 1426]. Calender [1280, 1303]. calibrating [3064]. Calibration [1858, 1934, 2848, 1713, 2746, 2014]. Caliński [1570]. call [633, 640, 646, 663, 671, 683, 714, 726, 747, 769, 792, 800, 812, 819, 831, 841, 874, 884, 893, 902, 911, 921, 930, 939, 951, 961, 972, 985, 995, 1016, 1028, 1040, 1049, 1060, 1070, 1080, 1090, 1100, 1111, 1125, 1137, 1151, 1163, 1177, 1190, 1203, 1218, 1231, 1243, 1255, 1268, 1280, 1290, 1303, 1328, 1426]. calving [2093]. Cam [805]. camera [2125]. Campbell [971, 1526]. Canadian [503]. cancer [2831, 2553, 1841]. canonical [2732]. capability [1677]. capacities [1148, 1146]. capital [2296]. capture [2878]. capture-recapture [2878]. cardinal [450]. cards [1588]. Carlo [2844, 1184, 2819, 260, 2271, 1170, 1033]. Carnap [404]. cas [586]. Casals [2162]. case [586, 1897, 359, 3018, 976, 2446, 2912, 1431, 1779, 1105, 2066, 818]. case-cohort [1431]. case-control [1779]. categorial [765]. categorical [771, 1995, 1669, 2107]. categories [1077]. Cauchy [1889, 2808, 2828, 1010, 2995]. causality [1921, 2389, 611]. caused [2435]. cautionary [1248]. CDF [1560]. CE [1274]. Celebration [1570]. cell [729]. censored [3054, 648, 1928, 2081, 2321, 2656, 242, 946, 2209, 1853, 2233, 2879, 1467, 2143, 1128, 1257, 1348, 2073, 1754, 1948, 1647, 1633, 1620, 2386, 2527, 513, 2158, 2522, 3048, 2339, 808, 1186, 2000, 1517, 2346, 477, 1481, 1859, 1515, 2600, 979, 2857, 2868, 1958]. censoring [821, 410, 2411, 2960, 2220, 2337, 934, 2530, 1671, 1692, 2892, 2305, 2247, 1876, 2346, 1535, 2030, 2201, 1609, 314, 1625, 2024, 2948, 2511]. censorship [1416]. center [2727, 1555]. center-outward [2727]. centered [2039]. central [2580, 1676, 1973, 681]. Centre [2822]. Centre-free [2822]. certain [650, 675]. Cerulli [2113]. CES [629]. Chain [1384, 924, 1734, 1116]. chains [30, 926, 2050, 944]. Chakraborti [1984]. Chambers [2197]. Chandan [2089]. Change [1034, 2664, 2455, 2998, 2454, 1359, 2866, 2459, 1980, 2463, 629, 644, 2171, 1093, 1695, 1930, 2689, 2864, 2883, 2748, 1597, 2362, 2394, 2460, 2516, 2164, 2458, 2957, 2648, 1655, 1342, 1679, 2297, 1823, 2951, 2935, 2919, 1895, 1622, 1508]. Change-of-variance [1359]. Change-point [2455, 2998, 2171, 1930, 2689, 2957, 1655, 1342, 2951, 1508]. change-points [1695, 2864, 2648]. changed [2457]. Changepoint [2456, 2717]. Changes [3003, 727, 2335, 2883, 1429, 1737, 2717, 3031, 2319, 271, 1896]. changing [2853]. channel [21]. Chapman [2198, 2147, 2219, 2218, 2259, 2197, 2217, 2556, 2160]. character [1287]. Characteristic [1645, 605, 2790, 777, 2739, 2122, 2347, 2750]. characteristics [2749, 2277, 877, 2276]. Characterization [2621, 1305, 1042, 997, 987, 486, 1719, 897, 1445, 1399, 1173, 578, 1718, 660, 1540, 1067, 1293, 1043, 414, 983, 1035, 1541, 920, 2377, 1325, 861, 815, 1106]. Characterizations [838, 615, 2058, 1096, 2022, 2165, 1121, 889, 1085, 2202, 574, 1098, 858, 1419, 2817, 753, 1311, 839, 1859]. characterize [2559]. characterized [810]. characterizing [1120]. Charakterisierung [403]. Charakterisierungen [485, 527, 537]. chart [2918, 1377, 1967, 2974, 1794, 1510, 876, 1105, 1508]. charting [1831]. charts [695, 1734, 1836, 2244, 1294, 1064, 1868, 2827, 2034, 2332, 1378, 1238, 1171, 2216, 1698]. Charu [2089]. Chatterjee [3001]. Chebyshev [2403]. check [2583]. checkerboard [3027]. checks [1825, 3029]. chemical [2098]. chemischen [146]. Chernoff [2250, 1740]. Chi [386, 2294, 2970, 1557]. Chi-square [2294, 2970, 1557]. children [688]. China [143, 200]. chisquares [329]. choice [1259, 1366, 2354, 2040, 826, 1861, 2302, 450]. choices [1077, 1196]. Cholesky [3017]. Choosing [1367, 729]. Choquet [1147, 1148, 1139, 1142, 1143]. Chow [1685, 180]. Christensen [1404, 1423]. Christoph [293]. Christopher [2556]. Chung [1685]. Circuits [2696]. circular [1721, 2580, 2638, 717, 2743, 2304, 2704, 1790, 2325, 1322, 2839, 2371]. Class [1362, 1932, 1317, 710, 1180, 494, 1148, 2082, 1055, 3045, 1717, 849, 867, 2006, 396, 2541, 1755, 2153, 1031, 2255, 1672, 2434, 1947, 2123, 1198, 2041, 1480, 1273, 837, 2309, 1587, 1195, 944, 842, 1287, 2801, 2867, 2227]. classes [1638, 1214, 2820, 1990, 2967]. Classical [2255, 1955, 2697, 1435, 2527, 1106, 2942, 1434]. Classification [197, 2315, 1723, 2326, 1832, 2354, 1912, 1109, 2079, 2431, 2789, 2187, 2141, 1987, 2947]. classifiers [2568]. classifying [2720]. Claus [2198, 2217]. clear [1476, 2357]. clickstream [2990]. Clinical [1685, 2643, 2981, 2935]. closed [2447, 1767, 2076, 2800]. closed-form [2800]. closeness [1947, 1480]. closer [915]. closure [1214, 938]. clouds [2274]. CLT [1988, 2511, 2661]. Cluster [1496, 1913, 2110]. clustered [1913, 3066, 2884]. Clustering [2831, 2284, 2027, 3061, 2834, 2733, 2167, 2501, 2726, 907, 2089, 2855, 2618]. clusters [3035]. clusterwise [2753]. Co [1686, 416]. Co-editor [1686]. co-efficients [416]. coal [27]. Cobb [2268]. Cochran [1927]. coefficient [2226, 2279, 1954, 2355, 1402, 3041, 2679, 1603, 1221, 1909, 2955, 2185, 2244, 2245, 1181, 2472, 396, 2333, 999, 1678, 3060, 2136, 2889, 2810, 2917, 1133, 1198, 2009, 1273, 2992, 2103, 2306, 2760, 1801, 2049, 2026, 2223, 2963, 2978, 2064, 2361, 2436, 2476, 1297, 3010, 2631, 3001, 2872, 2065, 3007]. coefficients [422, 455, 530, 1403, 1496, 2743, 762, 891, 650, 2040, 2475, 966, 1641, 2264, 1741, 970, 1310, 1758]. Coggeshall [2088]. Cohen [2923]. coherence [684]. Coherent [2969, 1972, 1848, 2016, 1214, 938, 2322, 2409, 2235]. cohort [1431]. coincidence [1540, 1274]. Cointegrated [1705]. cointegrating [1065]. Cointegration [1173, 1192, 906, 2594, 1044]. cold [548]. Coleman [1834]. collecting [2856]. collection [2435]. collision [2690]. column [2657, 2945]. column-orthogonal [2657]. columns [1581]. COM [1802, 1560]. combination [2047, 1457, 2300, 1855]. combinations [2540, 1357, 2290]. combined [740, 1390, 1756]. combined-optimal [1390]. Combining [1797, 1993, 1637, 854]. Comment [1472, 1544, 1523, 652]. commentary [2698]. Comments [1728, 525, 1588, 1607]. Commission [421]. commodities [522, 556]. Common [2787, 1461, 2415, 1333, 2496, 2524, 2967]. Communauté [12]. Communication [2982, 3039]. Communication-efficient [2982, 3039]. community [2909, 12, 33]. compact [559, 1798]. Companion [1708]. comparaison [12]. Comparative [2583, 1934, 2491, 788]. Comparing [1722, 1283, 1759, 1775, 2612, 2733, 1811, 2894, 301, 2692]. Comparison [1572, 695, 1600, 890, 1383, 798, 850, 1516, 1119, 1054, 2051, 2195, 1827, 598, 468, 2137, 828, 12, 1156, 1129, 1678, 2594, 2345, 806, 2040, 421, 1375, 1391, 2597, 1939, 1056, 2972, 2939, 3016, 450]. Comparisons [1480, 761, 1331, 2208, 1167, 243, 2191, 2235, 1020, 966, 2938, 779, 2067]. competing [2252, 2421, 1535, 2336]. competition [439]. complementary [2039]. Complete [2652, 1757, 2020, 2444, 2543, 2221, 1697, 242, 1576, 3067, 808, 3043, 3001]. Complex [1642, 1389, 1409, 1453, 1195, 1250, 1617, 2044]. complexity [2721]. Component [2417, 2940, 1797, 2195, 430, 1799, 2389, 1024, 429, 1998, 1993, 2627, 2409, 829, 2138, 2496, 1586, 2855, 2320]. Component-wise [2417]. components [669, 761, 2819, 1474, 2100, 2322, 2283, 2577, 1583, 1376, 1455, 1249, 2938, 2781, 909, 1022, 1183, 1433, 2227, 1818, 1476]. Componentwise [1801]. Composite [2552, 2999, 1078, 2185, 2068, 2373, 1628, 2961, 715, 2681, 2062, 2701]. Compositional [2755, 2433, 2528, 1400]. compositions [2755]. compound [2525, 1159, 916, 2744, 1128, 948, 2973, 2599]. comprehensive [1140, 2152, 2477]. Computation [998, 1036, 1813, 1683, 1869, 743]. Computational [2836, 1525, 1681, 1833]. computer [2797, 601, 2848]. Computereinsatz [291]. computermäßigen [601]. Computerprogramms [191]. Computing [1286, 1613, 2775, 735, 688, 2010, 472, 2404, 2932]. concave [768, 2538, 2900]. Concentration [813, 228, 2069, 1790, 253, 318, 817, 2563]. concept [1140, 490, 228, 2755, 366, 2, 2646]. conceptions [1295]. concepts [1052, 1072, 1694, 805]. Concerning [785]. concomitant [1395, 1446]. Concomitants [1837, 1462, 1490, 1855]. condition [1830, 1810, 2223]. Conditional [2482, 2464, 991, 1200, 2750, 1275, 1300, 761, 494, 1550, 2157, 749, 1293, 2869, 987, 1096, 538, 699, 770, 2570, 2598, 1307, 1719, 2057, 1477, 1716, 1877, 2881, 2076, 2289, 2400, 2203, 2395, 2247, 300, 352, 1485, 1445, 2359, 709, 1323, 1085, 2507, 2376, 2623, 2997, 2401, 1106, 1226]. Conditionally [1619, 2250, 1836]. Conditions [2862, 1274, 2277, 3043, 2398]. conductivity [2264]. cone [2513]. Conference [1570]. Confidence [1158, 2402, 2666, 2133, 2916, 2683, 2972, 2973, 1233, 1564, 1611, 1012, 1114, 1638, 1221, 723, 1690, 656, 260, 798, 2484, 847, 420, 2698, 2522, 1905, 690, 2269, 229, 1815, 1677, 2453, 2715]. configurations [2125]. confirmatory [1922]. Conformal [3008]. conforming [1794]. Congdon [1891]. Conical [656]. conjecture [732, 791, 1404, 1423]. Connectedness [1576]. connection [2054, 644, 1420, 2312]. connections [2252]. consequences [2668]. consideration [1323]. considerations [205, 2766, 949]. considering [407]. Consistency [1909, 2498, 1511, 909, 1752, 2398, 2734, 2083, 3067, 2977, 2247, 1845, 475, 624, 2020, 2149, 2549, 2253, 2430, 2221, 2901]. Consistent [2568, 1898, 2319, 1641, 3001, 2841, 993]. consisting [1109]. constancy [2355]. constant [2960, 548, 2260]. constant-stress [2960]. constants [3042]. constrained [2840, 1885, 945, 2639, 2453]. constraint [2030, 970, 1625]. constraints [1537, 993, 2611, 1441, 497, 2134]. construct [2893]. constructed [2434]. Constructing [1038, 2725, 556, 1132, 2572, 2632, 690, 1968]. Construction [2492, 1538, 2423, 2756, 2657, 2453, 2586, 522, 2633, 2818, 1474, 1168, 3050, 229]. constructions [2770]. consulting [1305]. Contacts [7, 14, 23, 32, 38, 566, 45, 56, 67, 76, 85, 94, 101, 110, 118, 125, 133, 141, 149, 158, 168, 177, 186, 193, 201, 212, 222, 238, 247, 256, 266, 275, 284, 294, 303, 310, 321, 332, 341, 350, 357, 373, 383, 391, 400, 408, 418, 427, 435, 446, 457, 465, 473, 481, 488, 492, 501, 510, 519, 534, 541, 549, 557, 575, 582, 592, 604, 609, 617, 634, 655, 676, 697, 718, 739, 760, 781, 802, 823, 844, 865, 886, 894, 903, 913, 922, 931, 940]. Contacts [952, 962, 973, 986, 996, 1006, 1017, 1029, 1041, 1050, 1061, 1071, 1081, 1091, 1101, 1112, 1126, 1138, 1152, 1164, 1178, 1191, 1204, 1219, 1232, 1244, 1256, 1269, 1281, 1291, 1304, 1316, 1329, 1339, 1352, 1363, 1373, 1382, 1394, 1411, 1427, 1439, 1452, 1473, 1489, 1506, 1527, 1548, 1569, 1590, 1606, 1627, 1646, 1667, 1688, 1709, 1730, 1751, 1772, 1793, 1814, 1835, 1850, 1871, 1892, 1907, 1923, 1944, 1965, 1986]. contagion [782]. containing [1380, 1476, 2357]. contaminated [2426, 2753]. Contemporary [1504]. content [668]. context [171, 2376]. Contingency [936, 599, 1247, 1670]. continuation [2821]. continuous [2621, 2783, 3062, 2851, 1407, 2378, 2151, 2817, 2891, 1746, 1025, 1109, 1556, 2087, 1798, 2736]. continuous-time [2891]. contoured [1553, 815]. contours [2125]. contrast [2911]. contrasts [2128]. contribution [171]. Contributions [682]. Control [2244, 1571, 1831, 2918, 695, 2794, 916, 1377, 1734, 1836, 1794, 1294, 1868, 1510, 2034, 1779, 1157, 935, 774, 1171, 1105, 2216, 1698, 679, 2236, 3036, 2877, 1508]. controlled [2635]. controversies [43]. Convergence [2828, 2457, 2721, 2031, 2330, 1600, 2652, 1207, 1883, 1616, 2260, 1700, 1757, 1884, 2444, 2543, 3043, 2602, 2468, 1697]. convergent [1760]. Convex [1694, 1042, 2590]. convexity [2687, 2686]. Convolution [795, 2667]. convolutions [416]. Conway [1764, 2827]. cooperation [37]. coordinate [1564, 2886]. coordinates [1611]. Copula [2710, 2192, 2740, 2413, 2169, 1722, 2284, 1886, 3028, 2412, 2152, 2659, 2937, 2235, 2843, 2319, 2336, 3027, 2582, 2948]. Copula-based [2710, 2192, 2740, 1886, 2412, 2843]. copulas [2952, 2690, 1977, 2796, 2912, 1782, 1880, 1940, 2997]. core [1144]. corrected [2209, 1949, 1597, 2181]. Correcting [2375]. Correction [2369, 2847, 3069, 2687, 2996, 2731, 2350, 2830, 1095, 2691, 1346]. corrections [1610, 1734, 2858, 2996]. Correlated [308, 2424, 2902, 3065, 1118, 1881, 1108, 1578, 1557, 2271, 1767, 2704, 2002, 2956, 2906, 2196, 1117, 876, 984, 1068, 1932, 2781, 909, 1022, 1183, 3071, 1209]. Correlation [1863, 455, 1851, 1403, 1496, 2743, 2365, 2609, 2679, 1221, 2816, 1909, 2100, 1181, 650, 339, 396, 2644, 2833, 1458, 848, 643, 2400, 1133, 1585, 1273, 641, 2352, 2992, 1939, 2735, 2953, 2026, 3001, 2371, 2168, 2625]. correlations [1897, 2424]. corresponding [2901]. cost [553, 2030, 3049, 2980, 1625, 2746]. costs [478]. COTE [1227]. count [2667, 2969, 2915, 2634, 1935, 2157, 1482, 2175, 2366, 3070, 2888, 2788, 2884, 901]. counting [2525, 976]. countries [727, 12, 584, 98]. country [586]. counts [904, 2542, 2153, 2332]. coupled [2586]. Course [2199, 2554, 2003]. covariables [2134]. Covariance [1949, 2944, 814, 2440, 2540, 2876, 2316, 2828, 2744, 2883, 2567, 1910, 777, 1031, 1580, 1621, 748, 1942, 2617, 920, 2096, 2843, 2084, 605, 2269, 2507, 2156, 1988, 2563, 2599, 2661, 2991]. covariances [752, 775]. Covariate [2807, 2391, 2248, 2519, 2762, 2933]. covariate-adjusted [2519]. covariates [2807, 2535, 2620, 1787, 598, 2923, 2488, 2581, 1109, 2422, 3052, 2927, 2575, 2600, 2062, 2480, 2370, 2414, 2320]. coverage [2522]. coverages [668, 2818]. cows [2093].

Cox

[1921, 2525, 1652, 2813, 2751, 2341, 2130]. crack [1915]. Craig [971]. Cramér [895, 2783, 701]. Cramér-von [895]. Cramér’s [2469]. Cranmer [2674]. Crawley [1545]. CRC [2675, 2198, 2147, 2219, 2218, 2259, 2197, 2217, 2556, 2160]. credibility [2168]. Cressie [2556]. crisp [2508, 1604]. criteria [2572, 691, 2691, 735, 1309, 2311, 2306, 2139]. criterion [3064, 2850, 2688, 1982, 1391, 1655, 1516, 1480, 2138, 423, 2721, 3058, 2759, 2064, 2720, 2925]. Critical [1165, 957, 2974, 1179, 1000, 783, 2565]. criticism [1521]. critique [612, 128]. cross [1901, 2864, 1371, 1117, 3007]. cross-sectional [1901, 2864, 3007]. crossing [2279, 1009, 2263]. crossover [2816]. Csiszar [1116]. cubes [2755]. cum [891]. cumulants [948]. Cumulative [2171, 331, 2387, 2946, 2000, 2327, 2205, 2025]. cure [1764, 2813, 2751, 2066]. Current [445, 1812, 1431, 1543, 1386]. curve [898, 2385, 2441, 2584, 1931, 241, 1861, 3071]. Curves [1645, 2279, 2620, 2447, 1746, 2894, 1510, 2040]. CUSUM [2827, 935, 2216, 1253, 773]. Cusum-schemes [773]. CVaR [2418].

D [757, 1486, 1708, 1963, 2109, 2104, 672, 1682, 1983, 1891, 2238, 780, 789, 1684, 2125]. D. [2055, 2109]. Dale [1870]. Dalenius [505]. damage [847, 1581, 331]. damit [368]. Daniels [1707]. dans [586, 12]. Darstellung [208, 337]. Darstelung [614]. DasGupta [1662]. Data [2675, 2555, 1706, 1642, 2217, 2368, 1673, 2160, 3074, 3021, 2186, 630, 2830, 1233, 3054, 2237, 405, 1731, 1918, 1464, 2317, 771, 2234, 2553, 1928, 1260, 2540, 1374, 1777, 2321, 2791, 2656, 744, 2667, 904, 900, 941, 359, 1707, 2915, 2634, 2406, 2834, 1693, 1114, 2171, 2515, 2983, 1842, 2724, 1073, 2856, 1722, 2354, 2157, 1717, 864, 2435, 2896, 2864, 2904, 1853, 2233, 1383, 1762, 2412, 2879, 1624, 2560, 2869, 1970, 1467, 1594, 3044, 2985, 2596, 1128, 736, 1934, 1257, 1501, 2206, 1482, 2245, 693, 1657, 2175, 2112, 2500, 2499, 2491, 2403, 1975, 1283, 2642, 1669, 2292, 1349]. data [765, 1168, 2314, 2941, 2073, 1492, 2999, 1156, 1484, 1754, 1319, 1743, 1343, 2527, 1703, 199, 1580, 2745, 793, 1652, 1321, 1912, 2261, 2158, 2522, 2002, 2107, 1789, 2517, 2057, 2906, 2912, 2662, 2366, 2494, 2889, 2878, 3013, 3005, 2998, 2094, 2339, 1431, 2460, 3055, 2764, 808, 2203, 1890, 1962, 2577, 963, 1945, 2140, 2187, 3026, 1186, 1857, 1913, 2141, 2096, 1604, 1726, 2000, 2433, 2753, 1686, 2224, 1801, 2714, 1808, 2173, 2812, 2132, 2888, 2359, 876, 2681, 2341, 477, 1920, 909, 1022, 2855, 1105, 3058, 3066, 3059, 2978, 1817, 3071, 2990, 794, 1515, 2044, 2361, 2436, 2601, 2528, 2750, 2788]. data [2884, 2139, 901, 979, 1400, 2757, 2994, 1275, 1300, 3039, 2134, 2338, 2399, 2872, 2065, 1677, 2857, 2967, 2824, 2701, 1917, 2605, 2545, 2533, 2371, 2324, 2625, 1822, 2414, 2868, 3007, 1958, 2948, 2661, 2239, 1727, 1681, 2089, 1870, 1746, 1665, 789, 1524]. data-dredging [2642]. data-driven [2139]. data-specific [1808]. dataset [2396]. Datasets [1546, 1642, 2593, 2954, 2496]. Daten [479, 580, 610, 182, 199]. datenanalyse [840, 1503]. Datenmaterial [191]. Datenqualität [912]. Dave [1834]. David [1746, 1526, 2331, 1642, 2751]. DD [2187]. DD-classification [2187]. decay [2843]. d’échange [12]. Decision [478, 1682, 1140, 462, 551, 1142, 795, 205, 824, 1309, 1145, 1749]. decision- [205]. decision-making [795]. decision-theoretic [824]. Decisions [30, 879]. Decomposing [599]. decomposition [706, 2782, 2726, 2261]. decompositions [483, 2077, 1642]. deconvolution [2700]. Decreasing [2189, 2286, 2672]. deep [3008, 3051]. deeper [3000]. deepest [2561].

Deficiency [2922]. defined [1470]. defining [441]. definite [920, 1444]. Definition [198, 2802, 450, 174]. Definitionsgleichungen [441]. degradation [2604]. Degree [2720, 762, 1827, 2651]. Degrees [2611, 744, 1095, 2061]. Delampady [1547]. delayed [2461, 1536]. deletions [2525]. delta [2650, 2183]. demand [27]. demographic [306]. Demographie [626]. Demographische [283]. demonstrated [130]. demonstration [1987]. denoised [1689]. dense [2596]. densities [1540, 2782, 1208, 1608]. Density [3029, 1702, 3004, 2552, 2169, 1467, 1766, 977, 1511, 2663, 3002, 2057, 1696, 3011, 1559, 917, 1656, 1729, 1499, 2721, 945, 1752, 2942, 2024, 2511]. density-weighted [1696]. Dental [2147]. departures [1596]. dependence [1901, 1936, 2012, 2679, 2167, 1886, 2864, 2027, 2174, 2955, 2853, 817, 608, 1966, 2336, 2043, 1694, 1697]. dependencies [1323, 1783]. dependency [1240, 2067]. dependent [455, 2870, 1558, 2652, 1658, 2284, 2498, 1298, 1883, 2322, 2534, 2057, 2456, 1945, 2648, 1485, 1757, 1979, 2149, 2444, 2600, 2183, 2951, 2468, 2221, 2901, 2661]. dependently [1762, 2412]. Depth [1976, 2803, 1909, 1710, 2668, 1912, 2752, 1358, 2315, 2559]. derivation [279, 1405]. derivative [2596]. Derivatives [1940, 1766, 1696]. derived [2975]. Deriving [1308]. dermoscopic [2467]. descent [2010, 2886]. descriptive [1703]. Design [2769, 2299, 2730, 695, 2902, 2118, 891, 2797, 430, 2849, 887, 2741, 2308, 2880, 1577, 2581, 2698, 1868, 1390, 2305, 2613, 1296, 2311, 2762, 2972, 2767, 1861, 2759, 2587, 3043, 2757, 1463, 2824, 2807, 1983, 1487]. design-cum-model [891]. designed [171]. designing [2262, 2427]. designs [1305, 1347, 2117, 2771, 2272, 1389, 1409, 2303, 1992, 1092, 2004, 1265, 2492, 2595, 216, 2816, 2513, 2633, 2526, 2572, 1576, 1575, 2304, 2696, 2308, 2135, 2654, 2268, 2569, 1995, 2423, 2756, 2137, 3009, 2006, 2483, 2567, 2795, 2258, 2589, 2875, 1579, 2038, 2704, 1581, 1324, 1398, 2448, 2820, 1469, 2894, 1991, 1248, 2300, 1584, 2765, 2597, 2039, 2784, 2306, 2760, 2836, 2309, 2758, 2852, 1277, 2128, 2312, 2098, 1421, 2936, 2720, 2358, 2989, 2176, 2518, 2558, 2725, 2565, 1476, 2357, 2586, 2576]. Desmarais [2674]. detect [2262, 1966]. Detecting [1737, 2362, 2516, 2827, 1937, 2132, 3035, 2374, 1510, 2779, 2319]. Detection [2864, 1342, 1971, 2578, 2689, 916, 2869, 2664, 1429, 788, 2417, 657, 2717, 1263, 611, 2394, 2909, 2460, 2457, 2735, 2454, 2462, 2951, 2459, 2919]. determinants [1808]. determination [981, 738, 2139]. determined [924]. determining [989, 1131]. deterministic [1413, 927, 2129]. Deutschland [147, 155, 464, 480, 71, 467, 130]. Deutschlands [92]. develop [1734]. developed [727]. Development [632, 243, 98]. developments [429, 1364]. deviance [2827]. deviation [2702, 1769]. deviations [1695, 2541, 1874, 1994, 1616]. device [1617]. DGFR [2438]. DHD [2090]. diagnosis [1418]. Diagnostic [1825, 1193, 3032, 2044, 1591]. diagnostics [2693, 1650, 2296, 1415, 2066, 3033, 1902, 1803, 1951]. diagonal [2727]. dichotome [616]. dichotomisches [199]. dichotomization [2898]. dichotomous [1349, 199, 2672, 2189, 2676]. Dickey [1592]. Dickinson [1984]. Diehl [512]. Dienstleistungsbereichen [9]. Dieser [194, 558, 47, 57, 68, 77, 86, 95, 102, 111, 119, 126, 134, 142, 150, 159, 169, 178, 187, 202, 213, 223, 239, 248, 257, 267, 276, 285, 295, 304, 311, 322, 333, 342, 351, 358, 401, 374, 384, 392, 409, 419, 428, 436, 447, 458, 466, 474, 482, 493, 502, 511, 520, 535, 542, 550, 567, 583, 593]. Dieter [2298, 2432, 2298]. Dietmar [2217, 480]. Dietrich [349]. Difference [1618, 1158, 2364, 1611, 2620, 2083, 1522, 1333, 2854, 1308, 2954, 2972, 1416, 1022, 229, 2798, 1758]. Difference-based [1618, 2364]. differences [2684, 2906, 1756]. different [2172, 744, 1496, 243, 2137, 1509, 850, 2086, 2099, 1196, 237, 2558]. differential [2873, 122]. Differentialgleichungen [122]. differentially [2651]. Differenzen [433]. Differenzenfiltern [397]. Differenzierung [209]. difficulties [882]. diffusion [2776, 1874, 1899, 3038, 2265]. digital [2125]. dilation [1043]. Dimension [2523, 2947, 2474, 2882, 2650, 2887, 2928, 2639, 2414]. dimensional [2317, 2540, 2926, 2965, 2913, 2876, 599, 2864, 2560, 2815, 2869, 2159, 2185, 2688, 2659, 1942, 2079, 2538, 3006, 2449, 2959, 2789, 1964, 1919, 2496, 2927, 2309, 352, 3057, 2623, 2031, 2353, 2587, 2994, 1929, 2480, 2857, 2877, 2071, 2563, 2625, 2680, 2414, 2637, 2661]. dimensionalen [48]. Dimensionality [3056, 3044]. dimensions [2905, 2966, 1045, 3001]. diminishing [3011]. Dimitris [2554]. direct [1405, 1453, 1808]. directed [2488, 2210]. direction [2580, 2325, 2405]. Directional [843, 2565]. Direkter [75]. Dirichlet [777, 2875, 3026, 1376, 3000, 605]. disaggregate [1775]. disaster [3070]. discontinuity [2194]. Discordancy [2018]. discovery [2794, 3036]. discrepancy [2473, 2039, 2363, 2585, 2989]. Discrete [2783, 2826, 759, 1798, 858, 1992, 768, 1765, 2045, 1491, 900, 1419, 1293, 692, 987, 1096, 1938, 693, 706, 1336, 1307, 2718, 2434, 1109, 2366, 1556, 3052, 1195, 1250, 826, 2388]. discrete-valued [2045]. discretely [1898]. Discriminant [1580, 1843]. discriminate [3004]. discriminating [763]. Discrimination [2770, 2117, 1777, 991, 2958, 2709]. discriminatory [152]. discussion [440]. disease [2773, 1902]. diseases [2964]. diskreten [368]. diskreter [252]. Diskussionsbeitrag [440]. disorders [2462]. Disparitäts [365]. Disparitäts- [365]. disparity [744]. dispersion [2356, 2738, 1270, 882, 1844, 824, 2311, 1626, 2132]. Dispersive [1356]. distance [2226, 2578, 639, 1379, 2365, 2195, 1118, 2250, 2385, 2739, 2400, 1740, 2954, 1778, 2953, 1174, 2994, 1689, 2625]. distance-based [2226]. distances [2783, 2214]. distinct [1674, 1122]. distinguishable [2833]. distortion [2503, 2485, 2547]. Distributed [3058, 1286, 1613, 561, 752, 664, 448, 1959, 2005, 1492, 2999, 2322, 1692, 2119, 2982, 588]. Distribution [2226, 308, 1276, 354, 1819, 1956, 3011, 2235, 2310, 1543, 1446, 681, 1341, 1529, 1846, 1831, 3017, 1736, 578, 615, 1718, 924, 1165, 1742, 1224, 2042, 2737, 2722, 2046, 2655, 2847, 648, 687, 1653, 2008, 1860, 1889, 2081, 2382, 1012, 1540, 2228, 1088, 1634, 1534, 2013, 2321, 530, 1030, 628, 439, 546, 638, 226, 287, 1800, 1159, 639, 1615, 1738, 1462, 162, 242, 1046, 1114, 705, 2821, 215, 1802, 1842, 1856, 1720, 2849, 1854, 1459, 2435, 470, 916, 1538, 3037, 2747, 2808, 2233, 1762, 1298, 2879, 2660, 2960, 832, 1470, 2944, 1761, 1862, 653, 1938]. distribution [820, 2007, 2005, 1457, 1676, 1657, 396, 1768, 316, 2349, 2478, 846, 260, 1235, 836, 948, 1847, 1522, 2941, 2073, 852, 1699, 2962, 1336, 1475, 899, 2200, 2484, 2806, 1651, 1691, 2875, 1981, 2177, 1319, 2069, 667, 730, 1633, 1620, 1878, 2213, 1904, 554, 2386, 2627, 259, 2409, 977, 1915, 1465, 346, 1790, 513, 1511, 1811, 1942, 2240, 3048, 144, 263, 675, 1057, 2732, 983, 1035, 2421, 2428, 808, 1541, 2203, 2315, 2981, 2811, 806, 920, 1844, 751, 2479, 2825, 2957, 1857, 1955, 1559, 1560, 1888, 2777, 1311, 3031, 1913, 1026, 1820, 1969, 2922, 956, 1997, 2265, 1519]. distribution [2529, 300, 352, 1876, 2347, 1250, 1325, 1518, 1933, 1535, 416, 1036, 1855, 477, 1228, 2030, 2201, 2264, 2202, 2939, 340, 1449, 2216, 2673, 679, 157, 945, 1174, 2685, 815, 2074, 743, 979, 1388, 1625, 666, 2330, 2645, 2101, 2211, 2800, 2334, 1715, 2991]. Distribution-free [2235, 1543, 705, 1915, 346]. distribution-preserving [920]. Distributional [2200, 2924]. Distributions [1781, 1970, 675, 1497, 1428, 1721, 1120, 1335, 1365, 2621, 838, 2940, 1317, 710, 997, 1928, 1054, 2053, 1103, 933, 821, 2783, 2826, 1088, 3004, 2926, 2282, 898, 587, 1765, 2438, 483, 1837, 2743, 1055, 1220, 1435, 722, 975, 1299, 2574, 1507, 1246, 1603, 1828, 1533, 2666, 1419, 2091, 2863, 890, 2822, 1293, 1624, 1407, 867, 1594, 1096, 414, 2143, 1882, 1663, 2331, 2817, 1509, 2754, 777, 2210, 2068, 1733, 1647, 2415, 2255, 795, 1599, 2976, 1025, 489, 2379, 1714, 2325, 1719, 2434, 1498, 2739, 1477, 1716, 2878, 1947, 1322, 1885, 2606, 1616, 2466, 2916, 1816, 2087, 1376]. distributions [753, 2559, 1740, 850, 1121, 1804, 2196, 1585, 2085, 1598, 2646, 1361, 1223, 1182, 2575, 1838, 1195, 605, 978, 1445, 1058, 2043, 1843, 2939, 2052, 839, 1859, 2276, 2509, 1106, 1098, 2967, 1803, 1951, 2344]. disturbances [853, 1023, 2082, 1108, 1227, 790, 1210]. Divergence [1975, 1321, 2777, 3004, 2552, 2151, 2695, 3012, 1511, 1441, 2373, 866, 1531, 1345, 1223]. Divergence-based [1975, 1321, 2777]. divergences [2257, 1116, 2311]. diverging [2647, 2443]. DIVISIA [612, 507, 685, 496, 612]. DIVISIA-indices [612]. divisibility [889]. divisible [2943, 2939]. DLR [1214]. Dmitrienko [1771]. DMTTF [3045]. do [1903]. Dodge [929, 1440, 1612]. does [3064, 1014, 1754, 2559, 1680]. DOI [1728]. dominance [2374, 2009, 917]. domination [1619]. dose [2262, 2773, 2236]. dose-escalation [2236]. dose-finding [2773]. Double [1023, 1936, 2435, 2179, 1377, 1169, 766, 1362, 1198, 1240, 1769, 1929]. double-hurdle [1240]. Double-length [1023]. double-truncation [2179]. doubling [2633]. Doubly [2494, 648, 2081, 1765, 1180, 2028, 1633, 513, 1251, 2341]. Douglas [1834, 1660, 2268]. Dr. [323]. dredging [2642]. drei [371, 485, 527, 97]. dreier [403]. drift [2001, 715, 2080]. driven [2993, 2388, 2139, 2060]. drop [699]. drug [2300]. Drygas [471]. dual [2257, 2722, 1362, 2615]. Dudoit [1568]. due [1206]. duplication [1469]. duration [1921, 2012, 2076, 2289, 2339, 2395, 2359]. Durbin [298]. durch [255, 356, 217, 196, 166]. Durchschnitte [297]. Durchschnitten [397]. Dynamic [2467, 2167, 1918, 2387, 716, 554, 2845, 2203, 2946, 1387]. dynamical [2767]. Dynamik [499]. Dynamische [833]. dynamischen [88, 250].

each [1198, 2009]. early [2772]. Easy [2147]. EBLUP [1513]. eBook [2197]. econometric [2791, 1894, 171, 205, 2113, 1748]. Econometrics [1708, 682, 971, 24, 1504, 720, 2112, 1704, 2145]. Economic [12, 33, 818, 695, 717, 727, 686, 1429, 525, 2113, 429, 361, 1703, 524, 195, 189, 25, 440, 421]. Economics [787, 1143, 2108]. economies [243]. économique [12]. economy [503]. Ed [1725, 2056, 1686]. EDF [2384]. EDF-based [2384]. edge [3067]. Edgeworth [1690]. Edition [2198, 2147, 2218, 1726, 2055, 2105, 2367, 2349, 1526, 1685, 1962]. Editor [1906, 1774, 1686, 1502, 1773, 1408]. Editorial [719, 2116, 2299, 2514, 1113, 2454, 914, 1051, 2757, 1]. editors [1661, 862, 1893]. edn [1770]. eds [1771, 1890, 2089, 2217, 2110, 799]. education [154, 71]. educational [2791]. Edward [2002]. EEG [1902]. EFARIMA [2012]. Effect [1413, 197, 2902, 1648, 2797, 430, 2520, 2215, 829, 2393, 1528, 1587, 2557, 560, 2405, 1677]. Effective [2264, 2633, 2079]. effectiveness [2779]. Effects [855, 1918, 1957, 3065, 2832, 1664, 2451, 1482, 1154, 2137, 2499, 1551, 2333, 2988, 2248, 2956, 2617, 2252, 2930, 2128, 2812, 1196, 1966, 2610, 2963, 2978, 324, 325, 326, 1222, 189, 2835, 2357, 1636]. Effekts [207]. Efficiencies [870]. Efficiency [1054, 1776, 2191, 1827, 761, 2045, 2296, 1369, 1227, 643, 2374, 2731, 2742, 1022, 861]. Efficient [2042, 2913, 393, 1990, 2591, 1074, 2436, 2835, 1558, 2983, 1019, 964, 334, 2380, 1333, 2982, 2361, 2601, 3039, 2480, 2800, 2807]. efficients [416]. Effizienzvergleiche [269]. EFTA [83]. Eggenberger [782]. Eichhorn [810]. Eigenschaften [564]. eigenvalues [2747]. eigenvector [2330]. eigenvectors [1988]. Eignung [41, 49]. Einbeziehung [337]. einem [585, 467, 273]. Einfache [590]. einfacher [434, 678]. Einfachheitspostulat [568]. Einführung [626, 281, 179]. Einheiten [347, 80]. Einige [552, 273, 80, 529, 595, 606, 54, 372, 365, 362, 563, 319, 424, 298, 441, 480, 11, 17, 347, 451, 291, 258]. einiger [62, 345]. Einkommensdisparitäten [262]. Einkommensverteilung [108]. Einkommensverteilungen [348]. Einteilung [84]. Einzelgleichungen [171]. Einzelhandelspreise [91]. einzelnen [220]. either [2924, 2526]. EIV [2403]. Ekstrøm [2198]. elasticity [1925]. election [1586]. electromagnetic [2264]. elementary [2108]. Eliciting [1635]. Elizabeth [1660]. Elliott [2218]. ellipsoidal [2453]. Elliptical [2391, 2226, 1193, 1066, 1716, 1933, 2949]. elliptically [1621, 815]. ellipticity [2607]. EM-algorithm [2195]. embedding [2574]. Emma [2106]. Empirical [2535, 1840, 1696, 2422, 2888, 1416, 3059, 1792, 885, 2011, 2746, 1702, 1978, 2021, 2626, 2464, 2815, 767, 2590, 2499, 2390, 1910, 1597, 584, 1170, 2136, 2854, 2998, 2809, 2122, 2040, 2123, 1033, 2934, 2615, 2929, 2231, 3046, 3014, 3034, 2692, 2330, 2640, 2025, 1958, 1758]. empirically [2833]. Empiricist [1708]. empiriebezogenen [614]. empirisch [376]. Empirische [480, 508]. empirischen [153]. empirischer [254]. Ende [5]. Endlichkeitskorrekturen [678]. endlichstufige [412]. Endnachfrage [255]. endogeneity [2546]. endogenous [1478]. endpoints [2424, 2565]. Energieverbrauchsprognosen [139]. energy [1237, 2766]. Engpaßbereichen [368]. Enhanced [2995, 2565]. enrichment [2771]. entropies [500, 768]. Entropy [1396, 462, 1341, 2975, 1635, 1211, 2892, 1945, 2327, 1759, 2025]. Entscheidungen [375, 544, 29, 121]. Entscheidungsbegriff [15]. Entscheidungsmodellen [651, 49]. Entscheidungsmodells [88]. Entscheidungsprinzip [596]. Entscheidungsproblem [594, 18, 551]. Entscheidungssituationen [204]. Entscheidungsspielen [17]. entscheidungstheoretischen [590]. entscheidungstheoretisches [532]. Entscheidungstheorie [58]. Entstehung [3]. Entwicklung [288, 98, 277, 261, 88]. Entwicklungsländer [140]. Entwicklungsländern [480]. Entwicklungsniveau [243]. Entwicklungsprognose [200]. Entwicklungsstand [515, 514, 521]. environment [1574, 2155]. environmental [1665]. environments [955]. Epidemic [2717, 2164]. Epps [2747]. Epsilon [1738]. equal [1014, 1994]. Equalities [1941, 1741, 2156]. Equality [3025, 455, 1076, 1108, 1520, 992, 2295, 910, 497, 2646, 3014, 2440]. equation [822, 1358, 560]. equations [2273, 2004, 171, 441, 2494, 122, 680, 1444, 2601, 229, 949]. equi [1117, 984, 1068, 1209]. equi-correlated [984, 1068, 1209]. equi-cross-correlated [1117]. equicorrelation [1076]. equidispersion [2542]. equilibrium [1498]. equineighbored [2704]. Equivalence [1456, 2630, 937, 1602, 3025, 2759, 2780, 2948]. Equivalent [3043, 1265, 749]. equivariant [1330, 974, 2316, 2579]. era [2856]. Ereignissen [677]. Ergebnisse [362, 338, 163, 208, 156, 211]. Erhebung [338]. Erhebungen [580]. Erhebungseinheiten [41]. Erhebungsfehler [912]. erklärenden [217]. Erlang [1286, 2663]. Erlernung [204]. Ernesto [2199]. Erneuerung [335]. Errata [1300]. Erratum [1849, 734, 1964, 1729]. Errechnung [506]. Erroneous [2369]. Error [1109, 580, 2830, 1197, 2184, 669, 761, 1936, 637, 1153, 2391, 2734, 1002, 1900, 2863, 755, 3047, 2815, 1799, 1087, 2232, 3015, 2068, 3068, 856, 1621, 977, 1630, 1678, 2804, 2718, 2348, 2136, 2477, 1094, 2150, 3031, 1516, 2347, 438, 1641, 1924, 909, 1022, 1183, 2583, 2921, 2188, 2980, 779, 2267, 2064, 1763, 2399, 2101, 1818, 1803, 2511, 2334]. error-based [2718]. Errors [1777, 1259, 1127, 2902, 359, 1631, 1650, 2083, 3065, 2964, 1592, 2550, 1840, 2503, 1118, 2670, 2570, 740, 2985, 1369, 2499, 2403, 923, 1998, 2258, 1767, 1553, 643, 868, 1272, 1714, 2956, 2136, 3032, 1845, 751, 2813, 1391, 641, 817, 1701, 2126, 353, 382, 2102, 411, 1117, 1277, 2288, 2340, 1932, 2781, 1387, 1745, 2149, 2476, 2933, 1792, 2450, 2074, 2531, 2221, 1689, 2497, 2485, 2547, 3007, 1636]. errors-in-variables [1840, 2670, 740, 2499, 923, 2136, 751, 1701, 2126, 2340, 1792, 2450]. Erste [246]. ersten [433]. Erwartungswert [181]. Erweitertes [246]. Erweiterungen [552]. escalation [2236]. ESEMIFAR [2012]. Esenwein [626]. Esenwein-Rothe [626]. especially [574]. essays [24]. Esseen [2550]. ESTAR [1672, 2048]. Estimability [2130]. estimate [2035, 2262, 242, 3044, 2036]. estimated [669, 891, 1198, 2009, 1843]. estimates [1054, 764, 162, 1722, 2814, 137, 334, 775, 144, 1261, 1196, 854, 1608, 1686]. Estimating [2207, 1674, 1093, 2028, 1565, 2590, 2892, 1368, 1032, 3024, 751, 2458, 824, 2504, 2201, 1432, 1097, 2531, 3062, 722, 2446, 513, 1358, 2494, 2345, 2247, 892, 1453, 2128, 2610, 2601, 949]. Estimation [1846, 1224, 1395, 1957, 2092, 1447, 821, 2321, 1462, 1693, 2983, 2532, 1459, 2233, 1213, 1264, 1821, 197, 957, 1181, 2389, 1024, 1013, 1578, 1004, 923, 1349, 2703, 852, 2962, 2787, 2806, 1307, 1562, 2487, 999, 1320, 2530, 1620, 2213, 2627, 2409, 2956, 263, 2810, 3038, 2428, 1945, 1371, 2050, 1000, 956, 2628, 2714, 2297, 1208, 1122, 1632, 1659, 1209, 2618, 2933, 1075, 1549, 1818, 2547, 3021, 3007, 2603, 2831, 1365, 2940, 1317, 2799, 1330, 2042, 3063, 2237, 1259, 1731, 1954, 1918, 2317, 2950, 1318, 1001, 953, 1653, 1928, 974, 853, 1629, 2926, 2902, 2656, 1030, 1166, 628]. estimation [2045, 2411, 469, 937, 1800, 2512, 1837, 1002, 762, 2169, 1220, 1435, 891, 975, 1306, 2562, 3065, 2700, 946, 2832, 1879, 1695, 2209, 430, 2222, 2911, 1107, 2425, 2091, 2445, 2435, 3028, 2896, 2027, 2819, 2660, 1407, 1799, 1467, 2220, 2664, 976, 1881, 251, 2987, 2850, 2596, 282, 881, 650, 3015, 958, 2061, 316, 846, 260, 2229, 1062, 1766, 942, 959, 2258, 1561, 882, 899, 2248, 2001, 2380, 1156, 1536, 2385, 3020, 1319, 2069, 1647, 1380, 1633, 1671, 2255, 2386, 2527, 2441, 1064, 2976, 1898, 2659, 1170, 2119, 1714, 2325, 1654, 1321, 2261, 2158, 1512, 1994, 2348, 2519, 2517]. estimation [2538, 2057, 2895, 2410, 1477, 2494, 3005, 2289, 2861, 2094, 2418, 2584, 2839, 1844, 2813, 2281, 3011, 2140, 1133, 2765, 1955, 2150, 497, 540, 1628, 2124, 1273, 1997, 2524, 694, 2126, 1047, 2843, 2529, 1679, 1801, 2049, 2729, 2288, 2812, 2615, 1494, 2971, 393, 1535, 834, 877, 944, 1950, 1829, 1044, 1673, 1641, 2341, 1384, 1866, 2591, 43, 854, 1313, 2269, 3058, 2557, 2963, 1037, 1823, 807, 2887, 1285, 970, 1595, 1979, 2204, 2436, 2788, 2450, 2867, 1873, 1297, 1310, 1478, 189, 2059, 2639, 2942, 1275, 1300, 666, 2835, 2343, 2480, 1689, 2024, 2324, 2398, 2669, 2637, 728]. estimations [2245, 2906, 1033, 1741, 2429]. estimative [2721]. Estimator [2328, 2364, 1042, 2172, 1573, 2734, 3018, 2552, 3053, 2082, 1787, 1784, 2083, 2171, 2195, 1949, 1909, 2550, 2498, 991, 1776, 2570, 740, 2135, 2955, 3067, 1087, 2232, 2151, 1501, 988, 2100, 2805, 1794, 2988, 1998, 1370, 1522, 1993, 1207, 1532, 993, 848, 2242, 1333, 1621, 2015, 977, 1790, 2215, 2481, 1511, 1614, 1788, 2420, 3048, 926, 2076, 2641, 1094, 1198, 1391, 2009, 1964, 1919, 859, 917, 1011, 2899, 1480, 2571, 2713, 1345, 2310, 837, 2874, 1779, 1838, 2205, 1494, 2048, 2953, 2388, 1442, 1932, 2781, 1022, 1618, 1038, 1744, 2551, 2997, 1210, 2020, 2361, 2549, 1752, 2256, 2430, 1563, 2518, 2544, 2723]. estimator [2471, 2901, 949]. Estimators [1221, 2344, 1258, 1637, 2437, 2553, 669, 2013, 2791, 2952, 2931, 2578, 1159, 639, 1002, 1132, 1797, 1379, 1461, 1063, 1007, 870, 2465, 1717, 755, 2277, 2920, 2225, 3047, 2179, 1211, 723, 1118, 1108, 1474, 2478, 2090, 2537, 1910, 2295, 1807, 1602, 1332, 1129, 1272, 1362, 1915, 1678, 2718, 2283, 776, 1947, 2606, 1441, 2466, 806, 1531, 1845, 2719, 2702, 3025, 1249, 1585, 2731, 2742, 2372, 1528, 875, 1516, 1701, 1194, 1778, 837, 966, 475, 2579, 2340, 2510, 700, 1228, 1287, 1384, 1481, 2801, 1990, 1931, 2301, 2156, 779, 2440, 1160, 945, 797, 2149, 2381, 2612, 1119, 2253, 2062]. estimators [2221, 2168, 2800, 2060, 2511, 2991, 589, 2407]. Estrada [2199]. Euclidean [1778, 1795]. EUR [2182]. EUR181 [2019]. europäische [6, 290]. Europäischen [347, 93, 124, 107]. Europas [98]. Europe [1686, 421, 98]. European [468, 12, 584, 33, 1586]. européenne [12]. EV [1845, 2223, 2543]. evaluate [2405]. evaluation [359, 2699, 2113, 2276]. Evaluations [1601]. event [693, 237, 2414]. events [633, 640, 646, 663, 671, 683, 703, 714, 726, 747, 769, 792, 800, 812, 819, 831, 841, 874, 884, 893, 902, 911, 921, 930, 939, 951, 961, 972, 985, 995, 1016, 1028, 1040, 1049, 1060, 1070, 1080, 1090, 1100, 1111, 1125, 1137, 1151, 1163, 1177, 1190, 1203, 1218, 1231, 1243, 1255, 1268, 1280, 1290, 1303, 1328, 1426, 2525]. every [157]. Evidence [1429, 3064, 1383, 2711]. Evidential [1464]. EViews [1706]. EWG [230, 244, 83, 218, 34, 75, 336, 91, 120]. EWG-Länder [75]. EWG-Ländern [34, 91]. EWMA [2974, 1294, 2827, 2034, 1105]. EWMA- [1105]. EWNA [1378]. ex [163]. ex-post-Input-Output-Tabelle [163]. Exact [2707, 245, 2636, 1612, 1025, 2379, 2240, 1144, 1885, 2957, 1376, 1292, 1656, 237, 695, 2513, 977, 2034, 2758, 1855, 1924, 945, 1729]. exakte [362]. exakten [616]. Exaktheitsanalyse [139]. examination [915]. example [1076, 130]. examples [946, 1356, 2433]. Exceedance [2968, 667, 3069]. exceedances [1103]. exceeding [1712]. Excel [1566]. Excess [2268, 2569, 1053, 2051]. exchange [737, 12, 407]. exchangeability [1639, 1668]. exchangeable [1538, 1455, 1855]. exciting [2886]. exist [1744]. Existence [157, 857, 629, 1854, 2914, 1390, 837]. exogeneity [584]. exogenous [560]. expansion [1690]. Expansions [1608, 2101, 2200]. expectancy [1634]. Expectation [1745, 2736, 1298, 1014, 856, 1719, 2247, 1445, 1968, 1106]. expectations [2438, 987, 1096, 1881, 1014, 982, 1942, 1490, 1085]. Expected [1140, 1147, 1032, 1825]. expectile [2983, 2689, 2637]. expectiles [2738]. expensive [2581]. Experimental [2004, 1487, 2262, 2816, 2633, 2836, 830, 2767, 2989]. Experimente [17]. Experimenten [645]. experiments [1571, 1635, 1648, 2118, 2797, 887, 2513, 2741, 2427, 2769, 2006, 2875, 2658, 1983, 966, 854, 1172]. explanatory [2629, 2993]. explicit [2407]. Explorative [840]. exploratory [896, 1962]. explosive [762]. Exponential [469, 1542, 1529, 1831, 578, 615, 1742, 1224, 2042, 2092, 1447, 1928, 821, 1540, 2228, 2707, 3004, 1046, 1205, 1802, 2465, 1867, 1459, 890, 2179, 2960, 1956, 934, 1024, 1509, 260, 1236, 989, 1647, 1633, 2379, 2522, 2663, 675, 2018, 1541, 454, 850, 956, 2504, 1517, 2190, 363, 1056, 1238, 3049, 839, 2774, 2685, 1297, 1098, 1131, 1625, 2645, 2780, 2530]. exponential-type [839]. exponentiality [1067, 2636, 2123, 2070, 221, 2384, 2025]. Exponentially [1698, 561, 2918, 1492]. exponentials [1847]. exponentialverteilte [674]. Exponentialverteilung [498]. exponentiated [2918, 2974, 1671, 2255, 1888]. export [2339]. expression [2641, 1656, 1729]. extended [1615, 2652, 1856, 2498, 2719, 2221]. Extending [2197]. Extension [3061, 1862, 2387, 736, 1509, 2037, 2152, 1960, 1640, 1820, 1969]. extensions [2194, 1018, 1691]. extra [797]. extrapolation [423]. Extremal [2635, 2127]. extreme [1529, 1395, 287, 975, 2608, 2424, 2178, 3042, 1106]. Extremes [1436, 1406, 1224, 1600, 354, 2562, 2200, 1446, 2101]. extropy [2817, 2946].

F [2182, 2002, 1504, 2110].

Fächerwahlverhalten [491]. Factor [1493, 1901, 2308, 2137, 2795, 2229, 896, 2659, 2732, 1985, 2959, 2882, 1922, 2600, 2168, 1476]. Factorial [666, 1347, 623, 507, 553, 644, 1092, 2513, 2755, 1469, 661, 1946]. factorials [1305, 2473, 2632, 2363, 2585]. factorization [2909, 605]. factors [2629, 1367, 2526, 2624, 2139, 2925]. failure [784, 1957, 1447, 1634, 2699, 1848, 2560, 2960, 820, 693, 2023, 2286, 2636, 2892, 1333, 2622, 2809, 2041, 2751, 2278, 1251, 548, 2510, 2030]. failures [1781, 1885]. Faktorantithesen [564]. Faktoren [485, 527, 537].

Faktorenanalyse [424, 673]. Faktorenmodell [319]. Faktorenwerten [319]. Faktorumkehrteest [518]. fallacy [114]. Falsche [377]. falschen [432]. false [2794, 3036]. Familie [165]. Familien [81]. families [722, 1236, 2937, 2606, 2466, 1946, 1297, 2780]. family [639, 1738, 1837, 2222, 2863, 2179, 1211, 1407, 976, 1804, 2294, 1058, 1446, 1068, 1228, 1481, 2470, 2774]. family-wise [2863]. Fast [1294, 1963, 1912, 2187, 3049, 1828]. fat [2598]. fat-tail [2598]. fatigue [331]. faulty [657]. FDR [2877]. feasibility [505]. feasible [517]. Feature [2560, 3036, 2365, 2855, 2750, 2994, 2919, 2625]. features [1294]. Federal [71, 468, 130]. fehlenden [649]. Fehler [182]. Fehlerfortpflanzung [394]. Fehlern [787, 642]. fehlspezifizierten [547]. Felder [456]. Fellegi [1323]. Fensterfunktionen [385]. Fertilität [581]. Fertility [591]. Festlegung [171]. few [1248]. FGM [2152, 2336]. Fibre [708, 1957]. fidelity [2797]. Fiducial [79]. fields [1979]. Fieller [2111]. Figini [1726]. filling [2875, 2820]. Film [413]. filter [2028, 2767]. filtering [720]. Filzmoser [2433]. final [992]. finance [2598, 2269, 1873]. Financial [1725, 932, 2053, 2463, 1722, 2167, 1594, 971, 2837, 2359, 1966, 1171, 1285, 130]. finanzielle [315]. finanziellen [140]. Finanzierungsrechnung [130]. FIND [137, 2358, 137]. FIND-maximum [137]. FIND-Maximum-Likelihood [137]. finders [518]. Finding [1596, 2773]. Finetti [783]. Finite [1736, 1787, 1948, 691, 2509, 2831, 1001, 1928, 2965, 1234, 764, 1132, 1265, 1974, 2091, 1961, 864, 1213, 1821, 1142, 2644, 2878, 1032, 751, 2125, 1261, 879, 1038]. finite-dimensional [2965]. Finite-sample [1787, 1234]. finiteness [2862]. Finn [1682]. First [2199, 2776, 2911, 2960, 2304, 2308, 3040, 1307, 2892, 2093, 2554, 2003, 2102, 1775, 2030, 1022, 1313, 1037, 1863, 2798]. first-failure [2960, 2892, 2030]. first-order [2911, 2304, 2308, 3040, 2102]. first-passage-time [2776]. fisher [1312, 2903, 1158, 660, 711, 486, 172, 1283, 2069, 672, 1940, 685]. Fisherschen [616]. Fit [389, 1728, 1067, 2282, 2690, 2885, 1593, 2808, 2904, 2412, 152, 260, 2588, 942, 959, 87, 919, 2240, 2706, 2122, 2395, 3052, 1782, 1020, 2000, 2564, 1223, 2439, 2025, 2582, 2948]. fitness [3029]. Fitting [2912, 2237, 2117, 1828, 2195, 489, 363]. fixed [1918, 2054, 2033, 2499, 2988, 2956, 1587, 3027, 2963, 2978, 324, 325, 326, 2587, 2639, 1636]. fixed-effects [1918, 1636]. Fleischer [912]. Flexible [2926, 2634, 2558, 2427, 2246]. Flexible-dimensional [2926]. flood [2856]. flow [2117]. flows [1537]. focus [2525]. focused [2064]. foldover [2363, 2613]. Folgerungen [273]. follow [2427]. follow-up [2427]. Following [2392]. Football [2555]. Forderung [532]. forecast [1127, 361]. Forecasting [1084, 1180, 2525, 2728, 3070, 2163, 2452, 3019, 243, 888, 2899]. forecasts [1414, 1775]. foreign [1808]. form [546, 638, 1615, 597, 2076, 612, 2800]. formation [1214, 938, 472, 3]. formula [404, 526]. formulae [653, 2265]. formulations [1560]. Formulierung [585]. Forschung [124]. Forsmann [912]. Fortran [688]. forward [2693, 2273, 2733, 2375, 1922]. Foster [2106]. foundation [1703]. Foundations [2217, 3074, 494, 2427, 2367, 1985, 2088]. four [2448, 2798]. four-level [2448]. Fourier [2885, 2249, 2040, 2479, 2176]. FPCA [2512]. FPCA-based [2512]. FQ [1208]. fraction [1794, 608]. Fractional [1897, 2537, 2242, 2594, 1469, 2388, 2544]. fractionally [925, 2531, 2866]. fractions [517]. Frage [195]. frailty [1509, 2285, 2148]. frame [2735, 2615]. frames [843]. framework [1140, 490, 2785, 2048, 2495]. Frank [1771]. Frans [1686]. Frauen [165].

Fréchet [3071]. free [705, 2822, 2671, 1915, 346, 2400, 2811, 2235, 1543, 2994, 2932, 2625]. freedom [1095, 2061, 2611]. Fremdinformationen [171]. French [586, 12]. frequencies [3070]. frequency [3067, 2446, 2899, 2359, 3059]. Frequentist [1257, 2975, 1673]. Friedensforschung [214]. frontier [1558]. full [2287, 2176]. Fuller [1592]. function [1258, 2207, 1042, 687, 1447, 1629, 1067, 1417, 2013, 2902, 2656, 1765, 3062, 494, 1078, 1220, 1367, 2849, 2445, 2790, 878, 1211, 820, 988, 3015, 980, 845, 2390, 279, 334, 2588, 1522, 1520, 1711, 1405, 2687, 2686, 144, 2428, 2122, 330, 2719, 1521, 2150, 875, 917, 1011, 2036, 956, 2504, 2413, 2347, 2205, 2070, 681, 2510, 855, 2781, 1843, 3058, 3049, 2788, 2893, 1106, 2932, 2024]. Functional [3019, 1988, 2563, 1731, 2931, 2902, 2965, 2512, 2834, 2445, 2596, 1934, 2206, 2988, 2417, 2865, 2362, 2410, 2662, 2810, 2315, 2577, 2839, 2984, 2187, 3031, 2103, 2628, 709, 2855, 2142, 2523, 2928, 2951, 2793, 2442, 2897, 2561]. functional-circular [2839]. functionals [1144, 2127, 1399]. functions [1335, 1552, 857, 1012, 494, 2356, 1007, 2284, 2846, 1976, 3028, 278, 1538, 2828, 1264, 1355, 1819, 598, 1755, 1602, 977, 2516, 2315, 2740, 1009, 880, 2096, 1309, 1047, 1778, 523, 1668, 3059, 945, 2183, 2893, 1097, 1694, 2582, 3016, 728]. fundamental [2732, 1525]. Fundierung [590]. Funktional [612]. Funktionalform [612]. Funktionen [252, 191]. funnel [2730]. Fürst [127]. Further [2054, 2840, 2046, 415, 530, 727, 2733, 888, 652]. fusion [2900]. fussy [2246]. future [2228, 2707, 2829, 1553, 1543]. futures [1937]. Fuzzy [953, 543, 1604, 2508, 736, 1989, 795, 2501, 1032, 1791, 2280, 2908, 2413, 1937, 1422, 794, 2095].

G [816, 1660, 912, 805, 440]. G. [1749, 625]. Gain [608]. game [652]. games [791]. Gamma [849, 2233, 1691, 1947, 838, 2437, 664, 354, 2712, 1114, 1828, 215, 2664, 2337, 2119, 2663, 2777, 901, 869, 489]. Gamma-minimax [849]. Gao [1643]. Garantiekurve [241]. GARCH [2502, 2452, 1084, 2375, 1533, 2420, 2786, 2665, 1375, 2785, 3033]. GARCH-type [1375]. Garcia [2162]. Gauss [1270, 1520, 2719, 1047, 1483]. Gaussian [838, 1165, 1471, 1034, 2885, 2816, 786, 2378, 957, 2833, 2898, 2271, 2393, 1196, 935, 2066, 1021, 2075, 2342, 2770, 2497]. GBP [2197]. Gebietsveränderungen [65]. Gedächtnis [364]. GEE [2157, 2352]. gegebenem [191]. Gegenbauer [2703]. gegenseitige [54]. Gegenwart [25]. Gegenwärtige [89]. gegliederten [336]. Gehaltsstruktur [480]. Geleit [1]. Geleitwort [618]. Gelman [1727]. Gemeinden [293]. Gemeindestatistiken [131]. Gemeinschaften [347]. gene [2294]. General [1139, 1236, 1287, 2967, 2535, 1078, 946, 2850, 1270, 2756, 3012, 1520, 2068, 2295, 1755, 2038, 2255, 2392, 2668, 2434, 2079, 2517, 2719, 3025, 497, 2785, 2504, 1182, 1047, 1778, 842, 1741, 1817, 2440, 679, 2101, 1677, 2544, 1996, 2901, 461]. Generalised [2821, 2764, 390, 2687, 2686]. generalization [1742, 1938, 1475, 926, 776, 1026, 1864]. generalizations [782]. Generalized [2364, 2952, 2712, 2378, 1118, 1594, 2007, 1710, 1284, 875, 2227, 1677, 3021, 1831, 1347, 2730, 924, 2042, 2722, 1197, 2172, 2553, 1889, 1928, 1140, 1053, 1992, 2228, 3004, 1534, 2051, 530, 2862, 228, 2915, 2630, 2512, 1220, 1306, 1299, 1603, 1842, 1949, 1367, 2863, 2896, 2513, 2660, 2682, 2850, 3067, 1956, 3015, 1760, 1768, 1235, 2333, 2688, 2589, 1699, 2844, 1633, 1878, 1621, 2119, 2215, 2726, 3048, 2810, 2894, 1334, 2289, 2858, 2996, 1541, 2641, 2719, 2479, 1344, 1969, 2571, 2713, 2138, 1296, 1656, 1729, 1420, 1468, 741, 1601, 1933, 1445, 1199, 2336, 2610, 2921, 2188, 2052, 1859, 1021, 1210]. generalized [1619, 2031, 2204, 2353, 2361, 2893, 1106, 1098, 666, 1929, 2065, 949, 1715, 1664, 2002]. generating [1335, 1417, 762, 2390, 1711]. generation [712, 2645]. generators [712]. genetic [2713]. Genomics [1568]. Geoffrey [1660]. Geometric [2943, 1718, 2940, 1271, 983, 753, 2099, 2202, 2052, 2350, 2506, 1098, 526]. geometrical [1252]. geometry [1579, 1249]. George [1890]. Geostatistical [2393]. Gerhard [127]. German [231, 571, 658, 6, 41, 73, 176, 401, 533, 568, 148, 288, 20, 262, 246, 165, 131, 208, 254, 113, 365, 508, 491, 378, 610, 4, 15, 551, 596, 8, 139, 649, 631, 413, 218, 230, 244, 48, 585, 362, 434, 99, 44, 121, 105, 243, 293, 338, 51, 182, 364, 385, 545, 136, 69, 366, 626, 52, 72, 190, 261, 433, 636, 161, 677, 235, 236, 9, 255, 232, 281, 252, 367, 5, 146, 233, 368, 356, 514, 521, 425, 479, 369, 137, 103, 272, 171, 442, 34, 59, 337, 529, 138, 344, 181]. German [651, 117, 563, 516, 456, 552, 601, 616, 659, 576, 100, 297, 375, 444, 531, 319, 424, 97, 547, 2055, 678, 317, 259, 199, 674, 62, 63, 106, 164, 220, 388, 532, 512, 153, 75, 298, 345, 399, 397, 441, 147, 155, 188, 539, 579, 627, 123, 394, 467, 590, 673, 156, 206, 2, 3, 16, 19, 29, 565, 40, 42, 49, 46, 88, 109, 167, 192, 336, 370, 360, 395, 402, 536, 515, 581, 591, 509, 140, 207, 612, 180, 203, 613, 621, 642, 185, 124, 320, 506, 499, 412, 645, 122, 81, 292, 309, 335, 183, 35]. German [495, 10, 464, 480, 603, 50, 387, 11, 376, 217, 528, 184, 498, 196, 22, 17, 31, 39, 58, 250, 371, 437, 377, 432, 614, 26, 64, 273, 347, 460, 417, 452, 476, 210, 312, 92, 241, 296, 174, 270, 305, 115, 580, 283, 268, 431, 93, 83, 195, 219, 280, 89, 36, 883, 290, 451, 564, 403, 426, 443, 463, 518, 163, 84, 108, 143, 200, 191, 234, 214, 291, 348, 65, 54, 55, 66, 61, 74, 98, 80, 91, 112, 130, 107, 116, 127, 135, 151, 166, 160, 179, 173, 170, 258]. German [249, 264, 289, 277, 315, 96, 104, 28, 302, 225, 198, 698, 60, 372, 240, 18, 307, 25, 440, 485, 527, 537, 555, 71, 209].

Germany [71, 2487, 130, 885]. Gerß [464]. Gesamt [246]. Gesamteuropäische [90]. Gesamtprodukt [289]. Gesamtproduktion [394]. Gesamtrechnung [130, 268]. Gesamtrechnungen [281]. Gesamtrechnungsdaten [315]. Geschäftspolitik [176]. Geschäftsstatistiken [64]. Geschichte [3]. geschichteter [185, 307]. Geschlechts [81]. Geschlechts- [81]. Gesellschaft [165]. gesellschaftliche [289]. gesellschaftlicher [595, 606]. Gesetz [388]. Gestaltung [320]. Gewalten [50]. Geweke [1504]. Gewerbe [120]. gewerblichen [135, 170]. gezeigt [337, 130]. Ghosh [1547]. Gibbons [1984]. Gibbs [1247, 1443]. Gini [1428, 1402, 2316, 253, 318, 2953, 2318, 2798]. Gini-based [2318]. Ginikoeffizienten [610]. Giovanni [2113]. Girolami [2003]. Giudici [1726]. given [2482, 1564, 2796, 2976]. Glättung [636]. Gleichheit [552]. Gleichverteilung [434]. gleichzeitig [612]. gleitenden [297, 397]. Gliederung [289]. GLIM [771]. Global [1245, 2620, 2215, 3071, 1649, 1896]. globaler [595, 606]. GLS [1108, 992, 910]. GMM [941, 3065, 993, 2861, 1033, 1478]. GMM-Monte [1033]. Goliath [2751]. Gönen [1645]. Good [1566, 1184, 776]. Goodness [152, 942, 959, 2122, 2395, 3052, 2564, 2582, 1728, 2282, 2690, 2885, 1593, 2808, 2904, 2412, 260, 2588, 87, 2240, 2706, 1020, 2000, 1223, 2025, 2948]. Goodness-of-fit [2122, 2395, 3052, 2564, 2582, 2690, 2885, 2808, 2904, 2412, 2588, 2706, 2000, 1223, 2025, 2948]. goodness-of-fit-statistics [1020]. gOS [2438]. GPED [1276, 1417]. Grade [1045]. gradient [1564, 2010]. Gradual [3003, 2319]. gradually [2853]. Graduiertenentwicklung [283]. Graeco [1588]. grafischen [1503]. Graphical [793, 2160]. Graphics [1546]. graphs [2488, 1682]. Gray [2218, 2568]. gray-level [2568]. GRCA [2570]. Greene [1748]. Greenfield [1834]. Grenzen [254]. grid [964]. Größe [394]. großen [388, 26]. Group [1872, 2726, 3050, 2774, 1471, 2291, 1319, 2837, 2691, 2886, 2836, 2312, 2031, 2353, 1072, 1625, 2841, 2071, 2324, 2565]. group-censoring [1625]. grouped [1954, 744, 1114, 2246, 1156, 1319]. growth [243, 1429, 334, 2441, 1915, 1931, 1861]. Grubbs [1415, 1640, 1951]. Grundausstattung [338]. Grunddaten [258]. Grundlage [136]. Grundlagen [208, 442, 464, 60, 440, 883, 200]. Grundmodell [58]. Grundprobleme [424]. Gründung [6]. guarantee [241]. guaranteed [2818, 3024]. Günter [682, 2217, 619]. Günther [1833]. Gusti [1706]. Gut [246]. Gut-Modell [246]. Güterproduktion [147, 155]. Güterverteilungsmatrizen [627].

H [682, 2112, 1661, 471, 2055, 1472, 1544, 833, 912, 787, 1748]. Hadamard [2970, 1134, 1130, 1371]. half [1938, 1620, 1820, 1969]. half-logistic [1620]. half-normal [1938, 1820, 1969]. Halfspace [2668, 2559, 2404]. Hall [2198, 2147, 2219, 2218, 2259, 2197, 2217, 2556, 2160]. Hall/CRC [2198, 2147, 2219, 2218, 2259, 2197, 2217, 2556, 2160]. Hamiltonian [2844]. Hand [1746]. Handbook [799, 1663, 2110]. Handbooks [2110]. handling [2141]. Hankel [2021, 2998]. Hansheng [1685]. Hardcover [2675, 2112, 2259, 2556, 2161]. Harmless [1708]. harmonic [3009]. harmonischen [426]. Harris [926]. Harris-recurrent [926]. Harville [2331]. Hassler [1567, 2108]. hazard [1165, 900, 2222, 820, 1551, 2687, 2686, 1498, 2517, 2516, 1009, 1953, 1817, 1925]. hazards [2813, 2751]. Health [1526, 1935]. healthcare [2144]. heaped [1260]. heaping [2830]. Heath [684]. Heavy [3003, 2249, 2570, 1714, 2575, 1679, 1803, 1951, 2637]. heavy-tailed [2249, 2570, 1714, 2575, 1679, 1803, 1951]. Heckman [1754, 2653]. Hector [2161]. Heft [912]. Hefte [309]. Heidelberg [440, 441, 395, 402]. Heidelberger [571, 441, 395, 402, 440]. Heike [1546]. Heilig [440]. Heinz [2873]. Heisenberg [490]. Heisenberg-like [490]. Hellinger [2994]. Help [7, 14, 23, 32, 38, 566, 45, 56, 67, 76, 85, 94, 101, 110, 118, 125, 133, 141, 149, 158, 168, 177, 186, 193, 201, 212, 222, 238, 247, 256, 266, 275, 284, 294, 303, 310, 321, 332, 341, 350, 357, 373, 383, 391, 400, 408, 418, 427, 435, 446, 457, 465, 473, 481, 488, 492, 501, 510, 519, 534, 541, 549, 557, 575, 582, 592, 604, 609, 617, 634, 655, 676, 697, 718, 739, 760, 781, 802, 823, 844, 865, 886, 894, 903, 913, 922, 931, 940]. Help [952, 962, 973, 986, 996, 1006, 1017, 1029, 1041, 1050, 1061, 1071, 1081, 1091, 1101, 1112, 1126, 1138, 1152, 1164, 1178, 1191, 1204, 1219, 1232, 1244, 1256, 1269, 1281, 1291, 1304, 1316, 1329, 1339, 1352, 1363, 1373, 1382, 1394, 1411, 1427, 1439, 1452, 1473, 1489, 1506, 1527, 1548, 1569, 1590, 1606, 1627, 1646, 1667, 1688, 1709, 1730, 1751, 1772, 1793, 1814, 1835, 1850, 1871, 1892, 1907, 1923, 1944, 1965, 1986, 2730]. Hendrik [1644]. Hennig [2110]. Henry [2218]. Herausgeber [46]. Hermite [2950]. Hessian [2949]. heterogeneity [1657, 2421, 963, 2637]. Heterogeneous [2546, 3054, 1484, 2618, 2074]. Heteroscedastic [1589, 3063, 1840, 2795, 856, 2792, 917, 2036, 1277, 2223]. Heteroscedasticity [1650, 2351, 2061, 2279, 2068]. heteroskedastic [1836, 1192, 1477, 1716, 1877, 1701, 2014]. Heteroskedasticity [758, 2397, 749, 1799, 2389, 2677, 1485]. heuristic [1366, 2925]. hidden [2700, 478, 279, 2575, 2990]. Hierarchical [748, 430, 2567, 1727, 1891, 1591]. Hierarchie [34]. hierarchisch [452]. hierarchisch-agglomeratives [452]. Hierarchische [577]. hierarchischer [476]. Hiernaux [2162]. High [2815, 643, 2807, 2317, 2540, 2905, 2913, 2876, 2864, 2869, 2185, 1794, 2688, 2659, 1942, 2079, 2538, 3006, 2449, 2959, 2789, 2966, 2496, 2927, 2359, 3057, 2650, 3059, 2623, 2031, 2353, 2587, 2639, 2736, 3001, 1929, 2857, 2877, 2071, 2563, 2680, 2414]. high-dimension [2650]. High-dimensional [2815, 2317, 2540, 2913, 2876, 2864, 2869, 2185, 2688, 2659, 1942, 2538, 3006, 2449, 2959, 2496, 2927, 3057, 2031, 2353, 2587, 1929, 2857, 2877, 2071, 2563, 2680, 2414]. high-frequency [2359, 3059]. high-order [2807, 2736]. high-quality [1794]. Higher [1205, 1360, 2165, 2647, 2845, 3060, 2683, 1225]. higher-order [2647, 2845, 3060]. Highest [1467, 1745]. Highly [2010, 2728, 1687]. Hilbert [1710]. Hilfe [148, 585, 103, 140, 35, 191]. Hilfsmittel [246]. Hill [1727]. Hiroshi [1681]. historical [2814]. historische [15]. history [693, 1102, 3]. Hochschulen [338, 64]. Hochschulstatistik [64]. Hochstand [170]. Hofmann [1546]. Hogan [1707]. Höheren [117]. Hölder [2457]. Holm [1354]. Holt [249]. Homogeneity [2906, 1466, 1927, 1155, 1975, 2739, 2085, 1157, 2075, 2399, 2857]. homogeneity-pursuit [2857]. homogeneous [2829]. homoscedasticity [3030]. homoskedastischen [103]. Honda [1154]. hormesis [2262]. hospitalizations [2964]. Hotelling [1377, 2966]. household [2487]. Hron [2433]. Hrsg [682]. Huber [2611]. humidity [2271]. Hungary [232]. hurdle [1240]. Hurst [2544]. hybrid [2321, 2337, 934, 2073, 2530, 1692, 2527, 2810, 2036, 1876, 1817]. hybrids [1978]. hyper [2132, 2453]. hyper-plane [2453]. hyper-Poisson [2132]. hypercube [2595, 2423, 2756, 2936, 2558]. Hypercubic [1421]. hyperellipsoidal [2243]. hyperexponential [1311]. hypergeometric [2508, 1235, 777, 601, 2922, 605]. hypergeometrischen [356]. hypergeometrischer [601]. Hyperkorridore [614]. hyperpolic [2053]. hypersphere [1470]. hyperspherical [2904]. Hypothesentests [60]. Hypotheses [1713, 2508, 329, 2373, 2122, 2961, 735, 1604, 1791, 2280, 715, 1422, 229, 2095, 2599]. Hypothesis [2917, 2134, 2678, 587, 2905, 2508, 2741, 2865, 2392, 3006, 2916, 1939, 871, 2026, 2189, 2080]. hysteretic [3034].

i.i.d [2084]. i.i.d. [1187, 1215, 1973]. IASC [472]. idea [1691]. ideal [629, 486, 685]. idealen [485, 527, 537]. identical [1104, 2938]. identically [752, 2322]. identifiable [2833]. Identification [1911, 1847, 1251, 1222, 2467, 546, 638, 2846, 868, 205, 2361, 2476, 2343]. Identifying [1808, 2720]. identische [372]. identities [2736]. IFR [2534]. ignorable [2032, 2605]. Igor [2217]. ihre [44, 211, 627, 1503, 84, 315, 372]. ihrem [514, 521, 515]. ihren [281, 3]. II [2081, 821, 638, 946, 1853, 1620, 1671, 2522, 2087, 2000, 2346, 477]. III [1476]. illustrated [1588]. illustration [882, 1597]. illustrations [1170]. ILR [1214]. im [491, 413, 48, 171, 138, 504, 319, 547, 654, 394, 3, 642, 184, 66, 112, 120, 277, 315]. image [2568]. images [2467, 2125]. imbalanced [2750]. impact [2248, 1687]. imperfect [2676]. Implementing [1262]. Implications [803]. Implicit [2895]. implied [404, 871]. imprecise [282, 1145, 1105]. improbable [1687]. improve [1782]. Improved [2397, 2013, 2791, 1461, 1654, 2899, 1701, 2063, 1160, 1297, 2317, 872, 2987, 1501, 700].

Improvement [1370, 2958]. improvements [2345]. Improving [1261]. impulse [1905]. Imputation [1239, 780, 1460]. impute [2842]. imputed [2635, 2533]. inaccuracy [2387, 2022, 2946]. inactivity [2165]. inadequacy [2643]. INAR [2383, 1824, 2870, 2542, 2153, 943, 2332, 2086, 2566, 2350, 2506, 2323]. including [3018, 2733, 2833]. inclusion [625, 600, 908]. Incoherent [2236]. income [1030, 1904, 694, 2692, 861]. incomplete [2303, 2576, 216, 2985, 2140, 774, 1313]. incomplete-block [2303, 2576]. incorporating [2488]. increasing [2949]. increasing-Hessian [2949]. increment [827]. independence [630, 2790, 2744, 2491, 2754, 3068, 2881, 920, 2266, 2650, 3001, 2545]. independent [1613, 752, 1534, 1854, 1882, 2389, 2644, 1522, 2663, 827, 1518, 1285, 1863]. indeterminacy [1904]. Index [63, 485, 527, 537, 1428, 2603, 623, 660, 553, 629, 644, 717, 2532, 486, 2245, 2958, 2068, 445, 2170, 2956, 2662, 2837, 3013, 2345, 2917, 2927, 661, 523, 574, 1285, 2618, 2933, 2835, 2442, 2547, 685, 526]. index-addendum [661]. indexes [522, 507, 644, 2969, 524]. Indexiffern [54]. Indexmodelle [659]. Indexproblems [463]. India [2491, 27]. indicators [688, 2560, 195, 2511]. indices [556, 496, 1765, 727, 1648, 686, 525, 767, 810, 2481, 918, 612, 574, 2692, 1677, 564]. Indikatoren [235, 195]. Indirect [1170, 2779]. indirekten [100]. indirekter [75]. Individual [1571, 2684, 1154, 1198, 2009, 2760]. Indizes [612]. induced [727, 2849, 2534, 2778]. Inductive [404]. industrial [2272, 2769, 421]. Industrie [480]. Industrie- [480]. industrielle [66]. industriellen [72, 147, 155, 579, 65, 54]. industrieller [148]. Industrien [220]. Industrieproduktion [105]. industriestatistischer [41].

Industriestruktur [72]. industry [1873, 1726, 1834]. Inequalities [880, 1010, 2180, 2156, 1795]. Inequality [767, 918, 3045, 2465, 701, 1087, 993, 1542, 694, 2995, 945, 2692, 2134]. Inference [2046, 1648, 2724, 2960, 934, 2891, 1551, 2416, 2023, 822, 1811, 1570, 1640, 1913, 1876, 3010, 2014, 1529, 3017, 2903, 2722, 2643, 1464, 648, 687, 2234, 2535, 2975, 1558, 572, 620, 226, 2620, 1738, 1046, 313, 1185, 2249, 2179, 2412, 692, 704, 2503, 1514, 2883, 1649, 538, 699, 770, 2461, 2695, 1732, 227, 339, 343, 2333, 2073, 1492, 1454, 2853, 1994, 2878, 2502, 2786, 808, 1816, 2050, 1296, 353, 382, 2238, 2888, 1786, 2223, 340, 2188, 1086, 2302, 3034, 869, 1388, 2640, 2746, 1636, 2778, 2334, 1869]. inferences [2047, 287, 1867, 2007, 420, 3026, 2929, 1984]. Inferential [2674]. Inferenz [371]. inferior [2903]. infinite [1299, 2159, 889, 1194]. infinite-armed [1299]. infinitely [629, 2943, 2939]. Inflated [1598, 2870, 2490, 2175, 2548, 2132]. inflation [2899, 2323, 2168]. Influence [2351, 1415, 1951, 2693, 2225, 1193, 2296, 1934, 1064, 1714, 1066, 2066, 3033]. influential [1971, 788]. Information [596, 1507, 594, 531, 2691, 1146, 2761, 2948, 2621, 1318, 953, 462, 3064, 744, 602, 811, 216, 1565, 1514, 152, 1995, 171, 1283, 1669, 2688, 1982, 1320, 1875, 543, 1654, 205, 1655, 1312, 2000, 2138, 2874, 1940, 774, 1641, 1866, 3027, 1038, 797, 2064, 2139, 2932, 651]. Information-based [2761]. information-theoretical [205]. informational [1237]. Informationen [545]. Informationsmaßen [545]. Informationsprogrammen [121]. Informationstheoretisch [262]. Informationstheoretische [206]. Informative [404, 2772, 1913, 2888, 2287]. informatively [2868]. informativity [462]. informed [2300]. infrared [2735]. Infrastruktur [337]. Infrastrukturen [277]. Ingeborg [626]. Ingenieurwissenschaftler [2109]. inhomogeneous [2749]. INID [1205, 1360]. initial [1810, 1294]. inner [1812, 668]. innerhalb [362]. innovational [1720]. innovations [2249, 1194, 1679]. Inoue [1749]. Input [41, 148, 123, 336, 370, 124, 128, 13, 417, 452, 36, 290, 451, 163, 28, 1537, 33, 421, 2310, 37, 154, 271]. Input-output [128, 1537, 33, 421, 37, 154, 271]. Inputs [283]. insbesondere [281, 440]. insights [2771, 2706, 3000]. inspection [928, 657, 929, 1440, 1612, 2305, 2267]. inspired [2824]. Instandhaltung [20]. Instituts [6, 591]. Instrument [673]. Instrumental [2832, 1799, 2610]. insurance [1935]. integer [2355, 2993, 2682, 1479, 2420, 1877, 1825, 2475, 2254]. integer-valued [2993, 2682, 1479, 2420, 1877, 1825, 2475, 2254]. integers [2606]. Integral [507, 1139, 1143]. integrals [1142, 1458]. Integrated [1705, 925, 2958, 2487, 2531, 2866, 2511]. Integrating [2856]. Integration [51, 1567, 33, 1897, 888, 927, 3006, 1225, 2610, 1074, 2736]. integrative [2857]. Intensitätsstufen [84]. intensities [2028, 834]. intensity [2525, 2911, 798]. inter [1464, 579]. inter-industriellen [579]. inter-record [1464]. Interaction [1347, 1809, 1554, 1476]. interactions [2308, 2597, 2927, 325, 326]. Interaktionsanalyse [577]. interarrival [1959]. interblock [216]. intercept [1320, 1630, 1324, 2419]. interest [1084, 1368, 2591, 2867, 885]. interference [1576]. intergroup [216]. International [421, 2, 3, 11, 28, 472, 1570, 1686, 632]. internationale [8, 105, 370, 173]. Internationalen [591, 4, 231, 5, 2, 3, 603, 277, 225]. internationaler [139, 10, 9, 290, 55]. interpoint [2214]. Interpolation [636]. Interpretation [2055, 555, 1957, 507, 2684, 1252, 776, 42, 1887, 526]. Interpreting [355, 1703]. interquantile [1756]. interregionalen [480]. intersection [2741]. intersection-union [2741]. Interval [2787, 1366, 2411, 2406, 1159, 1114, 1722, 1257, 2906, 2305, 2961, 2265, 1798, 2201, 1386, 1609, 970, 2868, 2715]. interval-censored [1257, 2868]. interval-valued [2406]. intervalldatierten [677]. Intervallschätzung [307]. intervals [1340, 1233, 1812, 1158, 2402, 1012, 2707, 1403, 1972, 1167, 668, 1638, 1221, 2666, 1853, 723, 1690, 2133, 2032, 260, 798, 245, 2698, 346, 2683, 1543, 773, 690, 2972, 2063, 2269, 2612, 229, 1968, 1677]. intra [216, 396, 1031, 1273]. intra-class [396, 1031, 1273]. intraclass [2567, 2441]. intrinsic [3044, 2010]. Introducing [1353]. Introduction [1660, 2182, 1139, 2108, 1983, 1833, 1547, 1893, 2161, 1704, 1486, 1566, 1724]. introductory [783]. Invariant [1179, 2207, 1412, 2905, 598, 1474, 1583, 1249, 2126]. Invarianz [378]. inventory [964]. Inverse [2120, 2066, 838, 1165, 2722, 1471, 1693, 2474, 786, 1448, 957, 2962, 2255, 2281, 2922, 1779, 1920, 2201, 1843, 2507, 1021, 2142, 2528, 2075, 2342, 2897, 461, 1715]. Inverse-Gaussian [2066]. Inversen [579]. Inversion [2671, 730]. Inversion-free [2671]. inverted [2960, 1376]. invertibility [415]. Investigating [2486]. Investitionsentscheidungen [516]. Investitionstätigkeit [100]. Investitionsvolumens [98]. investment [1808, 98]. involving [300, 352]. irregularities [772]. irregularly [2577]. ISBN [2675, 2019, 2198, 2147, 2219, 2112, 2218, 2259, 2182, 2197, 2217, 2199, 2556, 2160, 2161]. isomorphism [2720]. isotonic [1760, 1947]. Issue [2369, 2116, 1139, 2299, 2454, 2757]. issues [2920, 2530, 2374, 1157]. item [2366]. iterativ [563]. iterativ-multiplen [563].

Iterative [2124, 2713]. iterativer [191]. IV [1900, 1476]. Izenman [1832, 1723].

J [771, 1707, 2182, 1746, 2056, 2674, 1526, 1545, 1568, 789]. J. [2105]. Jackknife [1007, 2136, 2809, 3014, 1971, 807, 2692, 949]. jackknife-after-bootstrap [1971]. jackknifed [1776]. Jahrbuch [166]. Jahrbücher [53]. Jahrbücher-zwei [53]. Jahre [66, 367]. Jahresabschluß [240]. Jahrgang [46]. Jahrhunderts [464]. Jambu [840]. James [2106, 1704, 2088, 1686]. Jannach [2217]. Japan [272]. Japanese [1429]. Japans [135]. Jason [2674]. Jayanta [1547]. Jean [1984]. Jeffrey [1962]. Jelke [2219]. Jennifer [1727]. Jensen [1682]. Jerez [2162]. Jim [1661, 1683]. Jiti [1643]. Joarder [1472, 1544]. John [1664, 1665, 323, 2197, 971, 1504]. joinpoint [2121]. Joint [975, 2210, 3005, 1981, 2271, 730, 2481, 1855]. Jones [1724, 1774]. Jorion [1725]. Jörn [2104, 1708]. Jörn-Steffen [2104]. Jose [2162, 2106]. Joseph [1707]. Joshua [1708, 2104]. judgement [2425]. Judgment [2614, 2251]. Julian [1832, 1723]. Julio [1906]. jump [1898, 877, 1171]. jump-preserving [1171]. jumps [1899]. Jun [1685]. Jürgen [2432].

K. [1663, 2392]. Kahneman [1963]. Kakosyan [732]. Kalman [2838, 2028, 1914, 720, 2767]. Kantorovich [1010]. Kantorovich-type [1010]. kapitalistischen [176, 65]. kapitalistischer [190]. Kaplan [926]. Karel [2433]. Katsuto [2349]. Kaufkraftparitäten [467]. Kendall [2853, 3027]. Kenngrößen [72]. Kenntnis [579]. Kernel [1340, 2354, 3002, 2140, 877, 2557, 1595, 2035, 846, 1710, 2418, 2702, 2984, 2141, 2476, 1752, 2669]. kernels [2574]. Ketten [59, 252, 387]. Kevlar [1957]. Kibria [1999]. Kimball [129].

Klarstellung [433]. Klasse [545, 181]. klassierten [610]. Klassifikation [514, 521, 515]. Klassifikationsproblemen [425]. Klassifikationsverfahren [452, 476]. Klassifizierung [72, 198]. Klebanov [732]. kleinem [506]. kleinen [26]. Kleinst [217]. Knight [2199]. knots [2852]. known [1461, 999, 1025, 1678, 1811, 1769, 1297, 1549]. Koeffizienten [529]. Koeffizientenschätzung [573]. kohärenten [365]. Kohlenstatistik [90]. Kolmogoroff [444]. Kolmogorov [998, 1989, 1025, 1036]. Komponenten [297]. Komponentenmodelle [442]. Konfidenzintervalle [375]. Konfidenzintervallen [506]. Koning [1661]. Konjunktur [246]. Konjunktur- [246]. Konjunkturelle [367]. konjunktureller [235, 195]. Konjunkturforschung [25]. Konsequenz [532]. konstanter [297]. Konstanz [367]. Konstruktion [673]. Konstruktionsverfahren [412]. Konsumentenkaufkraft [312]. Konsumfunktion [367]. Konsumtionsmitteln [289]. kontemporär [642]. Kontingenztabellen [539]. Kontrolle [335]. Kontrollrechnung [585]. Konzentration [218, 230, 244]. Konzentrationsmaße [381]. Konzentrationsmessung [302]. Konzentrationsphänomenen [365]. Konzept [258]. Konzeption [320]. Koordination [64]. Korrelationszahl [296]. korrelierten [642]. Korrespondenzanalyse [569]. Krämer [1503]. kriging [2590]. Krishnaiah [799]. Krishnamoorthy [1663]. Kriterien [29]. Kritik [612, 437, 431, 320, 432, 431]. Kritische [398, 380, 399, 337, 555, 437]. Kriz [1018]. Kronecker [1580, 2650]. Krzanowski [1746]. Küchenhoff [787]. Kühn [480]. Kullback [2000]. Kulturausgaben [293]. Kumaraswamy [2940, 2213, 2386, 1844, 1857, 1955]. kumulative [255]. Kurtosis [1461, 2822, 1821, 1555, 801, 1549]. kurzer [35]. kybernetischer [198].

L [1749, 1870, 2218, 1962, 805]. Laan [1568]. Lack [2439, 3029]. lack-of-fitness [3029]. LAD [2394, 2639]. Lag [2382, 664, 1936, 448, 588, 573]. Lag-Modellen [573]. Lagerhaltung [59]. Lagetheorem [368]. Länder [293, 336, 54, 190, 62, 75, 290, 107, 116]. Ländern [218, 230, 244, 72, 514, 521, 515, 98, 34, 124, 65, 91]. Landes [131]. Landes- [131]. landwirtschaftliche [107]. Laplace [1288, 1653, 946, 1507, 1599, 1465, 1271, 2374, 2123, 1804, 2265, 1238, 2211]. Large [1874, 1546, 316, 2541, 2823, 3030, 2671, 2632, 1616, 2612, 1988, 2330, 2480, 2471, 2661]. large-dimensional [2480]. larger [1572, 2987]. largest [860]. Lars [1503]. Laspeyres [727]. lasso [2312, 1930, 2291, 2394, 2886, 2031, 2353, 2071]. latent [2629, 942, 959, 2844, 1170]. Latin [2595, 2423, 2756, 1588, 2936, 2558]. lattice [1743, 1652]. lattices [500]. laufenden [183]. law [2056, 3072]. laws [1556, 897]. LBI [916]. lead [1355]. League [3]. Learning [1149, 1723, 1832, 3041, 2769, 2568, 2977, 2907, 3051, 3005, 1747, 1943, 2770, 3016, 2367, 2431, 2003]. Least [2950, 762, 2242, 2388, 2867, 2544, 1637, 669, 727, 1306, 1784, 2700, 1695, 2135, 334, 2403, 1207, 848, 643, 1621, 2289, 2641, 1480, 2860, 2288, 475, 1044, 1931, 825, 908, 1346, 1131, 2518]. least-squares [1784, 2288, 1931, 1346]. Lebensdauerverteilungen [498]. Lebensniveau [231]. Lebensniveauunterschiede [81]. LeCam [1174]. lecture [413]. Lee [1664, 2473]. left [2035, 2734, 2445, 2412, 2976, 2057, 743]. left-truncated [2412, 2976, 2057, 743]. Legendre [2169]. Lehmann [667, 1115]. Lehrbücher [192]. Lehrbüchern [320].

Leibler [2000]. length [838, 1023, 1936, 2517, 1322, 1712, 889, 876, 679]. length-biased [838, 2517]. length-biasing [889]. lengths [2210, 1981]. Lenkungsmethoden [232]. Leone [1761]. Leontief [579, 394]. Lernmodelle [136]. lesion [2467]. Let [1903]. Letter [1502, 1773, 1408, 1774]. Level [2654, 1897, 2520, 243, 2633, 2307, 2308, 2568, 2137, 2212, 2448, 2632, 2624, 1469, 2613, 1987, 1485, 1698, 2585, 2357]. Level-augmented [2654]. level-effect [2520]. levels [2427, 656]. leverage [1966]. Lévy [978, 2388]. Li [1890]. liars [2779]. Lieferungen [579]. Life [1634, 1959, 1692, 1120, 1335, 1365, 731, 710, 721, 756, 1067, 898, 1055, 1220, 2960, 988, 2073, 793, 808, 2377, 1623, 1251, 990, 1953, 1786]. life-testing [1055]. lifetime [1341, 1331, 2208, 1972, 1842, 2435, 1848, 2409, 2534, 1997, 1535, 2202, 1817, 2723, 2746, 1958]. lifetimes [2707, 1753, 1692, 2235]. light [507]. like [490, 2075]. Likelihood [2222, 137, 2871, 460, 1056, 1786, 1422, 2643, 2397, 857, 1928, 2535, 448, 2293, 2626, 1412, 3064, 628, 711, 2552, 1787, 2515, 946, 2464, 1949, 1909, 1867, 3028, 2179, 2412, 704, 1840, 699, 1355, 835, 2451, 1881, 736, 881, 227, 343, 2478, 2499, 2229, 2537, 2754, 3020, 2415, 2976, 2659, 2745, 1405, 2434, 3048, 1284, 1696, 2136, 2422, 2878, 2854, 2373, 2809, 330, 2813, 1521, 1964, 1919, 2934, 1701, 2711, 1194, 2874, 2729, 2615, 879, 1196, 2888, 2929, 1416, 2231, 2341, 3014, 3059, 2301, 2963, 1609, 2401, 3034, 1792, 2692, 3010, 2640, 2095, 2669, 2011, 2746, 1958, 1758, 1664]. likelihood-based [2643, 2422]. likelihood-ratio [3064, 711, 2745, 2711]. Likelihoodfunktion [44]. Limit [1605, 2619, 2849, 2747, 1973]. limited [2874]. Limiting [1088, 2754, 2986, 3037, 1581, 2479, 1585]. Limits [517, 705, 1319, 919, 2034, 1769]. Lindley [2234]. Lindsay [2151]. Lindsey [771]. Linear [162, 1879, 1660, 1649, 2143, 2447, 316, 2764, 1570, 411, 2211, 3021, 2364, 1259, 1731, 1999, 2054, 953, 1197, 2553, 2535, 974, 2540, 1573, 2694, 1992, 853, 1023, 1600, 2931, 2902, 2656, 1971, 1412, 2391, 2734, 2915, 2118, 2630, 1403, 2512, 2905, 2712, 242, 1078, 891, 1306, 2532, 2964, 2689, 1949, 1367, 2749, 2465, 2550, 1961, 278, 1664, 2896, 2513, 2225, 3047, 597, 1840, 2815, 2503, 1118, 1575, 1193, 2985, 1819, 2232, 2159, 3040, 1108, 1474, 2840, 1457, 1154, 145, 172, 1941, 2033, 2408, 2499, 2567, 1364, 379, 788, 2090, 2537, 2988, 1998, 1370, 2589, 882, 1207, 2001, 2295]. linear [3068, 1562, 2131, 2844, 559, 993, 2865, 1320, 1031, 1357, 848, 716, 543, 643, 790, 1170, 2119, 977, 2392, 1652, 1252, 2726, 2290, 2002, 1614, 1788, 2283, 2538, 2647, 3060, 3006, 3048, 2410, 1066, 2136, 2611, 2894, 3013, 1334, 2252, 2861, 2449, 2858, 2996, 1541, 1582, 1983, 2431, 1441, 1531, 1583, 2719, 1391, 2683, 497, 3031, 917, 1342, 2017, 2571, 2713, 2103, 2753, 1296, 871, 1778, 122, 2049, 2628, 2714, 2309, 353, 382, 2102, 1117, 1626, 741, 475, 709, 984, 1068, 1855, 3057, 1924, 1932, 1931, 825, 2767, 2921, 1618, 2188, 560, 1741, 2156, 1037, 1172, 2328, 970, 797, 1359, 2031, 2064, 2353]. linear [2361, 2436, 2587, 2601, 2933, 2539, 1792, 2450, 1075, 2405, 2893, 1763, 189, 229, 2256, 2342, 1929, 2227, 2338, 1563, 2480, 2442, 2709, 2761, 3036, 2429, 1818, 2077, 2071, 2154, 2497, 2547, 2669, 1622, 3007, 1996, 1636, 2320, 1758, 728, 2060, 2661, 2407, 2331]. lineare [122, 270, 36, 883]. linearen [529, 547, 217]. linearer [594, 136, 651]. Linearity [2080, 2495, 334]. linearization [1156]. Linearkombination [96, 104]. linearly [286, 343]. lines [1377]. LINEX [1258, 1629, 1087, 875, 1011]. link [2356, 863, 1367, 2033, 2740, 2788, 2893]. linkage [1323]. linked [1624]. Linnik [1800]. LINSTAT’2008 [1570]. Lipovetsky [2486]. Lipschitz [3062]. Lipschitz-continuous [3062]. LISE [137, 137]. LISE- [137]. list [1735]. lists [2968, 3069]. Little [789]. Liu [1776, 2225, 3047, 1370, 1993, 2015, 2119, 1788, 2328, 1563]. Liu-type [1776, 1993, 2015, 1788]. lives [1288, 1711]. living [553, 727]. LLMM [2211]. Llorente [1906]. load [2829, 469, 1875]. load-sharing [2829]. load-strength [469]. Local [2693, 2656, 1221, 2749, 2351, 2225, 2296, 1934, 1942, 588, 624, 2342, 2237, 2382, 2902, 2734, 2832, 2220, 1911, 1649, 2589, 2362, 1714, 1066, 1596, 1157, 3058, 2204]. Locally [2988, 2894, 2309, 978, 2268, 2619, 1583]. location [1330, 1860, 1540, 1083, 162, 242, 1107, 890, 538, 2987, 1536, 2069, 667, 2466, 850, 2504, 2766, 1823, 1160]. location-scale [1330, 538, 2466]. location-scale-rotation [2466]. locations [1093, 1008]. Log [745, 2405, 1023, 629, 644, 354, 2712, 215, 2724, 878, 1819, 2478, 1319, 559, 2415, 2976, 1057, 2093, 3024, 2150, 1841, 2971, 2264, 1608, 2497]. log-beta [1841]. log-change [629, 644]. log-gamma [2712, 215]. Log-linear [2405, 1023, 559, 2497]. log-logistic [2478, 1319, 2976, 1057, 2093]. log-normal [354, 1819, 2415, 2264]. Log-ratios [745]. log-survival [878]. logarithmic [628, 745, 2101, 2530]. logging [2264]. Logik [99]. Logistic [2328, 1860, 2053, 3053, 162, 242, 2776, 2924, 1457, 316, 2478, 334, 1319, 1620, 2671, 2976, 868, 1057, 2093, 1531, 2124, 1345, 1238, 2551, 2877]. logit [1077, 1787, 1095, 3055, 1673, 787]. Logit- [787]. logit-skew-normal [3055]. loglinear [1441]. lognormal [2776, 852, 2030, 1443, 489]. Lohn [499, 480]. Lohn- [480]. Lohnstatistik [464]. Lohntheorie [246]. Lomax [1104]. Long [1282, 1127, 2012, 2638, 1153, 1737, 2891, 2683, 1528, 817, 2882, 1783, 1359, 1595, 1979, 2439, 1622]. long-memory [2638, 1153, 2683, 2882]. long-range [2012, 1783]. long-term [817]. Longitudinal [2773, 2186, 1707, 1495, 864, 2596, 2245, 1870, 2314, 2649, 3013, 2802, 2577, 2300, 1801, 2557, 1524, 2361, 2436, 2601, 2065, 2029]. look [2694, 860]. loss [1258, 2207, 1042, 1552, 1447, 1629, 1076, 278, 1211, 1264, 1087, 3015, 1755, 1602, 2611, 2719, 2150, 875, 1011, 956, 2504, 1309, 2270]. loss-functions [278]. Lösung [425]. low [2945, 1045, 1858]. lower [2013, 2438, 2222, 2572, 2473, 2710, 2448, 1308, 2039, 2980]. LPI [490]. LPI-concept [490]. LPRE [2913]. LPRE-based [2913]. LR [2908]. LRD [2297]. LS [1845, 2310, 2340]. LSE [2398]. LTPD [1612]. LTRC [2515]. Lungenkrebssterbefälle [115]. Lusin [1148]. Luxembourg [586]. Lyle [2238].

M [2197, 1545, 840, 2110, 1813, 816, 798]. M. [2112, 1774]. M/E [798]. M/G/1 [816]. Machin [1526]. machine [2367, 2769, 2568, 2431, 2003, 818, 1943]. machines [2945]. MacKinleay [971]. Macroeconomic [2019]. Made [2147]. Madsen [1644]. Mahalanobis [2578, 1710]. majority [2344]. make [1168]. making [1140, 1142, 795, 1145]. makroökonometrischer [533]. makroökonomischen [367]. makroökonomischer [156]. malignant [2467]. malnutrition [2491]. malus [1550]. Managing [1725]. Mangion [2556]. Manifold [1723, 1832]. manifolds [3071]. Manisera [2675]. Mann [3018]. manpower [154]. Manuel [2106]. manufactured [2267]. many [629, 2173]. mapping [2294]. Marc [1525]. Marginal [448, 2799, 2464, 770, 343, 1913, 2099, 2623, 3010]. marginalized [2314]. marginally [2586]. marginals [936, 1742, 1491, 1271, 943, 2190, 2319]. Maria [2106]. Marica [2675]. Mark [2003, 1568]. market [954, 1533, 232]. market-steering [232]. markets [738, 971]. Markoff [252]. Markoffsche [59]. Markov [387, 924, 2452, 811, 2700, 1185, 1298, 1734, 1270, 1520, 30, 2177, 1898, 926, 1116, 2719, 2672, 2050, 1047, 2575, 944, 2189, 1483, 2990]. Markovian [2275, 306, 478]. Markt [232]. Markt- [232]. Marktentscheidungen [176]. Marshal [2278]. Marshall [1615, 1856, 1475, 1651, 1733]. Marsland [1943]. Martin [1486, 1546]. Martínez [2182, 1962]. martingale [2083]. Maschinenbau [138]. Masse [366]. massive [2983, 2593, 2681, 3058, 2750, 3039, 2396]. Mastering [2104]. matching [2856, 2415, 1026, 2470]. materiellen [258]. Matérn [2828]. Mathematical [787, 2145, 2298]. Mathematics [2555, 1102]. Mathematiker [2109]. Mathematische [2109, 209]. mathematischen [372]. MATLAB [1962]. Matloff [2431]. matrices [3017, 2970, 2540, 1577, 379, 1580, 2617, 1716, 2811, 1130, 2156, 763, 1988, 2330, 2932, 2661, 2259]. Matrix [712, 2111, 579, 1642, 2921, 1795, 1042, 1330, 1507, 2609, 1949, 2316, 2331, 1270, 777, 1621, 1942, 2909, 1010, 2873, 2998, 2641, 540, 824, 2945, 2084, 1626, 605, 680, 2507, 779, 1444, 2587, 2563, 2948, 461, 2991]. Matrix-variate [2921]. Matrizen [35]. Matthias [2433]. Max [596, 1141, 2630, 832, 1651, 1983]. Max-min [1141, 1651]. maxeminoptimaler [658]. maxima [1401, 2052]. Maximal [1412, 1635, 2702]. maximin [2795, 2852]. maximising [652]. maximizers [857]. maximizing [2127]. Maximum [1928, 628, 946, 1881, 137, 2478, 2976, 512, 2871, 1194, 1815, 2669, 2047, 2975, 1787, 1949, 2179, 881, 2229, 2537, 3048, 920, 330, 1964, 1919, 2263, 1701, 2874, 2729, 879, 2341, 2301, 2963, 2401]. Maxwell [2827, 1764, 1497]. Maxwell-Poisson [1764]. MBT [1981]. McElroy [2554]. MCMC [1375]. MD [2340]. MDPs [2907]. Mean [721, 756, 1153, 1414, 2787, 1678, 2420, 779, 2060, 1529, 3017, 731, 1224, 2008, 1447, 1471, 1012, 1067, 2926, 3062, 2620, 1002, 1132, 2776, 1007, 991, 786, 1264, 2010, 2296, 2596, 988, 860, 339, 809, 1807, 1532, 2286, 245, 1129, 2636, 2415, 748, 2325, 3006, 745, 2809, 1371, 1997, 1623, 2084, 2297, 1378, 1251, 2070, 1786, 1783, 1122, 1228, 1442, 1632, 2980, 1038, 1432, 797, 2884, 2183, 1297, 2531, 2134, 2798, 2919, 2544, 2723, 2680, 2715, 2778, 580]. mean-reverting [3062]. mean-shift [2919]. Mean-variance [1414, 2296]. Means [197, 1158, 1395, 1540, 2905, 430, 1638, 1867, 2133, 598, 2805, 1206, 1357, 2290, 1994, 776, 1308, 1782, 2196, 2954, 2524, 301, 271, 2216]. measure [1053, 3004, 1851, 2679, 2573, 366, 1355, 3012, 2022, 2241, 2946, 1826, 2711]. Measurement [744, 2391, 228, 3047, 2503, 2232, 1369, 2106, 1272, 1630, 1714, 2804, 817, 1641, 1924, 2583, 2921, 2188, 2064, 2476, 2933, 1689, 1818, 1803, 2485, 2547, 2334]. measurements [1831]. Measures [1852, 1639, 801, 1880, 2621, 997, 1141, 2816, 1976, 2738, 2387, 2695, 1675, 334, 942, 959, 2710, 559, 2038, 420, 1555, 2192, 1580, 2668, 866, 1121, 2127, 2311, 1798, 842, 2336, 2676, 560, 1119]. Measuring [727, 1246, 559, 2739, 2413]. mechanism [2535]. Media [413]. Median [278, 2474, 1207, 2404, 1844, 1384]. median-dispersion [1844]. Medians [368, 449]. mediation [2406, 2802]. medical [2746, 1526]. Medienverbund [413]. medizinisch [377, 432]. medizinisch-wissenschaftlicher [432]. medizinischer [431]. mehrdimensionalen [528]. Mehrwertsteuer [281]. Mehryar [2367]. Meier [926]. Meila [2110]. Meinung [50]. Melamed [732]. melanoma [2467]. membership [1045]. memoriam [682, 323]. memory [1127, 2638, 1153, 1737, 2891, 2683, 1528, 2882, 879, 1282, 1359, 1595, 1979, 2439, 1622]. memoryless [414]. Mences [512]. Mengen [504]. Mengenindexproblems [443].

Mengentheoretischen [198]. Menges [440, 619, 682]. Merkmale [674, 616]. Messen [539]. Messens [376, 31]. Messung [262, 365, 337, 100, 319, 140, 203, 174, 348, 160]. Meßwerten [649]. Meta [1155, 1927, 1333, 2607]. Meta-analysis [1155, 1927]. meta-analytic [1333]. meta-ellipticity [2607]. method [2830, 2047, 556, 853, 2952, 1971, 1093, 1828, 3044, 1766, 2380, 1156, 2671, 672, 2726, 2394, 2632, 2649, 2289, 421, 2966, 892, 2141, 2096, 2138, 2882, 1323, 2650, 1460, 2067, 2528, 2183, 1968, 1839, 2910, 2480, 1758, 2407]. method-of-moments [2952]. Methode [100, 297, 156, 512]. Methoden [9, 467, 377, 432, 431, 10]. Methodik [73]. Methodische [464]. Methodologie [536]. methodologies [1637]. methodology [2869]. Methods [2614, 2555, 2345, 1566, 1813, 1999, 737, 448, 2975, 1092, 2552, 1930, 2733, 2763, 1019, 243, 2819, 2664, 770, 232, 2206, 1247, 2144, 2455, 2703, 1872, 2417, 445, 729, 1597, 1581, 3050, 1916, 2124, 2934, 1796, 2297, 1346, 324, 2162, 1770, 1643, 1547, 1891, 2110]. metrics [2104]. mex [832]. Michael [1707, 1526]. Microsoft [1566]. migration [1686]. Miguel [2162]. Mike [1869]. mild [2398]. Miller [2056]. Million [1546]. Mills [3023]. min [1141, 2630, 1651, 596]. min-Entscheidungsprinzip [596]. min-max [2630]. minification [1603, 2472, 1468]. minima [1401, 2052]. Minimax [1780, 2907, 2280, 2524, 715, 1763, 953, 722, 849, 1169, 791, 1332, 2538, 2150, 2306, 1769]. minimax-criteria [2306]. Minimierung [234]. minimization [2766]. minimizing [2783, 1511]. Minimum [2911, 386, 2385, 2522, 144, 1441, 1345, 1305, 1347, 639, 2552, 2195, 723, 1118, 1649, 2151, 2484, 2481, 1398, 2363, 1420, 1779, 3027, 1896, 1689]. minimum-variance [1896]. Mining [1642, 1726, 1890]. misclassification [1777]. misclassified [1321]. Mises [2069, 895, 2783, 2743, 1878, 1917]. Mises-type [2783]. misleading [2084]. Missing [2662, 3021, 2603, 2535, 2620, 2354, 2560, 1501, 2977, 1767, 2581, 2261, 2912, 2422, 2094, 1248, 2141, 2564, 1535, 2929, 2231, 789, 1037, 1209, 1460, 2134, 2835, 2062, 2370, 2545, 2320, 2511, 1707]. missingness [2032, 2605]. misspecific [1127]. misspecification [1626]. Misspecified [963, 905]. Mithat [1645]. mitigate [2708]. mitochondrial [2563]. Mitsunori [1890]. Mitte [367, 464]. Mittels [426]. Mixed [2383, 1117, 1305, 2364, 415, 2054, 2391, 2630, 937, 1574, 2155, 2082, 1495, 2307, 2451, 1474, 881, 2033, 2408, 1551, 2333, 2844, 2038, 2446, 2120, 1614, 2632, 1583, 1009, 2899, 2190, 2836, 1798, 1587, 1626, 2557, 1210, 1839, 2227, 2835, 2211, 1818, 461, 2407]. mixed-frequency [2446, 2899]. mixed-model [2038]. mixing [1259, 1840, 1766, 1511, 2937, 2418, 1884, 3059, 2459, 1917, 1926]. Mixture [3000, 2831, 1736, 2940, 3054, 1165, 2008, 1928, 2826, 1507, 864, 1624, 957, 2006, 2941, 2875, 3020, 1948, 2878, 1991, 1466, 3026, 2085, 2751, 2224, 1838, 1056, 1843, 2990, 1433, 2585, 1822, 2344]. mixture-of-mixture [2006]. Mixtures [2426, 2600, 1120, 2195, 1961, 2143, 1847, 2898, 1719, 3055, 2509, 1098]. ML [775]. MLE [1088, 2520, 2957, 2398]. MLEs [1860]. MM [2380]. MNAR [2535]. mODa [2757]. mODa12 [2299]. modal [2011]. modality [2737]. mode [2035]. Modeforschung [254]. Model [602, 1930, 597, 355, 197, 2299, 2622, 2400, 2623, 2990, 2994, 2631, 2625, 2471, 3054, 2799, 2452, 1353, 2437, 2117, 1731, 1999, 2054, 953, 1197, 2355, 1564, 2535, 2694, 669, 761, 1901, 1921, 2791, 2035, 2965, 1147, 2667, 758, 900, 570, 2829, 469, 1574, 2155, 1631, 1764, 3053, 762, 2082, 1078, 891, 1034, 1787, 2083, 2515, 2773, 3065, 2532, 1695, 2689, 430, 1063, 1095, 1935, 2520, 1722, 870, 2550, 1550, 2445, 2157, 1867, 755, 864, 2498, 1840, 1118, 1576, 1575, 2304, 1735, 888, 1355, 740, 2296, 2850, 1128, 1108, 2268, 2569, 1270, 2245, 957, 881, 958, 1348, 1509, 145]. model [286, 1941, 2408, 1004, 334, 2090, 1998, 1370, 2314, 2073, 1520, 1018, 2716, 2285, 2295, 982, 3068, 559, 993, 822, 856, 1320, 1553, 2530, 2038, 2153, 1031, 1380, 1397, 1479, 1671, 2255, 2213, 848, 716, 2386, 2527, 2441, 2627, 2892, 790, 1874, 1272, 977, 1415, 1630, 2804, 2534, 441, 1512, 1614, 1788, 2283, 2420, 2519, 2517, 2647, 2956, 2845, 3002, 3060, 3051, 3048, 2732, 2662, 2548, 2889, 3013, 2093, 2076, 2289, 2861, 1431, 2764, 2665, 2203, 1582, 1960, 2641, 2839, 1531, 395, 402, 1845, 1375, 2300, 2719, 751, 2813, 2672, 1640, 1466, 3026, 2393, 1186, 1094, 1391, 2886, 497, 1249, 1628, 2731, 2742]. model [2653, 917, 1596, 1841, 3029, 1701, 2086, 2190, 2475, 871, 1047, 1778, 2766, 2628, 1808, 2173, 154, 2278, 2132, 2260, 635, 1966, 2493, 1323, 2189, 1953, 1641, 2681, 2721, 2340, 2341, 2287, 1924, 855, 1932, 909, 1183, 1931, 2921, 2066, 1861, 1618, 1021, 1483, 1741, 2156, 2978, 1817, 1713, 970, 1433, 2020, 2149, 2381, 2361, 2436, 2618, 2788, 2933, 3034, 3033, 2867, 1222, 1075, 314, 1763, 901, 869, 2757, 2256, 2253, 2430, 3039, 2338, 1591, 2211, 1677, 3035, 1689, 2442, 2709, 3036, 2678, 2029, 1951, 2077, 2497, 1822, 2414, 2746, 2868, 1508, 2947, 1758, 2014, 440, 71, 2842, 3061, 3000]. Model-free [2400, 2994, 2625]. Model-Oriented [2299, 2757]. Modelibildung [532]. Modeling [1533, 1657, 2153, 1813, 2670, 2175, 1948, 2424, 2410, 3055, 1844, 2196, 2557, 2401, 2600, 2528, 2967, 1524]. Modell [246, 585, 532, 394, 499, 273, 440, 71, 246, 395, 402]. Modellannahmen [156]. Modellbauers [26]. Modellbildung [621, 206]. Modelle [344, 360, 13, 698, 209, 883, 576]. Modellen [573]. Modelling [2012, 1482, 817, 2275, 1537, 2941, 611, 1869, 2376]. Modells [571, 529, 441]. Models [620, 2351, 2764, 1456, 3021, 1340, 2603, 784, 2831, 2193, 2364, 1330, 3063, 2505, 2279, 1259, 1954, 1918, 2317, 2397, 1957, 2553, 1260, 1611, 974, 1992, 853, 1824, 448, 2273, 2012, 1971, 1412, 2391, 1084, 2969, 494, 2915, 2630, 2634, 2155, 1650, 2512, 2712, 2004, 1077, 2097, 1055, 1306, 1784, 2913, 3041, 2652, 2700, 2964, 1879, 2464, 2291, 2209, 1949, 2375, 1533, 1610, 1367, 1495, 1935, 2465, 2851, 2943, 1886, 1664, 2896, 2513, 1185, 2249, 2225, 3047, 1959, 692, 1799, 306, 2503, 2682, 538, 708, 2308, 2570, 2451, 1193, 2143, 2985, 2232, 2159, 2151, 2250, 2331, 1369]. models [1934, 1827, 1836, 1474, 2840, 2185, 1003, 2447, 1809, 1870, 934, 607, 2033, 2483, 2499, 2567, 2795, 429, 361, 1364, 379, 2588, 1551, 923, 2333, 2537, 2988, 942, 959, 2677, 2589, 882, 1207, 2248, 2380, 782, 1385, 1562, 3020, 1767, 2844, 1948, 2023, 1982, 2148, 1530, 643, 2659, 868, 1170, 2119, 748, 2120, 1714, 1652, 2726, 2002, 1324, 2170, 2348, 2538, 2617, 2895, 691, 1066, 1477, 1716, 2136, 2649, 2810, 2894, 2502, 2786, 3038, 1334, 2252, 2421, 2449, 2858, 2996, 2959, 2802, 2431, 1441, 963, 2426, 2266, 2395, 3052, 3032, 2740, 1375, 1583, 2458, 2917, 1584, 1009, 331, 3031, 1655, 2274, 2899]. models [1342, 2536, 2785, 2017, 2571, 2713, 2036, 641, 1345, 1825, 2751, 1194, 3070, 2103, 1517, 1296, 588, 1686, 2306, 2836, 2575, 1801, 2309, 2852, 1587, 2102, 1946, 1277, 1626, 2288, 741, 1196, 1262, 1251, 1500, 1535, 2359, 1775, 990, 1044, 1673, 1056, 984, 1068, 2223, 3057, 2336, 1240, 2591, 825, 908, 1313, 2063, 1387, 2767, 2583, 2848, 2188, 1346, 2774, 2963, 1037, 1172, 1209, 1074, 1443, 2990, 1045, 1285, 324, 325, 1274, 2031, 2064, 2353, 2361, 2444, 2439, 2476, 2587, 2543, 2549, 2601, 2884, 1792, 2450, 1873, 2323, 2130, 2405, 2893, 1478, 189, 2358, 2631, 2443, 1839, 1929, 2176, 2227, 2399]. models [2872, 2835, 2065, 2221, 2343, 2480, 2640, 2919, 2370, 2770, 2761, 2725, 2429, 1818, 1803, 2071, 2154, 2485, 2547, 2582, 2669, 3007, 1636, 2320, 2948, 1434, 728, 2599, 2334, 2407, 2019, 1727, 1983]. moderate [1874]. Modern [1832, 1723, 1486, 2110, 1770, 1487]. modification [863]. Modifications [1354, 1810]. Modified [1529, 872, 3048, 2697, 1776, 2993, 1836, 2415, 2627, 1672, 2825, 1617, 2953, 1036, 2929, 1442, 1037, 1210]. Modifying [965]. modulus [1815]. möglicher [156]. Möglichkeit [191]. Möglichkeiten [97, 254]. Mohan [1547]. Mohri [2367]. Moment [1716, 2180, 1335, 2652, 3045, 2390, 1711, 1904, 2076, 1560, 2265, 2444, 2543, 3043, 666]. Moments [1761, 1882, 1557, 1286, 1613, 2207, 2081, 2952, 2862, 1205, 1360, 878, 948, 2598, 2165, 1057, 1200, 2376, 1449, 2736, 2942, 2798]. monetary [2626]. monetary-unit [2626]. Mónica [2368]. Monitoring [1794, 2216, 2230, 2087, 1623, 1378, 1171, 2269, 2144]. monotone [1141]. monotonic [1952]. Monotonicity [3023]. Monte [2844, 1184, 2819, 260, 2271, 1170, 1033]. Montgomery [1834]. Montgomory [1660]. monthly [522]. Mood [3066]. Morgan [2674]. Morgenstern [1837, 1446]. Morris [1983]. Mortalität [495]. Moss [2145]. most [1830, 978, 2489, 2859]. Mostly [1708]. motion [2463, 2735]. motivated [2587]. moving [1529, 1224, 1395, 2918, 1353, 390, 415, 2562, 2263, 2178, 1775, 1698, 2029, 461]. MPS [315]. MSE [991, 1094, 2009, 917, 2713, 1513]. MSE-based [991]. much [1572]. Mulaik [1985]. Mulero [2182]. Multi [2303, 2855, 1083, 1574, 2155, 2797, 2776, 1354, 2755, 2795, 2212, 2627, 1623, 352, 1287, 2759, 2324, 2357, 2576]. multi-character [1287]. multi-component [2627]. multi-dimensional [352]. multi-environment [1574, 2155]. multi-factor [2795]. multi-factorial [2755].

Multi-feature [2855]. multi-fidelity [2797]. multi-group [2324]. multi-level [2212]. multi-objective [2759]. Multi-part [2303, 2576]. multi-sample [1083]. multi-sigmoidal [2776]. multi-way [1354]. multicollinear [3053]. Multicollinearity [438, 189, 359, 282, 2827]. multicomponent [2962, 2255, 2213, 2386, 2527]. multidimensional [827, 2106]. multidimensionale [453, 471]. multifactor [2399]. Multikollinearität [217]. multilevel [2314, 1727, 2802]. multilevel/hierarchical [1727]. Multinomial [2214, 2799, 1077, 1095, 777, 729, 1811, 691, 1441, 866, 605, 1673, 1434, 2344]. multiparametric [1318]. Multiple [2406, 2633, 1810, 2460, 2961, 2935, 1980, 821, 1806, 1245, 1674, 1167, 1221, 2465, 2986, 2864, 1429, 251, 2061, 856, 1553, 2415, 2534, 2477, 1308, 1584, 2496, 1052, 2836, 879, 2129, 2462, 2067, 2539, 2531, 2565, 199, 1771, 1568, 780]. multiple-group [2836]. multiple-sensitive [1806]. multiplen [563]. multipler [96, 104]. multiples [479]. Multiplicative [2485, 2505, 2913, 2517, 2889, 3032, 2024, 2669]. multiplier [1019]. multiply [1467, 2600, 2533]. multiregional [37]. Multiregionale [13]. multiresponse [2449]. multiscale [2264]. multistage [2658, 2428, 2584]. Multitude [1599]. Multivariate [2505, 2272, 2578, 1073, 1762, 2796, 742, 1723, 2957, 2575, 2938, 2166, 1831, 2226, 1317, 1742, 2237, 1653, 2008, 1889, 1402, 2952, 422, 587, 637, 2733, 701, 2378, 1575, 2944, 1911, 2206, 980, 1760, 607, 1179, 2455, 2748, 777, 2417, 982, 1743, 856, 1553, 1530, 1031, 1343, 1580, 2120, 1630, 1252, 2534, 1512, 2786, 801, 1510, 2466, 2426, 2825, 3056, 1094, 1391, 1371, 2050, 2731, 2742, 2954, 2922, 2310, 2319, 2607, 2084, 2882, 605, 301, 2566, 423, 1855, 2855, 813, 1444, 797, 2509, 2528, 3033, 698, 2949, 3036, 1803, 2948, 2991]. multivariate-GARCH-BEKK [2786]. multiwavelets [2169]. Murtagh [2110]. Music [3074, 1890, 2217]. Mustererkennung [484]. Mutual [1669].

N [2111, 2554]. N. [1684]. nach

[146, 514, 521, 297, 532, 336, 515, 84, 258, 289]. Nachfrage [146]. Nachfrageanalyse [153]. naive [2881]. Nassim [1687]. National [468, 130]. nationale [53]. Nations [265, 3]. Natur [2109]. Natur- [2109]. natural [3070, 523, 1483, 685]. NBAFR [710]. NBUC [1331, 1430]. NBUFR [938]. NBUL [1728, 1593]. ND [2020]. Near [927, 2058]. Near-integration [927]. near-order [2058]. nearest [2220]. nearly [2164, 2784, 2758]. nearness [2850, 1391, 1516, 966, 1119]. Necessary [765]. negative [838, 2282, 1002, 2666, 2993, 1295, 2350, 2506, 2254, 1697, 2800]. negatively [2652, 2498, 3067, 1883, 1542, 2180, 1757, 2444, 2602, 2221]. neglected [1413, 1206]. neighbors [2220]. Nelder [1664]. Nested [2936, 2979, 2595, 1935, 2423, 1579, 1581, 2348, 2617, 735]. nested-error [2348]. Nettoproduktion [184, 54]. Network [413, 2199, 2674]. Networks [1682, 2749, 3040, 2651, 2765]. Neudecker [2873]. neue [192, 188]. Neuer [170]. neuere [536, 467]. Newton [2671]. Neyman [2903]. NGG [1957]. Ngurah [1706]. niche [2646]. Nicholas [1687]. nicht [504]. nichtlineare [270]. nichtregulärer [181]. Nielsen [1682]. Nigel [2147]. Niveau [277]. Niveaus [72]. NMAR [2702]. NN [2410]. no [326, 2067]. Noel [2556]. noise [2950, 3062, 1192, 2534, 2672, 2189, 1075, 2866, 2154]. noises [1309]. noisy [2673]. Nomenklaturen [80]. Non [3062, 1972, 1935, 1676, 2672, 837, 1535, 2778, 2279, 1952, 752, 2707, 1002, 1104, 2508, 2816, 1538, 1639, 2027, 1211, 2032, 1154, 334, 2001, 2999, 2271, 2286, 1320, 1875, 1479, 543, 742, 919, 1915, 2481, 1477, 2456, 2479, 2309, 1914, 2938, 393, 475, 1632, 1659, 855, 2942, 2761, 2605, 2565]. non-additivity [742]. Non-asymptotic [2778]. Non-central [1676]. non-classical [2942]. non-crossing [2279]. non-directional [2565]. non-exchangeability [1639]. non-exchangeable [1538]. Non-existence [837]. non-Gaussian [2816, 2271]. non-identical [1104, 2938]. non-identically [752]. non-ignorable [2032, 2605]. non-linear [1154, 2309, 475, 2761]. non-linearity [334]. non-monotonic [1952]. non-negative [1002]. non-normal [1477]. non-normality [855]. non-normalized [2481]. non-null [2479]. Non-parametric [1972, 2672, 1535, 2027, 2286, 393]. non-propagating [1914]. non-random [2707]. non-randomized [2508]. non-randomly [2999]. non-regular [1211]. non-response [1632, 1659]. non-sample [1320, 1875]. non-stationary [2001, 1479, 1915, 2456]. non-stochastic [543]. noncentered [2496]. noncentral [1361, 681]. nonexchangeability [1436, 1406]. nonhomogeneous [1298, 704, 834, 990]. nonidentically [1286, 1613]. nonignorable [2896, 2494, 2649]. Noninformative [1357, 1539]. noninvertible [2349]. Nonlinear [1900, 2397, 2273, 1631, 1650, 1695, 1610, 2880, 2451, 923, 2248, 2380, 2002, 2889, 2310, 1843, 2163, 2872, 1643]. nonlocal [2841]. nonmonotone [1012]. nonmonotonic [2636]. Nonnegative [2538, 2837, 1444, 1970, 2909, 920]. nonnegative-definite [920]. nonnormal [1210]. nonnormality [801]. Nonparametric [1233, 2184, 1092, 2277, 2955, 3040, 988, 2491, 1766, 3068, 2158, 2094, 2418, 2839, 2789, 3026, 834, 2329, 1171, 324, 325, 326, 1508, 1259, 895, 2638, 900, 1083, 2083, 1418, 2464, 3045, 2550, 2498, 1643, 3008, 2804, 1912, 2519, 3051, 2458, 1816, 3056, 1596, 2934, 2319, 2607, 2347, 3058, 3059, 2020, 2549, 2430, 2631, 2459, 2251, 2669, 1984]. nonregular [1407]. nonresponse [1921, 2896, 1821, 2494, 2649, 780]. nonsationary [925]. Nonsense [1206, 2644]. nonstationarity [1485]. nonstationary [905, 2349, 829, 2164, 1916, 1528, 3070]. norm [2403, 3056, 2992, 1795]. Normal [1533, 1457, 2787, 1702, 3017, 1395, 2655, 2847, 2008, 1012, 354, 1631, 1462, 2195, 2863, 1854, 2699, 916, 1762, 1624, 2660, 1264, 2143, 1819, 1882, 1862, 1938, 860, 1768, 2754, 1699, 2200, 1484, 3020, 1767, 1129, 1357, 1539, 2415, 1790, 2290, 1994, 1477, 3055, 1712, 2426, 806, 2825, 2957, 540, 2085, 859, 1011, 1820, 1969, 824, 1000, 2524, 301, 423, 1499, 1855, 1228, 2264, 2939, 1021, 2216, 1432, 797, 2509, 1608, 3010, 1822, 2334, 2991]. normal/independent [1882]. normality [2035, 1974, 2520, 2747, 2790, 1179, 2390, 2090, 2598, 2823, 2057, 1845, 2792, 2953, 855, 1839]. normalization [1600, 2609, 2200, 2101]. normalized [2541, 2481]. normalizing [3042]. Normalverteilung [103, 356, 317, 296]. Normalverteilungen [460]. Norman [2431]. norms [2984, 562]. Note [587, 958, 2500, 1133, 1068, 390, 415, 862, 895, 623, 496, 1331, 530, 1002, 851, 1299, 1927, 2365, 1720, 2573, 1214, 486, 625, 282, 1003, 396, 2974, 981, 379, 1520, 1492, 848, 554, 775, 1324, 1057, 1805, 454, 1616, 1248, 2247, 1311, 1997, 661, 1623, 1420, 300, 1485, 2181, 301, 600, 2098, 1174, 1478, 1131, 2253, 1591]. notes [2131]. notion [1498]. notions [1952]. novel [2913, 2037, 2632, 2843, 2720]. nuclear [3056]. nuisance [1575, 976, 2717, 1582, 2128]. null [2747, 1630, 2479, 2961, 1939, 2026, 2080]. null-hypothesis [2026]. number [1572, 1093, 989, 2823, 3030, 730, 2647, 2040, 2205, 1386, 560, 1433, 2139, 1131, 2443, 2925]. numbers [2058, 629, 644, 2210, 2908, 1587, 685]. Numerical [327, 359, 598, 1086, 2670, 279, 1440]. numerisches [658, 246]. Núñez [1870]. NUOPT [1486]. nutritional [688]. Nutzenaxiomatik [31]. NWAFR [710].

ø-divergence [1223]. O. [1703]. object [2735]. Objective [2290, 597, 1168, 2759, 2334]. objects [1572]. observation [2760, 2888, 1745]. Observations [308, 644, 1971, 1495, 2851, 2284, 1383, 704, 2664, 2133, 2461, 2833, 2455, 1283, 788, 1785, 1536, 1767, 2704, 1380, 1187, 1215, 1865, 2789, 2648, 1791, 2388, 2098, 1037, 1209, 1745, 2183, 2835, 2967]. Observed [656, 1898]. obtained [2783]. obtaining [2407]. occurring [2749]. odds [1675, 2958, 3024, 1779, 2971, 1500, 2868]. Oehler [293]. off [1735, 1680]. Offene [621]. offenen [394]. öffentlichen [50]. Office [1566, 468]. Ogihara [1890]. OK [2713]. Ökonometrie [360, 344, 568, 40, 621, 912]. Ökonometrische [698, 70, 78]. ökonometrischen [585, 614, 26, 273]. ökonometrischer [171, 344, 516, 206]. ökonometrisches [499]. ökonomischer [613]. Olaf [1503]. old [1588, 1589]. Olkin [1615, 1856, 1475, 1651, 1733, 2278]. OLS [1108, 992, 910, 1387]. OLSE [1941, 1520, 1227, 2766]. omitted [2890, 1391, 917, 1516, 1744]. Omnibus [3032]. One [1853, 821, 2155, 1927, 1407, 998, 860, 1155, 2175, 245, 2548, 2419, 2760, 2762, 1946, 855, 237, 324, 1222, 1297, 1670]. One- [1853, 821]. one-inflated [2175]. one-parameter [1297]. one-sided [2419, 1670]. one-stage [2155]. one-way [1927, 1155, 855, 324, 1222]. Online [2838, 2708]. Operating [1645]. operation [897]. Operational [109]. operationelle [49]. Operations [19]. operative [736]. operator [2993, 2246, 2350, 2506]. operators [785]. Opinion [3073, 2219]. opposite [30, 2043]. Optimal [2117, 1629, 2118, 2741, 1575, 3022, 2308, 3009, 2258, 2898, 1553, 2305, 2606, 2466, 2641, 2597, 2363, 1296, 2306, 2760, 2762, 2128, 37, 2681, 2551, 1625, 2793, 2313, 2701, 2780, 1042, 2580, 3062, 1076, 851, 2763, 2513, 2427, 2633, 2880, 2135, 406, 1577, 2268, 2569, 2521, 2483, 2567, 2795, 1454, 1581, 672, 2215, 1899, 1324, 1390, 2894, 2040, 1583, 2765, 2613, 2311, 2309, 2758, 2852, 2860, 2312, 2346, 505, 517, 2493, 2098, 2129, 2612, 2358, 2176, 2518, 2761, 2725]. optimale [544, 113, 250, 614, 219]. optimalen [585, 204, 368]. optimaler [29, 498]. Optimality [2304, 2704, 2592, 2100, 1561, 2589, 2984, 1157, 1277, 449, 2759, 3039]. Optimierung [121, 196, 234]. Optimierungsaufgabe [69]. Optimierungsproblemen [250]. optimization [1486, 2590, 1056]. Optimized [737]. Optimum [1001, 242, 964, 1319, 1991, 665, 695, 837, 1277, 608, 2757]. options [1453]. oracle [2071]. Order [2463, 1104, 1961, 1364, 1286, 1613, 1812, 2482, 2058, 2722, 2807, 2081, 1103, 752, 2228, 2051, 2862, 1784, 1205, 1360, 2749, 2911, 2513, 1043, 2304, 987, 2308, 2135, 2850, 1761, 3040, 316, 981, 1522, 1307, 2177, 2165, 1719, 1498, 2647, 2845, 3060, 1057, 1541, 2374, 1322, 2809, 1945, 1116, 1916, 2597, 1312, 2327, 2102, 2288, 1601, 2938, 1196, 2346, 1445, 1262, 1775, 1739, 1855, 1225, 3042, 1313, 2470, 1859, 1449, 1037, 157, 1619, 3034, 1863, 2736, 1463, 2640, 2518]. ordered [1605, 2707, 1245, 2741, 2191, 1881, 1819, 2805, 1947, 2792, 2524, 785, 2329]. ordering [731, 1288, 721, 756, 629, 1356, 1711, 2148, 2745]. orderings [898, 2822, 2949]. orders [1401, 2051, 2738, 1295, 1399, 2182]. ordinal [2500, 2941, 1170, 2900, 1966]. ordinalist [114]. Ordinary [908, 643, 1480, 1931, 825]. Organisationen [4, 139, 5]. organisatorische [464]. Oriented [2299, 1573, 2757]. origin [852]. Ornstein [2950, 3062, 2001, 2388, 2544]. orthant [2444, 2468, 2901]. orthogonal [2053, 2592, 2307, 2756, 2782, 2624, 2657, 2909, 3050, 1946, 2205, 2720, 2979, 2586]. orthogonalen [319]. orthogonality [809]. Osband [2846]. Other [2555, 171]. out-of-control [1510]. Out-of-sample [1127]. outcome [2032]. outcomes [1073]. Outer [1812]. outlier [2578, 2733, 1720, 2869, 2417, 245, 2966]. outliers [3054, 1824, 2697, 2375, 2763, 1911, 835, 989, 2379, 1914, 1222, 1131]. outlying [1263]. outlyingness [2778]. Output [336, 124, 13, 290, 163, 28, 1537, 33, 421, 2084, 128, 37, 154, 417, 271, 451, 41, 148, 123, 370, 417, 452, 36, 451]. Output-Analyse [28, 41, 36]. Output-Forschung [124]. Output-Modelle [13]. Output-Rechnungen [148]. Output-Tabellen [336, 290, 123, 370, 417, 452, 451]. outstanding [730]. outward [2727]. over- [2132]. over-parameterized [2840]. overall [2972]. overdispersed [2915, 2634, 2173]. overdispersion [1482, 2542, 2323]. overdose [2236]. overlap [2646]. overlapping [2909, 1322, 2270, 2801]. overstate [3064]. overview [3063, 887, 1282, 779]. Oxford [2112, 2199, 2161]. ozone [2096].

P [1770, 1487, 799, 853, 2712, 2430]. P-C [2430]. P-spline [2712]. p=1 [237]. p=1/2 [237]. Pace [1704]. packages [954]. pädagogisch [2055]. pädagogisch-psychologischer [2055]. Pagination [2369]. pair [762, 2384]. paired [2133, 2455, 2854, 2597]. pairwise [1883, 2900]. Palella [1525]. Palma [2146]. Panel [904, 2335, 1918, 1374, 1901, 900, 941, 2171, 3065, 1073, 1900, 2864, 2112, 2499, 2988, 1484, 2956, 963, 1939, 2812, 2888, 909, 1022, 1183, 2963, 2978, 2788, 2884, 2399, 3007]. panels [2456]. Panik [1813]. Paola [2106, 2675]. Paolo [1726]. paper [525, 625, 1472, 1544, 1523, 2161, 440]. Paperback [1566]. Papers [1728, 1607, 633, 640, 646, 663, 671, 683, 714, 726, 747, 769, 792, 800, 812, 819, 831, 841, 874, 884, 893, 902, 911, 921, 930, 939, 951, 961, 972, 985, 995, 1016, 1028, 1040, 1049, 1060, 1070, 1080, 1090, 1100, 1111, 1125, 1137, 1151, 1163, 1177, 1190, 1203, 1218, 1231, 1243, 1255, 1268, 1280, 1290, 1303, 1328, 1426, 1906]. parabolic [2273]. paradigm [1097]. paradox [928]. parallel [2699, 2972]. Parameter [260, 899, 2001, 1156, 2871, 2529, 944, 498, 191, 1873, 2799, 648, 2355, 821, 639, 2803, 1837, 242, 1797, 1046, 1114, 1407, 976, 699, 2987, 1998, 809, 852, 2295, 1536, 999, 3020, 1320, 1620, 1878, 1064, 2659, 489, 1790, 513, 675, 2018, 2150, 2504, 2297, 1481, 2486, 1979, 2436, 1297, 869, 1388, 2544, 153]. Parameter-Restriktionen [153]. parameterisation [810]. parameterization [2942]. parameterizations [1844]. parameterized [2840]. Parameters [2833, 2787, 1846, 1330, 415, 687, 2092, 1860, 974, 359, 1462, 162, 975, 2562, 946, 2354, 1459, 890, 2233, 1213, 1967, 1575, 2850, 2032, 227, 343, 316, 2478, 334, 882, 2416, 2806, 847, 2069, 1982, 993, 822, 1647, 1633, 1025, 2409, 1654, 1065, 2647, 144, 2895, 1947, 808, 1582, 751, 850, 2713, 956, 2086, 2099, 1838, 1876, 1195, 1250, 1036, 2201, 1618, 1823, 1160, 189, 229, 2075, 2443, 2344]. Parameterschätzungen [547]. Parametric [1317, 2700, 863, 2412, 1012, 915, 1083, 2356, 1972, 2027, 2879, 2883, 2333, 1520, 2286, 2581, 2122, 2672, 1596, 1047, 1778, 2843, 393, 1535, 2583, 2154, 2582, 2320, 728]. parametrischen [60]. parametrized [2447, 2470]. Paretian [1533, 1074]. Pareto [1846, 2207, 933, 2707, 1030, 439, 2562, 1603, 1467, 2007, 958, 1236, 899, 989, 2484, 1647, 144, 263, 808, 1556, 1455, 2150, 2529, 1468, 1859, 979, 1106]. Parmigiani [1749]. part [1259, 2303, 2009, 2576]. Partial [1469, 2414, 1978, 1259, 1731, 1999, 462, 898, 602, 2712, 2964, 1638, 2896, 543, 2165, 2410, 2984, 1277, 2376, 1618, 1741, 2361, 2436, 2601, 2468, 2154, 2547]. Partially [2102, 2364, 2620, 2512, 762, 1306, 2532, 1840, 2960, 2191, 2133, 2499, 2090, 2073, 2380, 2647, 3060, 2136, 2889, 2810, 3013, 2861, 2103, 2049, 2714, 3057, 2064, 2618, 2933, 1792, 2450, 1310, 2338, 2872, 2480, 2442, 2429, 3007, 1636, 2320]. partially-accelerated [2073]. particles [2117]. particular [439, 870, 440]. Partielle [531]. Partiellen [596]. partieller [594, 545, 651]. partitioned [2295, 1562, 2077]. Partitioning [2220, 2467]. parts [2267, 2736]. passage [2776]. past [2387, 1997, 1623, 2723, 25]. pathway [1691]. pattern [2775, 1966, 2585, 1996]. Patterned [2259]. patterns [2463, 1781, 1937]. Paul [2144]. Pawitan [1664]. pay [1680]. pays [586, 12]. PCA [2688]. PDFs [1559]. Pearson [2970, 997, 2392, 1334, 2735, 1058, 1199]. Peck [1660]. penalization [3044, 3056]. Penalized [1954, 1695, 2499, 2878, 2639, 2062, 2186, 3028, 2229, 1368, 1511, 1964, 1919, 2401, 2163]. penalizing [3058]. penalty [2538, 2900, 2139]. perceivers [1308]. percentage [998, 1025, 681]. percentile [2453]. Perfect [2178, 1400, 1879, 2251]. Performance [2437, 1999, 2283, 2940, 2902, 1234, 1355, 1087, 2144, 1094, 1011, 2017, 774, 1105]. Periode [362]. Periodic [888, 1937, 2566, 2544]. Periodically [2196, 2635, 2271]. periodicities [279]. periodogram [828]. periodograms [2196]. permitting [1453]. permittivity [2264]. Permutation [967, 968, 2819, 2490, 2137, 1385, 2038, 2192, 1823, 1895]. Perron [1720].

Perron-type [1720]. persistence [2335, 1783]. persistent [905, 2728]. Personelle [108]. personeller [348]. persons [1674]. perspective [1246, 2842, 2451, 1747, 1943, 1684]. Perspectives [1570]. Perspektiven [112]. perturbed [1127]. Pesaran [2112]. Peter [1724, 2433, 1891, 1724, 265]. petit [586]. Pfaff [1705]. PH [2813]. pharmaceutical [1771]. phase [2525, 1821, 2087, 1632, 2236]. phi [2257, 1441, 1531]. phi-divergence [1441, 1531]. Philip [2199]. Philippe [1725]. Phillip [1566, 1566]. Photographie [631]. Physica [349]. piece [2879, 1699, 3048, 1933]. Piegorsch [1665]. Piesch [381]. Pilz [2432]. Pischke [1708, 2104]. pitfalls [386]. Pitman [2850, 1947, 1391, 1516, 1480, 966, 1119]. pivotal [620]. plan [2604, 786, 232, 71]. plane [2453]. planning [154]. plans [1169, 1692, 929, 1440, 1612, 2624, 2363, 2613, 1517, 2346, 2030, 1609, 1769, 1625]. Planung [52]. playing [1588]. plot [1571, 2492, 2326, 2925]. plots [2730]. PLS [576, 1572, 2141]. Plug [1259]. Plug-in [1259]. PLUS [1566]. plutonium [817]. PMP [259, 259]. Point [2411, 1159, 2787, 1830, 722, 1034, 2171, 1930, 2689, 2749, 2911, 1722, 1107, 2664, 2883, 3040, 2455, 2748, 1440, 2976, 2394, 2998, 2460, 2957, 1655, 1342, 824, 2454, 2951, 2866, 2459, 2919, 1622, 1508]. points [1980, 2463, 1093, 1695, 2864, 998, 957, 2796, 1025, 1960, 2466, 2458, 2648, 681, 505, 1861, 2705, 2645, 2935, 2910]. pointwise [1454]. Poisson [2383, 2282, 2667, 1180, 1764, 1159, 1435, 1802, 2028, 948, 1272, 2366, 2548, 2878, 2827, 1560, 943, 2332, 2132, 834, 990, 2973, 2066, 2401, 1174, 2715]. polar [2660]. polar-generalized [2660]. polarimetric [2777]. polarization [2276]. Polens [62]. Policy [647, 2903, 1331, 816, 171]. Politik [533]. Politis [2554]. Polling [3073]. Polls [2219]. Polya [782, 2324]. polygon [3067].

Polynomial [500, 2788, 2237, 762, 2768, 2762, 3058]. polynomials [2053]. pooled [937]. Pooling [1077, 1095, 1343]. Population [2615, 2207, 1001, 2208, 1132, 1265, 1974, 991, 1213, 2133, 1821, 1349, 1807, 1532, 245, 1654, 1942, 892, 2150, 2504, 2294, 1261, 301, 1442, 1659, 1038, 2216]. Populations [2787, 2092, 1777, 1461, 916, 987, 2823, 3030, 1994, 2079, 1032, 866, 1554, 2524, 1157, 1601, 1038, 763, 3016]. Portefeuille [362]. Portfolio [2507, 1486, 2230, 1332]. portfolios [1649, 1414, 1896]. Portfoliotheorie [659]. Portmanteau [2682, 2665, 2505]. positive [2175, 2889, 2009, 416, 1432]. positive-part [2009]. positively [2534]. possible [3068, 157]. possibly [2620, 2570, 1270, 245, 237]. Post [2614, 2425, 991, 163, 2251]. post-stratification [991]. postage [1588]. Posterior [2772, 1443, 1467, 3015, 1616, 1074]. pot [2326]. pot-pot [2326]. potency [1578]. poverty [2106, 842]. Power [2533, 687, 1600, 3004, 1076, 872, 1083, 2552, 1046, 2629, 2849, 998, 129, 598, 279, 1192, 1847, 2200, 1597, 1511, 1065, 144, 2665, 2981, 416, 1056, 1239, 1442, 855, 3042, 1106, 2101, 2877, 2565]. power-function [144]. power-related [2981]. powerful [1830, 1810, 2752, 978, 2489, 2859]. Powers [2787]. pp [2675, 2198, 2147, 2219, 2112, 2218, 2259, 2182, 2197, 2217, 2199, 2556, 2160, 2161]. Practical [949, 2920, 2930, 3000, 1433]. practice [1834]. practitioners [1796]. prädeterminiertem [642]. Prado [1869]. Praxis [280]. pre [1087, 1129, 977, 859, 917, 1011, 1047, 1680, 945]. pre-test [1087, 1129, 977, 859, 917, 1011, 1047, 945]. pre-testing [1680]. Precedence [1115]. precise [1105]. Precision [668, 824, 2524, 608, 2991]. predator [1554]. predecessors [3]. predict [1841]. Prediction [2228, 2851, 1021, 1172, 1340, 1365, 1812, 1957, 1197, 2707, 1153, 1972, 1078, 1167, 2666, 1853, 3022, 145, 607, 3008, 2120, 2768, 1186, 1212, 2243, 2760, 1543, 438, 984, 2781, 2063, 2623, 1515, 2612, 1763, 2631, 3039, 2029, 2605]. predictions [2829, 825, 908]. Predictive [2722, 2017, 835, 2088, 477]. predictor [851, 2453]. predictors [1068]. Preface [1908, 1570]. preference [629]. preferences [624]. Preis [443, 499]. Preis- [443]. Preis-Dynamik [499]. Preisindizes [612]. Preisniveauindizes [673]. Preisrechnung [52]. Preisstrukturen [174]. Preisumkehrtests [555]. Preisvergleich [467]. Preliminary [1647, 1531, 2372, 2172, 1796]. premiums [1550]. preparations [1578]. prequential [1735]. prescribed [936]. presence [3062, 570, 2697, 2923, 339, 2717, 1109, 1309, 475, 1632, 1659, 2074, 1625, 2545, 2948]. present [2134, 25]. Preservation [731, 1288, 1711, 1401, 1295, 1430]. preserving [920, 1171, 2857]. Press [2675, 2112, 2182, 2197, 2199, 2161]. Pretest [2799, 3047, 2429]. prey [1554]. price [522, 556, 623, 660, 717, 727, 686, 525, 810, 612, 661, 524, 271, 861]. price-induced [727]. pricing [2296]. prime [1534, 1946]. Primer [2198]. Principal [1960, 2138, 1797, 2100, 1998, 1993, 2726, 2283, 2577, 2466, 2496, 1586, 2781, 2855]. principle [2975, 490, 2846, 330, 918]. principles [1749]. print [2182, 2197]. prior [2940, 953, 1299, 2849, 2277, 1514, 1995, 3012, 1320, 1875, 2415, 776, 774, 1641, 2470, 1021, 1086]. Priorités [586]. Priorities [586]. priority [1337]. priors [1635, 722, 849, 2815, 1357, 1539, 2841, 2324, 2334]. Privacy [2857, 2045, 1827, 1995]. Privacy-preserving [2857]. privacy-protecting [1995]. private [2651]. prize [439]. prize-competition-distribution [439]. probabilistic [2046, 404]. probabilities [2775, 891, 2699, 3022, 2710, 601, 2718, 2522, 263, 2263, 1939, 2278]. Probability [1417, 512, 1140, 1629, 2783, 1662, 494, 2574, 2091, 2277, 1293, 1525, 625, 1181, 404, 79, 559, 554, 2122, 42, 2087, 2559, 1504, 600, 2470, 783, 237, 404]. PROBASTAT [2116]. probit [1478, 1434]. Probitregression [787]. Problem [1123, 1135, 1161, 1175, 1188, 1201, 1216, 1241, 1350, 1393, 1424, 1438, 1451, 1488, 1505, 2114, 2115, 631, 235, 140, 1301, 1326, 1371, 528, 1229, 225, 1158, 551, 711, 1083, 251, 2521, 1760, 172, 1332, 1008, 2432, 505, 2264, 2129, 1359, 2067]. Problematik [220, 156, 183]. Probleme [231, 595, 10, 603, 11, 89, 80, 606, 288, 376, 151]. problems [1299, 2818, 2490, 2461, 1771, 205, 2929]. Problemsection [1266, 1278, 1314]. procedure [1731, 1806, 1800, 1354, 2819, 964, 981, 1561, 804, 1754, 1865, 871, 879, 1866, 1313, 2848, 2401, 1131]. Procedures [1568, 1629, 1245, 1055, 2986, 1565, 934, 2079, 3032, 879, 1037, 2302, 1072]. Proceedings [1570]. Process [478, 390, 2950, 551, 2525, 2776, 2028, 2550, 2993, 2292, 1336, 2001, 955, 2542, 1875, 1271, 1868, 1877, 2164, 3026, 300, 352, 1468, 2888, 715, 2973, 3046, 2350, 2506, 1173, 2401, 2254, 2459, 1677, 2770, 1622, 461]. Processes [1724, 1978, 415, 1127, 933, 2383, 2275, 2870, 1153, 1180, 2525, 1603, 2885, 2749, 2911, 306, 2378, 976, 2461, 414, 2914, 3040, 2472, 845, 2891, 981, 1794, 2703, 1651, 1733, 847, 2619, 2108, 1898, 1915, 1465, 1899, 2534, 2728, 1133, 2087, 2683, 1528, 943, 827, 2294, 2457, 1485, 1378, 2566, 834, 877, 990, 935, 2388, 1238, 1359, 2544, 1996, 2238]. processing [199]. product [1956, 1557, 1580, 1376, 1559, 1134, 1182, 1497, 1659, 818]. production [2317, 12, 421]. products [1344, 1130, 1371, 1698]. Produkten [146]. Produktion [72, 107, 258]. Produktions [289]. Produktions- [289]. Produktionsfunktionen [156]. Produktionsindizes [65]. Produktivitätsmessung [11]. produzierende [120]. Professor [2873]. Professoren [338]. Profile [2813, 2963, 2415, 2827, 1521]. profiling [2895]. Prognose [146, 103]. Prognoseeignung [137]. Prognosegüte [344]. Prognosemethoden [243]. Prognosen [70, 78, 250]. Prognoserechnung [97]. Prognoseuntersuchung [509]. Prognoseverfahrens [614]. prognosis [137]. program [586, 413]. Programme [36, 586]. Programmierens [22]. programming [2118, 2513]. programs [2113]. progressive [2081, 2411, 2960, 1671, 2527, 2892, 2030, 2201, 1609, 1625]. progressively [2321, 1853, 2233, 2073, 1620, 2386, 1186, 1517, 2346, 477, 1859]. project [71]. Projection [2784, 2595, 2561]. projections [2931]. Projective [2125]. projector [1573]. Projektes [71]. proof [791]. propagating [2975, 1914]. propensity [2649, 2708]. proper [1635]. Properties [2684, 1154, 1455, 825, 2645, 2518, 784, 1736, 2482, 731, 1288, 615, 2046, 2655, 2847, 1653, 1396, 1738, 1379, 1787, 215, 749, 755, 2815, 2660, 2923, 881, 1155, 2408, 2285, 1711, 2322, 1295, 2617, 2937, 2235, 1134, 1130, 2377, 2536, 943, 2327, 2340, 2510, 2376, 2301, 1744, 2381, 2602, 2468, 2497, 2471]. property [2258, 1430, 2071]. prophet [791]. prophet-proof [791]. proportion [1693, 1565, 806, 1308, 892]. proportional [2222, 625, 1181, 1551, 2813, 2041, 2751, 1500, 1953, 600, 1463, 2868]. proportionality [1418]. proportions [1674]. propos [1]. proposals [3063]. prostate [2553, 1841]. protecting [1995]. protection [1827]. providing [553]. proxy [2659, 977, 825]. Pseudo [2537, 942, 959, 3020, 2341, 2515, 3028, 712]. Pseudo-likelihood [3020, 3028]. Pseudo-maximum [2537]. pseudo-random [712]. Pseudolikelihood [1785]. psychologischer [2055]. Public [3073, 2219]. publication [647]. publications [1750]. Pulley [2747]. pure [1336, 877]. pursuit [2622, 2857]. Pyramidal [907].

QMLE [2497]. quadrant [1883]. Quadrate [217]. Quadratic [1474, 1564, 1611, 1007, 540, 1249, 1309, 1047]. quadratischen [250]. Qualität [580]. Qualitative [1917, 2807, 2526]. qualitativen [833]. Qualitätsregelkarten [873]. Quality [2034, 1794, 1796, 774, 679]. quality-control [679]. quantal [470]. Quantenmechanik [161]. quantification [2481]. Quantifizierung [207]. Quantifying [2642]. Quantile [932, 2279, 2679, 2451, 2231, 2338, 2872, 2186, 639, 162, 2365, 2291, 2209, 2920, 2879, 2185, 2068, 2593, 2999, 1555, 2501, 2804, 3013, 2339, 2377, 2564, 1261, 2070, 2681, 2623, 1595, 2436, 2396, 2793, 2062, 2701]. quantiles [2402, 3022, 2500, 2806, 2854, 1416, 2329, 1926]. quantitative [2940, 2807, 1717, 2526, 2624, 2366, 1122, 71]. quantitativen [833]. quantitatives [71]. quasi [835, 1284, 2265, 2963]. quasi-Bayesian [835]. quasi-likelihood [1284]. quasi-maximum [2963]. quasi-stationary [2265]. quaternions [2970]. question [1827, 2212, 1397, 1617, 195]. questioning [2779]. questions [1806]. queue [21]. queueing [798, 2260]. queues [816]. quincunx [1102]. quotient [1956, 2005, 1182]. Quotienten [460]. Quotienten-Test [460].

R [1486, 2198, 1545, 1962, 1566, 2110, 912, 799, 789, 2105, 2161, 440, 2675, 2218, 2197, 2556, 1750, 2160, 3075, 2239, 2432, 2111, 1705, 1833, 2433, 1683, 2368]. R/S [1566]. R/S-PLUS [1566]. radar [2777]. radial [1852, 1977]. radiation [2276]. radii [920]. Rahlf [2239]. Rahmen [171, 315]. Rande [192]. Random [197, 2472, 1877, 1613, 2603, 1401, 2482, 561, 2402, 1957, 2184, 2355, 752, 2228, 2707, 2931, 2902, 2035, 2411, 1265, 1787, 2876, 3065, 2652, 1104, 1205, 1360, 1603, 430, 2885, 1852, 2445, 1854, 1664, 1639, 3037, 1383, 2560, 712, 2005, 1457, 1809, 1154, 934, 2644, 928, 1989, 1283, 1551, 2259, 2754, 1883, 2716, 1385, 730, 1187, 1215, 2022, 1973, 3006, 2617, 2662, 2252, 1032, 1805, 2981, 2393, 1134, 1130, 880, 1542, 2908, 2475, 1519, 2306, 2760, 2575, 2413, 772, 1518, 1251, 416, 1497, 2180, 2231, 1668, 1866, 855, 1990, 1386, 1700, 1757, 1884, 2052, 1745, 1979, 2204, 2444, 3043]. random [1222, 2602, 2736, 2468, 2935, 2080, 1677, 2949, 2511]. random-effects [2252]. randomisation [1168]. Randomization [830, 2302, 2658]. Randomization-based [2302]. Randomized [1001, 1389, 1409, 2508, 2212, 1806, 2045, 1002, 1674, 1693, 1565, 1397, 1248, 2961, 1453, 1617, 1122]. randomly [2652, 1128, 2891, 2037, 2999, 730, 3043, 2338, 2872]. randomness [955]. range [1276, 2012, 1756, 1783, 1979]. Rangsummentest [292, 309]. Rangtests [649]. Rank [1008, 1585, 2343, 1402, 1851, 3018, 2191, 2688, 2422, 978, 680, 1864, 1310]. Rank-based [2343, 3018]. rank-ordered [2191]. rank-reduced [1310]. ranked [1529, 1846, 2831, 1224, 1395, 2707, 2562, 2608, 1872, 1807, 1532, 1789, 2241, 2428, 2584, 2706, 2529, 2178, 2384, 2723]. ranked-based [2723]. ranking [2727, 2178, 2251]. rankness [2945]. Rao [799, 2552, 701, 1334]. Rao-type [2552]. Raquel [1869]. rare [2773]. Rasch [2109, 2432, 2298]. Rate [1899, 784, 1165, 3062, 2794, 1764, 820, 1207, 1498, 2477, 2041, 2945, 1251, 548, 2260, 2972, 2510, 2591, 2066, 1925, 2330, 3036]. Rate-optimal [1899]. Rates [54, 1600, 1084, 2863, 2023, 1368, 1109, 1616, 2899, 3049, 2867, 885, 2414]. rather [2856]. Ratio [308, 1532, 1622, 2599, 2397, 1534, 2293, 3064, 711, 870, 1854, 1867, 1734, 1355, 736, 1675, 227, 1014, 1785, 2754, 1557, 1807, 1539, 2745, 1362, 2434, 253, 318, 2809, 1583, 2711, 837, 1779, 1261, 2971, 1056, 1497, 1786, 1384, 3014, 1422, 1460, 2216, 1609, 3023, 2095]. ratio-type [1261]. rational [2569]. Rationalization [2462]. ratios [1994, 745, 1344]. Rauchen [495]. Raum [48, 112, 120]. räumlichen [368]. raw [405]. ray [2668]. Rayleigh [631, 648, 1842, 1128, 2892, 1874, 513, 1719, 1497]. Raymer [1686]. re [1844]. re-parameterizations [1844]. Real [2689, 2924, 1429, 584, 861]. Real-time [2689, 1429]. RealGARCH [2497]. realizations [1213]. realized [1897]. recalculation [2301]. recapture [2878]. received [724, 750]. Receiver [1645]. Rechnungen [148]. rechtliche [464]. rechtlichen [208]. Rechtsgrundlagen [211, 92]. Rechtsvorschriften [208]. reciprocal [1736]. recommended [2526]. reconsideration [2155]. reconsidered [717, 253, 318]. Reconstruction [632]. Record [561, 1233, 1365, 615, 1464, 2234, 1103, 1396, 2013, 1462, 1837, 2222, 1063, 1383, 1096, 2817, 1283, 1857, 1955, 753, 1543, 1759, 1323, 1200, 1386, 1106, 1098, 1388]. records [1812, 2621, 1718, 1419, 1386]. recovery [216]. rectangular [1036]. Recurrence [2081, 1449, 1057, 1841]. recurrent [926, 2414]. recursion [653]. recursions [2838]. Recursive [879, 2467, 853]. Reddy [2089]. Reduced [2971, 1562, 1310, 2176]. Reducing [1905, 1265]. reduction [2474, 313, 3056, 2882, 560, 2523, 2928, 2414, 2947]. reduction-based [2947]. redundancy [1669]. Redundant [1187, 1215, 1166].

Referat [72]. Refined [2565]. reflect [1168]. reflection [2192]. reflections [589]. reflective [2580]. Reflexionen [398, 380, 399]. regime [2097]. regime-switching [2097]. region [2803, 2818, 2308, 2484, 856, 1991, 2453]. Regional [480]. Regional- [480]. regionalstatistischen [347]. regions [656, 847, 1553, 2522, 2243]. registrierten [509]. Regression [2237, 2194, 1660, 1537, 693, 2175, 1723, 1739, 2884, 2186, 2799, 1637, 3063, 2437, 2279, 1259, 2693, 1999, 2317, 2397, 953, 1197, 2184, 1564, 1611, 2694, 761, 853, 2791, 2931, 2902, 2656, 2638, 2965, 1971, 2734, 2667, 904, 1832, 2620, 1631, 3053, 1002, 2082, 1797, 1078, 891, 1306, 1784, 2083, 2983, 2652, 2474, 2209, 2924, 1610, 863, 1073, 1935, 2445, 2851, 2498, 2879, 1624, 2815, 2503, 2670, 1690, 2880, 835, 2451, 1193, 2143, 2985, 251, 2032, 1087, 2159, 1108, 2185, 459, 3009, 1348, 1509, 1760, 145, 172, 828, 2483, 2795, 846, 379, 2229, 2403, 788, 2537, 1847, 1370, 1993]. regression [1207, 2068, 2999, 804, 1767, 1948, 2581, 1982, 2865, 856, 1320, 1553, 2671, 848, 3008, 1493, 790, 1727, 199, 868, 1272, 977, 2804, 1252, 2158, 1512, 1614, 1788, 2283, 2348, 2519, 2538, 3002, 3048, 2422, 2611, 2548, 3013, 2093, 2094, 2449, 2339, 2121, 454, 1747, 1770, 2431, 2641, 2426, 2839, 2040, 1531, 2740, 1845, 1844, 2458, 2140, 2702, 2984, 2917, 892, 3056, 1094, 1198, 1391, 2009, 540, 2731, 2742, 2372, 1655, 875, 917, 1516, 1596, 1841, 1342, 2017, 2124, 2036, 641, 1345, 2927, 1194, 2103, 2564, 2753, 562, 2306, 2760, 2575, 2049, 2714, 2852, 966, 353, 382, 2762, 2729, 2347, 1117]. regression [2173, 2288, 2132, 423, 984, 1068, 1641, 2681, 2231, 1920, 3057, 2043, 1384, 1932, 909, 1022, 1183, 825, 854, 908, 1313, 1387, 3058, 2066, 3059, 560, 1741, 3071, 1037, 1172, 1209, 2623, 1744, 1443, 1210, 970, 2020, 2044, 2149, 2142, 2381, 2436, 2439, 2587, 2543, 2549, 2523, 2528, 2539, 1075, 1119, 1310, 2358, 2059, 2256, 2253, 2430, 2342, 2176, 2338, 2396, 2872, 2793, 1563, 2221, 2459, 2841, 2313, 2518, 1689, 2442, 2701, 2877, 2725, 2485, 2669, 2011, 2471, 2637, 2746, 2897, 1636, 1758, 2561, 1813, 563]. Regressionen [96, 104]. regressions [1023, 1936, 2356, 762, 1961, 1154, 1206, 2593, 1065, 1584, 363]. Regressionsanalyse [199]. Regressionskoeffizienten [552]. Regressionsmodell [547, 642]. Regressionsmodellen [217]. Regressor [642]. regressors [1170, 1391, 475, 1478]. Regular [1577, 1211, 601, 2357]. Regularities [772]. Regularization [2121, 2653, 1954, 2630, 2138]. regularized [2611]. related [2775, 762, 2920, 2490, 286, 343, 1498, 2374, 2981, 1157, 605, 1694, 2932]. Relatioinships [686]. relation [752, 723, 803]. relations [2081, 607, 1364, 1057, 1449]. relationship [878, 709, 1925]. relationships [459, 559]. relativ [302]. Relative [1197, 2016, 1227, 1022, 2184, 761, 637, 2734, 1264, 1355, 1369, 1578, 2271, 2148, 1763]. relevance [2930]. relevant [2331, 1391]. Relevanz [272]. Reliability [1166, 407, 548, 2030, 1952, 997, 469, 2604, 1220, 1435, 1732, 958, 1024, 1013, 2962, 2255, 2213, 2386, 2527, 2627, 2241, 1121, 1517, 1058, 1199, 2676, 1684]. remainder [664]. remaining [2208]. remark [625]. remarks [2054, 1435, 538, 2923, 2840, 796, 611, 441, 690, 340, 451]. REML [1405]. removal [2411]. removed [2307]. renewable [2278]. renewal [784, 414, 2672, 2189]. Renyi [1341, 1945]. repair [548]. repairable [1166]. repeated [637, 2816, 2157, 2038, 1580]. Replacement [2716, 1331, 1448, 1181, 2281]. Replicated [1924, 1934, 1308, 966, 854, 1172]. Replication [2698]. Replik [399]. Reply [318, 1774]. report [632, 440]. representation [2257, 1141, 2782, 1805, 1361, 969, 2673, 1926]. representations [2826, 2097, 2337]. representative [2645, 2910]. representativeness [1976]. representing [1055]. reproducing [1710]. Republic [130, 71]. requirements [2270]. Resampling [1566, 2910]. research [1168, 1581, 25, 1685, 19, 109]. Reservelagern [335]. residual [1341, 731, 1288, 721, 756, 1067, 1848, 2387, 988, 1753, 1711, 1553, 2827, 1945, 2946, 3025, 2235, 2377, 2000, 1623, 1251, 2070, 1786, 2202, 2025]. residual-based [2827]. Residuals [459, 942, 959, 2164, 1825, 1591]. resistance [588]. resistant [2010, 2128]. resolution [1476]. Resolving [329]. resource [2270]. respect [2016, 2759]. respiratory [2964]. Respondent [2045]. response [2643, 1001, 1389, 1409, 2535, 1806, 2045, 1002, 1674, 1693, 1565, 692, 2159, 2212, 1397, 1294, 2119, 2900, 2889, 2094, 3052, 2300, 2702, 2984, 2758, 1617, 2972, 2929, 2231, 1122, 1632, 1659, 1172, 1209, 2523, 2134]. response-adaptive [2643]. responses [2603, 2799, 1992, 1674, 2816, 1495, 470, 1578, 286, 343, 2662, 1905, 1453, 2564, 984, 1068, 1209, 3036]. restraints [415]. Restricted [2484, 1637, 2082, 2435, 2308, 2850, 2232, 2283, 1991, 2731, 2742, 2571, 2713, 2524, 1778, 2223, 1931, 1823, 2551, 2450, 1563]. restrictions [2172, 2295, 2658, 1531, 871, 826, 1924, 680, 1483, 2328, 2077, 2859]. Restriktionen [153]. Restringierte [573]. result [1003]. results [1978, 2469, 1605, 664, 2051, 2794, 849, 888, 729, 2148, 584, 1430, 2404, 827, 1277, 605, 779, 1285, 325, 326, 885, 3023, 2357, 2901, 2344, 685]. RETRACTED [2746]. retrial [2260]. return [1285]. returns [2053, 1558]. reversal [717, 486, 525, 524]. Reversed [1401, 1295, 1498, 2222, 2041, 1925]. reversibility [2318]. reverting [3062]. Review [771, 1704, 381, 1749, 757, 1707, 1984, 2367, 1660, 1664, 626, 682, 1486, 1643, 1525, 1663, 2331, 2675, 211, 1834, 1870, 2019, 2113, 2144, 2198, 2147, 2219, 1567, 1708, 2112, 2146, 2218, 2349, 2259, 2239, 1661, 2182, 2197, 971, 1727, 1703, 1746, 1963, 2109, 2104, 2106, 2108, 2432, 2674, 349, 1526, 1645, 1682, 1685, 1771, 2002, 2107, 2555, 1545, 2111, 2145, 1705, 1706, 1985, 1546, 1642, 1665, 1681, 1770, 1890, 1983, 1962, 2089, 2088, 2217, 2431, 2554, 1833, 1943, 2003, 265, 2199, 833, 1726, 2433, 2556, 1504, 1547, 1566, 1683, 1686, 1687, 1891, 1869, 1503, 2368, 2238, 2298, 1487]. Review [464, 2110, 912, 1568, 799, 1725, 805, 780, 789, 787, 783, 1684, 2160, 1524, 1748, 1813, 1724, 2105, 2161, 2162, 3019, 405, 553, 1245, 2194, 755, 328, 1179, 2491, 1062, 824, 1000, 1796, 2824]. reviews [1027, 1150, 1048, 746, 1189, 1202, 1039, 1079, 950, 1099, 1069, 1110, 1381, 778, 1136, 1217, 1230, 960, 1015, 1059, 754, 994, 2072, 1242, 1267, 1289, 1302, 1315, 1327, 1392, 1124, 1425, 1437, 725, 1351, 1410, 733, 1089, 1450, 1279, 1254, 1176, 1338, 1372, 1162, 1005]. revisited [2521, 2033]. Revisiting [3000]. revival [967, 968]. RGW [83, 112, 107, 116]. RGW-Länder [107, 116]. riboflavin [2317]. Richard [18, 1747]. Ridge [571, 2553, 2429, 2991, 2364, 1637, 2437, 1999, 1690, 2232, 1614, 776, 3025, 2372, 875, 2571, 1494, 700, 1618, 2440, 1744, 2059, 2471]. Ridge-type [2553, 2429, 2991, 776]. right [2081, 1104, 1459, 2879, 1492, 1057, 2346, 1515, 1958]. right-truncated [1104]. Ripley [757]. Riquelme [2182]. Risiko [16]. Risikoanalyse [516]. Risk [1391, 859, 1011, 1725, 3015, 2588, 1333, 2906, 1431, 2641, 1145, 1047, 1535, 624, 1684, 2168, 2901, 1725]. risks [2296, 2421, 2336]. RKHS [2574]. Rob [2432]. Robert [1704, 1524]. ROBOUT [2869]. Robust [2603, 3063, 1918, 2697, 1784, 2474, 1550, 2670, 1732, 2185, 1755, 2069, 896, 2501, 1512, 2394, 1358, 2577, 2049, 2860, 2729, 2288, 2812, 1914, 1922, 2887, 2142, 2361, 2601, 2951, 2637, 1552, 1921, 2578, 1234, 1828, 2924, 2819, 2696, 1910, 2380, 1602, 3008, 1915, 1790, 2325, 2752, 2158, 2732, 2494, 2764, 2606, 2372, 2457, 2882, 1950, 2767, 2835, 2480]. robustification [2630]. robustified [1083]. robustifying [2417, 1916]. Robustness [1083, 2805, 2561, 1246, 1221, 3012, 871, 1917]. ROC [2385, 1746, 2584]. Rocci [2110]. Roche [2106]. Rogers [2003]. role [2053, 404]. Rolf [364]. Roling [2899]. Romig [929, 1440, 1612]. root [2382, 1234, 1720, 1413, 1900, 1810, 828, 1484, 1672, 1262, 2048, 2074]. roots [941, 829]. Rost [2055]. Rostamizadeh [2367]. rotation [2466]. Rothe [626]. Rounded [1789, 1877, 1822]. Rounding [1019]. row [2945]. rows [1581]. Rubin [789, 780]. Rudolph [2217, 404]. Rukhin [1020]. rule [1078, 406, 1332, 1391, 2129, 825, 908]. rules [1967, 2315, 450]. Run [1658, 1967, 2632, 876, 679]. runs [2592, 1298, 2307, 2210, 1981, 1322, 1712]. Rüstungen [151]. Ruud [1661]. Ryan [1770, 1487].

S [1486, 1525, 2056, 1985, 2554]. S-PLUS [1566]. S. [349]. s00362-007-0113-0 [1728]. Saarbrücken [6]. Saarlandes [6]. Sabina [2106]. Sachgütererzeugung [135, 170]. saddlepoint [2075]. Sage [1704]. Saisonbereinigung [636, 233]. Saisonbereinigungsverfahren [236]. Saisonfigur [297]. Saisonschwankungen [147, 155]. Saito [1681]. Samanta [1547]. same [762, 1827]. Sample [705, 925, 346, 2482, 2272, 1127, 664, 2228, 2021, 1276, 3064, 1083, 1234, 764, 1265, 1787, 1093, 1063, 1007, 2091, 749, 755, 991, 1853, 998, 1155, 845, 2500, 1989, 1794, 765, 667, 245, 1320, 1875, 1620, 2671, 2386, 1008, 1789, 691, 2854, 2811, 751, 1887, 2954, 1020, 2934, 1656, 1729, 1680, 773, 2347, 978, 2721, 2801, 3066, 2470, 2301, 1460, 1744, 157, 2612, 2183, 1131, 2942, 1988, 1926, 2471, 1758, 2661, 1685]. sample-splitting [1093]. sampled [1674, 2891, 1431]. sampler [1443]. samples [2402, 648, 637, 872, 242, 2498, 1514, 766, 3067, 1348, 316, 2644, 1872, 1522, 2417, 989, 1647, 1633, 2379, 513, 675, 2018, 2706, 2648, 2510, 1285, 1619, 229, 2901]. Sampling [1463, 1529, 1846, 2831, 1224, 1395, 2707, 737, 2626, 3062, 1830, 2604, 811, 891, 1693, 2562, 1638, 991, 786, 1448, 1734, 1169, 766, 2191, 1821, 1087, 1247, 860, 1181, 2608, 964, 906, 1807, 1532, 1692, 929, 1612, 1362, 2241, 2428, 2584, 1441, 2281, 1517, 2529, 1779, 2178, 1759, 773, 517, 608, 774, 2287, 600, 1122, 1384, 1481, 1632, 1659, 1866, 2030, 1990, 1609, 1769, 1433, 1226, 2384, 625]. Sampling. [799]. samplings [1032]. Sandrine [1568]. sandwich [2061]. Santos [2106]. SAR [3065]. SAS [1503, 2002, 1645, 1813]. Sato [525]. saturated [2569]. Saunders [1738, 2359]. Sawitzki [1833]. SB [1790, 2325]. SB-robust [1790, 2325]. scalar [2159]. Scale [2787, 1846, 2207, 1330, 2092, 1860, 1054, 2905, 162, 1114, 1107, 538, 2143, 227, 999, 1620, 1008, 513, 1947, 1756, 2811, 2466, 850, 2085, 2504, 2509, 2075]. scale-free [2811]. scale-invariant [2905]. scale-like [2075]. scale-shape [2509]. scaled [1319]. scaling [2684]. scatter [2811]. scattered [2276]. Schätz [614]. Schätz- [614]. Schätzer [529, 642]. Schätzfehler [387]. Schätzfunktionen [181, 270]. Schätzung [156, 191, 297]. Schätzungen [654]. Schätzverfahren [512]. Schätzwerte [217]. Schätzwerten [137]. Schema [22]. scheme [1734, 1475, 774, 1632]. schemes [907, 2084, 1759, 773, 935, 679]. Scherer [1486]. Schichtung [196]. Schindler [1292]. Schluß [44]. Schmeidler [969]. Schmidt [664]. Schneeberger [912]. Schneeweiß [682]. Schoffer [1503]. Schott [2109, 2298]. Schranken [317]. Schreiber [246].

Schriften [480]. Schwankungsbereiche [394]. Schwarz [1010, 1655, 2995]. Schwierigkeiten [26]. Science [2675, 2368, 830, 2949]. Sciences [1526]. scientist [330]. Score [2708, 2821, 1610, 1369, 2649, 1334]. scree [2925]. screening [2365, 2526, 2560, 2400, 2750, 2994, 3036, 2545, 2625, 818]. search [2693, 2733, 406, 964, 1922]. searches [814]. searching [279]. Seasonal [522, 635, 2505, 556, 1366, 706, 1227, 2350, 2506]. seasonality [1413]. Seat [1456]. secant [2053]. Sechs [51]. Second [2198, 2218, 1726, 2470, 1784, 2367, 2749, 2513, 2135, 2598, 1962, 1916, 2288, 2518, 440]. Second-order [2470, 1784, 2749, 2135, 2288, 2518]. Section [2115, 1301, 1123, 1135, 1161, 1175, 1188, 1201, 1216, 1241, 1350, 1393, 1424, 1438, 1451, 1488, 1505, 2114, 1326, 1229]. sectional [1901, 2864, 3007]. SeDuMi [2358]. Seemingly [2753, 1333, 2641]. segment [2457]. segmentation [2568]. seiner [289]. seit [5, 464]. Sektoren [336]. Selected [2206, 2207, 316, 2150, 1160]. selecting [860]. Selection [1982, 2610, 2092, 2913, 1930, 2291, 1935, 1961, 2708, 1735, 888, 2246, 2185, 2688, 1332, 2804, 2726, 2718, 2170, 2538, 2647, 3051, 691, 2622, 2810, 2861, 2449, 2691, 1375, 1964, 1919, 2653, 2713, 2352, 2138, 1808, 2205, 2287, 2507, 2990, 1745, 2361, 2436, 2539, 2443, 1929, 2338, 2065, 2480, 2841, 2919, 2709, 2761, 2547, 2868]. selectors [846]. self [2609, 2570, 2886]. self-exciting [2886]. self-normalization [2609]. self-weighted [2570]. selling [2129]. Semi [2333, 2581, 3016, 933, 1412, 1603, 2879, 1465, 2909, 1468, 2223, 2154, 2582, 2947]. semi-Burr [1603]. semi-linear [1412]. semi-orthogonal [2909]. Semi-parametric [2333, 2581, 2879, 2154, 2582]. semi-Pareto [933, 1603, 1468]. Semi-supervised [3016, 2947]. semi-varying [2333, 2223]. semicontinuous [2969]. Semifoldover [2624]. Semiparametric [2896, 3060, 2188, 2317, 1306, 2515, 3065, 3041, 2652, 2464, 2225, 3047, 1643, 2594, 2956, 2845, 2895, 2122, 1528, 2049, 2628, 2557, 2963, 2149, 2381, 2253, 2343, 2948]. semistable [1491]. semivaluations [500]. semivariance [3014]. sensitive [1806, 2045, 1674, 1693, 2924, 1565, 1717, 1995, 2212, 1349, 2366]. sensitivity [2215]. sensor [2269]. separated [1781]. separation [1916]. sequence [1658, 2635, 1981, 730, 2651, 2957, 772, 2020, 1926]. Sequences [3003, 1766, 2418, 2735]. Sequential [1055, 2797, 976, 699, 2230, 1536, 2705, 1471, 1629, 786, 2880, 2461, 1753, 1872, 1561, 729, 775, 2938, 715, 2026, 2972, 2774, 2980, 1072, 2919, 2565]. sequentially [2531]. sequentiellen [16, 645]. serial [3068, 2266, 1133, 1939]. serially [909, 1022, 1183]. Series [1644, 2198, 2218, 1705, 1706, 932, 3019, 405, 2463, 2035, 2012, 2638, 628, 1491, 762, 905, 2167, 2445, 2027, 1643, 2682, 2664, 2590, 925, 2112, 2146, 2349, 429, 279, 361, 2455, 2537, 2748, 2271, 1263, 2153, 803, 2362, 1064, 2662, 1477, 2649, 3005, 2460, 2728, 2554, 2479, 2458, 1916, 3070, 1939, 1504, 1869, 2099, 2843, 2319, 2882, 2205, 416, 1966, 2318, 1171, 2928, 2163, 2951, 3075, 2582, 2162]. servers [2260]. set [1529, 1846, 2831, 1224, 1395, 1564, 1611, 2707, 2562, 2191, 2608, 1872, 1807, 1532, 1789, 2241, 2428, 2584, 2706, 2886, 1249, 2529, 2178, 2384]. SETAR [2063]. Seth [2106]. sets [2172, 543]. settings [2764]. Several [361, 2820, 1648, 1213, 1981]. Severini [2555]. Shakil [1523]. Shannon [500]. Shao [1685]. Shape [1910, 1846, 1046, 2010, 2447, 2125, 1481, 2509, 1388]. shaped [1634]. Shapleyschen [210]. Sharing [1147, 2829]. Sharp [2078, 2438, 2572, 2541]. Shein [1685]. Shein-Chung [1685]. Shewhart [1836, 1064, 876]. shift [2919]. shifted [2969, 852]. shifts [1897, 2827, 1783]. Shirley [1834]. Shiryaev [2265]. Shock [784, 1959, 2672, 2189]. short [1737, 1263, 2683, 1939]. shortest [723, 2715]. shot [2534]. shot-noise [2534]. shoulder [1830]. showing [2542]. Shrinkage [3047, 1829, 2269, 2799, 2553, 2791, 2578, 2246, 1332, 1129, 2718, 2009, 2372, 1228, 1481, 3034, 2612, 2343, 2640, 2429, 2637]. shrinking [2507]. shrunk [1310]. Sicht [626, 10]. sided [705, 998, 2419, 1769, 1670, 2565]. siebziger [367]. Siegel [526]. Sigma [1796]. sigmoidal [2776]. sign [695, 2752, 1823]. sign-change [1823]. signal [2887, 1075]. signal-to-noise [1075]. signals [1510, 2084, 2673]. Signature [2360, 2962]. Signature-based [2360]. signed [1864]. signed-rank [1864]. Significance [641, 656, 2287]. Silvia [1726]. Simcer [72]. Simes [2174]. similar [327]. similarity [2010]. Simon [2003]. Simple [1265, 1263, 2752, 2643, 556, 660, 853, 2667, 572, 2633, 740, 172, 2286, 2079, 751, 1866, 2344]. simplicial [2127, 2311]. simplicity [1867, 389]. simplification [570]. Simulated [1429, 2267, 1434, 2952, 1905]. Simulation [1637, 2491, 942, 959, 729, 2271, 2424, 1844, 217, 1285, 757]. simulation-based [1844]. Simulations [2829, 755, 2695, 2908]. Simulationsstudie [398, 380, 399]. Simultaneous [1330, 2047, 974, 1167, 1638, 847, 2781, 1861, 1823, 2525, 2004, 1107, 822, 2084, 2486, 2361, 2453, 2599]. simultaneously [612]. Sine [1721]. Sine-skewed [1721]. Singer [833]. Singh [625]. Single [171, 2068, 2442, 2603, 2532, 2245, 2212, 1597, 929, 1612, 2170, 2956, 2662, 3013, 2927, 1587, 21, 2618, 2933, 2835, 2547]. single-channel [21]. Single-index [2068, 2442, 2532, 2245, 2956, 3013, 2927, 2933, 2835, 2547]. Singpurwalla [1684]. singular [1270, 1562, 871, 1483, 2256]. Sinne [504]. sinusoidal [2673]. SIRS [2464]. Situation [165]. situations [462, 861]. Sitzung [591]. Six [1796, 870, 584, 2612]. Size [1192, 1685, 2228, 2707, 1276, 1648, 1063, 625, 1181, 1794, 1597, 829, 1913, 2721, 600, 2470, 2301]. sizes [2402, 705, 765, 346, 2632, 773, 738, 2558]. Skalierung [453, 471, 528]. skew [1702, 2737, 2655, 2847, 1653, 2008, 1860, 1631, 1862, 1768, 2608, 1699, 2200, 1767, 2844, 1630, 2732, 3055, 2825, 1640, 1816, 1804, 1933, 1499, 2509, 2949]. skew- [2732]. skew-elliptical [1933, 2949]. skew-normal [1702, 2655, 2847, 1631, 1768, 1699, 2200, 2825, 1499, 2509]. skew-symmetric [1816]. skew-symmetric-Laplace [1804]. skew-symmetry [2608]. skew-t [1630]. skewed [1721, 3026]. skewness [2738, 801, 807]. Skillicorn [1642]. skin [2467]. skip [906]. Skyler [2674]. slash [1862]. slice [2897]. Sliced [1920, 2528, 2474, 2423, 2756, 2142, 2558]. slices [2558]. slicing [2897]. slope [2399]. Slow [1963]. SLS [2258]. slug [1581]. Small [664, 749, 755, 1155, 2348, 1020, 1744, 586, 2950, 2272, 3064, 872, 1184, 1794, 2333, 503, 1950, 1285, 229, 3035]. Small-sample [1020, 3064]. smaller [2805]. smallest [354]. Smeeton [2147]. Smirnov [998, 1989, 1036]. Smith [1724]. Smooth [2584, 2834, 2502]. Smoothed [2854, 1340, 1595, 2692]. smoothers [2342]. smoothing [932, 2712, 2354, 759, 1914, 2848, 2476, 2631, 2370]. Sobol’ [2481]. social [450]. socialist [98]. socio [2113, 1703]. socio-economic [2113, 1703]. Softly [1310]. softmax [2313]. softplus [2893]. software [954, 2120, 2162]. Solka [1962]. solution [2643, 923]. solutions [2432, 680, 1444]. Some [2317, 2655, 644, 2051, 483, 764, 1435, 1419, 2920, 538, 653, 2923, 2817, 386, 2408, 429, 796, 1018, 1356, 982, 1458, 657, 782, 2131, 2148, 611, 584, 1581, 1430, 2404, 205, 753, 1130, 2039, 943, 827, 2327, 1277, 690, 43, 805, 1588, 2376, 340, 839, 1741, 2156, 1607, 2468, 2251, 2357, 589, 1120, 1335, 1365, 615, 2469, 2437, 2931, 915, 727, 755, 1293, 2818, 1734, 306, 2490, 934, 1810, 379, 1192, 729, 2581, 559, 1170, 2325, 1654, 1295, 1311, 1554, 2731, 2742, 880, 1916, 605, 1058, 2995, 449, 700, 908, 560, 1209, 3043, 2798, 2847, 441, 451]. Sonia [2162]. Sotoca [2162]. source [1916]. Sowjetunion [249]. sozialen [209]. Sozialforschung [833]. sozialistischen [52, 72, 124, 65, 98]. sozialistischer [190]. Sozialstatistik [10, 388, 603, 179]. Sozialwissenschaften [376]. SP06 [1371]. Space [2875, 2097, 1495, 2091, 1710, 2820, 956, 2162]. Space-filling [2875, 2820]. spaced [2577]. spaces [2010, 2159, 1142, 2782, 2457]. spacing [242]. spacings [578, 1605, 2801]. Spaltensummen [579]. spares [407]. Sparse [2834, 2488, 2229, 2768, 2496, 2596, 2988, 2726, 2538, 2449, 2678]. sparsity [2797, 2945, 2031]. Spatial [2351, 1743, 1950, 1704, 2184, 1936, 3041, 2532, 1935, 2911, 1482, 1975, 2292, 1767, 2844, 1652, 2647, 2845, 3060, 3051, 2861, 2536, 2785, 2049, 2714, 2963, 2978, 2276, 2933, 2443, 2605, 3007]. spatial-angular [2276]. spatially [1108, 1183]. spatio [2749, 2911, 2956, 2096, 2556]. spatio-temporal [2749, 2911, 2956, 2096, 2556]. spatiotemporal [3065, 2536]. SPD [3017]. Special [2116, 1139, 2179, 2299, 944, 2454, 2757]. specific [361, 2292, 1808]. Specification [1009, 2842, 1550, 814, 3038, 205, 1769]. specified [2250, 2210]. Spectral [3061, 1942, 2330, 2661]. spectrum [2728]. Spektral [385]. Spektraltheorie [161]. Spezifikationen [456]. Spezifikationsfehler [539]. spherical [2826, 483, 2697, 2904, 1970, 3009]. spherically [1553]. SPI [531]. Spline [573, 2712, 3028, 2337, 2782, 2170, 3002, 3011, 1801, 2852, 2848, 2788, 2631]. Splinefunktionen [636]. splines [1368, 2163]. split [1571, 2492]. split-plot [1571, 2492]. splitting [1093]. SPLUS [1486]. Sports [2555, 2555, 1661]. spreads [1937]. Springer [2019, 2108, 1666]. spurious [1380, 641]. Square [386, 580, 2970, 1153, 1002, 2403, 1557, 1678, 2294, 1228, 2511]. Squared [2718, 1197, 1012, 3015, 920, 2150, 2980, 779, 1763]. squared-radii [920]. squares [1637, 2950, 669, 762, 1306, 1784, 2700, 2135, 334, 1553, 848, 643, 2242, 1621, 2289, 2641, 3025, 1480, 2860, 2288, 475, 1044, 2388, 1931, 825, 908, 1588, 1346, 2867, 1131, 2518, 2544]. Staaten [84, 249]. Staatssektors [123]. Stabilitätsproblem [344, 40]. Stability [1325, 2452, 716]. StabKoff [2877]. stable [2931, 1533, 416, 1074]. Städte [467]. stage [1318, 2155, 1638, 860, 964, 1561, 1780, 1044, 2370]. Stahlexport [75]. stamps [1588]. standard [727, 1167, 1592, 919, 2120, 1811, 1994, 2813, 1769]. standardisierter [290]. Standardmodell [362]. Standardstrukturen [261]. standby [1166, 548]. Standorten [368]. STAR [2502, 2495]. starshaped [918]. Start [1987]. Start-up [1987]. starter [2554]. State [1495, 924, 2097, 1898, 2765, 2278, 2162]. state-space [2097, 2162]. states [1095, 478]. static [2770]. stationarity [415, 872]. Stationary [905, 2335, 414, 2914, 2703, 2001, 2619, 1479, 1915, 2456, 827, 2265, 2190, 1378]. statischen [394]. statistic [2482, 2970, 895, 2926, 327, 279, 2748, 2362, 2516, 920, 2479, 1292, 1036, 3066, 807, 157, 1463, 2095]. Statistical [1728, 2234, 954, 1558, 1046, 1379, 2776, 1185, 2503, 2695, 791, 468, 2073, 672, 472, 1570, 1723, 1747, 1857, 1134, 2908, 2929, 2223, 43, 1388, 1636, 2730, 586, 2975, 494, 570, 1832, 1984, 2620, 2856, 1383, 366, 1649, 770, 843, 686, 2461, 1663, 2206, 2923, 2447, 420, 2192, 1252, 1868, 1358, 2959, 866, 27, 1130, 2377, 2945, 1796, 2995, 2110, 774, 813, 560, 2444, 2602, 2468, 2565, 440, 1834, 2144, 1661, 2431, 1906, 789, 1607, 25]. Statistics [413, 2147, 1526, 2555, 2111, 2648, 2556, 1566, 1286, 1613, 1812, 2058, 2469, 2722, 2272, 2081, 1103, 752, 2228, 1396, 2051, 2862, 1662, 2525, 1104, 1205, 1360, 1610, 1658, 1019, 682, 987, 129, 1761, 316, 2754, 1522, 2541, 445, 667, 1530, 2968, 3069, 2432, 1672, 2745, 1719, 1057, 2145, 1334, 1541, 2164, 3, 1833, 1020, 1312, 2711, 1223, 1504, 1543, 1759, 1601, 2938, 2346, 1445, 1739, 1855, 3042, 805, 1859, 1449, 1444, 1619, 1795, 2951, 2161, 1917, 1996, 2661, 799, 1974, 2, 1703, 1771, 1681, 2298]. Statistik [6, 4, 413, 51, 5, 211, 272, 117, 2109, 3, 536, 183, 912, 92, 116, 151, 1666, 626, 2, 19, 192, 360, 581, 495, 214, 291, 2105, 372]. Statistiken [250]. Statistikers [50]. statistiques [586]. statistisch [532]. statistisch-entscheidungstheoretisches [532]. Statistische [366, 1503, 309, 53, 166, 123, 595, 606, 883, 80, 160, 25, 381, 231, 337, 642]. Statistischen [591, 73, 165, 44, 121, 211, 536, 140, 403, 277]. statistischer [377, 432, 580, 431]. statistisches [594, 264]. status [688, 1431]. steady [2278]. steering [232]. Steffen [2104, 1708]. stein [825, 1078, 1062, 1198, 1391, 2372, 859, 1011, 1516, 966, 854, 908]. stein-rule [825, 1391, 908]. Stein-type [2372, 859, 966, 854]. Stein-variance [1011]. Steland [1666]. Stellung [135]. Stellungnahme [399, 337]. step [2880, 2245, 2073, 1754, 2023, 1953, 2855]. step-stress [2073, 2023, 1953]. Stephen [1526, 2088, 1943]. stepwise [2986, 804, 879]. Stewardson [1834]. Stichprobe [185, 307]. Stichproben [479, 912]. Stichprobenentscheidungen [16]. Stichprobenproblem [649].

Stichprobentheorie [196, 270, 305]. Stichprobenumfang [506]. Stichprobenumfangs [185]. Stichprobenverfahren [269, 590]. stick [2292]. stigmatized [1122]. stigmatizing [1453]. Stochastic [2273, 2738, 2285, 2322, 2182, 2274, 2536, 2450, 1401, 1952, 2172, 1084, 1180, 602, 2208, 2028, 2249, 2378, 2232, 543, 1295, 1614, 3038, 2571, 122, 1679, 2238, 2938, 635, 475, 1641, 2591, 1399, 2376, 2328, 2551, 1619, 229, 1563, 3035, 2678, 2495, 757, 2108, 1724]. stochastische [122, 136, 1567, 531]. Stochastischen [596, 425, 22]. stochastischer [545, 456]. stock [1533]. stopping [2772, 2521, 1454, 2129]. Störgrößenprozessen [547]. strata [1946]. Strategie [171]. Strategien [113]. strategies [2317, 1226]. strategy [2613]. Stratification [2614, 2425, 991, 964, 505, 517, 608, 665, 2251]. Stratified [766, 1638, 1397, 2906, 1990]. streams [2237]. Strecker [682, 912]. Streifzug [166]. strength [469, 1886, 1732, 958, 2255, 2213, 2386, 2527, 2627, 2360]. stress [1886, 2960, 1732, 958, 1024, 2073, 2023, 2255, 2213, 2386, 2527, 2627, 2360, 1953, 1817]. stress-strength [1886, 1732, 958, 2255, 2213, 2386, 2527, 2627, 2360]. Streuung [348]. string [1781, 2258]. strings [1885]. Strong [2734, 1207, 2984, 475, 2602, 2901, 1259, 705, 2657, 3050, 1390, 2247, 1845, 1700, 1884, 2204, 2979]. strongly [2870, 2459, 1917]. Structural [648, 687, 287, 355, 286, 607, 808, 503, 2802, 353, 382, 410, 1894, 602, 620, 226, 313, 1900, 1429, 650, 2389, 2642, 1192, 1597, 822, 2362, 1044, 477, 1282, 340, 1713, 314, 1896, 189, 1895, 2014]. Structure [2476, 2816, 2744, 2944, 2883, 699, 3040, 2567, 814, 1031, 2441, 1368, 748, 2274, 2352, 2294, 2440, 2361, 885, 2399, 2599]. Structured [1946, 1580]. structures [2423, 171, 12, 2853, 1681, 920, 421, 1966]. Struktur [91, 371, 273]. Struktur-Inferenz [371]. Strukturen [171]. Strukturierung [444]. Strukturwahrscheinlichkeit [103]. Strukturwandlung [115]. strutural [396]. Student [308, 327, 2924, 1533, 856, 1559, 1342]. Student- [856, 1342]. Studenten [491]. Studentization [947, 979]. Studien [2055]. Studienbücher [833]. Studienfachkombinationen [508]. Studienleistung [369]. Studienprogramm [413]. studies [1921, 2262, 1033, 1779, 2583, 1707]. study [2831, 1637, 413, 2773, 2733, 2850, 129, 2491, 260, 788, 1910, 1170, 2698, 1652, 866]. Stuvel [574]. sub [2242, 1587]. sub-class [1587]. subclasses [326]. Subdata [2709, 2761]. subexponential [2204]. subglobal [1056]. Subgroup [2900]. Subhabrata [1984]. subject [2391, 497]. subjected [1024]. submission [647]. subpopulation [2029]. subroutine [688]. subsampling [2983, 2763, 2671, 2764, 2860, 2762, 2681, 2793, 2313, 2824, 2701]. subsequences [1973]. subsequent [1885]. subsets [562]. subspace [2616]. subspaces [1474]. Substitutionsaxiom [167]. subtraction [2735]. success [1629, 1298, 1981, 1322]. successes [1781, 1885]. successive [329, 1659]. such [1168]. Suchverfahrens [280]. Sudderth [684]. sufficiency [2408, 796, 2131, 1582]. Sufficient [1971, 1530, 2414]. suffiziente [250]. suggestions [727]. suitability [137]. sum [2171, 1553, 1519, 1518, 2052]. Suman [2106]. summands [2204]. Sums [1344, 1978, 2652, 1970, 1883, 2663, 1973, 3025, 2263, 1700, 1757, 1884, 2543, 3043, 2602, 2468, 1697]. Sung [2543]. Sunter [1323]. superfluous [908]. superharmonischen [252]. superpopulation [1628]. supersaturated [1992, 2492]. Supervised [2928, 2977, 3016, 2947]. supplement [439]. support [1307, 2542, 2945, 2705]. supremum [2204]. Sure [2545, 1883, 1004, 2887]. surveillance [2034]. Survey [591, 1389, 1409, 1374, 954, 1265, 891, 1995, 1364, 1010, 1888, 1292, 1453, 1052, 720, 1384, 2044, 1752]. surveys [1001, 2045, 1002, 2708, 2487, 2366, 2615, 1617, 1287, 1460, 780]. survival [878, 2560, 1128, 980, 1657, 2962, 2517, 1009, 2393, 2510, 1920, 2935, 1694, 2625, 2948]. suspected [1865]. swan [1687]. switching [2452, 2097]. SYMARMA [2203]. symmetric [2621, 2397, 2926, 1360, 2964, 1610, 867, 1142, 2817, 1621, 2428, 2203, 1816, 1804, 2102, 2599]. symmetrical [1347, 1650, 2985, 245, 2936]. symmetrically [2308]. Symmetriebegriffe [305]. Symmetry [1668, 2226, 2580, 587, 570, 2727, 2744, 1977, 2608, 2616, 1670]. synchronized [2137]. Synthetic [1628, 2777]. System [469, 468, 1208, 997, 1166, 2016, 1013, 2962, 2999, 2487, 2409, 1623, 2278, 1199, 407]. systematic [3019, 2296]. Systematiken [80]. Systematisierung [80, 127]. Systeme [190, 833, 53, 198]. Systemen [69]. Systemgedanke [61]. Systems [787, 315, 768, 2273, 1972, 2699, 1848, 1214, 938, 1753, 798, 2322, 2446, 1455, 2235, 2360, 37, 2270, 548, 2486, 288].

t [1630]. Tabelle [163]. Tabellen [336, 290, 123, 370, 417, 452, 451]. table [1537]. tables [936, 599, 1247, 421, 417, 813, 271, 451, 1670]. tabulations [327]. Tadeusz [1570]. Tail [1285, 2679, 2167, 2027, 2955, 2588, 2598, 2710, 806, 2917, 753, 1939, 1880, 2901]. Tailed [3003, 2249, 2570, 1714, 2575, 1679, 1803, 1951]. tails [863, 1555, 2637]. Takayuki [1681]. Taking [420]. Taleb [1687]. Talwalkar [2367]. Tamhane [1771]. Tanaka [2349]. tangent [708]. Tao [1890].

Tapas [1547]. tapering [1528]. target [1078, 2781]. targeting [2420]. Taschenbuch [264]. tau [2853]. tau-based [2853]. taxonomic [421]. Taxonomy [417, 451]. Technical [265]. Technik [99]. technique [1674, 2633, 1591]. Techniques [1723, 1389, 1409, 1832, 2212, 2366, 2121, 1616, 2779, 690]. Technischen [338]. technologischer [35]. Teil [165, 230, 244]. Teilen [92]. Teilmedian [368]. teletraffic [1594]. temperature [2271]. Templ [2433]. temporal [2749, 2911, 2864, 803, 2956, 2203, 2096, 2556]. tensor [2261]. term [1368, 817, 885]. terms [1305, 3068, 1094, 2205, 207]. Terms-of-trade-Effekts [207]. Terrel [1035]. Test [1611, 2540, 434, 3006, 498, 460, 296, 463, 3003, 630, 1980, 695, 895, 2172, 2382, 1471, 1901, 2051, 455, 587, 1245, 915, 1830, 2905, 2604, 1927, 2690, 2876, 1418, 2609, 1720, 2842, 2924, 1592, 2747, 2790, 2960, 2174, 998, 525, 2296, 1087, 736, 1154, 2390, 1989, 260, 279, 1785, 2754, 2748, 2073, 2598, 1484, 1385, 87, 955, 2286, 667, 1129, 1647, 1692, 919, 1025, 1780, 1672, 977, 2392, 2240, 3038, 808, 2665, 2374, 1756, 2809, 1531, 1308, 2966, 2672, 1009, 2372, 2954, 859, 917, 1011, 2280, 1237, 2000, 2934, 871, 1047, 2178, 1680, 1796, 2347, 524, 1036, 2026]. test [2318, 2287, 2676, 3014, 1864, 1444, 945, 2439, 2587, 1072, 2074, 2489, 3001, 2134, 2025, 2563, 2582, 1622, 2948, 2599, 552, 298]. test-based [1796]. Testing [1728, 2355, 1901, 1936, 2890, 941, 1615, 2727, 1593, 2744, 980, 2608, 2616, 2823, 3030, 1875, 716, 584, 2686, 2164, 2266, 2792, 3031, 2085, 1554, 2930, 1791, 2646, 1939, 1157, 2070, 1568, 1783, 2323, 229, 2075, 1839, 2399, 2680, 2320, 1921, 1076, 1412, 1055, 2983, 2508, 1900, 2986, 2741, 2904, 2819, 2490, 3040, 2137, 942, 959, 2865, 1187, 1215, 919, 1874, 1865, 1321, 2290, 1771, 2739, 2419, 2809, 1556, 2916, 1116, 2477, 2917, 1466, 2961, 1604, 1680, 1485, 990, 2189, 1953, 2650, 3046, 3014, 1422, 1173, 1713, 2067, 2095, 2678, 2780, 2599, 2687]. Testmodell [479]. Tests [1952, 1067, 637, 622, 2629, 2677, 180, 1587, 3057, 1225, 221, 2978, 1670, 2193, 2580, 2505, 2397, 2272, 2021, 2282, 3064, 717, 872, 1894, 1083, 2697, 1234, 2552, 2335, 2885, 3045, 1413, 749, 2741, 916, 2808, 2944, 2682, 656, 843, 486, 152, 967, 968, 1977, 1369, 598, 1809, 459, 1810, 1155, 828, 1179, 2491, 1975, 2588, 2642, 1192, 906, 1168, 657, 1597, 3068, 245, 2636, 1343, 1008, 742, 2192, 2379, 1915, 1252, 2752, 1899, 2594, 1065, 2906, 2018, 829, 801, 2706, 2373, 2122, 2395, 3052, 1308, 1583, 2123, 2777, 1292, 2648, 1655, 641, 2564, 1052, 2319]. tests [2607, 1987, 978, 1262, 2048, 449, 715, 1056, 2329, 2774, 1422, 518, 237, 2692, 1896, 2227, 2080, 2166, 2251, 2384, 1253, 1895, 2495, 2565, 2859, 1434, 1115, 412]. Textbook [2147, 1526]. texts [2108]. Textual [2368]. TFR [1274]. th [1848]. Their [189, 838, 2463, 644, 1837, 723, 767, 1458, 775, 2718, 2522, 3050, 2252, 1144, 2981, 3025, 1585, 2039, 1855, 2180, 1795, 2468, 2324]. theorem [1148, 1621, 2668, 1973, 2377, 684, 969, 2759, 794]. theorems [2619]. theoretic [824]. theoretical [2920, 767, 205]. theoretischer [191]. Theorie [31, 89, 60, 280]. Theories [572, 620]. Theory [2056, 1813, 2989, 1952, 1662, 1749, 717, 1306, 2113, 445, 2432, 2002, 1308, 3072, 751, 612, 331, 1547, 2566, 2612, 743, 2162, 450, 2199, 783, 2349]. theory-principles [1749]. there [1744]. Theus [1546]. thinking [1661, 1963]. thinning [2355, 2086, 2350, 2506, 2254]. Third [2147, 2597]. third-order [2597]. Thomas [1682, 2239, 2555, 1770, 1487]. Thöni’schen [432]. Thorn [2198]. Three [1610, 525, 1272, 24, 639, 430, 599, 486, 1878, 489, 2632, 2624, 2613, 1987, 1698, 1670, 2344, 450]. three-level [2632, 2624, 2613, 1987, 1698]. three-parameter [639, 1878, 489]. three-way [1670]. threshold [2262, 2993, 2914, 1712, 2886, 1817, 2254]. thresholding [2568]. ties [3018]. Time [1491, 900, 1644, 2218, 906, 2871, 1705, 314, 932, 1165, 3019, 405, 1447, 2463, 2051, 2035, 2638, 717, 2525, 1366, 2829, 2775, 762, 2689, 905, 2776, 2167, 2445, 2435, 2027, 2378, 1643, 2682, 1429, 693, 2590, 1206, 2891, 706, 429, 928, 279, 361, 2455, 2537, 2748, 2716, 2271, 1263, 2286, 2636, 2153, 803, 2362, 1064, 2165, 793, 2662, 1477, 2622, 2649, 3005, 2093, 2460, 2728, 2809, 2458, 1916, 2751, 3070, 1939, 2099, 2843, 2319, 407, 2882, 879, 1966, 2318, 1171, 2507, 1861, 2978, 3071, 2928, 2163, 2951, 3075, 2582, 3007, 2162, 2112, 2146, 2349, 1706, 2554, 1869]. Time-discrete [900]. time-homogeneous [2829]. time-restricted [2435]. time-series [2455, 2271]. time-varying [1206, 2507, 2978, 3007]. times [1464, 1957, 1959, 1454, 2305, 1535, 1200, 901]. Tobit [758]. today [682]. Toeplitz [2617]. Toeplitz-type [2617]. tolerance [1233, 668, 705, 2666, 856, 1553, 346, 1968]. Toleranzgrenzen [674]. Tony [1834]. too [2173]. tool [2262, 2044]. toolbox [1895]. Tools [1145, 793]. topologischen [456]. Topp [1761]. torus [2743, 2371]. total [1001, 2051, 2403, 2215, 1453, 2610, 2029]. total-effect [2215]. tracking [2837, 1914]. trade [1735, 207]. trade-off [1735]. traditional [570]. traffic [798]. train [2261]. transect [1830]. Transfer [3041, 3005]. Transferfunktionen [397]. transfers [918]. transform [1288, 1236, 2998, 2123, 2265]. Transformation [96, 104, 405, 2515, 2677, 1652, 1239, 1442, 1863, 1839, 1694]. transformations [2448]. transformed [1132, 2724, 1562, 2348]. transforms [2021, 1782]. transient [816]. transienter [252]. transition [2502]. Translationsinvarianz [97]. transmuted [2093]. Transvariation [48]. treatment [1571, 2773, 2832, 916, 2930, 2927, 1157, 2128, 2812, 783]. treatments [2741, 2658]. tree [2467, 2324]. tree-based [2467]. trees [2210]. Trend [706, 3003, 2516, 1679, 2129, 2495]. Trendbereinigung [433]. Trends [1570, 2012, 927]. tri [2909]. tri-factorization [2909]. trials [2643, 1574, 2155, 1658, 1298, 2981, 2236, 2935]. triangular [2004, 605, 2026]. triangularly [2005]. Triangulation [336]. trigamma [2932]. trigamma-free [2932]. trigonometric [1970, 2483]. trimean [2078]. trimmed [1514, 1964, 1919]. Trimming [328]. Trindade [2162]. triparametric [1250]. trivial [3062]. true [623, 553, 661]. truncated [1428, 1889, 2081, 1540, 1777, 1765, 1104, 1459, 878, 1762, 2412, 1882, 2805, 1492, 2976, 1025, 2022, 2057, 1057, 1740, 956, 1251, 2341, 743, 2338, 2872]. Truncating [1679, 2685]. truncation [2382, 664, 2035, 2734, 2445, 2435, 3037, 2179, 2976, 1025]. TRV [1274]. Tsallis [2327]. Tschiersch [1503]. Tucker [2554]. Tukey [1594, 2404]. Tukey-type [1594]. tuning [1982, 2713]. twice [2656]. twist [1903, 992, 910]. Two [2021, 722, 2879, 1470, 197, 2245, 1989, 2787, 2252, 1044, 2370, 1758, 924, 2272, 857, 1611, 821, 1648, 2620, 2155, 1797, 2876, 1507, 705, 1638, 1781, 1854, 890, 1185, 1853, 2633, 2308, 998, 1821, 2005, 1003, 1809, 2137, 2805, 1578, 964, 1998, 1522, 1561, 1699, 1557, 1754, 667, 2148, 803, 2796, 1008, 1780, 2079, 2448, 2632, 3048, 1469, 675, 2018, 2854, 1947, 3025, 1559, 850, 1455, 1371, 1249, 2648, 2954, 2934, 2524, 1838, 2413, 1587, 1680, 1759, 2347, 1946, 978, 1933, 2260, 665, 2650, 1632, 1843, 3066, 2486, 2440, 1769, 325, 326, 1619, 869, 2565, 3016, 1476, 2357]. two- [2632]. two-armed [1648]. two-factor [2137, 1476]. two-level [2633, 2308, 2137, 2448, 1469, 2357]. two-parameter [821, 1797, 1998, 675, 2018, 869]. two-phase [1821]. Two-piece [2879, 1699, 3048, 1933]. Two-sample [2021, 1989, 1758, 1853, 667, 1008, 2954, 1680, 2347, 978, 3066]. two-samples [2648]. two-sided [705, 998, 1769, 2565]. two-stage [2155, 1638, 1561, 1780]. two-state [924]. two-step [1754]. Two-Way [197, 1470, 2272, 1809, 1587, 325, 326]. Type [1671, 1736, 1158, 1980, 2469, 2553, 821, 2783, 2321, 2525, 2411, 287, 2552, 2700, 946, 1720, 1610, 2550, 1776, 1594, 2589, 1993, 1475, 2636, 1380, 1620, 1064, 2015, 2522, 1788, 2283, 2617, 1010, 776, 2094, 1375, 2372, 859, 2000, 966, 1261, 477, 1384, 854, 839, 2216, 1285, 1595, 2543, 2429, 2991, 1534, 410, 1853, 2073, 2346, 2201, 2081]. type-I [2321, 2411, 2073, 2201]. type-II [821, 946, 2522, 1853, 2346, 2081]. Typen [41, 403]. Tzanetakis [1890].

Übergangsratenmodelle [677].

Überlegungen [590]. Übersicht [698]. UdSSR [337, 163, 52]. Uhlenbeck [2950, 3062, 2001, 2388, 2544]. Ulrich [211]. Ultrahigh [1919, 2560, 2994, 2625, 2637, 1964]. ultrahigh-dimensional [2560]. ultrastructural [1641]. Umsatzsteuer [255]. UMVU [1220]. UMVUE [1088]. umweltstatistischen [288]. unabhängige [479]. Unabhängigkeitstests [616]. Unbalanced [1374, 1809, 2137, 2407]. unbekannte [506]. unbeobachtbaren [576]. unbestimmte [444]. Unbestimmtheiten [444]. Unbiased [1807, 694, 984, 2364, 1573, 1830, 3018, 851, 1132, 870, 2277, 723, 2215, 2481, 2283, 144, 1583, 2281, 497, 540, 2332, 1779, 1068, 1931, 1038, 2059, 2256, 2060]. Unbiasedly [1453]. unbiasedness [790, 909]. uncertain [1320, 2310, 970, 2867]. Uncertainty [2032, 2481, 1353, 2975, 490, 313, 1367, 1810, 2248]. unconditional [538]. under-dispersion [2132]. underdispersed [2667, 2634, 2821]. underdispersion [2542]. Understand [2555]. Understanding [2644, 2257, 2219, 1642]. undesirable [1265]. unemployment [1073]. unequal [2092, 891, 1008, 326, 763]. unexpected [2312]. Ungarischen [265]. ungarischer [337]. Ungarn [312, 232]. Ungarns [264]. Ungenauigkeiten [182]. Ungewißheit [113, 375, 29]. unification [1734]. Unified [2692, 2655, 2847, 1664, 2904, 2850, 1556, 866, 1968]. Uniform [2595, 1605, 1088, 2734, 722, 2572, 867, 2654, 2632, 1035, 2247, 1519, 1518, 1745, 2685]. Uniformity [2585, 2572, 919, 2784, 1237, 1420, 3046, 2166]. Uniformly [1830, 2859, 1959, 723, 3067, 1779, 2489, 2060]. unilateral [2906]. unimodal [849, 820, 1816, 1208]. unimodality [858, 2293]. union [2741, 1586]. uniqueness [857, 406]. Unit [828, 2382, 1331, 2626, 941, 1234, 1720, 1413, 1900, 1810, 1484, 1672, 829, 1262, 2048, 2074]. Unitarily [2126]. United [265]. units [1265, 1122]. univariate [2195, 3037, 1193, 3020, 1129, 2087, 2843, 2163, 2736, 2582]. Universität [6]. Universitäten [338]. University [2112, 2199, 2161]. unknown [2580, 2815, 852, 1008, 919, 1025, 2665, 850, 2504, 2852, 1876, 1262, 1036, 3042, 1433, 2788]. unobserved [2421]. unraveling [3000]. unrelated [1827, 1397, 2641, 2753, 1617]. unrestricted [1778]. unscharfe [504]. Unternehmensanalyse [138]. Unternehmensgewinne [183]. Unternehmenskonzentration [203]. Unternehmungen [176]. Untersuchungen [337]. Unverfälschte [674]. unvollständiger [579]. Unwin [1546, 2160]. upcrossings [2345]. Updatable [3021]. Updating [1843]. Upper [2260, 777, 2710, 1955, 1098, 1388]. urn [1018]. urns [1149]. Ursachen [156]. Ursprung [109]. USA [83]. usable [1168]. use [2814, 386, 171, 1510, 2981, 1198, 2493, 417, 451]. used [980, 379]. Useful [878, 1560, 2667, 982, 1458, 1199]. uses [720]. Using [1318, 1995, 1143, 2555, 1868, 1706, 2893, 2358, 1529, 1846, 2830, 2621, 3054, 1224, 1395, 1728, 1954, 771, 1901, 1936, 1245, 900, 941, 1462, 1496, 1132, 1674, 2562, 2474, 2222, 1533, 1367, 1593, 864, 2277, 2819, 1624, 1514, 766, 835, 1414, 1247, 2389, 2590, 3012, 1794, 2212, 2962, 1691, 1532, 2362, 1727, 1368, 2434, 2002, 2517, 3055, 3026, 1955, 2961, 1596, 1825, 2646, 956, 3070, 562, 965, 1586, 1251, 2070, 774, 1953, 1641, 1739, 1122, 1442, 1632, 1866, 2781, 2855, 2767, 1346, 1021, 1038, 2216, 807, 1443, 2163, 1119, 2139, 2236, 2942, 2342, 1839, 2841]. usual [1516]. Utility [1355, 1140, 1147, 330, 523]. Uwe [1567, 2108].

V [72, 1682, 265]. vague [1635, 1791]. validation [2830, 2568, 2460]. validity [727, 2865]. valuation [420]. value [2794, 287, 975, 2378, 245, 1032, 1604, 1791, 1309, 1386, 1106, 2227, 2901, 1725]. value-at-risk [2901]. valued [1428, 2045, 2406, 2993, 2682, 1479, 2420, 1877, 1825, 2475, 2254]. values [1233, 1365, 561, 615, 2535, 1103, 2013, 1245, 3064, 2829, 1462, 1837, 2222, 2635, 1096, 2817, 777, 2977, 2796, 1868, 1248, 1955, 753, 2961, 605, 1172, 1460, 1589, 1106, 1098, 1388]. VAR [2446, 1173]. variability [2934]. Variable [2356, 2708, 2804, 2647, 3051, 2449, 2443, 1929, 2065, 2868, 1395, 1132, 2913, 2832, 1734, 1967, 1799, 2246, 2185, 730, 2119, 2726, 2718, 2170, 2538, 2622, 2810, 2861, 2802, 1964, 1919, 2653, 1453, 773, 416, 1122, 2583, 1209, 1745, 2361, 2436, 2539, 1463, 2338, 2480, 2547]. Variablen [576, 217, 787, 96, 104]. Variablenprüfpläne [674]. Variables [2604, 1286, 1613, 2830, 561, 1572, 2184, 752, 2890, 1534, 2045, 2652, 1104, 1205, 1360, 1854, 1840, 2993, 1169, 2670, 740, 1881, 1995, 2005, 1457, 2499, 1989, 923, 1883, 1557, 2844, 730, 929, 1440, 1612, 1187, 1215, 2659, 868, 584, 2022, 977, 1109, 1973, 675, 2136, 1032, 1510, 2981, 751, 2702, 2984, 1344, 880, 1542, 917, 1516, 1701, 2126, 1519, 2413, 411, 1518, 1251, 1497, 2180, 2340, 2610, 1668, 825, 908, 1700, 1757, 1884, 2052, 560, 1744, 1769, 2444, 2600, 3043, 1792, 2450, 2405, 2602, 2736, 2468, 2357, 1636, 2453]. Variance [2008, 858, 390, 3018, 1574, 1461, 1007, 1900, 1293, 2819, 2815, 723, 2220, 1735, 1649, 2296, 2955, 1087, 1414, 1474, 881, 1539, 977, 1678, 2215, 2481, 144, 2136, 1583, 2281, 1887, 1198, 540, 1249, 859, 1011, 1516, 1000, 2036, 2524, 1194, 1656, 1729, 1779, 393, 1122, 1866, 1359, 763, 1549, 1896, 2866, 2227, 1818, 1622, 2637, 949]. Variance-mean [2008]. variances [669, 891, 2590, 880]. Varianzen [460]. variate [822, 2921]. variates [2663, 665]. Variation [2939, 2226, 2463, 422, 530, 2244, 999, 1678, 3011, 1297, 3010]. variation-diminishing [3011]. variety [1574, 2155]. variogram [2267]. variogram-based [2267]. various [1055, 2572, 1602, 2358, 3016]. VARMA [2505, 2097]. VARs [2890]. Varying [2871, 2103, 2279, 1954, 3041, 2185, 2245, 1206, 2333, 3060, 2136, 2889, 2810, 2917, 1801, 2049, 2223, 2507, 2963, 2978, 2064, 2361, 2436, 2476, 2631, 2872, 2065, 3007]. varying-coefficient [2279, 2245, 2136, 2963, 2064, 2631, 2872]. Vas [265]. Vasicek [2867]. Vatolkin [2217]. vector [1428, 2193, 2799, 851, 3037, 1414, 712, 1553, 1805, 2266, 3032, 2084, 2680]. vector-autoregressive [1414]. vectorautoregressive [981]. vectorial [2455]. vectors [2876, 1852, 1639, 2754, 3006, 1134, 1371, 2650, 2949]. verallgemeinerten [456]. Verallgemeinerung [616, 444, 210, 443, 180]. Verantwortlichkeit [50]. Verarbeitenden [480]. verbesserten [280]. Verbesserung [385]. Verbrauchsniveau [93]. verbundene [649]. Verdooren [2432]. vereinfachte [100]. vereinfachten [97]. vereinfachtes [412]. Vereinfachung [463]. Vereinigten [249]. Verfahren [658, 533, 516, 1503, 345]. Verflechtungen [143]. Vergangenheit [25]. Vergleich [105, 190, 9, 252, 272, 34, 62, 654, 370, 460, 83, 290, 55, 345, 312, 173, 277]. Vergleichbarkeit [225]. vergleichenden [11, 28]. Vergleichs [231, 99]. Verhalten [181]. Verhaltens [204]. Verhaltensweisen [219]. Verhältniszahlen [403]. Verkehrsproblemen [480]. Verlag [349, 265]. verschiedenartige [53]. verschiedenen [372]. verschiedener [41, 243]. version [1702, 895, 1402, 2316, 1938, 2946]. versions [1025, 1780]. Versuch [3]. Versuche [337, 432, 431, 377]. versus [1897, 2508, 389, 1597, 621, 1282, 1460, 698]. Verteilung [631, 368, 356, 259]. verteilungsfreie [622]. Verteilungsfunktionen [372]. vertex [2589]. vertices [1611]. Verwendung [44, 171, 372]. very [316]. verzerrte [529]. Verzerrungen [156]. verzögerten [642]. via [838, 1992, 1067, 2118, 2365, 1664, 2513, 2741, 2027, 1862, 2958, 2229, 814, 2380, 1982, 2804, 2448, 2394, 3006, 3051, 2400, 1616, 3056, 2713, 2138, 2496, 2934, 3029, 2729, 2202, 2887, 2401, 2910, 3036, 2877, 2625]. Vicente [1870]. video [2735]. vier [537]. view [1440, 824]. Views [770, 1589]. VIF [2181]. Vining [1660]. violations [2924]. Visualisation [2239]. Visualizing [1546]. vitality [988]. Voeller [810]. Vol [787, 799]. volatilities [1084]. Volatility [1033, 1084, 1533, 2249, 2389, 3038, 2591]. Völkerbundes [3]. Völkerbundszeit [4]. Volkseinkommens [62]. Volksrepublik [143, 200]. Volkswirtschaft [246, 258]. Volkswirtschaften [243]. Volkswirtschaftliche [130]. Volkswirtschaftlichen [281, 268, 89]. vollständige [621]. Volume [2369, 2484, 98]. volumes [2127]. Vonesh [2002]. vor [50]. Voraussetzung [97]. Vorläufern [3]. Vorlesung [413]. Vorratsänderungen [160]. Vorschläge [320]. Vorwort [515, 46]. vs [2061, 2751].

W [1707, 2146, 971, 1703, 2674, 1665, 1748, 1724]. W. [672]. Wachstums [246]. Wachstums- [246]. Wachstumsvergleiche [243]. Wagenführ [364]. Wahrscheinlichkeit [506, 48]. Wahrscheinlichkeiten [444]. Wahrscheinlichkeitsinterpretationen [42]. Wahrscheinlichkeitssubjektivismus [437]. wahrscheinlichkeitstheoretische [508, 442]. Wahrscheinlichkeitstransformationen [651]. Wahrscheinscheinlichkeiten [601]. waiting [2775, 901]. Wald [18, 735]. walk [2204, 2080]. walks [2644]. Walsh [323]. Walter [381, 1665, 1503]. Walters [1526]. Wang [1685]. Waring [666]. Warner [1674, 1693]. warning [1377]. Watson [1470, 298]. Watson-Test [298]. Wavelet [1979, 2498, 1766, 2036, 2549]. wavelet-based [2036]. Way [197, 2272, 1927, 1354, 1470, 1809, 1155, 1587, 855, 324, 325, 326, 1222, 1670]. Wayne [2218]. Weak [2097, 1718, 2505, 1419, 796, 1398, 2859]. weakly [2648, 2579, 2149]. wealth [2051]. web [2262]. web-based [2262]. Weibull [2722, 2081, 1260, 3004, 226, 639, 1615, 1856, 2233, 597, 2960, 1514, 1369, 1348, 1509, 1657, 2073, 2962, 2416, 2806, 1651, 1733, 1671, 1692, 2627, 489, 1466, 1888, 1841, 1325, 1953, 477, 1481, 2201, 1843, 340, 743, 1388, 1715]. Weibullized [1534]. Weiche [360]. weicher [532]. weighing [2098]. weight [2907]. Weighted [2745, 2339, 1037, 1689, 2364, 1742, 2918, 1447, 2652, 2832, 1246, 2375, 2550, 2570, 2232, 2695, 2185, 3015, 2037, 2748, 1883, 2285, 1696, 2422, 2651, 2122, 2946, 1121, 2263, 2874, 2860, 785, 1931, 1700, 1757, 1884, 1698, 2543, 3043, 2602, 2062, 2678, 1697]. weighting [2794, 897, 1275, 1300]. Weihs [2217]. Weiss [1524, 1292]. Weißbach [2105]. Welch [2924]. Welfe [2019]. Welt [84, 164, 135, 170]. Weltbevölkerung [106]. Weltkrieges [5]. Weltnahrungsmittelproduktion [106]. Weltproduktion [66]. Weltsachgüterproduktion [63]. Wendy [1962]. Wertes [210]. Wertindextreue [463]. Wertrechnung [52]. Wertschöpfung [9]. West [1869]. Westeuropa [139]. westeuropäischer [62]. where [1564]. whether [980]. White [1592, 2154]. Whitney [3018]. whose [416]. wichtigsten [281, 163]. widely [2468, 2901]. wider [2309]. Widerspruch [378]. Wiegert [912]. Wielandt [1795]. Wien [349]. Wiener [2461, 715]. Wigner [2330]. Wiilekens [1686]. Wikle [2556]. Wilcoxon [309, 2362, 292, 1864]. Wild [2193, 1901, 2951]. Wiley [1504]. Wilfried [246]. Wilks [2966]. Willi [293]. windows [954]. Winkler [1703]. Winsorization [328]. winsorized [2778]. Wintgen [273]. Wirtschaft [533]. wirtschaftlicher [261]. Wirtschafts [388, 833, 603, 179]. Wirtschafts- [833, 603, 179]. Wirtschafts-und [388]. Wirtschaftsakademien [117]. Wirtschaftsfachschulen [117]. Wirtschaftsforschung [28]. Wirtschaftsgemeinschaft [93]. Wirtschaftsgruppen [34]. Wirtschaftshilfe [54]. Wirtschaftsintegration [112]. wirtschaftspolitischen [272]. Wirtschaftsstatistik [61, 127, 80]. Wirtschaftsstatistiken [281]. wirtschaftsstatistischen [182, 190, 440]. Wirtschaftsstrukturen [290]. wirtschaftstheoretischen [273]. Wirtschaftswissenschaftler [2105]. wirtschaftswissenschaftlichen [240]. Wirtschaftszweigen [220]. wise [2863, 2417]. Wishart [982]. Wissenschaften [265]. wissenschaftlicher [377, 432]. Wissenschaftstheorie [568]. within [1140, 490, 2487, 1386]. without [1448, 1809, 1324, 2281, 2173, 797]. Wladyslaw [2019]. WOD [2510]. Wohlfahrt [108]. Wohlstands [89]. Wohnbaukosten [55]. Wojtek [1746]. Wolfgang [464]. Woodward [2218]. work [1754, 33]. worked [2433]. Working [2352]. World [591, 632, 2924]. wrap [2363]. wrap-around [2363]. Wrocław [417, 451]. Wrocław-Taxonomy [417, 451]. Wu [2088]. Würzburg [349]. Würzburg-Wien [349].

X [1380]. xii [2219, 1736, 2321, 1475]. xiv [2147]. xvii [2198]. xxi [2259]. xxvii [2019].

Yadohisa [1681]. Yang [805]. Yates [965]. York [2182]. Youngjo [1664]. Yudi [1664]. Yule [300, 352].

Zadehs [504]. Zahlen [388, 83]. Zammit [2556]. Zammit-Mangion [2556]. ZDF [413, 413]. Zeilen [579]. Zeilen- [579]. Zeit [564, 518]. Zeit- [564, 518]. Zeitkontinuierliche [833]. Zeitreihe [297, 509]. Zeitreihen [636, 442, 613, 35]. Zeitreihenanalyse [161]. zeitreihendaten [698]. Zeitreihenmodelle [698]. Zeitreihenmodellierung [1567]. Zeitreihensätzen [233]. Zerlegung [613]. Zero [2548, 2870, 2490, 2858, 2996, 1939, 2132, 2323]. zero-adjusted [2858, 2996]. Zero-and-one-inflated [2548]. zero-inflated [2870, 2490]. zeros [2173]. Zimmerman [1870]. Zoltán [265]. zonoid [2803]. Zschocke [349]. Zuccolotto [2675]. Zufallsstichproben [678]. Zum [413, 165, 15, 235, 344, 388, 298, 46, 167, 291, 364, 366, 190, 1, 2, 109, 140, 292, 309, 225]. Zur [204, 658, 365, 596, 9, 514, 521, 425, 171, 563, 100, 319, 467, 673, 3, 515, 833, 320, 499, 437, 452, 210, 451, 28, 73, 533, 131, 610, 121, 236, 356, 103, 272, 456, 601, 678, 220, 394, 207, 203, 506, 183, 432, 476, 241, 431, 195, 426, 348, 98, 91, 127, 198, 555]. Zurückweisung [432]. Zuverlässigkeit [20]. Zwei [649, 53, 192, 485, 527, 537, 53]. zweidimensionalen [491, 296]. zweier [252, 460]. zweiseitigen [674, 296]. zweiten [5]. zweiter [440]. Zwer [440]. zwischen [217, 269, 83, 372]. Zwischenbericht [61].

References

M:1960:GAP

[1] G. M. Zum Geleit, Avant-propos, Editorial. Statistical Papers, 1(1):1–7, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF02922251.

Menges:1960:BIS

[2] Günter Menges. Zum Begriff der internationalen Statistik. (German) [The concept of international statistics]. Statistical Papers, 1(1):8–21, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF02922252.

Menges:1960:VGI

[3] Günter Menges. Versuch einer Geschichte der internationalen Statistik von ihren Vorläufern im Altertum bis zur Entstehung des Völkerbundes. (German) [Attempting a history of international statistics from its predecessors in antiquity to the formation of the League of Nations]. Statistical Papers, 1(1):22–64, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922253.

Bott:1960:SIO

[4] Dietrich Bott. Die Statistik in den internationalen Organisationen der Völkerbundszeit. (German) []. Statistical Papers, 1(1):65–103, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922254.

Gossmann:1960:SIO

[5] Jens Goßmann. Die Statistik in den internationalen Organisationen seit dem Ende des zweiten Weltkrieges. (German) []. Statistical Papers, 1(1): 104–163, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922255.

Anonymous:1960:BGI

[6] Anonymous. Bericht von der Gründung des Instituts für europäische Statistik an der Universität des Saarlandes, Saarbrücken. (German) []. Statistical Papers, 1(1):164–171, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922256.

Anonymous:1960:HC

[7] Anonymous. Help & contacts. Statistical Papers, 1(1):??, December 1960. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Brendow:1961:IBG

[8] Klaus Brendow. Der internationale Budgetvergleich. (German) []. Statistical Papers, 2(1):1–15, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922264.

Gehrecke:1961:IVM

[9] Siegfried Gehrecke. Internationaler Vergleich der Methoden zur Berechnung der Wertschöpfung in den Dienstleistungsbereichen. (German) []. Statistical Papers, 2(1):16–32, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922265.

Sange:1961:MPS

[10] Norbert Sange. Methoden und Probleme der Sozialstatistik in internationaler Sicht. (German) []. Statistical Papers, 2(1):33–92, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922266.

Schikowski:1961:EPI

[11] Wolfgang Schikowski. Einige Probleme der international vergleichenden Produktivitätsmessung. (German) []. Statistical Papers, 2(1):93–116, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922267.

Houssiaux:1961:CSP

[12] Jacques Houssiaux. La comparaison des structures de production et d’échange dans les pays de la Communauté économique européenne. (French) [the comparison of production and exchange structures in the countries of the European Economic Community]. Statistical Papers, 2(1): 117–169, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922268.

Rutsch:1961:MIO

[13] Martin Rutsch. Multiregionale input– output-modelle. Statistical Papers, 2 (1):171–184, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922269.

Anonymous:1961:HC

[14] Anonymous. Help & contacts. Statistical Papers, 2(1):??, December 1961. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bott:1962:AHB

[15] Dietrich Bott. Allgemeine und historische Betrachtungen zum Entscheidungsbegriff. (German) []. Statistical Papers, 3(1):1–38, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927680.

Menges:1962:BRS

[16] Günter Menges and Minaketan Behara. Das Bayes’sche Risiko bei sequentiellen Stichprobenentscheidungen. (German) []. Statistical Papers, 3(1):39–61, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927681.

Schneeweiss:1962:EEE

[17] Hans Schneeweiß. Einige Experimente mit Entscheidungsspielen. (German) []. Statistical Papers, 3(1):62–78, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927682.

Zschocke:1962:EAW

[18] Dietrich Zschocke. Das Entscheidungsproblem bei Abraham Wald und Richard Bellman. (German) []. Statistical Papers, 3(1):79–98, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927683.

Menges:1962:SOR

[19] Günter Menges. Statistik und Operations Research. (German) []. Statistical Papers, 3(1):99–106, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927684.

Baumgarten:1962:ZIG

[20] Erwin Baumgarten. Zuverlässigkeit und Instandhaltung. (German) []. Statistical Papers, 3(1):107–123, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927685.

Stoller:1962:ASC

[21] David S. Stoller. An application of the single-channel queue. Statistical Papers, 3(1):124–130, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927686.

Schneeweiss:1962:ASS

[22] Hans Schneeweiß. Ein allgemeines Schema des stochastischen Programmierens. (German) []. Statistical Papers, 3(1):131–157, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927687.

Anonymous:1962:HC

[23] Anonymous. Help & contacts. Statistical Papers, 3(1):??, December 1962. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Menges:1963:TEE

[24] Günter Menges. Three essays in econometrics. Statistical Papers, 4 (1):1–37, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923041.

Zwer:1963:SKV

[25] Reiner Zwer. Die statistische Konjunkturforschung in Vergangenheit und Gegenwart. (German) [Statistical economic research in the past and present]. Statistical Papers, 4(1):38–79, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923042.

Schulmann:1963:KGS

[26] Horst Schulmann. Von kleinen und großen Schwierigkeiten des ökonometrischen Modellbauers. (German) []. Statistical Papers, 4(1):80–114, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923043.

Narayanan:1963:SAD

[27] R. Narayanan. A statistical analysis of demand for coal in India. Statistical Papers, 4(1):115–134, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923044.

Weis:1963:BIO

[28] Hans Weis. Beitragsmöglichkeiten der Input–Output-Analyse zur international vergleichenden Wirtschaftsforschung. (German) []. Statistical Papers, 4(1):135–150, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923045.

Menges:1963:KOE

[29] Günter Menges. Kriterien optimaler Entscheidungen unter Ungewißheit. (German) []. Statistical Papers, 4(1): 151–171, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923046.

Kale:1963:DOM

[30] B. K. Kale. Decisions opposite Markov chains. Statistical Papers, 4 (1):172–177, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923047.

Schneeweiss:1963:NTM

[31] Hans Schneeweiß. Nutzenaxiomatik und Theorie des Messens. (German) []. Statistical Papers, 4(1):178–220, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923048.

Anonymous:1963:HC

[32] Anonymous. Help & contacts. Statistical Papers, 4(1):??, December 1963. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Menges:1964:IIO

[33] Günter Menges. Integration and inputoutput work in the European Economic Community. Statistical Papers, 5(1–2):5–18, December 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922817.

Helmstadter:1964:VHW

[34] Ernst Helmstädter. Ein Vergleich der Hierarchie der Wirtschaftsgruppen in den EWG-Ländern. (German) []. Statistical Papers, 5(1–2):19–37, December 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922818.

Rutsch:1964:ATM

[35] Martin Rutsch. Die Akkomodation technologischer Matrizen mit Hilfe kurzer Zeitreihen. (German) []. Statistical Papers, 5(1–2):39–49, December 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922819.

Tintner:1964:LPI

[36] Gerhard Tintner. Lineare Programme und Input–Output-Analyse. (German) []. Statistical Papers, 5(1–2):50–55, December 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922820.

Rutsch:1964:OCM

[37] Martin Rutsch. Optimal cooperation in multiregional input-output systems. Statistical Papers, 5(1–2):56–70, December 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922821.

Anonymous:1964:HCa

[38] Anonymous. Help & contacts. Statistical Papers, 5(1–2):??, December 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Schneeweiss:1965:AG

[39] Hans Schneeweiß. Das Aggregationsproblem. (German) []. Statistical Papers, 6(1):1–26, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922284.

Menges:1965:SOG

[40] Günter Menges and Helmut Diehl. Das Stabilitätsproblem in der Ökonometrie.

(German) []. Statistical Papers, 6 (1):27–52, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922285.

Anonymous:1965:EVT

[41] Anonymous. Die Eignung verschiedener Typen industriestatistischer Erhebungseinheiten für die Input– Output-Analyse. (German) []. Statistical Papers, 6(1):53–79, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922286.

Menges:1965:WGI

[42] Günter Menges. Über Wahrscheinlichkeitsinterpretationen. (German) [On the interpretation of probability]. Statistical Papers, 6(1):81–96, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922287.

Sprott:1965:SES

[43] David A. Sprott. Statistical estimation — some approaches and controversies. Statistical Papers, 6(1):97–111, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922288.

Diehl:1965:LIV

[44] Helmut Diehl and David A. Sprott. Die Likelihoodfunktion und ihre Verwendung beim statistischen Schluß. (German) []. Statistical Papers, 6(1): 112–134, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922289.

Anonymous:1965:HC

[45] Anonymous. Help & contacts. Statistical Papers, 6(1):??, December 1965. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Menges:1966:VHJ

[46] G. Menges and R. Wagenführ. Vorwort der Herausgeber zum 7. Jahrgang. (German) []. Statistical Papers, 7(1– 2):1, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02922841.

Anonymous:1966:DAa

[47] Anonymous. Zu dieser ausgabe. Statistical Papers, 7(1–2):2, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922842.

Dagum:1966:WAT

[48] Camilo Dagum. Wahrscheinlichkeit und Ausmaß der Transvariation im n-dimensionalen Raum. (German) []. Statistical Papers, 7(1–2):3–29, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922843.

Menges:1966:OEE

[49] G. Menges and H. Diehl. Über die operationelle Eignung von Entscheidungsmodellen. (German) []. Statistical Papers, 7(1–2):30–41, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922844.

Sauvy:1966:VSV

[50] Alfred Sauvy. Die Verantwortlichkeit des Statistikers vor der öffentlichen Meinung und den öffentlichen Gewalten. (German) []. Statistical Papers, 7(1–2):42–50, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922845.

Erfahrungsbericht:1966:SIS

[51] Ein Erfahrungsbericht and Rolf Wagenführ. Die Statistik in der Integration der Sechs. (German) []. Statistical Papers, 7(1–2):51–73, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922846.

Frenzel:1966:WPS

[52] Gottfried Frenzel. Wertrechnung und Preisrechnung in der sozialistischen Planung der UdSSR. (German) []. Statistical Papers, 7(1–2):74–112, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922847.

Wagenfuhr:1966:ZNS

[53] Rolf Wagenführ. Zwei nationale statistische Jahrbücher-zwei verschiedenartige statistische systeme. Statistical Papers, 7(1–2):113–116, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922848.

Wagenfuhr:1966:IIB

[54] Rolf Wagenführ. Indexiffern der industriellen Brutto- und Nettoproduktion für einige Länder des Rates für gegenseitige Wirtschaftshilfe. (German) []. Statistical Papers, 7(1– 2):117–119, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922849.

Wagenfuhr:1966:IVW

[55] Rolf Wagenführ. Internationaler Vergleich der Wohnbaukosten. (German) []. Statistical Papers, 7(1–2):120–122, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922850.

Anonymous:1966:HCa

[56] Anonymous. Help & contacts. Statistical Papers, 7(1–2):??, March 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1966:DAb

[57] Anonymous. Zu dieser ausgabe. Statistical Papers, 7(3–4):123–124, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922951.

Schneeweiss:1966:GEG

[58] Hans Schneeweiß. Das Grundmodell der Entscheidungstheorie. (German) []. Statistical Papers, 7(3–4): 125–137, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922952.

Henn:1966:MKL

[59] Rudolf Henn. Markoffsche Ketten bei der Lagerhaltung. (German) []. Statistical Papers, 7(3–4):138–147, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922953.

Zahlen:1966:GTP

[60] J. P. Zahlen. Über die Grundlagen der Theorie der parametrischen Hypothesentests. (German) []. Statistical Papers, 7(3–4):148–174, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922954.

Wagenfuhr:1966:SWZ

[61] Rolf Wagenführ and Siegfried Maaß. Der Systemgedanke in der Wirtschaftsstatistik — Ein Zwischenbericht. (German) []. Statistical Papers, 7(3–4): 175–211, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922955.

Kseckovska:1966:VVP

[62] E. Kseckovska. Über den Vergleich des Volkseinkommens Polens und einiger westeuropäischer Länder. (German) []. Statistical Papers, 7(3–4): 212–221, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922956.

Kuczynski:1966:IWB

[63] Jürgen Kuczynski. Ein Index der Weltsachgüterproduktion, 1850 bis 1965. (German) []. Statistical Papers, 7(3– 4):222–237, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922957.

Schulz:1966:KAH

[64] Udo Schulz. Die Koordination der amtlichen Hochschulstatistik mit den Geschäftsstatistiken der Hochschulen. (German) []. Statistical Papers, 7(3– 4):238–251, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922958.

Wagenfuhr:1966:AGI

[65] Rolf Wagenführ. Die Auswirkungen von Gebietsveränderungen auf die industriellen Produktionsindizes von kapitalistischen und sozialistischen Ländern. (German) []. Statistical Papers, 7(3–4):252–253, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922959.

Wagenfuhr:1966:IWI

[66] Rolf Wagenführ. Die industrielle Weltproduktion im Jahre 1965. (German) []. Statistical Papers, 7(3–4): 253–254, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02922960.

Anonymous:1966:HCb

[67] Anonymous. Help & contacts. Statistical Papers, 7(3–4):??, September 1966. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1967:DAa

[68] Anonymous. Zu dieser ausgabe. Statistical Papers, 8(1):1, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922859.

Ferschl:1967:OAS

[69] F. Ferschl. Über eine Optimierungsaufgabe in Assemblage–Systemen. (German) []. Statistical Papers, 8(1):2–17, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922860.

Menges:1967:OPa

[70] G. Menges. Ökonometrische prognosen. Statistical Papers, 8(1):18–31, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922861.

vonWeizsacker:1967:PPQ

[71] C. C. von Weizsäcker and H. L. Freytag. Plan eines Projektes: “Ein quantitatives Modell des Bildungswesens in der Bundesrepublik Deutschland”. (German) [Plan of a project:

“A quantitative model of education in the Federal Republic of Germany”]. Statistical Papers, 8(1):32–49, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922862.

Frenzel:1967:RVS

[72] G. Frenzel. Referat über V. Simcer: Besonderheiten der Berechnung des Niveaus der industriellen Produktion in den sozialistischen Ländern und Klassifizierung und Kenngrößen der Industriestruktur. (German) []. Statistical Papers, 8(1):50–65, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922863.

Anonymous:1967:MSA

[73] Anonymous. Zur Methodik der statistischen Aggregation. (German) []. Statistical Papers, 8(1):66–78, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922864.

Wagenfuhr:1967:AAG

[74] R. Wagenführ. Arbeitsmittel und Arbeitsgegenstände. (German) []. Statistical Papers, 8(1):79–81, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922865.

Lauer:1967:DIS

[75] B. Lauer. Direkter und indirekter Stahlexport der EWG-Länder. (German) []. Statistical Papers, 8(1):82–84, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922866.

Anonymous:1967:HCa

[76] Anonymous. Help & contacts. Statistical Papers, 8(1):??, March 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1967:DAb

[77] Anonymous. Zu dieser ausgabe. Statistical Papers, 8(2):85–86, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923491.

Menges:1967:OPb

[78] G. Menges. Ökonometrische prognosen. Statistical Papers, 8(2):87–98, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923492.

Kalbfleisch:1967:FP

[79] J. G. Kalbfleisch and D. A. Sprott. Fiducial probability. Statistical Papers, 8(2):99–109, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923493.

Wagenfuhr:1967:NSS

[80] R. Wagenführ and W. Kunhardt. Nomenklaturen, Systematiken, statistische Einheiten — Einige Probleme der Systematisierung der Wirtschaftsstatistik. (German) []. Statistical Papers, 8(2):110–128, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923494.

Rajcin:1967:BLB

[81] V. Ja. Rajcin and E. M. Poberezskaja. Berechnungsmethodik für Lebensniveauunterschiede unter Berücksichtigung von Geschlechts- und Altersstruktur der Familien. (German) []. Statistical Papers, 8(2):129–138, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923495.

Hampe:1967:B

[82] A. Hampe. Buchbesprechung. Statistical Papers, 8(2):139–144, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923496.

Tiede:1967:ZAV

[83] M. Tiede. Zahlen der Arbeitskräfte Ein Vergleich zwischen EWG, EFTA, RGW und USA. (German) []. Statistical Papers, 8(2):145–150, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923497.

Voss:1967:SWI

[84] W. Voß. Die Staaten der Welt und ihre Einteilung nach Intensitätsstufen. (German) []. Statistical Papers, 8 (2):151–162, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923498.

Anonymous:1967:HCb

[85] Anonymous. Help & contacts. Statistical Papers, 8(2):??, June 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1967:DAc

[86] Anonymous. Zu dieser ausgabe. Statistical Papers, 8(3):163–164, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922867.

Kale:1967:TGF

[87] B. K. Kale and V. P. Godambe. A test of goodness of fit. Statistical Papers, 8(3):165–172, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922868.

Menges:1967:EAD

[88] G. Menges and H. Diehl. Entwicklung eines allgemeinen dynamischen Entscheidungsmodells. (German) []. Statistical Papers, 8(3):173– 182, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922869.

Tinbergen:1967:GPT

[89] J. Tinbergen. Gegenwärtige Probleme der Theorie des volkswirtschaftlichen Wohlstands. (German) []. Statistical Papers, 8(3):183–192, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922870.

Brendow:1967:GK

[90] K. Brendow. Gesamteuropäische kohlenstatistik. Statistical Papers, 8 (3):193–209, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922871.

Wagenfuhr:1967:SEE

[91] R. Wagenführ. Zur Struktur der Einzelhandelspreise in den EWG-Ländern. (German) []. Statistical Papers, 8(3): 210–215, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922872.

Stalf:1967:RAS

[92] E. Stalf. Rechtsgrundlagen der amtlichen Statistik in beiden Teilen Deutschlands. (German) []. Statistical Papers, 8(3):216–222, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922873.

Tiede:1967:VAE

[93] M. Tiede. Das Verbrauchsniveau der Arbeiter in der Europäischen Wirtschaftsgemeinschaft 1963/64. (German) []. Statistical Papers, 8(3): 223–225, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922874.

Anonymous:1967:HCc

[94] Anonymous. Help & contacts. Statistical Papers, 8(3):??, September 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1967:DAd

[95] Anonymous. Zu dieser ausgabe. Statistical Papers, 8(4):227, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922577.

Weber:1967:BBM

[96] E. Weber. Biometrische Bearbeitung multipler Regressionen unter besonderer Berücksichtigung der Auswahl, der Transformation und der Linearkombination von Variablen. (German) []. Statistical Papers, 8(4):228–251, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922578.

Kozak:1967:DMV

[97] J. Kozák. Drei Möglichkeiten der vereinfachten Prognoserechnung unter der Voraussetzung der Translationsinvarianz. (German) []. Statistical Papers, 8(4):252–262, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922579.

Wagenfuhr:1967:EIS

[98] R. Wagenführ. Zur Entwicklung des Investitionsvolumens in den sozialistischen Ländern Europas. (German) [On the development of investment volume in the socialist countries of Europe]. Statistical Papers, 8(4):263–278, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922580.

Dieckmann:1967:LTV

[99] B. Dieckmann and D. Berstecher. Logik und Technik des Vergleichs. (German) []. Statistical Papers, 8(4): 279–298, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922581.

Koch:1967:VMI

[100] W. Koch. Eine vereinfachte Methode zur indirekten Messung der Investitionstätigkeit. (German) []. Statistical Papers, 8(4):299–304, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922582.

Anonymous:1967:HCd

[101] Anonymous. Help & contacts. Statistical Papers, 8(4):??, December 1967. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1968:DAa

[102] Anonymous. Zu dieser ausgabe. Statistical Papers, 9(1):1–2, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922884.

Haq:1968:PHN

[103] M. Safiul Haq. Zur Prognose aus einer homoskedastischen Normalverteilung mit Hilfe der Strukturwahrscheinlichkeit. (German) []. Statistical Papers, 9(1):3–12, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922885.

Weber:1968:BBM

[104] E. Weber. Biometrische Bearbeitung multipler Regressionen unter besonderer Berücksichtigung der Auswahl, der Transformation und der Linearkombination von Variablen. (German) []. Statistical Papers, 9(1):13–33, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922886.

Drechsler:1968:IVI

[105] Lászlo Drechsler. Der internationale Vergleich der Industrieproduktion. (German) []. Statistical Papers, 9 (1):34–50, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922887.

Kuczynski:1968:WWG

[106] Jürgen Kuczynski. Weltbevölkerung und Weltnahrungsmittelproduktion. (German) []. Statistical Papers, 9 (1):51–57, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922888.

Wagenfuhr:1968:LPE

[107] Rolf Wagenführ. Die landwirtschaftliche Produktion der europäischen RGW-Länder. (German) []. Statistical Papers, 9(1):58–59, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02922889.

Voss:1968:PEW

[108] W. Voß. Personelle Einkommensverteilung und Wohlfahrt. (German) []. Statistical Papers, 9(1):60–65, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922890.

Menges:1968:UOR

[109] Günter Menges. Zum Ursprung von Operational Research. (German) []. Statistical Papers, 9(1):66, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922891.

Anonymous:1968:HCa

[110] Anonymous. Help & contacts. Statistical Papers, 9(1):??, March 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1968:DAb

[111] Anonymous. Zu dieser ausgabe. Statistical Papers, 9(2):69–70, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923540.

Wagenfuhr:1968:BPW

[112] Rolf Wagenführ and Gottfried Frenzel. Bestandsaufnahme und Perspektiven der Wirtschaftsintegration im RGW– Raum. (German) []. Statistical Papers, 9(2):71–81, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923541.

Bernard:1968:OSU

[113] Georges Bernard. Optimale Strategien unter Ungewißheit. (German) []. Statistical Papers, 9(2):82–100, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923542.

Enzer:1968:OF

[114] Hermann Enzer. The ordinalist fallacy. Statistical Papers, 9(2):101–107, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923543.

Stowe:1968:SAL

[115] Heinz Stöwe and Rolf Rodiek. Strukturwandlung in der Altersverteilung der Lungenkrebssterbefälle. (German) []. Statistical Papers, 9(2):108–132, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923544.

Wagenfuhr:1968:SRL

[116] Rolf Wagenführ and Gottfried Frenzel. Die Statistik der RGW-Länder. (German) []. Statistical Papers, 9 (2):133–162, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923545.

Huttner:1968:SHW

[117] Manfred Hüttner. “Statistik” an Höheren Wirtschaftsfachschulen bzw. Wirtschaftsakademien. (German) []. Statistical Papers, 9(2):163–165, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923546.

Anonymous:1968:HCb

[118] Anonymous. Help & contacts. Statistical Papers, 9(2):??, June 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1968:DAc

[119] Anonymous. Zu dieser ausgabe. Statistical Papers, 9(3):167–168, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923649.

Wagenfuhr:1968:PGI

[120] Rolf Wagenführ. Das produzierende gewerbe im EWG–Raum. Statistical Papers, 9(3):169–175, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923650.

Diehl:1968:OIS

[121] Helmut Diehl and Stephanus L. Louwes. Zur Optimierung von Informationsprogrammen bei statistischen Entscheidungen. (German) []. Statistical Papers, 9(3):176–188, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923651.

Precht:1968:LSD

[122] Manfred Precht. Über lineare stochastische Differentialgleichungen. (German) [On linear stochastic differential equations]. Statistical Papers, 9 (3):189–209, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923652.

Maass:1968:SBS

[123] Siegfried Maaß. Die statistische Behandlung des Staatssektors in Input– Output-Tabellen. (German) []. Statistical Papers, 9(3):210–234, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923653.

Petrova:1968:IOF

[124] V. Petrova. Über Input–Output-Forschung in den europäischen sozialistischen Ländern. (German) []. Statistical Papers, 9(3):235–241, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923654.

Anonymous:1968:HCc

[125] Anonymous. Help & contacts. Statistical Papers, 9(3):??, September 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1968:DAd

[126] Anonymous. Zu dieser ausgabe. Statistical Papers, 9(4):243, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02927703.

Wagenfuhr:1968:SWA

[127] Rolf Wagenführ. Zur Systematisierung der Wirtschaftsstatistik — Antwort an Gerhard Fürst. (German) []. Statistical Papers, 9(4):244–254, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927704.

Rosenbluth:1968:IOA

[128] Gideon Rosenbluth. Input-output analysis: A critique. Statistical Papers, 9 (4):255–268, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927705.

Gebert:1968:PSK

[129] James R. Gebert. A power study of Kimball’s statistics. Statistical Papers, 9(4):269–273, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927706.

Wagenfuhr:1968:FVG

[130] Rolf Wagenführ and Erich Stalf. Finanzierungsrechnung und Volkswirtschaftliche Gesamtrechnung — gezeigt am Beispiel der Bundesrepublik Deutschland. (German) [Financial accounts and national accounts — demonstrated by the example of the Federal Republic of Germany]. Statistical Papers, 9(4):274–295, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927707.

Bender:1968:ABL

[131] Ulrich Bender. Zur Abgrenzung von Bundes-, Landes- und Gemeindestatistiken. (German) []. Statistical Papers, 9(4):296–300, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927708.

Tiede:1968:B

[132] Manfred Tiede, Werner Voß, and Gert Elstermann. Buchbesprechung. Statistical Papers, 9(4):301–306, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927709.

Anonymous:1968:HCd

[133] Anonymous. Help & contacts. Statistical Papers, 9(4):??, December 1968. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1969:DAa

[134] Anonymous. Zu dieser ausgabe. Statistical Papers, 10(1):1, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02927740.

Wagenfuhr:1969:JSG

[135] Rolf Wagenführ. Japans Stellung in der gewerblichen Sachgütererzeugung in der Welt. (German) []. Statistical Papers, 10(1):2–4, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927741.

Feichtinger:1969:SAG

[136] Gustav Feichtinger. Stochastische Automaten als Grundlage linearer Lernmodelle. (German) []. Statistical Papers, 10(1):5–21, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927742.

Hansen:1969:PLF

[137] Gerd Hansen. Die Prognoseeignung von LISE- und FIND-Maximum-Likelihood–Schätzwerten. (German) [The prognosis suitability of Lise and FIND-maximum likelihood estimates]. Statistical Papers, 10(1):22–45, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927743.

Hofmann:1969:UIM

[138] Rolf Hofmann. Unternehmensanalyse im Maschinenbau. (German) []. Statistical Papers, 10(1):46–65, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927744.

Brendow:1969:EIO

[139] Klaus Brendow. Energieverbrauchsprognosen internationaler Organisationen für Westeuropa — eine Exaktheitsanalyse. (German) []. Statistical Papers, 10(1):66–85, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927745.

Mordi:1969:PSM

[140] O. Mordi. Zum Problem der statistischen Messung der finanziellen Hilfe für Entwicklungsländer. (German) []. Statistical Papers, 10(1):86–95, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927746.

Anonymous:1969:HCa

[141] Anonymous. Help & contacts. Statistical Papers, 10(1):??, March 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1969:DAb

[142] Anonymous. Zu dieser ausgabe. Statistical Papers, 10(2):97, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923078.

Voss:1969:AVV

[143] Werner Voß. Die außenwirtschaftlichen Verflechtungen der Volksrepublik China. (German) []. Statistical Papers, 10 (2):98–103, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923079.

Likes:1969:MVU

[144] Jiří Likeš, CSc. Minimum variance unbiased estimates of the parameters of power-function and Pareto’s distribution. Statistical Papers, 10(2):104–110, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923080.

Haq:1969:PLR

[145] M. Safiul Haq. On prediction from a linear regression model. Statistical Papers, 10(2):111–118, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923081.

Gossmann:1969:PNN

[146] Jens Goßmann and Günter Menges. Die Prognose der Nachfrage nach chemischen Produkten. (German) []. Statistical Papers, 10(2):119–134, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923082.

Litz:1969:SIGa

[147] Hans-Peter Litz and Manfred Tiede. Die Saisonschwankungen in der industriellen Güterproduktion der Bundesrepublik Deutschland 1954–1966. (German) []. Statistical Papers, 10 (2):135–170, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923083.

Bartelmus:1969:BIB

[148] Peter Bartelmus. Bereinigung industrieller Bruttoproduktionswerte mit Hilfe von Input–Output-Rechnungen. (German) []. Statistical Papers, 10 (2):171–181, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923084.

Anonymous:1969:HCb

[149] Anonymous. Help & contacts. Statistical Papers, 10(2):??, June 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1969:DAc

[150] Anonymous. Zu dieser ausgabe. Statistical Papers, 10(3):183, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923094.

Wagenfuhr:1969:PSR

[151] R. Wagenführ. Probleme einer Statistik der Rüstungen. (German) []. Statistical Papers, 10(3):184–191, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923095.

Gebert:1969:GFT

[152] J. R. Gebert and B. K. Kale. Goodness of fit tests based on discriminatory information. Statistical Papers, 10 (3):192–200, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923096.

Lauenstein:1969:PRE

[153] H. Lauenstein. Parameter-Restriktionen in der empirischen Nachfrageanalyse. (German) []. Statistical Papers, 10(3): 201–211, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923097.

Rutsch:1969:IOM

[154] M. Rutsch. An input-output model for education and manpower planning. Statistical Papers, 10(3):212– 222, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923098.

Litz:1969:SIGb

[155] Hans-Peter Litz and Manfred Tiede. Die Saisonschwankungen in der industriellen Güterproduktion der Bundesrepublik Deutschland 1954–1966. (German) []. Statistical Papers, 10(3): 223–247, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923099.

Meisel:1969:PMM

[156] Peter W. Meisel. Die Problematik der Modellannahmen und der Methode bei der Schätzung makroökonomischer Produktionsfunktionen und die Ursachen möglicher Verzerrungen der Ergebnisse. (German) []. Statistical Papers, 10(3):248–258, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923100.

Walsh:1969:EEP

[157] John E. Walsh. Existence of every possible distribution for any sample order statistic. Statistical Papers, 10(3): 259–260, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02923101.

Anonymous:1969:HCc

[158] Anonymous. Help & contacts. Statistical Papers, 10(3):??, September 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1969:DAd

[159] Anonymous. Zu dieser ausgabe. Statistical Papers, 10(4):263, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922832.

Wagenfuhr:1969:SMV

[160] Rolf Wagenführ. Die statistische Messung der Vorratsänderungen. (German) []. Statistical Papers, 10(4): 264–270, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922833.

Fries:1969:ASZ

[161] Christian Fries. Die Ausprägungen der Spektraltheorie in Zeitreihenanalyse und Quantenmechanik. (German) [].

Statistical Papers, 10(4):271–276, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922834.

Chan:1969:LQE

[162] Lai K. Chan. Linear quantile estimates of the location and scale parameters of the logistic distribution. Statistical Papers, 10(4):277–282, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922835.

Volodarskij:1969:WEE

[163] L. Volodarskij and M. Ejdel’man. Die wichtigsten Ergebnisse der expost-Input-Output-Tabelle der UdSSR für 1966. (German) []. Statistical Papers, 10(4):283–294, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922836.

Kuczynski:1969:AWG

[164] Jürgen Kuczynski. Die Arbeitsteilung in der Welt. (German) []. Statistical Papers, 10(4):295–298, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922837.

Beker:1969:AST

[165] Gabriele Beker. Anmerkungen zum statistischen Teil des Berichts der Bundesregierung über die Situation der Frauen in Familie, Beruf und Gesellschaft. (German) []. Statistical Papers, 10(4):299–307, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922838.

Wagenfuhr:1969:SDS

[166] Rolf Wagenführ. Streifzug durch das Statistische Jahrbuch für die BRD 1969. (German) []. Statistical Papers, 10(4):308–313, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922839.

Menges:1969:BSB

[167] Günter Menges and Martin Rutsch. Bemerkungen zum Substitutionsaxiom des Bernoullinutzens. (German) []. Statistical Papers, 10(4):314–315, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922840.

Anonymous:1969:HCd

[168] Anonymous. Help & contacts. Statistical Papers, 10(4):??, December 1969. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1970:DAa

[169] Anonymous. Zu dieser ausgabe. Statistical Papers, 11(1):1, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922290.

Wagenfuhr:1970:NHG

[170] Rolf Wagenführ. Neuer Hochstand der gewerblichen Sachgütererzeugung der Welt. (German) []. Statistical Papers, 11(1):2–5, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922291.

Heike:1970:BVF

[171] Hans-Dieter Heike. Ein Beitrag zur Verwendung von Fremdinformationen im Rahmen einer Strategie zur Festlegung ökonometrischer Strukturen — Einzelgleichungen. (German) [A contribution to the use of other information in the context of a policy of designed econometric structures — single equations]. Statistical Papers, 11(1):6–21, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922292.

Haq:1970:SLR

[172] M. Safiul Haq. On a simple linear regression analogue of the Behrens– Fisher problem. Statistical Papers, 11(1):22–29, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922293.

Wagenfuhr:1970:IBV

[173] Rolf Wagenführ. Der internationale bildungsstatistische Vergleich. (German) []. Statistical Papers, 11(1):30–56, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922294.

Steiger:1970:DMP

[174] Hans-Hermann Steiger. Definition und Messung von Preisstrukturen. (German) []. Statistical Papers, 11(1):57–63, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922295.

Elstermann:1970:B

[175] G. Elstermann and H. Bühlmann. Buchbesprechungen. Statistical Papers, 11(1):64–66, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922296.

Anonymous:1970:GMK

[176] Anonymous. Geschäftspolitik und Marktentscheidungen der kapitalistischen Unternehmungen (Auszug). (German) []. Statistical Papers, 11(1):67–71, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922297.

Anonymous:1970:HCa

[177] Anonymous. Help & contacts. Statistical Papers, 11(1):??, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1970:DAb

[178] Anonymous. Zu dieser ausgabe. Statistical Papers, 11(2):143, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922309.

Wagenfuhr:1970:EWS

[179] Rolf Wagenführ. Einführung in die Wirtschafts- und Sozialstatistik.

(German) []. Statistical Papers, 11 (2):144–152, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922310.

Munnich:1970:VTC

[180] F. E. Münnich. Verallgemeinerung eines Tests von Chow. (German) []. Statistical Papers, 11(2):153–159, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922311.

Hulsmann:1970:AVK

[181] Jochen Hülsmann and Volker Steinmetz. Das asymptotische Verhalten einer Klasse nichtregulärer Schätzfunktionen für den Erwartungswert. (German) []. Statistical Papers, 11 (2):160–165, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922312.

Esenwein-Rothe:1970:UFW

[182] Ingeborg Esenwein-Rothe. Ungenauigkeiten und Fehler in wirtschaftsstatistischen Daten. (German) []. Statistical Papers, 11(2):166–193, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922313.

Rodter:1970:PAL

[183] Gerhard Rödter. Zur Problematik des Aufbaus einer laufenden Statistik der Unternehmensgewinne. (German) []. Statistical Papers, 11(2):194–225, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922314.

Schmidt:1970:NIB

[184] Hans-Joachim Schmidt. Die Nettoproduktion im Bankensektor. (German) []. Statistical Papers, 11(2):226–233, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922315.

Petho:1970:ABS

[185] Szilvester Pethö and Z. Szarka. Eine Art der Bestimmung des Stichprobenumfangs bei geschichteter Stichprobe. (German) []. Statistical Papers, 11 (2):234–236, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922316.

Anonymous:1970:HCb

[186] Anonymous. Help & contacts. Statistical Papers, 11(2):??, June 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1970:DAc

[187] Anonymous. Zu dieser ausgabe. Statistical Papers, 11(3):73, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922892.

Litz:1970:NAB

[188] Hans-Peter Litz. Neue Ansätze in der Bildungsstatistik. (German) []. Statistical Papers, 11(3):74–85, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922893.

Winckler:1970:MBT

[189] Klaus Winckler. Multicollinearity and blockcollinearity — their effects on the estimation of structural parameters in linear economic models. Statistical Papers, 11(3):86–112, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922894.

Frenzel:1970:VWS

[190] Gottfried Frenzel and Friedrich J. Hartmann. Zum Vergleich der wirtschaftsstatistischen Systeme kapitalistischer und sozialistischer Länder. (German) []. Statistical Papers, 11 (3):113–129, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922895.

Voss:1970:MSP

[191] Werner Voß. Die Möglichkeit der Schätzung der Parameter theoretischer Funktionen bei gegebenem Datenmaterial mit Hilfe eines Computerprogramms iterativer Approximation. (German) []. Statistical Papers, 11 (3):130–138, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922896.

Menges:1970:ZNL

[192] Günter Menges. Zwei neue Lehrbücher am Rande der Statistik. (German)

[]. Statistical Papers, 11(3):139–140, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922897.

Anonymous:1970:HCc

[193] Anonymous. Help & contacts. Statistical Papers, 11(3):??, March 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1970:DAd

[194] Anonymous. Zu dieser ausgabe. Statistical Papers, 11(4):237, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922935.

Tiede:1970:FKI

[195] Manfred Tiede. Zur Frage konjunktureller Indikatoren. (German) [On the question of economic indicators]. Statistical Papers, 11(4):238–241, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922936.

Schneeberger:1970:OSD

[196] Hans Schneeberger. Optimierung in der Stichprobentheorie durch Schichtung. (German) []. Statistical Papers, 11 (4):242–253, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922937.

Giri:1970:BEM

[197] Narayan Giri. Bayesian estimation of means of a two-way classification with random effect model. Statistical Papers, 11(4):254–260, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922938.

Wintgen:1970:MDK

[198] Georg Wintgen. Zur Mengentheoretischen Definition und Klassifizierung kybernetischer Systeme. (German) []. Statistical Papers, 11(4):261–298, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922939.

Krug:1970:MRD

[199] Walter Krug. Multiple Regressionsanalyse oder dichotomisches Aufbereiten von Daten?. (German) [Multiple regression analysis or dichotomous processing of data?]. Statistical Papers, 11(4):299–313, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922940.

Voss:1970:GEV

[200] Werner Voß. Grundlagen einer Entwicklungsprognose für die Volksrepublik China. (German) []. Statistical Papers, 11(4):314–323, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922941.

Anonymous:1970:HCd

[201] Anonymous. Help & contacts. Statistical Papers, 11(4):??, December 1970. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1971:DAa

[202] Anonymous. Zu dieser ausgabe. Statistical Papers, 12(1):1, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923608.

Paschen:1971:MBU

[203] Herbert Paschen and Raphael Buyse. Zur Messung der Betriebs- und Unternehmenskonzentration. (German) []. Statistical Papers, 12(1):2–13, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923609.

Feichtinger:1971:EOV

[204] Gustav Feichtinger. Zur erlernung optimalen verhaltens in entscheidungssituationen. Statistical Papers, 12 (1):14–21, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923610.

Menges:1971:SDI

[205] Günter Menges. Some decision- and information-theoretical considerations about the econometric problems of specification and identification. Statistical Papers, 12(1):22–31, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923611.

Meissner:1971:IAO

[206] Werner Meißner. Informationstheoretische Analyse ökonometrischer Modellbildung. (German) []. Statistical Papers, 12(1):32–36, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923612.

Mordi:1971:QTT

[207] Obi Mordi. Zur Quantifizierung des Terms-of-trade-Effekts. (German) []. Statistical Papers, 12(1):37–46, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923613.

Bender:1971:RGB

[208] Ulrich Bender. Die rechtlichen Grundlagen der Bildungsstatistik. Darstellung der Rechtsvorschriften und der Auswirkungen auf bildungsstatistische Ergebnisse. (German) []. Statistical Papers, 12(1):47–71, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923614.

vonderLippe:1971:MMS

[209] Peter von der Lippe. Mathematische Modelle der sozialen Differenzierung. (German) []. Statistical Papers, 12 (1):72–83, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923615.

Skala:1971:BVS

[210] Heinz J. Skala. Bemerkungen zur Verallgemeinerung des Shapleyschen Wertes. (German) []. Statistical Papers, 12(1):84–87, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923616.

Gunzert:1971:BRU

[211] Rudolf Gunzert. Book review: Ulrich Bender, Die Rechtsgrundlagen der Statistik in der BRD und ihre Auswirkungen auf die statistischen Ergebnisse. Statistical Papers, 12 (1):87–88, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02923617.

Anonymous:1971:HCa

[212] Anonymous. Help & contacts. Statistical Papers, 12(1):??, March 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1971:DAb

[213] Anonymous. Zu dieser ausgabe. Statistical Papers, 12(2):91–92, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922942.

Voss:1971:SFG

[214] Werner Voß. Statistik in der Friedensforschung. (German) []. Statistical Papers, 12(2):93–99, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922943.

Consul:1971:LGD

[215] P. C. Consul and G. C. Jain. On the log-gamma distribution and its properties. Statistical Papers, 12(2):100–106, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922944.

Cleroux:1971:RII

[216] Robert Cléroux. On the recovery of interblock information in intra and intergroup balanced incomplete block designs. Statistical Papers, 12(2):107–113, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922945.

Schips:1971:BAM

[217] Bernd Schips and W. Stier. Bestimmung der Auswirkung von Multikollinearität zwischen den erklärenden Variablen in linearen Regressionsmodellen auf Kleinst–Quadrate– Schätzwerte durch Simulation. (German) []. Statistical Papers, 12 (2):114–126, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922946.

Buyse:1971:KALa

[218] Raphael Buyse and Herbert Paschen. Die Konzentration in der Automobilindustrie in den Ländern der EWG. (German) []. Statistical Papers, 12 (2):127–142, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922947.

Tiede:1971:OVB

[219] Manfred Tiede. Optimale Verhaltensweisen bei Blackjack. (German) []. Statistical Papers, 12(2):143–154, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922948.

Kuczynski:1971:PBA

[220] Jürgen Kuczynski. Zur Problematik der Berechnung der Arbeitsproduktivität in einzelnen Industrien und Wirtschaftszweigen. (German) []. Statistical Papers, 12(2):155–157, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922949.

Srinivasan:1971:TE

[221] R. Srinivasan. Tests for exponentiality. Statistical Papers, 12(2):157–160, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922950.

Anonymous:1971:HCb

[222] Anonymous. Help & contacts. Statistical Papers, 12(2):??, June 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1971:DAc

[223] Anonymous. Zu dieser ausgabe. Statistical Papers, 12(3–4):161–162, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923560.

Wagenfuhr:1971:AB

[224] Rolf Wagenführ. Der Aktuelle Beitrag. Statistical Papers, 12(3–4): 163–166, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923561.

Werner:1971:PIV

[225] Heinz Werner. Zum Problem der internationalen Vergleichbarkeit von Arbeitslosenquoten. (German) []. Statistical Papers, 12(3–4):167–176, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923562.

Bury:1971:SIA

[226] K. V. Bury and B. Bernholtz. On structural inference applied to the Weibull distribution. Statistical Papers, 12(3– 4):177–192, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923563.

Haq:1971:LIR

[227] M. Safiul Haq. On likelihood inference for the ratio of scale parameters. Statistical Papers, 12(3–4):193–203, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923564.

Bruckmann:1971:GCC

[228] Gerhart Bruckmann. A generalized concept of concentration and its measurement. Statistical Papers, 12(3–4): 204–223, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923565.

Winckler:1971:THC

[229] Klaus Winckler. Testing hypotheses and the construction of confidence intervals for the parameters of stochastic linear difference equations in small samples. Statistical Papers, 12(3–4):224–244, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923566.

Buyse:1971:KALb

[230] Raphael Buyse and Herbert Paschen. Die Konzentration in der Automobilindustrie in den Ländern der EWG Teil 2. (German) []. Statistical Papers, 12(3–4):245–269, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923567.

Freytag:1971:SPI

[231] Hans Ludwig Freytag. Statistische Probleme des internationalen Lebensniveau–Vergleichs. (German) []. Statistical Papers, 12(3–4):270– 284, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923568.

Gergely:1971:PML

[232] István Gergely. Plan und Markt-Lenkungsmethoden in Ungarn. (German) [Plan and market-steering methods in Hungary]. Statistical Papers, 12(3–4):285–306, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923569.

Granzer:1971:ASZ

[233] Randolf Gränzer. Das Additionsproblem bei der Saisonbereinigung von Zeitreihensätzen. (German) []. Statistical Papers, 12(3–4):307–313, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923570.

Voss:1971:MOG

[234] Werner Voß. Minimierung und Optimierung. (German) []. Statistical Papers, 12(3–4):314–318, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923571.

Garbers:1971:APK

[235] Hermann Garbers. Eine Anmerkung zum Problem konjunktureller Indikatoren. (German) []. Statistical Papers, 12(3–4):319–322, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923572.

Garbers:1971:BSG

[236] Hermann Garbers. Zur Bewertung von Saisonbereinigungsverfahren.

(German) []. Statistical Papers, 12(3– 4):323–324, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02923573.

Walsh:1971:ETB

[237] John E. Walsh and Grace J. Kelleher. Exact tests of binomial p=1/ 2 when one binomial event possibly has different probability. Statistical Papers, 12(3–4):325–328, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923574.

Anonymous:1971:HCc

[238] Anonymous. Help & contacts. Statistical Papers, 12(3–4):??, September 1971. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1972:DAa

[239] Anonymous. Zu dieser ausgabe. Statistical Papers, 13(1):1–2, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923655.

Zimmermann:1972:JWA

[240] Karl Zimmermann. Der Jahresabschluß unter wirtschaftswissenschaftlichen Aspekten. (German) []. Statistical Papers, 13(1):3–14, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923656.

Stange:1972:BGG

[241] K. Stange. Zur Beurteilung einer Garantiekurve. (German) [Assessing a guarantee curve]. Statistical Papers, 13(1):15–40, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923657.

Chan:1972:OSA

[242] Lai K. Chan and Smiley Cheng. Optimum spacing for the asymptotically best linear estimate of the location parameter of the logistic distribution when samples are complete or censored. Statistical Papers, 13(1):41–57, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923658.

Ehrlich:1972:EVV

[243] Eva Ehrlich. Das Entwicklungsniveau verschiedener Volkswirtschaften — Wachstumsvergleiche und Prognosemethoden. (German) [The level of development of different economies — comparisons of growth and forecasting methods]. Statistical Papers, 13(1):58–95, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923659.

Buyse:1972:KAL

[244] Raphael Buyse and Herbert Paschen. Die Konzentration in der Automobilindustrie in den Ländern der EWG (Teil 3). (German) []. Statistical Papers, 13(1):96–101, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923660.

Kelleher:1972:EIT

[245] Grace J. Kelleher and John E. Walsh. Exact intervals and tests for mean of symmetrical population when one “sample” value possibly an outlier. Statistical Papers, 13(1):102–105, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923661.

Bedau:1972:WSA

[246] Klaus-Dietrich Bedau. Wilfried Schreiber: Ein analytisch-numerisches Gesamt–Modell der Volkswirtschaft als Hilfsmittel der Wachstums-, Konjunkturund Lohntheorie. Erste Ausbaustufe: Erweitertes Ein–Gut-Modell. (German) []. Statistical Papers, 13(1): 106–108, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923662.

Anonymous:1972:HCa

[247] Anonymous. Help & contacts. Statistical Papers, 13(1):??, March 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1972:DAb

[248] Anonymous. Zu dieser ausgabe. Statistical Papers, 13(2):109–110, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02927762.

Wagenfuhr:1972:HSV

[249] Rolf Wagenführ. Holt die Sowjetunion die Vereinigten Staaten von Amerika ein?. (German) []. Statistical Papers, 13(2):111–115, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927763.

Schneeweiss:1972:OPS

[250] Christoph Schneeweiß. Optimale Prognosen und suffiziente Statistiken in quadratischen dynamischen Optimierungsproblemen. (German) []. Statistical Papers, 13(2):116–129, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927764.

Garg:1972:EPM

[251] Mohan L. Garg, B. Raja Rao, and Sati Mazumdar. On an estimation problem in multiple regression. Statistical Papers, 13(2):130–144, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927765.

Goldstein:1972:VSF

[252] Bernd Goldstein. Über den Vergleich der superharmonischen Funktionen zweier diskreter transienter Markoff– Ketten. (German) []. Statistical Papers, 13(2):145–159, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927766.

Marfels:1972:GRC

[253] Christian Marfels. The Gini ratio of concentration reconsidered. Statistical Papers, 13(2):160–179, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927767.

Bender:1972:MGE

[254] Ulrich Bender. Möglichkeiten und Grenzen empirischer Modeforschung. (German) []. Statistical Papers, 13 (2):180–190, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02927768.

Gerber:1972:BED

[255] Klaus Gerber. Die Belastung der Endnachfrage durch die kumulative Brutto–Allphasen-Umsatzsteuer. (German) []. Statistical Papers, 13 (2):191–202, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02927769.

Anonymous:1972:HCb

[256] Anonymous. Help & contacts. Statistical Papers, 13(2):??, June 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1972:DAc

[257] Anonymous. Zu dieser ausgabe. Statistical Papers, 13(3):203–204, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

com/accesspage/article/10.1007/
BF02923499.

Wagenfuhr:1972:EGV

[258] Rolf Wagenführ and Mitarbeiter. Einige Grunddaten der Volkswirtschaft der BRD, berechnet nach dem Konzept der materiellen Produktion. (German) []. Statistical Papers, 13(3): 205–210, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923500.

Kriz:1972:PVG

[259] Jürgen Kriz. Die PMP–Verteilung. (German) [The PMP distribution]. Statistical Papers, 13(3):211–224, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923501.

Heuts:1972:PEE

[260] R. M. J. Heuts. Parameter estimation in the exponential distribution, confidence intervals and a Monte Carlo study for a goodness of fit test. Statistical Papers, 13(3):225–246, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923502.

Frenzel:1972:BSW

[261] Gottfried Frenzel and Siegfried Vetter. Bezugssysteme und Standardstrukturen wirtschaftlicher Entwicklung. (German) []. Statistical Papers, 13(3): 247–269, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923503.

Bedau:1972:IBM

[262] Klaus-Dietrich Bedau. Informationstheoretisch begründete Messung von Einkommensdisparitäten. (German) []. Statistical Papers, 13(3): 270–290, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923504.

Likes:1972:EPP

[263] Jiří Likeš, CSc. Estimation of probabilities Pr(a < x < b) for a Pareto distribution. Statistical Papers, 13(3): 291–297, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923505.

Wagenfuhr:1972:STU

[264] Rolf Wagenführ. Statistisches Taschenbuch Ungarns 1972. Budapest 1972. (German) []. Statistical Papers, 13(3): 298–299, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02923506.

Mordi:1972:BRP

[265] Obi Mordi. Book review: Péter Vas–Zoltán: United Nations Technical Assistance Verlag der Ungarischen Adademie der Wissenschaften, Budapest V. Statistical Papers, 13(3): 300–305, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923507.

Anonymous:1972:HCc

[266] Anonymous. Help & contacts. Statistical Papers, 13(3):??, September 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1972:DAd

[267] Anonymous. Zu dieser ausgabe. Statistical Papers, 13(4):307–308, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922317.

Suppanz:1972:AVG

[268] Karl Suppanz. Von Aggregaten der Volkswirtschaftlichen Gesamtrechnung. (German) []. Statistical Papers, 13(4):309–313, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02922318.

Stenger:1972:EZS

[269] Horst Stenger. Effizienzvergleiche zwischen stichprobenverfahren. Statistical Papers, 13(4):314–334, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922319.

Stenger:1972:LNS

[270] Horst Stenger. Lineare und nichtlineare Schätzfunktionen in der Stichprobentheorie. (German) []. Statistical Papers, 13(4):335–353, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922320.

Szilagyi:1972:APC

[271] György Szilágyi. Analyses of price changes by means of input-output tables. Statistical Papers, 13(4): 354–362, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922321.

Hauser:1972:WRA

[272] Siegfried Hauser and Siegfried Lörcher. Zur wirtschaftspolitischen Relevanz der Amtlichen Statistik Ein Vergleich BRD-Japan. (German) []. Statistical Papers, 13(4):363–387, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922322.

Schulze:1972:BEA

[273] Peter M. Schulze. Bemerkungen zu einem Aufsatz von Wintgen: Einige Folgerungen für die wirtschaftstheoretischen und ökonometrischen Begriffe Modell und Struktur. (German) []. Statistical Papers, 13(4):388–390, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922323.

Leiner:1972:B

[274] Bernd Leiner. Buchbesprechung. Statistical Papers, 13(4):391–392, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/accesspage/article/10.1007/
                                        BF02922324.

Anonymous:1972:HCd

[275] Anonymous. Help & contacts. Statistical Papers, 13(4):??, December 1972. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1973:DAa

[276] Anonymous. Zu dieser ausgabe. Statistical Papers, 14(1):1–2, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922402.

Wagenfuhr:1973:NEI

[277] Rolf Wagenführ. Niveau und Entwicklung der Infrastrukturen im internationalen statistischen Vergleich. (German) []. Statistical Papers, 14(1):3–5, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922403.

Drygas:1973:MLL

[278] Hilmar Drygas. Median and linear loss-functions. Statistical Papers, 14(1):6–12, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922404.

Heuts:1973:NTS

[279] R. M. J. Heuts. A new test statistic for searching hidden periodicities in time series and the derivation and numerical calculation of its power function. Statistical Papers, 14(1):13–38, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922405.

Tiede:1973:TPV

[280] Manfred Tiede. Theorie und Praxis eines verbesserten Suchverfahrens. (German) []. Statistical Papers, 14 (1):39–58, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922406.

Gerling:1973:EMB

[281] Bernd Gerling. Die Einführung der Mehrwertsteuer in der BRD in ihren Auswirkungen auf die wichtigsten Wirtschaftsstatistiken — insbesondere auf die Angaben der Volkswirtschaftlichen Gesamtrechnungen. (German) []. Statistical Papers, 14 (1):59–83, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922407.

Gupta:1973:NMI

[282] R. P. Gupta. A note on multicollinearity and imprecise estimation. Statistical Papers, 14(1):84–87, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922408.

Stumpf:1973:DIG

[283] Horst Stumpf. Demographische Inputs und Graduiertenentwicklung. (German) []. Statistical Papers, 14(1):88–97, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922409.

Anonymous:1973:HCa

[284] Anonymous. Help & contacts. Statistical Papers, 14(1):??, March 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1973:DAb

[285] Anonymous. Zu dieser ausgabe. Statistical Papers, 14(2):99–100, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922875.

Haq:1973:SAB

[286] M. Safiul Haq. Structural analysis for the bilinear model with linearly related responses. Statistical Papers, 14(2):101–110, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922876.

Bury:1973:SIT

[287] K. V. Bury. Structural inferences on the type I extreme value distribution. Statistical Papers, 14(2):111–122, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922877.

Bartelmus:1973:PEU

[288] Peter Bartelmus. Probleme der Entwicklung eines umweltstatistischen Systems. (German) []. Statistical Papers, 14(2):123–148, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922878.

Wagenfuhr:1973:GGB

[289] Rolf Wagenführ and Mitarbeiter. Das “gesellschaftliche Gesamtprodukt” der BRD in seiner Gliederung nach Produktions- und Konsumtionsmitteln. (German) []. Statistical Papers, 14(2):149–161, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922879.

Vetter:1973:IVW

[290] Siegfried Vetter. Internationaler Vergleich von Wirtschaftsstrukturen auf der Basis standardisierter Input– Output-Tabellen für 8 europäische Länder. (German) []. Statistical Papers, 14(2):162–180, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922880.

Voss:1973:EBC

[291] Werner Voß. Einige Beispiele zum Computereinsatz in der beschreibenden Statistik. (German) []. Statistical Papers, 14(2):181–192, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922881.

Riedwyl:1973:RWG

[292] Hans Riedwyl. Zum Rangsummentest von Wilcoxon. (German) []. Statistical Papers, 14(2):193–202, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922882.

Elstermann:1973:WAC

[293] Gert Elstermann. Willi Albert und Christoph Oehler: Die Kulturausgaben der Länder, des Bundes und der Gemeinden 1950 bis 1967. (German) []. Statistical Papers, 14(2):203–204, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922883.

Anonymous:1973:HCb

[294] Anonymous. Help & contacts. Statistical Papers, 14(2):??, June 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1973:DAc

[295] Anonymous. Zu dieser ausgabe. Statistical Papers, 14(3):205, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02929694.

Stange:1973:ZTK

[296] Kurt Stange. Über einen zweiseitigen Test für die Korrelationszahl einer zweidimensionalen Normalverteilung. (German) []. Statistical Papers, 14(3):206–236, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02929695.

Kockelkorn:1973:SKZ

[297] Ulrich Kockelkorn and Bernhard Rüger. Schätzung der Komponenten einer Zeitreihe mit konstanter Saisonfigur nach der Methode der gleitenden Durchschnitte. (German) []. Statistical Papers, 14(3):237–270, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929696.

Leiner:1973:EBD

[298] Bernd Leiner. Einige Bemerkungen zum Durbin–Watson-Test. (German) []. Statistical Papers, 14(3): 271–273, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02929697.

Moeschlin:1973:B

[299] O. Moeschlin and Jochen Schwarze. Buchbesprechung. Statistical Papers, 14(3):274–276, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02929698.

Rao:1973:NCD

[300] B. Raja Rao, Sati Mazumdar, and C. C. Li. A note on a conditional distribution involving a Yule process. Statistical Papers, 14(3):277–281, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929699.

Samanta:1973:NCM

[301] M. Samanta. A note on comparing the means of a multivariate normal population. Statistical Papers, 14(3):282–286, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02929700.

Weissker:1973:KAR

[302] Jürgen Weißker. Konzentrationsmessung — absolut und relativ. (German) []. Statistical Papers, 14(3): 287–290, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02929701.

Anonymous:1973:HCc

[303] Anonymous. Help & contacts. Statistical Papers, 14(3):??, September 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1973:DAd

[304] Anonymous. Zu dieser ausgabe. Statistical Papers, 14(4):291, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923064.

Stenger:1973:SSG

[305] H. Stenger. Symmetriebegriffe in der Stichprobentheorie. (German) []. Statistical Papers, 14(4):292–309, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923065.

Feichtinger:1973:MMS

[306] G. Feichtinger. Markovian models for some demographic processes. Statistical Papers, 14(4):310–334, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923066.

Zseby:1973:BIG

[307] S. Zseby. Bayessche Intervallschätzung bei geschichteter Stichprobe. (German) []. Statistical Papers, 14(4): 335–349, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923067.

Ali:1973:DSR

[308] M. M. Ali. Distribution of “Student” ratio for correlated observations. Statistical Papers, 14(4):350–356, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923068.

Riedwyl:1973:RWS

[309] H. Riedwyl. Zum Rangsummentest von Wilcoxon, Statistische Hefte. (German) []. Statistical Papers, 14(4):357, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923069.

Anonymous:1973:HCd

[310] Anonymous. Help & contacts. Statistical Papers, 14(4):??, December 1973. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1974:DAa

[311] Anonymous. Zu dieser ausgabe. Statistical Papers, 15(1):1–2, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923639.

Stadlbauer:1974:VKU

[312] Josef Stadlbauer. Vergleich der Konsumentenkaufkraft in Ungarn und der BRD. (German) []. Statistical Papers, 15(1):3–13, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923640.

Cunningham:1974:RUS

[313] A. A. Cunningham. The reduction of uncertainty by structural inference. Statistical Papers, 15(1):14–26, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923641.

Whitney:1974:TCS

[314] J. B. Whitney and Ch. E. Minder. Time censoring and the structural model. Statistical Papers, 15(1):27–35, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923642.

Wagenfuhr:1974:FBI

[315] Rolf Wagenführ and Mitarbeiter. Die “finanzielle Bilanz” im Rahmen des MPS–Systems und ihre Anwendbarkeit auf die Gesamtrechnungsdaten der BRD. (German) []. Statistical Papers, 15(1):36–64, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923643.

Hassanein:1974:LEP

[316] K. M. Hassanein. Linear estimation of the parameters of the logistic distribution by selected order statistics for very large samples. Statistical Papers, 15(1):65–70, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923644.

Kreyszig:1974:SNG

[317] E. Kreyszig. Schranken für die Normalverteilung. (German) []. Statistical Papers, 15(1):71–74, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923645.

Marfels:1974:GRC

[318] Christian Marfels. The Gini ratio of concentration reconsidered: Reply. Statistical Papers, 15(1):75–80, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923646.

Kohle:1974:EBM

[319] Dieter Köhle. Einige Bemerkungen zur Messung von Faktorenwerten im orthogonalen Faktorenmodell. (German) []. Statistical Papers, 15(1):81–87, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923647.

Pfeilsticker:1974:KVK

[320] Arne Pfeilsticker and Hans-Udo Schmidt. Kritik und Vorschläge zur Konzeption und Gestaltung von Lehrbüchern. (German) []. Statistical Papers, 15(1):88–102, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923648.

Anonymous:1974:HCa

[321] Anonymous. Help & contacts. Statistical Papers, 15(1):??, March 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1974:DAb

[322] Anonymous. Zu dieser ausgabe. Statistical Papers, 15(2–3):103, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922898.

Minton:1974:MDJ

[323] P. D. Minton. In memoriam Dr. John E. Walsh 1919–1972. Statistical Papers, 15(2–3):104–105, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02922899.

Walsh:1974:NMM

[324] J. E. Walsh and G. J. Kelleher. Nonparametric models and methods for one-way ANOVA with fixed effects. Statistical Papers, 15(2–3):106–115, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922900.

Walsh:1974:NMR

[325] J. E. Walsh and G. J. Kelleher. Nonparametric models and results for twoway ANOVA with fixed effects and interactions. Statistical Papers, 15 (2–3):116–129, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922901.

Walsh:1974:NRT

[326] J. E. Walsh and G. J. Kelleher. Nonparametric results for two-way ANOVA with fixed effects, no interactions, and unequal subclasses. Statistical Papers, 15(2–3):130–141, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922902.

Birnbaum:1974:NTS

[327] Z. W. Birnbaum and H. J. Friedman. Numerical tabulations for a statistic similar to Student’s t. Statistical Papers, 15(2–3):143–156, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922903.

Dixon:1974:TWR

[328] W. J. Dixon and K. K. Yuen. Trimming and Winsorization: A review. Statistical Papers, 15(2–3):157–170, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922904.

Drane:1974:RHS

[329] J. W. Drane and R. Harrist. Resolving hypotheses with successive chisquares. Statistical Papers, 15(2–3):171–180, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922905.

Menges:1974:SUF

[330] G. Menges and B. Jacke. The scientist’s utility function and the principle of maximum likelihood. Statistical Papers, 15(2–3):181–201, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922906.

Murthy:1974:CFD

[331] V. K. Murthy and G. B. Swartz. Cumulative fatigue damage — theory and models. Statistical Papers, 15 (2–3):202–231, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922907.

Anonymous:1974:HCb

[332] Anonymous. Help & contacts. Statistical Papers, 15(2–3):??, June 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1974:DAc

[333] Anonymous. Zu dieser. ausgabe. Statistical Papers, 15(4):233, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922908.

Heuts:1974:EAC

[334] R. M. J. Heuts. An efficient algorithm to calculate the least squares estimates of the parameters in the logistic growth function and the application of Beale’s measures of nonlinearity for this model. Statistical Papers, 15(4):234–255, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922909.

Rinne:1974:KER

[335] H. Rinne. Kontrolle und Erneuerung von Reservelagern. (German) []. Statistical Papers, 15(4):256–275, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922910.

Menges:1974:TNS

[336] Günter Menges and Peter Beutel. Triangulation der nach 11 Sektoren gegliederten Input–Output-Tabellen der Länder der alten EWG. (German) []. Statistical Papers, 15(4): 276–302, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922911.

Herbel:1974:SVM

[337] Norbert Herbel. Statistische Versuche der Messung der Infrastruktur, gezeigt am Beispiel ungarischer Untersuchungen unter Einbeziehung der UdSSR. — Darstellung und kritische Stellungnahme. (German) []. Statistical Papers, 15(4):303–323, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922912.

Elstermann:1974:GPU

[338] Gert Elstermann. Über die Grundausstattung von Professoren an Universitäten und Technischen Hochschulen — Ergebnisse einer Erhebung. (German) []. Statistical Papers, 15(4): 324–328, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922913.

Haq:1974:IAM

[339] M. S. Haq. On inference about the mean in the presence of correlation. Statistical Papers, 15(4):329–334, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922914.

Tan:1974:SRS

[340] P. Tan and Aly Sherif. Some remarks on structural inference applied to Weibull distribution. Statistical Papers, 15(4):335–341, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922915.

Anonymous:1974:HCc

[341] Anonymous. Help & contacts. Statistical Papers, 15(4):??, December 1974. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1975:DAa

[342] Anonymous. Zu dieser ausgabe. Statistical Papers, 16(1):1, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923049.

Haq:1975:MLI

[343] M. Safiul Haq. On marginal likelihood inference about the parameters of linearly related responses. Statistical Papers, 16(1):2–13, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923050.

Hujer:1975:POM

[344] R. Hujer. Prognosegüte ökonometrischer Modelle — Ein Beitrag zum “Stabilitätsproblem der ökonometrie”. (German) []. Statistical Papers, 16 (1):14–28, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923051.

Leiner:1975:VEA

[345] B. Leiner. Vergleich einiger autoprojektiver Verfahren. (German) []. Statistical Papers, 16(1):29–38, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923052.

Laga:1975:SSD

[346] Josef Laga and Jiri Likes. Sample sizes for distribution-free tolerance intervals. Statistical Papers, 16(1):39–56, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923053.

Schulze:1975:EBR

[347] Peter M. Schulze. Einige Bemerkungen zu den regionalstatistischen Einheiten der Europäischen Gemeinschaften. (German) []. Statistical Papers, 16 (1):57–60, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923054.

Voss:1975:MSP

[348] Werner Voß. Zur Messung der Streuung personeller Einkommensverteilungen. (German) []. Statistical Papers, 16(1):61–68, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923055.

Leiner:1975:BRD

[349] Bernd Leiner. Book review: Dietrich Zschocke: Betriebsökonometrie Physica Verlag. Würzburg-Wien 1974. 287 S. Statistical Papers, 16(1):69–70, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923056.

Anonymous:1975:HCa

[350] Anonymous. Help & contacts. Statistical Papers, 16(1):??, March 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1975:DAb

[351] Anonymous. Zu dieser ausgabe. Statistical Papers, 16(2):73, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02922916.

Rao:1975:CDI

[352] B. R. Rao and S. Mazumdar. A conditional distribution involving a multidimensional Yule process with an application. Statistical Papers, 16(2):74–84, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922917.

Reiser:1975:SIL

[353] B. Reiser. Structural inference for linear regression with autocorrelated errors. Statistical Papers, 16(2):85–104, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922918.

Bury:1975:DSL

[354] K. V. Bury. Distribution of smallest log-normal and gamma extremes. Statistical Papers, 16(2):105–114, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922919.

Farewell:1975:ISM

[355] V. T. Farewell and R. L. Prentice. Interpreting the structural model. Statistical Papers, 16(2):115–122, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922920.

Haagen:1975:AHV

[356] K. Haagen and W. Schweitzer. Zur Approximation der hypergeometrischen Verteilung durch die Normalverteilung. (German) []. Statistical Papers, 16 (2):123–127, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02922921.

Anonymous:1975:HCb

[357] Anonymous. Help & contacts. Statistical Papers, 16(2):??, June 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1975:DAc

[358] Anonymous. Zu dieser ausgabe. Statistical Papers, 16(3):129, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02922995.

Broeckx:1975:NEB

[359] F. Broeckx, L. D’Hooge, M. Goovaerts, and J. van den Broeck. Numerical evaluation of bounded Bayesian parameters in case of autocorrelated errors and multicollinearity in data. Statistical Papers, 16(3):130–143, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02922996.

Menges:1975:WMO

[360] G. Menges. Weiche Modelle in Ökonometrie und Statistik. (German) []. Statistical Papers, 16(3): 144–156, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922997.

Heuts:1975:SFM

[361] R. M. J. Heuts and P. J. Rens. Several forecast models applied to a specific economic time series. Statistical Papers, 16(3):157–187, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02922998.

Davies:1975:EEA

[362] L. Davies and G. Ronning. Einige exakte und asymptotische Ergebnisse für das Standardmodell der Portefeuille– Auswahl innerhalb einer Periode. (German) []. Statistical Papers, 16(3): 188–212, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02922999.

Saleh:1975:FER

[363] A. E. Saleh and G. H. Choudhry. On fitting exponential regressions. Statistical Papers, 16(3):213–222, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923000.

Esenwein-Rothe:1975:GRW

[364] I. Esenwein-Rothe. Zum Gedächtnis an Rolf Wagenführ 1905–1975. (German) []. Statistical Papers, 16(3): 227–232, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923001.

Bihn:1975:EBK

[365] W. R. Bihn. Einige Betrachtungen zur kohärenten Messung von Disparitätsund Konzentrationsphänomenen. (German) []. Statistical Papers, 16(3): 233–249, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923002.

Ferschl:1975:BSM

[366] F. Ferschl. Zum Begriff “Statistische Masse”. (German) [On the concept of “statistical measure”]. Statistical Papers, 16(3):250–255, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923003.

Gollnick:1975:KAK

[367] H. Gollnick. Konjunkturelle Auswirkungen einer Konstanz der makroökonomischen Konsumfunktion Mitte der siebziger Jahre. (German) []. Statistical Papers, 16(3):256–267, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02923004.

Gulicher:1975:LTM

[368] H. Gülicher. Ein Lagetheorem für einen Teilmedian eines k-Medians und damit eines von k optimalen Standorten einer diskreten räumlichen Verteilung mit Engpaßbereichen. (German) []. Statistical Papers, 16(3):268–285, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923005.

Hampe:1975:ASG

[369] A. Hampe. Ausbildungsförderung und Studienleistung. (German) []. Statistical Papers, 16(3):286–295, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923006.

Menges:1975:IVA

[370] G. Menges and R. Zwer. Der internationale Vergleich aufgrund von Input– Output-Tabellen. (German) []. Statistical Papers, 16(3):296–315, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923007.

Schneeweiss:1975:SID

[371] H. Schneeweiß. Struktur-Inferenz an drei Beispielen. (German) []. Statistical Papers, 16(3):316–326, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923008.

Zahlen:1975:EIB

[372] J. P. Zahlen. Über einige identische Beziehungen zwischen verschiedenen Verteilungsfunktionen und ihre Verwendung in der mathematischen Statistik. (German) []. Statistical Papers, 16 (3):327–341, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923009.

Anonymous:1975:HCc

[373] Anonymous. Help & contacts. Statistical Papers, 16(3):??, September 1975. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1976:DAa

[374] Anonymous. Zu dieser ausgabe. Statistical Papers, 17(1):1, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923308.

Kofler:1976:KEU

[375] E. Kofler. Konfidenzintervalle in Entscheidungen bei Ungewißheit. (German) []. Statistical Papers, 17(1):2–21, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923309.

Schildkamp-Kundiger:1976:PME

[376] E. Schildkamp-Kündiger. Probleme des Messens in empirisch arbeitenden Sozialwissenschaften. (German) []. Statistical Papers, 17(1):22–32, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923310.

Schriever:1976:FAS

[377] K.-H. Schriever. Falsche Anwendung statistischer Methoden bei der Auswertung medizinisch wissenschaftlicher Versuche. (German) []. Statistical Papers, 17(1):33–42, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02923311.

Bomsdorf:1976:IAW

[378] E. Bomsdorf. Invarianz und Additivität, ein Widerspruch?. (German) []. Statistical Papers, 17(1):43–49, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923312.

Hild:1976:NSM

[379] C. Hild. A note on some matrices used in linear regression models. Statistical Papers, 17(1):50–56, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923313.

Leiner:1976:KRS

[380] B. Leiner. Kritische reflexionen zu einer simulationsstudie. Statistical Papers, 17(1):57–60, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923314.

Bedau:1976:BRW

[381] K. D. Bedau. Book review: Walter Piesch: Statistische Konzentrationsmaße. Statistical Papers, 17(1):61–62, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923315.

Reiser:1976:SIL

[382] B. Reiser. Structural inference for linear regression with autocorrelated errors. Statistical Papers, 17(1):64, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923316.

Anonymous:1976:HCa

[383] Anonymous. Help & contacts. Statistical Papers, 17(1):??, March 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1976:DAb

[384] Anonymous. Zu dieser ausgabe. Statistical Papers, 17(2):65, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02923057.

Fahrion:1976:VBS

[385] R. Fahrion. Verbesserung der Bandbreite für Spektral–Fensterfunktionen. (German) []. Statistical Papers, 17(2):66–80, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02923058.

Harsaae:1976:SPU

[386] Erik Harsaae. Some pitfalls in the use of minimum chi square. Statistical Papers, 17(2):81–104, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02923059.

Schaich:1976:SMK

[387] E. Schaich. Schätzfehler bei Markov– Ketten. (German) []. Statistical Papers, 17(2):105–130, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923060.

Kuczynski:1976:BGG

[388] Thomas Kuczynski. Bemerkung zum Gesetz der großen Zahlen in der Wirtschafts-und Sozialstatistik. (German) []. Statistical Papers, 17 (2):131–135, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923061.

Ehrenberg:1976:FVS

[389] A. S. C. Ehrenberg. Fit versus simplicity. Statistical Papers, 17(2):136–145, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923062.

Anderson:1976:NAG

[390] O. D. Anderson. A note on the asymptotic generalised variance for a moving average process. Statistical Papers, 17(2):146–151, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923063.

Anonymous:1976:HCb

[391] Anonymous. Help & contacts. Statistical Papers, 17(2):??, June 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1976:DAc

[392] Anonymous. Zu dieser ausgabe. Statistical Papers, 17(3):153, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923070.

Samanta:1976:ENP

[393] M. Samanta. Efficient non-parametric estimation in the analysis of variance. Statistical Papers, 17(3):154– 172, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923071.

Maass:1976:FIS

[394] Siegfried Maaß and Hugo Ch. Rieß. Zur Fehlerfortpflanzung im statischen offenen Leontief–Modell: Schwankungsbereiche für die Größe Gesamtproduktion. (German) []. Statistical Papers, 17 (3):173–193, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923072.

Menges:1976:HMGa

[395] Günter Menges, Bernd Heilig, and Reiner Zwer. Das Heidelberger Modell. (German) [The Heidelberg model]. Statistical Papers, 17(3):194–204, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923073.

Haq:1976:NSD

[396] M. Safiul Haq and V. Ming Ng. A note on strutural distribution of the intra-class correlation coefficient. Statistical Papers, 17(3):205–210, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923074.

Leiner:1976:TDG

[397] Bernd Leiner. Transferfunktionen von Differenzenfiltern und gleitenden Durchschnitten. (German) []. Statistical Papers, 17(3):211–222, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923075.

Birkenfeld:1976:KRS

[398] Wolfgang Birkenfeld. Kritische reflexionen zu einer simulationsstudie. Statistical Papers, 17(3):223–225, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923076.

Leiner:1976:RBS

[399] Bernd Leiner. Replik zu Birkenfelds Stellungnahme betreffend “Kritische Reflexionen zu einer Simulationsstudie”. (German) []. Statistical Papers, 17(3):226–227, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02923077.

Anonymous:1976:HCc

[400] Anonymous. Help & contacts. Statistical Papers, 17(3):??, September 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1976:DAG

[401] Anonymous. Zu dieser Ausgabe. (German) []. Statistical Papers, 17 (4):229, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02923034.

Menges:1976:HMGb

[402] Günter Menges, Reiner Zwer, and Bernd Heilig. Das Heidelberger Modell. (German) [The Heidelberg model]. Statistical Papers, 17(4):230–266, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923035.

Vogt:1976:CDT

[403] Arthur Vogt. Charakterisierung dreier Typen von statistischen Verhältniszahlen. (German) []. Statistical Papers, 17 (4):267–277, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923036.

Harsaae:1976:IPR

[404] E. Harsaae. Informative probability and its role in the probabilistic arguments implied by Rudolph Carnap’s “Inductive probability” and Bayes’ formula. Statistical Papers, 17(4): 278–284, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02923037.

Anderson:1976:TRT

[405] O. D. Anderson. On the transformation of raw time series data: A review. Statistical Papers, 17(4):285–289, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923038.

Gottinger:1976:UOS

[406] Hans W. Gottinger. On uniqueness of an optimal search rule. Statistical Papers, 17(4):290–294, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923039.

Ramanarayanan:1976:RAS

[407] R. Ramanarayanan. Reliability analysis of the 2-out-of-k: F. System with spares considering exchange time. Statistical Papers, 17(4):295–298, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02923040.

Anonymous:1976:HCd

[408] Anonymous. Help & contacts. Statistical Papers, 17(4):??, December 1976. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1977:DAa

[409] Anonymous. Zu dieser ausgabe. Statistical Papers, 18(1):1, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932900.

Bernholtz:1977:TCS

[410] B. Bernholtz. Type I censoring and the structural approach. Statistical Papers, 18(1):2–12, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932901.

Ringwald:1977:LAE

[411] Karl Ringwald. Linear aggregation and errors in variables. Statistical Papers, 18(1):13–25, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932902.

Pfuff:1977:VKE

[412] Franz Pfuff. Ein vereinfachtes Konstruktionsverfahren für endlichstufige Tests. (German) []. Statistical Papers, 18(1):26–40, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932903.

Buning:1977:FVZ

[413] H. Büning. Film oder Vorlesung? Zum ZDF–Studienprogramm “Statistik im Medienverbund”. (German) [Film or lecture? On the ZDF study program “Statistics in the Media Network”]. Statistical Papers, 18(1):41–45, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932904.

Garagorry:1977:CSR

[414] Fernando L. Garagorry and Mohammad Ahsanullah. A characterization of stationary renewal processes and of memoryless distributions. Statistical Papers, 18(1):46–48, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932905.

Anderson:1977:FNS

[415] O. D. Anderson. A further note on the stationarity and invertibility restraints on the parameters of mixed autoregressive moving average processes. Statistical Papers, 18(1):49–52, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932906.

Sankaranarayanan:1977:PSW

[416] G. Sankaranarayanan and R. Ramanarayanan. On a power series whose co-efficients are convolutions of a stable distribution of a positive random variable. Statistical Papers, 18(1):53–57, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932907.

Sherif:1977:AWT

[417] S. A. Sherif. Die Anwendung der “Wrocław-Taxonomy” auf Input– Output-Tabellen. (German) [The use of the “Wrocław-Taxonomy” in input– output tables]. Statistical Papers, 18(1):58–64, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.

1007/BF02932908. See remarks [451].

Anonymous:1977:HCa

[418] Anonymous. Help & contacts. Statistical Papers, 18(1):??, March 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1977:DAb

[419] Anonymous. Zu dieser ausgabe. Statistical Papers, 18(2):69, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932741.

Kofler:1977:TAC

[420] Edward Kofler. Taking into account confidence measures in the valuation of statistical inferences. Statistical Papers, 18(2):70–82, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932742.

Menges:1977:ICI

[421] Günter Menges and Sherif A. Sherif. International comparison of industrial production structures: Application of a taxonomic method to the input-output tables of the Economic Commission for Europe. Statistical Papers, 18(2):83–122, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932743.

Bennett:1977:MCV

[422] B. M. Bennett. On multivariate coefficients of variation. Statistical Papers, 18(2):123–128, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932744.

Scobey:1977:CEM

[423] P. Scobey and D. G. Kabe. On a criterion for extrapolation in multivariate normal regression. Statistical Papers, 18(2):129–132, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932745.

Kohle:1977:EGF

[424] D. Köhle. Einige Grundprobleme der Faktorenanalyse. (German) []. Statistical Papers, 18(2):133–141, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932746.

Hamerle:1977:SAL

[425] A. Hamerle and H. Pape. Über einen stochastischen Ansatz zur Lösung von Klassifikationsproblemen. (German) []. Statistical Papers, 18(2):142–146, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932747.

Vogt:1977:AHM

[426] A. Vogt. Zur Anwendung des harmonischen Mittels. (German) []. Statistical Papers, 18(2):147–151, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932748.

Anonymous:1977:HCb

[427] Anonymous. Help & contacts. Statistical Papers, 18(2):??, June 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1977:DAc

[428] Anonymous. Zu dieser ausgabe. Statistical Papers, 18(3):153, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02933106.

Heiler:1977:SRD

[429] Siegfried Heiler. Some recent developments in the analysis of component models for economic time series. Statistical Papers, 18(3):154–180, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02933107.

Clement:1977:BEM

[430] B. Clement and N. Giri. Bayesian estimation of means in a three component hierarchical design with random effect model. Statistical Papers, 18(3): 181–192, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02933108.

Thoni:1977:ASM

[431] H. Thöni. Zur Anwendung statistischer Methoden bei der Auswertung medizinischer Versuche: Eine Kritik an der Kritik. (German) []. Statistical Papers, 18(3):193–197, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02933109.

Schriever:1977:FAS

[432] K.-H Schriever. Zur falschen Anwendung statistischer Methoden bei der Auswertung medizinisch-wissenschaftlicher Versuche: Eine Zurückweisung der Thöni’schen Kritik. (German) []. Statistical Papers, 18(3):198–202, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02933110.

Friedmann:1977:TED

[433] Ralph Friedmann. Trendbereinigung mit ersten Differenzen — Eine Klarstellung. (German) []. Statistical Papers, 18(3):203–208, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02933111.

Dickmann:1977:ETG

[434] Heinrich Dickmann. Ein einfacher Test auf Gleichverteilung. (German) []. Statistical Papers, 18(3):209–214, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02933112.

Anonymous:1977:HCc

[435] Anonymous. Help & contacts. Statistical Papers, 18(3):??, September 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1977:DAd

[436] Anonymous. Zu dieser ausgabe. Statistical Papers, 18(4):217, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

com/accesspage/article/10.1007/
BF02932781.

Schneeweiss:1977:KBK

[437] H. Schneeweiß. Kritische Bemerkungen zur Kritik am Wahrscheinlichkeitssubjektivismus. (German) []. Statistical Papers, 18(4):218–232, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932782.

Seifert:1977:MPE

[438] H. G. Seifert. Multicollinearity and the prediction error. Statistical Papers, 18 (4):233–253, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932783.

Bomsdorf:1977:PCD

[439] E. Bomsdorf. The prize-competition-distribution a particular L-distribution as a supplement to the Pareto distribution. Statistical Papers, 18(4): 254–264, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932784.

vNatzmer:1977:DGM

[440] W. v. Natzmer. Ein Diskussionsbeitrag zu “G. Menges / R. Zwer / B. Heilig: Das Heidelberger Modell, Ein zweiter Bericht, insbesondere über die wirtschaftsstatistischen Grundlagen” [1]. (German) [A discussion paper on “G. Menges / R. Zwer / B. Heilig: The Heidelberg model, a second report”, in particular on economic statistical basics]. Statistical Papers, 18(4):265–277, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932785.

Leiner:1977:EBD

[441] B. Leiner. Einige Bemerkungen zu den Definitionsgleichungen des Heidelberger Modells. (German) [Some remarks on the defining equations of the Heidelberg model]. Statistical Papers, 18(4):278–280, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932786.

Heiler:1977:WGK

[442] S. Heiler. Wahrscheinlichkeitstheoretische Grundlagen für Komponentenmodelle beiökonomischen Zeitreihen. (German) []. Statistical Papers, 18(4): 281–286, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932787.

Vogt:1977:VPM

[443] A. Vogt. Eine Verallgemeinerung des Preis- und Mengenindexproblems. (German) []. Statistical Papers, 18 (4):287–297, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932788.

Kofler:1977:SUV

[444] E. Kofler and G. Menges. Die Strukturierung von Unbestimmtheiten und eine Verallgemeinerung des Axiomensystems von Kolmogoroff für unbestimmte Wahrscheinlichkeiten. (German) []. Statistical Papers, 18(4): 298–302, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932789.

Joiner:1977:CIS

[445] B. L. Joiner. 1976 current index to statistics. applications, methods and theory. Statistical Papers, 18 (4):303, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02932790.

Anonymous:1977:HCd

[446] Anonymous. Help & contacts. Statistical Papers, 18(4):??, December 1977. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1978:DAa

[447] Anonymous. Zu dieser ausgabe. Statistical Papers, 19(1):1, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932757.

Bellhouse:1978:MLM

[448] D. R. Bellhouse. Marginal likelihood methods for distributed lag models. Statistical Papers, 19(1):2–14, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932758.

Schlittgen:1978:OST

[449] R. Schlittgen. The optimality of some tests for medians. Statistical Papers, 19(1):15–24, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932759.

vonRothkirch:1978:ADC

[450] Christoph von Rothkirch. Axiomatic definition and comparison of three aggregation rules in cardinal social choice theory. Statistical Papers, 19(1):25–44, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932760.

Vogel:1978:EAA

[451] Friedrich Vogel. Einige Anmerkungen zur Anwendung der “Wrocław Taxonomy” auf Input–Output-Tabellen. (German) [Some remarks on the use of the “Wrocław-Taxonomy” in input– output tables]. Statistical Papers, 19(1):45–52, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932761. See [417].

Sherif:1978:HAK

[452] S. A. Sherif. Ein hierarchischagglomeratives Klassifikationsverfahren zur Anwendung auf Input–Output-Tabellen. (German) []. Statistical Papers, 19(1):53–62, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02932762.

Drygas:1978:MS

[453] Hilmar Drygas. Über multidimensionale skalierung. Statistical Papers, 19(1):63–66, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932763.

Mansfield:1978:NER

[454] Edward R. Mansfield and J. Wanzer Drane. A note on exponential regression. Statistical Papers, 19(1):67–70, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932764.

Bennett:1978:TED

[455] B. M. Bennett. On a test for equality of dependent correlation coefficients. Statistical Papers, 19(1):71–76, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932765.

Karmann:1978:TBS

[456] A. Karmann. Zur topologischen Beschreibung stochastischer Felder mit verallgemeinerten Spezifikationen. (German) []. Statistical Papers, 19 (1):77–81, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932766.

Anonymous:1978:HCa

[457] Anonymous. Help & contacts. Statistical Papers, 19(1):??, March 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1978:DAb

[458] Anonymous. Zu dieser ausgabe. Statistical Papers, 19(2):83, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932713.

Hackl:1978:RTA

[459] P. Hackl and W. Katzenbeisser. Residuals in tests for adequacy of regression relationships. Statistical Papers, 19(2):84–98, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932714.

Schweitzer:1978:LQT

[460] Walter Schweitzer. Ein Likelihood– Quotienten-Test für den Vergleich der Varianzen zweier Normalverteilungen. (German) []. Statistical Papers, 19 (2):99–113, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932715.

deGooijer:1978:IAM

[461] J. G. de Gooijer. On the inverse of the autocovariance matrix for a general mixed autoregressive moving average process. Statistical Papers, 19(2):114–123, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932716.

Behara:1978:EID

[462] M. Behara, E. Kofler, and G. Menges. Entropy and informativity in decision situations under partial information. Statistical Papers, 19(2):124–130, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932717.

Vogt:1978:WTV

[463] Arthur Vogt. Der Wertindextreue– Test und eine Vereinfachung des Indexproblems. (German) []. Statistical Papers, 19(2):131–139, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932718.

Sangmeister:1978:BRW

[464] Hartmut Sangmeister. Book review: Wolfgang Gerß: Lohnstatistik in Deutschland. Methodische, rechtliche und organisatorische Grundlagen seit der Mitte des 19. Jahrhunderts. (German) []. Statistical Papers, 19 (2):140–143, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932719.

Anonymous:1978:HCb

[465] Anonymous. Help & contacts. Statistical Papers, 19(2):??, June 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1978:DAc

[466] Anonymous. Zu dieser ausgabe. Statistical Papers, 19(3):145, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932799.

Maass:1978:NMB

[467] Siegfried Maaß. Neuere Methoden zur Berechnung von Kaufkraftparitäten mit einem Preisvergleich für ausgewählte Städte der Bundesrepublik Deutschland. (German) []. Statistical Papers, 19(3):146–166, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932800.

Gottstein:1978:ESN

[468] Doris Gottstein. The European System of National Accounts in comparison with the system of national accounts of the Federal Statistical Office. Statistical Papers, 19(3):167–196, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932801.

Bury:1978:SRE

[469] K. V. Bury. System reliability estimation for the exponential loadstrength model. Statistical Papers, 19 (3):197–203, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932802.

Drane:1978:BDQ

[470] J. Wanzer Drane, Donald B. Owen, and Guy Burton Seibert. The Burr distribution and quantal responses. Statistical Papers, 19(3):204–210, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932803.

Kirsch:1978:BHD

[471] Arnold Kirsch. Bemerkung zu H. Drygas, “Über multidimensionale Skalierung”. Statistical Papers, 19(3): 211–212, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02932804.

Latter:1978:AFI

[472] R. Latter. Announcement of formation of International Association for Statistical computing (IASC). Statistical Papers, 19(3):213–214, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932805.

Anonymous:1978:HCc

[473] Anonymous. Help & contacts. Statistical Papers, 19(3):??, September 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1978:DAd

[474] Anonymous. Zu dieser ausgabe. Statistical Papers, 19(4):217, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932720.

Schmid:1978:SCN

[475] Friedrich Schmid. Strong consistency of non-linear least squares estimators in the presence of stochastic regressors. Statistical Papers, 19(4):218–230, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932721.

Sitterberg:1978:AHK

[476] Georg Sitterberg. Zur Anwendung hierarchischer Klassifikationsverfahren. (German) []. Statistical Papers, 19 (4):231–246, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932722.

Sherif:1978:SPD

[477] A. Sherif and P. Tan. On structural predictive distribution with type II progressively censored Weibull data. Statistical Papers, 19(4):247–255, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932723.

Gottinger:1978:MDP

[478] H. W. Gottinger. A Markovian decision process with hidden states and hidden costs. Statistical Papers, 19 (4):256–261, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932724.

Hamerle:1978:MTB

[479] A. Hamerle and P. Kemény. Über ein multiples Testmodell für binäre Daten und k unabhängige Stichproben. (German) []. Statistical Papers, 19(4): 262–267, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932725.

Sangmeister:1978:EBE

[480] Hartmut Sangmeister. Einige Bemerkungen über die “Empirische Analyse der interregionalen Lohn- und Gehaltsstruktur in der Verarbeitenden Industrie der Bundesrepublik Deutschland”, von Dietmar Kühn, Schriften zu Regional- und Verkehrsproblemen in Industrie- und Entwicklungsländern, Band 21. (German) []. Statistical Papers, 19(4):268–271, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932726.

Anonymous:1978:HCd

[481] Anonymous. Help & contacts. Statistical Papers, 19(4):??, December 1978. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1979:DAa

[482] Anonymous. Zu dieser ausgabe. Statistical Papers, 20(1):1, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932767.

Bishop:1979:SDS

[483] L. Bishop, D. A. S. Fraser, and K. W. Ng. Some decompositions of spherical distributions. Statistical Papers, 20(1):2–21, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932768.

Menges:1979:AM

[484] G. Menges. Adaptive Mustererkennung. Statistical Papers, 20(1):22–38, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932769.

vonBorries:1979:ZCI

[485] Dietrich F. W. von Borries. Zwei Charakterisierungen eines “idealen” Index für drei Faktoren. (German) []. Statistical Papers, 20(1):39–53, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932770.

Funke:1979:CFI

[486] H. Funke and J. Voeller. Characterization of Fisher’s “ideal index” by three reversal tests — a note. Statistical Papers, 20(1):54–60, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932771.

Anonymous:1979:B

[487] Anonymous. Bücherbesprechung. Statistical Papers, 20(1):61–64, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932772.

Anonymous:1979:HCa

[488] Anonymous. Help & contacts. Statistical Papers, 20(1):??, March 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Kubler:1979:FTP

[489] Heiner Kübler. On the fitting of the three-parameter distributions Lognormal, Gamma, and Weibull. Statistical Papers, 20(2):68–125, June 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932451.

Brand:1979:HLU

[490] P. Brand. A Heisenberg-like uncertainty principle within the framework of the LPI-concept. Statistical Papers, 20(2):126–136, June 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932452.

Bocker:1979:FSI

[491] Franz Böcker. Das Fächerwahlverhalten von Studenten im zweidimensionalen Affinitätsmodell. (German) []. Statistical Papers, 20(2):137–146, June 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932453.

Anonymous:1979:HCb

[492] Anonymous. Help & contacts. Statistical Papers, 20(2):??, June 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1979:DAb

[493] Anonymous. Zu dieser ausgabe. Statistical Papers, 20(3):147, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/accesspage/article/10.1007/
BF02932773.

Brenner:1979:FCP

[494] David Brenner and D. A. S. Fraser. On foundations for conditional probability with statistical models — when is a class of functions a function? Statistical Papers, 20(3):148–159, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932774.

Rutsch:1979:MRS

[495] Martin Rutsch. Mortalität, Rauchen und Statistik. (German) []. Statistical Papers, 20(3):160–171, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932775.

Banerjee:1979:NDI

[496] K. S. Banerjee. A note on the divisia indices. Statistical Papers, 20(3): 172–175, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932776.

Neudecker:1979:BLU

[497] H. Neudecker. Best linear unbiased estimation of β subject to linear equality constraints in the general linear model. Statistical Papers, 20(3): 176–182, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932777.

Schmidt:1979:OTP

[498] H. Schmidt. Ein optimaler Test für den Parameter der Exponentialverteilung und andere Lebensdauerverteilungen. (German) []. Statistical Papers, 20(3): 183–190, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932778.

Pflaumer:1979:OML

[499] Peter Pflaumer. Ein ökonometrisches Modell zur Lohn–Preis-Dynamik. (German) []. Statistical Papers, 20(3): 191–196, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932779.

Behara:1979:SPE

[500] M. Behara and J. M. S. Chawla. Shannon and polynomial entropies as semivaluations on lattices. Statistical Papers, 20(3):197–203, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932780.

Anonymous:1979:HCc

[501] Anonymous. Help & contacts. Statistical Papers, 20(3):??, September 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1979:DAc

[502] Anonymous. Zu dieser ausgabe. Statistical Papers, 20(4):205, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/accesspage/article/10.1007/
BF02932791.

Marfels:1979:SAS

[503] Christian Marfels. Structural aspects of small business in the Canadian economy. Statistical Papers, 20(4): 206–236, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932792.

Kofler:1979:UMI

[504] Eduard Kofler and Günter Menges. Über unscharfe Mengen (nicht im Sinne Zadehs). Statistical Papers, 20(4): 237–249, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932793.

Schneeberger:1979:PFO

[505] H. Schneeberger and W. Goller. On the problem of the feasibility of optimal stratification points according to Dalenius. Statistical Papers, 20(4): 250–256, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932794.

Pflaumer:1979:AEK

[506] Peter Pflaumer. Zur approximativen Errechnung von Konfidenzintervallen für eine unbekannte Wahrscheinlichkeit bei kleinem Stichprobenumfang. (German) []. Statistical Papers, 20(4): 257–260, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932795.

Banerjee:1979:IFI

[507] K. S. Banerjee. An interpretation of the factorial indexes in the light of divisia integral indexes. Statistical Papers, 20(4):261–269, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932796.

Bocker:1979:ESW

[508] Franz Böcker. Empirische Studienfachkombinationen — Eine wahrscheinlichkeitstheoretische analyse. (German) []. Statistical Papers, 20(4): 270–275, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932797.

Mohr:1979:PZR

[509] Walter Mohr. Prognoseuntersuchung für die Zeitreihe der registrierten Arbeitslosen in der BRD. (German) []. Statistical Papers, 20(4):276–283, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932798.

Anonymous:1979:HCd

[510] Anonymous. Help & contacts. Statistical Papers, 20(4):??, December 1979. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1980:DAa

[511] Anonymous. Zu dieser ausgabe. Statistical Papers, 21(1):1, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932806.

Kuss:1980:SMD

[512] Uwe Kuß. Ein Schätzverfahren von Mences und Diehl und die Maximum Probability–Methode. (German) []. Statistical Papers, 21(1):2–13, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932807.

Lee:1980:ESP

[513] Kwan R. Lee, C. H. Kapadia, and Dwight B. Brock. On estimating the scale parameter of the Rayleigh distribution from doubly censored samples. Statistical Papers, 21(1):14–29, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932808.

Hadi:1980:BKLa

[514] Fathi Abdel Hadi, Patricia Pietsch, Christoph V. Rothkirch, and Hartmut Sangmeister. Ein Beitrag zur Klassifikation von Ländern nach ihrem Entwicklungsstand. (German) []. Statistical Papers, 21(1):30–48, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932809.

Menges:1980:BKL

[515] Günter Menges. Bemerkungen zur Klassifikation von Ländern (Vorwort zu “Ein Beitrag zur Klassifikation von Ländern nach ihrem Entwicklungsstand”). (German) []. Statistical Papers, 21(1):49–52, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932810.

Jockel:1980:AOV

[516] Karl-Heinz Jöckel and Peter Pflaumer. Die Anwendung ökonometrischer Verfahren bei der Risikoanalyse für Investitionsentscheidungen. (German) []. Statistical Papers, 21(1):53–60, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932811.

Schneeberger:1980:LFS

[517] H. Schneeberger and D. Drefahl. Limits of feasible sampling fractions in optimal stratification. Statistical Papers, 21(1):61–65, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932812.

Vogt:1980:ZFF

[518] Arthur Vogt. Der Zeit- und der Faktorumkehrteest als “finders of tests”. (German) []. Statistical Papers, 21 (1):66–71, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932813.

Anonymous:1980:HCa

[519] Anonymous. Help & contacts. Statistical Papers, 21(1):??, March 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1980:DAb

[520] Anonymous. Zu dieser ausgabe. Statistical Papers, 21(2):73–74, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932727.

Hadi:1980:BKLb

[521] Fathi Abdel Hadi, Patricia Pietsch, Christoph von Rothkirch, and Hartmut Sangmeister. Ein Beitrag zur Klassifikation von Ländern nach ihrem Entwicklungsstand. (German) []. Statistical Papers, 21(2):75–109, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932728.

Balk:1980:SCC

[522] Bert M. Balk. Seasonal commodities and the construction of annual and monthly price indexes. Statistical Papers, 21(2):110–116, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932729.

Sato:1980:NIU

[523] Kazuo Sato. The “natural” index and utility functions. Statistical Papers, 21(2):117–126, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932730.

Sato:1980:PRT

[524] Kazuo Sato. The price reversal test and economic indexes. Statistical Papers, 21(2):127–130, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932731.

Funke:1980:TCS

[525] H. Funke and J. Voeller. Three comments on Sato’s paper: The price reversal test and economic indices. Statistical Papers, 21(2):131–132, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932732.

vonBorries:1980:GII

[526] Dietrich F. W. von Borries. A “geometric” interpretation of the index formula of Siegel. Statistical Papers, 21(2):133–139, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932733.

vonBorries:1980:ZCIa

[527] Dietrich F. W. von Borries. Zwei Charakterisierungen eines “idealen” Index für drei Faktoren. (German) []. Statistical Papers, 21(2):139, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932734.

Schmid:1980:PMS

[528] Friedrich Schmid. Über ein Problem der mehrdimensionalen Skalierung. (German) []. Statistical Papers, 21 (2):140–144, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932735.

Hild:1980:EVS

[529] Claus Hild. Über einige verzerrte Schätzer für die Koeffizienten eines linearen Modells. (German) []. Statistical Papers, 21(2):145–155, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932736.

Bennett:1980:FND

[530] B. M. Bennett. A further note on the distribution of generalized coefficients of variation. Statistical Papers, 21(2):156–159, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932737.

Kofler:1980:SPI

[531] E. Kofler, G. Menges, R. Fahrion, S. Huschens, and U. Kuß. Stochastische Partielle Information (SPI). (German) []. Statistical Papers, 21(2):160–167, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932738.

Kuss:1980:ASE

[532] Uwe Kuß. Ein allgemeines statistischentscheidungstheoretisches Modell als Konsequenz der Ätialität und der Forderung nach weicher Modelibildung. (German) []. Statistical Papers, 21 (2):168–173, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932739.

Anonymous:1980:AMV

[533] Anonymous. Zur Anwendung makroökonometrischer Verfahren in Wirtschaft und Politik. (German) []. Statistical Papers, 21 (2):174–178, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932740.

Anonymous:1980:HCb

[534] Anonymous. Help & contacts. Statistical Papers, 21(2):??, June 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1980:DAc

[535] Anonymous. Zu dieser ausgabe. Statistical Papers, 21(3):181, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932613.

Menges:1980:ASB

[536] Günter Menges. Adaptive Statistik (Bemerkungen über neuere Bestrebungen in der statistischen Methodologie). (German) []. Statistical Papers, 21(3): 182–208, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932614.

vonBorries:1980:ZCIb

[537] Dietrich F. W. von Borries. Zwei Charakterisierungen eines “idealen” Index für vier Faktoren. (German) []. Statistical Papers, 21(3):209–223, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932615.

Fraser:1980:SRC

[538] D. A. S. Fraser and Philip Mc-Dunnough. Some remarks on conditional and unconditional inference for location-scale models. Statistical Papers, 21(3):224–231, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932616.

Lorenzen:1980:SBM

[539] Gunter Lorenzen. Spezifikationsfehler beim Messen der Abhängigkeit in Kontingenztabellen. (German) []. Statistical Papers, 21(3):232–238, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932617.

Neudecker:1980:BQU

[540] Heinz Neudecker. Best quadratic unbiased estimation of the variance matrix in normal regression. Statistical Papers, 21(3):239–243, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932618.

Anonymous:1980:HCc

[541] Anonymous. Help & contacts. Statistical Papers, 21(3):??, September 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1980:DAd

[542] Anonymous. Zu dieser ausgabe. Statistical Papers, 21(4):245, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932884.

Kofler:1980:FSN

[543] Eduard Kofler and Günter Menges. Fuzzy sets and non-stochastic linear partial information. Statistical Papers, 21(4):246–260, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932885.

Kuss:1980:COE

[544] Uwe Kuß. C-optimale entscheidungen. Statistical Papers, 21(4):261–279, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932886.

Fahrion:1980:KIB

[545] R. Fahrion. Über eine Klasse von Informationsmaßen für die Bewertung stochastischer (partieller) Informationen. (German) []. Statistical Papers, 21 (4):280–295, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932887.

Brenner:1980:IDF

[546] David Brenner and D. A. S. Fraser. The identification of distribution form. Statistical Papers, 21(4):296–304, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932888.

Kramer:1980:PIL

[547] Walter Krämer. Parameterschätzungen im linearen Regressionsmodell bei fehlspezifizierten Störgrößenprozessen. (German) []. Statistical Papers, 21 (4):305–314, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932889.

Sankaranarayanan:1980:RAC

[548] G. Sankaranarayanan and R. Ramanarayanan. Reliability and availability of 1-out-of-n:G cold standby systems with arbitrary failure rate and constant repair rate. Statistical Papers, 21(4):315–320, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932890.

Anonymous:1980:HCd

[549] Anonymous. Help & contacts. Statistical Papers, 21(4):??, December 1980. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1981:DAa

[550] Anonymous. Zu dieser ausgabe. Statistical Papers, 22(1):1, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932814.

Bott:1981:AEG

[551] Dietrich Bott. Adäquationsproze und Entscheidungsproblem. (German) [the adequations process and decision problem]. Statistical Papers, 22(1):2–24, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932815.

Katzenbeisser:1981:TGR

[552] Walter Katzenbeisser. Test auf Gleichheit von Regressionskoeffizienten: Einige Erweiterungen. (German) []. Statistical Papers, 22(1):25–39, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932816.

Banerjee:1981:RFA

[553] Kali S. Banerjee. A review of the factorial approach providing the true index of cost of living. Statistical Papers, 22(1):40–57, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932817.

Koebel:1981:NDP

[554] Michael Koebel. A note on a dynamic probability distribution. Statistical Papers, 22(1):58–66, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932818.

vonBorries:1981:IPK

[555] Dietrich F. W. von Borries. Zur Interpretation des “Preisumkehrtests” — Eine kritische Anmerkung. (German)

[]. Statistical Papers, 22(1):67–71, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932819.

Balk:1981:SMC

[556] Bert M. Balk. A simple method for constructing price indices for seasonal commodities. Statistical Papers, 22(1):72–78, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932820.

Anonymous:1981:HCa

[557] Anonymous. Help & contacts. Statistical Papers, 22(1):??, March 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1964:DA

[558] Anonymous. Zu dieser ausgabe. Statistical Papers, 22(2):81, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02933545.

Kawasaki:1964:MRL

[559] Seiichi Kawasaki and Klaus F. Zimmermann. Measuring relationships in the log-linear probability model by some compact measures of association. Statistical Papers, 22(2):82–109, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02933546.

Thiel:1964:ERN

[560] Norbert Thiel. The effect of a reduction of the number of exogenous variables in a linear regression equation on some statistical measures. Statistical Papers, 22(2):110–120, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02933547.

Ahsanullah:1964:RVE

[561] Mohammad Ahsanullah. Record values of exponentially distributed random variables. Statistical Papers, 22 (2):121–127, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02933548.

Precht:1964:BSR

[562] Manfred Precht and H. Redinger. Best subsets regression using L_{P} - norms with 1 ≤ p < ∞. Statistical Papers, 22(2):128–134, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02933549.

Jockel:1964:EBI

[563] Kari-Heinz Jöckel and Peter Pflaumer. Einige Bemerkungen zur “iterativmultiplen” regression. (German) []. Statistical Papers, 22(2):135–141, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02933550.

Vogt:1964:ZFE

[564] Arthur Vogt. Die Zeit- und die Faktorantithesen von Eigenschaften von Indices. (German) []. Statistical Papers, 22(2):142–143, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02933551.

Menges:1964:AAG

[565] Günter Menges. Ätialität und Adäquation. (German) []. Statistical Papers, 22(2):144–149, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02933552.

Anonymous:1964:HCb

[566] Anonymous. Help & contacts. Statistical Papers, 22(2):??, June 1964. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1981:DAb

[567] Anonymous. Zu dieser ausgabe. Statistical Papers, 22(3):151, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02934640.

Assenmacher:1981:EWO

[568] Walter Assenmacher and Günther E. Braun. Das Einfachheitspostulat in Wissenschaftstheorie und Ökonometrie. (German) []. Statistical Papers, 22(3): 152–175, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02934641.

Lorenzi:1981:K

[569] René Lorenzi. Korrespondenzanalyse. Statistical Papers, 22(3):176– 194, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02934642.

Brenner:1981:STS

[570] David Brenner and D. A. S. Fraser. A simplification of the traditional statistical model in the presence of symmetry. Statistical Papers, 22(3): 195–206, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02934643.

Abel:1981:ARA

[571] Peter Abel and Günter Menges. Aktualisierung und Ridge–Analyse des Heidelberger Modells. (German) []. Statistical Papers, 22(3):207–230, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02934644.

Brenner:1981:TIS

[572] David Brenner, D. A. S. Fraser, and G. Monette. Theories of inference or simple additives. Statistical Papers, 22(3):231–233, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02934645.

Kocklauner:1981:RKS

[573] Gerhard Kockläuner. Restringierte Koeffizientenschätzung in Spline–Lag-Modellen. Statistical Papers, 22(3): 234–240, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02934646.

Vogt:1981:CIE

[574] Arthur Vogt. Characterizations of indices, especially of the Stuvel and the Banerjee index. Statistical Papers, 22 (3):241–245, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02934647.

Anonymous:1981:HCb

[575] Anonymous. Help & contacts. Statistical Papers, 22(3):??, September 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Knepel:1981:MUV

[576] Helmut Knepel. Modelle mit unbeobachtbaren Variablen — Der PLS– Ansatz. (German) []. Statistical Papers, 22(4):248–279, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02932468.

Scholing:1981:HI

[577] Eberhard Scholing. Die hierarchische interaktionsanalyse. Statistical Papers, 22(4):280–315, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932469.

Ahsanullah:1981:CED

[578] M. Ahsanullah. On characterization of the exponential distribution by spacings. Statistical Papers, 22(4): 316–320, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932470.

Lorenzen:1981:AZS

[579] Gunter Lorenzen. Abschätzungen für die Zeilen- und Spaltensummen der Leontief–Inversen bei unvollständiger Kenntnis der Matrix der inter-industriellen Lieferungen. (German) []. Statistical Papers, 22(4): 321–328, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932471.

Strecker:1981:MSE

[580] Heinrich Strecker. Mean Square Error und Qualität statistischer Daten in Erhebungen. (German) []. Statistical Papers, 22(4):329–333, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932472.

Menges:1981:SFG

[581] Günter Menges and Christel Menges. Statistik und Fertilität. (German) []. Statistical Papers, 22(4):334–344, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932473.

Anonymous:1981:HCc

[582] Anonymous. Help & contacts. Statistical Papers, 22(4):??, December 1981. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1982:DAa

[583] Anonymous. Zu dieser ausgabe. Statistical Papers, 23(1):1, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932699.

Kugler:1982:TRV

[584] Peter Kugler. Testing real variables exogeneity: Some empirical results for six European countries. Statistical Papers, 23(1):2–11, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932700.

Danzer:1982:FBE

[585] Klaus Danzer and Hans-Dieter Heike. Formulierung des Bankensektors in einem ökonometrischen Modell mit Hilfe der optimalen Kontrollrechnung. (German) []. Statistical Papers, 23 (1):12–25, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932701.

Als:1982:PPS

[586] G. Als. Priorités et programme statistiques dans un petit pays — le cas de Luxembourg. (French) [Priorities and statistical program in a small country — the case of Luxembourg].

Statistical Papers, 23(1):26–38, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932702.

Bennett:1982:NTH

[587] B. M. Bennett. Note on a test of the hypothesis of symmetry in multivariate distributions. Statistical Papers, 23(1):39–43, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932703.

Polasek:1982:LRD

[588] Wolfgang Polasek. Local resistance in distributed lag models. Statistical Papers, 23(1):44–51, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932704.

vanderPutten:1982:SRA

[589] R. van der Putten and H. Neudecker. Some reflections on Almon estimators. Statistical Papers, 23(1):52–56, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932705.

Magg:1982:EUE

[590] Alexander Magg. Einfache Überlegungen zu einer entscheidungstheoretischen Fundierung von Stichprobenverfahren. (German) []. Statistical Papers, 23 (1):57–61, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932706.

Menges:1982:WFS

[591] G. Menges and Ch. Menges. Der World Fertility Survey auf der 43. Sitzung des Internationalen Statistischen Instituts. (German) []. Statistical Papers, 23(1):62, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02932707.

Anonymous:1982:HCa

[592] Anonymous. Help & contacts. Statistical Papers, 23(1):??, March 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1982:DAb

[593] Anonymous. Zu dieser ausgabe. Statistical Papers, 23(2):65, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932821.

Ehemann:1982:SEL

[594] Klaus Ehemann. Ein statistisches entscheidungsproblem mit linearer partieller information. Statistical Papers, 23(2):67–83, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932822.

Menges:1982:ESAa

[595] Günter Menges. Über einige statistische aspekte globaler gesellschaftlicher probleme. Statistical Papers, 23 (2):84–105, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932823.

Brachinger:1982:ASP

[596] H.-W Brachinger. Eine Anmerkung zur Stochastischen Partiellen Information und dem Max e min-Entscheidungsprinzip. (German) []. Statistical Papers, 23(2):106–109, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932824.

Evans:1982:WMO

[597] Michael Evans, Donald A. S. Fraser, and Hélène Massam. The Weibull model, objective form, and linear analysis. Statistical Papers, 23(2):110–115, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932825.

Giri:1982:NCP

[598] N. Giri and O. Mouqadem. Numerical comparison of power functions of invariant tests for means with covariates. Statistical Papers, 23(2):116–121, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932826.

Drane:1982:DTD

[599] J. W. Drane and T. A. Hua. Decomposing three dimensional contingency tables. Statistical Papers, 23(2):122–127, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932827.

Singh:1982:NIP

[600] G. Singh. A note on inclusion probability proportional to size sampling. Statistical Papers, 23(2):128–130, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02932828. See

remark [625].

Keel:1982:CBH

[601] Alex Keel. Zur computermäßigen Berechnung hypergeometrischer Wahrscheinscheinlichkeiten. (German) [On regular computer calculation of hypergeometric probabilities]. Statistical Papers, 23(2):131–133, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932829.

Brenner:1982:MAS

[602] David Brenner, Donald A. S. Fraser, Günter Menges, and Erika Rost. Model analysis with structural and stochastic partial information. Statistical Papers, 23(2):134–141, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932830.

Sangmeister:1982:API

[603] Hartmut Sangmeister. Aktuelle Probleme der internationalen Wirtschaftsund Sozialstatistik. (German) []. Statistical Papers, 23(2):142–149, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932831.

Anonymous:1982:HCb

[604] Anonymous. Help & contacts. Statistical Papers, 23(2):??, June 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Ronning:1982:CVT

[605] Gerd Ronning. Characteristic values and triangular factorization of the covariance matrix for multinomial, Dirichlet and multivariate hypergeometric distributions and some related results. Statistical Papers, 23(3): 152–176, September 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02933049.

Menges:1982:ESAb

[606] Günter Menges. Über einige statistische aspekte globaler gesellschaftlicher probleme. Statistical Papers, 23(3): 177–217, September 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02933050.

Haq:1982:SRP

[607] M. Safiul Haq. Structural relations and prediction for the multivariate models. Statistical Papers, 23(3): 218–227, September 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02933051.

Schneeberger:1982:GPO

[608] H. Schneeberger and D. Drefahl. Gain in precision by optimum stratification and optimum allocation in dependence on the sampling fraction. Statistical Papers, 23(3):228–237, September 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02933052.

Anonymous:1982:HCc

[609] Anonymous. Help & contacts. Statistical Papers, 23(3):??, September 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bomsdorf:1982:AGK

[610] Eckart Bomsdorf. Zur Abschätzung des Ginikoeffizienten bei klassierten Daten. (German) []. Statistical Papers, 23 (4):240–257, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932444.

Kugler:1982:SRC

[611] Peter Kugler. Some remarks on causality detection by autoregressive modelling. Statistical Papers, 23(4): 258–274, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932445.

Mundlos:1982:APF

[612] Bernd Mundlos. Eine Axiomatik für Preisindizes in Funktionalform — gleichzeitig eine Kritik der Kritik an DIVISIA–Indizes. (German) [An axiomatic theory of price indices in funktional form — simultaneously a critique of the critique of DIVISIA-indices]. Statistical Papers, 23(4):275–290, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932446.

Pauly:1982:ZAO

[613] Ralf Pauly. Zerlegung und Analyse ökonomischer Zeitreihen. (German) []. Statistical Papers, 23(4):291–303, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932447.

Schuler:1982:OHB

[614] Klaus Schüler. Optimale Hyperkorridore– Begründung und Darstelung eines empiriebezogenen ökonometrischen Schätz- und Prognoseverfahrens. (German) []. Statistical Papers, 23(4): 304–325, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932448.

Ahsanullah:1982:CED

[615] Mohammad Ahsanullah. Characterizations of the exponential distribution by some properties of the record values. Statistical Papers, 23(4):326–332, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932449.

Keel:1982:VEF

[616] Alex Keel. Eine Verallgemeinerung des exakten Fisherschen Unabhängigkeitstests von 2 auf r dichotome Merkmale. (German) []. Statistical Papers, 23(4): 333–341, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932450.

Anonymous:1982:HCd

[617] Anonymous. Help & contacts. Statistical Papers, 23(4):??, December 1982. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Herausgeber:1983:G

[618] Die Herausgeber. Geleitwort. Statistical Papers, 24(1):1, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932485.

Grohmann:1983:GM

[619] Heinz Grohmann. Günter Menges. Statistical Papers, 24(1):3–5, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932486.

Brenner:1983:MTI

[620] D. Brenner, D. A. S. Fraser, and G. Monette. On models and theories of inference; structural or pivotal analysis. Statistical Papers, 24(1):7–19, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932487.

Pauly:1983:OVV

[621] R. Pauly. Offene versus vollständige Modellbildung in der Ökonometrie. (German) []. Statistical Papers, 24 (1):21–45, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932488.

Buning:1983:AVT

[622] Herbert Büning. Adaptive verteilungsfreie tests. Statistical Papers, 24 (1):47–67, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932489.

Balk:1983:NTF

[623] B. M. Balk. A note on the true factorial price index. Statistical Papers, 24(1):69–72, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932490.

Spreemann:1983:LCR

[624] K. Spreemann and G. Bamberg. Local consistency of risk preferences. Statistical Papers, 24(1):73–83, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932491.

Gabler:1983:RGS

[625] S. Gabler. A remark to G. Singh’s paper: A note on inclusion probability proportional to size sampling. Statistical Papers, 24(1):85–86, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932492. See [600].

Ferschl:1983:BRE

[626] F. Ferschl. Book review: Esenwein-Rothe, Ingeborg: Einführung in die Demographie. Bevölkerungsstruktur und Bevölkerungsprozeß aus der Sicht der Statistik. (German) []. Statistical Papers, 24(1):87–91, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932493.

Lorenzen:1983:GIB

[627] Gunter Lorenzen. Güterverteilungsmatrizen und ihre Bewertung. (German) []. Statistical Papers, 24(1):93–119, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932494.

Bohning:1983:MLE

[628] Dankmar Böhning. Maximum likelihood estimation of the logarithmic series distribution. Statistical Papers, 24(1):121–140, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932495.

Banerjee:1983:EIM

[629] K. S. Banerjee. On the existence of infinitely many ideal log-change index numbers associated with the CES preference ordering. Statistical Papers, 24(1):141–148, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932496.

Ahmad:1983:TBI

[630] M. Ahmad and N. C. Giri. A test of bivariate independence with additional data. Statistical Papers, 24(1): 149–154, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932497.

Buhler:1983:ARV

[631] Erhard Bühler. Eine Anwendung der Rayleigh–Verteilung auf ein Problem der Photographie. (German) []. Statistical Papers, 24(1):155–159, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932498.

Sangmeister:1983:IBR

[632] Hartmut Sangmeister. International Bank for Reconstruction and Development / The World Bank: World development report 1982. Statistical Papers, 24(1):161–166, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932499.

Anonymous:1983:CECa

[633] Anonymous. Calendar of events and call for papers. Statistical Papers, 24 (1):167–170, December 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932500.

Anonymous:1983:HCa

[634] Anonymous. Help & contacts. Statistical Papers, 24(1):??, December 1983.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Schlicht:1983:SAS

[635] Ekkehart Schlicht. Seasonal adjustment in a stochastic model. Statistical Papers, 25(1):1–12, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932388.

Friedrich:1983:IGS

[636] Dieter Friedrich. Interpolation, Glättung und Saisonbereinigung von Zeitreihen mit Splinefunktionen. (German) []. Statistical Papers, 25(1):13–51, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932389.

Bennett:1983:TRE

[637] B. M. Bennett. Tests for relative error in repeated multivariate samples. Statistical Papers, 25(1):53–60, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932390.

Brenner:1983:IDF

[638] D. Brenner, M. Evans, D. A. S. Fraser, H. Massam, and E. Rost. The identification of distribution form II. Statistical Papers, 25(1):61–68, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932391.

Carmody:1983:FMQ

[639] T. J. Carmody, R. L. Eubank, and V. N. LaRiccia. A family of minimum quantile distance estimators for the three-parameter Weibull distribution. Statistical Papers, 25(1):69–82, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932392.

Anonymous:1983:CECb

[640] Anonymous. Calendar of events and call for papers. Statistical Papers, 25(1):83–86, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932393.

Palm:1983:STS

[641] F. C. Palm and J. M. Sneek. Significance tests and spurious correlation in regression models with autocorrelated errors. Statistical Papers, 25(1):87–105, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932394.

Pauly:1983:SAA

[642] R. Pauly. Statistische Analyse ausgewählter Schätzer im kontemporär korrelierten Regressionsmodell mit prädeterminiertem verzögerten Regressor und autokorrelierten Fehlern. (German) []. Statistical Papers, 25 (1):107–134, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932395.

Kramer:1983:HCA

[643] Walter Krämer. High correlation among errors and the efficiency of ordinary least squares in linear models. Statistical Papers, 25(1):135–142, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932396.

Banerjee:1983:SOL

[644] K. S. Banerjee. Some observations on the log-change index numbers and their connection with the factorial indexes. Statistical Papers, 25 (1):143–158, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932397.

Pfuff:1983:BSE

[645] Franz Pfuff. Bayeslösungen bei sequentiellen Experimenten. (German) []. Statistical Papers, 25(1):159–170, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932398.

Anonymous:1983:CECc

[646] Anonymous. Calendar of events and call for papers. Statistical Papers, 25 (1):171–174, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932399.

Anonymous:1983:PSA

[647] Anonymous. Policy for the submission, acceptance and publication of algorithms. Statistical Papers, 25 (1):175–179, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932400.

Ariyawansa:1983:SIP

[648] K. A. Ariyawansa and J. G. C. Templeton. Structural inference on the parameter of the Rayleigh distribution from doubly censored samples. Statistical Papers, 25(1):181–199, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932401.

Brunner:1983:RVZ

[649] E. Brunner and N. Neumann. Rangtests für das verbundene Zwei–Stichprobenproblem mit fehlenden Meßwerten. (German) []. Statistical Papers, 25(1):201–210, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932402.

Hag:1983:SEC

[650] M. Safiul Hag and A. K. Md. E. Saleh. On the structural estimation of certain correlation coefficients. Statistical Papers, 25(1):211–218, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932403.

Huschens:1983:WEL

[651] S. Huschens. Wahrscheinlichkeitstransformationen in Entscheidungsmodellen bei linearer partieller Information. (German) []. Statistical Papers, 25 (1):219–230, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932404.

Kramer:1983:GMF

[652] W. Krämer. On the game of maximising R̄^{2}: A further comment: A further comment. Statistical Papers, 25(1):231–233, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932405.

Gleissner:1983:SRF

[653] W. Gleissner. Some recursion formulae for the beta distribution. Statistical Papers, 25(1):235–236, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02932406.

Leiner:1983:ASI

[654] B. Leiner. Arima–Schätzungen im vergleich. Statistical Papers, 25 (1):237–244, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932407.

Anonymous:1983:HCb

[655] Anonymous. Help & contacts. Statistical Papers, 25(1):??, March 1983. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Fraser:1985:CTO

[656] D. A. S. Fraser and H. Massam. Conical tests: Observed levels of significance and confidence regions. Statistical Papers, 26(1):1–17, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932515.

Johnson:1985:STD

[657] N. L. Johnson and S. Kotz. Some tests for detection of faulty inspection. Statistical Papers, 26(1):19–29, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932516.

Abel:1985:NVB

[658] P. Abel. Ein numerisches Verfahren zur Bestimmung maxeminoptimaler Aktionen. (German) []. Statistical Papers, 26(1):31–41, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932517.

Kischka:1985:BIP

[659] P. Kischka. Bayessche Indexmodelle in der Portfoliotheorie. (German) []. Statistical Papers, 26(1):43–57, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932518.

Balk:1985:SCF

[660] B. M. Balk. A simple characterization of Fisher’s price index. Statistical Papers, 26(1):59–63, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932519.

Pfingsten:1985:NTF

[661] A. Pfingsten. A note on the true factorial price index-addendum to a note by balk. Statistical Papers, 26 (1):65–69, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932520.

Arminger:1985:B

[662] G. Arminger, U. Küsters, D. Plachky, K. A. Schäffer, and B. Leiner. Buchbesprechungen. Statistical Papers, 26 (1):71–84, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932521.

Anonymous:1985:CECa

[663] Anonymous. Calendar of events and call for papers. Statistical Papers, 26 (1):85–86, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02932522.

Baltagi:1985:SSR

[664] H. Baltagi and J. R. Ferry. Small sample results on Schmidt’s truncation remainder to the gamma distributed lag. Statistical Papers, 26(1):87–95, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932523.

Schneeberger:1985:OST

[665] H. Schneeberger and J. P. Pöllot. Optimum stratification with two variates. Statistical Papers, 26(1):97–113, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932524.

Xekalaki:1985:FME

[666] E. Xekalaki. Factorial moment estimation for the bivariate generalized Waring distribution. Statistical Papers, 26(1):115–129, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932525.

Katzenbeisser:1985:DTS

[667] W. Katzenbeisser. The distribution of two-sample location exceedance test statistics under Lehmann alternatives. Statistical Papers, 26(1):131–138, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932526.

Chou:1985:PCC

[668] Y. M. Chou and D. B. Owen. On the precision of the coverages of βcontent inner tolerance intervals. Statistical Papers, 26(1):139–146, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932527.

Baltagi:1985:BEV

[669] B. H. Baltagi. The bias of the estimated variances of least squares estimators in an error components model. Statistical Papers, 26(1):147–155, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932528.

Drygas:1985:B

[670] H. Drygas, L. Fahrmeir, B. Leiner, and H. Strasser. Buchbesprechungen. Statistical Papers, 26(1):157–166, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932529.

Anonymous:1985:CECb

[671] Anonymous. Calendar of events and call for papers. Statistical Papers, 26 (1):167–169, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932530.

Laisney:1985:SAW

[672] François Laisney and Karl Ringwald. Statistical aspects of W. D. Fisher’s method of optimal aggregation. Statistical Papers, 26(1):171–197, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932531.

Meinischmidt:1985:FIK

[673] Gerhard Meinischmidt. Faktorenanalyse als ein Instrument zur Konstruktion von Preisniveauindizes. (German) []. Statistical Papers, 26(1): 199–210, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932532.

Krumbholz:1985:UVE

[674] W. Krumbholz and C. D. Ludwig. Unverfälschte Variablenprüfpläne für exponentialverteilte Merkmale mit zweiseitigen Toleranzgrenzen. (German) []. Statistical Papers, 26(1):211–223, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932533.

Likes:1985:DCV

[675] Jiří Likeš, CSc. Distributions of certain variables in samples from twoparameter exponential distribution. Statistical Papers, 26(1):225–236, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932534.

Anonymous:1985:HC

[676] Anonymous. Help & contacts. Statistical Papers, 26(1):??, December 1985. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Galler:1986:UIE

[677] Heinz P. Galler. Übergangsratenmodelle bei intervalldatierten Ereignissen. (German) []. Statistical Papers, 27 (1):1–22, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932552.

Kreienbrock:1986:AEA

[678] L. Kreienbrock. Zur Auswirkung von Endlichkeitskorrekturen bei der Analyse einfacher Zufallsstichproben. (German) []. Statistical Papers, 27 (1):23–35, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932553.

Waldmann:1986:BDR

[679] K. H. Waldmann. Bounds to the distribution of the run length in general quality-control schemes. Statistical Papers, 27(1):37–56, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932554.

Sondermann:1986:BAS

[680] Dieter Sondermann. Best approximate solutions to matrix equations under rank restrictions. Statistical Papers, 27(1):57–66, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932555.

Schader:1986:DFP

[681] M. Schader and F. Schmid. Distribution function and percentage points for the central and noncentral F - distribution. Statistical Papers, 27 (1):67–74, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932556.

Ferschl:1986:BRS

[682] F. Ferschl. Book review: Schneeweiß H. und Strecker, H. (Hrsg): Contributions to econometrics and statistics today. In memoriam Günter Menges. Statistical Papers, 27(1):75–80, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932557.

Anonymous:1986:CEC

[683] Anonymous. Calendar of events and call for papers. Statistical Papers, 27 (1):81–82, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02932558.

Skala:1986:CTH

[684] Heinz J. Skala. On σ-coherence and a theorem of Heath and Sudderth. Statistical Papers, 27(1):83–88, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932559.

vanYzeren:1986:FII

[685] J. van Yzeren. Fisher’s ideal index numbers as natural divisia results. Statistical Papers, 27(1):89–99, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932560.

Fuchs-Seliger:1986:RBE

[686] Susanne Fuchs-Seliger. Relatioinships between economic and statistical price indices. Statistical Papers, 27(1): 101–115, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932561.

Ariyawansa:1986:SIP

[687] K. A. Ariyawansa and J. G. C. Tempelton. Structural inference for parameters of a power function distribution. Statistical Papers, 27(1):117–139, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932562.

Bohning:1986:FSC

[688] D. Böhning and F. P. Schelp. A Fortran subroutine for computing indicators of the nutritional status of children and adolescents. Statistical Papers, 27(1):141–150, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932563.

Heyer:1986:B

[689] H. Heyer, H. Schneeweiß, A. Tobik, F. Baur, and L. Knüsel. Buchbesprechungen. Statistical Papers, 27 (1):151–164, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932564.

Rothe:1986:SRB

[690] Günter Rothe. Some remarks on bootstrap techniques for constructing confidence intervals. Statistical Papers, 27(1):165–172, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932565.

Linhart:1986:FSS

[691] H. Linhart and W. Zucchini. Finite sample selection criteria for multinomial models. Statistical Papers, 27 (1):173–178, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932566.

Fahrmeir:1986:AID

[692] Ludwig Fahrmeir and Heinz Kaufmann. Asymptotic inference in discrete response models. Statistical Papers, 27 (1):179–205, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932567.

Hamerle:1986:RAD

[693] Alfred Hamerle. Regression analysis for discrete event history or failure time data. Statistical Papers, 27 (1):207–225, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932568.

Perez:1986:UEI

[694] R. Pérez, C. Caso, and M. A. Gil. Unbiased estimation of income inequality. Statistical Papers, 27(1):227–237, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932569.

Arnold:1986:CAE

[695] B. F. Arnold. Comparison of the approximate and exact optimum economic design of control charts basing on the sign test. Statistical Papers, 27(1):239–241, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932570.

Hudec:1986:B

[696] M. Hudec, H. Strasser, F. Ferschl, and G. Uebe. Buchbesprechungen. Statistical Papers, 27(1):243–252, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932571.

Anonymous:1986:HC

[697] Anonymous. Help & contacts. Statistical Papers, 27(1):??, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wolters:1987:OMZ

[698] Jürgen Wolters. Ökonometrische Modelle bei zeitreihendaten versus multivariate Zeitreihenmodelle–Eine Übersicht. (German) []. Statistical Papers, 28(1):1–25, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932588.

Fraser:1987:SPS

[699] D. A. S. Fraser. Sequential parameter structure, conditional inference, and likelihood drop. Statistical Papers, 28(1):27–52, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932589.

Singh:1987:SIR

[700] Balvir Singh and Yogendra P. Chaubey. On some improved ridge estimators. Statistical Papers, 28(1):53–67, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932590.

Drygas:1987:MRC

[701] Hilmar Drygas. On the multivariate Rao–Cramér inequality. Statistical Papers, 28(1):69–71, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932591.

Haagen:1987:B

[702] K. Haagen, R. D. Reiss, O. Krafft, and H. Schmidbauer. Buchbesprechungen. Statistical Papers, 28(1):73–78, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932592.

Anonymous:1987:CE

[703] Anonymous. Calendar of events. Statistical Papers, 28(1):79–80, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02932593.

Fahrmeir:1987:ALI

[704] Ludwig Fahrmeir. Asymptotic likelihood inference for nonhomogeneous observations. Statistical Papers, 28 (1):81–116, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932594.

Chou:1987:SSS

[705] Youn-Min Chou and Gary M. Johnson. Sample sizes for strong two-sided distribution-free tolerance limits. Statistical Papers, 28(1):117–131, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932595.

Hebbel:1987:TSD

[706] Hartmut Hebbel and Siegfried Heiler. Trend and seasonal decomposition in discrete time. Statistical Papers, 28 (1):133–158, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932596.

Rieder:1987:B

[707] H. Rieder and A. Ungerer. Buchbesprechungen. Statistical Papers, 28 (1):159–161, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932597.

Fraser:1987:FAT

[708] D. A. S. Fraser. Fibre analysis and tangent models. Statistical Papers, 28(1):163–181, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932598.

Schneeweiss:1987:ACA

[709] H. Schneeweiß, D. A. Sprott, and R. Viveros. An approximate conditional analysis of the linear functional relationship. Statistical Papers, 28(1): 183–202, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932599.

Alzaid:1987:NNC

[710] A. A. Alzaid, A. N. Ahmed, and M. Al-Osh. On the NBAFR (NWAFR) class of life distributions. Statistical Papers, 28(1):203–216, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932600.

Bozdogan:1987:ALR

[711] Hamparsum Bozdogan and Donald E. Ramirez. An adjusted likelihoodratio algorithm for the Behrens–Fisher problem. Statistical Papers, 28(1): 217–231, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932601.

Grothe:1987:MGP

[712] Holger Grothe. Matrix generators for pseudo-random vector generation. Statistical Papers, 28(1):233–238, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02932602.

Pfanzagl:1987:B

[713] J. Pfanzagl, W. Wefelmeyer, G. Pflug, H. Drygas, and W. Krämer. Buchbesprechungen. Statistical Papers, 28(1): 239–243, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02932603.

Anonymous:1987:CEC

[714] Anonymous. Calendar of events and call for papers. Statistical Papers, 28 (1):245–246, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02932604.

Schmitz:1987:MST

[715] N. Schmitz. Minimax sequential tests of composite hypotheses on the drift of a Wiener process. Statistical Papers, 28(1):247–261, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02932605.

Kodde:1987:TSL

[716] David A. Kodde and Franz C. Palm. Testing the stability of a linear dynamic model. Statistical Papers, 28 (1):263–270, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932606.

Bossert:1987:CTR

[717] Walter Bossert and Andreas Pfingsten. The circular and time reversal tests reconsidered in economic price index theory. Statistical Papers, 28 (1):271–284, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02932607.

Anonymous:1987:HC

[718] Anonymous. Help & contacts. Statistical Papers, 28(1):??, December 1987. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Anonymous:1988:E

[719] Anonymous. Editorial. Statistical Papers, 29(1):1, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02924507.

Schneider:1988:AUK

[720] Wolfgang Schneider. Analytical uses of Kalman filtering in econometrics — a survey. Statistical Papers, 29 (1):3–33, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924508.

Alzaid:1988:MRL

[721] Abdulhamid A. Alzaid. Mean residual life ordering. Statistical Papers, 29 (1):35–43, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924509.

Chen:1988:TPP

[722] Laxiang Chen and Jürgen Eichenauer. Two point priors and Γ-minimax estimating in families of uniform distributions. Statistical Papers, 29(1): 45–57, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924510.

Ferentinos:1988:SCI

[723] K. K. Ferentinos. On shortest confidence intervals and their relation with uniformly minimum variance unbiased estimators. Statistical Papers, 29(1):59–75, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924511.

Anonymous:1988:BR

[724] Anonymous. Books received. Statistical Papers, 29(1):76–77, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02924512.

Schal:1988:BR

[725] M. Schäl, H. Stenger, H. Drygas, Lütkepohl, and K. Jacobs. Book reviews. Statistical Papers, 29(1): 77–82, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924513.

Anonymous:1988:CEC

[726] Anonymous. Calendar of events and call for papers. Statistical Papers, 29 (1):83–84, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02924514.

Brachinger:1988:MPI

[727] Hans Wolfgang Brachinger and Renate Schubert. Measuring price-induced changes in standard of living in least developed countries: economic validity of Laspeyres price indices and some further suggestions. Statistical Papers, 29(1):85–112, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924515.

Zmyslony:1988:AEP

[728] Roman Zmyślony and Hilmar Drygas. On admissible estimation for parametric functions in linear models. Statistical Papers, 29(1):113–123, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924516.

Jones:1988:BSM

[729] P. W. Jones and S. A. Madhi. Bayesian sequential methods for choosing the best multinomial cell: some simulation results. Statistical Papers, 29(1): 125–132, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924517.

Katzenbeisser:1988:JDR

[730] Walter Katzenbeisser. On the joint distribution of the random variables ‘number of inversions’ and ‘number of outstanding variables’ in a randomly arranged sequence. Statistical Papers, 29(1):133–141, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924518.

Ahmed:1988:PPM

[731] Abdul-Hadi N. Ahmed. Preservation properties for the mean residual life ordering. Statistical Papers, 29(1): 143–150, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924519.

Ahsanullah:1988:CKK

[732] M. Ahsanullah. On a conjecture of Kakosyan, Klebanov and Melamed. Statistical Papers, 29(1):151–157, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924520.

Schneeweiss:1988:BR

[733] H. Schneeweiß, E. v. Collani, I. Klein, U. Kockelkorn, and K. Weichselberger. Book reviews. Statistical Papers, 29(1):159–168, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924521.

Miller:1988:E

[734] Rupert G. Miller. Erratum. Statistical Papers, 29(1):168, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02924522.

Palm:1988:CWC

[735] F. C. Palm and D. A. Kodde. Computing Wald criteria for nested hypotheses. Statistical Papers, 29(1):169–190, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924523.

Gil:1988:OEL

[736] M. A. Gil and M. R. Casals. An operative extension of the likelihood ratio test from fuzzy data. Statistical Papers, 29(1):191–203, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924524.

Bausch:1988:OSE

[737] Thomas Bausch. Optimized sampling by exchange methods. Statistical Papers, 29(1):205–218, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924525.

Schlittgen:1988:DSM

[738] Rainer Schlittgen. On the determination of sizes of markets. Statistical Papers, 29(1):219–225, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924526.

Anonymous:1988:HC

[739] Anonymous. Help & contacts. Statistical Papers, 29(1):??, December 1988. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Fusek:1989:CES

[740] I. Fusek and L. Fusková. A combined estimator in the simple errorsin-variables model. Statistical Papers, 30(1):1–15, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924304.

Rothe:1989:BGL

[741] Günter Rothe. Bootstrap for generalized linear models. Statistical Papers, 30(1):17–26, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924305.

Koziol:1989:MTN

[742] James A. Koziol. Multivariate tests for non-additivity. Statistical Papers, 30 (1):27–37, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924306.

Wingo:1989:LTW

[743] Dallas R. Wingo. The left-truncated Weibull distribution: theory and computation. Statistical Papers, 30(1): 39–48, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924307.

Bomsdorf:1989:MDG

[744] Eckart Bomsdorf. Measurement of disparity from grouped data with different degrees of information. Statistical Papers, 30(1):49–60, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924308.

Lorenzen:1989:LRL

[745] Gunter Lorenzen. Log-ratios and the logarithmic mean. Statistical Papers, 30(1):61–75, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924309.

Ferschl:1989:BR

[746] F. Ferschl and K. W. Gaede. Book reviews. Statistical Papers, 30(1): 76–80, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924310.

Anonymous:1989:CEC

[747] Anonymous. Calendar of events and call for papers. Statistical Papers, 30(1):81, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02924311.

Kusters:1989:HMC

[748] Ulrich Küsters. Hierarchical mean and covariance structure models. Statistical Papers, 30(1):83–104, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924312.

Diebold:1989:SSP

[749] Francis X. Diebold and Peter Pauly. Small sample properties of asymptotically equivalent tests for autoregressive conditional heteroskedasticity. Statistical Papers, 30(1):105–131, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924313.

Anonymous:1989:BR

[750] Anonymous. Books received. Statistical Papers, 30(1):132, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02924314.

Mittag:1989:EPS

[751] H.-J. Mittag. Estimating parameters in a simple errors-in-variables model: a new approach base on finite sample distribution theory. Statistical Papers, 30(1):133–140, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924315.

Balakrishnan:1989:RCO

[752] N. Balakrishnan. A relation for the covariances of order statistics from n independent and non-identically distributed random variables. Statistical Papers, 30(1):141–146, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924316.

Nagaraja:1989:SCG

[753] H. N. Nagaraja, Pali Sen, and R. C. Srivastava. Some characterizations of geometric tail distributions based on record values. Statistical Papers, 30 (1):147–155, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924317.

Krug:1989:BR

[754] W. Krug, R. Pauly, and H. Lütkepohl. Book reviews. Statistical Papers, 30 (1):156–162, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02924318.

Dielman:1989:SSP

[755] Terry E. Dielman and Roger C. Pfaffenberger. Small sample properties of estimators in the autocorrelated error model: a review and some additional simulations. Statistical Papers, 30(1):163–183, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924319.

Alzaid:1989:MRL

[756] A. A. Alzaid. Mean residual life ordering. Statistical Papers, 30(1):184, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02924320.

Breitung:1989:BRR

[757] K. Breitung. Book review: Ripley, B. D., Stochastic simulation. Statistical Papers, 30(1):184, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02924321.

Brannas:1989:HTM

[758] Kurt Brännäs and Thomas Laitila. Heteroskedasticity in the tobit model. Statistical Papers, 30(1):185–196, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02924322.

Linhart:1989:DS

[759] H. Linhart. Discrete smoothing. Statistical Papers, 30(1):197–211, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924323.

Anonymous:1989:HC

[760] Anonymous. Help & contacts. Statistical Papers, 30(1):??, December 1989. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Baltagi:1990:ECR

[761] Badi H. Baltagi. The error components regression model: conditional relative efficiency comparisons. Statistical Papers, 31(1):1–13, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924669.

Chandra:1990:LSE

[762] K. Suresh Chandra and P. Dhanavanthan. Least squares estimation of the coefficients of a partially explosive model, with polynomial regressions of same degree, generating a pair of related time series. Statistical Papers, 31(1):15–31, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924670.

Weiss:1990:ADP

[763] Günter Weiß. An algorithm for discriminating populations with unequal variance matrices. Statistical Papers, 31(1):33–39, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924671.

Cepar:1990:SAB

[764] Drago Cepar and Zoran Radalj. Some asymptotic behaviour of the bootstrap estimates on a finite sample. Statistical Papers, 31(1):41–46, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924672.

Huschens:1990:NSS

[765] Stefan Huschens. Necessary sample sizes for categorial data. Statistical Papers, 31(1):47–53, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924673.

Fleischer:1990:SSU

[766] Karlheinz Fleischer. Stratified sampling using double samples. Statistical Papers, 31(1):55–63, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924674.

Giorgi:1990:IIT

[767] Giovanni M. Giorgi and Andrea Pallini. Inequality indices: theoretical and empirical aspects of their asymptotic behaviour. Statistical Papers, 31(1): 65–76, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924675.

Behara:1990:CED

[768] M. Behara and Z. Dudek. On concave entropies of discrete systems. Statistical Papers, 31(1):77–80, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924676.

Anonymous:1990:CEC

[769] Anonymous. Calendar of events and call for papers. Statistical Papers, 31 (1):81–82, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02924677.

Fraser:1990:VCM

[770] D. A. S. Fraser. Views on conditional and marginal methods of statistical inference. Statistical Papers, 31 (1):83–93, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924678.

Arminger:1990:BRL

[771] G. Arminger. Book review: Lindsey, J.: The analysis of categorical data using GLIM. Statistical Papers, 31(1):94, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02924679.

Revesz:1990:RIR

[772] P. Révész. Regularities and irregularities in a random 0,1 sequence. Statistical Papers, 31(1):95–101, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924680.

Rendtel:1990:CSV

[773] U. Rendtel. Cusum-schemes with variable sampling intervals and sample sizes. Statistical Papers, 31(1): 103–118, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924681.

Seidel:1990:PSS

[774] W. Seidel. On the performance of a sampling scheme in statistical quality control using incomplete prior information. Statistical Papers, 31(1): 119–130, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924682.

Kusters:1990:NSM

[775] U. Küsters. A note on sequential ML estimates and their asymptotic covariances. Statistical Papers, 31(1): 131–145, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924683.

Loesgen:1990:GBI

[776] K. H. Loesgen. A generalization and Bayesian interpretation of ridge-type estimators with good prior means. Statistical Papers, 31(1):147–154, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02924684.

Huschens:1990:UBC

[777] S. Huschens. On upper bounds for the characteristic values of the covariance matrix for multinomial, Dirichlet and multivariate hypergeometric distributions. Statistical Papers, 31(1): 155–159, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924685.

Kimura:1990:BR

[778] K. Kimura, R. Göb, and H. Daduna. Book reviews. Statistical Papers, 31 (1):160–164, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924686.

Trenkler:1990:MSE

[779] G. Trenkler and H. Toutenburg. Mean squared error matrix comparisons between biased estimators — an overview of recent results. Statistical Papers, 31 (1):165–179, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02924687.

Toutenburg:1990:BRR

[780] H. Toutenburg. Book review: Rubin, D. B.: Multiple imputation for nonresponse in surveys. Statistical Papers, 31(1):180, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02924688.

Anonymous:1990:HC

[781] Anonymous. Help & contacts. Statistical Papers, 31(1):??, December 1990. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Johnson:1991:SGB

[782] N. L. Johnson and S. Kotz. Some generalizations of Bernoulli and Polya– Eggenberger contagion models. Statistical Papers, 32(1):1–17, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925473.

Viertl:1991:BRF

[783] R. Viertl. Book review: de Finetti, Bruno: Theory of probability (A critical introductory treatment). Statistical Papers, 32(1):18, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02925474.

Abouammoh:1991:SMR

[784] A. M. Abouammoh and M. I. Hendi. Shock models with renewal failure rate properties. Statistical Papers, 32 (1):19–34, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925475.

Skala:1991:COW

[785] H. J. Skala. Concerning ordered weighted averaging aggregation operators. Statistical Papers, 32(1): 35–44, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925476.

Edgeman:1991:SSP

[786] R. L. Edgeman and P. M. Salzberg. A sequential sampling plan for the inverse Gaussian mean. Statistical Papers, 32(1):45–53, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925477.

Tutz:1991:BRK

[787] G. Tutz. Book review: Küchenhoff, H.: Logit- und Probitregression mit Fehlern in den Variablen. Mathematical Systems in Economics Vol. 117. Statistical Papers, 32(1):54, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02925478.

Hossain:1991:CSD

[788] A. Hossain and D. N. Naik. A comparative study on detection of influential observations in linear regression. Statistical Papers, 32(1):55–69, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925479.

Toutenburg:1991:BRL

[789] H. Toutenburg. Book review: Little, R. J. A. and D. B. Rubin: Statistical analysis with missing data. Statistical Papers, 32(1):70, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/accesspage/article/10.1007/
                                        BF02925480.

Kramer:1991:AUL

[790] W. Krämer. The asymptotic unbiasedness of S^{2} in the linear regression model with AR(1)-disturbances. Statistical Papers, 32(1):71–73, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925481.

Godde:1991:SGA

[791] M. Gödde. Statistical games against a prophet-proof of a minimax conjecture. Statistical Papers, 32(1):75–81, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925482.

Anonymous:1991:CECa

[792] Anonymous. Calendar of events and call for papers. Statistical Papers, 32 (1):82–83, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02925483.

Kunitz:1991:GTL

[793] H. Kunitz and H. Pamme. Graphical tools for life time data analysis. Statistical Papers, 32(1):85–113, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925484.

Viertl:1991:BTF

[794] R. Viertl and H. Hule. On Bayes’ theorem for fuzzy data. Statistical Papers, 32(1):115–122, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925485.

Kofler:1991:CFD

[795] E. Kofler and P. Zweifel. Convolution of fuzzy distributions in decision-making. Statistical Papers, 32(1):123–136, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925486.

Heyer:1991:SRW

[796] H. Heyer and T. Kusama. Some remarks on weak sufficiency. Statistical Papers, 32(1):137–154, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925487.

Wang:1991:ALE

[797] L. Wang. Admissible linear estimators of the multivariate normal mean without extra information. Statistical Papers, 32(1):155–165, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925488.

Jain:1991:CCI

[798] S. Jain. Comparison of confidence intervals of traffic intensity for M/E_{k}/1 queueing systems. Statistical Papers, 32(1):167–174, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925489.

Skala:1991:BRK

[799] H. J. Skala. Book review: Krishnaiah, P. R. and C. R. Rao (eds): Sampling. Handbook of Statistics, vol. 6. Statistical Papers, 32(1):175–176, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02925490.

Anonymous:1991:CECb

[800] Anonymous. Calendar of events and call for papers. Statistical Papers, 32 (1):177–178, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02925491.

Lutkepohl:1991:MMS

[801] H. Lütkepohl and B. Theilen. Measures of multivariate skewness and kurtosis for tests of nonnormality. Statistical Papers, 32(1):179–193, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925492.

Anonymous:1991:HC

[802] Anonymous. Help & contacts. Statistical Papers, 32(1):??, December 1991. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Kirchgassner:1992:ITA

[803] G. Kirchgässner and J. Wolters. Implications of temporal aggregation on the relation between two time series. Statistical Papers, 33(1):1–19, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925307.

Johnsson:1992:PSR

[804] T. Johnsson. A procedure for stepwise regression analysis. Statistical Papers, 33(1):21–29, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925308.

Strasser:1992:BRC

[805] H. Strasser. Book review: Le Cam, L, and G. L. Yang: Asymptotics in statistics: Some basic concepts. Statistical Papers, 33(1):30–32, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925309.

Md-Yusof:1992:CEP

[806] S. Md-Yusof and S. E. Rigdon. A comparison of estimators for the proportion in the tail of a normal distribution. Statistical Papers, 33(1):33–38, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925310.

Tu:1992:ESS

[807] D. Tu and L. Zhang. On the estimation of skewness of a statistic using the jackknife and the bootstrap. Statistical Papers, 33(1):39–56, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925311.

Mahmoud:1992:SIP

[808] M. Mahmoud and M. S. Maswadah. Structural inference on the parameters of the Pareto distribution from complete and censored life test data. Statistical Papers, 33(1):57–68, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925312.

Hurlimann:1992:POM

[809] W. Hürlimann. On parameter orthogonality to the mean. Statistical Papers, 33(1):69–74, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925313.

Gleissner:1992:PPI

[810] W. Gleißner. A parameterisation of the price indices characterized by the Eichhorn–Voeller axioms. Statistical Papers, 33(1):75–82, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925314.

Chandra:1992:MSA

[811] K. Suresh Chandra and S. Sampath. Markov sampling with auxiliary information. Statistical Papers, 33(1): 83–91, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925315.

Anonymous:1992:CECa

[812] Anonymous. Calendar of events and call for papers. Statistical Papers, 33 (1):93–94, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02925316.

Strasser:1992:CMS

[813] H. Strasser. Concentration of multivariate statistical tables. Statistical Papers, 33(1):95–117, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925317.

Homburg:1992:CSA

[814] Ch. Homburg and A. Dobratz. Covariance structure analysis via specification searches. Statistical Papers, 33 (1):119–142, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925318.

Wesolowski:1992:CBE

[815] J. Wesolowski. A characterization of the bivariate elliptically contoured distribution. Statistical Papers, 33(1): 143–149, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925319.

Bohm:1992:TAM

[816] W. Böhm. A transient analysis of M/G/1 queues with N-policy. Statistical Papers, 33(1):151–157, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925320.

Paparoditis:1992:MLT

[817] E. Paparoditis. Modelling long-term dependence in measurement errors of plutonium concentration. Statistical Papers, 33(1):159–170, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925321.

vCollani:1992:EMA

[818] E. v. Collani and Ch. Weigand. Economic machine adjustment in the case of product screening. Statistical Papers, 33(1):171–184, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925322.

Anonymous:1992:CECb

[819] Anonymous. Calendar of events and call for papers. Statistical Papers, 33 (1):185–186, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02925323.

Greenwich:1992:UHR

[820] M. Greenwich. A unimodal hazard rate function and its failure distribution. Statistical Papers, 33(1):187–202, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925324.

Balasubramanian:1992:EOT

[821] K. Balasubramanian and N. Balakrishnan. Estimation for one- and two-parameter exponential distributions under multiple type-II censoring.

Statistical Papers, 33(1):203–216, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925325.

Khan:1992:IAP

[822] S. Khan. Inference about the parameters of a bi-variate simultaneous equation model: structural approach. Statistical Papers, 33(1):217–225, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925326.

Anonymous:1992:HC

[823] Anonymous. Help & contacts. Statistical Papers, 33(1):??, December 1992. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Pal:1993:END

[824] N. Pal. Estimating the normal dispersion matrix and the precision matrix from a decision-theoretic point of view: a review. Statistical Papers, 34(1):1–26, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925524.

Srivastava:1993:POL

[825] V. K. Srivastava and M. Dube. Properties of the ordinary least squares and stein-rule predictions in linear regression models with proxy variables. Statistical Papers, 34(1):27–41, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925525.

Schneider:1993:BDC

[826] H. Schneider. Binary discrete choice under asymmetric restrictions. Statistical Papers, 34(1):43–57, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925526.

Paul:1993:SRM

[827] V. Paul and B. Chandrasekar. Some results for multidimensional stationary independent increment processes. Statistical Papers, 34(1):59–65, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925527.

Hassler:1993:URT

[828] U. Hassler. Unit root tests: the autoregressive approach in comparison with the periodogram regression. Statistical Papers, 34(1):67–82, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925528.

Liu:1993:SNC

[829] P. C. Liu and J. Praschnik. The size of the nonstationary component and its effect on tests for unit roots. Statistical Papers, 34(1):83–88, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925529.

Sprott:1993:RES

[830] D. A. Sprott and V. T. Farewell. Randomization in experimental science.

Statistical Papers, 34(1):89–94, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925530.

Anonymous:1993:CECa

[831] Anonymous. Calendar of events and call for papers. Statistical Papers, 34 (1):95–96, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02925531.

Ferguson:1993:ADM

[832] T. S. Ferguson. On the asymptotic distribution of max and mex. Statistical Papers, 34(1):97–111, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925532.

Nowak:1993:BRS

[833] E. Nowak. Book review: Singer H.: Zeitkontinuierliche Dynamische Systeme. Studienbücher zur qualitativen und quantitativen Wirtschaftsund Sozialforschung, Band 3. Statistical Papers, 34(1):112, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02925533.

Schabe:1993:NEI

[834] H. Schäbe. Nonparametric estimation of intensities of nonhomogeneous Poisson processes. Statistical Papers, 34(1):113–131, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925534.

Fung:1993:QBA

[835] W. K. Fung. A quasi-Bayesian analysis of regression outliers using Akaike’s predictive likelihood. Statistical Papers, 34(1):133–141, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925535.

Huang:1993:DA

[836] M. L. Huang and K. Y. Fung. Ddistribution and its applications. Statistical Papers, 34(1):143–159, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925536.

Prabhu-Ajgaonkar:1993:NEO

[837] S. G. Prabhu-Ajgaonkar. Nonexistence of an optimum estimator in a class of ratio estimators. Statistical Papers, 34(1):161–165, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925537.

Ahmed:1993:CGI

[838] A. N. Ahmed and A. M. Abouammoh. Characterizations of gamma, inverse Gaussian, and negative binomial distributions via their length-biased distributions. Statistical Papers, 34(1): 167–173, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925538.

Targhetta:1993:SCD

[839] M. L. Targhetta. Some characterizations of distributions of the exponential-type. Statistical Papers, 34 (1):175–180, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925539.

Polasek:1993:JME

[840] W. Polasek. Jambu, m.: Explorative datenanalyse. Statistical Papers, 34 (1):181–186, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925540.

Anonymous:1993:CECb

[841] Anonymous. Calendar of events and call for papers. Statistical Papers, 34 (1):187–188, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02925541.

Schmid:1993:GCP

[842] F. Schmid. A general class of poverty measures. Statistical Papers, 34(1): 189–211, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925542.

Fraser:1993:DTS

[843] D. A. S. Fraser. Directional tests and statistical frames. Statistical Papers, 34 (1):213–236, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925543.

Anonymous:1993:HC

[844] Anonymous. Help & contacts. Statistical Papers, 34(1):??, December 1993. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hassler:1994:SAF

[845] U. Hassler. The sample autocorrelation function of I(1) processes. Statistical Papers, 35(1):1–16, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926395.

Herrmann:1994:ADB

[846] E. Herrmann. Asymptotic distribution of bandwidth selectors in kernel regression estimation. Statistical Papers, 35(1):17–26, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926396.

Kahle:1994:SCR

[847] W. Kahle. Simultaneous confidence regions for the parameters of damage processes. Statistical Papers, 35 (1):27–41, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02926397.

Knautz:1994:NCB

[848] Henning Knautz and Götz Trenkler. A note on the correlation between S^{2} and the least squares estimator in the linear regression model. Statistical Papers, 35(1):42, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926398.

Eichenauer-Herrmann:1994:GMR

[849] J. Eichenauer-Herrmann, K. Ickstadt, and E. Weiß. Gamma-minimax results for the class of unimodal priors. Statistical Papers, 35(1):43–56, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926399.

Nassrallah:1994:CLP

[850] B. Nassrallah and A. K. Md. Ehsanes Saleh. Comparison of location parameters of two exponential distributions when scale parameters are different and unknown. Statistical Papers, 35(1):57–69, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926400.

Chandrasekar:1994:NOV

[851] B. Chandrasekar and T. Edwin Prabakaran. A note on optimal vector unbiased predictor. Statistical Papers, 35(1):71–80, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926401.

Iwase:1994:EPL

[852] K. Iwase and K. Kanefuji. Estimation for 3-parameter lognormal distribution with unknown shifted origin.

Statistical Papers, 35(1):81–90, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926402.

Baltagi:1994:SRE

[853] Badi H. Baltagi and Qi Li. A simple recursive estimation method for linear regression models with AR(p) disturbances. Statistical Papers, 35 (1):93–100, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926403.

Srivastava:1994:AST

[854] V. K. Srivastava and H. Toutenburg. Application of Stein-type estimation in combining regression estimates from replicated experiments. Statistical Papers, 35(1):101–112, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926404.

Singhal:1994:ENN

[855] R. A. Singhal and H. Sahai. Effects of non-normality on the power function in a one-way random model. Statistical Papers, 35(1):113–125, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926405.

Khan:1994:ETR

[856] S. Khan. β-expectation tolerance region for the heteroscedastic multiple regression model with multivariate Student-t error. Statistical Papers, 35(1):127–138, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926406.

Ariyawansa:1994:EUM

[857] K. A. Ariyawansa. On the existence and uniqueness of maximizers of two likelihood functions. Statistical Papers, 35(1):139–150, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926407.

Abouammoh:1994:CVB

[858] A. M. Abouammoh, A. M. Ali, and A. F. Mashhour. On characterizations and variance bounds of discrete α-unimodality. Statistical Papers, 35 (1):151–161, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926408.

Ohtani:1994:RBP

[859] K. Ohtani. Risk behavior of a pretest estimator for normal variance with the Stein-type estimator. Statistical Papers, 35(1):163–168, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926409.

Gupta:1994:BLA

[860] S. S. Gupta and K. J. Miescke. Bayesian look ahead one stage sampling allocations for selecting the largest normal mean. Statistical Papers, 35(1):169–177, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02926410.

Weinhardt:1994:EPI

[861] Ch. Weinhardt. The efficiency of price income situations, the real average income — a characterization. Statistical Papers, 35(1):179–186, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926411.

Anonymous:1994:EN

[862] Anonymous. Editors’ note. Statistical Papers, 35(1):187, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02926412.

Czado:1994:PLM

[863] C. Czado. Parametric link modification of both tails in binary regression. Statistical Papers, 35(1):189–201, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926413.

Dietz:1994:ALD

[864] E. Dietz and D. Böhning. Analysis of longitudinal data using a finite mixture model. Statistical Papers, 35(1): 203–210, December 1994. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02926414.

Anonymous:1994:HC

[865] Anonymous. Help & contacts. Statistical Papers, 35(1):??, December 1994.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Menendez:1995:ABS

[866] M. L. Menéndez, D. Morales, L. Pardo, and M. Salicrú. Asymptotic behaviour and statistical applications of divergence measures in multinomial populations: a unified study. Statistical Papers, 36(1):1–29, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926015.

Ferguson:1995:CSB

[867] T. S. Ferguson. A class of symmetric bivariate uniform distributions. Statistical Papers, 36(1):31–40, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926016.

Kuchenhoff:1995:ILR

[868] H. Küchenhoff. The identification of logistic regression models with errors in the variables. Statistical Papers, 36(1):41–47, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926017.

Wong:1995:AIT

[869] A. Wong. On approximate inference for the two-parameter gamma model. Statistical Papers, 36(1):49–59, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926018.

Dalabehera:1995:ESA

[870] M. Dalabehera and L. N. Sahoo. Efficiencies of six almost unbiased ratio estimators under a particular model. Statistical Papers, 36(1):61–67, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926019.

Pordzik:1995:RLH

[871] P. R. Pordzik. On the robustness of the linear hypothesis test procedure in the singular linear model with implied restrictions. Statistical Papers, 36(1):69–75, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926020.

Breitung:1995:MST

[872] J. Breitung. Modified stationarity tests with improved power in small samples. Statistical Papers, 36(1):77–95, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926021.

Arnold:1995:Q

[873] B. F. Arnold. Qualitätsregelkarten. Statistical Papers, 36(1):96, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02926022.

Anonymous:1995:CECa

[874] Anonymous. Calendar of events and call for papers. Statistical Papers, 36 (1):97–98, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926023.

Ohtani:1995:GRR

[875] K. Ohtani. Generalized ridge regression estimators under the LINEX loss function. Statistical Papers, 36(1): 99–110, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926024.

Schmid:1995:RLS

[876] W. Schmid. On the run length of a Shewhart chart for correlated data. Statistical Papers, 36(1):111–130, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926025.

Schabe:1995:KEC

[877] H. Schäbe and J. Tiedge. Kernel estimation for characteristics of pure jump processes. Statistical Papers, 36(1): 131–144, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926026.

El-Arishy:1995:URB

[878] S. El-Arishy. Useful relationship between the log-survival function and truncated moments, with applications. Statistical Papers, 36(1):145–154, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926027.

Rukhin:1995:RPM

[879] A. L. Rukhin and J. Shi. Recursive procedures for multiple decisions: finite time memory and stepwise maximum likelihood procedure. Statistical Papers, 36(1):155–162, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926028.

Nollau:1995:IVS

[880] V. Nollau. Inequalities for variances of some functions of random variables. Statistical Papers, 36(1):163–174, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926029.

Guven:1995:APM

[881] B. Güven. Asymptotic properties of maximum likelihood estimation in the mixed analysis of variance model. Statistical Papers, 36(1):175–182, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926030.

Infante:1995:DED

[882] A. M. Infante. The difficulties of estimation of dispersion parameters in linear models — an illustration. Statistical Papers, 36(1):183–189, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926031.

Trenkler:1995:LMA

[883] G. Trenkler. Lineare modelle. Algebraische grundlagen und statistische anwendungen. (German) []. Statistical Papers, 36(1):190, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02926032.

Anonymous:1995:CECb

[884] Anonymous. Calendar of events and call for papers. Statistical Papers, 36 (1):191–192, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926033.

Wolters:1995:TSI

[885] J. Wolters. On the term structure of interest rates — empirical results for Germany. Statistical Papers, 36(1):193–214, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926034.

Anonymous:1995:HC

[886] Anonymous. Help & contacts. Statistical Papers, 36(1):??, December 1995. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Draper:1996:ODE

[887] N. R. Draper and F. Pukelsheim. An overview of design of experiments. Statistical Papers, 37(1):1–32, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926157.

Franses:1996:PIF

[888] Philip Hans Franses and Richard Paap. Periodic integration: further results on model selection and forecasting. Statistical Papers, 37(1):33–52, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926158.

Pakes:1996:CLBa

[889] Anthony G. Pakes, T. Sapatinas, and E. B. Fosam. Characterizations, length-biasing, and infinite divisibility. Statistical Papers, 37(1):53–69, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926159.

Dominguez:1996:CLP

[890] J. Armando Dominguez and D. A. Sprott. Comparison of location parameters of two exponential distributions. Statistical Papers, 37(1):71–78, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926160.

Chaudhuri:1996:ADM

[891] Arijit Chaudhuri and Tapabrata Maiti. Asymptotic design-cum-model based estimation of variances of estimated linear regression coefficients in survey sampling with unequal probabilities. Statistical Papers, 37(1):79–84, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926161.

Naik:1996:RME

[892] V. D. Naik and P. C. Gupta. On regression method for estimating a population proportion. Statistical Papers, 37(1):85–92, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926162.

Anonymous:1996:CECa

[893] Anonymous. Calendar of events and call for papers. Statistical Papers, 37(1):93–94, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926163.

Anonymous:1996:HCa

[894] Anonymous. Help & contacts. Statistical Papers, 37(1):??, March 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Arrenberg:1996:NNT

[895] Jutta Arrenberg. A note on the nonparametric test based on the L_{1}- version of the Cramér-von Mises statistic. Statistical Papers, 37(2):95–104, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926575.

Kosfeld:1996:REF

[896] Reinhold Kosfeld. Robust exploratory factor analysis. Statistical Papers, 37(2):105–122, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926576.

Pakes:1996:CLBb

[897] Anthony G. Pakes. Characterization of laws by balancing weighting against a binary operation. Statistical Papers, 37(2):123–140, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926577.

Belzunce:1996:ACP

[898] F. Belzunce, J. Candel, and J. M. Ruiz. The ageing curve and partial orderings of life distributions. Statistical Papers, 37(2):141–152, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926578.

Jeevanand:1996:PEB

[899] E. S. Jeevanand and Veenus Padamadan. Parameter estimation for a bivariate Pareto distribution. Statistical Papers, 37(2):153–164, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02926579.

Brecht:1996:TDN

[900] Beatrix Brecht and Leo Brecht. Timediscrete nonparametric hazard model using panel data. Statistical Papers, 37(2):165–176, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02926580.

Winkelmann:1996:CDM

[901] Rainer Winkelmann. A count data model for gamma waiting times. Statistical Papers, 37(2):177–187, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926581.

Anonymous:1996:CECb

[902] Anonymous. Calendar of events and call for papers. Statistical Papers, 37(2):189–190, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926582.

Anonymous:1996:HCb

[903] Anonymous. Help & contacts. Statistical Papers, 37(2):??, June 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Brannas:1996:PDR

[904] K. Brännäs and P. Johansson. Panel data regression for counts. Statistical Papers, 37(3):191–213, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926583.

Crato:1996:SPT

[905] Nuno Crato and Howard M. Taylor. Stationary persistent time series misspecified as nonstationary ARIMA. Statistical Papers, 37(3): 215–223, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926584.

Hu:1996:TAS

[906] Wanhong Hu. Time aggregation and skip sampling in cointegration tests. Statistical Papers, 37(3):225– 234, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926585.

Lasch:1996:PCS

[907] Rainer Lasch. Pyramidal clustering schemes. Statistical Papers, 37(3): 235–251, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926586.

Srivastava:1996:OLS

[908] V. K. Srivastava, M. Dube, and Virender Singh. Ordinary least squares and Stein-rule predictions in regression models under inclusion of some superfluous variables. Statistical Papers, 37 (3):253–265, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926587.

Song:1996:CAU

[909] Seuck Heun Song. Consistency and asymptotic unbiasedness of S^{2} in the serially correlated error components regression model for panel data. Statistical Papers, 37(3):267–275, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926588.

Kramer:1996:ATE

[910] Walter Krämer, Robert Bartels, and Denzil G. Fiebig. Another twist on the equality of OLS and GLS. Statistical Papers, 37(3):277–281, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926589.

Anonymous:1996:CECc

[911] Anonymous. Calendar of events and call for papers. Statistical Papers, 37 (3):283–284, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926590.

Schmidt:1996:BRS

[912] J. Schmidt. Book review: Strecker, H. und R. Wiegert: (mit Beiträgen von G. Forsmann, H. Schneeberger und K. Fleischer): Stichproben, Erhebungsfehler, Datenqualität. Angewandte Statistik und Ökonometrie, Heft 36. Statistical Papers, 37(3): 285–287, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926591.

Anonymous:1996:HCc

[913] Anonymous. Help & contacts. Statistical Papers, 37(3):??, September 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Trenkler:1996:E

[914] G. Trenkler.  Editorial.  Statis-

tical Papers, 37(4):289, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02926109.

Beuckelaer:1996:CES

[915] A. De Beuckelaer. A closer examination on some parametric alternatives to the ANOVA F -test. Statistical Papers, 37(4):291–305, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926110.

Durairajan:1996:LTD

[916] T. M. Durairajan and K. J. Raman. LBI tests for the detection of compound normal distribution in control and treatment populations. Statistical Papers, 37(4):307–321, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926111.

Ohtani:1996:DFM

[917] Kazuhiro Ohtani and Judith A. Giles. The density function and the MSE dominance of the pre-test estimator in a heteroscedastic linear regression model with omitted variables. Statistical Papers, 37(4):323–342, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926112.

Mosler:1996:IIS

[918] Karl Mosler and Pietro Muliere. Inequality indices and the starshaped principle of transfers. Statistical Papers, 37(4):343–364, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926113.

Krumbholz:1996:NST

[919] Wolf Krumbholz and Friedrich Schmid. A non standard χ^{2}-test of fit for testing uniformity with unknown limits. Statistical Papers, 37(4):365–373, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926114.

Meaux:1996:CND

[920] Lauri M. Meaux, Dean M. Young, and John W. Seaman. A characterization of nonnegative-definite independence distribution-preserving covariance structures for the maximum squared-radii statistic. Statistical Papers, 37(4):375–382, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926115.

Anonymous:1996:CECd

[921] Anonymous. Calendar of events and call for papers. Statistical Papers, 37 (4):383–384, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02926116.

Anonymous:1996:HCd

[922] Anonymous. Help & contacts. Statistical Papers, 37(4):??, December 1996. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hsiao:1997:ENE

[923] C. Hsiao, L. Wang, and Q. Wang. Estimation of nonlinear errors-in-variables models: an approximate solution. Statistical Papers, 38(1):1–25, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925212.

Akahira:1997:GBD

[924] M. Akahira, H. Kashima, and K. Takahashi. A generalized binomial distribution determined by a two-state Markov chain and a distribution by the Bayesian approach. Statistical Papers, 38(1):27–42, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925213.

Hassler:1997:SAN

[925] U. Hassler. Sample autocorrelations of nonsationary fractionally integrated series. Statistical Papers, 38(1):43–62, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925214.

Liebscher:1997:GKM

[926] E. Liebscher and H. Schäbe. A generalization of the Kaplan–Meier estimator to Harris-recurrent Markov chains. Statistical Papers, 38(1):63–75, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925215.

Haldrup:1997:NID

[927] N. Haldrup and S. Hylleberg. Nearintegration and deterministic trends. Statistical Papers, 38(1):77–101, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925216.

Herff:1997:IPR

[928] W. Herff, B. Jochems, and U. Kamps. The inspection paradox with random time. Statistical Papers, 38(1):103–110, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925217.

Klufa:1997:DRA

[929] J. Klufa. Dodge–Romig AOQL single sampling plans for inspection by variables. Statistical Papers, 38 (1):111–119, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925218.

Anonymous:1997:CECa

[930] Anonymous. Calendar of events and call for papers. Statistical Papers, 38 (1):121–123, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925219.

Anonymous:1997:HCa

[931] Anonymous. Help & contacts. Statistical Papers, 38(1):??, March 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Abberger:1997:QSF

[932] Klaus Abberger. Quantile smoothing in financial time series. Statistical Papers, 38(2):125–148, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925220.

Balakrishna:1997:BSP

[933] N. Balakrishna and K. Jayakumar. Bivariate semi-Pareto distributions and processes. Statistical Papers, 38 (2):149–165, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925221.

Hanagal:1997:IPS

[934] David D. Hanagal. Inference procedures in some bivariate exponential models under hybrid random censoring. Statistical Papers, 38(2):167–189, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925222.

Schmid:1997:CCS

[935] W. Schmid. CUSUM control schemes for Gaussian processes. Statistical Papers, 38(2):191–217, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02925223.

Aebi:1997:CTP

[936] Robert Aebi. Contingency tables with prescribed marginals. Statistical Papers, 38(2):219–229, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925224.

Buse:1997:EPM

[937] A. Buse. On the equivalence of pooled and mixed estimation. Statistical Papers, 38(2):231–241, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925225.

Gohout:1997:NCU

[938] Wolfgang Gohout and Ingo Kuhnert. NBUFR closure under the formation of coherent systems. Statistical Papers, 38(2):243–248, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925226.

Anonymous:1997:CECb

[939] Anonymous. Calendar of events and call for papers. Statistical Papers, 38(2):249–251, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925227.

Anonymous:1997:HCb

[940] Anonymous. Help & contacts. Statistical Papers, 38(2):??, June 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Breitung:1997:TUR

[941] Jörg Breitung. Testing for unit roots in panel data using a GMM approach. Statistical Papers, 38(3): 253–269, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925268.

Hubler:1997:PLMa

[942] Olaf Hübler. Pseudo latent models: Goodness of fit measures, residuals, estimation, testing, and simulation. Statistical Papers, 38(3):271–285, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925269.

Park:1997:SAP

[943] YouSung Park and Chan Wook Oh. Some asymptotic properties in INAR(1) processes with Poisson marginals. Statistical Papers, 38(3): 287–302, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925270.

Schabe:1997:PES

[944] Hendrik Schäbe. Parameter estimation for a special class of Markov chains. Statistical Papers, 38(3): 303–327, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925271.

Wan:1997:EDD

[945] Alan T. K. Wan. The exact density and distribution functions of the inequality constrained and pre-test estimators. Statistical Papers, 38(3): 329–341, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925272.

Childs:1997:MLE

[946] Aaron Childs and N. Balakrishnan. Maximum likelihood estimation of Laplace parameters based on general type-II censored examples. Statistical Papers, 38(3):343–349, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925273.

Fraser:1997:AAS

[947] D. A. S. Fraser and A. C. M. Wong. On the accuracy of approximate Studentization. Statistical Papers, 38(3):351–356, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925274.

Huang:1997:MCC

[948] Mei Ling Huang and Karen Y. Fung. On moments and cumulants of the D compound Poisson distribution. Statistical Papers, 38(3):357–361, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925275.

Ziegler:1997:PCJ

[949] Andreas Ziegler. Practical considerations of the jackknife estimator of variance for generalized estimating equations. Statistical Papers, 38(3): 363–369, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925276.

Hassler:1997:BR

[950] Uwe Hassler and Wolfgang Härdle. Book reviews. Statistical Papers, 38 (3):370–372, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925277.

Anonymous:1997:CECc

[951] Anonymous. Calendar of events and call for papers. Statistical Papers, 38 (3):373–376, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925278.

Anonymous:1997:HCc

[952] Anonymous. Help & contacts. Statistical Papers, 38(3):??, September 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Arnold:1997:FPI

[953] Bernhard F. Arnold and Peter Stahlecker. Fuzzy prior information and minimax estimation in the linear regression model. Statistical Papers, 38 (4):377–391, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925995.

Bankhofer:1997:SSP

[954] Udo Bankhofer and Andreas Hilbert. Statistical software packages for windows: A market survey. Statistical Papers, 38(4):393–407, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925996.

Kale:1997:TRE

[955] Mohan Kale and T. V. Ramanathan. A test for randomness of the environments in a branching process. Statistical Papers, 38(4):409–421, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925997.

Parsian:1997:EPE

[956] A. Parsian and N. Sanjari Farsipour. Estimation of parameters of exponential distribution in the truncated space using asymmetric loss function. Statistical Papers, 38(4):423–443, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925998.

Gupta:1997:ECP

[957] Ramesh C. Gupta and Olcay Akman. Estimation of critical points in the mixture inverse Gaussian model. Statistical Papers, 38(4):445–452, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925999.

Hanagal:1997:NER

[958] David D. Hanagal. Note on estimation of reliability under bivariate Pareto stress-strength model. Statistical Papers, 38(4):453–459, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926000.

Hubler:1997:PLMb

[959] O. Hübler. Pseudo latent models: Goodness of fit measures, residuals, estimation, testing and simulation. Statistical Papers, 38(4):461–463, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926001.

Kramer:1997:BR

[960] Walter Krämer, Christian Kleiber, and Sabine Warschburger. Book reviews. Statistical Papers, 38(4):465–468, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926002.

Anonymous:1997:CECd

[961] Anonymous. Calendar of events and call for papers. Statistical Papers, 38 (4):469–471, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02926003.

Anonymous:1997:HCd

[962] Anonymous. Help & contacts. Statistical Papers, 38(4):??, December 1997. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Matyas:1998:MHP

[963] Lázló Mátyás and Pierre Blanchard. Misspecified heterogeneity in panel data models. Statistical Papers, 39 (1):1–27, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925369.

Heike:1998:OSA

[964] Hans-Dieter Heike and Wolfgang Jaspers. Optimum stratification and allocation in inventory sampling: An efficient two stage grid search procedure. Statistical Papers, 39(1):29–40, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925370.

Riedwyl:1998:MUY

[965] Hans Riedwyl. Modifying and using Yates’ algorithm. Statistical Papers, 39(1):41–60, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925371.

Rao:1998:PNC

[966] C. R. Rao, V. K. Srivastava, and H. Toutenburg. Pitman nearness comparisons of Stein-type estimators for regression coefficients in replicated experiments. Statistical Papers, 39(1):61–74, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925372.

Gebhard:1998:PTRa

[967] Jens Gebhard and Norbert Schmitz. Permutation tests — a revival?! Statistical Papers, 39(1):75–85, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925373.

Gebhard:1998:PTRb

[968] Jens Gebhard and Norbert Schmitz. Permutation tests — a revival?! Statistical Papers, 39(1):87–96, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925374.

Skala:1998:RTS

[969] Heinz J. Skala. On a representation theorem of Schmeidler. Statistical Papers, 39(1):97–107, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925375.

Wan:1998:BEL

[970] Alan T. K. Wan and William E. Griffiths. Bayesian estimation of the linear regression model with an uncertain interval constraint on coefficients. Statistical Papers, 39(1):109–118, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925376.

Kramer:1998:BRJ

[971] Walter Krämer. Book review: John Y. Campbell, Andrew W. Lo, A. Craig MacKinleay: The econometrics of financial markets. Statistical Papers, 39(1):119–121, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925377.

Anonymous:1998:CECa

[972] Anonymous. Calendar of events and call for papers. Statistical Papers, 39 (1):122–124, January 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925378.

Anonymous:1998:HCa

[973] Anonymous. Help & contacts. Statistical Papers, 39(1):??, January 1998.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bai:1998:SEE

[974] S. Kalpana Bai and T. M. Durairajan. Simultaneous equivariant estimation of the parameters of linear models. Statistical Papers, 39(2):125–134, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925401.

Chen:1998:JEP

[975] Zhenmin Chen. Joint estimation for the parameters of the extreme value distributions. Statistical Papers, 39 (2):135–146, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925402.

Franz:1998:SEF

[976] Jürgen Franz and Ryszard Magiera. Sequential estimation for a family of counting processes in the nuisance parameter case. Statistical Papers, 39 (2):147–162, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925403.

Kurumai:1998:EDD

[977] Hiroko Kurumai and Kazuhiro Ohtani. The exact distribution and density functions of a pre-test estimator of the error variance in a linear regression model with proxy variables. Statistical Papers, 39(2):163–177, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925404.

Runde:1998:LMP

[978] Ralf Runde. Locally most powerful two-sample rank tests for Lévy distributions. Statistical Papers, 39 (2):179–188, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925405.

Wong:1998:ASP

[979] A. Wong. Approximate Studentization for Pareto distribution with application to censored data. Statistical Papers, 39(2):189–201, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925406.

Hanagal:1998:TWS

[980] David D. Hanagal. Testing whether the survival function is multivariate new better than used. Statistical Papers, 39(2):203–211, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925407.

Heintel:1998:NBO

[981] Markus Heintel. A note on a Bayesian order determination procedure for vectorautoregressive processes. Statistical Papers, 39(2):213–221, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925408.

Joarder:1998:SUW

[982] Anwar H. Joarder. Some useful Wishart expectations based on the multivariate t-model. Statistical Papers, 39(2):223–229, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925409.

Lopez-Blazquez:1998:CGD

[983] Fernando López-Blázquez and Begoña Salamanca Miño. A characterization of the geometric distribution. Statistical Papers, 39(2):231–236, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925410.

Shalabh:1998:UPL

[984] Shalabh. Unbiased prediction in linear regression models with equi-correlated responses. Statistical Papers, 39 (2):237–244, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925411.

Anonymous:1998:CECb

[985] Anonymous. Calendar of events and call for papers. Statistical Papers, 39(2):245–248, April 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925412.

Anonymous:1998:HCb

[986] Anonymous. Help & contacts. Statistical Papers, 39(2):??, April 1998.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Franco:1998:CDP

[987] Manuel Franco and José M. Ruiz. Characterization of discrete populations through conditional expectations of order statistics. Statistical Papers, 39(3):249–262, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02929702.

Guillamon:1998:NEM

[988] A. Guillamón, J. Navarro, and J. M. Ruiz. Nonparametric estimator for mean residual life and vitality function. Statistical Papers, 39(3):263–276, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929703.

Jeevanand:1998:DNO

[989] E. S. Jeevanand and N. Unnikrishnan Nair. On determining the number of outliers in exponential and Pareto samples. Statistical Papers, 39(3):277–290, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929704.

Schabe:1998:ALT

[990] Hendrik Schäbe. Accelerated life testing models for nonhomogeneous Poisson processes. Statistical Papers, 39 (3):291–312, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02929705.

Dorfleitner:1998:CMB

[991] Gregor Dorfleitner. Conditional MSEbased discrimination of the sample mean and the post-stratification estimator in population sampling. Statistical Papers, 39(3):313–319, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929706.

Jaeger:1998:FTE

[992] Andreas Jaeger and Walter Krämer. A final twist on the equality of OLS and GLS. Statistical Papers, 39(3):321–324, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929707.

Kazarian:1998:CBG

[993] Levon Kazarian. A consistent bootstrapped GMM estimator for the linear model with arbitrary inequality constraints on parameters. Statistical Papers, 39(3):325–333, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929708.

Kusters:1998:BR

[994] Ulrich Küsters. Book reviews. Statistical Papers, 39(3):335–342, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929709.

Anonymous:1998:CECc

[995] Anonymous. Calendar of events and call for papers. Statistical Papers, 39(3):343–346, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02929710.

Anonymous:1998:HCc

[996] Anonymous. Help & contacts. Statistical Papers, 39(3):??, July 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Asadl:1998:CPS

[997] Majid Asadl. Characterization of the Pearson system of distributions based on reliability measures. Statistical Papers, 39(4):347–360, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927098.

Friedrich:1998:CPP

[998] Thomas Friedrich and Helmut Schellhaas. Computation of the percentage points and the power for the two-sided Kolmogorov–Smirnov one sample test. Statistical Papers, 39(4):361–375, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927099.

Kanefuji:1998:ESP

[999] Koji Kanefuji and Kosei Iwase. Estimation for a scale parameter with known coefficient of variation. Statistical Papers, 39(4):377–388, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927100.

Pal:1998:ENV

[1000] Nabendu Pal, Chiahua Ling, and Jyh-Jiuan Lin. Estimation of a normal variance — a critical review. Statistical Papers, 39(4):389–404, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927101.

Arnab:1998:RRS

[1001] Raghunath Arnab. Randomized response surveys: Optimum estimation of a finite population total. Statistical Papers, 39(4):405–408, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927102.

Chadhury:1998:NNN

[1002] Arijit Chadhury, Arun K. Adhikary, and Tapabrata Maiti. A note on non-negative mean square error estimation of regression estimators in randomized response surveys. Statistical Papers, 39(4):409–415, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927103.

Gupta:1998:NRT

[1003] A. K. Gupta and D. G. Kabe. A note on a result for two SUR models. Statistical Papers, 39(4):417–421, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927104.

Heijmans:1998:ESM

[1004] Risto Heijmans and Heinz Neudecker. Estimation of the SURE model. Statistical Papers, 39(4):423–430, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927105.

Yagoub:1998:BR

[1005] Yasir Yagoub, Walter Krämer, and Andreas Handl. Book reviews. Statistical Papers, 39(4):431–433, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927106.

Anonymous:1998:HCd

[1006] Anonymous. Help & contacts. Statistical Papers, 39(4):??, October 1998. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Cubeddu:1999:QAJ

[1007] C. Cubeddu and M. L. Targhetta. A quadratic approximation for jackknife estimators of the variance of sample mean functions. Statistical Papers, 40(1):1–12, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927107.

Kossler:1999:RTT

[1008] Wolfgang Kössler. Rank tests in the two-sample scale problem with unequal and unknown locations. Statistical Papers, 40(1):13–35, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927108.

Mosler:1999:SAM

[1009] Karl Mosler and Tomas Philipson.

Specification analysis in mixed hazard models and a test of crossing survival functions. Statistical Papers, 40 (1):37–54, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02927109.

Liu:1999:SCS

[1010] Shuangzhe Liu and Heinz Neudecker.

A survey of Cauchy–Schwarz and Kantorovich-type matrix inequalities. Statistical Papers, 40(1):55–73, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927110.

Ohtani:1999:RPP

[1011] Kazuhiro Ohtani. Risk performance of a pre-test estimator for normal variance with the Stein-variance estimator under the LINEX loss function. Statistical Papers, 40(1):75–87, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02927111.

Bar-Lev:1999:CIN

[1012] Shaul K. Bar-Lev and Benjamin Reiser.

On confidence intervals for nonmonotone parametric functions and an application to the squared mean of the normal distribution. Statistical Papers, 40(1):89–98, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927112.

Hanagal:1999:ESR

[1013] David D. Hanagal. Estimation of system reliability. Statistical Papers, 40 (1):99–106, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927113.

Heijmans:1999:WDE

[1014] Risto Heijmans. When does the expectation of a ratio equal the ratio of expectations? Statistical Papers, 40(1):107–115, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927114.

Kramer:1999:BR

[1015] Walter Krämer, Christian Kleiber, and Gerd Ronning. Book reviews. Statistical Papers, 40(1):117–120, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02927115.

Anonymous:1999:CECa

[1016] Anonymous. Calendar of events and call for papers. Statistical Papers, 40 (1):121–124, January 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02927116.

Anonymous:1999:HCa

[1017] Anonymous. Help & contacts. Statistical Papers, 40(1):??, January 1999.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Jackson:1999:SEA

[1018] J. Bradford Jackson. Some extensions and applications of Kriz’s (1972) urn model. Statistical Papers, 40 (2):125–142, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925513.

Dorfleitner:1999:RMM

[1019] Gregor Dorfleitner and Thomas Klein.

Rounding with multiplier methods: An efficient algorithm and applications in statistics. Statistical Papers, 40(2):143–157, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925514.

Pardo:1999:SSC

[1020] J. A. Pardo and M. C. Pardo.

Small-sample comparisons for the Rukhin goodness-of-fit-statistics. Statistical Papers, 40(2):159–174, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925515.

Thabane:1999:PNM

[1021] L. Thabane and M. Safiul Haq. Prediction from a normal model using a generalized inverse Gaussian prior. Statistical Papers, 40(2):175–184, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925516.

Song:1999:REF

[1022] Seuck Heun Song and Dietmar Stemann. Relative efficiency of first difference estimator in panel data regression with serially correlated error components. Statistical Papers, 40(2):185–198, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925517.

Baltagi:1999:DLR

[1023] Badi H. Baltagi. Double-length regressions for linear and log-linear regressions with AR(1) disturbances. Statistical Papers, 40(2):199–209, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925518.

Hanagal:1999:ERC

[1024] David D. Hanagal. Estimation of reliability of a component subjected to bivariate exponential stress. Statistical Papers, 40(2):211–220, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925519.

Krumbholz:1999:EPP

[1025] Wolf Krumbholz and Rainer Lassahn. Exact percentage points for the Kolmogorov test on truncated versions of known continuous distributions with unknown truncation parameters. Statistical Papers, 40(2):221–231, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925520.

Niermann:1999:GMD

[1026] Stefan Niermann. A generalization of the matching distribution. Statistical Papers, 40(2):233–238, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925521.

Bartels:1999:BR

[1027] Robert Bartels, Andreas Handl, Hilmar Drygas, and Norbert Schmitz. Book reviews. Statistical Papers, 40(2): 239–242, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925522.

Anonymous:1999:CECb

[1028] Anonymous. Calendar of events and call for papers. Statistical Papers, 40(2):243–245, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925523.

Anonymous:1999:HCb

[1029] Anonymous. Help & contacts. Statistical Papers, 40(2):??, April 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bhattacharya:1999:BEP

[1030] Samir K. Bhattacharya, Anoop Chaturvedi, and Nand Kishore Singh. Bayesian estimation for the Pareto income distribution. Statistical Papers, 40(3):247–262, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/BF02929874.

Kibria:1999:MLM

[1031] B. M. Golam Kibria and M. Safiul Haq. The multivariate linear model with multivariate t and intra-class covariance structure. Statistical Papers, 40 (3):263–276, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02929875.

Lubiano:1999:EEV

[1032] M. Asunción Lubiano and M. Angeles Gil. Estimating the expected value of fuzzy random variables in random samplings from finite populations. Statistical Papers, 40(3):277–295, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929876.

Nagel:1999:VGM

[1033] Hartmut Nagel and Rainer Schöbel. Volatility and GMM-Monte Carlo studies and empirical estimations. Statistical Papers, 40(3):297–321, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929877.

Chen:1999:CPA

[1034] Jie Chen and A. K. Gupta. Change point analysis of a Gaussian model. Statistical Papers, 40(3):323–333, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

springer.com/article/10.1007/BF02929878.

Lopez-Blazquez:1999:TCU

[1035] F. López-Blázquez and B. Salamanca Miño. On Terrel’s characterization of uniform distribution. Statistical Papers, 40(3):335–342, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929879.

Schellhaas:1999:MKS

[1036] Helmut Schellhaas. A modified Kolmogorov–Smirnov test for a rectangular distribution with unknown parameters: Computation of the distribution of the test statistic. Statistical Papers, 40(3):343–349, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929880.

Toutenburg:1999:WMF

[1037] Helge Toutenburg, Andreas Fieger, and V. K. Srivastava. Weighted modified first order regression procedures for estimation in linear models with missing X-observations. Statistical Papers, 40(3):351–361, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02929881.

Tracy:1999:CUE

[1038] Derrick S. Tracy, Housila P. Singh, and Rajesh Singh. Constructing an unbiased estimator of population mean in finite populations using auxiliary information. Statistical Papers, 40(3):363–368, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02929882.

Handl:1999:BR

[1039] Andreas Handl, Christian Kleiber, Walter Krämer, and Gertrud Moosmüller. Book reviews. Statistical Papers, 40 (3):369–373, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02929883.

Anonymous:1999:CECc

[1040] Anonymous. Calendar of events and call for papers. Statistical Papers, 40 (3):374–376, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02929884.

Anonymous:1999:HCc

[1041] Anonymous. Help & contacts. Statistical Papers, 40(3):??, September 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Alexander:1999:COM

[1042] T. Leo Alexander and B. Chandrasekar. Characterization of an optimal matrix estimator under convex loss function. Statistical Papers, 40 (4):377–391, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02934632.

Fagiuoli:1999:CDO

[1043] Enrico Fagiuoli, Franco Pellerey, and Moshe Shaked. A characterization of the dilation order and its applications. Statistical Papers, 40(4):393–406, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02934633.

Schroer:1999:TSL

[1044] Gunar Schröer. Two stage least squares estimation in structural cointegration models. Statistical Papers, 40 (4):407–438, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02934634.

Wachter:1999:GMM

[1045] Kenneth W. Wachter. Grade of membership models in low dimensions. Statistical Papers, 40(4):439–457, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02934635.

Chen:1999:SIA

[1046] Zhenmin Chen. Statistical inference about the shape parameter of the exponential power distribution. Statistical Papers, 40(4):459–468, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02934636.

Pordzik:1999:PTE

[1047] Paweł Pordzik. On pre-test estimation of parametric functions in the general Gauss–Markov model with quadratic risk. Statistical Papers, 40(4):469–478, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02934637.

Drygas:1999:BR

[1048] Hilmar Drygas, Siegfried Heiler, Günter Bamberg, Christian Kleiber, and Walter Krämer. Book reviews. Statistical Papers, 40(4):479–484, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02934638.

Anonymous:1999:CECd

[1049] Anonymous. Calendar of events and call for papers. Statistical Papers, 40 (4):485–487, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02934639.

Anonymous:1999:HCd

[1050] Anonymous. Help & contacts. Statistical Papers, 40(4):??, October 1999. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Trenkler:2000:E

[1051] Götz Trenkler.  Editorial.  Sta-

tistical Papers, 41(1):1, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02925673.

Pigeot:2000:BCM

[1052] Iris Pigeot. Basic concepts of multiple tests — a survey. Statistical Papers, 41(1):3–36, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925674.

Bachmaier:2000:AGE

[1053] Martin Bachmaier and Volker Guiard.

An alternative and generalized excess measure and its advantages. Statistical Papers, 41(1):37–52, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925675.

Bachmaier:2000:ECE

[1054] Martin Bachmaier. Efficiency comparison of M-estimates for scale at tdistributions. Statistical Papers, 41 (1):53–64, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925676.

Chaturvedi:2000:STP

[1055] Ajit Chaturvedi, Ajay Kumar, and K. Surinder. Sequential testing procedures for a class of distributions representing various life-testing models. Statistical Papers, 41(1):65–84, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925677.

Seidel:2000:LRT

[1056] Wilfried Seidel, Karl Mosler, and Manfred Alker. Likelihood ratio tests based on subglobal optimization: A power comparison in exponential mixture models. Statistical Papers, 41(1):85–98, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925678.

Lin:2000:NRR

[1057] Chien-Tai Lin. A note on the recurrence relations between moments of order statistics from right truncated log-logistic distribution. Statistical Papers, 41(1):99–107, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925679.

Sankaran:2000:SRA

[1058] P. G. Sankaran and N. Unnikrishnan Nair. On some reliability aspects of Pearson family of distributions. Statistical Papers, 41(1):109–117, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925680.

Kramer:2000:BR

[1059] Walter Krämer, Kepher Henry Makambi, Wolfgang Polasek, and Götz Trenkler. Book reviews. Statistical Papers, 41 (1):119–123, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925681.

Anonymous:2000:CECa

[1060] Anonymous. Calendar of events and call for papers. Statistical Papers, 41 (1):124–126, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925682.

Anonymous:2000:HCa

[1061] Anonymous. Help & contacts. Statistical Papers, 41(1):??, January 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hoffmann:2000:SER

[1062] Kurt Hoffmann. Stein estimation — a review. Statistical Papers, 41 (2):127–158, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926100.

Cramer:2000:AES

[1063] Erhard Cramer. Asymptotic estimators of the sample size in a record model. Statistical Papers, 41(2):159–171, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926101.

Kramer:2000:IPE

[1064] Holger Kramer and Wolfgang Schmid.

The influence of parameter estimation on the ARL of Shewhart type charts for time series. Statistical Papers, 41(2):173–196, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926102.

Li:2000:PBB

[1065] Hongyi Li. The power of bootstrap based tests for parameters in cointegrating regressions. Statistical Papers, 41(2):197–210, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926103.

Liu:2000:LIE

[1066] Shuangzhe Liu. On local influence for elliptical linear models. Statistical Papers, 41(2):211–224, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926104.

Baringhaus:2000:TFE

[1067] Ludwig Baringhaus and Norbert Henze. Tests of fit for exponentiality based on a characterization via the mean residual life function. Statistical Papers, 41(2):225–236, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926105.

Shalabh:2000:NFU

[1068] Shalabh. Note on a family of unbiased predictors for the equi-correlated responses in linear regression models. Statistical Papers, 41(2):237–241, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926106.

Heiler:2000:BR

[1069] Siegfried Heiler, Walter Krämer, Christian Kleiber, Philipp Sibbertsen, Kepher Makambi, and Thomas Wenzel. Book reviews. Statistical Papers, 41(2):243–248, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926107.

Anonymous:2000:CECb

[1070] Anonymous. Calendar of events and call for papers. Statistical Papers, 41(2):249–252, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926108.

Anonymous:2000:HCb

[1071] Anonymous. Help & contacts. Statistical Papers, 41(2):??, April 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wassmer:2000:BCG

[1072] Gernot Wassmer. Basic concepts of group sequential and adaptive group sequential test procedures. Statistical Papers, 41(3):253–279, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925923.

Czado:2000:MRA

[1073] Claudia Czado. Multivariate regression analysis of panel data with binary outcomes applied to unemployment data. Statistical Papers, 41(3):281–304, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925924.

Tsonias:2000:EPI

[1074] Efthymios G. Tsonias. Efficient posterior integration in stable Paretian models. Statistical Papers, 41(3):305–325, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925925.

Wencheko:2000:ESN

[1075] Eshetu Wencheko. Estimation of the signal-to-noise in the linear regression model. Statistical Papers, 41(3):327–343, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925926.

Bhatti:2000:OTE

[1076] M. Ishaq Bhatti. On optimal testing for the equality of equicorrelation: An example of loss in power. Statistical Papers, 41(3):345–352, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925927.

Caudill:2000:PCC

[1077] Steven B. Caudill. Pooling choices or categories in multinomial logit models. Statistical Papers, 41(3):353–358, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925928.

Chaturvedi:2000:SRP

[1078] Anoop Chaturvedi and Shri Prakash Singh. Stein rule prediction of the composite target function in a general linear regression model. Statistical Papers, 41(3):359–367, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925929.

Hansen:2000:BR

[1079] Gerd Hansen, Ricardo Maronna, Götz Trenkler, Andreas Handl, and Evelyn Korn. Book reviews. Statistical Papers, 41(3):369–374, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925930.

Anonymous:2000:CECc

[1080] Anonymous. Calendar of events and call for papers. Statistical Papers, 41(3):375–378, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925931.

Anonymous:2000:HCc

[1081] Anonymous. Help & contacts. Statistical Papers, 41(3):??, July 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Trenkler:2000:A

[1082] G. Trenkler. Announcement. Statistical Papers, 41(4):379, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02925758.

Buning:2000:RPP

[1083] Herbert Büning. Robustness and power of parametric, nonparametric, robustified and adaptive tests — the multi-sample location problem. Statistical Papers, 41(4):381–407, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925759.

Boscher:2000:FIR

[1084] Hans Boscher, Eva-Maria Fronk, and Iris Pigeot. Forecasting interest rates volatilities by GARCH (1,1) and stochastic volatility models. Statistical Papers, 41(4):409–422, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925760.

Su:2000:CBC

[1085] Jyh-Cherng Su and Wen-Jang Huang. Characterizations based on conditional expectations. Statistical Papers, 41 (4):423–435, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02925761.

Tsionas:2000:NBI

[1086] Efthymios G. Tsionas. Numerical Bayesian inference with arbitrary prior. Statistical Papers, 41(4):437–451, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925762.

Geng:2000:SPI

[1087] W. J. Geng and Alan T. K. Wan. On the sampling performance of an inequality pre-test estimator of the regression error variance under LINEX loss. Statistical Papers, 41(4):453–472, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02925763.

Barranco-Chamorro:2000:LDM

[1088] L. Barranco-Chamorro, J. L. Moreno-Rebollo, and E. López-Blázquez. Limiting distributions of MLE and UMVUE in the biparametric uniform distribution. Statistical Papers, 41 (4):473–484, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925764.

Schneeweiss:2000:BR

[1089] Hans Schneeweiß, Siegfried Heiler, Christian Kleiber, Ricardo Maronna, and Shuangzhe Liu. Book reviews. Statistical Papers, 41(4):485–490, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02925765.

Anonymous:2000:CECd

[1090] Anonymous. Calendar of events and call for papers. Statistical Papers, 41 (4):491–494, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02925766.

Anonymous:2000:HCd

[1091] Anonymous. Help & contacts. Statistical Papers, 41(4):??, October 2000. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Brunner:2001:NMF

[1092] Edgar Brunner and Madan L. Puri.

Nonparametric methods in factorial designs. Statistical Papers, 42 (1):1–52, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000039.

Chong:2001:ELN

[1093] Terence Tai-Leung Chong. Estimating the locations and number of change points by the samplesplitting method. Statistical Papers, 42(1):53–79, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000040.

Namba:2001:MPE

[1094] Akio Namba. MSE performance of the 2SHI estimator in a regression model with multivariate t error terms. Statistical Papers, 42 (1):81–96, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000041.

Cramer:2001:PSM

[1095] Jan S. Cramer and G. Ridder. Pooling states in the multinomial logit model: degrees of freedom. A correction. Statistical Papers, 42(1):97–99, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620000042.

Franco:2001:CDD

[1096] Manuel Franco and José M. Ruiz. Characterizations of discrete distributions based on conditional expectations of record values. Statistical Papers, 42 (1):101–110, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000043.

William:2001:EFB

[1097] Martin L. William and T. M. Durairajan. Estimating functions in the Bayesian paradigm. Statistical Papers, 42(1):111–122, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000044.

Wu:2001:CGM

[1098] Jong-Wuu Wu. Characterizations of generalized mixtures of geometric and exponential distributions based on upper record values. Statistical Papers, 42(1):123–133, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000045.

Hauser:2001:BR

[1099] Michael A. Hauser, Götz Trenkler, Ricardo Maronna, Walter Krämer, Peter Hackl, and Walter Krämer. Book reviews. Statistical Papers, 42(1): 134–138, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000046.

Anonymous:2001:CECa

[1100] Anonymous. Calendar of events and call for papers. Statistical Papers, 42 (1):139–142, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620000047.

Anonymous:2001:HCa

[1101] Anonymous. Help & contacts. Statistical Papers, 42(1):??, January 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Kunert:2001:QHM

[1102] Joachim Kunert, Astrid Montag, and Sigrid Pöhlmann. The quincunx: history and mathematics. Statistical Papers, 42(2):143–169, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s003620100048.

Bairamov:2001:DEA

[1103] Ismihan G. Bairamov and Samuel Kotz. On distributions of exceedances associated with order statistics and record values for arbitrary distributions. Statistical Papers, 42 (2):171–185, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100049.

Childs:2001:OSN

[1104] Aaron Childs, N. Balakrishnan, and Mohamed Moshref. Order statistics from non-identical right-truncated Lomax random variables with applications. Statistical Papers, 42 (2):187–206, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100050.

Stemann:2001:ECC

[1105] Dietmar Stemann and Claus Weihs. The EWMA-X-S-control chart and its performance in the case of precise and imprecise data. Statistical Papers, 42(2):207–223, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100051.

Wu:2001:CGE

[1106] Jong-Wuu Wu and Wen-Chuan Lee.

On the characterization of generalized extreme value, power function, generalized Pareto and classical Pareto distributions by conditional expectation of record values. Statistical Papers, 42(2):225–242, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100052.

Dasiou:2001:BPS

[1107] Despina Dasiou and Chronis Moyssiadis. The 50% breakdown point in simultaneous M-estimation of location and scale. Statistical Papers, 42(2):243–252, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100053.

Gotu:2001:EOG

[1108] Butte Gotu. The equality of OLS and GLS estimators in the linear regression model when the disturbances are spatially correlated. Statistical Papers, 42(2):253–263, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100054.

Leung:2001:ERC

[1109] Chi-Ying Leung. Error rates in classification consisting of discrete and continuous variables in the presence of covariates. Statistical Papers, 42 (2):265–273, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100055.

Heiler:2001:BR

[1110] Siegfried Heiler, Wolfgang Polasek, Christian Kleiber, and Gerd Ronning. Book reviews. Statistical Papers, 42 (2):274–278, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100056.

Anonymous:2001:CECb

[1111] Anonymous. Calendar of events and call for papers. Statistical Papers, 42(2):279–282, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100057.

Anonymous:2001:HCb

[1112] Anonymous. Help & contacts. Statistical Papers, 42(2):??, April 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Neudecker:2001:E

[1113] H. Neudecker and Risto Heijmans. Editorial. Statistical Papers, 42(3):283, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/accesspage/article/10.1007/
s003620100058.

Chen:2001:ACI

[1114] Zhenmin Chen and Jie Mi. An approximate confidence interval for the scale parameter of the gamma distribution based on grouped data. Statistical Papers, 42(3):285–299, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100059.

vanderLaan:2001:PTL

[1115] Paul van der Laan and Subha Chakraborti. Precedence tests and Lehmann alternatives. Statistical Papers, 42(3):301–312, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100060.

Menendez:2001:CDT

[1116] M. L. Menéndez, J. A. Pardo, and L. Pardo. Csiszar’s φ-divergences for testing the order in a Markov chain. Statistical Papers, 42(3):313–328, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100061.

Rodel:2001:MLR

[1117] Egmar Rödel. Mixed linear regression with equi-cross-correlated errors. Statistical Papers, 42(3):329–351, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100062.

Fernandez:2001:GMD

[1118] J. M. Vilar Fernández and W. González Manteiga. Generalized minimum distance estimators of a linear model with correlated errors. Statistical Papers, 42(3):353–373, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100063.

Wencheko:2001:CRE

[1119] Eshetu Wencheko. Comparison of regression estimators using Pitman measures of nearness. Statistical Papers, 42(3):375–386, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100064.

Abraham:2001:CMS

[1120] Bovas Abraham and N. Unnikrishnan Nair. On characterizing mixtures of some life distributions. Statistical Papers, 42(3):387–393, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s003620100065.

Navarro:2001:CTR

[1121] J. Navarro, Y. del Aguila, and J. M. Ruiz. Characterizations through reliability measures from weighted distributions. Statistical Papers, 42 (3):395–402, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100066.

Singh:2001:EMV

[1122] Sarjinder Singh, Munir Mahmood, and D. S. Tracy. Estimation of mean and variance of stigmatized quantitative variable using distinct units in randomized response sampling. Statistical Papers, 42(3):403–411, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100067.

Anonymous:2001:PSa

[1123] Anonymous. Problem section. Statistical Papers, 42(3):412–413, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s003620100068.

Pigeot:2001:BR

[1124] Iris Pigeot, Götz Trenkler, and Götz Trenkler. Book reviews. Statistical Papers, 42(3):415–418, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100069.

Anonymous:2001:CECc

[1125] Anonymous. Calendar of events and call for papers. Statistical Papers, 42(3):419–422, July 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100070.

Anonymous:2001:HCc

[1126] Anonymous. Help & contacts. Statistical Papers, 42(3):??, July 2001.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Arranz:2001:SFE

[1127] Miguel A. Arranz and Francesc Marmol. Out-of-sample forecast errors in misspecific perturbed long memory processes. Statistical Papers, 42 (4):423–436, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100071.

Ghitany:2001:CRS

[1128] M. E. Ghitany. A compound Rayleigh survival model and its application to randomly censored data. Statistical Papers, 42(4):437–450, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100072.

Khan:2001:CPT

[1129] Shahjahan Khan and A. K. Md. E. Saleh. On the comparison of the pre-test and shrinkage estimators for the univariate normal mean. Statistical Papers, 42(4):451–473, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s003620100073.

Neudecker:2001:SSP

[1130] Heinz Neudecker and Shuangzhe Liu. Some statistical properties of Hadamard products of random matrices. Statistical Papers, 42(4): 475–487, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100074.

Wu:2001:NDN

[1131] Jong-Wuu Wu. A note on determining the number of outliers in an exponential sample by least squares procedure. Statistical Papers, 42(4): 489–503, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100075.

Chang:2001:CAU

[1132] Horng-Jinh Chang and Kuo-Chung Huang. On constructing almost unbiased estimators of finite population mean using transformed auxiliary variable. Statistical Papers, 42 (4):505–515, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100076.

Mudelsee:2001:NBE

[1133] Manfred Mudelsee. Note on the bias in the estimation of the serial correlation coefficient of AR(1) processes. Statistical Papers, 42(4): 517–527, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s003620100077.

Neudecker:2001:SPH

[1134] Heinz Neudecker and Shuangzhe Liu.

Statistical properties of the Hadamard product of random vectors. Statistical Papers, 42(4):529–533, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s003620100078.

Anonymous:2001:PSb

[1135] Anonymous. Problem section. Statistical Papers, 42(4):535–536, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s003620100079.

Kleiber:2001:BR

[1136] Christian Kleiber, Gabriele Widmann, Walter Krämer, and Christian Kleiber. Book reviews. Statistical Papers, 42 (4):537–541, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100080.

Anonymous:2001:CECd

[1137] Anonymous. Calendar of events and call for papers. Statistical Papers, 42 (4):543–546, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s003620100081.

Anonymous:2001:HCd

[1138] Anonymous. Help & contacts. Statistical Papers, 42(4):??, October 2001. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Chateauneuf:2002:GIS

[1139] Alain Chateauneuf, Michéle Cohen, and Dieter Denneberg. General introduction to this special issue on Choquet integral and applications. Statistical Papers, 43(1):1–3, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-001-
0082-7.

Augustin:2002:EUW

[1140] Thomas Augustin. Expected utility within a generalized concept of probability — a comprehensive framework for decision making under ambiguity. Statistical Papers, 43(1):5–22, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-001-
0083-6.

Bruning:2002:MMA

[1141] Martin Brüning and Dieter Denneberg.

Max-min (σ-)additive representation of monotone measures. Statistical Papers, 43(1):23–35, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-001-0084-5.

Grabisch:2002:SAC

[1142] Michel Grabisch and Christophe Labreuche. The symmetric and asymmetric Choquet integrals on finite spaces for decision making. Statistical Papers, 43(1):37–52, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-001-
0085-4.

Heilpern:2002:UCI

[1143] Stanislaw Heilpern. Using Choquet integral in economics. Statistical Papers, 43(1):53–73, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-001-0086-3.

Maasss:2002:EFT

[1144] Sebastian Maasß. Exact functionals and their core. Statistical Papers, 43 (1):75–93, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-001-0087-2.

Philippe:2002:TDM

[1145] Fabrice Philippe. Tools for decision making under imprecise risk. Statistical Papers, 43(1):95–110, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-001-
                                        0088-1.

Vergnaud:2002:IC

[1146] Jean-Christophe Vergnaud. Information and capacities. Statistical Papers, 43(1):111–125, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-001-0089-0.

Billot:2002:SBA

[1147] Antoine Billot, Alain Chateauneuf, Itzhak Gilboa, and Jean-Marc Tallou. Sharing beliefs and the absence of betting in the Choquet expected utility model. Statistical Papers, 43 (1):127–136, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-001-0090-7.

Castaldo:2002:LTC

[1148] Adriana Castaldo and Massimo Marinacci. A Lusin theorem for a class of Choquet capacities. Statistical Papers, 43(1):137–142, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-001-0091-6.

Marinacci:2002:LAU

[1149] Massimo Marinacci. Learning from ambiguous urns. Statistical Papers, 43 (1):143–151, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-001-0092-5.

Christian:2002:BR

[1150] Kleiber Christian, Ricardo Maronna, and Leonhard Knorr-Held. Book reviews. Statistical Papers, 43(1): 153–156, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-001-0093-4.

Anonymous:2002:CECa

[1151] Anonymous. Calendar of events and call for papers. Statistical Papers, 43 (1):157–160, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-001-0094-3.

Anonymous:2002:HCa

[1152] Anonymous. Help & contacts. Statistical Papers, 43(1):??, January 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bisaglia:2002:MSP

[1153] Luisa Bisaglia and Silvano Bordignon. Mean square prediction error for longmemory processes. Statistical Papers, 43(2):161–175, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0095-x.

Haggstrom:2002:PHT

[1154] Erling Häggström. Properties of Honda’s test of random individual effects in non-linear regressions. Statistical Papers, 43(2):177–196, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0096-9.

Hartung:2002:SSP

[1155] Joachim Hartung, Dogan Argaç, and Kepher H. Makambi. Small sample properties of tests on homogeneity in one-way anova and meta-analysis. Statistical Papers, 43(2):197–235, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0097-8.

Johnk:2002:PEG

[1156] Max D. Jöhnk and Stefan Niermann. Parameter estimation with grouped data according to the linearization method — a comparison with alternative approaches. Statistical Papers, 43(2):237–255, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0098-7.

Raman:2002:THC

[1157] K. J. Raman and T. M. Surairajan. Testing homogeneity of control and treatment populations — local optimality and related issues. Statistical Papers, 43(2):257–271, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0099-6.

Akahira:2002:CID

[1158] Masafumi Akahira. Confidence intervals for the difference of means: application to the behrens–Fisher type problem. Statistical Papers, 43(2):273–284, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0100-4.

CaridadyOcerin:2002:PIE

[1159] José M. Caridad y Ocerin and José Diz Pérez. Point and interval estimators in a binomial-Poisson compound distribution. Statistical Papers, 43(2):285–290, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0101-3.

Vellaisamy:2002:IES

[1160] P. Vellaisamy and Abraham P. Punnen.

Improved estimators for the selected location parameters. Statistical Papers, 43(2):291–299, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-002-0102-2.

Anonymous:2002:PSa

[1161] Anonymous. Problem section. Statistical Papers, 43(2):301–302, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-002-0103-1.

Widmann:2002:BR

[1162] Gabriele Widmann, Peter Hackl, Ricardo Maronna, and Christian Kleiber. Book reviews. Statistical Papers, 43 (2):303–306, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-002-0104-0.

Anonymous:2002:CECb

[1163] Anonymous. Calendar of events and call for papers. Statistical Papers, 43(2):307–310, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-002-0105-z.

Anonymous:2002:HCb

[1164] Anonymous. Help & contacts. Statistical Papers, 43(2):??, April 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Akman:2002:CTD

[1165] Olcay Akman and Prashant Sansgiry. Critical time distribution of the mixture inverse Gaussian hazard rate. Statistical Papers, 43(3):311–322, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0106-y.

Bhuyan:2002:RER

[1166] Pradip Bhuyan and Pranita Sarmah.

Reliability estimation of a repairable standby redundant system. Statistical Papers, 43(3):323–336, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0107-x.

Cheung:2002:SPI

[1167] Siu Hung Cheung, Ka Ho Wu, and Siok Ping Lim. Simultaneous prediction intervals for multiple comparisons with a standard. Statistical Papers, 43(3):337–347, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0108-9.

Hutchinson:2002:CDR

[1168] T. P. Hutchinson, D. Cairns, and E. Chekaluk. The construction of data to reflect the research objective, and how randomisation tests make such data usable. Statistical Papers, 43(3):349–359, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0109-8.

Feldmann:2002:AMD

[1169] Benno Feldmann and Wolf Krumbholz. ASN-minimax double sampling plans for variables. Statistical Papers, 43(3):361–377, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0110-2.

Kukuk:2002:IEL

[1170] Martin Kukuk. Indirect estimation of (latent) linear models with ordinal regressors a Monte Carlo study and some empirical illustrations. Statistical Papers, 43(3):379–399, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0111-1.

Steland:2002:NMF

[1171] Ansgar Steland. Nonparametric monitoring of financial time series by jump-preserving control charts. Statistical Papers, 43(3):401–422, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0112-0.

Toutenburg:2002:PRV

[1172] H. Toutenburg and Shalabh. Prediction of response values in linear regression models from replicated experiments. Statistical Papers, 43(3):423–433, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0113-z.

Triacca:2002:CVP

[1173] Umberto Triacca. Cointegration in VAR(1) process: Characterization and testing. Statistical Papers, 43 (3):435–443, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0114-y.

Weba:2002:NLB

[1174] Michael Weba. A note on LeCam’s bound for the distance between the Poisson binomial and the Poisson distribution. Statistical Papers, 43 (3):445–452, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0115-x.

Anonymous:2002:PSb

[1175] Anonymous. Problem section. Statistical Papers, 43(3):453–457, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0116-9.

Trenkler:2002:BR

[1176] Götz Trenkler, Christian Kleiber, Peter Hackl, and Ricardo Maronna. Book reviews. Statistical Papers, 43 (3):459–462, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0117-8.

Anonymous:2002:CECc

[1177] Anonymous. Calendar of events and call for papers. Statistical Papers, 43(3):463–466, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-002-0118-7.

Anonymous:2002:HCc

[1178] Anonymous. Help & contacts. Statistical Papers, 43(3):??, July 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Henze:2002:ITM

[1179] Norbert Henze. Invariant tests for multivariate normality: a critical review. Statistical Papers, 43(4):467–506, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-002-
                                        0119-6.

Bouzas:2002:FCD

[1180] P. R. Bouzas, A. M. Aguilera, and M. J. Valderrama. Forecasting a class of doubly stochastic Poisson processes. Statistical Papers, 43(4):507–523, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-002-
                                        0120-0.

Gupta:2002:ECC

[1181] Jai P. Gupta. Estimation of the correlation coefficient in probability proportional to size with replacement sampling. Statistical Papers, 43(4): 525–536, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-002-0121-z.

Pham-Gia:2002:PQG

[1182] T. Pham-Gia and N. Turkkan. The product and quotient of general beta distributions. Statistical Papers, 43 (4):537–550, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0122-y.

Song:2002:BPR

[1183] Seuck Heun Song and Byoung Cheol Jung. BLUP in the panel regression model with spatially and serially correlated error components. Statistical Papers, 43(4):551–566, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0123-x.

Diaz-Emparanza:2002:SMC

[1184] Ignacio Díaz-Emparanza. Is a small Monte Carlo analysis a good analysis? Statistical Papers, 43(4):567–577, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0124-9.

Dumitrescu:2002:SIT

[1185] Monica Dumitrescu. Statistical inference for two Markov binomial models with applications. Statistical Papers, 43(4):579–585, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0125-8.

Mousa:2002:BPP

[1186] M. A. M. Ali Mousa and Z. F. Jaheen.

Bayesian prediction for progressively censored data from the Burr model.

Statistical Papers, 43(4):587–593, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0126-7.

Konopka:2002:ROT

[1187] Jörg Konopka and Norbert Schmitz. Redundant observations at testing i.i.d. random variables. Statistical Papers, 43 (4):595–602, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0127-6.

Anonymous:2002:PSc

[1188] Anonymous. Problem section. Statistical Papers, 43(4):603–606, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0128-5.

Gross:2002:BR

[1189] Jürgen Groß, Olaf Schoffer, and Christian Kleiber. Book reviews. Statistical Papers, 43(4):607–609, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0129-4.

Anonymous:2002:CECd

[1190] Anonymous. Calendar of events and call for papers. Statistical Papers, 43 (4):611–614, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0130-y.

Anonymous:2002:HCd

[1191] Anonymous. Help & contacts. Statistical Papers, 43(4):??, October 2002. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hoglund:2003:SPS

[1192] Rune Höglund and Ralf Östermark.

Size and power of some cointegration tests under structural breaks and heteroskedastic noise. Statistical Papers, 44(1):1–22, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0131-x.

Galea:2003:IDU

[1193] Manuel Galea, Gilberto A. Paula, and Miguel Uribe-Opazo. On influence diagnostic in univariate elliptical linear regression models. Statistical Papers, 44(1):23–45, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0132-9.

Paulaauskas:2003:MLE

[1194] Vygantas Paulaauskas and Svetlozar T.

Rachev. Maximum likelihood estimators in regression models with infinite variance innovations. Statistical Papers, 44(1):47–65, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0133-8.

Rodriguez-Avi:2003:NCD

[1195] José Rodríguez-Avi, Antonio Conde-Sánchez, and Antonio José Sáez-Castillo. A new class of discrete distributions with complex parameters. Statistical Papers, 44(1):67–88, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0134-7.

Salau:2003:EDC

[1196] M. O. Salau. The effects of different choices of order for autoregressive approximation on the Gaussian likelihood estimates for ARMA models. Statistical Papers, 44(1):89–105, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0135-6.

Arnold:2003:RSE

[1197] Bernhard F. Arnold and Peter Stahlecker. Relative squared error prediction in the generalized linear regression model. Statistical Papers, 44 (1):107–115, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0136-5.

Namba:2003:USV

[1198] Akio Namba. On the use of the Stein variance estimator in the double kclass estimator when each individual regression coefficient is estimated. Statistical Papers, 44(1):117–124, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0137-4.

Sankaran:2003:GPS

[1199] P. G. Sankaran, N. Unnikrishnan Nair, and T. K. Sindhu. A generalized Pearson system useful in reliability analysis. Statistical Papers, 44(1):125–130, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0138-3.

Stepanov:2003:CMR

[1200] A. Stepanov. Conditional moments of record times. Statistical Papers, 44 (1):131–140, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-002-0139-2.

Anonymous:2003:PSa

[1201] Anonymous. Problem section. Statistical Papers, 44(1):141–143, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-002-
0140-9.

Hackl:2003:BR

[1202] Peter Hackl, Krämer, and Ricardo Maronna. Book reviews. Statistical Papers, 44(1):144–146, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03036368.

Anonymous:2003:CECa

[1203] Anonymous. Calendar of events and call for papers. Statistical Papers, 44 (1):147–150, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-002-0142-7.

Anonymous:2003:HCa

[1204] Anonymous. Help & contacts. Statistical Papers, 44(1):??, January 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Childs:2003:HOM

[1205] Aaron Childs. Higher order moments of order statistics from INID exponential random variables. Statistical Papers, 44(2):151–167, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-003-0143-1.

Hassler:2003:NRD

[1206] Uwe Hassler. Nonsense regressions due to neglected time-varying means. Statistical Papers, 44(2):169–182, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-003-
                                        0144-0.

Ip:2003:SCR

[1207] W. C. Ip, Ying Yang, P. Y. K. Kwan, and Y. K. Kwan. Strong convergence rate of the least median absolute estimator in linear regression models. Statistical Papers, 44(2):183–201, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-003-
                                        0145-z.

Scheffner:2003:EUD

[1208] Axel Scheffner and Ralf Runde. Estimation of unimodal densities based on the FQ–system. Statistical Papers, 44(2):203–216, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0146-y.

Toutenburg:2003:ERM

[1209] H. Toutenburg and Shalabh. Estimation of regression models with equicorrelated responses when some observations on the response variable are missing. Statistical Papers, 44 (2):217–232, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0147-x.

Verma:2003:MGM

[1210] Sanjay Verma and R. Karan Singh.

A modified generalized mixed regression estimator when disturbances are nonnormal. Statistical Papers, 44 (2):233–248, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0148-9.

Farsipour:2003:AEN

[1211] N. Sanjari Farsipour. Admissibility of estimators in the non-regular family under entropy loss function. Statistical Papers, 44(2):249–256, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-003-
0149-8.

Neudecker:2003:BAP

[1212] Heinz Neudecker and Albert Satorra. On best affine prediction. Statistical Papers, 44(2):257–266, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-003-
                                        0150-2.

Espejo:2003:EFP

[1213] M. Ruiz Espejo, M. Delgado Pineda, and S. Nadarajah. Estimation of finite population parameters with several realizations. Statistical Papers, 44(2):267–278, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-003-0151-1.

Franco:2003:NCI

[1214] Manuel Franco, M. Carmen Ruiz, and José M. Ruiz. A note on closure of the ILR and DLR classes under formation of coherent systems. Statistical Papers, 44(2):279–288, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-003-0152-0.

Konopka:2003:ROT

[1215] Jörg Konopka and Norbert Schmitz. Redundant observations at testing i.i.d. random variables. Statistical Papers, 44(2):289, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF03036359.

Anonymous:2003:PSb

[1216] Anonymous. Problem section. Statistical Papers, 44(2):290–292, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03036360.

Kleiber:2003:BRa

[1217] Christian Kleiber, Heinz J. Skala, and Götz Trenkler. Book reviews. Statistical Papers, 44(2):293–296, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03036361.

Anonymous:2003:CECb

[1218] Anonymous. Calendar of events and call for papers. Statistical Papers, 44(2):297–300, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03036362.

Anonymous:2003:HCb

[1219] Anonymous. Help & contacts. Statistical Papers, 44(2):??, April 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Chaturvedi:2003:UER

[1220] Ajit Chaturvedi and Sanjeev K. Tomer.

UMVU estimation of the reliability function of the generalized life distributions. Statistical Papers, 44 (3):301–313, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0157-8.

Croux:2003:EMC

[1221] Cristophe Croux and Catherine Dehon. Estimators of the multiple correlation coefficient: Local robustness and confidence intervals. Statistical Papers, 44(3):315–334, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0158-7.

Wellmann:2003:IOO

[1222] Jürgen Wellmann and Ursula Gather. Identification of outliers in a oneway random effects model. Statistical Papers, 44(3):335–348, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-003-
0159-6.

Perez:2003:AAD

[1223] T. Pérez and J. A. Pardo. Asymptotic approximations for the distributions of the Kø-divergence goodnessof-fit statistics. Statistical Papers, 44(3):349–366, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0160-0.

Al-Saleh:2003:EME

[1224] Mohammad Fraiwan Al-Saleh and Said Ali Al-Hadhrami. Estimation of the mean of the exponential distribution using moving extremes ranked set sampling. Statistical Papers, 44 (3):367–382, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0161-z.

Shin:2003:TOI

[1225] Dong Wan Shin and Man-Suk Oh.

Tests for the order of integration against higher order integration. Statistical Papers, 44(3):383–396, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-003-
0162-y.

Wywial:2003:CSS

[1226] Janusz Wywial. On conditional sampling strategies. Statistical Papers, 44(3):397–419, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0163-x.

Jeske:2003:REO

[1227] Roland Jeske and Seuck Heun Song.

Relative efficiency of OLSE and COTE for seasonal autoregressive disturbances. Statistical Papers, 44 (3):421–432, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-003-0164-9.

Singh:2003:FSE

[1228] Housila P. Singh and Sushil K. Shukla.

A family of shrinkage estimators for the square of mean in normal distribution. Statistical Papers, 44(3):433–442, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-003-
0165-8.

Schmitz:2003:PS

[1229] Norbert Schmitz. Problem section. Statistical Papers, 44(3):443–446, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03036371.

Kleiber:2003:BRb

[1230] Christian Kleiber, Ricardo Maronna, and Hans Daduna. Book reviews. Statistical Papers, 44(3):447–450, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03036372.

Anonymous:2003:CECc

[1231] Anonymous. Calendar of events and call for papers. Statistical Papers, 44(3):451–454, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03036373.

Anonymous:2003:HCc

[1232] Anonymous. Help & contacts. Statistical Papers, 44(3):??, July 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Ahmadi:2003:NCT

[1233] J. Ahmadi and N. R. Arghami. Nonparametric confidence and tolerance intervals from record values data. Statistical Papers, 44(4):455–468, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926004.

Carstensen:2003:FSP

[1234] Kai Carstensen. The finite-sample performance of robust unit root tests. Statistical Papers, 44(4):469–482, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926005.

Hida:2003:AGH

[1235] Eisuke Hida and Masafumi Akahira.

An approximation to the generalized hypergeometric distribution. Statistical Papers, 44(4):483–497, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926006.

Hurlimann:2003:GAT

[1236] Werner Hürlimann. General affine transform families: why is the Pareto an exponential transform? Statistical Papers, 44(4):499–518, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926007.

Pardo:2003:TUB

[1237] M. C. Pardo. A test for uniformity based on informational energy. Statistical Papers, 44(4):521–534, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02926008.

Sim:2003:CEL

[1238] C. H. Sim and W. K. Wong. R-charts for the exponential, Laplace and logistic processes. Statistical Papers, 44 (4):535–554, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02926009.

Singh:2003:IPT

[1239] Sarjinder Singh and Balbinder Deo. Imputation by power transformation. Statistical Papers, 44(4):555–579, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926010.

Smith:2003:DDH

[1240] Murray D. Smith. On dependency in double-hurdle models. Statistical Papers, 44(4):581–595, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926011.

Anonymous:2003:PSc

[1241] Anonymous. Problem section. Statistical Papers, 44(4):597–599, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926012.

Maronna:2003:BR

[1242] Ricardo Maronna, Philipp Sibbertsen, and Olaf Hübler. Book reviews. Statistical Papers, 44(4):601–604, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02926013.

Anonymous:2003:CECd

[1243] Anonymous. Calendar of events and call for papers. Statistical Papers, 44 (4):605–608, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02926014.

Anonymous:2003:HCd

[1244] Anonymous. Help & contacts. Statistical Papers, 44(4):??, October 2003. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bernhard:2004:GMT

[1245] Gudrun Bernhard, Markus Klein, and Gerhard Hommel. Global and multiple test procedures using ordered pvalues — a review. Statistical Papers, 45(1):1–14, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02778266.

Chung:2004:MRW

[1246] Younshik Chung and Chansoo Kim.

Measuring robustness for weighted distributions: Bayesian perspective. Statistical Papers, 45(1):15–31, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778267.

Green:2004:BMC

[1247] Paul E. Green and Taesung Park.

Bayesian methods for contingency tables using Gibbs sampling. Statistical Papers, 45(1):33–50, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778268.

Mehrotra:2004:CNA

[1248] Devan V. Mehrotra. A cautionary note on the analysis of randomized block designs with a few missing values. Statistical Papers, 45(1):51–66, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778269.

Neumann:2004:GSA

[1249] Konrad Neumann and Stefan Zontek. On geometry of the set of admissible invariant quadratic estimators in balanced two variance components model. Statistical Papers, 45(1):67–80, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778270.

Rodriguez-Avi:2004:TDD

[1250] José Rodríguez-Avi, Antonio Conde-Sánchez, Antonio José Sáez-Castillo, and María José Olmo-Jiménez. A triparametric discrete distribution with complex parameters. Statistical Papers, 45(1):81–95, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02778271.

Sankaran:2004:IMU

[1251] P. G. Sankaran and S. M. Sunoj. Identification of models using failure rate and mean residual life of doubly truncated random variables. Statistical Papers, 45(1):97–109, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778272.

Langsrud:2004:GIS

[1252] Øyvind Langsrud. The geometrical interpretation of statistical tests in multivariate linear regression. Statistical Papers, 45(1):111–122, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778273.

Zeileis:2004:ABC

[1253] Achim Zeileis. Alternative boundaries for CUSUM tests. Statistical Papers, 45(1):123–131, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778274.

Sturtz:2004:BR

[1254] Sibylle Sturtz, Ricardo Maronna, and Peter Hackl. Book reviews. Statistical Papers, 45(1):133–135, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02778275.

Anonymous:2004:CECa

[1255] Anonymous. Calendar of events and call for papers. Statistical Papers, 45 (1):136–138, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02778276.

Anonymous:2004:HCa

[1256] Anonymous. Help & contacts. Statistical Papers, 45(1):??, January 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Gomez:2004:FBA

[1257] Guadalupe Gómez, M. Luz Calle, and Ramon Oller. Frequentist and Bayesian approaches for interval-censored data. Statistical Papers, 45(2):139–173, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02777221.

Akdeniz:2004:NBE

[1258] Fikri Akdeniz. New biased estimators under the LINEX loss function. Statistical Papers, 45(2):175–190, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02777222.

Aneiros-Perez:2004:PBC

[1259] Germán Aneiros-Pérez. Plug-in bandwidth choice for estimation of nonparametric part in partial linear regression models with strong mixing errors. Statistical Papers, 45(2):191–210, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02777223.

Augustin:2004:BAW

[1260] Thomas Augustin and Joachim Wolff. A bias analysis of Weibull models under heaped data. Statistical Papers, 45(2):211–229, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777224.

Rueda:2004:IRT

[1261] M. Rueda and A. Arcos. Improving ratio-type quantile estimates in a finite population. Statistical Papers, 45(2):231–248, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777225.

Sanchez:2004:IUR

[1262] Ismael Sánchez. Implementing unit root tests in ARMA models of unknown order. Statistical Papers, 45 (2):249–266, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777226.

Karioti:2004:SDO

[1263] Vassiliki Karioti and Chrysseis Caroni. Simple detection of outlying short time series. Statistical Papers, 45(2):267–278, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777227.

Farsipour:2004:ENM

[1264] N. Sanjari Farsipour and A. Asgharzadeh. Estimation of a normal mean relative to balanced loss functions. Statistical Papers, 45 (2):279–286, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777228.

Chang:2004:SRS

[1265] Horng-Jinh Chang, Chih-Li Wang, and Kuo-Chung Huang. Simple random sample equivalent survey designs reducing undesirable units from a finite population. Statistical Papers, 45(2):287–295, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02777229.

Neudecker:2004:Pa

[1266] Heinz Neudecker. Problemsection. Statistical Papers, 45(2):297–301, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02777230.

Maronna:2004:BRa

[1267] Ricardo A. Maronna, Christian Kleiber, and Olaf Schoffer. Book reviews. Statistical Papers, 45(2): 303–307, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02777231.

Anonymous:2004:CECb

[1268] Anonymous. Calendar of events and call for papers. Statistical Papers, 45(2):308–310, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02777232.

Anonymous:2004:HCb

[1269] Anonymous. Help & contacts. Statistical Papers, 45(2):??, April 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Gross:2004:GGM

[1270] Jürgen Groß. The general Gauss– Markov model with possibly singular dispersion matrix. Statistical Papers, 45(3):311–336, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777575.

Lekshmi:2004:APG

[1271] V. Seetha Lekshmi and K. K. Jose.

An autoregressive process with geometric α-Laplace marginals. Statistical Papers, 45(3):337–350, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02777576.

Kukush:2004:TEP

[1272] Alexander Kukush, Hans Schneeweiß, and Roland Wolf. Three estimators for the Poisson regression model with measurement errors. Statistical Papers, 45(3):351–368, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777577.

Pal:2004:ICC

[1273] Nabendu Pal and Wooi K. Lim.

On intra-class correlation coefficient estimation. Statistical Papers, 45 (3):369–392, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777578.

Wang:2004:CCT

[1274] Ronghua Wang and Heliang Fei. Conditions for the coincidence of the TFR, TRV and CE models. Statistical Papers, 45(3):393–412, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02777579.

Wywial:2004:CEA

[1275] Janusz Wywial. Conditional estimation of average on the basis of weighting data. Statistical Papers, 45(3):413–431, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.

1007/BF02777580. See errata [1300].

Bazargan-Lari:2004:DRW

[1276] A. Bazargan-Lari. Distribution of the range when sample size has a GPED_{1}. Statistical Papers, 45(3):433–438, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02777581.

Rodriguez:2004:SRO

[1277] Carmelo Rodríguez, Isabel Ortiz, and Ignacio Martínez. Some results on optimality in models with heteroscedastic errors from partial optimum designs. Statistical Papers, 45(3):439–450, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02777582.

Neudecker:2004:Pb

[1278] Heinz Neudecker and Götz Trenkler. Problemsection. Statistical Papers, 45(3):451–456, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02777583.

Sibbertsen:2004:BR

[1279] Philipp Sibbertsen, Roland Schultze, and Ricardo Maronna. Book reviews. Statistical Papers, 45(3):457–460, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02777584.

Anonymous:2004:CECc

[1280] Anonymous. Calender of events and call for papers. Statistical Papers, 45(3):461–464, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02777585.

Anonymous:2004:HCc

[1281] Anonymous. Help & contacts. Statistical Papers, 45(3):??, July 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Sibbertsen:2004:LMV

[1282] Philipp Sibbertsen. Long memory versus structural breaks: An overview. Statistical Papers, 45(4):465–515, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760564.

Hofmann:2004:CFI

[1283] Glenn Hofmann. Comparing the Fisher information in record data and random observations. Statistical Papers, 45(4):517–528, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760565.

Lin:2004:GQL

[1284] Lu Lin. Generalized quasi-likelihood. Statistical Papers, 45(4):529–544, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760566.

Wagner:2004:TIE

[1285] Niklas Wagner and Terry A. Marsh. Tail index estimation in small samples. Simulation results for independent and ARCH-type financial return models. Statistical Papers, 45(4):545–561, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760567.

Abdelkader:2004:CMO

[1286] Yousry H. Abdelkader. Computing the moments of order statistics from nonidentically distributed Erlang variables. Statistical Papers, 45 (4):563–570, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02760568.

Singh:2004:GCE

[1287] Sarjinder Singh, I. S. Grewal, and Anwar Joarder. General class of estimators in multi-character surveys. Statistical Papers, 45(4):571–582, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760569.

Ahmed:2004:PPL

[1288] Hadi Ahmed and Mohamed Kayid. Preservation properties for the Laplace transform ordering of residual lives. Statistical Papers, 45(4):583–590, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760570.

Maronna:2004:BRb

[1289] Ricardo Maronna, Peter Hackl, and Christian Kleiber. Book reviews. Statistical Papers, 45(4):591–593, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02760571.

Anonymous:2004:CECd

[1290] Anonymous. Calendar of events and call for papers. Statistical Papers, 45 (4):594–596, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02760572.

Anonymous:2004:HCd

[1291] Anonymous. Help & contacts. Statistical Papers, 45(4):??, October 2004. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Neuhauser:2005:ETB

[1292] Markus Neuhäuser. Exact tests based on the Baumgartner–Weiß–Schindler statistic — a survey. Statistical Papers, 46(1):1–29, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762032.

El-Arishy:2005:CVC

[1293] Samia El-Arishy. A conditional variance characterization of some discrete probability distributions. Statistical Papers, 46(1):31–45, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762033.

Knoth:2005:FIR

[1294] Seven Knoth. Fast initial response features for EWMA control charts. Statistical Papers, 46(1):47–64, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762034.

Li:2005:RPP

[1295] Xiaohu Li and Richard C. M. Yam. Reversed preservation properties of some negative aging conceptions and stochastic orders. Statistical Papers, 46(1):65–78, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762035.

Pettersson:2005:ODA

[1296] Hans Pettersson. Optimal design in average for inference in generalized linear models. Statistical Papers, 46(1):79–99, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762036.

Wencheko:2005:IEM

[1297] E. Wencheko and P. Wijekoon. Improved estimation of the mean in oneparameter exponential families with known coefficient of variation. Statistical Papers, 46(1):101–115, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762037.

Eryilmaz:2005:DES

[1298] Serkan Eryilmaz. On the distribution and expectation of success runs in nonhomogeneous Markov dependent trials. Statistical Papers, 46(1):117–128, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762038.

Chen:2005:NIA

[1299] Kung-Yu Chen and Chien-Tai Lin. A note on infinite-armed Bernoulli bandit problems with generalized beta prior distributions. Statistical Papers, 46 (1):129–140, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762039.

Wywial:2005:CEA

[1300] Janusz Wywial. Conditional estimation of average on the basis of weighting data — errata. Statistical Papers, 46(1):141–142, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02762040. See

[1275].

Neudecker:2005:PSa

[1301] Heinz Neudecker. Problem section.

Statistical Papers, 46(1):143–146, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762041.

Maronna:2005:BRa

[1302] Ricardo Maronna, Karsten Webel, and Christian Kleiber. Book reviews. Statistical Papers, 46(1):147–152, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02762042.

Anonymous:2005:CECa

[1303] Anonymous. Calender of events and call for papers. Statistical Papers, 46 (1):153–155, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02762043.

Anonymous:2005:HCa

[1304] Anonymous. Help & contacts. Statistical Papers, 46(1):??, January 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Ai:2005:CMA

[1305] Mingyao Ai and Runchu Zhang. Characterization of minimum aberration mixed factorials in terms of consulting designs. Statistical Papers, 46 (2):157–171, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02762966.

Chen:2005:ATS

[1306] Gemai Chen and Jinhong You. An asymptotic theory for semiparametric generalized least squares estimation in partially linear regression models. Statistical Papers, 46(2):173–193, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762967.

Jung:2005:ECF

[1307] Robert C. Jung, Gerd Ronning, and A. R. Tremayne. Estimation in conditional first order autoregression with discrete support. Statistical Papers, 46(2):195–224, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762968.

Meyners:2005:DLB

[1308] Michael Meyners. Deriving a lower bound for the proportion of perceivers in replicated difference tests by means of multiple test theory. Statistical Papers, 46(2):225–246, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762969.

Pham-Gia:2005:BDC

[1309] T. Pham-Gia and N. Turkhan.

Bayesian decision criteria in the presence of noises under quadratic and absolute value loss functions. Statistical Papers, 46(2):247–266, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762970.

Wencheko:2005:SSP

[1310] Eshetu Wencheko. Softly shrunk and partially shrunk rank-reduced estimation of the regression coefficients. Statistical Papers, 46(2):267–279, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02762971.

Nassar:2005:NSC

[1311] Manal M. Nassar. A note on some characterizations of the hyperexponential distribution. Statistical Papers, 46(2):281–292, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02762972.

Park:2005:CFI

[1312] Sangun Park. On calculating the fisher information in order statistics. Statistical Papers, 46(2):293–301, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02762973.

Srivastava:2005:FOR

[1313] V. K. Srivastava and H. Toutenburg. On the first order regression procedure of estimation for incomplete regression models. Statistical Papers, 46(2):303–307, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02762974.

Neudecker:2005:P

[1314] Heinz Neudecker and Götz Trenkler. Problemsection. Statistical Papers, 46(2):309–312, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02762975.

Maronna:2005:BRb

[1315] Ricardo A. Maronna, Christian Kleiber, and Olaf Schoffer. Book reviews. Statistical Papers, 46(2): 313–317, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762976.

Anonymous:2005:HCb

[1316] Anonymous. Help & contacts. Statistical Papers, 46(2):??, April 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Al-Hussaini:2005:PEU

[1317] Essam K. Al-Hussaini and Saieed F.

Ateya. Parametric estimation under a class of multivariate distributions. Statistical Papers, 46(3):321–338, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762837.

Arcos:2005:UMA

[1318] A. Arcos, M. Rueda, and M. D.

Martínez-Miranda. Using multiparametric auxiliary information at the estimation stage. Statistical Papers, 46(3):339–358, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762838.

Kantam:2005:OGL

[1319] R. R. L. Kantam, A. Vasudeva Roa, and G. Srinivasa Roa. Optimum group limits for estimation in scaled loglogistic distribution from a grouped data. Statistical Papers, 46(3):359–377, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762839.

Khan:2005:EIP

[1320] Shahjahan Khan, Zahirul Hoque, and A. K. Md. E. Saleh. Estimation of the intercept parameter for linear regression model with uncertain non-sample prior information. Statistical Papers, 46(3):379–395, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762840.

Landaburu:2005:DBE

[1321] E. Landaburu, D. Morales, and L. Pardo. Divergence-based estimation and testing with misclassified data. Statistical Papers, 46(3):397–409, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762841.

Makri:2005:BCB

[1322] Frosso S. Makri and Andreas N. Philippou. On binomial and circular binomial distributions of order k for l-overlapping success runs of length k. Statistical Papers, 46(3):411–432, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762842.

Schurle:2005:MCC

[1323] Josef Schürle. A method for consideration of conditional dependencies in the Fellegi and Sunter model of record linkage. Statistical Papers, 46(3):433–449, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762843.

Li:2005:NOD

[1324] Kim-Hung Li, Tai-Shing Lau, and Chongqi Zhang. A note on D-optimal designs for models with and without an intercept. Statistical Papers, 46(3):451–458, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02762844.

Romanas:2005:SCW

[1325] Yanushkevichius Romanas and Yanushkevichiene Olga. Stability of characterization of Weibull distribution. Statistical Papers, 46(3):459–468, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762845.

Neudecker:2005:PSb

[1326] Heinz Neudecker and Michel van de Velden. Problem section. Statistical Papers, 46(3):469–472, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762846.

Maronna:2005:BRc

[1327] Ricardo Maronna, Roland Schultze, and Christian Kleiber. Book reviews. Statistical Papers, 46(3):473–476, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02762847.

Anonymous:2005:CECb

[1328] Anonymous. Calendar of events and call for papers. Statistical Papers, 46(3):477–481, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02762848.

Anonymous:2005:HCc

[1329] Anonymous. Help & contacts. Statistical Papers, 46(3):??, July 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Alexander:2005:SEE

[1330] T. Leo Alexander and B. Chandrasekar. Simultaneous equivariant estimation of the parameters of matrix scale and matrix location-scale models. Statistical Papers, 46(4):483–507, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02763001.

Belzunce:2005:NRP

[1331] Félix Belzunce, Eva-María Ortega, and José M. Ruiz. A note on replacement policy comparisons from NBUC lifetime of the unit. Statistical Papers, 46(4):509–522, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02763002.

Kashima:2005:AMB

[1332] Hiroyuki Kashima. An application of a minimax Bayes rule and shrinkage estimators to the portfolio selection problem under the Bayesian approach. Statistical Papers, 46(4):523–540, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02763003.

Kuhnert:2005:FMA

[1333] Ronny Kuhnert and Dankmar Böhning.

The failure of meta-analytic asymptotics for the seemingly efficient estimator of the common risk difference. Statistical Papers, 46(4):541–554, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02763004.

Lovison:2005:RSP

[1334] Gianfranco Lovison. On Rao score and Pearson X^{2} statistics in generalized linear models. Statistical Papers, 46 (4):555–574, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02763005.

Ahmad:2005:BMG

[1335] Ibrahim A. Ahmad and A. R. Mugdadi.

Bounds of moment generating functions of some life distributions. Statistical Papers, 46(4):575–585, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02763006.

Janardan:2005:DDA

[1336] Konanur G. Janardan. A discrete distribution associated with a pure birth process. Statistical Papers, 46 (4):587–597, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02763007.

Zontek:2005:AP

[1337] Stefan Zontek and Konrad Neumann. Acknowledgement of priority. Statistical Papers, 46(4):599, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02763008.

Webel:2005:BR

[1338] Karsten Webel and W. Urfer. Book reviews. Statistical Papers, 46(4): 601–604, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02763009.

Anonymous:2005:HCd

[1339] Anonymous. Help & contacts. Statistical Papers, 46(4):??, October 2005. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Abberger:2006:KSP

[1340] Klaus Abberger. Kernel smoothed prediction intervals for ARMA models. Statistical Papers, 47(1):1–15, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-005-
                                        0269-4.

Abraham:2006:RER

[1341] B. Abraham and P. G. Sankaran. Renyi’s entropy for residual lifetime distribution. Statistical Papers, 47 (1):17–29, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-005-0270-y.

Osorio:2006:DCP

[1342] Felipe Osorio and Manuel Galea. Detection of a change-point in Studentt linear regression models. Statistical Papers, 47(1):31–48, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0271-x.

Kibria:2006:PMD

[1343] B. M. Golam Kibria and A. K. Ms. E.

Saleh. Pooling multivariate data under W, LR and LM tests. Statistical Papers, 47(1):49–68, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0272-9.

Nadarajah:2006:SPR

[1344] Saralees Nadarajah. Sums, products and ratios of generalized beta variables. Statistical Papers, 47(1):69–90, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0273-8.

Pardo:2006:MDE

[1345] Julio Angel Pardo, Leandro Pardo, and María del Carmen Pardo. Minimum Φ-divergence estimator in logistic regression models. Statistical Papers, 47 (1):91–108, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-005-0274-7.

Tanizaki:2006:LSB

[1346] Hisashi Tanizaki, Shigeyuki Hamori, and Yoichi Matsubayashi. On leastsquares bias in the AR(p) models: Bias correction using the bootstrap methods. Statistical Papers, 47(1):109–124, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-005-
                                        0275-6.

Ai:2006:IBS

[1347] Mingyao Ai and Shuyuan He. Interaction balance for symmetrical factorial designs with generalized minimum aberration. Statistical Papers, 47 (1):125–135, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-005-0276-5.

Hanagal:2006:BWR

[1348] David D. Hanagal. Bivariate Weibull regression model based on censored samples. Statistical Papers, 47(1): 137–147, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-005-0277-4.

Huang:2006:ESD

[1349] Kuo-Chung Huang. Estimation of sensitive data from a dichotomous population. Statistical Papers, 47(1):149–156, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-005-
0278-3.

Anonymous:2006:PS

[1350] Anonymous. Problem section. Statistical Papers, 47(1):157–158, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-005-0279-2.

Schaller:2006:BR

[1351] Mathias Schaller, Ricardo Maronna, and Uwe Ligges. Book reviews. Statistical Papers, 47(1):159–162, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0280-9.

Anonymous:2006:HCa

[1352] Anonymous. Help & contacts. Statistical Papers, 47(1):??, January 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Alonso:2006:IMU

[1353] Andrés M. Alonso, Daniel Peña, and Juan Romo. Introducing model uncertainty by moving blocks bootstrap. Statistical Papers, 47(2):167–179, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0282-7.

Didelez:2006:MBH

[1354] Vanessa Didelez, Iris Pigeot, and Patricia Walter. Modifications of the Bonferroni–Holm procedure for a multi-way ANOVA. Statistical Papers, 47(2):181–209, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-005-0283-6.

Friedman:2006:UFL

[1355] Craig Friedman and Sven Sandow. Utility functions that lead to the likelihood ratio as a relative model performance measure. Statistical Papers, 47(2):211–225, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-005-0284-5.

Jeon:2006:DOS

[1356] Jongwoo Jeon, Subhash Kochar, and Chul Gyu Park. Dispersive ordering — some applications and examples. Statistical Papers, 47(2):227–247, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0285-4.

Kim:2006:NPL

[1357] Dal Ho Kim, Sang Gil Kang, and Woo Dong Lee. Noninformative priors for linear combinations of the normal means. Statistical Papers, 47 (2):249–262, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-005-0286-3.

Lin:2006:REE

[1358] Lu Lin and Minghua Chen. Robust estimating equation based on statistical depth. Statistical Papers, 47 (2):263–278, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-005-0287-2.

Wang:2006:CVP

[1359] Lihong Wang and Jinde Wang.

Change-of-variance problem for linear processes with long memory. Statistical Papers, 47(2):279–298, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0288-1.

Childs:2006:HOM

[1360] Aaron Childs. Higher order moments of order statistics from INID symmetric random variables. Statistical Papers, 47(2):299–310, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-005-0289-0.

Pe:2006:ARN

[1361] Than Pe and Hilmar Drygas. An alternative representation of noncentral beta and F distributions. Statistical Papers, 47(2):311–318, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-005-
0290-7.

Kumar:2006:CDR

[1362] Manoj Kumar and Shashi Bahl. Class of dual to ratio estimators for double sampling. Statistical Papers, 47 (2):319–326, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-005-0291-6.

Anonymous:2006:HCb

[1363] Anonymous. Help & contacts. Statistical Papers, 47(2):??, March 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Heyer:2006:ORL

[1364] Herbert Heyer. Order relations for linear models: A survey on recent developments. Statistical Papers, 47 (3):331–372, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0293-z.

Ahmadi:2006:BEP

[1365] Jafar Ahmadi and M. Doostparast. Bayesian estimation and prediction for some life distributions based on record values. Statistical Papers, 47 (3):373–392, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0294-y.

Bruno:2006:CTI

[1366] Giancarlo Bruno and Edoardo Otranto. The choice of time interval in seasonal adjustment: A heuristic approach. Statistical Papers, 47(3):393–417, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0295-x.

Czado:2006:CLF

[1367] Claudia Czado and Adrian E. Raftery.

Choosing the link function and accounting for link uncertainty in generalized linear models using Bayes factors. Statistical Papers, 47(3):419–442, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0296-9.

Krivobokova:2006:ETS

[1368] Tatyana Krivobokova, Göran Kauermann, and Theofanis Archontakis. Estimating the term structure of interest rates using penalized splines. Statistical Papers, 47(3):443–459, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0297-8.

Gimenez:2006:ARE

[1369] Patricia Giménez, Enrico A. Colosimo, and Heleno Bolfarine. Asymptotic relative efficiency of score tests in Weibull models with measurement errors. Statistical Papers, 47(3):461–470, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0298-7.

Hubert:2006:ILE

[1370] M. H. Hubert and P. Wijekoon.

Improvement of the Liu estimator in linear regression model. Statistical Papers, 47(3):471–479, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0300-4.

Neudecker:2006:PSE

[1371] Heinz Neudecker and Götz Trenkler. Problem 4/SP06: Estimation of the Hadamard and cross products of two mean vectors in multivariate analysis. Statistical Papers, 47(3):481–485, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0301-3.

Webel:2006:BR

[1372] Karsten Webel and Rainer Schlittgen. Book reviews. Statistical Papers, 47 (3):487–488, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-006-0302-
                                        2.

Anonymous:2006:HCc

[1373] Anonymous. Help & contacts. Statistical Papers, 47(3):??, June 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Baltagi:2006:UPD

[1374] Badi H. Baltagi and Seuck Heun Song. Unbalanced panel data: A survey. Statistical Papers, 47(4):493–523, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0304-0.

Miazhynskaia:2006:CBM

[1375] Tatiana Miazhynskaia and Georg Dorffner. A comparison of Bayesian model selection based on MCMC with an application to GARCH-type models. Statistical Papers, 47(4):525–549, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0305-z.

Nadarajah:2006:EAD

[1376] Saraless Nadarajah. Exact and approximate distributions for the product of inverted Dirichlet components. Statistical Papers, 47(4):551–568, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0306-y.

Faraz:2006:HCC

[1377] Alireza Faraz and Ahmad Parsian.

Hotelling’s T^{2} control chart with double warning lines. Statistical Papers, 47(4):569–593, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0307-x.

Rosolowski:2006:ECM

[1378] M. Rosolowski and W. Schmid. EWNA charts for monitoring the mean and the autocovariances of stationary processes. Statistical Papers, 47(4): 595–630, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0308-9.

Chen:2006:SPD

[1379] Dechang Chen, Michael Fries, and Xiuzhen Cheng. Statistical properties of distance estimators. Statistical Papers, 47(4):631–641, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0309-8.

Kim:2006:BEB

[1380] Chansoo Kim and Younshik Chung. Bayesian estimation of P (Y < X) from Burr-type X model containing spurious observations. Statistical Papers, 47 (4):643–651, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0310-2.

Kibria:2006:BR

[1381] B. M. Golam Kibria, Matthias Arnold, and Rainer Schlittgen. Book reviews. Statistical Papers, 47(4):653–658, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0311-1.

Anonymous:2006:HCd

[1382] Anonymous. Help & contacts. Statistical Papers, 47(4):??, October 2006. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Emadi:2007:CRD

[1383] M. Emadi, J. Ahmadi, and N. R. Arghami. Comparison of record data and random observations based on statistical evidence. Statistical Papers, 48(1):1–21, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0313-z.

Singh:2007:CRR

[1384] Sarjinder Singh, Housila P. Singh, and Lakshmi N. Upadhyaya. Chain ratio and regression type estimators for median estimation in survey sampling. Statistical Papers, 48(1):23–46, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0314-y.

Jung:2007:NRP

[1385] Byoung Cheol Jung, Myoungshic Jhun, and Seuck Heun Song. A new random permutation test in ANOVA models. Statistical Papers, 48(1):47–62, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0315-x.

Stepanov:2007:NRW

[1386] A. Stepanov. The number of records within a random interval of the current record value. Statistical Papers, 48(1):63–79, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0316-9.

Stocker:2007:ABO

[1387] Toni Stocker. On the asymptotic bias of OLS in dynamic regression models with autocorrelated errors. Statistical Papers, 48(1):81–93, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0317-8.

Wu:2007:SIA

[1388] Jong-Wuu Wu and Hsiao-Chiao Tseng. Statistical inference about the shape parameter of the Weibull distribution by upper record values. Statistical Papers, 48(1):95–129, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0318-7.

Arnab:2007:RRTa

[1389] Raghunath Arnab and Georg Dorffner. Randomized response techniques for complex survey designs. Statistical Papers, 48(1):131–141, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0319-6.

Liau:2007:ESC

[1390] Pen-Hwang Liau. The existence of the strong combined-optimal design. Statistical Papers, 48(1):143–150, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0320-0.

Namba:2007:RCS

[1391] Akio Namba and Kazuhiro Ohtani. Risk comparison of the Stein-rule estimator in a linear regression model with omitted relevant regressors and multivariate t errors under the Pitman nearness criterion. Statistical Papers, 48 (1):151–162, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0321-z.

Maronna:2007:BR

[1392] Ricardo Maronna, Matthias Fischer, Jürgen Groß, and Andreas Karlsson. Book reviews. Statistical Papers, 48 (1):163–170, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0322-y.

Anonymous:2007:PSa

[1393] Anonymous. Problem section. Statistical Papers, 48(1):171–173, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0323-x.

Anonymous:2007:HCa

[1394] Anonymous. Help & contacts. Statistical Papers, 48(1):??, January 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Al-Saleh:2007:EMB

[1395] Mohammad Fraiwan Al-Saleh and Ahmad Mohammad Al-Ananbeh. Estimation of the means of the bivariate normal using moving extreme ranked set sampling with concomitant variable. Statistical Papers, 48(2):179–195, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0325-8.

Baratpour:2007:EPR

[1396] S. Baratpour, J. Ahmadi, and N. R. Arghami. Entropy properties of record statistics. Statistical Papers, 48(2):197–213, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0326-7.

Kim:2007:SUQ

[1397] Jong-Min Kim and Matthew E. Elam. A stratified unrelated question randomized response model. Statistical Papers, 48(2):215–233, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0327-6.

Li:2007:DMA

[1398] Peng-Fei Li, Min-Qian Liu, and Run-Chu Zhang. 2^{m}4^{1} designs with minimum aberration or weak minimum aberration. Statistical Papers, 48 (2):235–248, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0328-5.

Sordo:2007:CSO

[1399] Miguel A. Sordo and Héctor M. Ramos. Characterization of stochastic orders by L-functionals. Statistical Papers, 48(2):249–263, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0329-4.

Wong:2007:PAB

[1400] Tzu-Tsung Wong. Perfect aggregation of Bayesian analysis on compositional data. Statistical Papers, 48(2):265–282, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0330-y.

Ahmad:2007:RPS

[1401] Ibrahim Ahmad and Mohamed Kayid.

Reversed preservation of stochastic orders for random minima and maxima with applications. Statistical Papers, 48(2):283–293, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0331-x.

Behboodian:2007:MVG

[1402] Javad Behboodian, Ali Dolati, and Manuel Úbeda-Flores. A multivariate version of Gini’s rank association coefficient. Statistical Papers, 48(2):295–304, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0332-9.

Canal:2007:AIL

[1403] Luisa Canal and Rocco Micciolo. Admissibility intervals for linear correlation coefficients. Statistical Papers, 48(2):305–311, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0333-8.

Gupta:2007:CCa

[1404] K. Gupta, T. Nguyen, and L. Pardo. On Christensen’s conjecture. Statistical Papers, 48(2):313–319, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0334-7.

LaMotte:2007:DDR

[1405] Lynn Roy LaMotte. A direct derivation of the REML likelihood function. Statistical Papers, 48(2):321–327, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0335-6.

Nelson:2007:EN

[1406] Roger B. Nelson. Extremes of nonexchangeability. Statistical Papers, 48 (2):329–336, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0336-5.

Farsipour:2007:AEO

[1407] N. Sanjari Farsipour. Admissible estimation in an one parameter nonregular family of absolutely continuous distributions. Statistical Papers, 48(2):337–345, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0337-4.

Nadarajah:2007:LE

[1408] Saralees Nadarajah and Samuel Kotz. Letter to the Editor. Statistical Papers, 48(2):347–348, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-006-0338-
3.

Arnab:2007:RRTb

[1409] Raghunath Arnab. Randomized response techniques for complex survey designs. Statistical Papers, 48(2):349, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-006-0339-2.

Schlittgen:2007:BR

[1410] Rainer Schlittgen and Gero Szepannek. Book reviews. Statistical Papers, 48(2):351–352, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-006-0340-
9.

Anonymous:2007:HCb

[1411] Anonymous. Help & contacts. Statistical Papers, 48(2):??, April 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Bhowmik:2007:MIL

[1412] Jahar L. Bhowmik and Maxwell L.

King. Maximal invariant likelihood based testing of semi-linear models. Statistical Papers, 48(3):357–383, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0342-7.

Demetrescu:2007:END

[1413] Matei Demetrescu and Uwe Hassler. Effect of neglected deterministic seasonality on unit root tests. Statistical Papers, 48(3):385–402, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0343-6.

Gohout:2007:MVP

[1414] Wolfgang Gohout and Katja Specht. Mean-variance portfolios using Bayesian vector-autoregressive forecasts. Statistical Papers, 48(3):403–418, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0344-5.

Lachos:2007:IDG

[1415] Víctor H. Lachos, Filidor Vilca, and Manuel Galea. Influence diagnostics for the Grubbs’s model. Statistical Papers, 48(3):419–436, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0345-4.

Shen:2007:ELD

[1416] Junshan Shen and Shuyuan He. Empirical likelihood for the difference of quantiles under censorship. Statistical Papers, 48(3):437–457, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0346-3.

Bazargan-Lari:2007:PGF

[1417] A. Bazargan-Lari. Probability generating function of GPED_{2}. Statistical Papers, 48(3):459–466, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0347-2.

Choi:2007:NTD

[1418] Jong-Hoo Choi and Hyo-Il Park. A nonparametric test for diagnosis of the proportionality assumption. Statistical Papers, 48(3):467–477, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0348-1.

Danielak:2007:SCD

[1419] Katarzyna Danielak and Anna Dembińska.

Some characterizations of discrete distributions based on weak records. Statistical Papers, 48(3):479–489, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0349-0.

Qin:2007:NCB

[1420] Hong Qin and Mingyao Ai. A note on the connection between uniformity and generalized minimum aberration. Statistical Papers, 48(3):491–502, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0350-7.

Thannippara:2007:HDA

[1421] A. Thannippara, B. Kurian, D. K. Ghosh, S. C. Bagui, and S. Mandal. Hypercubic designs and applications. Statistical Papers, 48(3):503–508, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0351-6.

Torabi:2007:LRT

[1422] Hamzeh Torabi and Javad Behboodian. Likelihood ratio tests for fuzzy hypotheses testing. Statistical Papers, 48(3):509–522, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0352-5.

Gupta:2007:CCb

[1423] A. K. Gupta, T. Nguyen, and L. Pardo. On Christensen’s conjecture. Statistical Papers, 48(3):523, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-007-0379-
                                        2.

Anonymous:2007:PSb

[1424] Anonymous. Problem section. Statistical Papers, 48(3):525–534, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0354-3.

Polasek:2007:BRa

[1425] Wolfgang Polasek, W. Urfer, and B. M. Golam Kibria. Book reviews.

Statistical Papers, 48(3):535–538, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0353-4.

Anonymous:2007:CEC

[1426] Anonymous. Calendar of events and call for papers. Statistical Papers, 48 (3):539–542, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0355-2.

Anonymous:2007:HCc

[1427] Anonymous. Help & contacts. Statistical Papers, 48(3):??, ???? 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Abdul-Sathar:2007:VVB

[1428] E. I. Abdul-Sathar, R. P. Suresh, and K. R. Muraleedharan Nair. A vector valued bivariate Gini index for truncated distributions. Statistical Papers, 48(4):543–557, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0356-1.

Fukuda:2007:SRT

[1429] Kosei Fukuda. Simulated real-time detection of multiple structural changes: Evidence from Japanese economic growth. Statistical Papers, 48(4): 559–580, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0357-0.

Li:2007:SPR

[1430] Xiaohu Li and Guoxin Qiu. Some preservation results of NBUC aging property with applications. Statistical Papers, 48(4):581–594, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0358-z.

Ma:2007:ARM

[1431] Shuangge Ma. Additive risk model with case-cohort sampled current status data. Statistical Papers, 48(4): 595–608, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0359-y.

Tripathi:2007:EPN

[1432] Yogesh Mani Tripathi and Somesh Kumar. Estimating a positive normal mean. Statistical Papers, 48 (4):609–629, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0360-5.

Wang:2007:PSA

[1433] Liqun Wang and James C. Fu. A practical sampling approach for a Bayesian mixture model with unknown number of components. Statistical Papers, 48 (4):631–653, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0361-4.

Ziegler:2007:SCT

[1434] Andreas Ziegler. Simulated classical tests in multinomial probit models. Statistical Papers, 48(4):655–681, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0362-3.

Chaturvedi:2007:SRC

[1435] Ajit Chaturvedi, Neeraj Tiwari, and Sanjay Kumar. Some remarks on classical and Bayesian reliability estimation of binomial and Poisson distributions. Statistical Papers, 48(4):683–693, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0363-2.

Nelsen:2007:EN

[1436] Roger B. Nelsen. Extremes of nonexchangeability. Statistical Papers, 48(4): 695, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s00362-007-0380-9.

Polasek:2007:BRb

[1437] Wolfgang Polasek, Uwe Ligges, and Karsten Luebke. Book reviews. Statistical Papers, 48(4):697–701, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0366-z.

Anonymous:2007:PSc

[1438] Anonymous. Problem section. Statistical Papers, 48(4):703–706, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0367-y.

Anonymous:2007:HCd

[1439] Anonymous. Help & contacts. Statistical Papers, 48(4):??, October 2007. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Klufa:2008:DRA

[1440] J. Klufa. Dodge–Romig AOQL plans for inspection by variables from numerical point of view. Statistical Papers, 49(1):1–13, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0369-9.

Martin:2008:MPD

[1441] N. Martín and L. Pardo. Minimum phi-divergence estimators for loglinear models with linear constraints and multinomial sampling. Statistical Papers, 49(1):15–36, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0370-3.

Singh:2008:MEP

[1442] Housila P. Singh, Rajesh Tailor, Sarjinder Singh, and Jong-Min Kim. A modified estimator of population mean using power transformation. Statistical Papers, 49(1):37–58, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00362-006-
0371-2.

Upadhyay:2008:PAL

[1443] S. K. Upadhyay and M. Peshwani.

Posterior analysis of lognormal regression models using the Gibbs sampler. Statistical Papers, 49(1):59–85, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0372-1.

Vaish:2008:NDS

[1444] Akhil K. Vaish and N. Rao Chaganty. Nonnegative definite solutions to matrix equations with applications to multivariate test statistics. Statistical Papers, 49(1):87–99, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0373-0.

Samuel:2008:CDC

[1445] Philip Samuel. Characterization of distributions by conditional expectation of generalized order statistics. Statistical Papers, 49(1):101–108, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0364-1.

Scaria:2008:DEC

[1446] Johny Scaria and N. Unnikrishnan Nair. Distribution of extremes of r^{th} concomitant from the Morgenstern family. Statistical Papers, 49 (1):109–119, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0365-0.

Asgharzadeh:2008:EEM

[1447] A. Asgharzadeh and N. Sanjari Farsipour. Estimation of the exponential mean time to failure under a weighted balanced loss function. Statistical Papers, 49(1):121–131, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0375-y.

Espejo:2008:ISR

[1448] M. Ruiz Espejo, H. P. Singh, and S. Saxena. On inverse sampling without replacement. Statistical Papers, 49(1):133–137, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0376-x.

Thomas:2008:RRM

[1449] P. Yageen Thomas and Philip Samuel. Recurrence relations for the moments of order statistics from a beta distribution. Statistical Papers, 49(1):139–146, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0377-9.

Shalabh:2008:BR

[1450] Shalabh and W. Polasek. Book reviews. Statistical Papers, 49(1): 147–149, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0381-8.

Anonymous:2008:PSa

[1451] Anonymous. Problem section. Statistical Papers, 49(1):151–156, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0382-7.

Anonymous:2008:HCa

[1452] Anonymous. Help & contacts. Statistical Papers, 49(1):??, March 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Pal:2008:UET

[1453] Sanghamitra Pal. Unbiasedly estimating the total of a stigmatizing variable from a complex survey on permitting options for direct or randomized responses. Statistical Papers, 49 (2):157–164, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0001-z.

Jokiel-Rokita:2008:APO

[1454] Alicja Jokiel-Rokita. Asymptotically pointwise optimal and asymptotically optimal stopping times in the Bayesian inference. Statistical Papers, 49(2):165–175, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0002-y.

Navarro:2008:PST

[1455] Jorge Navarro, Jose M. Ruiz, and Carlos J. Sandoval. Properties of systems with two exchangeable Pareto components. Statistical Papers, 49 (2):177–190, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0003-x.

Schwingenschlogl:2008:AES

[1456] Udo Schwingenschlögl. Asymptotic equivalence of seat bias models. Statistical Papers, 49(2):191–200, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0005-8.

Gupta:2008:NLR

[1457] Arjun K. Gupta and Saralees Nadarajah. Normal and logistic random variables: distribution of the linear combination. Statistical Papers, 49 (2):201–209, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0006-7.

Joarder:2008:SUI

[1458] Anwar H. Joarder. Some useful integrals and their applications in correlation analysis. Statistical Papers, 49(2):211–224, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0007-6. See com-

ment [1472, 1544].

Dixit:2008:EPR

[1459] Ulhas J. Dixit and Parviz N. Nasiri. Estimation of parameters of a right truncated exponential distribution. Statistical Papers, 49(2):225–236, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0008-5.

Toutenburg:2008:AVI

[1460] H. Toutenburg, V. K. Srivastava, and Shalabh. Amputation versus imputation of missing values through ratio method in sample surveys. Statistical Papers, 49(2):237–247, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0009-4.

Chipoyera:2008:IEC

[1461] Honest W. Chipoyera and Eshetu Wencheko. Improved estimators of common variance of p-populations when kurtosis is known. Statistical Papers, 49(2):249–262, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0010-y.

Chacko:2008:EPB

[1462] Manoj Chacko and P. Yageen. Thomas.

Estimation of parameters of bivariate normal distribution using concomitants of record values. Statistical Papers, 49(2):263–275, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0011-x.

Wywial:2008:SDP

[1463] Janusz L. Wywial. Sampling design proportional to order statistic of auxiliary variable. Statistical Papers, 49 (2):277–289, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0012-9.

Arashi:2008:EIB

[1464] M. Arashi and M. Emadi. Evidential inference based on record data and inter-record times. Statistical Papers, 49(2):291–301, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0013-8.

Kuttykrishnan:2008:BSL

[1465] A. P. Kuttykrishnan and K. Jayakumar. Bivariate semi α-Laplace distribution and processes. Statistical Papers, 49(2):303–313, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0014-7.

Mosler:2008:HTW

[1466] Karl Mosler and Christoph Scheicher. Homogeneity testing in a Weibull mixture model. Statistical Papers, 49 (2):315–332, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0015-6.

Fernandez:2008:HPD

[1467] Arturo J. Fernández. Highest posterior density estimation from multiply censored Pareto data. Statistical Papers, 49(2):333–341, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0016-5.

Ristic:2008:GSP

[1468] Miroslav M. Ristić. A generalized semi-Pareto minification process. Statistical Papers, 49(2):343–351, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0017-4.

Liau:2008:PDT

[1469] Pen-Hwang Liau. Partial duplication in two-level fractional factorial designs. Statistical Papers, 49(2):353–361, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0018-3.

Figueiredo:2008:TWA

[1470] Adelaide Figueiredo. Two-way ANOVA for the Watson distribution defined on the hypersphere. Statistical Papers, 49(2):363–376, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0019-2.

Bacanli:2008:GST

[1471] Sevil Bacanli and Yaprak Parlak Demirhan. A group sequential test for the inverse Gaussian mean. Statistical Papers, 49(2):377–386, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0020-9.

Nadarajah:2008:CPH

[1472] Saralees Nadarajah. Comment on the paper by A. H. Joarder. Statistical Papers, 49(2):387–389, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-

0068-1. See [1458].

Anonymous:2008:HCb

[1473] Anonymous. Help & contacts. Statistical Papers, 49(2):??, April 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Grzadziel:2008:QSC

[1474] Mariusz Grzadziel. Quadratic subspaces and construction of Bayes invariant quadratic estimators of variance components in mixed linear models. Statistical Papers, 49(3):399–419, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0021-8.

Jayakumar:2008:GMO

[1475] K. Jayakumar and Thomas Mathew. On a generalization to Marshall–Olkin scheme and its application to Burr type XII distribution. Statistical Papers, 49(3):421–439, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0024-5.

Zhao:2008:DRI

[1476] S. Zhao and R. Zhang. 2^{m}4^{n} designs with resolution III or IV containing clear two-factor interaction components. Statistical Papers, 49 (3):441–454, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0025-4.

Liu:2008:ECH

[1477] Shuangzhe Liu and Chris C. Heyde.

On estimation in conditional heteroskedastic time series models under non-normal distributions. Statistical Papers, 49(3):455–469, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0026-3.

Wilde:2008:NGE

[1478] Joachim Wilde. A note on GMM estimation of probit models with endogenous regressors. Statistical Papers, 49(3):471–484, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0027-2.

Kim:2008:NSI

[1479] Hee-Young Kim and Yousung Park.

A non-stationary integer-valued autoregressive model. Statistical Papers, 49(3):485–502, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0028-1.

Ozkale:2008:CCE

[1480] M. Revan Özkale and Selahattin Kaçiranlar. Comparisons of the r − k class estimator to the ordinary least squares estimator under the Pitman’s closeness criterion. Statistical Papers, 49(3):503–512, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0029-0.

Singh:2008:FSE

[1481] Housila Prasad Singh, Sharad Saxena, and Harshada Joshi. A family of shrinkage estimators for Weibull shape parameter in censored sampling. Statistical Papers, 49(3):513–529, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0030-7.

Gschlossl:2008:MCD

[1482] Susanne Gschlößl and Claudia Czado. Modelling count data with overdispersion and spatial effects. Statistical Papers, 49(3):531–552, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-006-
                                        0031-6.

Tian:2008:NRS

[1483] Yongge Tian, M. Beisiegel, E. Dagenais, and C. Haines. On the natural restrictions in the singular Gauss– Markov model. Statistical Papers, 49(3):553–564, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0032-5.

Jonsson:2008:ANA

[1484] Kristian Jönsson. The accuracy of normal approximation in a heterogeneous panel data unit root test. Statistical Papers, 49(3):565–579, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0033-4.

Rodrigues:2008:NTN

[1485] Paulo M. M. Rodrigues and Antonio Rubia. A note on testing for nonstationarity in autoregressive processes with level dependent conditional heteroskedasticity. Statistical Papers, 49(3):581–593, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0035-2.

Fischer:2008:BRS

[1486] Matthias Fischer. Book review:

Scherer, B. and Martin, R. D.: Introduction to modern portfolio optimization with NUOPT and SPLUS and S+Bayes. Statistical Papers, 49 (3):595–596, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-007-0076-
1.

Sailer:2008:BRT

[1487] Oliver Sailer. Book review: Thomas P. Ryan: Modern Experimental Design. Statistical Papers, 49(3):597–598, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-007-0090-3.

Anonymous:2008:PSb

[1488] Anonymous. Problem section. Statistical Papers, 49(3):599–601, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0094-z.

Anonymous:2008:HCc

[1489] Anonymous. Help & contacts. Statistical Papers, 49(3):??, July 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Okolewski:2008:BEC

[1490] Andrzej Okolewski and Marek Kaluszka. Bounds for expectations of concomitants. Statistical Papers, 49(4): 603–618, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0041-4.

Bouzar:2008:TSD

[1491] Nadjib Bouzar and K. Jayakumar. Time series with discrete semistable marginals. Statistical Papers, 49(4): 619–635, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0040-5.

Jiang:2008:NIR

[1492] L. Jiang and A. C. M. Wong. A note on inference for P (X < Y) for right truncated exponentially distributed data. Statistical Papers, 49(4):637–651, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0034-3.

Kosfeld:2008:FAR

[1493] Reinhold Kosfeld and Jørgen Lauridsen. Factor analysis regression. Statistical Papers, 49(4):653–667, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0036-1.

Sakallioglu:2008:NBE

[1494] Sadullah Sakallioglu and Selahattin Kaçiranlar. A new biased estimator based on ridge estimation. Statistical Papers, 49(4):669–689, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0037-0.

Czado:2008:SSM

[1495] Claudia Czado and Peter X.-K. Song.

State space mixed models for longitudinal observations with binary and binomial responses. Statistical Papers, 49 (4):691–714, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0039-y.

Chae:2008:CAU

[1496] Seong S. Chae, Chansoo Kim, Jong-Min Kim, and William D. Warde.

Cluster analysis using different correlation coefficients. Statistical Papers, 49(4):715–727, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-006-0043-2.

Shakil:2008:DPR

[1497] M. Shakil, B. M. Golam Kibria, and Kuang-Chao Chang. Distributions of the product and ratio of Maxwell and Rayleigh random variables. Statistical Papers, 49(4):729–747, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0052-9.

Li:2008:RHR

[1498] Xiaohu Li and Maochao Xu. Reversed hazard rate order of equilibrium distributions and a related aging notion. Statistical Papers, 49(4):749–767, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0046-7.

Sharafi:2008:BSN

[1499] M. Sharafi and J. Behboodian. The Balakrishnan skew-normal density. Statistical Papers, 49(4):769–778, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0038-z.

Sankaran:2008:POM

[1500] P. G. Sankaran and K. Jayakumar. On proportional odds models. Statistical Papers, 49(4):779–789, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-006-
0042-3.

Gonzalez:2008:IEA

[1501] S. González, M. Rueda, and A. Arcos. An improved estimator to analyse missing data. Statistical Papers, 49 (4):791–796, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0045-8.

Franco:2008:LE

[1502] Manuel Franco and Juana-María Vivo.

Letter to the Editor. Statistical Papers, 49(4):797–798, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-008-0126-
3.

Richter:2008:BRW

[1503] Bernd Richter. Book review: Walter Krämer, Olaf Schoffer, Lars Tschiersch: Datenanalyse mit SAS — Statistische Verfahren und ihre grafischen Aspekte. Statistical Papers, 49(4):799–800, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-007-0091-2.

Polasek:2008:BRJ

[1504] Wolfgang Polasek. Book review: John F. Geweke (2005): Contemporary Bayesian econometrics and statistics (Wiley series in probability and statistics). Statistical Papers, 49(4):801–802, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-007-0098-8.

Anonymous:2008:PSc

[1505] Anonymous. Problem section. Statistical Papers, 49(4):803–806, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0119-2.

Anonymous:2008:HCd

[1506] Anonymous. Help & contacts. Statistical Papers, 49(4):??, October 2008. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Choi:2009:IMM

[1507] Dongseok Choi and Saralees Nadarajah. Information matrix for a mixture of two Laplace distributions. Statistical Papers, 50(1):1–12, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0053-8.

Zhou:2009:NCC

[1508] Chunguang Zhou, Changliang Zou, Yujuan Zhang, and Zhaojun Wang. Nonparametric control chart based on change-point model. Statistical Papers, 50(1):13–28, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0054-7.

Hanagal:2009:WEB

[1509] David D. Hanagal. Weibull extension of bivariate exponential regression model with different frailty distributions. Statistical Papers, 50(1):29–49, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0057-4.

Maravelakis:2009:UAC

[1510] Petros E. Maravelakis and Sotirios Bersimis. The use of Andrews curves for detecting the out-of-control variables when a multivariate control chart signals. Statistical Papers, 50(1):51–65, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0060-9.

Lee:2009:CMP

[1511] Taewook Lee and Sangyeol Lee. Consistency of minimizing a penalized density power divergence estimator for mixing distribution. Statistical Papers, 50(1):67–80, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0062-7.

Li:2009:REM

[1512] Jiantao Li and Min Zheng. Robust estimation of multivariate regression model. Statistical Papers, 50 (1):81–100, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0063-6.

Zadlo:2009:ME

[1513] Tomasz Zadlo. On MSE of EBLUP. Statistical Papers, 50(1):101–118, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0066-3.

Fernandez:2009:WIU

[1514] Arturo J. Fernández. Weibull inference using trimmed samples and prior information. Statistical Papers, 50 (1):119–136, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0067-2.

Wang:2009:BPB

[1515] Lichun Wang and Noël Veraverbeke. Bayes prediction based on right censored data. Statistical Papers, 50 (1):137–149, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-006-0044-1.

Ohtani:2009:CSU

[1516] Kazuhiro Ohtani and Alan T. K. Wan. Comparison of the Stein and the usual estimators for the regression error variance under the Pitman nearness criterion when variables are omitted. Statistical Papers, 50(1):151–160, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0047-6.

Perez-Gonzalez:2009:AAP

[1517] Carlos J. Pérez-González and Arturo J.

Fernández. Accuracy of approximate progressively censored reliability sampling plans for exponential models. Statistical Papers, 50(1):161–170, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0048-5.

Sadooghi-Alvandi:2009:DSI

[1518] S. M. Sadooghi-Alvandi, A. R. Nematollahi, and R. Habibi. On the distribution of the sum of independent uniform random variables. Statistical Papers, 50(1):171–175, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0049-4.

Potuschak:2009:MDS

[1519] Heinrich Potuschak and Werner G.

Müller. More on the distribution of the sum of uniform random variables. Statistical Papers, 50(1):177–183, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0050-y.

Isotalo:2009:NEO

[1520] Jarkko Isotalo and Simo Puntanen. A note on the equality of the OLSE and the BLUE of the parametric function in the general Gauss–Markov model. Statistical Papers, 50(1):185–193, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-007-
                                        0055-6.

Montoya:2009:CPL

[1521] José A. Montoya, Eloísa Díaz-Francés, and David A. Sprott. On a criticism of the profile likelihood function. Statistical Papers, 50(1):195–202, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0056-5.

Hutson:2009:DFE

[1522] Alan D. Hutson. A distribution function estimator for the difference of order statistics from two independent samples. Statistical Papers, 50 (1):203–208, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0058-3.

Nadarajah:2009:CPS

[1523] Saralees Nadarajah. Comment on the paper by Shakil et al. Statistical Papers, 50(1):209–211, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0069-0.

Wang:2009:BRR

[1524] Liqun Wang. Book review: Robert E. Weiss (2006): Modeling longitudinal data. Statistical Papers, 50(1):213–214, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-007-0099-7.

Fried:2009:BRM

[1525] Roland Fried. Book review: Marc S. Palella: Fundamental probability. A computational approach. Statistical Papers, 50(1):215–216, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-007-0106-z.

Lenz:2009:BRM

[1526] Sylvia Tamara Lenz. Book review:

Michael J. Campbell, David Machin and Stephen J. Walters (2007): Medical Statistics, a Textbook for the Health Sciences, 4th edition. Statistical Papers, 50(1):217–218, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-007-0109-9.

Anonymous:2009:HCa

[1527] Anonymous. Help & contacts. Statistical Papers, 50(1):??, January 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Nouira:2009:ETS

[1528] Leïla Nouira, Mohamed Boutahar, and Vêlayoudom Marimoutou. The effect of tapering on the semiparametric estimators for nonstationary long memory processes. Statistical Papers, 50 (2):225–248, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0071-6.

Abu-Dayyeh:2009:MIA

[1529] Walid Abu-Dayyeh and Esam Al Sawi. Modified inference about the mean of the exponential distribution using moving extreme ranked set sampling. Statistical Papers, 50(2):249–259, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0072-5.

Kharrati-Kopaei:2009:SSM

[1530] M. Kharrati-Kopaei, A. R. Nematollahi, and Z. Shishebor. On the sufficient statistics for multivariate ARMA models: approximate approach. Statistical Papers, 50(2):261–276, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0077-0.

Menendez:2009:PPD

[1531] M. L. Menéndez, L. Pardo, and M. C. Pardo. Preliminary phi-divergence test estimators for linear restrictions in a logistic regression model. Statistical Papers, 50(2):277–300, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0078-z.

Kadilar:2009:REP

[1532] Cem Kadilar, Yesim Unyazici, and Hulya Cingi. Ratio estimator for the population mean using ranked set sampling. Statistical Papers, 50 (2):301–309, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0079-y.

Curto:2009:MSM

[1533] José Dias Curto, José Castro Pinto, and Gonçalo Nuno Tavares. Modeling stock markets’ volatility using GARCH models with normal, Student’s t and stable Paretian distributions. Statistical Papers, 50(2):311–321, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0080-5.

Bekker:2009:TDR

[1534] Andriëtte Bekker, Jacobus Roux, and Thu Pham-Gia. The Type I distribution of the ratio of independent “Weibullized” generalized betaprime variables. Statistical Papers, 50 (2):323–338, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0083-2.

Sankaran:2009:NPE

[1535] P. G. Sankaran and Ansa Alphonsa Antony. Non-parametric estimation of lifetime distribution of competing risk models when censoring times are missing. Statistical Papers, 50(2):339–361, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0086-z.

Jokiel-Rokita:2009:SEL

[1536] Alicja Jokiel-Rokita and Agnieszka Stepień. Sequential estimation of a location parameter from delayed observations. Statistical Papers, 50 (2):363–372, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0059-2.

Falocci:2009:RMF

[1537] Nicola Falocci, Renato Paniccià, and Elena Stanghellini. Regression modelling of the flows in an input-output table with accounting constraints. Statistical Papers, 50(2):373–382, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0061-8.

Durante:2009:CNE

[1538] Fabrizio Durante. Construction of non-exchangeable bivariate distribution functions. Statistical Papers, 50 (2):383–391, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0064-5.

Kim:2009:NPN

[1539] D. H. Kim, S. G. Kang, and W. D. Lee. Noninformative priors for the normal variance ratio. Statistical Papers, 50(2):393–402, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-007-0065-4.

Bar-Lev:2009:CED

[1540] Shaul K. Bar-Lev and Benzion Boukai. A characterization of the exponential distribution by means of coincidence of location and truncated densities. Statistical Papers, 50(2):403–405, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0070-7.

Mahmoud:2009:GOS

[1541] M. A. W. Mahmoud and H. Sh. Al-Nagar. On generalized order statistics from linear exponential distribution and its characterization. Statistical Papers, 50(2):407–418, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0073-4.

Nooghabi:2009:EIN

[1542] H. Jabbari. Nooghabi and H. A.

Azarnoosh. Exponential inequality for negatively associated random variables. Statistical Papers, 50 (2):419–428, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0081-4.

Raqab:2009:DFP

[1543] Mohammad Z. Raqab. Distributionfree prediction intervals for the future current record statistics. Statistical Papers, 50(2):429–439, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0082-3.

Nadarajah:2009:CPH

[1544] Saralees Nadarajah. Comment on the paper by A. H. Joarder. Statistical Papers, 50(2):441–443, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        007-0110-3.pdf. See [1458].

Ligges:2009:BRC

[1545] Uwe Ligges. Book review: Crawley, M. J. (2007) The R Book. Statistical Papers, 50(2):445–446, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-008-0118-3.

Maronna:2009:BRA

[1546] Ricardo Maronna. Book review: Antony Unwin, Martin Theus and Heike Hofmann: Graphics of Large Datasets: Visualizing a Million. Statistical Papers, 50(2):447–448, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-008-0120-9.

Polasek:2009:BRG

[1547] Wolfgang Polasek. Book review: Ghosh, Jayanta K., Delampady, Mohan, Samanta, Tapas: An introduction to Bayesian analysis, theory and methods. Statistical Papers, 50(2):449–450, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-008-0121-8.

Anonymous:2009:HCb

[1548] Anonymous. Help & contacts. Statistical Papers, 50(2):??, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wencheko:2009:EVW

[1549] Eshetu Wencheko and Honest W.

Chipoyera. Estimation of the variance when kurtosis is known. Statistical Papers, 50(3):455–464, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0084-1.

Deniz:2009:RBB

[1550] Emilio Gómez Déniz, José María Sarabia, and F. José Vázquez Polo. Robust Bayesian bonus–malus premiums under the conditional specification model. Statistical Papers, 50(3):465–480, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0085-0.

Horny:2009:IMP

[1551] Guillaume Horny. Inference in mixed proportional hazard models with K random effects. Statistical Papers, 50(3):481–499, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0087-y.

Arias-Nicolas:2009:LFR

[1552] J. P. Arias-Nicolás, J. Martín, and A. Suárez-Llorens. L_{p} loss functions: a robust Bayesian approach. Statistical Papers, 50(3):501–509, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0093-0.

Khan:2009:OTR

[1553] Shahjahan Khan. Optimal tolerance regions for future regression vector and residual sum of squares of multiple regression model with multivariate spherically contoured errors. Statistical Papers, 50(3):511–525, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0095-y.

Nkurunziza:2009:TIS

[1554] Sévérien Nkurunziza. Testing interaction in some predator–prey populations. Statistical Papers, 50 (3):527–551, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0096-x.

Kotz:2009:AQM

[1555] Samuel Kotz and Edith Seier. An analysis of quantile measures of kurtosis: center and tails. Statistical Papers, 50(3):553–568, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0101-4.

Meintanis:2009:UAT

[1556] Simos G. Meintanis. A unified approach of testing for discrete and continuous Pareto laws. Statistical Papers, 50(3):569–580, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0103-2.

Joarder:2009:MPR

[1557] Anwar H. Joarder. Moments of the product and ratio of two correlated chisquare variables. Statistical Papers, 50(3):581–592, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-007-0105-0.pdf.

Bodnar:2009:SIE

[1558] Taras Bodnar, Wolfgang Schmid, and Taras Zabolotskyy. Statistical inference of the efficient frontier for dependent asset returns. Statistical Papers, 50(3):593–604, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0108-x.

Nadarajah:2009:PDD

[1559] Saralees Nadarajah. The product t density distribution arising from the product of two Student’s t PDFs. Statistical Papers, 50(3):605–615, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0088-x.

Nadarajah:2009:UMC

[1560] Saralees Nadarajah. Useful moment and CDF formulations for the COM– Poisson distribution. Statistical Papers, 50(3):617–622, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0089-9.

Hwang:2009:AOT

[1561] Leng-Cheng Hwang and Jeng-Fu Liu. Asymptotic optimality of a two-stage procedure in Bayes sequential estimation. Statistical Papers, 50(3):623–631, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0092-1.

Kala:2009:ESP

[1562] Radosław Kala and Paweł Pordzik. Estimation in singular partitioned, reduced or transformed linear models. Statistical Papers, 50(3):633–638, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0097-9.

Yang:2009:ASR

[1563] Hu Yang and Jianwen Xu. An alternative stochastic restricted Liu estimator in linear regression. Statistical Papers, 50(3):639–647, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-007-0102-3.

Bachmaier:2009:CSC

[1564] Martin Bachmaier. A confidence set for that x-coordinate where a quadratic regression model has a given gradient. Statistical Papers, 50(3):649–660, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0104-1.

Diana:2009:ESP

[1565] Giancarlo Diana and Pier Francesco Perri. Estimating a sensitive proportion through randomized response procedures based on auxiliary information. Statistical Papers, 50(3):661–672, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0107-y.

Polasek:2009:BRP

[1566] Wolfgang Polasek. Book review:

Phillip I. Good (2005): Introduction to Statistics Through Resampling Methods and R/S-PLUS (Paperback); Phillip I. Good (2005): Introduction to Statistics Through Resampling Methods and Microsoft Office Excel (Paperback). Statistical Papers, 50(3):673–675, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0131-6.

Hanck:2009:BRU

[1567] Christoph Hanck. Book review: Uwe Hassler (2007): Stochastische Integration und Zeitreihenmodellierung. Statistical Papers, 50(3):677–679, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
008-0143-2.pdf.

Schwender:2009:BRS

[1568] Holger Schwender. Book review: Sandrine Dudoit, Mark J. van der Laan (2008): Multiple Testing Procedures with Applications to Genomics. Statistical Papers, 50(3):681–682, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-008-0144-1.

Anonymous:2009:HCc

[1569] Anonymous. Help & contacts. Statistical Papers, 50(3):??, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Markiewicz:2009:PPI

[1570] Augustyn Markiewicz and Götz Trenkler. Preface to the proceedings of the International Conference on Trends and Perspectives in Linear Statistical Inference LINSTAT’2008 in Celebration of Tadeusz Caliński’s 80th Birthday. Statistical Papers, 50(4):693–695, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        009-0245-5.pdf.

Aastveit:2009:ICT

[1571] Are Halvor Aastveit, Trygve Almøy, Iwona Mejza, and Stanislaw Mejza. Individual control treatment in splitplot experiments. Statistical Papers, 50(4):697–710, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0253-5.

Alin:2009:CPA

[1572] Aylin Alin. Comparison of PLS algorithms when number of objects is much larger than number of variables. Statistical Papers, 50(4):711–720, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0251-7.

Baksalary:2009:POA

[1573] Oskar Maria Baksalary and Götz Trenkler. A projector oriented approach to the best linear unbiased estimator. Statistical Papers, 50(4):721–733, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0252-6.

Calinski:2009:MMA

[1574] T. Caliński, S. Czajka, Z. Kaczmarek, P. Krajewski, and W. Pilarczyk. A mixed model analysis of variance for multi-environment variety trials. Statistical Papers, 50(4):735–759, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0249-1.

Filipiak:2009:ODU

[1575] Katarzyna Filipiak, Augustyn Markiewicz, and Anna Szczepańska. Optimal designs under a multivariate linear model with additional nuisance parameters. Statistical Papers, 50(4):761–778, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-009-
                                        0250-8.

Filipiak:2009:CCB

[1576] Katarzyna Filipiak and Rafal Rózański. Connectedness of complete block designs under an interference model. Statistical Papers, 50(4):779–787, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0247-3.

Graczyk:2009:ROD

[1577] Małgorzata Graczyk. Regular Aoptimal design matrices X = (x_{ij}) with x_{ij} = −1, 0, 1. Statistical Papers, 50 (4):789–795, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0248-2.

Hanusz:2009:ERP

[1578] Zofia Hanusz and Monika Krajewska. Estimation of a relative potency of two preparations with correlated responses. Statistical Papers, 50(4):797–804, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0246-4.

Kala:2009:NBD

[1579] Radosław Kala. On nested block designs geometry. Statistical Papers, 50 (4):805–815, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0260-6.

Krzysko:2009:DAM

[1580] Mirosław Krzyśko and Michal Skorzybut. Discriminant analysis of multivariate repeated measures data with a Kronecker product structured covariance matrices. Statistical Papers, 50 (4):817–835, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0259-z.

Lacka:2009:SOB

[1581] Agnieszka Lacka, Maria Kozlowska, and Jan Kozlowski. Some optimal block designs with nested rows and columns for research on alternative methods of limiting slug damage. Statistical Papers, 50(4):837–846, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0258-0.

Markiewicz:2009:ALS

[1582] Augustyn Markiewicz and Simo Puntanen. Admissibility and linear sufficiency in linear model with nuisance parameters. Statistical Papers, 50 (4):847–854, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0256-2.

Michalski:2009:LOI

[1583] Andrzej Michalski. On locally optimal invariant unbiased tests for the variance components ratio in mixed linear models. Statistical Papers, 50 (4):855–868, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0257-1.

Moreira:2009:MMR

[1584] Elsa Moreira, João Tiago Mexia, Miguel Fonseca, and Roman Zmyślony. L models and multiple regressions designs. Statistical Papers, 50(4): 869–885, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0255-3.

Niemiro:2009:RCE

[1585] Wojciech Niemiro and Wojciech Rejchel. Rank correlation estimators and their limiting distributions. Statistical Papers, 50(4):887–893, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0263-3.

Rodrigues:2009:AEU

[1586] Paulo C. Rodrigues and Ana T. Lima. Analysis of an European union election using principal component analysis. Statistical Papers, 50(4):895–904, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0264-2.

Rasch:2009:TAM

[1587] Dieter Rasch, Thomas Rusch, Marie Simecková, Klaus D. Kubinger, Karl Moder, and Petr Simecek. Tests of additivity in mixed and fixed effect twoway ANOVA models with single subclass numbers. Statistical Papers, 50 (4):905–916, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0254-4.

Styan:2009:SCL

[1588] George P. H. Styan, Christian Boyer, and Ka Lok Chu. Some comments on Latin squares and on Graeco– Latin squares, illustrated with postage stamps and old playing cards. Statistical Papers, 50(4):917–941, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0261-5.

Volaufova:2009:HAO

[1589] Julia Volaufova. Heteroscedastic ANOVA: old p values, new views. Statistical Papers, 50(4):943–962, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0262-4.

Anonymous:2009:HCd

[1590] Anonymous. Help & contacts. Statistical Papers, 50(4):??, August 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Yan:2010:NBR

[1591] Guofen Yan and J. Sedransk. A note on Bayesian residuals as a hierarchical model diagnostic technique. Statistical Papers, 51(1):1–10, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-007-
                                        0111-2.

Demetrescu:2010:DFT

[1592] Matei Demetrescu. On the Dickey– Fuller test with white standard errors. Statistical Papers, 51(1):11–25, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0112-1.

Diab:2010:TNU

[1593] L. S. Diab. Testing for NBUL using goodness of fit approach with applications. Statistical Papers, 51(1):27–40, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-

0113-0. See comments [1728].

Fischer:2010:GTT

[1594] Matthias Fischer. Generalized Tukeytype distributions with application to financial and teletraffic data. Statistical Papers, 51(1):41–56, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0114-z.

Wang:2010:KTS

[1595] Lihong Wang. Kernel type smoothed quantile estimation under long memory. Statistical Papers, 51(1):57–67, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-007-
                                        0115-y.

Opsomer:2010:FLD

[1596] J. D. Opsomer and M. Francisco-Fernández. Finding local departures from a parametric model using nonparametric regression. Statistical Papers, 51(1):69–84, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-007-
0116-x.

Jouini:2010:BMS

[1597] Jamel Jouini. Bootstrap methods for single structural change tests: power versus corrected size and empirical illustration. Statistical Papers, 51 (1):85–109, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0123-6.

Ospina:2010:IBD

[1598] Raydonal Ospina and Silvia L. P. Ferrari. Inflated beta distributions. Statistical Papers, 51(1):111–126, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0125-4.

Kozubowski:2010:MLD

[1599] Tomasz J. Kozubowski and Saralees Nadarajah. Multitude of Laplace distributions. Statistical Papers, 51 (1):127–148, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0127-2.

Barakat:2010:CBR

[1600] H. M. Barakat, E. M. Nigm, and Magdy E. El-Adll. Comparison between the rates of convergence of extremes under linear and under power normalization. Statistical Papers, 51 (1):149–164, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0128-1.

Rychlik:2010:EGO

[1601] Tomasz Rychlik. Evaluations of generalized order statistics from bounded populations. Statistical Papers, 51 (1):165–177, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0129-0.

Kaminska:2010:EBR

[1602] Agnieszka Kamińska. The equivalence of Bayes and robust Bayes estimators for various loss functions. Statistical Papers, 51(1):179–191, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-008-
                                        0130-7.

Cifarelli:2010:GSP

[1603] D. Michele Cifarelli, R. P. Gupta, and K. Jayakumar. On generalized semi-Pareto and semi-Burr distributions and random coefficient minification processes. Statistical Papers, 51 (1):193–208, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0132-5.

Parchami:2010:FVT

[1604] Abbas Parchami, S. Mahmoud Taheri, and Mashaallah Mashinchi. Fuzzy p-value in testing fuzzy hypotheses with crisp data. Statistical Papers, 51 (1):209–226, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0133-4.

Bairamov:2010:LRO

[1605] Ismihan Bairamov, Alexandre Berred, and Alexei Stepanov. Limit results for ordered uniform spacings. Statistical Papers, 51(1):227–240, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0134-3.

Anonymous:2010:HCa

[1606] Anonymous. Help & contacts. Statistical Papers, 51(1):??, January 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Waldl:2010:SCS

[1607] Helmut Waldl. Some comments on Statistical Papers 46, 270–395 (2005). Statistical Papers, 51(2):241–246, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0332-7.

Withers:2010:ELD

[1608] Christopher S. Withers and Saralees Nadarajah. Expansions for log densities of asymptotically normal estimates. Statistical Papers, 51 (2):247–257, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0135-2.

Tsai:2010:ASP

[1609] Tzong-Ru Tsai and Chin-Wei Lin. Acceptance sampling plans under progressive interval censoring with likelihood ratio. Statistical Papers, 51(2):259–271, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0141-4.

Cysneiros:2010:TBT

[1610] Audrey H. M. A. Cysneiros, Katya S. P. Rodrigues, Gauss M. Cordeiro, and Silvia L. P. Ferrari. Three Bartletttype corrections for score statistics in symmetric nonlinear regression models. Statistical Papers, 51(2):273–284, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0158-8.

Bachmaier:2010:TCS

[1611] Martin Bachmaier. Test and confidence set for the difference of the x-coordinates of the vertices of two quadratic regression models. Statistical Papers, 51(2):285–296, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0159-7.

Klufa:2010:ECD

[1612] J. Klufa. Exact calculation of the Dodge–Romig LTPD single sampling plans for inspection by variables. Statistical Papers, 51(2):297–305, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0160-1.

Abdelkader:2010:CMO

[1613] Yousry H. Abdelkader. Computing the moments of order statistics from independent nonidentically distributed beta random variables. Statistical Papers, 51(2):307–313, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0161-0.

Li:2010:NSM

[1614] Yalian Li and Hu Yang. A new stochastic mixed ridge estimator in linear regression model. Statistical Papers, 51(2):315–323, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0169-5.

Caroni:2010:TMO

[1615] Chrys Caroni. Testing for the Marshall–Olkin extended form of the Weibull distribution. Statistical Papers, 51(2):325–336, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0172-x.

Martinelli:2010:NCR

[1616] Andrea Martinelli, Matteo Ruggiero, and Stephen G. Walker. A note on convergence rates for posterior distributions via large deviations techniques. Statistical Papers, 51(2):337–347, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-008-
                                        0180-x.

Saha:2010:MUQ

[1617] Amitava Saha. A modified unrelated question randomized response device for complex surveys. Statistical Papers, 51(2):349–355, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0182-8.

Tabakan:2010:DBR

[1618] Gülin Tabakan and Fikri Akdeniz. Difference-based ridge estimator of parameters in partial linear model. Statistical Papers, 51(2):357–368, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-008-
                                        0192-6.

Wang:2010:CSD

[1619] Yashi Wang, Weiwei Zhuang, and Taizhong Hu. Conditionally stochastic domination of generalized order statistics from two samples. Statistical Papers, 51(2):369–373, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-008-
0194-4.

Kim:2010:ESP

[1620] Chansoo Kim and Keunhee Han. Estimation of the scale parameter of the half-logistic distribution under progressively type II censored sample. Statistical Papers, 51(2):375–387, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0197-9.

Kurata:2010:TCM

[1621] Hiroshi Kurata. A theorem on the covariance matrix of a generalized least squares estimator under an elliptically symmetric error. Statistical Papers, 51(2):389–395, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0199-7.

Zhao:2010:RTV

[1622] Wenzhi Zhao, Zheng Tian, and Zhiming Xia. Ratio test for variance change point in linear process with long memory. Statistical Papers, 51(2):397–407, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0202-3.

Poursaeed:2010:NMP

[1623] M. H. Poursaeed. A note on the mean past and the mean residual life of a (n − k + 1)-out-of-n system under multi monitoring. Statistical Papers, 51(2):409–419, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0207-y.

Fallah:2010:BRA

[1624] Afshin Fallah and Mohsen Mohammadzadeh. Bayesian regression analysis with linked data using mixture normal distributions. Statistical Papers, 51(2):421–430, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0208-x.

Wu:2010:OPG

[1625] Shuo-Jye Wu and Syuan-Rong Huang. Optimal progressive group-censoring plans for exponential distribution in presence of cost constraint. Statistical Papers, 51(2):431–443, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0212-1.

Rong:2010:MDM

[1626] Jian-Ying Rong and Xu-Qing Liu. On misspecification of the dispersion matrix in mixed linear models. Statistical Papers, 51(2):445–453, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0213-0.

Anonymous:2010:HCb

[1627] Anonymous. Help & contacts. Statistical Papers, 51(2):??, June 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Niemiro:2010:SCE

[1628] Wojciech Niemiro and Jacek Wesolowski.

Synthetic and composite estimation under a superpopulation model. Statistical Papers, 51(3):497–509, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0136-1.

Baran:2010:OSE

[1629] Jerzy Baran and Ryszard Magiera.

Optimal sequential estimation procedures of a function of a probability of success under LINEX loss. Statistical Papers, 51(3):511–529, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0137-0.

Lachos:2010:BAS

[1630] Victor H. Lachos, Vicente G. Cancho, and Reiko Aoki. Bayesian analysis of skew-t multivariate null intercept measurement error model. Statistical Papers, 51(3):531–545, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0138-z.

Cancho:2010:NRM

[1631] Vicente G. Cancho, Víctor H. Lachos, and Edwin M. M. Ortega. A nonlinear regression model with skewnormal errors. Statistical Papers, 51 (3):547–558, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0139-y.

Singh:2010:EMP

[1632] Housila P. Singh and Sunil Kumar. Estimation of mean in presence of non-response using two phase sampling scheme. Statistical Papers, 51 (3):559–582, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0140-5.

Kim:2010:BEP

[1633] Chansoo Kim and Seongho Song. Bayesian estimation of the parameters of the generalized exponential distribution from doubly censored samples. Statistical Papers, 51(3): 583–597, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0142-3.

Bebbington:2010:LEB

[1634] Mark Bebbington, Chin-Diew Lai, and Ricardas Zitikis. Life expectancy of a bathtub shaped failure distribution. Statistical Papers, 51(3): 599–612, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0148-x.

Bousquet:2010:EVP

[1635] Nicolas Bousquet. Eliciting vague but proper maximal entropy priors in Bayesian experiments. Statistical Papers, 51(3):613–628, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0149-9.

Zhou:2010:SIF

[1636] Haibo Zhou, Jinhong You, and Bin Zhou. Statistical inference for fixedeffects partially linear regression models with errors in variables. Statistical Papers, 51(3):629–650, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0150-3.

Alkhamisi:2010:SSN

[1637] M. A. Alkhamisi. Simulation study of new estimators combining the SUR ridge regression and the restricted least squares methodologies. Statistical Papers, 51(3):651–672, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0151-2.

Cozzucoli:2010:SCI

[1638] Paolo C. Cozzucoli. Simultaneous confidence intervals on partial means of classes in the two-stage stratified sampling. Statistical Papers, 51(3): 673–685, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0152-1.

Durante:2010:MNE

[1639] Fabrizio Durante, Erich Peter Klement, Carlo Sempi, and Manuel Úbeda-Flores. Measures of nonexchangeability for bivariate random vectors. Statistical Papers, 51(3): 687–699, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0153-0.

Montenegro:2010:ISE

[1640] Lourdes C. Montenegro, Víctor H. Lachos, and Heleno Bolfarine. Inference for a skew extension of the Grubbs model. Statistical Papers, 51 (3):701–715, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0157-9.

Shalabh:2010:CER

[1641] Shalabh, Gaurav Garg, and Neeraj Misra. Consistent estimation of regression coefficients in ultrastructural measurement error model using stochastic prior information. Statistical Papers, 51(3):717–748, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0162-z.

Maronna:2010:BRD

[1642] Ricardo Maronna. Book review: David Skillicorn (2007): Understanding Complex Datasets: Data Mining with Matrix Decompositions. Statistical Papers, 51(3):749, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-008-0147-y.pdf.

Fischer:2010:BRJ

[1643] Matthias Fischer. Book review:

Jiti Gao: Nonlinear time series — semiparametric and nonparametric methods. Statistical Papers, 51 (3):751, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-008-0154-z.pdf.

Fischer:2010:HMT

[1644] Matthias Fischer. Hendrik Madsen (2007): Time series analysis. Statistical Papers, 51(3):753–754, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-008-0155-y.

Lenz:2010:BRG

[1645] Sylvia Tamara Lenz. Book review:

Gönen, Mithat (2007): Analyzing Receiver Operating Characteristic Curves with SAS. Statistical Papers, 51(3): 755–756, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-008-0156-
x.

Anonymous:2010:HCc

[1646] Anonymous. Help & contacts. Statistical Papers, 51(3):??, September 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Kibria:2010:PTE

[1647] B. M. Golam Kibria and A. K.

Md. E. Saleh. Preliminary test estimation of the parameters of exponential and Pareto distributions for censored samples. Statistical Papers, 51 (4):757–773, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0163-y.

BuHamra:2010:IES

[1648] Sana S. BuHamra, Noriah M. Al-Kandari, and S. E. Ahmed. Inference on effect size indices from several two-armed experiments. Statistical Papers, 51(4):775–787, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0165-9.

Frahm:2010:LSI

[1649] Gabriel Frahm. Linear statistical inference for global and local minimum variance portfolios. Statistical Papers, 51 (4):789–812, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0170-z.

Cao:2010:HAA

[1650] Chun-Zheng Cao, Jin-Guan Lin, and Li-Xing Zhu. Heteroscedasticity and/ or autocorrelation diagnostics in nonlinear models with AR(1) and symmetrical errors. Statistical Papers, 51 (4):813–836, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0171-y.

Jose:2010:MOW

[1651] K. K. Jose, Shanoja R. Naik, and Miroslav M. Ristić. Marshall–Olkin q-Weibull distribution and max-min processes. Statistical Papers, 51(4): 837–851, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0173-9.

Lai:2010:BCT

[1652] Dejian Lai. Box–Cox transformation for spatial linear models: a study on lattice data. Statistical Papers, 51(4): 853–864, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0178-4.

Arslan:2010:AMS

[1653] Olcay Arslan. An alternative multivariate skew Laplace distribution: properties and estimation. Statistical Papers, 51(4):865–887, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0183-7.

Laheetharan:2010:IEP

[1654] A. Laheetharan and P. Wijekoon.

Improved estimation of the population parameters when some additional information is available. Statistical Papers, 51(4):889–914, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0185-5.

Nosek:2010:SIC

[1655] Konrad Nosek. Schwarz information criterion based tests for a changepoint in regression models. Statistical Papers, 51(4):915–929, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-008-
                                        0186-4.

Pham-Gia:2010:EED

[1656] T. Pham-Gia and N. Turkkan. Exact expression of the density of the sample generalized variance and applications. Statistical Papers, 51(4): 931–945, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-008-0187-3. See erra-

tum [1729].

Hanagal:2010:MHB

[1657] David D. Hanagal. Modeling heterogeneity for bivariate survival data by the Weibull distribution. Statistical Papers, 51(4):947–958, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-008-
                                        0188-2.

Demir:2010:RSS

[1658] Sevcan Demir and Serkan Eryilmaz. Run statistics in a sequence of arbitrarily dependent binary trials. Statistical Papers, 51(4):959–973, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-008-
0191-7.

Singh:2010:EPP

[1659] Housila P. Singh and Sunil Kumar.

Estimation of population product in presence of non-response in successive sampling. Statistical Papers, 51(4): 975–996, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0193-5.

Drygas:2010:BRD

[1660] Hilmar Drygas. Book review: Douglas C. Montgomory, Elizabeth A. Peck, G. Geoffrey Vining (2006): An Introduction to Linear Regression Analysis. Statistical Papers, 51(4):997–998, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-008-0164-x.

Kaiser:2010:BRJ

[1661] Jonas Kaiser. Book review: Jim Albert and Ruud H. Koning (editors): Statistical thinking in sports. Statistical Papers, 51(4):999–1000, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-008-0166-8.

Bornkamp:2010:ADA

[1662] Björn Bornkamp. Anirban DasGupta:

Asymptotic theory of statistics and probability. Statistical Papers, 51(4): 1001–1003, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-008-0167-7.

Giles:2010:BRK

[1663] David E. Giles. Book review: K. Krishnamoorthy (2006): Handbook of statistical distributions with applications. Statistical Papers, 51(4): 1005–1006, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-008-0168-
6.

Drygas:2010:BRY

[1664] Hilmar Drygas. Book review: Youngjo Lee, John A. Nelder, Yudi Pawitan: Generalized linear models with random effects: unified analysis via Hlikelihood. Statistical Papers, 51(4): 1007–1008, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-008-0175-
7.

Maronna:2010:BRW

[1665] Ricardo Maronna. Book review: Walter W. Piegorsch and A. John Bailer (2005): Analyzing environmental data. Statistical Papers, 51(4): 1009–1010, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-008-0176-
6.

Webel:2010:SAB

[1666] Karsten Webel. Steland, ansgar: Basiswissen statistik, Springer, Berlin, 2007. Statistical Papers, 51(4):1011– 1012, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-008-0177-
5.

Anonymous:2010:HCd

[1667] Anonymous. Help & contacts. Statistical Papers, 51(4):??, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Siburg:2011:SFE

[1668] Karl Friedrich Siburg and Pavel A.

Stoimenov. Symmetry of functions and exchangeability of random variables. Statistical Papers, 52(1):1–15, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-008-
0195-3.

Hong:2011:MIR

[1669] Chong Sun Hong and Beom Jun Kim.

Mutual information and redundancy for categorical data. Statistical Papers, 52(1):17–31, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0196-x.

Ye:2011:TSO

[1670] Ping Ye and Bhaskar Bhattacharya.

Tests of symmetry with one-sided alternatives in three-way contingency tables. Statistical Papers, 52(1): 33–51, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0198-8.

Kim:2011:BEE

[1671] Chansoo Kim, Jinhyouk Jung, and Younshik Chung. Bayesian estimation for the exponentiated Weibull model under type II progressive censoring. Statistical Papers, 52(1):53–70, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0203-2.

Kruse:2011:NUR

[1672] Robinson Kruse. A new unit root test against ESTAR based on a class of modified statistics. Statistical Papers, 52(1):71–85, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0204-1.

Scott:2011:DAF

[1673] Steven L. Scott. Data augmentation, frequentist estimation, and the Bayesian analysis of multinomial logit models. Statistical Papers, 52(1): 87–109, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0205-0.

Chaudhuri:2011:ESPa

[1674] Arijit Chaudhuri, Mausumi Bose, and Kajal Dihidar. Estimating sensitive proportions by Warner’s randomized response technique using multiple randomized responses from distinct persons sampled. Statistical Papers, 52 (1):111–124, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0210-3.

Gupta:2011:BOR

[1675] Ramesh C. Gupta. Bivariate odds ratio and association measures. Statistical Papers, 52(1):125–138, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0211-2.

Gupta:2011:NCB

[1676] Arjun K. Gupta, Johanna Marcela Orozco-Castañeda, and Daya K. Nagar. Non-central bivariate beta distribution. Statistical Papers, 52(1): 139–152, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0215-y.

Ye:2011:GCI

[1677] Rendao Ye, Tiefeng Ma, and Songgui Wang. Generalized confidence intervals for the process capability indices in general random effect model with balanced data. Statistical Papers, 52 (1):153–169, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0216-x.

Laheetharan:2011:MSE

[1678] A. Laheetharan and P. Wijekoon.

Mean square error comparison among variance estimators with known coefficient of variation. Statistical Papers, 52 (1):171–201, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0221-0.

Qin:2011:TEC

[1679] Ruibing Qin, Zheng Tian, and Hao Jin. Truncating estimation for the change in stochastic trend with heavy-tailed innovations. Statistical Papers, 52 (1):203–217, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0223-y.

Rasch:2011:TST

[1680] Dieter Rasch, Klaus D. Kubinger, and Karl Moder. The two-sample t test: pre-testing its assumptions does not pay off. Statistical Papers, 52(1): 219–231, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0224-x.

Maronna:2011:BRT

[1681] Ricardo Maronna. Book review: Takayuki Saito, Hiroshi Yadohisa (2005): Data analysis of asymmetric structures: advanced approaches in computational statistics. Statistical Papers, 52(1):233–234, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-009-0200-5.

Lenz:2011:BRF

[1682] Hans-J. Lenz. Book review: Finn V. Jensen, Thomas D. Nielsen (2007):

Bayesian Networks and Decision Graphs. Statistical Papers, 52(1): 235–237, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0201-4.

Polasek:2011:BRJa

[1683] Wolfgang Polasek. Book review: Jim Albert (2007): Bayesian computation with R. Statistical Papers, 52 (1):239–240, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-009-0206-
z.

Viertl:2011:BRD

[1684] R. Viertl. Book review: N. D.

Singpurwalla (2006): Reliability and risk: a Bayesian perspective. Statistical Papers, 52(1):241, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
009-0217-9.pdf.

Lenz:2011:BRS

[1685] Sylvia Tamara Lenz. Book review:

Shein-Chung Chow, Jun Shao, Hansheng Wang (2008): Sample Size Calculations in Clinical Research, 2nd edition. Statistical Papers, 52(1): 243–244, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-009-0218-

Polasek:2011:BRJb

[1686] Wolfgang Polasek. Book review: James Raymer (ed), Frans Wiilekens (Coeditor): International migration in Europe: data, models and estimates. Statistical Papers, 52(1):245–246, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-009-0225-9.

Polasek:2011:BRN

[1687] Wolfgang Polasek. Book review: Nassim Nicholas Taleb: The black swan: the impact of the highly improbable. Statistical Papers, 52(1): 247–249, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0226-8.

Anonymous:2011:HCa

[1688] Anonymous. Help & contacts. Statistical Papers, 52(1):??, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

You:2011:WDM

[1689] Jinhong You, Xian Zhou, Lixing Zhu, and Bin Zhou. Weighted denoised minimum distance estimation in a regression model with autocorrelated measurement errors. Statistical Papers, 52(2):263–286, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0227-7.

Firinguetti:2011:ACI

[1690] Luis Firinguetti and Gladys Bobadilla.

Asymptotic confidence intervals in ridge regression based on the Edgeworth expansion. Statistical Papers, 52(2):287–307, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0229-5.

Joseph:2011:GDE

[1691] Dhannya P. Joseph. Gamma distribution and extensions by using pathway idea. Statistical Papers, 52(2):309–325, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0231-y.

Kim:2011:LTS

[1692] Min Kim and Bong-Jin Yum. Life test sampling plans for Weibull distributed lifetimes under accelerated hybrid censoring. Statistical Papers, 52(2):327–342, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0233-9.

Chaudhuri:2011:ESPb

[1693] Arijit Chaudhuri, Mausumi Bose, and Kajal Dihidar. Estimation of a sensitive proportion by Warner’s randomized response data through inverse sampling. Statistical Papers, 52(2):343–354, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0234-8.

Yilmaz:2011:CTS

[1694] Mehmet Yilmaz. Convex transformation on survival functions and related dependence concepts. Statistical Papers, 52(2):355–370, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0235-7.

Ciuperca:2011:PLA

[1695] Gabriela Ciuperca. Penalized least absolute deviations estimation for nonlinear model with change-points. Statistical Papers, 52(2):371–390, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0236-6.

Liu:2011:ELD

[1696] Wanrong Liu and Xuewen Lu. Empirical likelihood for density-weighted average derivatives. Statistical Papers, 52(2):391–412, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0237-5.

Zarei:2011:CCW

[1697] H. Zarei and H. Jabbari. Complete convergence of weighted sums under negative dependence. Statistical Papers, 52(2):413–418, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0238-4.

Tsai:2011:EWM

[1698] Tzong-Ru Tsai and Wen-Pin Yen. Exponentially weighted moving average control charts for three-level products. Statistical Papers, 52(2):419–429, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0239-3.

Jamalizadeh:2011:GST

[1699] A. Jamalizadeh, A. R. Arabpour, and N. Balakrishnan. A generalized skew two-piece skew-normal distribution. Statistical Papers, 52(2):431–446, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0240-x.

Sung:2011:SCW

[1700] Soo Hak Sung. On the strong convergence for weighted sums of random variables. Statistical Papers, 52(2):447–454, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0241-9.

Patriota:2011:IML

[1701] Alexandre G. Patriota, Artur J.

Lemonte, and Heleno Bolfarine. Improved maximum likelihood estimators in a heteroskedastic errors-in-variables model. Statistical Papers, 52(2):455–467, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0243-7.

Abtahi:2011:AEV

[1702] A. Abtahi, M. Towhidi, and J. Behboodian. An appropriate empirical version of skew-normal density. Statistical Papers, 52(2):469–489, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0244-6.

Kramer:2011:BRWa

[1703] Walter Krämer. Book review: O. W. Winkler: Interpreting socio-economic data — a foundation of descriptive statistics. Statistical Papers, 52(2):491– 492, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s00362-009-0242-8.

Arnold:2011:BRJ

[1704] Matthias Arnold. Book review: James Le Sage, Robert K. Pace: Introduction to spatial econometrics. Statistical Papers, 52(2):493–494, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-009-0267-z.

Lutkepohl:2011:BRB

[1705] Helmut Lütkepohl. Book review: Bernhard Pfaff (2006): Analysis of Integrated and Cointegrated Time Series with R. Statistical Papers, 52 (2):495–496, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/

article/10.1007/s00362-009-0278-
9.

Lutkepohl:2011:BRG

[1706] Helmut Lütkepohl. Book review: I Gusti Ngurah Agung (2009): Time Series Data Analysis Using EViews. Statistical Papers, 52(2):497–499, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0281-1.

Bucker:2011:BRM

[1707] Michael Bücker. Book review: Michael J. Daniels, Joseph W. Hogan: Missing data in longitudinal studies. Statistical Papers, 52(2):501–502, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-009-0282-0.

Hanck:2011:BRJ

[1708] Christoph Hanck. Book review: Joshua D. Angrist and Jörn–Steffen Pischke (2009): Mostly Harmless Econometrics: An Empiricist’s Companion. Statistical Papers, 52(2):503–504, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
009-0284-y.pdf.

Anonymous:2011:HCb

[1709] Anonymous. Help & contacts. Statistical Papers, 52(2):??, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hu:2011:GMD

[1710] Yonggang Hu, Yong Wang, Yi Wu, Qiang Li, and Chenping Hou. Generalized Mahalanobis depth in the reproducing kernel Hilbert space. Statistical Papers, 52(3):511–522, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0265-1.

Kayid:2011:PPM

[1711] M. Kayid. Preservation properties of the moment generating function ordering of residual lives. Statistical Papers, 52(3):523–529, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0266-0.

Makri:2011:RLE

[1712] Frosso S. Makri and Zaharias M. Psillakis. On runs of length exceeding a threshold: normal approximation. Statistical Papers, 52(3):531–551, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0268-y.

Vilca-Labra:2011:HTS

[1713] Filidor Vilca-Labra, Reiko Aoki, and Camila Borelli Zeller. Hypotheses testing for structural calibration model. Statistical Papers, 52(3):553–565, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0269-x.

Lachos:2011:ELI

[1714] V. H. Lachos, T. Angolini, and C. A.

Abanto-Valle. On estimation and local influence analysis for measurement errors models under heavy-tailed distributions. Statistical Papers, 52 (3):567–590, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0270-4.

deGusmao:2011:GIW

[1715] Felipe R. S. de Gusmão, Edwin M. M.

Ortega, and Gauss M. Cordeiro. The generalized inverse Weibull distribution. Statistical Papers, 52(3):591–619, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0271-3.

Liu:2011:MMC

[1716] Shuangzhe Liu, Chris C. Heyde, and Wing-Keung Wong. Moment matrices in conditional heteroskedastic models under elliptical distributions with applications in AR–ARCH models. Statistical Papers, 52(3):621–632, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0272-2.

Diana:2011:CEQ

[1717] Giancarlo Diana and Pier Francesco Perri. A class of estimators for quantitative sensitive data. Statistical Papers, 52(3):633–650, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-009-0273-1.

Ahsanullah:2011:CGD

[1718] Mohammad Ahsanullah and Fazil Aliev. A characterization of geometric distribution based on weak records. Statistical Papers, 52(3):651–655, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0274-0.

Lee:2011:CMR

[1719] Wen-Chuan Lee, Jong-Wuu Wu, and Chun-Te Li. Characterization of the mixtures of Rayleigh distributions by conditional expectation of order statistics. Statistical Papers, 52(3):657–675, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0275-z.

Costantini:2011:NAD

[1720] Mauro Costantini and Stephan Popp. A note on the asymptotic distribution of a Perron-type innovational outlier unit root test with a break. Statistical Papers, 52(3):677–682, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-009-
                                        0276-y.

Abe:2011:SSC

[1721] Toshihiro Abe and Arthur Pewsey. Sine-skewed circular distributions. Statistical Papers, 52(3):683–707, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0277-x.

Dakovic:2011:CPI

[1722] Rada Dakovic and Claudia Czado.

Comparing point and interval estimates in the bivariate t-copula model with application to financial data. Statistical Papers, 52(3):709–731, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0279-8.

Maronna:2011:AJI

[1723] Ricardo Maronna. Alan Julian izenman (2008): Modern multivariate statistical techniques: Regression, classification and manifold learning. Statistical Papers, 52(3):733–734, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-009-0291-z.

Wied:2011:BRP

[1724] Dominik Wied. Book review: Peter W. Jones and Peter Smith, Stochastic Processes: An Introduction. Statistical Papers, 52(3):735–736, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-009-0295-8.

Stoimenov:2011:BRP

[1725] Pavel Stoimenov. Book review:

Philippe Jorion, Value at Risk, 3rd Ed:

The New Benchmark for Managing Financial Risk. Statistical Papers, 52 (3):737–738, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-009-0296-
7.

Pardo:2011:BRP

[1726] J. A. Pardo. Book review: Paolo Giudici and Silvia Figini: Applied data mining for business and industry (Second Edition). Statistical Papers, 52 (3):739–740, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-010-0310-
0.

Kramer:2011:BRA

[1727] Walter Krämer. Book review: Andrew Gelman and Jennifer Hill: Data analysis using regression and multilevel/ hierarchical models. Statistical Papers, 52(3):741–742, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-010-0311-
z.

Anis:2011:CTN

[1728] M. Z. Anis. Comments on “Testing for NBUL using goodness of fit approach with applications” (Statistical Papers (2010) 51: 27–40, DOI 10.1007/s00362-007-0113-0). Statistical Papers, 52(3):743–747, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00362-010-

0322-9. See [1593].

Pham-Gia:2011:EEE

[1729] T. Pham-Gia and N. Turkkan. Erratum to: Exact expression of the density of the sample generalized variance and applications. Statistical Papers, 52(3):749, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
010-0319-4.pdf. See [1656].

Anonymous:2011:HCc

[1730] Anonymous. Help & contacts. Statistical Papers, 52(3):??, August 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Aneiros-Perez:2011:AEP

[1731] Germán Aneiros-Pérez and Philippe Vieu. Automatic estimation procedure in partial linear model with functional data. Statistical Papers, 52(4): 751–771, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0280-2.

Greco:2011:RIS

[1732] Luca Greco and Laura Ventura. Robust inference for the stress-strength reliability. Statistical Papers, 52(4): 773–788, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0286-9.

Jose:2011:MOB

[1733] K. K. Jose, Miroslav M. Ristić, and Ancy Joseph. Marshall–Olkin bivariate Weibull distributions and processes. Statistical Papers, 52(4): 789–798, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0287-8.

Faraz:2011:USC

[1734] Alireza Faraz and Erwin Saniga. A unification and some corrections to Markov chain approaches to develop variable ratio sampling scheme control charts. Statistical Papers, 52(4): 799–811, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0288-7.

Fokoue:2011:BVT

[1735] Ernest Fokoue and Bertrand Clarke. Bias-variance trade-off for prequential model list selection. Statistical Papers, 52(4):813–833, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0289-6.

Ahmad:2011:FMB

[1736] Khalaf E. Ahmad, Zeinhum F. Jaheen, and Heba S. Mohammed. Finite mixture of Burr type XII distribution and its reciprocal: properties and applications. Statistical Papers, 52 (4):835–845, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0290-0.

Hassler:2011:DCS

[1737] Uwe Hassler and Jan Scheithauer. Detecting changes from short to long memory. Statistical Papers, 52(4): 847–870, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0292-y.

Castillo:2011:EBS

[1738] Nabor O. Castillo, Héctor W. Gómez, and Heleno Bolfarine. Epsilon Birnbaum–Saunders distribution family: properties and inference. Statistical Papers, 52(4):871–883, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0293-x.

Sheikhi:2011:RAU

[1739] Ayyub Sheikhi and Ahad Jamalizadeh. Regression analysis using order statistics. Statistical Papers, 52 (4):885–892, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0294-9.

Nair:2011:CDT

[1740] K. R. Muraleedharan Nair, P. G.

Sankaran, and S. Smitha. Chernoff distance for truncated distributions. Statistical Papers, 52(4):893–909, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0297-6.

Tian:2011:SEE

[1741] Yongge Tian and Jieping Zhang. Some equalities for estimations of partial coefficients under a general linear regression model. Statistical Papers, 52(4): 911–920, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0298-5.

Al-Mutairi:2011:NBD

[1742] D. K. Al-Mutairi, M. E. Ghitany, and D. Kundu. A new bivariate distribution with weighted exponential marginals and its multivariate generalization. Statistical Papers, 52(4): 921–936, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-009-0300-2.

Kavousi:2011:SAA

[1743] Amir Kavousi, Mohammad Reza Meshkani, and Mohsen Mohammadzadeh. Spatial analysis of automultivariate lattice data. Statistical Papers, 52(4):937–952, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0302-0.

Uemukai:2011:SSP

[1744] Ryo Uemukai. Small sample properties of a ridge regression estimator when there exist omitted variables. Statistical Papers, 52(4):953–969, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-009-
0303-z.

VanBochove:2011:EUR

[1745] Cornelis A. Van Bochove. Expectation of a uniform random variable with uniform observation errors after selection of the highest observations. Statistical Papers, 52(4):971–977, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
009-0304-y.pdf.

Kramer:2011:BRWb

[1746] Walter Krämer. Book review: Wojtek J. Krzanowski and David J. Hand: ROC curves for continuous data. Statistical Papers, 52(4):979–980, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-010-0312-y.

Maronna:2011:RBS

[1747] Ricardo Maronna. Richard berk: Statistical learning from a regression perspective. Statistical Papers, 52(4): 981–982, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-010-0313-
x.

Webel:2011:BRG

[1748] Karsten Webel. Book review: Greene, W. H., Econometric analysis. Statistical Papers, 52(4):983–984, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-010-0315-8.

Bornkamp:2011:BRG

[1749] Björn Bornkamp. Book review: G. Parmigiani and L. Inoue: Decision theory-principles and approaches. Statistical Papers, 52(4):985–986, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-010-0318-5.

Polasek:2011:NPR

[1750] Wolfgang Polasek. New publications on R. Statistical Papers, 52(4): 987–988, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-010-0323-
                                        8.

Anonymous:2011:HCd

[1751] Anonymous. Help & contacts. Statistical Papers, 52(4):??, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wied:2012:CKD

[1752] Dominik Wied and Rafael Weißbach. Consistency of the kernel density estimator: a survey. Statistical Papers, 53(1):1–21, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0338-1.

Gurler:2012:RLS

[1753] S. Gurler. On residual lifetimes in sequential (n − k + 1)-out-of-n systems. Statistical Papers, 53(1):23–31, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0305-x.

Jonsson:2012:WDH

[1754] Robert Jonsson. When does Heckman’s two-step procedure for censored data work and when does it not? Statistical Papers, 53(1):33–49, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0306-9.

Jozani:2012:BRB

[1755] Mohammad Jafari Jozani, Éric Marchand, and Ahmad Parsian. Bayesian and robust Bayesian analysis under a general class of balanced loss functions. Statistical Papers, 53(1):51–60, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0307-8.

Marozzi:2012:CTD

[1756] Marco Marozzi. A combined test for differences in scale based on the interquantile range. Statistical Papers, 53(1):61–72, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0308-7.

Sung:2012:CCW

[1757] Soo Hak Sung. Complete convergence for weighted sums of negatively dependent random variables. Statistical Papers, 53(1):73–82, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-010-
                                        0309-6.

Zi:2012:TSE

[1758] Xuemin Zi, Changliang Zou, and Yukun Liu. Two-sample empirical likelihood method for difference between coefficients in linear regression model. Statistical Papers, 53(1):83–93, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-010-
                                        0314-9.

Razmkhah:2012:CTS

[1759] M. Razmkhah, H. Morabbi, and J. Ahmadi. Comparing two sampling schemes based on entropy of record statistics. Statistical Papers, 53 (1):95–106, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0316-7.

Hansohm:2012:CAG

[1760] Jürgen Hansohm and Xiaomi Hu. A convergent algorithm for a generalized multivariate isotonic regression problem. Statistical Papers, 53(1): 107–115, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0317-6.

Genc:2012:MOS

[1761] Ali I. Genç. Moments of order statistics of Topp–Leone distribution. Statistical Papers, 53(1):117–131, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0320-y.

Emura:2012:MND

[1762] Takeshi Emura and Yoshihiko Konno.

Multivariate normal distribution approaches for dependently truncated data. Statistical Papers, 53(1): 133–149, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0321-x.

Wilczynski:2012:MPL

[1763] Maciej Wilczyński. Minimax prediction in the linear model with a relative squared error. Statistical Papers, 53 (1):151–164, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0325-6.

Cancho:2012:BAC

[1764] Vicente G. Cancho, Mário de Castro, and Josemar Rodrigues. A Bayesian analysis of the Conway– Maxwell-Poisson cure rate model. Statistical Papers, 53(1):165–176, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00362-010-
0326-5.

Bernadic:2012:DTF

[1765] Milenko Bernadic and José Candel. The doubly truncated function of indices on discrete distributions. Statistical Papers, 53(1):177–193, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0327-4.

Hosseinioun:2012:NED

[1766] N. Hosseinioun, H. Doosti, and H. A. Nirumand. Nonparametric estimation of the derivatives of a density by the method of wavelet for mixing sequences. Statistical Papers, 53 (1):195–203, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0328-3.

Karimi:2012:BSR

[1767] Omid Karimi and Mohsen Mohammadzadeh. Bayesian spatial regression models with closed skew normal correlated errors and missing observations. Statistical Papers, 53(1): 205–218, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0329-2.

Hasanalipour:2012:NGB

[1768] P. Hasanalipour and M. Sharafi. A new generalized Balakrishnan skew-normal distribution. Statistical Papers, 53 (1):219–228, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0330-9.

Vangjeli:2012:AMD

[1769] Eno Vangjeli. ASN-minimax double sampling plans by variables for twosided specification limits when the standard deviation is known. Statistical Papers, 53(1):229–238, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0331-8.

Maronna:2012:BRT

[1770] Ricardo Maronna. Book review:

Thomas P. Ryan, Modern regression methods (2nd edn). Statistical Papers, 53(1):239–240, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-010-0347-0.

Lenz:2012:BRA

[1771] Sylvia Tamara Lenz. Book review:

Alex Dmitrienko, Ajit C. Tamhane, Frank Bretz (eds.): Multiple testing problems in pharmaceutical statistics. Statistical Papers, 53(1):241–242, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s00362-010-0350-5.

Anonymous:2012:HCa

[1772] Anonymous. Help & contacts. Statistical Papers, 53(1):??, February 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Jones:2012:LE

[1773] M. C. Jones. Letter to the Editor. Statistical Papers, 53(2):251, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-

012-0440-7.pdf. See reply [1774].

deGusmao:2012:RLE

[1774] F. R. S. de Gusmão, E. M. M. Ortega, and G. M. Cordeiro. Reply to the “Letter to the Editor” of M. C. Jones. Statistical Papers, 53(2):253–254, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-012-0441-6. See [1773].

Sbrana:2012:CAD

[1775] Giacomo Sbrana and Andrea Silvestrini. Comparing aggregate and disaggregate forecasts of first order moving average models. Statistical Papers, 53(2):255–263, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0333-6.

Duran:2012:EMJ

[1776] Esra Akdeniz Duran and Fikri Akdeniz. Efficiency of the modified jackknifed Liu-type estimator. Statistical Papers, 53(2):265–280, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-010-
                                        0334-5.

Batsidis:2012:EMD

[1777] Apostolos Batsidis. Errors of misclassification in discrimination with data from truncated t populations. Statistical Papers, 53(2):281–298, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0335-4.

Pordzik:2012:BED

[1778] Paweł R. Pordzik. A bound for the Euclidean distance between restricted and unrestricted estimators of parametric functions in the general linear model. Statistical Papers, 53(2):299–304, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-010-0336-3.pdf.

Quatto:2012:UMV

[1779] Piero Quatto and Antonella Zambon. The uniformly minimum variance unbiased estimator of odds ratio in case-control studies under inverse sampling. Statistical Papers, 53(2):305–309, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0337-2.

Krumbholz:2012:MVT

[1780] Wolf Krumbholz, Andreas Rohr, and Eno Vangjeli. Minimax versions of the two-stage t test. Statistical Papers, 53(2):311–321, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0339-0.

Dafnis:2012:DPT

[1781] Spiros D. Dafnis, Andreas N. Philippou, and Demetrios L. Antzoulakos. Distributions of patterns of two successes separated by a string of k−2 failures. Statistical Papers, 53(2):323–344, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-010-
                                        0340-7.

Michiels:2012:HIF

[1782] Frederik Michiels and Ann De Schepper. How to improve the fit of Archimedean copulas by means of transforms. Statistical Papers, 53 (2):345–355, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0341-6.

Sibbertsen:2012:TBP

[1783] Philipp Sibbertsen and Juliane Willert. Testing for a break in persistence under long-range dependencies and mean shifts. Statistical Papers, 53 (2):357–370, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0342-5.

Chen:2012:RSO

[1784] Xin Chen, Min Tsao, and Julie Zhou. Robust second-order least-squares estimator for regression models. Statistical Papers, 53(2):371–386, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0343-4.

Hu:2012:PRT

[1785] X. Joan Hu and Bin Zhang. Pseudolikelihood ratio test with biased observations. Statistical Papers, 53 (2):387–400, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0344-3.

Shen:2012:LRI

[1786] Junshan Shen, Wei Liang, and Shuyuan He. Likelihood ratio inference for mean residual life. Statistical Papers, 53(2):401–408, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0345-2.

Chen:2012:FSP

[1787] Qian Chen and David E. Giles. Finitesample properties of the maximum likelihood estimator for the binary logit model with random covariates. Statistical Papers, 53(2):409–426, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0348-z.

Li:2012:NLT

[1788] Yalian Li and Hu Yang. A new Liu-type estimator in linear regression model. Statistical Papers, 53(2):427–437, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0349-y.

Li:2012:RDA

[1789] Weiming Li, Tianqing Liu, and Zhidong Bai. Rounded data analysis based on ranked set sample. Statistical Papers, 53(2):439–455, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0351-4.

Laha:2012:SRE

[1790] Arnab Kumar Laha and K. C. Mahesh. SB-robust estimator for the concentration parameter of circular normal distribution. Statistical Papers, 53(2):457–467, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-010-0352-3.

Parchami:2012:TFH

[1791] Abbas Parchami, S. Mahmoud Taheri, and Mashaallah Mashinchi. Testing fuzzy hypotheses based on vague observations: a p-value approach. Statistical Papers, 53(2):469–484, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0353-2.

Wei:2012:ELP

[1792] Chuanhua Wei, Yubo Luo, and Xizhi Wu. Empirical likelihood for partially linear additive errors-in-variables models. Statistical Papers, 53(2):485–496, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0354-1.

Anonymous:2012:HCb

[1793] Anonymous. Help & contacts. Statistical Papers, 53(2):??, May 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Ho:2012:MCF

[1794] Linda Lee Ho, Roberto da Costa Quinino, Emílio Suyama, and Ruth Pereira Lourenço. Monitoring the conforming fraction of high-quality processes using a control chart p under a small sample size and an alternative estimator. Statistical Papers, 53(3):507–519, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0356-z.

Wang:2012:MEN

[1795] Litong Wang and Hu Yang. Matrix Euclidean norm Wielandt inequalities and their applications to statistics. Statistical Papers, 53(3):521–530, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-010-
0357-y.

Ravichandran:2012:RPT

[1796] J. Ravichandran. A review of preliminary test-based statistical methods for the benefit of six sigma quality practitioners. Statistical Papers, 53 (3):531–547, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-010-0359-9.

Chang:2012:CTP

[1797] Xinfeng Chang and Hu Yang. Combining two-parameter and principal component regression estimators. Statistical Papers, 53(3):549–562, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-011-
                                        0364-7.

Punzo:2012:DAC

[1798] Antonio Punzo and Alessandro Zini. Discrete approximations of continuous and mixed measures on a compact interval. Statistical Papers, 53 (3):563–575, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0365-6.

Fe:2012:IVE

[1799] Eduardo Fé. Instrumental variable estimation of heteroskedasticity adaptive error component models. Statistical Papers, 53(3):577–615, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-011-
                                        0366-5.

Cahoy:2012:EPL

[1800] Dexter O. Cahoy. An estimation procedure for the Linnik distribution. Statistical Papers, 53(3):617–628, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0367-4.

Qingguo:2012:CBS

[1801] Tang Qingguo and Cheng Longsheng.

Componentwise B-spline estimation for varying coefficient models with longitudinal data. Statistical Papers, 53 (3):629–652, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0369-2.

Cordeiro:2012:ECP

[1802] Gauss M. Cordeiro, Josemar Rodrigues, and Mário de Castro. The exponential COM–Poisson distribution. Statistical Papers, 53(3):653–664, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0370-9.

Zeller:2012:DMM

[1803] Camila B. Zeller, Rignaldo R. Carvalho, and Victor H. Lachos. On diagnostics in multivariate measurement error models under asymmetric heavytailed distributions. Statistical Papers, 53(3):665–683, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0371-8.

Nekoukhou:2012:FSS

[1804] V. Nekoukhou and M. H. Alamatsaz.

A family of skew-symmetric-Laplace distributions. Statistical Papers, 53 (3):685–696, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0372-7.

Lyhagen:2012:NRR

[1805] Johan Lyhagen. A note on the representation of E(x ⊗ xx′) and E(xx′ ⊗ xx′) for the random vector x. Statistical Papers, 53(3):697–701, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-011-
                                        0373-6.

Barabesi:2012:RRP

[1806] Lucio Barabesi, Sara Franceschi, and Marzia Marcheselli. A randomized response procedure for multiple-sensitive questions. Statistical Papers, 53(3): 703–718, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0374-5.

Jozani:2012:UAU

[1807] Mohammad Jafari Jozani, Saeed Majidi, and François Perron. Unbiased and almost unbiased ratio estimators of the population mean in ranked set sampling. Statistical Papers, 53 (3):719–737, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0376-3.

Reschenhofer:2012:IDF

[1808] Erhard Reschenhofer, Michael Schilde, Eva Oberecker, Ellen Payr, Hasan T. Tandogan, and Lea M. Wakolbinger.

Identifying the determinants of foreign direct investment: a data-specific model selection approach. Statistical Papers, 53(3):739–752, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0377-2.

Guven:2012:ATU

[1809] Bilgehan Güven. Approximate tests in unbalanced two-way random models without interaction. Statistical Papers, 53(3):753–766, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0378-1.

Hanck:2012:MUR

[1810] Christoph Hanck. Multiple unit root tests under uncertainty over the initial condition: some powerful modifications. Statistical Papers, 53(3): 767–774, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-011-0379-0.pdf.

Lee:2012:ICM

[1811] Chu-In Charles Lee, Wei Liu, Chul Gyu Park, and Jianan Peng. Inference for comparing a multinomial distribution with a known standard. Statistical Papers, 53(3):775–788, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0380-7.

Ahmadi:2012:OIP

[1812] Jafar Ahmadi and N. Balakrishnan. Outer and inner prediction intervals for order statistics intervals based on current records. Statistical Papers, 53 (3):789–802, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0383-4.

Webel:2012:BRP

[1813] Karsten Webel. Book review: Panik, M. (2009): Regression Modeling — Methods, Theory, and Computation with SAS. Statistical Papers, 53 (3):803–804, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-011-0368-
                                        3.

Anonymous:2012:HCc

[1814] Anonymous. Help & contacts. Statistical Papers, 53(3):??, August 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Withers:2012:MMC

[1815] Christopher S. Withers and Saralees Nadarajah. Maximum modulus confidence bands. Statistical Papers, 53 (4):811–819, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0384-3.

Mostofi:2012:BNI

[1816] A. Ghalamfarsa Mostofi and M. Kharrati-Kopaei. Bayesian nonparametric inference for unimodal skew-symmetric distributions. Statistical Papers, 53 (4):821–832, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0385-2.

Tojeiro:2012:GTS

[1817] Cynthia A. V. Tojeiro and Francisco Louzada. A general threshold stress hybrid hazard model for lifetime data. Statistical Papers, 53(4): 833–848, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0386-1.

Zare:2012:EVC

[1818] Karim Zare, Abdolrahman Rasekh, and Ali Akbar Rasekhi. Estimation of variance components in linear mixed measurement error models. Statistical Papers, 53(4):849–863, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0387-0.

Genc:2012:DLF

[1819] Ali I. Genç. Distribution of linear functions from ordered bivariate log-normal distribution. Statistical Papers, 53(4): 865–874, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0389-y.

Olmos:2012:EHN

[1820] Neveka M. Olmos, Héctor Varela, Héctor W. Gómez, and Heleno Bolfarine. An extension of the half-normal distribution. Statistical Papers, 53(4): 875–886, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0391-4.

Gamrot:2012:EFP

[1821] Wojciech Gamrot. Estimation of finite population kurtosis under twophase sampling for nonresponse. Statistical Papers, 53(4):887–894, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        011-0392-3.pdf.

Zhao:2012:ARD

[1822] Ningning Zhao and Zhidong Bai. Analysis of rounded data in mixture normal model. Statistical Papers, 53(4): 895–914, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0395-0.

Tsukuma:2012:SER

[1823] Hisayuki Tsukuma. Simultaneous estimation of restricted location parameters based on permutation and signchange. Statistical Papers, 53(4): 915–934, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0396-z.

Barczy:2012:AOI

[1824] Mátyás Barczy, Márton Ispány, Gyula Pap, Manuel Scotto, and Maria Eduarda Silva. Additive outliers in INAR(1) models. Statistical Papers, 53 (4):935–949, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0398-x.

Park:2012:DCI

[1825] Yousung Park and Hee-Young Kim.

Diagnostic checks for integer-valued autoregressive models using expected residuals. Statistical Papers, 53(4): 951–970, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0399-9.

Patil:2012:MA

[1826] P. N. Patil, P. P. Patil, and D. Bagkavos. A measure of asymmetry. Statistical Papers, 53(4): 971–985, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0401-6.

Giordano:2012:ECU

[1827] Sabrina Giordano and Pier Francesco Perri. Efficiency comparison of unrelated question models based on same privacy protection degree. Statistical Papers, 53(4):987–999, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0403-4.

Clarke:2012:FRM

[1828] Brenton R. Clarke, Peter L. McKinnon, and Geoff Riley. A fast robust method for fitting gamma distributions. Statistical Papers, 53(4): 1001–1014, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0404-3.

Schomaker:2012:SAE

[1829] Michael Schomaker. Shrinkage averaging estimation. Statistical Papers, 53(4):1015–1034, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00362-011-0405-
                                        2.

Borgoni:2012:UMP

[1830] Riccardo Borgoni and Piero Quatto. Uniformly most powerful unbiased test for shoulder condition in point transect sampling. Statistical Papers, 53(4): 1035–1044, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-011-0406-1.

Adamski:2012:GMB

[1831] K. Adamski, S. W. Human, and A. Bekker. A generalized multivariate beta distribution: control charting when the measurements are from an exponential distribution. Statistical Papers, 53(4):1045–1064, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-011-
                                        0407-0.

Bucker:2012:AJI

[1832] Michael Bücker. Alan Julian izenman: Modern statistical techniques: regression, classification, and manifold learning. Statistical Papers, 53(4): 1065–1066, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-011-0388-
z.

Mildenberger:2012:BRG

[1833] Thoralf Mildenberger. Book review:

Günther Sawitzki: Computational statistics. An introduction to R. Statistical Papers, 53(4):1067–1068, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-011-0393-2.

Hackl:2012:BRS

[1834] Peter Hackl. Book review: Shirley Coleman, Tony Greenfield, Dave Stewardson, Douglas C. Montgomery: Statistical practice in business and industry. Statistical Papers, 53(4): 1069, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-011-0394-1.pdf.

Anonymous:2012:HCd

[1835] Anonymous. Help & contacts. Statistical Papers, 53(4):??, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Goncalves:2013:AMS

[1836] Esmeralda Gonçalves, Joana Leite, and Nazaré Mendes-Lopes. The ARL of modified Shewhart control charts for conditionally heteroskedastic models.

Statistical Papers, 54(1):1–19, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0408-z.

Chacko:2013:CRV

[1837] Manoj Chacko and M. Shy Mary. Concomitants of k-record values arising from Morgenstern family of distributions and their applications in parameter estimation. Statistical Papers, 54 (1):21–46, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0409-y.

Quinino:2013:AEP

[1838] Roberto da Costa Quinino, Linda Lee Ho, and Emílio Suyama. Alternative estimator for the parameters of a mixture of two binomial distributions. Statistical Papers, 54(1):47–69, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0410-5.

Xu:2013:TNM

[1839] Wangli Xu, Yanwen Li, and Dawo Song. Testing normality in mixed models using a transformation method. Statistical Papers, 54(1):71–84, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0411-4.

Fan:2013:ELH

[1840] Guo-Liang Fan, Han-Ying Liang, and Jiang-Feng Wang. Empirical likelihood for heteroscedastic partially linear errors-in-variables model with αmixing errors. Statistical Papers, 54 (1):85–112, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0412-3.

Ortega:2013:LBW

[1841] Edwin M. M. Ortega, Gauss M.

Cordeiro, and Michael W. Kattan. The log-beta Weibull regression model with application to predict recurrence of prostate cancer. Statistical Papers, 54 (1):113–132, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0414-1.

Cordeiro:2013:BGR

[1842] Gauss M. Cordeiro, Cláudio T.

Cristino, Elizabeth M. Hashimoto, and Edwin M. M. Ortega. The beta generalized Rayleigh distribution with applications to lifetime data. Statistical Papers, 54(1):133–161, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0415-0.

Sultan:2013:UND

[1843] K. S. Sultan and A. S. Al-Moisheer.

Updating a nonlinear discriminant function estimated from a mixture of two inverse Weibull distributions. Statistical Papers, 54(1):163–175, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0416-z.

Mitnik:2013:KDM

[1844] Pablo A. Mitnik and Sunyoung Baek. The Kumaraswamy distribution: median-dispersion reparameterizations for regression modeling and simulation-based estimation. Statistical Papers, 54(1): 177–192, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0417-y.

Miao:2013:ANS

[1845] Yu Miao, Fangfang Zhao, Ke Wang, and Yanping Chen. Asymptotic normality and strong consistency of LS estimators in the EV regression model with NA errors. Statistical Papers, 54 (1):193–206, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0418-x.

Abu-Dayyeh:2013:ESS

[1846] Walid Abu-Dayyeh, Aissa Assrhani, and Kamarulzaman Ibrahim. Estimation of the shape and scale parameters of Pareto distribution using ranked set sampling. Statistical Papers, 54 (1):207–225, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0420-3.

Huang:2013:IPD

[1847] Wen-Jang Huang and Nan-Cheng Su.

Identification of power distribution mixtures through regression of exponentials. Statistical Papers, 54(1): 227–241, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0421-2.

Eryilmaz:2013:RLC

[1848] Serkan Eryilmaz. On residual lifetime of coherent systems after the rth failure. Statistical Papers, 54(1): 243–250, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0422-1.

Barranco-Chamorro:2013:E

[1849] I. Barranco-Chamorro, B. M. Golam Kibria, and A. K. Md. E. Saleh. Erratum. Statistical Papers, 54(1): 251–254, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0427-9.

Anonymous:2013:HCa

[1850] Anonymous. Help & contacts. Statistical Papers, 54(1):??, February 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Borroni:2013:NRC

[1851] Claudio Giovanni Borroni. A new rank correlation measure. Statistical Papers, 54(2):255–270, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-011-0423-0.

Dehgani:2013:MRA

[1852] Azam Dehgani, Ali Dolati, and Manuel Úbeda-Flores. Measures of radial asymmetry for bivariate random vectors. Statistical Papers, 54(2):271–286, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0425-y.

El-Din:2013:OTS

[1853] M. M. Mohie El-Din and A. R. Shafay. One- and two-sample Bayesian prediction intervals based on progressively Type-II censored data. Statistical Papers, 54(2):287–307, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-011-
0426-x.

Diaz-Frances:2013:ENA

[1854] Eloísa Díaz-Francés and Francisco J. Rubio. On the existence of a normal approximation to the distribution of the ratio of two independent normal random variables. Statistical Papers, 54(2):309–323, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0429-2.

Sheikhi:2013:EJD

[1855] Ayyub Sheikhi, Yaser Mehrali, and Mahbanoo Tata. On the exact joint distribution of a linear combination of order statistics and their concomitants in an exchangeable multivariate normal distribution. Statistical Papers, 54(2):325–332, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0430-9.

Cordeiro:2013:MOE

[1856] Gauss M. Cordeiro and Artur J.

Lemonte. On the Marshall–Olkin extended Weibull distribution. Statistical Papers, 54(2):333–353, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0431-8.

Nadar:2013:SAK

[1857] Mustafa Nadar, Alexander Papadopoulos, and Fatih Kizilaslan. Statistical analysis for Kumaraswamy’s distribution based on record data. Statistical Papers, 54(2):355–369, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0432-7.

Withers:2013:CLB

[1858] Christopher S. Withers and Saralees Nadarajah. Calibration with low bias. Statistical Papers, 54(2):371–379, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0433-6.

Tavangar:2013:CGP

[1859] Mahdi Tavangar and Marzieh Hashemi. On characterizations of the generalized Pareto distributions based on progressively censored order statistics. Statistical Papers, 54(2):381–390, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0434-5.

Asgharzadeh:2013:AML

[1860] A. Asgharzadeh, L. Esmaily, and S. Nadarajah. Approximate MLEs for the location and scale parameters of the skew logistic distribution. Statistical Papers, 54(2):391–411, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0436-3.

Szczepanska:2013:SCT

[1861] Anna Szczepańska. Simultaneous choice of time points and the block design in the growth curve model. Statistical Papers, 54(2):413–425, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        012-0438-1.pdf.

Genc:2013:SES

[1862] Ali I. Genç. A skew extension of the slash distribution via beta-normal distribution. Statistical Papers, 54 (2):427–442, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0439-0.

Withers:2013:CFO

[1863] Christopher S. Withers and Saralees Nadarajah. Correlation is first order independent of transformation. Statistical Papers, 54(2):443–456, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0442-5.

Taheri:2013:GWS

[1864] S. M. Taheri and G. Hesamian. A generalization of the Wilcoxon signedrank test and its applications. Statistical Papers, 54(2):457–470, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0443-4.

Kumar:2013:PTS

[1865] Nirpeksh Kumar. A procedure for testing suspected observations. Statistical Papers, 54(2):471–478, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0444-3.

Singh:2013:NPV

[1866] Housila P. Singh and Ramkrishna S.

Solanki. A new procedure for variance estimation in simple random sampling using auxiliary information. Statistical Papers, 54(2):479–497, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-012-
                                        0445-2.

Diaz-Frances:2013:SLB

[1867] Eloísa Díaz-Francés and José A. Montoya. The simplicity of likelihood based inferences for P (X < Y) and for the ratio of means in the exponential model. Statistical Papers, 54(2):499–522, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0446-1.

Li:2013:UVD

[1868] Zhonghua Li, Peihua Qiu, Snigdhansu Chatterjee, and Zhaojun Wang. Using p values to design statistical process control charts. Statistical Papers, 54(2):523–539, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0447-0.

Polasek:2013:BRR

[1869] Wolfgang Polasek. Book review: Raquel Prado and Mike West: Time series: modelling, computation and inference. Statistical Papers, 54(2):541–542, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-012-0428-3.

Hackl:2013:BRD

[1870] P. G. Hackl. Book review: Dale L. Zimmerman, Vicente A. Núñez–Antón: Antedependence models for longitudinal data. Statistical Papers, 54(2):543–544, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-012-0435-4.

Anonymous:2013:HCb

[1871] Anonymous. Help & contacts. Statistical Papers, 54(2):??, May 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hussein:2013:GSM

[1872] A. Hussein, H. A. Muttlak, and E. Al-Sawi. Group sequential methods based on ranked set samples. Statistical Papers, 54(3):547–562, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0448-z.

Weiss:2013:PEB

[1873] Christian H. Weiß and Hee-Young Kim.

Parameter estimation for binomial AR(1) models with applications in finance and industry. Statistical Papers, 54(3):563–590, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0449-y.

Kuang:2013:LMD

[1874] Nenghui Kuang and Huantian Xie.

Large and moderate deviations in testing Rayleigh diffusion model. Statistical Papers, 54(3):591–603, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0450-5.

Khan:2013:TBL

[1875] Shahjahan Khan and Budi Pratikno. Testing base load with non-sample prior information on process load. Statistical Papers, 54(3):605–617, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0451-4.

Rastogi:2013:IUP

[1876] Manoj Kumar Rastogi and Yogesh Mani Tripathi. Inference on unknown parameters of a Burr distribution under hybrid censoring. Statistical Papers, 54(3):619–643, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0452-3.

Liu:2013:RRI

[1877] Tianqing Liu and Xiaohui Yuan. Random rounded integer-valued autoregressive conditional heteroskedastic process. Statistical Papers, 54(3): 645–683, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0453-2.

Kim:2013:TPG

[1878] Sungsu Kim and Ashis SenGupta. A three-parameter generalized von Mises distribution. Statistical Papers, 54 (3):685–693, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0454-1.

Christensen:2013:LMA

[1879] Ronald Christensen and Yong Lin. Linear models that allow perfect estimation. Statistical Papers, 54(3):695–708, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0455-0.

Rosco:2013:MTA

[1880] J. F. Rosco and Harry Joe. Measures of tail asymmetry for bivariate copulas. Statistical Papers, 54(3):709–726, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0457-y.

Gamrot:2013:MLE

[1881] Wojciech Gamrot. Maximum likelihood estimation for ordered expectations of correlated binary variables. Statistical Papers, 54(3):727–739, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
012-0458-x.pdf.

Genc:2013:MTN

[1882] Ali I. Genç. Moments of truncated normal/independent distributions. Statistical Papers, 54(3):741–764, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0459-9.

Jabbari:2013:ASC

[1883] H. Jabbari. On almost sure convergence for weighted sums of pairwise negatively quadrant dependent random variables. Statistical Papers, 54 (3):765–772, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0460-3.

Sung:2013:SCW

[1884] Soo Hak Sung. On the strong convergence for weighted sums of ρ^{∗}-mixing random variables. Statistical Papers, 54(3):773–781, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0461-2.

Makri:2013:EDC

[1885] Frosso S. Makri and Zaharias M. Psillakis. Exact distributions of constrained (k, l) strings of failures between subsequent successes. Statistical Papers, 54(3):783–806, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0462-1.

Domma:2013:CBA

[1886] Filippo Domma and Sabrina Giordano. A copula-based approach to account for dependence in stress-strength models. Statistical Papers, 54(3):807–826, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0463-0.

Mukhopadhyay:2013:NIS

[1887] Nitis Mukhopadhyay and Bhargab Chattopadhyay. On a new interpretation of the sample variance. Statistical Papers, 54(3):827–837, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0465-y.

Nadarajah:2013:EWD

[1888] Saralees Nadarajah, Gauss M. Cordeiro, and Edwin M. M. Ortega. The exponentiated Weibull distribution: a survey. Statistical Papers, 54(3):839–877, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0466-x.

Ateya:2013:MTG

[1889] Saieed F. Ateya and Elham A. Madhagi. On multivariate truncated generalized Cauchy distribution. Statistical Papers, 54(3):879–897, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0467-9.

Maronna:2013:BRT

[1890] Ricardo Maronna. Book review:

Tao Li, Mitsunori Ogihara, George Tzanetakis (eds.). Music data mining. Statistical Papers, 54(3):899, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
012-0456-z.pdf.

Polasek:2013:BRP

[1891] Wolfgang Polasek. Book review: Peter D. Congdon: Applied Bayesian hierarchical methods. Statistical Papers, 54(3):901–902, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-012-0468-
                                        8.

Anonymous:2013:HCc

[1892] Anonymous. Help & contacts. Statistical Papers, 54(3):??, August 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Sibbertsen:2013:EI

[1893] Philipp Sibbertsen and Rafael Weißbach. Editors’ introduction. Statistical Papers, 54(4):907–909, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.com/
                                        content/pdf/10.1007/s00362-013-
                                        0542-x.pdf.

Breitung:2013:WBB

[1894] Jörg Breitung and Robinson Kruse. When bubbles burst: econometric tests based on structural breaks. Statistical Papers, 54(4):911–930, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0497-3.

Zeileis:2013:TPT

[1895] Achim Zeileis and Torsten Hothorn. A toolbox of permutation tests for structural change. Statistical Papers, 54 (4):931–954, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0503-4.

Wied:2013:ANT

[1896] Dominik Wied, Daniel Ziggel, and Tobias Berens. On the application of new tests for structural changes on global minimum-variance portfolios. Statistical Papers, 54(4):955–975, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0511-4.

Bertram:2013:FIV

[1897] Philip Bertram, Robinson Kruse, and Philipp Sibbertsen. Fractional integration versus level shifts: the case of realized asset correlations. Statistical Papers, 54(4):977–991, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0513-2.

Kremer:2013:CED

[1898] Alexander Kremer and Rafael Weißbach.

Consistent estimation for discretely observed Markov jump processes with an absorbing state. Statistical Papers, 54 (4):993–1007, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0515-0.

Lee:2013:ROT

[1899] Taesuk Lee, Mico Loretan, and Werner Ploberger. Rate-optimal tests for jumps in diffusion processes. Statistical Papers, 54(4):1009–1041, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0541-y.

Demetrescu:2013:NIP

[1900] Matei Demetrescu and Christoph Hanck. Nonlinear IV panel unit root testing under structural breaks in the error variance. Statistical Papers, 54(4):1043–1066, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0502-5.

Baltagi:2013:TCS

[1901] Badi H. Baltagi, Chihwa Kao, and Sanggon Na. Testing for cross-sectional dependence in a panel factor model using the wild bootstrap F test. Statistical Papers, 54(4):1067–1094, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0499-9.

Waser:2013:EDA

[1902] M. Waser, M. Deistler, H. Garn, T. Benke, P. Dal-Bianco, G. Ransmayr, D. Grossegger, and R. Schmidt. EEG in the diagnostics of Alzheimer’s disease. Statistical Papers, 54(4): 1095–1107, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0538-6.

Baksalary:2013:LUD

[1903] Oskar Maria Baksalary, Götz Trenkler, and Erkki Liski. Let us do the twist again. Statistical Papers, 54(4): 1109–1119, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-013-0512-3.pdf.

Kleiber:2013:MIB

[1904] Christian Kleiber. On moment indeterminacy of the Benini income distribution. Statistical Papers, 54(4): 1121–1130, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0535-9.

Lutkepohl:2013:RCB

[1905] Helmut Lütkepohl. Reducing confidence bands for simulated impulse responses. Statistical Papers, 54(4): 1131–1145, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0510-5.

Pardo:2013:JAP

[1906] Leandro Pardo and María del Carmen Pardo. Julio Angel Pardo Llorente, 1960–2013 (Associate Editor of Statistical Papers). Statistical Papers, 54(4):1147–1149, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0550-x.

Anonymous:2013:HCd

[1907] Anonymous. Help & contacts. Statistical Papers, 54(4):??, November 2013. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Fried:2014:P

[1908] Roland Fried, Sonja Kuhnt, and Christine H. Müller. Preface. Statistical Papers, 55(1):1–2, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-013-0570-6.

Denecke:2014:CLD

[1909] Liesa Denecke and Christine H. Müller. Consistency of the likelihood depth estimator for the correlation coefficient. Statistical Papers, 55(1):3–13, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0490-x.

Hubert:2014:SBR

[1910] M. Hubert, P. Rousseeuw, and K. Vakili. Shape bias of robust covariance estimators: an empirical study. Statistical Papers, 55(1):15–28, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0544-8.

Filzmoser:2014:ILM

[1911] Peter Filzmoser, Anne Ruiz-Gazen, and Christine Thomas-Agnan. Identification of local multivariate outliers.

Statistical Papers, 55(1):29–47, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0524-z.

Lange:2014:FNC

[1912] Tatjana Lange, Karl Mosler, and Pavlo Mozharovskyi. Fast nonparametric classification based on data depth. Statistical Papers, 55(1):49–69, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0488-4.

Nevalainen:2014:IMD

[1913] Jaakko Nevalainen, Somnath Datta, and Hannu Oja. Inference on the marginal distribution of clustered data with informative cluster size. Statistical Papers, 55(1):71–92, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0504-3.

Ruckdeschel:2014:RKT

[1914] Peter Ruckdeschel, Bernhard Spangl, and Daria Pupashenko. Robust Kalman tracking and smoothing with propagating and non-propagating outliers. Statistical Papers, 55(1): 93–123, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0496-4.

Kustosz:2014:ACG

[1915] Christoph P. Kustosz and Christine H. Müller. Analysis of crack growth with robust, distribution-free estimators and tests for non-stationary autoregressive processes. Statistical Papers, 55(1): 125–140, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0479-5.

Nordhausen:2014:RSS

[1916] Klaus Nordhausen. On robustifying some second order blind source separation methods for nonstationary time series. Statistical Papers, 55 (1):141–156, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0487-5.

Zahle:2014:QRM

[1917] Henryk Zähle. Qualitative robustness of von Mises statistics based on strongly mixing data. Statistical Papers, 55(1):157–167, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0478-6.

Aquaro:2014:RED

[1918] Michele Aquaro and Pavel Cízek. Robust estimation of dynamic fixedeffects panel data models. Statistical Papers, 55(1):169–186, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-013-
0545-7.

Neykov:2014:UDV

[1919] N. M. Neykov, P. Filzmoser, and P. N. Neytchev. Ultrahigh dimensional variable selection through the penalized maximum trimmed likelihood estimator. Statistical Papers, 55 (1):187–207, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0516-z. See erra-

tum [1964].

Shevlyakova:2014:SIR

[1920] Maya Shevlyakova and Stephan Morgenthaler. Sliced inverse regression for survival data. Statistical Papers, 55 (1):209–220, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0552-8.

Bednarski:2014:RCN

[1921] Tadeusz Bednarski. On robust causality nonresponse testing in duration studies under the Cox model. Statistical Papers, 55(1):221–231, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
013-0523-0.pdf.

Toman:2014:RCF

[1922] Ales Toman. Robust confirmatory factor analysis based on the forward search algorithm. Statistical Papers, 55 (1):233–252, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0525-y.

Anonymous:2014:HCa

[1923] Anonymous. Help & contacts. Statistical Papers, 55(1):??, February 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Singh:2014:RME

[1924] Sukhbir Singh, Kanchan Jain, and Suresh Sharma. Replicated measurement error model under exact linear restrictions. Statistical Papers, 55(2):253–274, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0469-7.

Veres-Ferrer:2014:RBR

[1925] Ernesto J. Veres-Ferrer and Jose M. Pavía. On the relationship between the reversed hazard rate and elasticity. Statistical Papers, 55(2):275–284, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0470-1.

Zhang:2014:BRS

[1926] Qinchi Zhang, Wenzhi Yang, and Shuhe Hu. On Bahadur representation for sample quantiles under αmixing sequence. Statistical Papers, 55(2):285–299, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0472-z.

Chen:2014:NCT

[1927] Zhongxue Chen, Hon Keung Tony Ng, and Saralees Nadarajah. A note on Cochran test for homogeneity in oneway ANOVA and meta-analysis. Statistical Papers, 55(2):301–310, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0475-9.

Ateya:2014:MLE

[1928] Saieed F. Ateya. Maximum likelihood estimation under a finite mixture of generalized exponential distributions based on censored data. Statistical Papers, 55(2):311–325, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0480-z.

Xu:2014:VSH

[1929] Dengke Xu, Zhongzhan Zhang, and Liucang Wu. Variable selection in high-dimensional double generalized linear models. Statistical Papers, 55 (2):327–347, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0481-y.

Ciuperca:2014:MSL

[1930] Gabriela Ciuperca. Model selection by LASSO methods in a changepoint model. Statistical Papers, 55 (2):349–374, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0482-x.

Song:2014:WLS

[1931] Guang Jing Song and Qing Wen Wang. On the weighted least-squares, the ordinary least-squares and the best linear unbiased estimators under a restricted growth curve model. Statistical Papers, 55(2):375–392, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0483-9.

Siray:2014:CEL

[1932] Gülesen Üstündağ Şiray, Selahattin Kaçiranlar, and Sadullah Sakallioğlu. r–k class estimator in the linear regression model with correlated errors. Statistical Papers, 55(2):393–407, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0484-8.

Salehi:2014:GST

[1933] Mahdi Salehi, Ahad Jamalizadeh, and Mahdi Doostparast. A generalized skew two-piece skew-elliptical distribution. Statistical Papers, 55(2):409–429, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0485-7.

Gimenez:2014:LIF

[1934] Patricia Giménez and María Laura Patat. Local influence for functional comparative calibration models with replicated data. Statistical Papers, 55(2):431–454, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0489-3.

Czado:2014:NNM

[1935] Claudia Czado, Holger Schabenberger, and Vinzenz Erhardt. Non nested model selection for spatial count regression models with application to health insurance. Statistical Papers, 55(2):455–476, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0491-9.

Baltagi:2014:TSL

[1936] Badi H. Baltagi and Long Liu. Testing for spatial lag and spatial error dependence using double length artificial regressions. Statistical Papers, 55(2):477–486, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-012-0492-8.

Reschenhofer:2014:DFP

[1937] Erhard Reschenhofer, Werner Ploberger, and Georg V. Lehecka. Detecting fuzzy periodic patterns in futures spreads. Statistical Papers, 55(2):487–496, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-012-
0493-7.

Gomez-Deniz:2014:DVH

[1938] E. Gómez-Déniz, F. J. Vázquez-Polo, and V. García-García. A discrete version of the half-normal distribution and its generalization with applications.

Statistical Papers, 55(2):497–511, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-012-
                                        0494-6.

Peiris:2014:TNH

[1939] Shelton Peiris. Testing the null hypothesis of zero serial correlation in short panel time series: a comparison of tail probabilities. Statistical Papers, 55(2):513–523, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-012-0495-5.

Schepsmeier:2014:DFI

[1940] Ulf Schepsmeier and Jakob Stöber. Derivatives and Fisher information of bivariate copulas. Statistical Papers, 55(2):525–542, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0498-x.

Haslett:2014:EBO

[1941] Stephen J. Haslett, Jarkko Isotalo, Yonghui Liu, and Simo Puntanen. Equalities between OLSE, BLUE and BLUP in the linear model. Statistical Papers, 55(2):543–561, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0500-7.

Li:2014:LEP

[1942] Weiming Li. Local expectations of the population spectral distribution of a high-dimensional covariance matrix. Statistical Papers, 55(2):563–573, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0501-6.

Mildenberger:2014:BRS

[1943] Thoralf Mildenberger. Book review:

Stephen Marsland: Machine learning. An algorithmic perspective. Statistical Papers, 55(2):575–576, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-012-0471-0.

Anonymous:2014:HCb

[1944] Anonymous. Help & contacts. Statistical Papers, 55(2):??, May 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Maya:2014:ERR

[1945] R. Maya, E. I. Abdul-Sathar, G. Rajesh, and K. R. Muraleedharan Nair. Estimation of the Renyi’s residual entropy of order α with dependent data. Statistical Papers, 55(3):585–602, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0506-1.

Rodrigues:2014:SOF

[1946] Paulo C. Rodrigues, Elsa E. Moreira, Vera M. Jesus, and João T. Mexia. Structured orthogonal families of one and two strata prime basis factorial models. Statistical Papers, 55(3):603–614, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0507-0.

Ma:2014:PCC

[1947] Tie Feng Ma and Shuangzhe Liu. Pitman closeness of the class of isotonic estimators for ordered scale parameters of two gamma distributions. Statistical Papers, 55(3):615–625, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0508-z.

Karlsson:2014:FMM

[1948] Maria Karlsson and Thomas Laitila. Finite mixture modeling of censored regression models. Statistical Papers, 55(3):627–642, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-013-0509-y.pdf.

Cordeiro:2014:CMB

[1949] Gauss M. Cordeiro, Denise A. Botter, Alexsandro B. Cavalcanti, and Lúcia P. Barroso. Covariance matrix of the bias-corrected maximum likelihood estimator in generalized linear models. Statistical Papers, 55(3):643–652, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0514-1.

Schmid:2014:SRS

[1950] Timo Schmid and Ralf T. Münnich. Spatial robust small area estimation.

Statistical Papers, 55(3):653–670, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0517-y.

Zeller:2014:IDG

[1951] Camila B. Zeller, Victor H. Lachos, and Filidor Vilca Labra. Influence diagnostics for Grubbs’s model with asymmetric heavy-tailed distributions. Statistical Papers, 55(3):671–690, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0519-9.

Anis:2014:TNM

[1952] M. Z. Anis. Tests of non-monotonic stochastic aging notions in reliability theory. Statistical Papers, 55 (3):691–714, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0520-3.

Sha:2014:BAS

[1953] Naijun Sha and Rong Pan. Bayesian analysis for step-stress accelerated life testing using Weibull proportional hazard model. Statistical Papers, 55 (3):715–726, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0521-2.

Antoniadis:2014:PEA

[1954] A. Antoniadis, I. Gijbels, and S. Lambert-Lacroix. Penalized estimation in additive varying coefficient models using grouped regularization. Statistical Papers, 55(3):727–750, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0522-1.

Nadar:2014:CBE

[1955] Mustafa Nadar and Fatih Kizilaslan. Classical and Bayesian estimation of P (X < Y) using upper record values from Kumaraswamy’s distribution. Statistical Papers, 55(3):751–783, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0526-x.

Genc:2014:DPQ

[1956] Ali I. Genç. Distribution of product and quotient of bivariate generalized exponential distribution. Statistical Papers, 55(3):785–803, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0527-9.

Argiento:2014:EPI

[1957] Raffaele Argiento, Alessandra Guglielmi, and Antonio Pievatolo. Estimation, prediction and interpretation of NGG random effects models: an application to Kevlar fibre failure times. Statistical Papers, 55(3):805–826, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0528-8.

Zheng:2014:AEL

[1958] Jiayin Zheng, Junshan Shen, and Shuyuan He. Adjusted empirical likelihood for right censored lifetime data. Statistical Papers, 55(3):827–839, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0529-7.

Eryilmaz:2014:LBS

[1959] Serkan Eryilmaz and Konul Bayramoglu. Life behavior of δ-shock models for uniformly distributed interarrival times. Statistical Papers, 55 (3):841–852, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0530-1.

Matsuura:2014:PPA

[1960] Shun Matsuura and Hiroshi Kurata.

Principal points for an allometric extension model. Statistical Papers, 55 (3):853–870, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0532-z.

Depraetere:2014:OSF

[1961] Nicolas Depraetere and Martina Vandebroek. Order selection in finite mixtures of linear regressions. Statistical Papers, 55(3):871–911, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0534-x.

Maronna:2014:BRW

[1962] Ricardo Maronna. Book review: Wendy Martínez, Angel R. Martínez, Jeffrey L. Solka: Exploratory data analysis with MATLAB^{(R)}, second edition. Statistical Papers, 55(3):913–914, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-013-0518-x.

Kramer:2014:BRK

[1963] Walter Krämer. Book review: Kahneman, D. (2011): Thinking, Fast and Slow. Statistical Papers, 55(3):915, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        013-0533-y.pdf.

Neykov:2014:EUD

[1964] N. M. Neykov, P. Filzmoser, and P. N. Neytchev. Erratum to: Ultrahigh dimensional variable selection through the penalized maximum trimmed likelihood estimator. Statistical Papers, 55(3):917–918, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        013-0531-0.pdf. See [1919].

Anonymous:2014:HCc

[1965] Anonymous. Help & contacts. Statistical Papers, 55(3):??, August 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Schnurr:2014:OPA

[1966] Alexander Schnurr. An ordinal pattern approach to detect and to model leverage effects and dependence structures between financial time series. Statistical Papers, 55(4):919–931, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0536-8.

Faraz:2014:VPC

[1967] Alireza Faraz, Giovanni Celano, Erwin Saniga, C. Heuchenne, and S. Fichera. The variable parameters T^{2} chart with run rules. Statistical Papers, 55(4): 933–950, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0537-7.

Withers:2014:UMC

[1968] Christopher S. Withers and Saralees Nadarajah. A unified method for constructing expectation tolerance intervals. Statistical Papers, 55(4): 951–965, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0543-9.

Olmos:2014:EGH

[1969] Neveka M. Olmos, Héctor Varela, Heleno Bolfarine, and Héctor W. Gómez. An extension of the generalized half-normal distribution. Statistical Papers, 55(4):967–981, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-013-
                                        0546-6.

Fernandez-Duran:2014:DSD

[1970] J. J. Fernández-Durán and M. M. Gregorio-Domínguez. Distributions for spherical data based on nonnegative trigonometric sums. Statistical Papers, 55(4):983–1000, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0547-5.

Beyaztas:2014:SJA

[1971] Ufuk Beyaztas and Aylin Alin. Sufficient jackknife-after-bootstrap method for detection of influential observations in linear regression models. Statistical Papers, 55(4):1001–1018, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0548-4.

Chahkandi:2014:NPP

[1972] M. Chahkandi, Jafar Ahmadi, and S. Baratpour. Non-parametric prediction intervals for the lifetime of coherent systems. Statistical Papers, 55(4): 1019–1034, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0549-3.

Li:2014:CLT

[1973] Deli Li, Oleg Klesov, and George Stoica. On the central limit theorem along subsequences of sums of i.i.d. random variables. Statistical Papers, 55(4): 1035–1045, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0551-9.

Ciginas:2014:ANF

[1974] Andrius Ciginas. On the asymptotic normality of finite population Lstatistics. Statistical Papers, 55(4): 1047–1058, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0553-7.

Hobza:2014:DBT

[1975] Tomás Hobza, Domingo Morales, and Leandro Pardo. Divergence-based tests of homogeneity for spatial data. Statistical Papers, 55(4):1059–1077, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0554-6.

Dong:2014:DFM

[1976] Ye Dong and Stephen M. S. Lee. Depth functions as measures of representativeness. Statistical Papers, 55(4): 1079–1105, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0555-5.

Genest:2014:TRS

[1977] Christian Genest and Johanna G.

Neslehová. On tests of radial symmetry for bivariate copulas. Statistical Papers, 55(4):1107–1119, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-013-
                                        0556-4.

Alvarez-Andrade:2014:ARH

[1978] Sergio Alvarez-Andrade and Salim Bouzebda. Asymptotic results for hybrids of empirical and partial sums processes. Statistical Papers, 55(4): 1121–1143, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0557-3.

Wang:2014:WEM

[1979] Lihong Wang and Jinde Wang. Wavelet estimation of the memory parameter for long range dependent random fields. Statistical Papers, 55(4): 1145–1158, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0558-2.

Al-Kandari:2014:ATT

[1980] Noriah M. Al-Kandari and Emad-Eldin A. A. Aly. An ANOVA-type test for multiple change points. Statistical Papers, 55(4):1159–1178, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0559-1.

Kamalja:2014:JDS

[1981] Kirtee K. Kamalja. On the joint distribution of success runs of several lengths in the sequence of MBT and its applications. Statistical Papers, 55(4): 1179–1206, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0560-8.

Kawano:2014:STP

[1982] Shuichi Kawano. Selection of tuning parameters in bridge regression models via Bayesian information criterion. Statistical Papers, 55(4): 1207–1223, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0561-7.

Maronna:2014:BRM

[1983] Ricardo Maronna. Book review:

Max D. Morris: Design of experiments: an introduction based on linear models. Statistical Papers, 55(4): 1225–1226, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-013-0539-
5.

Bucker:2014:BRJ

[1984] Michael Bücker. Book review:

Jean Dickinson Gibbons, Subhabrata Chakraborti: Nonparametric statistical inferences. Statistical Papers, 55(4): 1227–1228, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-013-0540-
z.

Lutkepohl:2014:BRM

[1985] Helmut Lütkepohl. Book review: Mulaik, S. A.: Foundations of factor analysis. Statistical Papers, 55(4): 1229–1230, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-013-0562-
                                        6.

Anonymous:2014:HCd

[1986] Anonymous. Help & contacts. Statistical Papers, 55(4):??, November 2014. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Rakitzis:2015:SDT

[1987] Athanasios C. Rakitzis and Demetrios L. Antzoulakos. Start-up demonstration tests with three-level classification. Statistical Papers, 56(1):1–21, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0564-4.

Xia:2015:FCE

[1988] Ningning Xia and Zhidong Bai. Functional CLT of eigenvectors for large sample covariance matrices. Statistical Papers, 56(1):23–60, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0565-3.

Hesamian:2015:TSK

[1989] Gholamreza Hesamian and Jalal Chachi. Two-sample Kolmogorov– Smirnov fuzzy test for fuzzy random variables. Statistical Papers, 56 (1):61–82, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0566-2.

Solanki:2015:ECE

[1990] Ramkrishna S. Solanki and Housila P.

Singh. Efficient classes of estimators in stratified random sampling. Statistical Papers, 56(1):83–103, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0567-1.

Mandal:2015:OMD

[1991] Nripes Kumar Mandal, Manisha Pal, Bikas K. Sinha, and Premadhis Das. Optimum mixture designs in a restricted region. Statistical Papers, 56 (1):105–119, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0568-0.

Balakrishnan:2015:ASD

[1992] N. Balakrishnan, C. Koukouvinos, and C. Parpoula. Analyzing supersaturated designs for discrete responses via generalized linear models. Statistical Papers, 56(1):121–145, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0569-z.

Inan:2015:CLT

[1993] Deniz Inan. Combining the Liu-type estimator and the principal component regression estimator. Statistical Papers, 56(1):147–156, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-013-
                                        0571-5.

Li:2015:EIE

[1994] Shuyou Li, Xu Lu, Ying Mi, and Wei Liu. The estimation and inference on the equal ratios of means to standard deviations of normal populations. Statistical Papers, 56(1):157–165, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-013-
                                        0572-4.

Groenitz:2015:UPI

[1995] Heiko Groenitz. Using prior information in privacy-protecting survey designs for categorical sensitive variables. Statistical Papers, 56(1): 167–189, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0573-3.

Zheng:2015:GPA

[1996] Fa mei Zheng and Qing pei Zang. A general pattern of asymptotic behavior of the R/S statistics for linear processes. Statistical Papers, 56 (1):191–204, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-013-0574-2.

Parvardeh:2015:NAD

[1997] Afshin Parvardeh. A note on the asymptotic distribution of the estimation of the mean past lifetime. Statistical Papers, 56(1):205–215, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-013-
0575-1.

Huang:2015:PCT

[1998] Jiewu Huang and Hu Yang. On a principal component two-parameter estimator in linear model with autocorrelated errors. Statistical Papers, 56 (1):217–230, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-013-0576-0.

Arashi:2015:PKM

[1999] M. Arashi and T. Valizadeh. Performance of Kibria’s methods in partial linear ridge regression model. Statistical Papers, 56(1):231–246, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0578-6.

Park:2015:CCR

[2000] Sangun Park and Johan Lim. On censored cumulative residual Kullback– Leibler information and goodness-of-fit test with type II censored data. Statistical Papers, 56(1):247–256, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0579-5.

Jiang:2015:PEN

[2001] Hui Jiang and Xing Dong. Parameter estimation for the non-stationary Ornstein–Uhlenbeck process with linear drift. Statistical Papers, 56(1):

257–268, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0580-z.

Lenz:2015:BRE

[2002] Sylvia Tamara Lenz. Book review: Edward F. Vonesh: Generalized linear and nonlinear models for correlated data, theory and applications using SAS^{(R)}. Statistical Papers, 56 (1):269–270, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-013-0563-
                                        5.

Mildenberger:2015:BRS

[2003] Thoralf Mildenberger. Book review: Simon Rogers and Mark Girolami: A first course in machine learning. Statistical Papers, 56(1):271, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        013-0577-z.pdf.

Casero-Alonso:2015:EDT

[2004] Víctor Casero-Alonso and Jesús López-Fidalgo. Experimental designs in triangular simultaneous equations models. Statistical Papers, 56(2):273–290, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0581-y.

Gunduz:2015:DQT

[2005] Selim Gündüz and Ali I. Genç. The distribution of the quotient of two triangularly distributed random variables. Statistical Papers, 56(2):291–310, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0582-x.

Hanna:2015:NCD

[2006] Hanen Hanna and Walter Tinsson. A new class of designs for mixture-of-mixture experiments. Statistical Papers, 56(2):311–331, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0583-9.

Gunasekera:2015:GIR

[2007] Sumith Gunasekera. Generalized inferences of R = Pr(X > Y) for Pareto distribution. Statistical Papers, 56(2):333–351, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0584-8.

Arslan:2015:VMM

[2008] Olcay Arslan. Variance-mean mixture of the multivariate skew normal distribution. Statistical Papers, 56(2):353–378, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0585-7.

Namba:2015:MDP

[2009] Akio Namba. MSE dominance of the positive-part shrinkage estimator when each individual regression coefficient is estimated. Statistical Papers, 56(2):379–390, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0586-6.

Fotouhi:2015:HRG

[2010] H. Fotouhi and M. Golalizadeh. Highly resistant gradient descent algorithm for computing intrinsic mean shape on similarity shape spaces. Statistical Papers, 56(2):391–410, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        014-0587-5.pdf.

Zhao:2015:ELB

[2011] Weihua Zhao, Riquan Zhang, Yukun Liu, and Jicai Liu. Empirical likelihood based modal regression. Statistical Papers, 56(2):411–430, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0588-4.

Beran:2015:MLR

[2012] Jan Beran, Yuanhua Feng, and Sucharita Ghosh. Modelling longrange dependence and trends in duration series: an approach based on EFARIMA and ESEMIFAR models. Statistical Papers, 56(2):431–451, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0590-x.

Belaghi:2015:IED

[2013] R. Arabi Belaghi, M. Arashi, and S. M. M. Tabatabaey. Improved estimators of the distribution function based on lower record values. Statistical Papers, 56(2):453–477, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0591-9.

deCastro:2015:ISH

[2014] Mário de Castro and Manuel Galea. Inference in a structural heteroskedastic calibration model. Statistical Papers, 56(2):479–494, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0592-8.

Kurnaz:2015:NLT

[2015] Fatma Sevinç Kurnaz and Kadri Ulas Akay. A new Liu-type estimator. Statistical Papers, 56(2):495–517, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0594-6.

Eryilmaz:2015:RBC

[2016] Serkan Eryilmaz and G. Yazgi Tutuncu. Relative behavior of a coherent system with respect to another coherent system. Statistical Papers, 56(2):519–529, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0595-5.

Ozkale:2015:PPL

[2017] M. Revan Özkale. Predictive performance of linear regression models. Statistical Papers, 56(2):531–567, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0596-4.

Lin:2015:DTT

[2018] Chien-Tai Lin and Shih-Chun Wang. Discordancy tests for two-parameter exponential samples. Statistical Papers, 56(2):569–582, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0597-3.

Hackl:2015:BRW

[2019] Peter Hackl. Book review: Wladyslaw Welfe: Macroeconomic models, Springer, 2013, xxvii + 425pp., EUR181,85, $229.00, £153.00, ISBN: 978-3-642-34467-1. Statistical Papers, 56(2):583–584, May 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-014-0589-
3.

Wang:2015:CCE

[2020] Xuejun Wang and Zeyu Si. Complete consistency of the estimator of nonparametric regression model under ND sequence. Statistical Papers, 56 (3):585–596, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0598-2.

Baringhaus:2015:TST

[2021] L. Baringhaus and D. Kolbe. Twosample tests based on empirical Hankel transforms. Statistical Papers, 56 (3):597–617, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0599-1.

Kundu:2015:CBM

[2022] Chanchal Kundu and Asok K. Nanda.

Characterizations based on measure of inaccuracy for truncated random variables. Statistical Papers, 56(3): 619–637, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0600-z.

Kateri:2015:ISS

[2023] Maria Kateri and Udo Kamps. Inference in step-stress models based on failure rates. Statistical Papers, 56 (3):639–660, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0601-y.

Zamini:2015:EDF

[2024] R. Zamini, V. Fakoor, and M. Sarmad. On estimation of a density function in multiplicative censoring. Statistical Papers, 56(3):661–676, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0602-x.

Zardasht:2015:ECR

[2025] V. Zardasht, S. Parsi, and M. Mousazadeh. On empirical cumulative residual entropy and a goodness-of-fit test for exponentiality. Statistical Papers, 56 (3):677–688, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0603-9;   http:
//link.springer.com/content/pdf/
10.1007/s00362-014-0603-9.pdf.

Schneider:2015:STT

[2026] Berthold Schneider, Dieter Rasch, Klaus D. Kubinger, and Takuya Yanagida. A sequential triangular test of a correlation coefficient’s nullhypothesis: 0 < ρ < ρ_{0}. Statistical Papers, 56(3):689–699, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0604-8.

Durante:2015:CTS

[2027] Fabrizio Durante, Roberta Pappadà, and Nicola Torelli. Clustering of time series via non-parametric tail dependence estimation. Statistical Papers, 56(3):701–721, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0605-7.

DeGenaro:2015:EDS

[2028] Alan De Genaro and Adilson Simonis. Estimating doubly stochastic Poisson process with affine intensities by Kalman filter. Statistical Papers, 56 (3):723–748, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0606-6.

Zadlo:2015:LMA

[2029] Tomasz Zadlo. On longitudinal moving average model for prediction of subpopulation total. Statistical Papers, 56 (3):749–771, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0607-5;   http:
//link.springer.com/content/pdf/
10.1007/s00362-014-0607-5.pdf.

Singh:2015:RSP

[2030] Sukhdev Singh and Yogesh Mani Tripathi. Reliability sampling plans for a lognormal distribution under progressive first-failure censoring with cost constraint. Statistical Papers, 56 (3):773–817, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0608-4.

Wang:2015:CSL

[2031] Lichun Wang, Yuan You, and Heng Lian. Convergence and sparsity of Lasso and group Lasso in highdimensional generalized linear models. Statistical Papers, 56(3):819–828, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0609-3.

Genback:2015:UIR

[2032] Minna Genbäck, Elena Stanghellini, and Xavier de Luna. Uncertainty intervals for regression parameters with non-ignorable missingness in the outcome. Statistical Papers, 56(3): 829–847, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0610-x;   http:
//link.springer.com/content/pdf/
10.1007/s00362-014-0610-x.pdf.

Haslett:2015:LBM

[2033] S. J. Haslett, S. Puntanen, and B. Arendacká. The link between the mixed and fixed linear models revisited. Statistical Papers, 56(3):849–861, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0611-9.

Morais:2015:QSE

[2034] Manuel Cabral Morais, Yarema Okhrin, and Wolfgang Schmid. Quality surveillance with EWMA control charts based on exact control limits. Statistical Papers, 56(3):863–885, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0612-8.

Benrabah:2015:KME

[2035] Ouafae Benrabah, Elias Ould Saïd, and Abdelkader Tatachak. A kernel mode estimate under random left truncation and time series model: asymptotic normality. Statistical Papers, 56 (3):887–910, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0613-7.

Palanisamy:2015:WBH

[2036] T. Palanisamy and J. Ravichandran.

A wavelet-based hybrid approach to estimate variance function in heteroscedastic regression models. Statistical Papers, 56(3):911–932, August 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0614-6.

Homei:2015:NER

[2037] Hazhir Homei. A novel extension of randomly weighted averages. Statistical Papers, 56(4):933–946, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0615-5.

Kherad-Pajouh:2015:GPA

[2038] Sara Kherad-Pajouh and Olivier Renaud. A general permutation approach for analyzing repeated measures ANOVA and mixed-model designs. Statistical Papers, 56(4): 947–967, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0617-3.

Ou:2015:SLB

[2039] Zujun Ou, Hong Qin, and Hongyi Li. Some lower bounds of centered L_{2}-discrepancy of 2^{s−k} designs and their complementary designs. Statistical Papers, 56(4):969–979, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0618-2.

Melas:2015:OCN

[2040] Viatcheslav B. Melas, Andrey Pepelyshev, Petr Shpilev, Luigi Salmaso, Livio Corain, and Rosa Arboretti. On the optimal choice of the number of empirical Fourier coefficients for comparison of regression curves. Statistical Papers, 56(4):981–997, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0619-1.

Oliveira:2015:PRF

[2041] Paulo E. Oliveira and Nuria Torrado. On proportional reversed failure rate class. Statistical Papers, 56(4): 999–1013, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0620-8.

Alizadeh:2015:EEG

[2042] M. Alizadeh, S. Rezaei, S. F. Bagheri, and S. Nadarajah. Efficient estimation for the generalized exponential distribution. Statistical Papers, 56(4): 1015–1031, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0621-7.

Siburg:2015:AOR

[2043] Karl Friedrich Siburg and Pavel A. Stoimenov. Almost opposite regression dependence in bivariate distributions. Statistical Papers, 56(4): 1033–1039, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0622-6.

Wang:2015:DTR

[2044] Zilin Wang and David Bellhouse. A diagnostic tool for regression analysis of complex survey data. Statistical Papers, 56(4):1041–1053, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0623-5.

Bose:2015:RPE

[2045] Mausumi Bose. Respondent privacy and estimation efficiency in randomized response surveys for discretevalued sensitive variables. Statistical Papers, 56(4):1055–1069, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0624-4.

Amiri:2015:IFP

[2046] Mehdi Amiri, Ahad Jamalizadeh, and Mina Towhidi. Inference and further probabilistic properties of the SUN_{n,2}- distribution. Statistical Papers, 56(4): 1071–1098, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0625-3.

Andres:2015:SIN

[2047] A. Martín Andrés and M. Álvarez Hernández. Simultaneous inferences: new method of maximum combination. Statistical Papers, 56(4):1099– 1113, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0627-1.

Sandberg:2015:EBU

[2048] Rickard Sandberg. M-estimator based unit root tests in the ESTAR framework. Statistical Papers, 56(4):1115– 1135, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0628-0.

Qingguo:2015:RES

[2049] Tang Qingguo. Robust estimation for spatial semiparametric varying coefficient partially linear regression. Statistical Papers, 56(4):1137–1161, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0629-z.

Nicolau:2015:EIM

[2050] João Nicolau and Flavio Ivo Riedlinger. Estimation and inference in multivariate Markov chains. Statistical Papers, 56(4):1163–1173, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0630-6.

Belzunce:2015:SRC

[2051] Félix Belzunce and Carolina Martínez-Riquelme. Some results for the comparison of generalized order statistics in the total time on test and excess wealth orders. Statistical Papers, 56(4): 1175–1190, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0631-5.

Tank:2015:DSM

[2052] Fatih Tank and Serkan Eryilmaz. The distributions of sum, minima and maxima of generalized geometric random variables. Statistical Papers, 56(4): 1191–1203, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0632-4.

Bagnato:2015:ROP

[2053] Luca Bagnato, Valerio Potì, and Maria Grazia Zoia. The role of orthogonal polynomials in adjusting hyperpolic secant and logistic distributions to analyse financial asset returns. Statistical Papers, 56(4):1205–1234, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0633-3.

Arendacka:2015:FRC

[2054] B. Arendacká and S. Puntanen. Further remarks on the connection between fixed linear model and mixed linear model. Statistical Papers, 56(4): 1235–1247, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0634-2.

Kramer:2015:DHR

[2055] Walter Krämer. D. H. Rost: Interpretation und Bewertung pädagogischpsychologischer Studien (3rd edition). (German) []. Statistical Papers, 56(4): 1249–1250, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-014-0616-
                                        4.

Kramer:2015:MJE

[2056] Walter Krämer. Miller, S. J. (ed.): Benford’s law. Theory and applications. Statistical Papers, 56(4):1251– 1252, November 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0717-
                                        8.

Liang:2016:ANC

[2057] Han-Ying Liang and Jong-Il Baek. Asymptotic normality of conditional density estimation with left-truncated and dependent data. Statistical Papers, 57(1):1–20, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0635-1.

Akbari:2016:CBN

[2058] M. Akbari, M. Fashandi, and Jafar Ahmadi. Characterizations based on the numbers of near-order statistics. Statistical Papers, 57(1):21–30, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-014-
0636-0.

Wu:2016:MUR

[2059] Jibo Wu and Hu Yang. More on the unbiased ridge regression estimation. Statistical Papers, 57(1):31–42, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0637-z.

Zmyslony:2016:MDB

[2060] Roman Zmyślony, João T. Mexia, Francisco Carvalho, and Inês J. Sequeira. Mean driven balance and uniformly best linear unbiased estimators. Statistical Papers, 57(1):43–53, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0638-y.

Hasler:2016:HMD

[2061] Mario Hasler. Heteroscedasticity: multiple degrees of freedom vs. sandwich estimation. Statistical Papers, 57(1):55–68, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0640-4.

Yang:2016:PWC

[2062] Hu Yang and Huilan Liu. Penalized weighted composite quantile estimators with missing covariates. Statistical Papers, 57(1):69–88, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00362-014-
0642-2.

Staszewska-Bystrova:2016:IBP

[2063] Anna Staszewska-Bystrova and Peter Winker. Improved bootstrap prediction intervals for SETAR models. Statistical Papers, 57(1):89–98, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0643-1.

Wang:2016:FIC

[2064] Hai Ying Wang, Xinjie Chen, and Nancy Flournoy. The focused information criterion for varying-coefficient partially linear measurement error models. Statistical Papers, 57(1): 99–113, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0645-z.

Yang:2016:VSG

[2065] Hu Yang, Chaohui Guo, and Jing Lv. Variable selection for generalized varying coefficient models with longitudinal data. Statistical Papers, 57(1):115–132, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0647-x.

Suzuki:2016:PIG

[2066] Adriano K. Suzuki, Vicente G. Cancho, and Francisco Louzada. The Poisson– inverse-Gaussian regression model with cure rate: a Bayesian approach and its case influence diagnostics. Statistical Papers, 57(1):133–159, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0649-8.

Wang:2016:NMT

[2067] Li Wang, Xingzhong Xu, and Yong A. New multiple testing method under no dependency assumption, with application to multiple comparisons problem. Statistical Papers, 57(1):161–183, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0650-2.

Jiang:2016:SIC

[2068] Rong Jiang, Wei-Min Qian, and Zhan-Gong Zhou. Single-index composite quantile regression with heteroscedasticity and general error distributions. Statistical Papers, 57 (1):185–203, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0646-y.

Kato:2016:REL

[2069] Shogo Kato and Shinto Eguchi. Robust estimation of location and concentration parameters for the von Mises– Fisher distribution. Statistical Papers, 57(1):205–234, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0648-9.

Sankaran:2016:TEU

[2070] P. G. Sankaran and N. N. Midhu. Testing exponentiality using mean residual quantile function. Statistical Papers, 57(1):235–247, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0651-1.

Zhang:2016:OPA

[2071] Caiya Zhang and Yanbiao Xiang. On the oracle property of adaptive group Lasso in high-dimensional linear models. Statistical Papers, 57(1):249–265, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-015-
                                        0684-0.

Lenz:2016:BR

[2072] Sylvia Tamara Lenz. Book reviews. Statistical Papers, 57(1): 267–269, March 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-014-0644-0.

Ismail:2016:SIS

[2073] Ali A. Ismail. Statistical inference for a step-stress partially-accelerated life test model with an adaptive Type-I progressively hybrid censored data from Weibull distribution. Statistical Papers, 57(2):271–301, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-014-
                                        0639-x.

Westerlund:2016:ADC

[2074] Joakim Westerlund. The asymptotic distribution of the CADF unit root test in the presence of heterogeneous AR(p) errors. Statistical Papers, 57 (2):303–317, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0655-x.

Wong:2016:THI

[2075] A. C. M. Wong. Testing homogeneity of inverse Gaussian scale-like parameters: a saddlepoint approach. Statistical Papers, 57(2):319–327, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0653-z.

Lu:2016:MCF

[2076] Wanbo Lu, Rui Ke, and Jingwen Liang. A moment closed form estimator for the autoregressive conditional duration model. Statistical Papers, 57(2):329–344, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-014-0652-0.

Zhang:2016:DBU

[2077] Xuan Zhang and Yongge Tian. On decompositions of BLUEs under a partitioned linear model with restrictions. Statistical Papers, 57(2):345–364, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-014-
0654-y.

Bieniek:2016:SBB

[2078] Mariusz Bieniek. Sharp bounds on the bias of trimean. Statistical Papers, 57(2):365–379, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-014-0641-3.pdf.

Li:2016:TSE

[2079] Zhaoyuan Li and Jianfeng Yao. On two simple and effective procedures for high dimensional classification of general populations. Statistical Papers, 57(2):381–405, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0660-8.

Yan:2016:LTU

[2080] Rongguo Yan and Lingxiang Zhang. Linearity tests under the null hypothesis of a random walk with drift. Statistical Papers, 57(2):407–418, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0659-1.

Athar:2016:RRB

[2081] Haseeb Athar and Zuber Akhter. Recurrence relations between moments of progressive type-II right censored order statistics from doubly truncated Weibull distribution. Statistical Papers, 57(2):419–428, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0658-2.

Chandra:2016:RCE

[2082] Shalini Chandra and Nityananda Sarkar. A restricted r − k class estimator in the mixed regression model with autocorrelated disturbances. Statistical Papers, 57(2):429–449, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0664-4.

Chen:2016:CEN

[2083] Zhiyong Chen, Haibin Wang, and Xuejun Wang. The consistency for the estimator of nonparametric regression model based on martingale difference errors. Statistical Papers, 57(2):451–469, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0662-6.

Ramos:2016:MSS

[2084] Patrícia Ferreira Ramos, Manuel Cabral Morais, António Pacheco, and Wolfgang Schmid. On the misleading signals in simultaneous schemes for the mean vector and covariance matrix of multivariate i.i.d. output. Statistical Papers, 57(2):471–498, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0663-5.

Niu:2016:THS

[2085] Xiaoqing Niu, Pengfei Li, and Peng Zhang. Testing homogeneity in a scale mixture of normal distributions. Statistical Papers, 57(2):499–516, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0665-3.

Popovic:2016:BIM

[2086] Predrag M. Popović. A bivariate INAR(1) model with different thinning parameters. Statistical Papers, 57(2):517–538, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0667-1.

Mukherjee:2016:PIM

[2087] Partha Sarathi Mukherjee. On phase II monitoring of the probability distributions of univariate continuous processes. Statistical Papers, 57 (2):539–562, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0668-0.

Maronna:2016:BRJ

[2088] Ricardo Maronna. Book review: James Wu and Stephen Coggeshall (2012): Foundations of predictive analytics. Statistical Papers, 57(2):563–564, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-014-0656-9.

Maronna:2016:BRC

[2089] Ricardo Maronna. Book review: Charu C. Aggarwal and Chandan K. Reddy (eds.): Data clustering: algorithms and applications. Statistical Papers, 57(2):565–566, April 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0661-
                                        7.

Hu:2016:AND

[2090] Hongchang Hu, Yu Zhang, and Xiong Pan. Asymptotic normality of DHD estimators in a partially linear model. Statistical Papers, 57(3): 567–587, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0666-
2.

Demirhan:2016:BAE

[2091] Haydar Demirhan and Kamil Demirhan.

A Bayesian approach for the estimation of probability distributions under finite sample space. Statistical Papers, 57 (3):589–603, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0669-
z.

Arshad:2016:EAS

[2092] Mohd. Arshad and Neeraj Misra. Estimation after selection from exponential populations with unequal scale parameters. Statistical Papers, 57(3): 605–621, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0670-
6.

Louzada:2016:TLL

[2093] Francisco Louzada and Daniele C. T.

Granzotto. The transmuted log-logistic regression model: a new model for time up to first calving of cows. Statistical Papers, 57(3):623–640, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-015-0671-5.

Luo:2016:NTR

[2094] Shuanghua Luo and Cheng yi Zhang. Nonparametric M-type regression estimation under missing response data. Statistical Papers, 57(3):641–664, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s00362-015-0672-4.

Yosefi:2016:NAT

[2095] Shima Yosefi, Mohsen Arefi, and Mohammad Ghasem Akbari. A new approach for testing fuzzy hypotheses based on likelihood ratio statistic. Statistical Papers, 57(3):665–688, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-015-0673-3.

Omidi:2016:NMB

[2096] Mehdi Omidi and Mohsen Mohammadzadeh. A new method to build spatio-temporal covariance functions: analysis of ozone data. Statistical Papers, 57(3):689–703, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-015-0674-2.

Cavicchioli:2016:WVR

[2097] Maddalena Cavicchioli. Weak VARMA representations of regime-switching state-space models. Statistical Papers, 57(3):705–720, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
accesspage/article/10.1007/s00362-
015-0675-1.

Smaga:2016:NOC

[2098] Lukasz Smaga. A note on D-optimal chemical balance weighing designs with autocorrelated observations. Statistical Papers, 57(3):721–730, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
015-0676-0.pdf.

Popovic:2016:GBT

[2099] Predrag M. Popović, Miroslav M.

Ristić, and Aleksandar S. Nastić. A geometric bivariate time series with different marginal parameters. Statistical Papers, 57(3):731–753, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0677-z.

Guo:2016:PCC

[2100] Wenxing Guo, Xiaohui Liu, and Shangli Zhang. The principal correlation components estimator and its optimality. Statistical Papers, 57(3): 755–779, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0678-
y.

Yang:2016:EEL

[2101] Geng Yang and Tingting Li. Expansions on extremes from logarithmic general error distribution under power normalization. Statistical Papers, 57 (3):781–793, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0679-
                                        x.

Relvas:2016:PLM

[2102] Carlos Eduardo M. Relvas and Gilberto A. Paula. Partially linear models with first-order autoregressive symmetric errors. Statistical Papers, 57 (3):795–825, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0680-
                                        4.

Peng:2016:VCP

[2103] Qing-Yan Peng, Jian-Jun Zhou, and Nian-Sheng Tang. Varying coefficient partially functional linear regression models. Statistical Papers, 57(3): 827–841, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0681-
                                        3.

Kramer:2016:BRJ

[2104] Walter Krämer. Book review: Joshua D. Angrist and Jörn-Steffen Pischke: Mastering metrics. Statistical Papers, 57(3):843, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-015-0685-z.pdf.

Wied:2016:BRJ

[2105] Dominik Wied. Book review: J.

Bleymüller and R. Weißbach: Statistik für Wirtschaftswissenschaftler (17th edition). Statistical Papers, 57(3): 845, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-015-0700-4.pdf.

Kramer:2016:BRS

[2106] Walter Krämer. Book review: Sabina Alkire, James Foster, Suman Seth, Maria Emma Santos, José Manuel Roche and Paola Ballon: Multidimensional poverty measurement and analysis. Statistical Papers, 57(3): 847–848, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0702-
2.

Lenz:2016:BRA

[2107] Sylvia Tamara Lenz. Book review:

Alan Agresti (2013): Categorical data analysis. Statistical Papers, 57(3): 849–850, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0733-
8.

Kramer:2016:BRU

[2108] Walter Krämer. Book review: Uwe Hassler (2016): Stochastic processes and calculus. An elementary introduction with applications, Springer texts in business and economics. Statistical Papers, 57(3):851, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
016-0746-y.pdf.

Kramer:2016:BRD

[2109] Walter Krämer. Book review: D. Rasch und D. Schott, Mathematische Statistik für Mathematiker, Naturund Ingenieurwissenschaftler. Statistical Papers, 57(3):853, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00362-
016-0752-0.pdf.

Schlittgen:2016:BRC

[2110] Rainer Schlittgen. Book review: C. Hennig, M. Meila, F. Murtagh and R. Rocci (eds.): Handbook of cluster analysis. Handbooks of modern statistical methods. Statistical Papers, 57 (3):855–856, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0785-
4;   http://link.springer.com/article/
10.1007/s00362-016-0785-4.

Liu:2016:BRF

[2111] Shuangzhe Liu. Book review: Fieller, N.: Basics of Matrix Algebra for Statistics with R. Statistical Papers, 57(3): 857–858, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0813-

4;   http://link.springer.com/article/
10.1007/s00362-016-0813-4.

Hassler:2016:BRM

[2112] Uwe Hassler. Book review: M. H.

Pesaran (2015): Time series and panel data econometrics. Oxford University Press, Oxford, 1104 pp, Hardcover 110.00 £££, ISBN: 978-0-19-873691-2. Statistical Papers, 57(3): 859–860, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0816-
1;   http://link.springer.com/article/
10.1007/s00362-016-0816-1.

Hackl:2016:BRG

[2113] Peter Hackl. Book review: Giovanni Cerulli: Econometric evaluation of socio-economic programs: theory and applications. Statistical Papers, 57(3): 861–862, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0817-
0;   http://link.springer.com/article/
10.1007/s00362-016-0817-0.

Anonymous:2016:PSa

[2114] Anonymous. Problem section. Statistical Papers, 57(3):863–866, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-014-0657-8;   http://link.
springer.com/article/10.1007/s00362-
014-0657-8.

Anonymous:2016:PSb

[2115] Anonymous. Problem section. Statistical Papers, 57(3):867–870, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-015-0716-9;   http://link.
                                        springer.com/article/10.1007/s00362-
                                        015-0716-9.

Harman:2016:ESI

[2116] Radoslav Harman and Werner G. Müller. Editorial for the special issue PROBASTAT 2015. Statistical Papers, 57(4):871–873, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0825-0;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        016-0825-0.pdf.

Amo-Salas:2016:ODM

[2117] Mariano Amo-Salas, Elvira Delgado-Márquez, Lenka Filová, and Jesús López-Fidalgo. Optimal designs for model discrimination and fitting for the flow of particles. Statistical Papers, 57 (4):875–891, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-016-0792-
                                        5;   http://link.springer.com/article/
                                        10.1007/s00362-016-0792-5.

Burclova:2016:ODE

[2118] Katarína Burclová and Andrej Pázman. Optimal design of experiments via linear programming. Statistical Papers, 57 (4):893–910, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0782-
7;   http://link.springer.com/article/
10.1007/s00362-016-0782-7.

Kurtoglu:2016:LEG

[2119] Fikriye Kurtoglu and M. Revan Özkale. Liu estimation in generalized linear models: application on gamma distributed response variable. Statistical Papers, 57(4): 911–928, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0814-
3;   http://link.springer.com/article/
10.1007/s00362-016-0814-3.

LaMotte:2016:IPM

[2120] Lynn R. LaMotte and Jeffrey D.

Wells. Inverse prediction for multivariate mixed models with standard software. Statistical Papers, 57(4): 929–938, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0815-
2;   http://link.springer.com/article/
10.1007/s00362-016-0815-2.

Maciak:2016:RTJ

[2121] Matús Maciak and Ivan Mizera. Regularization techniques in joinpoint regression. Statistical Papers, 57(4): 939–955, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0823-

                                        2;   http://link.springer.com/article/
                                        10.1007/s00362-016-0823-2.

Meintanis:2016:GFT

[2122] Simos G. Meintanis, James Allison, and Leonard Santana. Goodness-of-fit tests for semiparametric and parametric hypotheses based on the probability weighted empirical characteristic function. Statistical Papers, 57 (4):957–976, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-016-0760-
                                        0;   http://link.springer.com/article/
                                        10.1007/s00362-016-0760-0.

Milosevic:2016:NCE

[2123] Bojana Milosević and Marko Obradović. New class of exponentiality tests based on U-empirical Laplace transform. Statistical Papers, 57(4): 977–990, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-016-0818-
                                        z;   http://link.springer.com/article/
                                        10.1007/s00362-016-0818-z.

Ozkale:2016:IAB

[2124] M. Revan Özkale. Iterative algorithms of biased estimation methods in binary logistic regression. Statistical Papers, 57(4):991–1016, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-016-0780-9;   http://link.
                                        springer.com/article/10.1007/s00362-
                                        016-0780-9.

Patrangenaru:2016:PSA

[2125] Victor Patrangenaru, Robert Paige, K. David Yao, Mingfei Qiu, and David Lester. Projective shape analysis of contours and finite 3D configurations from digital camera images. Statistical Papers, 57(4):1017– 1040, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0824-
1;   http://link.springer.com/article/
10.1007/s00362-016-0824-1.

Pesta:2016:UIE

[2126] Michal Pesta. Unitarily invariant errors-in-variables estimation. Statistical Papers, 57(4):1041–1057, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-016-0800-9;   http://link.
springer.com/article/10.1007/s00362-
016-0800-9.

Pronzato:2016:EMM

[2127] Luc Pronzato, Henry P. Wynn, and Anatoly Zhigljavsky. Extremal measures maximizing functionals based on simplicial volumes. Statistical Papers, 57(4):1059– 1075, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0767-
6;   http://link.springer.com/article/
10.1007/s00362-016-0767-6.

Rosa:2016:OAD

[2128] Samuel Rosa and Radoslav Harman. Optimal approximate designs for estimating treatment contrasts resistant to nuisance effects. Statistical Papers, 57(4):1077– 1106, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0809-
0;   http://link.springer.com/article/
10.1007/s00362-016-0809-0.

Sofronov:2016:MOS

[2129] Georgy Yu. Sofronov. A multiple optimal stopping rule for a buying– selling problem with a deterministic trend. Statistical Papers, 57(4): 1107–1119, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-016-0776-
5;   http://link.springer.com/article/
10.1007/s00362-016-0776-5.

Wichitsa-nguan:2016:ECM

[2130] Korakot Wichitsa-nguan, Henning Läuter, and Hannelore Liero. Estimability in Cox models. Statistical Papers, 57(4):1121–1140, December 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-016-0755-x;   http://link.
springer.com/article/10.1007/s00362-
016-0755-x.

Kala:2017:SNL

[2131] Radosław Kala, Simo Puntanen, and Yongge Tian. Some notes on linear sufficiency. Statistical Papers, 58(1):1–17, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0682-2.

Saez-Castillo:2017:DUD

[2132] Antonio J. Sáez-Castillo and Antonio Conde-Sánchez. Detecting overand under-dispersion in zero inflated data with the hyper-Poisson regression model. Statistical Papers, 58(1):19–33, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0683-1.

Fuchs:2017:CIP

[2133] Nicole Fuchs, Werner Pölz, and Arne C. Bathke. Confidence intervals for population means of partially paired observations. Statistical Papers, 58(1):35–51, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00362-015-0686-
y.

Xu:2017:HTR

[2134] Hong-Xia Xu, Guo-Liang Fan, and Han-Ying Liang. Hypothesis test on response mean with inequality constraints under data missing when covariables are present. Statistical Papers, 58(1):53–75, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0687-x.

Gao:2017:ODB

[2135] Lucy L. Gao and Julie Zhou. D-optimal designs based on the second-order least squares estimator. Statistical Papers, 58(1):77–94, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0688-
                                        9.

Liu:2017:JEL

[2136] Ai-Ai Liu and Han-Ying Liang. Jackknife empirical likelihood of error variance in partially linear varyingcoefficient errors-in-variables models. Statistical Papers, 58(1):95–122, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-015-0689-8.

Hahn:2017:CDS

[2137] Sonja Hahn and Luigi Salmaso. A comparison of different synchronized permutation approaches to testing effects in two-level two-factor unbalanced ANOVA designs. Statistical Papers, 58(1):123–146, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0690-
                                        2.

Park:2017:PCS

[2138] Heewon Park and Sadanori Konishi. Principal component selection via adaptive regularization method and generalized information criterion. Statistical Papers, 58(1):147–160, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0691-1.

Westerlund:2017:DNF

[2139] Joakim Westerlund and Sagarika Mishra. On the determination of the number of factors using information criteria with data-driven penalty. Statistical Papers, 58(1):161–184, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0692-0.

Mojirsheibani:2017:KRE

[2140] Majid Mojirsheibani and Timothy Reese. Kernel regression estimation for incomplete data with applications. Statistical Papers, 58(1):185–209, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0693-z.

Nguyen:2017:KPB

[2141] Thuy Tuong Nguyen and Yury Tsoy. A kernel PLS based classification method with missing data handling. Statistical Papers, 58(1):211–225, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-015-0694-y.

Wang:2017:RFS

[2142] Guochang Wang, Jianjun Zhou, Wuqing Wu, and Min Chen. Robust functional sliced inverse regression. Statistical Papers, 58(1):227–245, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-015-0695-x.

Garay:2017:LCR

[2143] Aldo M. Garay, Victor H. Lachos, Heleno Bolfarine, and Celso R. B. Cabral. Linear censored regression models with scale mixtures of normal distributions. Statistical Papers, 58(1):247–278, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-015-0696-
                                        9.

Hackl:2017:BRA

[2144] Peter Hackl. Book review: Alex Bottle, Paul Aylin: Statistical methods for healthcare performance monitoring. Statistical Papers, 58(1):279–280, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00362-016-0836-x.

Liu:2017:BRM

[2145] Shuangzhe Liu. Book review: Moss, C. B.: Mathematical statistics for applied econometrics. Statistical Papers, 58(1):281–282, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00362-016-0857-
                                        5.

Hassler:2017:BRP

[2146] Uwe Hassler. Book review: Palma, W.: Time series analysis. Statistical Papers, 58(1):283–284, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-016-0858-4.

Hackl:2017:BRN

[2147] Peter Hackl. Book review: Nigel C. Smeeton, Dental Statistics Made Easy, Third Edition. Chapman and Hall/CRC Textbook, 2016, xiv + 198 pp., BP 38.00, ISBN 978-1-4987-7505-2. Statistical Papers, 58(1):285–286, March 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00362-017-0872-1.

Kayid:2017:SRR

[2148] M. Kayid, S. Izadkhah, and Ming J.

Zuo. Some results on the relative ordering of two frailty models. Statistical Papers, 58(2):287–301, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wang:2017:CES

[2149] Xuejun Wang, Xin Deng, Fengxi Xia, and Shuhe Hu. The consistency for the estimators of semiparametric regression model based on weakly dependent errors. Statistical Papers, 58(2):303–318, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Nematollahi:2017:AME

[2150] Nader Nematollahi. Admissible and minimax estimation of the parameter of the selected Pareto population under squared log error loss function. Statistical Papers, 58(2):319–339, June 2017.

CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Ghosh:2017:MDE

[2151] Abhik Ghosh and Ayanendranath Basu. The minimum S-divergence estimator under continuous models: the Basu–Lindsay approach. Statistical Papers, 58(2):341–372, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Hurlimann:2017:CEF

[2152] Werner Hürlimann. A comprehensive extension of the FGM copula. Statistical Papers, 58(2):373–392, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Khoo:2017:MTS

[2153] Wooi Chen Khoo, Seng Huat Ong, and Atanu Biswas. Modeling time series of counts with a new class of INAR(1) model. Statistical Papers, 58(2):393–416, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Zhang:2017:AWN

[2154] Tianyong Zhang, Demei Yuan, Jiali Ma, and Xuemei Hu. Assessing white noise assumption with semi-parametric additive partial linear models. Statistical Papers, 58(2):417–431, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Calinski:2017:MMA

[2155] T. Caliński, S. Czajka, Z. Kaczmarek, P. Krajewski, W. Pilarczyk, I. Siatkowski, and M. Siatkowski. On a mixed model analysis of multienvironment variety trials: a reconsideration of the one-stage and the twostage models and analyses. Statistical Papers, 58(2):433–465, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Tian:2017:SEI

[2156] Y. Tian. Some equalities and inequalities for covariance matrices of estimators under linear model. Statistical Papers, 58(2):467–484, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Dey:2017:CCM

[2157] Rajib Dey and M. Ataharul Islam.

A conditional count model for repeated count data and its application to GEE approach. Statistical Papers, 58(2):485–504, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Lemdani:2017:NRR

[2158] Mohamed Lemdani and Elias Ould Saïd. Nonparametric robust regression estimation for censored data. Statistical Papers, 58(2):505–525, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Ghiglietti:2017:LRM

[2159] Andrea Ghiglietti, Francesca Ieva, Anna Maria Paganoni, and Giacomo Aletti. On linear regression models in infinite dimensional spaces with scalar response. Statistical Papers, 58(2):527–548, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wachter:2017:BRU

[2160] Jasmin Wachter. Book review: Unwin, A., Graphical Data Analysis with R, Chapman and Hall/CRC, 2015, 310 pp., £49.99, ISBN 978-1-4987-1523-2. Statistical Papers, 58(2):549–551, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Xu:2017:BRA

[2161] Zheng Xu. Book review: Andy Hector (2015): The new statistics with R: an introduction for biologists. Oxford University Press, 224 pp., $125.00 (hardcover), $49.95 (paper), ISBN 978-0-19-872906-8. Statistical Papers, 58 (2):553–554, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

deMatosRibeiro:2017:BRJ

[2162] Patrick de Matos Ribeiro, Lukas Matuschek, and Martin Wagner. Book review: Jose Casals, Alfredo Garcia– Hiernaux, Miguel Jerez, Sonia Sotoca and A. Alexandre Trindade (2016): State-space methods for time series analysis: theory, applications and software. Statistical Papers, 58(2):555–556, June 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

Wegener:2017:FNU

[2163] Michael Wegener and Göran Kauermann. Forecasting in nonlinear univariate time series using penalized splines. Statistical Papers, 58(3): 557–576, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0711-1.

Markeviciute:2017:TEC

[2164] Jurgita Markeviciute, Alfredas Rackauskas, and Charles Suquet. Testing epidemic change in nearly nonstationary process with statistics based on residuals. Statistical Papers, 58(3): 577–606, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0712-0.

Kundu:2017:CBH

[2165] Chanchal Kundu and Kshirod Sarkar.

Characterizations based on higher order and partial moments of inactivity time. Statistical Papers, 58(3): 607–626, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0714-y.

Yang:2017:MTU

[2166] Mengta Yang and Reza Modarres. Multivariate tests of uniformity. Statistical Papers, 58(3):627–639, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0715-x.

DeLuca:2017:DTD

[2167] Giovanni De Luca and Paola Zuccolotto. Dynamic tail dependence clustering of financial time series. Statistical Papers, 58(3):641–657, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0718-7.

Zhang:2017:BCE

[2168] Qiang Zhang, Lijun Wu, and Qianqian Cui. The balanced credibility estimators with correlation risk and inflation factor. Statistical Papers, 58 (3):659–672, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0719-6.

Chatrabgoun:2017:LMA

[2169] O. Chatrabgoun, G. Parham, and R. Chinipardaz. A Legendre multiwavelets approach to copula density estimation. Statistical Papers, 58(3): 673–690, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0720-0.

Li:2017:BSV

[2170] Jianbo Li, Yuan Li, and Riquan Zhang. B spline variable selection for the single index models. Statistical Papers, 58 (3):691–706, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0721-z.

Chen:2017:CSE

[2171] Zhuoheng Chen and Yijun Hu. Cumulative sum estimator for change-point in panel data. Statistical Papers, 58 (3):707–728, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0722-y.

Arumairajan:2017:GPT

[2172] Sivarajah Arumairajan and Pushpakanthie Wijekoon. The generalized preliminary test estimator when different sets of stochastic restrictions are available. Statistical Papers, 58(3): 729–747, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0723-x.

Rodriguez-Avi:2017:RMO

[2173] José Rodríguez-Avi and María José Olmo-Jiménez. A regression model for overdispersed data without too many zeros. Statistical Papers, 58(3): 749–773, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0724-9.

Finner:2017:STU

[2174] H. Finner, M. Roters, and K. Strassburger. On the Simes test under dependence. Statistical Papers, 58(3): 775–789, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0725-8.

Hassanzadeh:2017:RMO

[2175] Fatemeh Hassanzadeh and Iraj Kazemi.

Regression modeling of one-inflated positive count data. Statistical Papers, 58(3):791–809, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s00362-015-0726-

Xu:2017:ODF

[2176] Xiaojian Xu and Xiaoli Shang. Doptimal designs for full and reduced Fourier regression models. Statistical Papers, 58(3):811–829, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-015-
                                        0727-6.

Kamalja:2017:MBD

[2177] K. K. Kamalja. Markov binomial distribution of order k and its applications. Statistical Papers, 58(3): 831–853, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0728-5.

Rahmani:2017:PRT

[2178] Hamid Rahmani and Mostafa Razmkhah. Perfect ranking test in moving extreme ranked set sampling. Statistical Papers, 58(3):855–875, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-015-
                                        0729-4.

Emura:2017:AIM

[2179] Takeshi Emura, Ya-Hsuan Hu, and Yoshihiko Konno. Asymptotic inference for maximum likelihood estimators under the special exponential family with double-truncation. Statistical Papers, 58(3):877–909, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-015-
0730-y.

Shen:2017:MIN

[2180] Aiting Shen, Yu Zhang, Benqiong Xiao, and Andrei Volodin. Moment inequalities for m-negatively associated random variables and their applications. Statistical Papers, 58(3): 911–928, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0731-x.

Salmeron:2017:NAC

[2181] R. Salmerón, J. García, C. B. García, and M. M. López Martín. A note about the corrected VIF. Statistical Papers, 58(3):929–945, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-015-
0732-9.

Kleiber:2017:BFM

[2182] Christian Kleiber. Book review:

Belzunce, F., Martínez–Riquelme, C. and J. Mulero: An Introduction to Stochastic Orders. Academic Press, New York, 2016, 174 pp., EUR 53.95 (print), ISBN 978-0-12-803768-3. Statistical Papers, 58(3):947–949, September 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0914-8.

Weba:2017:ADM

[2183] Michael Weba and Nora Dörmann.

Application of the delta method to functions of the sample mean when observations are dependent. Statistical Papers, 58(4):957–986, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-015-
                                        0734-7.

Attouch:2017:NRE

[2184] Mohammed Attouch, Ali Laksaci, and Nafissa Messabihi. Nonparametric relative error regression for spatial random variables. Statistical Papers, 58 (4):987–1008, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0735-6.

Guo:2017:RVS

[2185] Chaohui Guo, Hu Yang, and Jing Lv. Robust variable selection in highdimensional varying coefficient models based on weighted composite quantile regression. Statistical Papers, 58(4): 1009–1033, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0736-5.

Aghamohammadi:2017:BAP

[2186] A. Aghamohammadi and S. Mohammadi. Bayesian analysis of penalized quantile regression for longitudinal data. Statistical Papers, 58(4): 1035–1053, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-015-0737-4.

Mosler:2017:FDC

[2187] Karl Mosler and Pavlo Mozharovskyi.

Fast DD-classification of functional data. Statistical Papers, 58(4):1055– 1089, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-015-0738-3.

Tang:2017:SBI

[2188] Nian-Sheng Tang, De-Wang Li, and An-Min Tang. Semiparametric Bayesian inference on generalized linear measurement error models. Statistical Papers, 58(4):1091– 1113, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00362-016-0739-x.pdf;
http://link.springer.com/article/
10.1007/s00362-016-0739-x.

Sepehrifar:2017:DRD

[2189] Mohammad Sepehrifar and Shantia Yarahmadian. Decreasing renewal dichotomous Markov noise shock model with hypothesis testing applications. Statistical Papers, 58(4): 1115–1124, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0740-4.

Popovic:2017:MSA

[2190] Bozidar V. Popović, Miroslav M.

Ristić, and Narayana Balakrishna. A mixed stationary autoregressive model with exponential marginals. Statistical Papers, 58(4):1125–1148, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0741-3.

Frey:2017:ECP

[2191] Jesse Frey and Timothy G. Feeman. Efficiency comparisons for partially rank-ordered set sampling. Statistical Papers, 58(4):1149–1163, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0742-2.

Krupskii:2017:CBM

[2192] Pavel Krupskii. Copula-based measures of reflection and permutation asymmetry and statistical tests. Statistical Papers, 58(4):1165–1187, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0743-1.

Ahlgren:2017:WBT

[2193] Niklas Ahlgren and Paul Catani. Wild bootstrap tests for autocorrelation in vector autoregressive models. Statistical Papers, 58(4):1189–1216, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0744-0.

Choi:2017:RDR

[2194] Jin young Choi and Myoung jae Lee. Regression discontinuity: review with extensions. Statistical Papers, 58(4): 1217–1246, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0745-z.

Clarke:2017:CLM

[2195] Brenton R. Clarke, Thomas Davidson, and Robert Hammarstrand. A comparison of the L_{2} minimum distance estimator and the EM-algorithm when fitting k-component univariate normal mixtures. Statistical Papers, 58(4): 1247–1266, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0747-x.

Nematollahi:2017:PCM

[2196] A. R. Nematollahi, A. R. Soltani, and M. R. Mahmoudi. Periodically correlated modeling by means of the periodograms asymptotic distributions. Statistical Papers, 58(4): 1267–1278, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0748-9.

Kleiber:2017:BRJ

[2197] Christian Kleiber. Book review: John M. Chambers (2016): Extending R. Chapman and Hall/CRC Press, 364 pp., ISBN 978-1-4987-7571-7, GBP 44.99 (print), GBP 31.49 (eBook). Statistical Papers, 58(4):1279–1280, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0942-4.

Hackl:2017:BRC

[2198] Peter Hackl. Book review: Claus Thorn Ekstrøm (2017): The R Primer, Second Edition, Chapman & Hall/ CRC, The R Series, xvii + 408 pp., £39.99, ISBN 978-1-138-63197-7. Statistical Papers, 58(4):1281–1282, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0954-0.

Mosler:2017:BRE

[2199] Karl Mosler. Book review: Ernesto Estrada and Philip A. Knight (2015): A First Course in Network Theory, Oxford University Press, 272 pp., £29.99, ISBN 978-0-19-872646-3. Statistical Papers, 58(4):1283–1284, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0961-1.

Jiang:2018:DEE

[2200] Sha Jiang, Tingting Li, and Xin Liao. Distributional expansions on extremes from skew-normal distribution under power normalization. Statistical Papers, 59(1):1–20, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0749-8.

Singh:2018:EPI

[2201] Sukhdev Singh and Yogesh Mani Tripathi. Estimating the parameters of an inverse Weibull distribution under progressive Type-I interval censoring. Statistical Papers, 59(1):21–56, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0750-2.

Su:2018:CGD

[2202] Nan-Cheng Su and Wan-Ping Hung.

Characterizations of the geometric distribution via residual lifetime. Statistical Papers, 59(1):57–73, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0751-1.

Maior:2018:SND

[2203] Vinicius Q. S. Maior and Francisco José A. Cysneiros. SYMARMA: a new dynamic model for temporal data on conditional symmetric distribution. Statistical Papers, 59(1):75–97, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0753-z.

Wang:2018:LAE

[2204] Yuebao Wang, Hui Xu, Dongya Cheng, and Changjun Yu. The local asymptotic estimation for the supremum of a random walk with generalized strong subexponential summands. Statistical Papers, 59(1):99–126, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0754-y.

Saadi:2018:SNT

[2205] Nora Saadi, Smail Adjabi, and Ali Gannoun. The selection of the number of terms in an orthogonal series cumulative function estimator. Statistical Papers, 59(1):127–152, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0756-9.

Gorecki:2018:SSM

[2206] Tomasz Górecki, Mirosław Krzyśko, Łukasz Waszak, and Waldemar Wolyński. Selected statistical methods of data analysis for multivariate functional data. Statistical Papers, 59(1):153–182, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0757-8;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        016-0757-8.pdf.

Al-Mosawi:2018:EMS

[2207] Riyadh Rustam Al-Mosawi and Shahjahan Khan. Estimating moments of a selected Pareto population under asymmetric scale invariant loss function. Statistical Papers, 59(1):183–198, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0758-7.

Cha:2018:SCP

[2208] Ji Hwan Cha and Maxim Finkelstein. On stochastic comparisons for population age and remaining lifetime. Statistical Papers, 59(1):199–213, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0759-6.

Cizek:2018:BCQ

[2209] P. Cízek and S. Sadikoglu. Biascorrected quantile regression estimation of censored regression models. Statistical Papers, 59(1):215–247, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0761-z;   http://link.springer.
com/content/pdf/10.1007/s00362-
016-0761-z.pdf.

Inoue:2018:JDN

[2210] Kiyoshi Inoue and Sigeo Aki. Joint distributions of numbers of runs of specified lengths on directed trees. Statistical Papers, 59(1):249–269, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0762-y.

Yavuz:2018:LMM

[2211] Fulya Gokalp Yavuz and Olcay Arslan.

Linear mixed model with Laplace distribution (LLMM). Statistical Papers, 59(1):271–289, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0763-x.

Hsieh:2018:RRT

[2212] Shu-Hui Hsieh, Shen-Ming Lee, and Su-Hao Tu. Randomized response techniques for a multi-level attribute using a single sensitive question. Statistical Papers, 59(1):291–306, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0764-9.

Kizilaslan:2018:ERM

[2213] Fatih Kizilaslan and Mustafa Nadar. Estimation of reliability in a multicomponent stress-strength model based on a bivariate Kumaraswamy distribution. Statistical Papers, 59(1):307–340, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0765-8.

Modarres:2018:MID

[2214] Reza Modarres. Multinomial interpoint distances. Statistical Papers, 59 (1):341–360, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0766-7.

Lamboni:2018:GSA

[2215] Matieyendou Lamboni. Global sensitivity analysis: a generalized, unbiased and optimal estimator of totaleffect variance. Statistical Papers, 59 (1):361–386, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0768-5.

Tran:2018:MRP

[2216] Kim Phuc Tran, Philippe Castagliola, and Giovanni Celano. Monitoring the ratio of population means of a bivariate normal distribution using CUSUM type control charts. Statistical Papers, 59(1):387–413, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0769-4.

Maronna:2018:BRC

[2217] Ricardo Maronna. Book review: Claus Weihs, Dietmar Jannach, Igor Vatolkin and Günter Rudolph (eds.), Music Data Analysis: Foundations and Applications, Chapman & Hall/CRC, 2016, 676 pp., $89.95, ISBN 978-1-4987-1956-8. Statistical Papers, 59(1):415–416, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0976-7.

Hassler:2018:BRW

[2218] Uwe Hassler. Book review: Wayne A. Woodward, Henry L. Gray and Alan C. Elliott (2017): Applied Time Series Analysis with R, Second Edition, Chapman & Hall/CRC, 618 pp., $109.95, ISBN 978-1-4987-3422-6. Statistical Papers, 59(1):417–418, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0977-6.

Hackl:2018:BRJ

[2219] Peter Hackl. Book review: Jelke Bethlehem (2017): Understanding Public Opinion Polls, Chapman and Hall/ CRC, 286 pp. + xii, $59.95, ISBN 978-1-4987-6974-7. Statistical Papers, 59(1):419–420, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-0981-5.

Ferrario:2018:PEL

[2220] Paola Gloria Ferrario. Partitioning estimation of local variance based on nearest neighbors under censoring. Statistical Papers, 59(2):423–447, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0770-y.

Yang:2018:CCE

[2221] Wenzhi Yang, Haiyun Xu, Ling Chen, and Shuhe Hu. Complete consistency of estimators for regression models based on extended negatively dependent errors. Statistical Papers, 59(2):449–465, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0771-x.

Condino:2018:LBE

[2222] Francesca Condino, Filippo Domma, and Giovanni Latorre. Likelihood and Bayesian estimation of P (Y < X) using lower record values from a proportional reversed hazard family. Statistical Papers, 59(2):467–485, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0772-9.

Shi:2018:SIH

[2223] Jianhong Shi and Fanrong Zhao. Statistical inference for heteroscedastic semi-varying coefficient EV models under restricted condition. Statistical Papers, 59(2):487–511, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0773-8.

Qarmalah:2018:BMD

[2224] Najla M. Qarmalah, Jochen Einbeck, and Frank P. A. Coolen. k-boxplots for mixture data. Statistical Papers, 59(2):513–528, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0774-7;   http:
//link.springer.com/content/pdf/
10.1007/s00362-016-0774-7.pdf.

Emami:2018:LIL

[2225] Hadi Emami. Local influence for Liu estimators in semiparametric linear models. Statistical Papers, 59(2):529–544, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0775-6.

Aerts:2018:DUE

[2226] S. Aerts, G. Haesbroeck, and C. Ruwet.

Distribution under elliptical symmetry of a distance-based multivariate coefficient of variation. Statistical Papers, 59(2):545–579, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0777-4.

Xu:2018:GVT

[2227] Liwen Xu, Hongxia Guo, and Shenghua Yu. Generalized p value tests for variance components in a class of linear mixed models. Statistical Papers, 59(2):581–604, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0778-3.

Barakat:2018:PFG

[2228] H. M. Barakat, E. M. Nigm, Magdy E. El-Adll, and M. Yusuf. Prediction of future generalized order statistics based on exponential distribution with random sample size. Statistical Papers, 59(2):605–631, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0779-2.

Hirose:2018:SFR

[2229] Kei Hirose and Miyuki Imada. Sparse factor regression via penalized maximum likelihood estimation. Statistical Papers, 59(2):633–662, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0781-8.

Golosnoy:2018:SMP

[2230] Vasyl Golosnoy. Sequential monitoring of portfolio betas. Statistical Papers, 59(2):663–684, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0783-6.

Shen:2018:QRE

[2231] Yu Shen and Han-Ying Liang. Quantile regression and its empirical likelihood with missing response at random. Statistical Papers, 59(2):685–707, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0784-5.

Ghapani:2018:WRE

[2232] F. Ghapani, A. R. Rasekh, and B. Babadi. The weighted ridge estimator in stochastic restricted linear measurement error models. Statistical Papers, 59(2):709–723, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0786-3.

El-Sagheer:2018:EPW

[2233] Rashad M. El-Sagheer. Estimation of parameters of Weibull–gamma distribution based on progressively censored data. Statistical Papers, 59 (2):725–757, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0787-2;   http:
//link.springer.com/content/pdf/
10.1007/s00362-016-0787-2.pdf.

Asgharzadeh:2018:SIB

[2234] A. Asgharzadeh, A. Fallah, M. Z.

Raqab, and R. Valiollahi. Statistical inference based on Lindley record data. Statistical Papers, 59(2):759–779, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0788-1.

Navarro:2018:DFC

[2235] Jorge Navarro. Distribution-free comparisons of residual lifetimes of coherent systems based on copula properties. Statistical Papers, 59(2):781–800, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0789-0.

Wheeler:2018:IDE

[2236] Graham M. Wheeler. Incoherent dose-escalation in phase I trials using the escalation with overdose control approach. Statistical Papers, 59 (2):801–811, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0790-7;   http:
//link.springer.com/content/pdf/
10.1007/s00362-016-0790-7.pdf.

Amiri:2018:REL

[2237] Aboubacar Amiri and Baba Thiam. Regression estimation by local polynomial fitting for multivariate data streams. Statistical Papers, 59 (2):813–843, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0791-6.

Ristic:2018:BRL

[2238] Miroslav M. Ristić. Book review: Lyle D. Broemeling: Bayesian inference for stochastic processes. Statistical Papers, 59(2):845–846, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0991-3.

Husch:2018:BRT

[2239] Marc Hüsch. Book review: Thomas Rahlf: Data Visualisation with R. Statistical Papers, 59(2):847–848, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1003-3.

Li:2018:ETG

[2240] Benchong Li and Liya Fu. Exact test of goodness of fit for binomial distribution. Statistical Papers, 59(3): 851–860, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0793-4.

Mahdizadeh:2018:NRM

[2241] M. Mahdizadeh and Ehsan Zamanzade.

A new reliability measure in ranked set sampling. Statistical Papers, 59 (3):861–891, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0794-3.

Kuang:2018:LSE

[2242] Nenghui Kuang and Bingquan Liu.

Least squares estimator for α-sub-fractional bridges. Statistical Papers, 59(3):893–912, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.com/
article/10.1007/s00362-016-0795-
2.

Olive:2018:AHP

[2243] David J. Olive. Applications of hyperellipsoidal prediction regions. Statistical Papers, 59(3):913–931, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0796-1.

Guo:2018:CCC

[2244] Baocai Guo and Bing Xing Wang. Control charts for the coefficient of variation. Statistical Papers, 59(3): 933–955, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0797-0.

Guo:2018:TSE

[2245] Chaohui Guo, Hu Yang, and Jing Lv. Two step estimations for a single-index varying-coefficient model with longitudinal data. Statistical Papers, 59(3): 957–983, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0798-z.

Gao:2018:FSO

[2246] Xiaoli Gao. A flexible shrinkage operator for fussy grouped variable selection. Statistical Papers, 59(3): 985–1008, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0799-y.

Menni:2018:NEC

[2247] Nassira Menni and Abdelkader Tatachak. A note on estimating the conditional expectation under censoring and association: strong uniform consistency. Statistical Papers, 59(3): 1009–1030, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0801-8.

Jeliazkov:2018:IEU

[2248] Ivan Jeliazkov and Angela Vossmeyer.

The impact of estimation uncertainty on covariate effects in nonlinear models. Statistical Papers, 59(3):1031– 1042, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0802-7.

Ebner:2018:FIS

[2249] Bruno Ebner, Bernhard Klar, and Simos G. Meintanis. Fourier inference for stochastic volatility models with heavy-tailed innovations. Statistical Papers, 59(3):1043–1060, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0803-6.

Ghosh:2018:CDC

[2250] Amit Ghosh and Chanchal Kundu.

Chernoff distance for conditionally specified models. Statistical Papers, 59(3):1061–1083, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-016-
                                        0804-5.

Zamanzade:2018:SNT

[2251] Ehsan Zamanzade and Michael Vock. Some nonparametric tests of perfect judgment ranking for judgment post stratification. Statistical Papers, 59(3): 1085–1100, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0805-4.

Lu:2018:TCL

[2252] Changli Lu, Yuqin Sun, and Yongge Tian. Two competing linear randomeffects models and their connections. Statistical Papers, 59(3):1101– 1115, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0806-3.

Wu:2018:NCE

[2253] Yi Wu and Xuejun Wang. A note on the consistency for the estimators of semiparametric regression model. Statistical Papers, 59(3): 1117–1130, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0807-2.

Yang:2018:IVT

[2254] Kai Yang, Dehui Wang, Boting Jia, and Han Li. An integer-valued threshold autoregressive process based on negative binomial thinning. Statistical Papers, 59(3):1131–1160, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0808-1.

Kizilaslan:2018:CBE

[2255] Fatih Kizilaslan. Classical and Bayesian estimation of reliability in a multicomponent stress-strength model based on a general class of inverse exponentiated distributions. Statistical Papers, 59(3):1161–1192, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0810-7.

Wu:2018:BLU

[2256] Jibo Wu and Chaolin Liu. The best linear unbiased estimator in a singular linear regression model. Statistical Papers, 59(3):1193–1204, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0811-6.

AlMohamad:2018:TBU

[2257] Diaa Al Mohamad. Towards a better understanding of the dual representation of phi divergences. Statistical Papers, 59(3):1205–1253, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0812-5.

Huda:2018:ODS

[2258] S. Huda and Rahul Mukerjee. Optimal designs with string property under asymmetric errors and SLS estimation. Statistical Papers, 59(3): 1255–1268, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0819-y.

Hu:2018:BRA

[2259] Jiang Hu. Book review: Arup Bose (2018): Patterned random matrices, Chapman & Hall/CRC, 291 pp. + xxi, Hardcover, $129.95, ISBN:978-1-138-59146-2. Statistical Papers, 59(3): 1269–1270, September 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1024-y.

Satin:2018:UBR

[2260] Yacov Satin, Evsey Morozov, Ruslana Nekrasova, Alexander Zeifman, Ksenia Kiseleva, Anna Sinitcina, Alexander Sipin, Galina Shilova, and Irina Gudkova. Upper bounds on the rate of convergence for constant retrial rate queueing model with two servers. Statistical Papers, 59(4):1271–1282, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1014-0.

Lee:2018:BTT

[2261] Namgil Lee and Jong-Min Kim. Block tensor train decomposition for missing data estimation. Statistical Papers, 59(4):1283–1305, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1043-8.

Casero-Alonso:2018:WBT

[2262] Víctor Casero-Alonso, Andrey Pepelyshev, and Weng K. Wong. A webbased tool for designing experimental studies to detect hormesis and estimate the threshold dose. Statistical Papers, 59(4):1307–1324, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1038-5.

Noonan:2018:ABC

[2263] Jack Noonan and Anatoly Zhigljavsky.

Approximations of the boundary crossing probabilities for the maximum of moving weighted sums. Statistical Papers, 59(4):1325–1337, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1015-z;   http://link.springer.
com/content/pdf/10.1007/s00362-
018-1015-z.pdf.

Soboleva:2018:ECE

[2264] Olga N. Soboleva, Mikhail I. Epov, and Ekaterina P. Kurochkina. Effective coefficients in the electromagnetic logging problem with log-normal distribution, multiscale conductivity and permittivity. Statistical Papers, 59(4): 1339–1350, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1035-8.

Polunchenko:2018:AML

[2265] Aleksey S. Polunchenko and Andrey Pepelyshev. Analytic moment and Laplace transform formulae for the quasi-stationary distribution of the Shiryaev diffusion on an interval. Statistical Papers, 59(4):1351–1377, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1019-8.

Meintanis:2018:TSI

[2266] Simos G. Meintanis, Joseph Ngatchou-Wandji, and James Allison. Testing for serial independence in vector autoregressive models. Statistical Papers, 59(4):1379–1410, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1039-4.

Vicario:2018:SVB

[2267] Grazia Vicario and Giovanni Pistone. Simulated variogram-based error inspection of manufactured parts. Statistical Papers, 59(4):1411–1423, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1030-0.

Grigoriev:2018:ELO

[2268] Yu. D. Grigoriev, V. B. Melas, and P. V. Shpilev. Excess of locally D-optimal designs for Cobb–Douglas model. Statistical Papers, 59(4): 1425–1439, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1022-0.

Steland:2018:SCE

[2269] Ansgar Steland. Shrinkage for covariance estimation: asymptotics, confidence intervals, bounds and applications in sensor monitoring and finance. Statistical Papers, 59(4): 1441–1462, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1040-y.

Samouylov:2018:ALS

[2270] Konstantin Samouylov and Yuliya Gaidamaka. Analysis of loss systems with overlapping resource requirements. Statistical Papers, 59(4): 1463–1470, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1041-x.

Kargapolova:2018:MCS

[2271] Nina Kargapolova, Elena Khlebnikova, and Vasily Ogorodnikov. Monte Carlo simulation of the joint non-Gaussian periodically correlated timeseries of air temperature and relative humidity. Statistical Papers, 59(4): 1471–1481, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1031-z.

Arboretti:2018:MSS

[2272] Rosa Arboretti, Riccardo Ceccato, Livio Corain, Fabrizio Ronchi, and Luigi Salmaso. Multivariate small sample tests for two-way designs with applications to industrial statistics. Statistical Papers, 59(4):1483–1503, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1032-y.

Belopolskaya:2018:SMF

[2273] Yana Belopolskaya. Stochastic models for forward systems of nonlinear parabolic equations. Statistical Papers, 59(4):1505–1519, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1033-x.

Ogorodnikov:2018:SMA

[2274] Vasily A. Ogorodnikov, Evgeniya G. Kablukova, and Sergei M. Prigarin. Stochastic models of atmospheric clouds structure. Statistical Papers, 59(4):1521–1532, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1036-7.

Basharin:2018:MMA

[2275] Gely Basharin, Valeriy Naumov, and Konstantin Samouylov. On Markovian modelling of arrival processes. Statistical Papers, 59(4):1533–1540, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1042-9.

Tracheva:2018:ESA

[2276] Natalya Tracheva and Sergey Ukhinov. On the evaluation of spatial-angular distributions of polarization characteristics of scattered radiation. Statistical Papers, 59(4):1541–1557, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1034-9.

Dmitriev:2018:NEP

[2277] Yury G. Dmitriev and Gennady M.

Koshkin. Nonparametric estimators of probability characteristics using unbiased prior conditions. Statistical Papers, 59(4):1559–1575, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1044-7.

Rykov:2018:SSP

[2278] V. Rykov. On steady state probabilities of renewable system with Marshal–Olkin failure model. Statistical Papers, 59(4):1577–1588, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1037-6.

Andriyana:2018:QRV

[2279] Y. Andriyana, I. Gijbels, and A. Verhasselt. Quantile regression in varyingcoefficient models: non-crossing quantile curves and heteroscedasticity. Statistical Papers, 59(4):1589–1621, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0847-7.

Parchami:2018:MTF

[2280] Abbas Parchami, S. Mahmoud Taheri, Reinhard Viertl, and Mashaallah Mashinchi. Minimax test for fuzzy hypotheses. Statistical Papers, 59(4): 1623–1648, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0926-4.

Mohammadi:2018:VBU

[2281] Mohammad Mohammadi. A variance bound for unbiased estimation in inverse sampling without replacement. Statistical Papers, 59(4):1649– 1655, December 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0873-0.

Beltran-Beltran:2019:GFT

[2282] J. I. Beltrán-Beltrán and F. J. O’Reilly. On goodness of fit tests for the Poisson, negative binomial and binomial distributions. Statistical Papers, 60 (1):1–18, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0820-5.

Li:2019:PRA

[2283] Yalian Li and Hu Yang. Performance of the restricted almost unbiased type principal components estimators in linear regression model. Statistical Papers, 60(1):19–34, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-016-
0821-4.

DiLascio:2019:CDO

[2284] F. Marta L. Di Lascio and Simone Giannerini. Clustering dependent observations with copula functions. Statistical Papers, 60(1):35–51, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0822-3.

Jarrahiferiz:2019:SPW

[2285] J. Jarrahiferiz, M. Kayid, and S. Izadkhah. Stochastic properties of a weighted frailty model. Statistical Papers, 60(1):53–72, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0826-z.

Kattumannil:2019:SNP

[2286] Sudheesh K. Kattumannil and P. Anisha. A simple non-parametric test for decreasing mean time to failure. Statistical Papers, 60(1):73–87, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0827-y.

Sikov:2019:AFB

[2287] A. Sikov and J. M. Stern. Application of the full Bayesian significance test to model selection under informative sampling. Statistical Papers, 60 (1):89–104, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0828-x.

Rosadi:2019:RSO

[2288] D. Rosadi and P. Filzmoser. Robust second-order least-squares estimation for regression models with autoregressive errors. Statistical Papers, 60 (1):105–122, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0829-9.

Lu:2019:GLS

[2289] Wanbo Lu and Rui Ke. A generalized least squares estimation method for the autoregressive conditional duration model. Statistical Papers, 60 (1):123–146, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0830-3.

Lee:2019:OBT

[2290] Woo Dong Lee, Sang Gil Kang, and Yongku Kim. Objective Bayesian testing for the linear combinations of normal means. Statistical Papers, 60 (1):147–172, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0831-2.

Ciuperca:2019:AGL

[2291] Gabriela Ciuperca. Adaptive group LASSO selection in quantile models. Statistical Papers, 60(1):173– 197, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0832-1.

Hosseinpouri:2019:ASS

[2292] Mahdi Hosseinpouri and Majid Jafari Khaledi. An area-specific stick breaking process for spatial data. Statistical Papers, 60(1):199–221, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0833-0.

Belzunce:2019:ULR

[2293] Félix Belzunce and Carolina Martínez-Riquelme. On the unimodality of the likelihood ratio with applications. Statistical Papers, 60(1):223–237, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0834-z.

Rabier:2019:CSP

[2294] Charles-Elie Rabier, Jean-Marc Azaïs, Jean-Michel Elsen, and Céline Delmas. Chi-square processes for gene mapping in a population with family structure. Statistical Papers, 60 (1):239–271, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0835-y.

Jiang:2019:EEU

[2295] Bo Jiang and Yuqin Sun. On the equality of estimators under a general partitioned linear model with parameter restrictions. Statistical Papers, 60 (1):273–292, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0837-9.

Galea:2019:LID

[2296] Manuel Galea and Patricia Giménez. Local influence diagnostics for the test of mean-variance efficiency and systematic risks in the capital asset pricing model. Statistical Papers, 60(1): 293–312, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0838-8.

Rooch:2019:EML

[2297] Aeneas Rooch, Ieva Zelo, and Roland Fried. Estimation methods for the LRD parameter under a change in the mean. Statistical Papers, 60(1): 313–347, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0839-7.

Ristic:2019:BRD

[2298] Miroslav M. Ristić. Book review: Dieter Rasch and Dieter Schott, Mathematical statistics. Statistical Papers, 60 (1):349–350, February 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01085-z.

Harman:2019:ESI

[2299] Radoslav Harman, Werner G. Müller, and David Woods. Editorial for the special issue mODa12: Advances in model-oriented design and analysis. Statistical Papers, 60(2):351–354, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01101-2;   https://link.springer.
com/content/pdf/10.1007/s00362-
019-01101-2.pdf.

Mielke:2019:ADD

[2300] Tobias Mielke and Vladimir Dragalin.

Adaptive designs for drug combination informed by longitudinal model for the response. Statistical Papers, 60(2):355–371, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-01073-9.

Tarima:2019:APM

[2301] Sergey Tarima and Nancy Flournoy.

Asymptotic properties of maximum likelihood estimators with sample size recalculation. Statistical Papers, 60 (2):373–394, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01095-x.

Wang:2019:RBI

[2302] Yanying Wang, William F. Rosenberger, and Diane Uschner. Randomizationbased inference and the choice of randomization procedures. Statistical Papers, 60(2):395–404, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
01070-y.

Bailey:2019:MPB

[2303] R. A. Bailey and Peter J. Cameron. Multi-part balanced incomplete-block designs. Statistical Papers, 60(2): 405–426, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-01071-x;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-018-01071-x.pdf.

See appendage [2576].

Filipiak:2019:OBD

[2304] Katarzyna Filipiak, Razieh Khodsiani, and Augustyn Markiewicz. Optimality of block designs under the model with the first-order circular autoregression. Statistical Papers, 60(2):427–447, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01077-5;   https://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        018-01077-5.pdf.

Malevich:2019:ODI

[2305] Nadja Malevich and Christine H. Müller. Optimal design of inspection times for interval censoring. Statistical Papers, 60(2):449–464, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01067-7.

Prus:2019:ODM

[2306] Maryna Prus. Optimal designs for minimax-criteria in random coefficient regression models. Statistical Papers, 60(2):465–478, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-01072-w.

Fontana:2019:AML

[2307] Roberto Fontana and Fabio Rapallo.

On the aberrations of mixed level orthogonal arrays with removed runs. Statistical Papers, 60(2):479–493, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
01069-5.

Freise:2019:ODF

[2308] Fritjof Freise and Rainer Schwabe. Optimal designs for K-factor two-level models with first-order interactions on a symmetrically restricted design region. Statistical Papers, 60(2):495–513, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
01063-x.

Radloff:2019:LOD

[2309] Martin Radloff and Rainer Schwabe.

Locally D-optimal designs for a wider class of non-linear models on the kdimensional ball. Statistical Papers, 60(2):515–527, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-01078-4.

Pazman:2019:DMN

[2310] Andrej Pázman. Distribution of the multivariate nonlinear LS estimator under an uncertain input. Statistical Papers, 60(2):529–544, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01081-9.

Pronzato:2019:BDB

[2311] Luc Pronzato, Henry P. Wynn, and Anatoly Zhigljavsky. Bregman divergences based on optimal design criteria and simplicial measures of dispersion. Statistical Papers, 60(2):545–564, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01082-8.

Sagnol:2019:UCB

[2312] Guillaume Sagnol and Edouard Pauwels. An unexpected connection between Bayes A-optimal designs and the group lasso. Statistical Papers, 60(2):565–584, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01062-y.

Yao:2019:OSS

[2313] Yaqiong Yao and HaiYing Wang. Optimal subsampling for softmax regression. Statistical Papers, 60(2):585–599, April 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01068-6.

Inan:2019:MMM

[2314] Gul Inan and Ozlem Ilk. A marginalized multilevel model for bivariate longitudinal binary data. Statistical Papers, 60(3):601–628, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0840-1.

Makinde:2019:CRB

[2315] Olusola Samuel Makinde. Classification rules based on distribution functions of functional depth. Statistical Papers, 60(3):629–640, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0841-0.

Dang:2019:GCM

[2316] Xin Dang, Hailin Sang, and Lauren Weatherall. Gini covariance matrix and its affine equivariant version. Statistical Papers, 60(3):641–666, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0842-z.

Arashi:2019:SIE

[2317] M. Arashi and Mahdi Roozbeh. Some improved estimation strategies in highdimensional semiparametric regression models with application to riboflavin production data. Statistical Papers, 60(3):667–686, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0843-y.

Shelef:2019:GBT

[2318] Amit Shelef and Edna Schechtman.

A Gini-based time series analysis and test for reversibility. Statistical Papers, 60(3):687–716, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0845-9.

Quessy:2019:CNT

[2319] Jean-François Quessy. Consistent nonparametric tests for detecting gradual changes in the marginals and the copula of multivariate time series. Statistical Papers, 60(3):717–746, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0846-8.

Zhou:2019:TPC

[2320] Zhangong Zhou and Linjun Tang. Testing for parametric component of partially linear models with missing covariates. Statistical Papers, 60(3):747–760, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0848-6.

Belaghi:2019:EBP

[2321] R. Arabi Belaghi and M. Noori Asl. Estimation based on progressively type-I hybrid censored data from the Burr XII distribution. Statistical Papers, 60(3):761–803, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0849-5.

Kelkinnama:2019:SAP

[2322] M. Kelkinnama and M. Asadi. Stochastic and ageing properties of coherent systems with dependent identically distributed components. Statistical Papers, 60(3):805–821, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0850-z.

Weiss:2019:TZI

[2323] Christian H. Weiß, Annika Homburg, and Pedro Puig. Testing for zero inflation and overdispersion in INAR(1) models. Statistical Papers, 60(3):823–848, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0851-y.

Zhang:2019:PTP

[2324] Jianjun Zhang, Lei Yang, and Xianyi Wu. Polya tree priors and their estimation with multi-group data. Statistical Papers, 60(3):849–875, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0852-x.

Laha:2019:SRE

[2325] Arnab Kumar Laha, A. C. Pravida Raja, and K. C. Mahesh. SBrobust estimation of mean direction for some new circular distributions. Statistical Papers, 60(3):877–902, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0853-9.

Pokotylo:2019:CPP

[2326] Oleksii Pokotylo and Karl Mosler. Classification with the pot-pot plot. Statistical Papers, 60(3):903–931, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0854-8.

Rajesh:2019:SPC

[2327] G. Rajesh and S. M. Sunoj. Some properties of cumulative Tsallis entropy of order α. Statistical Papers, 60(3):933–943, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0855-7.

Varathan:2019:LLE

[2328] Nagarajah Varathan and Pushpakanthie Wijekoon. Logistic Liu estimator under stochastic linear restrictions. Statistical Papers, 60(3):945–962, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0856-6.

Soni:2019:NTO

[2329] Pooja Soni, Isha Dewan, and Kanchan Jain. Nonparametric tests for ordered quantiles. Statistical Papers, 60(3):963–981, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-016-0859-3.

Xia:2019:CRE

[2330] Ningning Xia and Zhidong Bai. Convergence rate of eigenvector empirical spectral distribution of large Wigner matrices. Statistical Papers, 60(3):983–1015, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0860-x.

Giles:2019:BRD

[2331] David E. Giles. Book review: David A.

Harville: Linear models and the relevant distributions and matrix algebra. Statistical Papers, 60(3):1017–1019, June 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01096-w.

Paulino:2019:AUC

[2332] Sofia Paulino, Manuel Cabral Morais, and Sven Knoth. On ARL-unbiased c-charts for INAR(1) Poisson counts. Statistical Papers, 60(4):1021–1038, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0861-9.

Hu:2019:SPS

[2333] Xuemei Hu and Weiming Yang. Semiparametric small area inference in generalized semi-varying coefficient mixed effects models. Statistical Papers, 60 (4):1039–1058, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0862-8.

deCastro:2019:BIM

[2334] Mário de Castro and Ignacio Vidal. Bayesian inference in measurement error models from objective priors for the bivariate normal distribution. Statistical Papers, 60(4):1059–1078, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0863-7.

Cerqueti:2019:PST

[2335] Roy Cerqueti, Mauro Costantini, Luciano Gutierrez, and Joakim Westerlund. Panel stationary tests against changes in persistence. Statistical Papers, 60(4):1079–1100, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0864-6;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        016-0864-6.pdf.

Shih:2019:BDM

[2336] Jia-Han Shih and Takeshi Emura. Bivariate dependence measures and bivariate competing risks models under the generalized FGM copula. Statistical Papers, 60(4):1101–1118, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-016-
                                        0865-5.

Gorny:2019:BSR

[2337] Julian Górny and Erhard Cramer. From B-spline representations to gamma representations in hybrid censoring. Statistical Papers, 60(4):

1119–1135, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0866-4.

Xu:2019:QRV

[2338] Hong-Xia Xu, Zhen-Long Chen, Jiang-Feng Wang, and Guo-Liang Fan. Quantile regression and variable selection for partially linear model with randomly truncated data. Statistical Papers, 60 (4):1137–1160, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-016-0867-3.

Lv:2019:WQR

[2339] Xiaofeng Lv, Gupeng Zhang, Xinkuo Xu, and Qinghai Li. Weighted quantile regression for censored data with application to export duration data. Statistical Papers, 60(4):1161–1192, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0868-2.

Shen:2019:APL

[2340] Aiting Shen. Asymptotic properties of LS estimators in the errors-in-variables model with MD errors. Statistical Papers, 60(4):1193–1206, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0869-1.

Shen:2019:PML

[2341] Pao sheng Shen and Yi Liu. Pseudo maximum likelihood estimation for the Cox model with doubly truncated data. Statistical Papers, 60(4):1207–1224, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-016-
0870-8.

Xiao:2019:LLS

[2342] Juxia Xiao, Xu Li, and Jianhong Shi. Local linear smoothers using inverse Gaussian regression. Statistical Papers, 60(4):1225–1253, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0871-2.

Yang:2019:RBS

[2343] Jing Yang, Hu Yang, and Fang Lu. Rank-based shrinkage estimation for identification in semiparametric additive models. Statistical Papers, 60 (4):1255–1281, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0874-z.

deMorais:2019:EPM

[2344] Ricardo Saldanha de Morais, Roberto da Costa Quinino, Emilio Suyama, and Linda Lee Ho. Estimators of parameters of a mixture of three multinomial distributions based on simple majority results. Statistical Papers, 60 (4):1283–1316, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0875-y.

Martins:2019:MEU

[2345] A. P. Martins and J. R. Sebastião.

Methods for estimating the upcrossings index: improvements and comparison. Statistical Papers, 60(4): 1317–1347, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0876-x.

Salemi:2019:OOC

[2346] Uoseph Hamdi Salemi, Sadegh Rezaei, and Saralees Nadarajah. A-optimal and D-optimal censoring plans in progressively Type-II right censored order statistics. Statistical Papers, 60 (4):1349–1367, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0877-9.

Rivas-Martinez:2019:TST

[2347] G. I. Rivas-Martínez, M. D. Jiménez-Gamero, and J. L. Moreno-Rebollo. A two-sample test for the error distribution in nonparametric regression based on the characteristic function. Statistical Papers, 60(4):1369–1395, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0878-8.

Li:2019:SAE

[2348] Huapeng Li, Yukun Liu, and Riquan Zhang. Small area estimation under transformed nested-error regression models. Statistical Papers, 60 (4):1397–1418, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0879-7.

Hassler:2019:BRK

[2349] Uwe Hassler. Book review: Katsuto Tanaka (2017): Time series analysis: nonstationary and noninvertible distribution theory, 2nd edition. Statistical Papers, 60(4):1419–1420, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-019-
                                        01114-x.

Tian:2019:CSG

[2350] Shengqi Tian, Dehui Wang, and Shuai Cui. Correction to: A seasonal geometric INAR process based on negative binomial thinning operator. Statistical Papers, 60(4):1421, August 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-019-
                                        01097-9;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        019-01097-9.pdf.

Dai:2019:BLI

[2351] Xiaowen Dai, Libin Jin, Maozai Tian, and Lei Shi. Bayesian local influence for spatial autoregressive models with heteroscedasticity. Statistical Papers, 60 (5):1423–1446, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0880-1.

Pardo:2019:WCS

[2352] María Carmen Pardo and Rosa Alonso. Working correlation structure selection in GEE analysis. Statistical Papers, 60 (5):1447–1467, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0881-0.

Wang:2019:AGL

[2353] Mingqiu Wang and Guo-Liang Tian.

Adaptive group Lasso for highdimensional generalized linear models. Statistical Papers, 60(5):1469–1486, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0882-z.

Demirdjian:2019:KCM

[2354] Levon Demirdjian and Majid Mojirsheibani. Kernel classification with missing data and the choice of smoothing parameters. Statistical Papers, 60 (5):1487–1513, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0883-y.

Awale:2019:TCT

[2355] Manik Awale, N. Balakrishna, and T. V. Ramanathan. Testing the constancy of the thinning parameter in a random coefficient integer autoregressive model. Statistical Papers, 60(5): 1515–1539, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0884-x.

Canterle:2019:VDB

[2356] Diego Ramos Canterle and Fábio Mariano Bayer. Variable dispersion beta regressions with parametric link functions. Statistical Papers, 60(5): 1541–1567, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0885-9.

Zhao:2019:SRT

[2357] Qianqian Zhao and Shengli Zhao. Some results on two-level regular designs with multi block variables containing clear effects. Statistical Papers, 60(5): 1569–1582, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0886-8.

Wong:2019:USF

[2358] Weng Kee Wong, Yue Yin, and Julie Zhou. Using SeDuMi to find various optimal designs for regression models. Statistical Papers, 60(5):1583–1603, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0887-7.

Saulo:2019:BSA

[2359] Helton Saulo, Jeremias Leão, Víctor Leiva, and Robert G. Aykroyd. Birnbaum–Saunders autoregressive conditional duration models applied to high-frequency financial data. Statistical Papers, 60(5):1605–1629, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0888-6.

Pakdaman:2019:SBA

[2360] Zohreh Pakdaman, Jafar Ahmadi, and Mahdi Doostparast. Signature-based approach for stress-strength systems. Statistical Papers, 60(5):1631–1647, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0889-5.

Wang:2019:REE

[2361] Kangning Wang and Lu Lin. Robust and efficient estimator for simultaneous model structure identification and variable selection in generalized partial linear varying coefficient models with longitudinal data. Statistical Papers, 60 (5):1649–1676, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0890-z.

Kosiorowski:2019:DSC

[2362] Daniel Kosiorowski, Jerzy P. Rydlewski, and Małgorzata Snarska. Detecting a structural change in functional time series using local Wilcoxon statistic. Statistical Papers, 60(5): 1677–1698, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0891-y.

Ou:2019:OFP

[2363] Zujun Ou and Hong Qin. Optimal foldover plans of asymmetric factorials with minimum wrap-around L_{2}- discrepancy. Statistical Papers, 60(5): 1699–1716, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0892-x.

Akdeniz:2019:GDB

[2364] Fikri Akdeniz and Mahdi Roozbeh. Generalized difference-based weighted mixed almost unbiased ridge estimator in partially linear models. Statistical Papers, 60(5):1717–1739, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0893-9.

Chen:2019:NQF

[2365] Xiaolin Chen, Xiaojing Chen, and Yi Liu. A note on quantile feature screening via distance correlation. Statistical Papers, 60(5):1741–1762, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0894-8.

Liu:2019:PPI

[2366] Yin Liu, Guo-Liang Tian, Qin Wu, and Man-Lai Tang. Poisson–Poisson item count techniques for surveys with sensitive discrete quantitative data. Statistical Papers, 60(5):1763–1791, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0895-7.

Chen:2019:BRM

[2367] Li-Pang Chen. Book review: Mehryar Mohri, Afshin Rostamizadeh, and Ameet Talwalkar: Foundations of machine learning, second edition. Statistical Papers, 60(5):1793–1795, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01124-9.

Rieger:2019:BRM

[2368] Jonas Rieger. Book review: Mónica Bécue–Bertaut (2019): Textual Data Science with R. Statistical Papers, 60 (5):1797–1798, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01126-7.

Anonymous:2019:CEP

[2369] Anonymous. Correction to: Erroneous pagination in volume 60, issue 2 and issue 3. Statistical Papers, 60 (5):1799–1802, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01131-w;   http:
//link.springer.com/content/pdf/
10.1007/s00362-019-01131-w.pdf.

Yoshida:2019:TSS

[2370] Takuma Yoshida. Two stage smoothing in additive models with missing covariates. Statistical Papers, 60(6): 1803–1826, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0896-6.

Zhan:2019:CCD

[2371] Xiaoping Zhan, Tiefeng Ma, Shuangzhe Liu, and Kunio Shimizu. On circular correlation for data on the torus. Statistical Papers, 60(6):1827–1847, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0897-5.

Norouzirad:2019:PTS

[2372] M. Norouzirad and M. Arashi. Preliminary test and Stein-type shrinkage ridge estimators in robust regression. Statistical Papers, 60(6): 1849–1882, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0899-3.

Martin:2019:DTC

[2373] N. Martín, L. Pardo, and K. Zografos. On divergence tests for composite hypotheses under composite likelihood. Statistical Papers, 60(6): 1883–1919, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0900-1.

Majumder:2019:TDL

[2374] Priyanka Majumder and Murari Mitra. A test for detecting Laplace order dominance and related Bahadur efficiency issues. Statistical Papers, 60(6): 1921–1937, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0901-0.

Crosato:2019:COG

[2375] Lisa Crosato and Luigi Grossi. Correcting outliers in GARCH models: a weighted forward approach. Statistical Papers, 60(6):1939–1970, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0903-y.

Sunoj:2019:SPC

[2376] S. M. Sunoj and N. Vipin. Some properties of conditional partial moments in the context of stochastic modelling. Statistical Papers, 60(6): 1971–1999, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0904-x.

Noughabi:2019:BQR

[2377] M. Shafaei Noughabi and M. Kayid.

Bivariate quantile residual life: a characterization theorem and statistical properties. Statistical Papers, 60(6): 2001–2012, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0905-9.

Fenoy:2019:GVM

[2378] Mar Fenoy, Pilar Ibarrola, and Juan B.

Seoane-Sepúlveda. Generalized p value for multivariate Gaussian stochastic processes in continuous time. Statistical Papers, 60(6):2013–2030, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0907-7.

Kumar:2019:EDT

[2379] Nirpeksh Kumar. Exact distributions of tests of outliers for exponential samples. Statistical Papers, 60(6): 2031–2061, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0908-6.

Jiang:2019:REE

[2380] Yunlu Jiang, Guo-Liang Tian, and Yu Fei. A robust and efficient estimation method for partially nonlinear models via a new MM algorithm. Statistical Papers, 60(6):2063–2085, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0909-5.

Wang:2019:APE

[2381] Xuejun Wang, Meimei Ge, and Yi Wu. The asymptotic properties of the estimators in a semiparametric regression model. Statistical Papers, 60(6): 2087–2108, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0910-z.

Aylar:2019:LTL

[2382] Emre Aylar, Stephan Smeekes, and Joakim Westerlund. Lag truncation and the local asymptotic distribution of the ADF test for a unit root. Statistical Papers, 60(6):2109–2118, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-

0911-y;   http://link.springer.
com/content/pdf/10.1007/s00362-
017-0911-y.pdf.

Barreto-Souza:2019:MPI

[2383] Wagner Barreto-Souza. Mixed Poisson INAR(1) processes. Statistical Papers, 60(6):2119–2139, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0912-x.

Zamanzade:2019:EBT

[2384] Ehsan Zamanzade. EDF-based tests of exponentiality in pair ranked set sampling. Statistical Papers, 60(6): 2141–2159, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0913-9.

Jokiel-Rokita:2019:MDE

[2385] Alicja Jokiel-Rokita and Rafał Topolnicki. Minimum distance estimation of the binormal ROC curve. Statistical Papers, 60(6):2161–2183, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0915-7;   http://link.springer.
com/content/pdf/10.1007/s00362-
017-0915-7.pdf.

Kohansal:2019:ERM

[2386] Akram Kohansal. On estimation of reliability in a multicomponent stressstrength model for a Kumaraswamy distribution based on progressively censored sample. Statistical Papers, 60(6):

2185–2224, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0916-6.

Ghosh:2019:BED

[2387] Amit Ghosh and Chanchal Kundu. Bivariate extension of (dynamic) cumulative residual and past inaccuracy measures. Statistical Papers, 60(6): 2225–2252, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0917-5.

Shen:2019:LSE

[2388] Guangjun Shen and Qian Yu. Least squares estimator for Ornstein–Uhlenbeck processes driven by fractional Lévy processes from discrete observations. Statistical Papers, 60(6):2253–2271, December 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0918-4.

Hai:2020:EVC

[2389] Tran Hoang Hai. Estimation of volatility causality in structural autoregressions with heteroskedasticity using independent component analysis. Statistical Papers, 61(1):1–16, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0919-3.

Henze:2020:TNB

[2390] Norbert Henze and Stefan Koch. On a test of normality based on the empirical moment generating function. Statistical Papers, 61(1):17–29, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0923-7.

Borssoi:2020:ELM

[2391] Joelmir A. Borssoi, Gilberto A. Paula, and Manuel Galea. Elliptical linear mixed models with a covariate subject to measurement error. Statistical Papers, 61(1):31–69, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0921-9.

LaMotte:2020:FKP

[2392] Lynn Roy LaMotte. Following K.

Pearson to test the general linear hypothesis. Statistical Papers, 61 (1):71–83, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0924-6.

Motarjem:2020:GSM

[2393] K. Motarjem, M. Mohammadzadeh, and A. Abyar. Geostatistical survival model with Gaussian random effect. Statistical Papers, 61(1):85–107, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0922-8.

Li:2020:RCP

[2394] Qiang Li and Liming Wang. Robust change point detection method via adaptive LAD–LASSO. Statistical Papers, 61(1):109–121, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0927-3.

Meintanis:2020:GFT

[2395] Simos G. Meintanis, Bojana Milosević, and Marko Obradović. Goodnessof-fit tests in conditional duration models. Statistical Papers, 61(1): 123–140, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0930-8.

Xu:2020:BAQ

[2396] Qifa Xu, Chao Cai, Cuixia Jiang, Fang Sun, and Xue Huang. Block average quantile regression for massive dataset. Statistical Papers, 61 (1):141–165, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0932-6.

Araujo:2020:IHL

[2397] Mariana C. Araújo, Audrey H. M. A. Cysneiros, and Lourdes C. Montenegro. Improved heteroskedasticity likelihood ratio tests in symmetric nonlinear regression models. Statistical Papers, 61 (1):167–188, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0933-5.

Zhang:2020:CML

[2398] Jin Zhang. Consistency of MLE, LSE and M-estimation under mild conditions. Statistical Papers, 61(1): 189–199, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0928-2.

Xu:2020:TSH

[2399] Feng Xu and Zekai He. Testing slope homogeneity in panel data models with a multifactor error structure. Statistical Papers, 61(1):201–224, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0929-1.

Lu:2020:MFC

[2400] Jun Lu and Lu Lin. Model-free conditional screening via conditional distance correlation. Statistical Papers, 61 (1):225–244, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0931-7.

Wang:2020:PAP

[2401] Xinyang Wang, Dehui Wang, and Haixiang Zhang. Poisson autoregressive process modeling via the penalized conditional maximum likelihood procedure. Statistical Papers, 61(1): 245–260, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0938-0.

Al-Mutairi:2020:CIQ

[2402] Jazaa S. Al-Mutairi and Mohammad Z. Raqab. Confidence intervals for quantiles based on samples of random sizes. Statistical Papers, 61(1): 261–277, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0935-3.

Hladik:2020:ERB

[2403] Milan Hladík, Michal Cerný, and Jaromír Antoch. EIV regression with bounded errors in data: total ‘least squares’ with Chebyshev norm. Statistical Papers, 61(1):279–301, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0939-z.

Liu:2020:SRC

[2404] Xiaohui Liu, Shihua Luo, and Yijun Zuo. Some results on the computing of Tukey’s halfspace median. Statistical Papers, 61(1):303–316, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0941-5.

Wiedermann:2020:LLM

[2405] Wolfgang Wiedermann and Alexander von Eye. Log-linear models to evaluate direction of effect in binary variables. Statistical Papers, 61(1): 317–346, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0936-2.

Calcagni:2020:MMA

[2406] Antonio Calcagnì, Luigi Lombardi, Lorenzo Avanzi, and Eduardo Pascali. Multiple mediation analysis for interval-valued data. Statistical Papers, 61(1):347–369, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0940-6.

vonRosen:2020:NMO

[2407] Tatjana von Rosen, Dietrich von Rosen, and Julia Volaufova. A new method for obtaining explicit estimators in unbalanced mixed linear models. Statistical Papers, 61(1): 371–383, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0937-1;   http:
//link.springer.com/content/pdf/
10.1007/s00362-017-0937-1.pdf.

Haslett:2020:SPL

[2408] S. J. Haslett, X. Q. Liu, A. Markiewicz, and S. Puntanen. Some properties of linear sufficiency and the BLUPs in the linear mixed model. Statistical Papers, 61(1):385–401, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0943-3.

Kulkarni:2020:EPC

[2409] M. G. Kulkarni and M. B. Rajarshi.

Estimation of parameters of component lifetime distribution in a coherent system. Statistical Papers, 61 (1):403–421, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0945-1.

Ling:2020:NEF

[2410] Nengxiang Ling, Germán Aneiros, and Philippe Vieu. k NN estimation in functional partial linear modeling. Statistical Papers, 61(1):423–444, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0946-0.

Budhiraja:2020:PIE

[2411] Sonal Budhiraja and Biswabrata Pradhan. Point and interval estimation under progressive type-I interval censoring with random removal. Statistical Papers, 61(1):445–477, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0948-y.

Emura:2020:PLI

[2412] Takeshi Emura and Chi-Hung Pan. Parametric likelihood inference and goodness-of-fit for dependently lefttruncated data, a copula-based approach. Statistical Papers, 61(1): 479–501, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0947-z.

Ranjbar:2020:CFF

[2413] Vahid Ranjbar and Gholamreza Hesamian. Copula function for fuzzy random variables: applications in measuring association between two fuzzy random variables. Statistical Papers, 61 (1):503–522, February 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0944-2.

Zhao:2020:PSD

[2414] Xiaobing Zhao and Xian Zhou. Partial sufficient dimension reduction on additive rates model for recurrent event data with high-dimensional covariates. Statistical Papers, 61(2):523–541, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0949-x.

Kim:2020:MPB

[2415] Yongku Kim, Woo Dong Lee, and Sang Gil Kang. A matching prior based on the modified profile likelihood for the common mean in multiple log-normal distributions. Statistical Papers, 61(2):543–573, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0950-4.

Jia:2020:IWP

[2416] Xiang Jia, Saralees Nadarajah, and Bo Guo. Inference on q-Weibull parameters. Statistical Papers, 61(2):575–593, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0951-3.

Ieva:2020:CWO

[2417] Francesca Ieva and Anna Maria Paganoni. Component-wise outlier detection methods for robustifying multivariate functional samples. Statistical Papers, 61(2):595–614, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0953-1.

Luo:2020:NKE

[2418] Zhongde Luo. Nonparametric kernel estimation of CVaR under αmixing sequences. Statistical Papers, 61(2):615–643, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0952-2.

Lu:2020:BIA

[2419] Zeng-Hua Lu. Bahadur intercept with applications to one-sided testing. Statistical Papers, 61(2):645–658, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0955-z.

Li:2020:MTE

[2420] Qi Li and Fukang Zhu. Mean targeting estimator for the integervalued GARCH(1, 401) model. Statistical Papers, 61(2):659–679, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-017-
0958-9.

Lu:2020:DUH

[2421] Yang Lu. The distribution of unobserved heterogeneity in competing risks models. Statistical Papers, 61 (2):681–696, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0956-y.

Liu:2020:ELB

[2422] Tianqing Liu and Xiaohui Yuan. Empirical likelihood-based weighted rank regression with missing covariates. Statistical Papers, 61(2):697–725, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0957-x.

Guo:2020:CLH

[2423] Bing Guo, Xue-Ping Chen, and Min-Qian Liu. Construction of Latin hypercube designs with nested and sliced structures. Statistical Papers, 61(2):727–740, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0959-8.

Leonov:2020:CES

[2424] Sergei Leonov and Bahjat Qaqish. Correlated endpoints: simulation, modeling, and extreme correlations. Statistical Papers, 61(2):741–766, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-017-
                                        0960-2.

Dastbaravarde:2020:EBJ

[2425] Ali Dastbaravarde and Ehsan Zamanzade. On estimation of P (X > Y) based on judgement post stratification. Statistical Papers, 61(2):767–785, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0962-0.

Mazza:2020:MMC

[2426] Angelo Mazza and Antonio Punzo. Mixtures of multivariate contaminated normal regression models. Statistical Papers, 61(2):787–822, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0964-y.

Elsawah:2020:NFD

[2427] A. M. Elsawah and Kai-Tai Fang. New foundations for designing U-optimal follow-up experiments with flexible levels. Statistical Papers, 61(2):823–849, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-017-
                                        0963-z.

Mahdizadeh:2020:ESD

[2428] M. Mahdizadeh and Ehsan Zamanzade. Estimation of a symmetric distribution function in multistage ranked set sampling. Statistical Papers, 61 (2):851–867, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0965-x.

Yuzbasi:2020:RTP

[2429] Bahadir Yüzbasi, S. Ejaz Ahmed, and Dursun Aydin. Ridge-type pretest and shrinkage estimations in partially linear models. Statistical Papers, 61 (2):869–898, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0967-8.

Wu:2020:CPC

[2430] Yi Wu, Xuejun Wang, and Narayanaswamy Balakrishnan. On the consistency of the P-C estimator in a nonparametric regression model. Statistical Papers, 61(2):899–915, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0966-9.

Maronna:2020:BRN

[2431] Ricardo Maronna. Book review: Norman Matloff (2017): Statistical regression and classification: from linear models to machine learning. Statistical Papers, 61(2):917–918, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01160-w.

Kramer:2020:BRD

[2432] Walter Krämer. Book review: Dieter Rasch, Rob Verdooren and Jürgen Pilz: Applied statistics: theory and problem solutions with R. Statistical Papers, 61(2):919–920, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01164-6;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-020-01164-6.pdf.

Pawlowsky-Glahn:2020:BRP

[2433] Vera Pawlowsky-Glahn. Book review: Peter Filzmoser, Karel Hron, Matthias Templ: Applied compositional data analysis, with worked examples in R. Statistical Papers, 61 (2):921–922, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01163-7.

Lee:2020:NGC

[2434] Hyunju Lee and Ji Hwan Cha. A new general class of discrete bivariate distributions constructed by using the likelihood ratio. Statistical Papers, 61(3):923–944, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0969-6.

Dorre:2020:BEL

[2435] Achim Dörre. Bayesian estimation of a lifetime distribution under double truncation caused by time-restricted data collection. Statistical Papers, 61(3):945–965, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0968-7.

Wang:2020:EPE

[2436] Kangning Wang and Xiaofei Sun. Efficient parameter estimation and variable selection in partial linear varying coefficient quantile regression model with longitudinal data. Statistical Papers, 61(3):967–995, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0970-0.

Amin:2020:PSR

[2437] Muhammad Amin, Muhammad Qasim, Muhammad Amanullah, and Saima Afzal. Performance of some ridge estimators for the gamma regression model. Statistical Papers, 61 (3):997–1026, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-017-0971-z.

Bieniek:2020:SLB

[2438] Mariusz Bieniek and Agnieszka Goroncy.

Sharp lower bounds on expectations of gOS based on DGFR distributions. Statistical Papers, 61(3):1027–1042, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0972-y;   http://link.springer.
com/content/pdf/10.1007/s00362-
017-0972-y.pdf.

Wang:2020:LFT

[2439] Lihong Wang. Lack of fit test for long memory regression models. Statistical Papers, 61(3):1043–1067, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-017-
0974-9.

Tsukuda:2020:CSA

[2440] Koji Tsukuda and Hiroshi Kurata. Covariance structure associated with an equality between two general ridge estimators. Statistical Papers, 61 (3):1069–1084, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0975-8.

Kopcova:2020:ISE

[2441] Veronika Kopcová and Ivan Zezula. On intraclass structure estimation in the growth curve model. Statistical Papers, 61(3):1085–1106, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-017-0973-x.

Yu:2020:SIP

[2442] Ping Yu, Jiang Du, and Zhongzhan Zhang. Single-index partially functional linear regression model. Statistical Papers, 61(3):1107–1123, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        0980-6.

Xie:2020:VSS

[2443] Tianfa Xie, Ruiyuan Cao, and Jiang Du. Variable selection for spatial autoregressive models with a diverging number of parameters. Statistical Papers, 61(3):1125–1145, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        0984-2.

Wang:2020:CMC

[2444] Xuejun Wang, Yi Wu, Shuhe Hu, and Nengxiang Ling. Complete moment convergence for negatively orthant dependent random variables and its applications in statistical models. Statistical Papers, 61(3):1147–1180, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0983-3.

Derrar:2020:ERF

[2445] Saliha Derrar, Ali Laksaci, and Elias Ould Saïd. M-estimation of the regression function under random left truncation and functional time series model. Statistical Papers, 61 (3):1181–1202, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0979-z.

Koelbl:2020:NAE

[2446] Lukas Koelbl and Manfred Deistler.

A new approach for estimating VAR systems in the mixed-frequency case. Statistical Papers, 61(3):1203–1212, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0985-1;   http://link.springer.
com/content/pdf/10.1007/s00362-
018-0985-1.pdf.

Gutierrez:2020:LMS

[2447] Luis Gutiérrez, Ramsés H. Mena, and Carlos Díaz-Avalos. Linear models for statistical shape analysis based on parametrized closed curves. Statistical Papers, 61(3):1213–1229, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0986-0.

Li:2020:NLB

[2448] Hongyi Li and Hong Qin. New lower bounds of four-level and two-level designs via two transformations. Statistical Papers, 61(3):1231–1243, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0987-z.

Luo:2020:VSH

[2449] Shan Luo. Variable selection in high-dimensional sparse multiresponse linear regression models. Statistical Papers, 61(3):1245–1267, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0989-x.

Wei:2020:SRE

[2450] Chuanhua Wei and Jin Yang. Stochastic restricted estimation in partially linear additive errors-in-variables models. Statistical Papers, 61(3):1269–1279, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0978-0.

Galarza:2020:QRN

[2451] Christian E. Galarza, Luis M. Castro, Francisco Louzada, and Victor H. Lachos. Quantile regression for nonlinear mixed effects models: a likelihood based perspective. Statistical Papers, 61(3):1281–1307, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0988-y.

Alemohammad:2020:MSA

[2452] N. Alemohammad, S. Rezakhah, and S. H. Alizadeh. Markov switching asymmetric GARCH model: stability and forecasting. Statistical Papers, 61 (3):1309–1333, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0992-2.

Zhou:2020:CSC

[2453] Sanyu Zhou, Defa Wang, and Jingjing Zhu. Construction of simultaneous confidence bands for a percentile hyperplane with predictor variables constrained in an ellipsoidal region. Statistical Papers, 61(3):1335–1346, June 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0990-4.

Sofronov:2020:ESI

[2454] Georgy Sofronov, Martin Wendler, and Volkmar Liebscher. Editorial for the special issue: Change point detection. Statistical Papers, 61(4): 1347–1349, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01199-9;   http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-020-01199-9.pdf.

Hlavka:2020:CPM

[2455] Zdenek Hlávka, Marie Husková, and Simos G. Meintanis. Change-point methods for multivariate time-series: paired vectorial observations. Statistical Papers, 61(4):1351–1383, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01175-3.

Maciak:2020:CDN

[2456] Matús Maciak, Michal Pesta, and Barbora Pestová. Changepoint in dependent and non-stationary panels. Statistical Papers, 61(4):1385–1407, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01180-6.

Rackauskas:2020:CPH

[2457] Alfredas Rackauskas and Martin Wendler. Convergence of U-processes in Hölder spaces with application to robust detection of a changed segment. Statistical Papers, 61(4):1409–1435, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01161-9;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        020-01161-9.pdf.

Mohr:2020:ECP

[2458] Maria Mohr and Leonie Selk. Estimating change points in nonparametric time series regression models. Statistical Papers, 61(4):1437–1463, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01162-8;   http://link.springer.
com/content/pdf/10.1007/s00362-
020-01162-8.pdf.

Yang:2020:CPD

[2459] Qing Yang, Yu-Ning Li, and Yi Zhang.

Change point detection for nonparametric regression under strongly mixing process. Statistical Papers, 61(4): 1465–1506, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01196-y.

Ma:2020:MCP

[2460] Lijing Ma, Andrew J. Grant, and Georgy Sofronov. Multiple change point detection and validation in autoregressive time series data. Statistical Papers, 61(4):1507–1528, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01198-w;   http://link.springer.
com/content/pdf/10.1007/s00362-
020-01198-w.pdf.

Gapeev:2020:PSS

[2461] Pavel V. Gapeev. On the problems of sequential statistical inference for Wiener processes with delayed observations. Statistical Papers, 61(4): 1529–1544, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01178-0;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-020-01178-0.pdf.

Szajowski:2020:RDM

[2462] Krzysztof J. Szajowski. Rationalization of detection of the multiple disorders. Statistical Papers, 61(4): 1545–1563, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01168-2;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-020-01168-2.pdf.

Bandt:2020:OPT

[2463] Christoph Bandt. Order patterns, their variation and change points in financial time series and Brownian motion. Statistical Papers, 61(4):1565–1588, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01171-7;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        020-01171-7.pdf.

Chu:2020:CSN

[2464] Yi Chu and Lu Lin. Conditional SIRS for nonparametric and semiparametric models by marginal empirical likelihood. Statistical Papers, 61(4): 1589–1606, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-0993-1.

Deng:2020:EIA

[2465] Xin Deng and Xuejun Wang. An exponential inequality and its application to M estimators in multiple linear models. Statistical Papers, 61(4): 1607–1627, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0994-0.

Matsuura:2020:OPP

[2466] Shun Matsuura and Thaddeus Tarpey.

Optimal principal points estimators of multivariate distributions of locationscale and location-scale-rotation families. Statistical Papers, 61(4): 1629–1643, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0995-z.

Aria:2020:DRT

[2467] Massimo Aria, Antonio D’Ambrosio, Carmela Iorio, Roberta Siciliano, and Valentina Cozza. Dynamic recursive tree-based partitioning for malignant melanoma identification in skin lesion dermoscopic images. Statistical Papers, 61(4):1645–1661, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0997-x.

Xi:2020:SCP

[2468] Mengmei Xi, Rui Wang, Zhaoyang Cheng, and Xuejun Wang. Some convergence properties for partial sums of widely orthant dependent random variables and their statistical applications. Statistical Papers, 61(4): 1663–1684, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0996-y.

Alvarez-Andrade:2020:CTR

[2469] Sergio Alvarez-Andrade and Salim Bouzebda. Cramér’s type results for some bootstrapped U-statistics. Statistical Papers, 61(4):1685–1699, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
0999-8.

Tanaka:2020:SOM

[2470] Toyoto Tanaka, Yoshihiro Hirose, and Fumiyasu Komaki. Second-order matching prior family parametrized by sample size and matching probability. Statistical Papers, 61(4):1701–1717, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1001-5.

Zhao:2020:MAE

[2471] Shangwei Zhao, Jun Liao, and Dalei Yu. Model averaging estimator in ridge regression and its large sample properties. Statistical Papers, 61(4): 1719–1739, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1002-4.

Han:2020:RCM

[2472] Lengyi Han, W. John Braun, and Jason Loeppky. Random coefficient minification processes. Statistical Papers, 61 (4):1741–1762, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1000-6.

Hu:2020:NLB

[2473] Liuping Hu, Kashinath Chatterjee, Jiaqi Liu, and Zujun Ou. New lower bound for Lee discrepancy of asymmetrical factorials. Statistical Papers, 61 (4):1763–1772, August 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-0998-9.

Christou:2020:RDR

[2474] Eliana Christou. Robust dimension reduction using sliced inverse median regression. Statistical Papers, 61(5): 1799–1818, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1007-z.

Popovic:2020:BIV

[2475] Predrag M. Popović and Hassan S.

Bakouch. A bivariate integer-valued bilinear autoregressive model with random coefficients. Statistical Papers, 61 (5):1819–1840, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1005-1.

Wang:2020:SIV

[2476] Mingqiu Wang, Peixin Zhao, and Xiaoning Kang. Structure identification for varying coefficient models with measurement errors based on kernel smoothing. Statistical Papers, 61(5): 1841–1857, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1009-x.

Meskaldji:2020:CER

[2477] Djalel-Eddine Meskaldji, Dimitri Van De Ville, Jean-Philippe Thiran, and Stephan Morgenthaler. A comprehensive error rate for multiple testing. Statistical Papers, 61(5):1859–1874, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1008-y.

He:2020:MLE

[2478] Xiaofang He, Wangxue Chen, and Wenshu Qian. Maximum likelihood estimators of the parameters of the log-logistic distribution. Statistical Papers, 61(5):1875–1892, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1011-3.

Miyazaki:2020:NNL

[2479] Ryo Miyazaki and Hidetoshi Murakami. The non-null limiting distribution of the generalized Baumgartner statistic based on the Fourier series approximation. Statistical Papers, 61 (5):1893–1909, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1012-2.

Yang:2020:REE

[2480] Hu Yang, Ning Li, and Jing Yang.

A robust and efficient estimation and variable selection method for partially linear models with large-dimensional covariates. Statistical Papers, 61(5): 1911–1937, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1013-1.

Lamboni:2020:UQM

[2481] Matieyendou Lamboni. Uncertainty quantification: a minimum variance unbiased (joint) estimator of the nonnormalized Sobol’ indices. Statistical Papers, 61(5):1939–1970, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1010-4.

Ahmadi:2020:CPR

[2482] Jafar Ahmadi and H. N. Nagaraja.

Conditional properties of a random sample given an order statistic. Statistical Papers, 61(5):1971–1996, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1016-y.

He:2020:ODT

[2483] Lei He and Rong-Xian Yue. Roptimal designs for trigonometric regression models. Statistical Papers, 61 (5):1997–2013, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-1017-x.

Jiang:2020:RMV

[2484] Fen Jiang, Junmei Zhou, and Jin Zhang. Restricted minimum volume confidence region for Pareto distribution. Statistical Papers, 61(5): 2015–2029, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-1018-9.

Zhang:2020:MRM

[2485] Jun Zhang, Junpeng Zhu, Yan Zhou, Xia Cui, and Tao Lu. Multiplicative regression models with distortion measurement errors. Statistical Papers, 61 (5):2031–2057, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-1020-2.

Toker:2020:ITP

[2486] Selma Toker. Investigating the two parameter analysis of Lipovetsky for simultaneous systems. Statistical Papers, 61(5):2059–2089, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1021-1.

Kamgar:2020:EWN

[2487] Saeideh Kamgar, Florian Meinfelder, Ralf Münnich, and Hamidreza Navvabpour. Estimation within the new integrated system of household surveys in Germany. Statistical Papers, 61(5):2091–2117, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1023-z.

Guo:2020:SDA

[2488] Xiao Guo and Hai Zhang. Sparse directed acyclic graphs incorporating the covariates. Statistical Papers, 61(5): 2119–2148, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1027-8.

Worz:2020:TUM

[2489] Sascha Wörz and Heinz Bernhardt. Towards an uniformly most powerful binomial test. Statistical Papers, 61(5): 2149–2156, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1029-6.

Finos:2020:ZIP

[2490] Livio Finos and Fortunato Pesarin.

On zero-inflated permutation testing and some related problems. Statistical Papers, 61(5):2157–2174, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1025-x.

Herwartz:2020:NTI

[2491] Helmut Herwartz and Simone Maxand.

Nonparametric tests for independence: a review and comparative simulation study with an application to malnutrition data in India. Statistical Papers, 61 (5):2175–2201, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-1026-9.

Chatterjee:2020:CSS

[2492] K. Chatterjee, C. Koukouvinos, and K. Mylona. Construction of supersaturated split-plot designs. Statistical Papers, 61(5):2203–2219, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1028-7.

Schomaker:2020:WWU

[2493] Michael Schomaker and Christian Heumann. When and when not to use optimal model averaging. Statistical Papers, 61(5):2221–2240, October 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1048-3.

Liu:2020:DRA

[2494] Tianqing Liu and Xiaohui Yuan. Doubly robust augmented-estimating-equations estimation with nonignorable nonresponse data. Statistical Papers, 61(6):2241–2270, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        1046-5.

Zhang:2020:LTS

[2495] Lingxiang Zhang. Linearity tests and stochastic trend under the STAR framework. Statistical Papers, 61(6): 2271–2282, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1047-4.

Park:2020:SCC

[2496] Heewon Park and Sadanori Konishi.

Sparse common component analysis for multiple high-dimensional datasets via noncentered principal component analysis. Statistical Papers, 61(6): 2283–2311, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1045-6.

Zhang:2020:APQ

[2497] Caiya Zhang, Kaihong Xu, and Lianfen Qian. Asymptotic properties of the QMLE in a log-linear RealGARCH model with Gaussian errors. Statistical Papers, 61(6):2313–2330, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1051-8.

Ding:2020:CWE

[2498] Liwang Ding, Ping Chen, and Yongming Li. Consistency for wavelet estimator in nonparametric regression model with extended negatively dependent samples. Statistical Papers, 61(6): 2331–2349, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1050-9.

He:2020:PEL

[2499] Bang-Qiang He, Xing-Jian Hong, and Guo-Liang Fan. Penalized empirical likelihood for partially linear errorsin-variables panel data models with fixed effects. Statistical Papers, 61(6): 2351–2381, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1049-2.

Hassler:2020:NSQ

[2500] Uwe Hassler. Note on sample quantiles for ordinal data. Statistical Papers, 61(6):2383–2391, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1054-5.

Lafuente-Rego:2020:RFC

[2501] B. Lafuente-Rego, P. D’Urso, and J. A. Vilar. Robust fuzzy clustering based on quantile autocovariances. Statistical Papers, 61(6):2393–2448, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1053-6.

Livingston:2020:BIS

[2502] Glen Livingston, Jr. and Darfiana Nur. Bayesian inference of smooth transition autoregressive (STAR)(k)-GARCH(l, 40m) models. Statistical Papers, 61(6): 2449–2482, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1056-3.

Feng:2020:SIL

[2503] Zhenghui Feng, Jun Zhang, and Qian Chen. Statistical inference for linear regression models with additive distortion measurement errors. Statistical Papers, 61(6):2483–2509, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1057-2.

Patra:2020:EFS

[2504] Lakshmi Kanta Patra, Suchandan Kayal, and Somesh Kumar. Estimating a function of scale parameter of an exponential population with unknown location under general loss function. Statistical Papers, 61(6):2511–2527, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
1052-7.

Amir:2020:MPT

[2505] Abdoulkarim Ilmi Amir and Yacouba Boubacar Maïnassara. Multivariate portmanteau tests for weak multiplicative seasonal VARMA models. Statistical Papers, 61(6):2529– 2560, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1055-4.

Tian:2020:SGI

[2506] Shengqi Tian, Dehui Wang, and Shuai Cui. A seasonal geometric INAR process based on negative binomial thinning operator. Statistical Papers, 61(6): 2561–2581, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1060-7.

Sun:2020:PSS

[2507] Ruili Sun, Tiefeng Ma, and Shuangzhe Liu. Portfolio selection: shrinking the time-varying inverse conditional covariance matrix. Statistical Papers, 61(6): 2583–2604, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1059-0.

Chukhrova:2020:RVN

[2508] Nataliya Chukhrova and Arne Johannssen. Randomized versus nonrandomized hypergeometric hypothesis testing with crisp and fuzzy hypotheses. Statistical Papers, 61(6): 2605–2641, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-1058-1.

Wang:2020:FMM

[2509] Wan-Lun Wang, Ahad Jamalizadeh, and Tsung-I Lin. Finite mixtures of multivariate scale-shape mixtures of skew-normal distributions. Statistical Papers, 61(6):2643–2670, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
01061-z.

Shen:2020:APE

[2510] Aiting Shen, Huiling Tao, and Xuejun Wang. The asymptotic properties for the estimators of the survival function and failure rate function based on WOD samples. Statistical Papers, 61(6):2671–2684, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
01064-w.

Zou:2020:CIS

[2511] Yu-Ye Zou and Han-Ying Liang. CLT for integrated square error of density estimators with censoring indicators missing at random. Statistical Papers, 61(6):2685–2714, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-018-
01065-9.

Cao:2020:FBE

[2512] Ruiyuan Cao, Jiang Du, Jianjun Zhou, and Tianfa Xie. FPCA-based estimation for generalized functional partially linear models. Statistical Papers, 61(6): 2715–2735, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-018-01066-8.

Duarte:2020:AEO

[2513] Belmiro P. M. Duarte and Guillaume Sagnol. Approximate and exact optimal designs for 2^{k} factorial experiments for generalized linear models via second order cone programming. Statistical Papers, 61(6):2737–2767, December 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-018-
                                        01075-7.

Muller:2021:E

[2514] Werner G. Müller, Carsten Jentsch, and Ulrike Schneider. Editorial. Statistical Papers, 62(1): 1–2, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01225-4.

Chen:2021:STM

[2515] Chyong-Mei Chen, Pao sheng Shen, and Yi Liu. On semiparametric transformation model with LTRC data: pseudo likelihood approach. Statistical Papers, 62(1):3–30, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-018-01080-
                                        w.

Majumder:2021:DTC

[2516] Priyanka Majumder and Murari Mitra. Detecting trend change in hazard functions — an L-statistic approach. Statistical Papers, 62(1): 31–52, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-018-01074-8.

Li:2021:EGA

[2517] Chengbo Li and Yong Zhou. The estimation for the general additivemultiplicative hazard model using the length-biased survival data. Statistical Papers, 62(1):53–74, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-018-01079-
3.

Yeh:2021:POR

[2518] Chi-Kuang Yeh and Julie Zhou. Properties of optimal regression designs under the second-order least squares estimator. Statistical Papers, 62 (1):75–92, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-018-01076-6.

Li:2021:AEC

[2519] Feng Li, Lu Lin, and Sanying Feng.

An adaptive estimation for covariateadjusted nonparametric regression model. Statistical Papers, 62(1): 93–115, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01084-0.

Dahl:2021:ANM

[2520] Christian M. Dahl and Emma M.

Iglesias. Asymptotic normality of the MLE in the level-effect ARCH model. Statistical Papers, 62(1): 117–135, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01086-y.

Guerra:2021:OSP

[2521] Manuel Guerra, Cláudia Nunes, and Carlos Oliveira. The optimal stopping problem revisited. Statistical Papers, 62(1):137–169, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-019-01088-
w.

Lennartz:2021:MAC

[2522] J. M. Lennartz, S. Bedbur, and U. Kamps. Minimum area confidence regions and their coverage probabilities for type-II censored exponential data. Statistical Papers, 62(1): 171–191, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01087-x.

Wang:2021:DRF

[2523] Guochang Wang, Beiting Liang, and Baojian Xie. Dimension reduction for functional regression with a binary response. Statistical Papers, 62 (1):193–208, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01083-1.

Patra:2021:MEC

[2524] Lakshmi Kanta Patra, Suchandan Kayal, and Somesh Kumar. Minimax estimation of the common variance and precision of two normal populations with ordered restricted means. Statistical Papers, 62(1): 209–233, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01090-2.

Bouzas:2021:FCT

[2525] Paula R. Bouzas, Nuria Ruiz-Fuentes, and Juan Eloy Ruiz-Castro. Forecasting counting and time statistics of compound Cox processes: a focus on intensity phase type process, deletions and simultaneous events. Statistical Papers, 62(1): 235–265, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01092-0.

Elsawah:2021:NRD

[2526] A. M. Elsawah, Kai-Tai Fang, and Xiao Ke. New recommended designs for screening either qualitative or quantitative factors. Statistical Papers, 62 (1):267–307, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01089-9.

Kohansal:2021:BCE

[2527] Akram Kohansal and Shirin Shoaee.

Bayesian and classical estimation of reliability in a multicomponent stress-strength model under adaptive hybrid progressive censored data. Statistical Papers, 62(1): 309–359, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01094-y.

Wang:2021:SIR

[2528] Huiwen Wang, Zhichao Wang, and Shanshan Wang. Sliced inverse regression method for multivariate compositional data modeling. Statistical Papers, 62(1):361–393, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-019-01093-
                                        z.

Qian:2021:PEP

[2529] Wenshu Qian, Wangxue Chen, and Xiaofang He. Parameter estimation for the Pareto distribution based on ranked set sampling. Statistical Papers, 62(1): 395–417, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01102-1.

Khan:2021:EIE

[2530] Ruhul Ali Khan and Murari Mitra. Estimation issues in the Exponential– Logarithmic model under hybrid censoring. Statistical Papers, 62(1): 419–450, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01100-3.

Xi:2021:EMB

[2531] Daiqing Xi and Tianxiao Pang. Estimating multiple breaks in mean sequentially with fractionally integrated errors. Statistical Papers, 62(1): 451–494, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01104-z.

Cheng:2021:EPL

[2532] Suli Cheng and Jianbao Chen. Estimation of partially linear single-index spatial autoregressive model. Statistical Papers, 62(1):495–531, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-019-01105-
y.

Zha:2021:PCM

[2533] Ruochen Zha and Ofer Harel. Power calculation in multiply imputed data. Statistical Papers, 62(1): 533–559, February 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01098-8.

Lee:2021:MIP

[2534] Hyunju Lee and Ji Hwan Cha.

On a multivariate IFR and positively dependent lifetime model induced by multiple shot-noise processes. Statistical Papers, 62(2): 561–590, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01099-7.

Bahari:2021:ELI

[2535] Fayyaz Bahari, Safar Parsi, and Mojtaba Ganjali. Empirical likelihood inference in general linear model with missing values in response and covariates by MNAR mechanism. Statistical Papers, 62 (2):591–622, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01103-0.

Otto:2021:SPS

[2536] Philipp Otto, Wolfgang Schmid, and Robert Garthoff. Stochastic properties of spatial and spatiotemporal ARCH models. Statistical Papers, 62(2):623–638, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01106-x.

Hu:2021:PML

[2537] Hongchang Hu, Weifu Hu, and Xinxin Yu. Pseudo-maximum likelihood estimators in linear regression models with fractional time series. Statistical Papers, 62(2): 639–659, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01091-1.

Li:2021:NEV

[2538] Ning Li and Hu Yang. Nonnegative estimation and variable selection under minimax concave penalty for sparse high-dimensional linear regression models. Statistical Papers, 62 (2):661–680, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01107-w.

Watagoda:2021:BML

[2539] Lasanthi C. R. Pelawa Watagoda and David J. Olive. Bootstrapping multiple linear regression after variable selection. Statistical Papers, 62 (2):681–700, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01108-9.

Bai:2021:TLC

[2540] Zhidong Bai, Jiang Hu, and Chao Zhang. Test on the linear combinations of covariance matrices in highdimensional data. Statistical Papers, 62(2):701–719, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01110-1.

Jiang:2021:SLD

[2541] Hui Jiang, Jin Shao, and Qingshan Yang. Sharp large deviations for a class of normalized L-statistics and applications. Statistical Papers, 62 (2):721–744, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01109-8.

Kang:2021:NIP

[2542] Yao Kang, Dehui Wang, and Kai Yang. A new INAR(1) process with bounded support for counts showing equidispersion, underdispersion and overdispersion. Statistical Papers, 62 (2):745–767, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01111-0.

Wang:2021:CMC

[2543] Yan Wang and Xuejun Wang. Complete f-moment convergence for Sung’s type weighted sums and its application to the EV regression models. Statistical Papers, 62(2):

769–793, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01112-z.

Yu:2021:LSE

[2544] Qian Yu. Least squares estimator of fractional Ornstein–Uhlenbeck processes with periodic mean for general Hurst parameter. Statistical Papers, 62(2):795–815, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01113-y.

Zambom:2021:SIS

[2545] Adriano Zanin Zambom and Gregory J. Matthews. Sure independence screening in the presence of missing data. Statistical Papers, 62 (2):817–845, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01115-w.

Ghosh:2021:HE

[2546] Pallab Ghosh, Kevin Grier, and Jaeho Kim. Heterogeneous endogeneity. Statistical Papers, 62 (2):847–886, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01116-9.

Zhang:2021:EVS

[2547] Jun Zhang. Estimation and variable selection for partial linear singleindex distortion measurement errors models. Statistical Papers, 62 (2):887–913, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01119-6.

Liu:2021:ZOI

[2548] Wenchen Liu, Yincai Tang, and Ancha Xu. Zero-and-one-inflated Poisson regression model. Statistical Papers, 62(2):915–934, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01118-7.

Wang:2021:CWE

[2549] Xuejun Wang, Yi Wu, and Shuhe Hu. On consistency of wavelet estimator in nonparametric regression models. Statistical Papers, 62 (2):935–962, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01117-8.

Deng:2021:BET

[2550] Xin Deng, Xuejun Wang, and Yi Wu.

The Berry–Esseen type bounds of the weighted estimator in a nonparametric model with linear process errors. Statistical Papers, 62 (2):963–984, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-019-01120-z.

Varathan:2021:OSR

[2551] Nagarajah Varathan and Pushpakanthie Wijekoon. Optimal stochastic restricted logistic estimator. Statistical Papers, 62(2):985–1002, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-019-01121-
                                        y.

Castilla:2021:CLM

[2552] E. Castilla, N. Martín, and K. Zografos. Composite likelihood methods: Rao-type tests based on composite minimum density power divergence estimator. Statistical Papers, 62 (2):1003–1041, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01122-x.

Asl:2021:RTS

[2553] M. Nooi Asl, H. Bevrani, and K. Mansson. Ridge-type shrinkage estimators in generalized linear models with an application to prostate cancer data. Statistical Papers, 62(2): 1043–1085, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01123-w.

Meyer:2021:BRT

[2554] Marco Meyer. Book review: Tucker S. McElroy, Dimitris N. Politis (2020): Time series: a first course with bootstrap starter. Statistical Papers, 62 (2):1087–1089, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01207-y.

Ley:2021:BRT

[2555] Christophe Ley. Book review: Thomas A. Severini (2020): Analytic Methods in Sports — Using Mathematics and Statistics to Understand Data from Baseball, Football, Basketball, and Other Sports. Statistical Papers, 62 (2):1091–1092, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-020-01220-1.

Pebesma:2021:BRC

[2556] Edzer Pebesma. Book review: Christopher K. Wikle, Andrew Zammit-Mangion and Noel Cressie (2019): Spatio-temporal Statistics with R. Chapman and Hall/CRC, 396 pp. $47.96 (Hardcover), ISBN 978-1-1387-1113-6. Statistical Papers, 62(2): 1093–1094, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01224-5.

Taavoni:2021:KES

[2557] M. Taavoni and M. Arashi. Kernel estimation in semiparametric mixed effect longitudinal modeling. Statistical Papers, 62(3):1095–1116, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01125-8.

Yuan:2021:FSL

[2558] Ru Yuan, Bing Guo, and Min-Qian Liu. Flexible sliced Latin hypercube designs with slices of different sizes. Statistical Papers, 62(3):1117–1134, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-019-
                                        01127-6.

Nagy:2021:HDD

[2559] Stanislav Nagy. Halfspace depth does not characterize probability distributions. Statistical Papers, 62 (3):1135–1139, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01130-x.

Fang:2021:FSU

[2560] Jianglin Fang. Feature screening for ultrahigh-dimensional survival data when failure indicators are missing at random. Statistical Papers, 62 (3):1141–1166, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01128-5.

Zuo:2021:RDP

[2561] Yijun Zuo. Robustness of the deepest projection regression functional. Statistical Papers, 62(3):1167–1193, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-019-
                                        01129-4.

Chen:2021:PPE

[2562] Wangxue Chen, Rui Yang, Dongsen Yao, and Chunxian Long. Pareto parameters estimation using moving extremes ranked set sampling. Statistical Papers, 62(3):1195–1211, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00362-019-
01132-9.

Zhang:2021:FTH

[2563] Tao Zhang, Zhiwen Wang, and Yanling Wan. Functional test for highdimensional covariance matrix, with application to mitochondrial calcium concentration. Statistical Papers, 62 (3):1213–1230, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01133-8.

Perez-Gonzalez:2021:GFT

[2564] Ana Pérez-González, Tomás R.

Cotos-Yáñez, Wenceslao González-Manteiga, and Rosa M. Crujeiras-Casais. Goodness-of-fit tests for quantile regression with missing responses. Statistical Papers, 62(3):1231–1264, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01135-6.

Zhang:2021:RCB

[2565] Fengqing Zhang and Jiangtao Gou. Refined critical boundary with enhanced statistical power for non-directional two-sided tests in group sequential designs with multiple endpoints. Statistical Papers, 62(3):1265–1290, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01134-7.

Santos:2021:TPM

[2566] Cláudia Santos, Isabel Pereira, and Manuel G. Scotto. On the theory of periodic multivariate INAR processes. Statistical Papers, 62(3):1291–1348, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01136-5.

He:2021:ODH

[2567] Lei He and Rong-Xian Yue. D-optimal designs for hierarchical linear models with intraclass covariance structure. Statistical Papers, 62(3):1349–1361, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01139-2.

Frigau:2021:CVG

[2568] Luca Frigau, Claudio Conversano, and Francesco Mola. Consistent validation of gray-level thresholding image segmentation algorithms based on machine learning classifiers. Statistical Papers, 62(3):1363–1386, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01138-3.

Grigoriev:2021:ESO

[2569] Yu. D. Grigoriev, V. B. Melas, and P. V. Shpilev. Excess and saturated D-optimal designs for the rational model. Statistical Papers, 62 (3):1387–1405, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01140-9.

Fu:2021:ACS

[2570] Ke-Ang Fu, Ting Li, Chang Ni, Wenkai He, and Renshui Wu. Asymptotics for the conditional self-weighted Mestimator of GRCA(1) models with possibly heavy-tailed errors. Statistical Papers, 62(3):1407–1419, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01141-8.

Ozkale:2021:SRR

[2571] M. Revan Özkale and Hans Nyquist.

The stochastic restricted ridge estimator in generalized linear models. Statistical Papers, 62(3):1421–1460, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01142-7.

Elsawah:2021:SLB

[2572] A. M. Elsawah, Kai-Tai Fang, Ping He, and Hong Qin. Sharp lower bounds of various uniformity criteria for constructing uniform designs. Statistical Papers, 62(3):1461–1482, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01143-6.

Eberl:2021:NMA

[2573] Andreas Eberl and Bernhard Klar.

A note on a measure of asymmetry. Statistical Papers, 62(3):1483–1497, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-019-
                                        01145-4.

Chen:2021:AKR

[2574] Liangzhi Chen, Thomas Hotz, and Haizhang Zhang. Admissible kernels for RKHS embedding of probability distributions. Statistical Papers, 62 (3):1499–1518, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01144-5.

Punzo:2021:MHM

[2575] Antonio Punzo, Salvatore Ingrassia, and Antonello Maruotti. Multivariate hidden Markov regression models: random covariates and heavy-tailed distributions. Statistical Papers, 62 (3):1519–1555, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-019-01146-3.

Bailey:2021:AMP

[2576] R. A. Bailey and Peter J. Cameron. Appendage to: Multi-part balanced incomplete-block designs. Statistical Papers, 62(3):1557–1558, June 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-019-
                                        01137-4;   http://link.springer.
                                        com/content/pdf/10.1007/s00362-
                                        019-01137-4.pdf. See [2303].

Maronna:2021:RFP

[2577] Ricardo A. Maronna. Robust functional principal components for irregularly spaced longitudinal data. Statistical Papers, 62(4):1563–1582, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01147-2.

Cabana:2021:MOD

[2578] Elisa Cabana, Rosa E. Lillo, and Henry Laniado. Multivariate outlier detection based on a robust Mahalanobis distance with shrinkage estimators. Statistical Papers, 62(4):1583–1609, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01148-1.

Shams:2021:WEE

[2579] M. Shams. On weakly equivariant estimators. Statistical Papers, 62(4): 1611–1650, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01149-0.

Ameijeiras-Alonso:2021:OTC

[2580] Jose Ameijeiras-Alonso, Christophe Ley, Arthur Pewsey, and Thomas Verdebout. On optimal tests for circular reflective symmetry about an unknown central direction. Statistical Papers, 62(4):1651–1674, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01150-7.

Kauermann:2021:SPR

[2581] Göran Kauermann and Mehboob Ali.

Semi-parametric regression when some (expensive) covariates are missing by design. Statistical Papers, 62(4): 1675–1696, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01152-5.

Zhang:2021:GFT

[2582] Shulin Zhang, Qian M. Zhou, and Huazhen Lin. Goodness-of-fit test of copula functions for semi-parametric univariate time series models. Statistical Papers, 62(4):1697–1721, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01153-4.

Sun:2021:CSA

[2583] Zhihua Sun, Dongshan Luo, Xiaohua Zhou, and Qingzhao Zhang. Comparative studies on the adequacy check of parametric measurement error models with auxiliary variable. Statistical Papers, 62(4):1723–1751, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01154-3.

Mahdizadeh:2021:SEA

[2584] M. Mahdizadeh and Ehsan Zamanzade. Smooth estimation of the area under the ROC curve in multistage ranked set sampling. Statistical Papers, 62 (4):1753–1776, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01151-6.

Wang:2021:UPL

[2585] Kang Wang, Zujun Ou, Jiaqi Liu, and Hongyi Li. Uniformity pattern of q-level factorials under mixture discrepancy. Statistical Papers, 62(4): 1777–1793, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-019-01155-2.

Zhou:2021:COM

[2586] Weiping Zhou, Jinyu Yang, and Min-Qian Liu. Construction of orthogonal marginally coupled designs. Statistical Papers, 62(4):1795–1820, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-019-
01156-1.

Wang:2021:BMT

[2587] Rui Wang and Xingzhong Xu. A Bayesian-motivated test for highdimensional linear regression models with fixed design matrix. Statistical Papers, 62(4):1821–1852, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01157-5.

Hoffmann:2021:RFG

[2588] Ingo Hoffmann and Christoph J.

Börner. The risk function of the goodness-of-fit tests for tail models. Statistical Papers, 62(4):1853–1869, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00362-020-
                                        01159-3.

Idais:2021:LOV

[2589] Osama Idais. On local optimality of vertex type designs in generalized linear models. Statistical Papers, 62(4): 1871–1898, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01158-4;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-020-01158-4.pdf.

Hancova:2021:EVT

[2590] Martina Hancová, Andrej Gajdos, Jozef Hanc, and Gabriela Vozáriková. Estimating variances in time series kriging using convex optimization and empirical BLUPs. Statistical Papers, 62 (4):1899–1938, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01165-5.

Song:2021:EEV

[2591] Yuping Song, Hangyan Li, and Yetong Fang. Efficient estimation for the volatility of stochastic interest rate models. Statistical Papers, 62(4): 1939–1964, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01166-4.

Chasiotis:2021:OOB

[2592] Vasilis Chasiotis, Stavros A. Chatzopoulos, Stratis Kounias, and Nikos Farmakis. On the optimality of orthogonal and balanced arrays with N ≡ 0(mod9) runs. Statistical Papers, 62 (4):1965–1980, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01167-3.

Jiang:2021:AQR

[2593] Rong Jiang, Wei wei Chen, and Xin Liu. Adaptive quantile regressions for massive datasets. Statistical Papers, 62 (4):1981–1995, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01170-8.

Leschinski:2021:CST

[2594] Christian Leschinski, Michelle Voges, and Philipp Sibbertsen. A comparison of semiparametric tests for fractional cointegration. Statistical Papers, 62 (4):1997–2030, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01169-1;   http:
//link.springer.com/content/pdf/
10.1007/s00362-020-01169-1.pdf.

Chen:2021:UPN

[2595] Hao Chen, Yan Zhang, and Xue Yang.

Uniform projection nested Latin hypercube designs. Statistical Papers, 62 (4):2031–2045, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01172-6.

Ghale-Joogh:2021:MDE

[2596] Hassan Sharghi Ghale-Joogh and S. Mohammad E. Hosseini-Nasab. On mean derivative estimation of longitudinal and functional data: from sparse to dense. Statistical Papers, 62(4): 2047–2066, August 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01173-5.

Nyarko:2021:OPC

[2597] Eric Nyarko. Optimal 2^{K} paired comparison designs for third-order interactions. Statistical Papers, 62(5): 2067–2082, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01174-4.

Jelito:2021:NFT

[2598] Damian Jelito and Marcin Pitera. New fat-tail normality test based on conditional second moments with applications to finance. Statistical Papers, 62 (5):2083–2108, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01176-2;   http:
//link.springer.com/content/pdf/
10.1007/s00362-020-01176-2.pdf.

Zmyslony:2021:RTT

[2599] Roman Zmyślony and Arkadiusz Kozioł. Ratio F test for testing simultaneous hypotheses in models with blocked compound symmetric covariance structure. Statistical Papers, 62 (5):2109–2118, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01182-4;   http:

//link.springer.com/content/pdf/
10.1007/s00362-020-01182-4.pdf.

Wang:2021:MAC

[2600] Wan-Lun Wang, Luis M. Castro, Wan-Chen Hsieh, and Tsung-I Lin. Mixtures of factor analyzers with covariates for modeling multiply censored dependent variables. Statistical Papers, 62 (5):2119–2145, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01177-1.

Wang:2021:REE

[2601] Kangning Wang, Mengjie Hao, and Xiaofei Sun. Robust and efficient estimating equations for longitudinal data partial linear models and its applications. Statistical Papers, 62(5): 2147–2168, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01181-5.

Wu:2021:SCP

[2602] Yi Wu, Xuejun Wang, and Aiting Shen. Strong convergence properties for weighted sums of m-asymptotic negatively associated random variables and statistical applications. Statistical Papers, 62(5):2169–2194, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01179-z.

Abebe:2021:RES

[2603] Ash Abebe, Huybrechts F. Bindele, Masego Otlaadisa, and Boikanyo Makubate. Robust estimation of single index models with responses missing at random. Statistical Papers, 62 (5):2195–2225, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01184-2.

Cha:2021:VAR

[2604] Ji Hwan Cha and F. G. Badía. Variables acceptance reliability sampling plan based on degradation test. Statistical Papers, 62(5):2227–2245, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01185-1.

Zahmatkesh:2021:BPS

[2605] Samira Zahmatkesh and Mohsen Mohammadzadeh. Bayesian prediction of spatial data with non-ignorable missingness. Statistical Papers, 62(5): 2247–2268, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01186-0.

Maronna:2021:ORE

[2606] Ricardo A. Maronna and Victor J. Yohai. Optimal robust estimators for families of distributions on the integers. Statistical Papers, 62(5): 2269–2281, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01187-z.

Quessy:2021:NTM

[2607] Jean-François Quessy. On nonparametric tests of multivariate meta-ellipticity. Statistical Papers, 62(5):2283–2310, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01189-x.

Hasanalipour:2021:TSS

[2608] Parisa Hasanalipour and Mostafa Razmkhah. Testing skew-symmetry based on extreme ranked set sampling. Statistical Papers, 62(5):2311–2332, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01183-3.

Choi:2021:SNB

[2609] Ji-Eun Choi and Dong Wan Shin. A self-normalization break test for correlation matrix. Statistical Papers, 62 (5):2333–2353, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01188-y.

Shingaki:2021:SIG

[2610] Ryusei Shingaki, Hiroshi Kanda, and Manabu Kuroki. Selection and integration of generalized instrumental variables for estimating total effects. Statistical Papers, 62(5):2355–2381, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01190-4.

Liu:2021:DFR

[2611] Yongxin Liu, Peng Zeng, and Lu Lin. Degrees of freedom for regularized regression with Huber loss and linear constraints. Statistical Papers, 62(5): 2383–2405, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01192-2.

Watagoda:2021:CSS

[2612] Lasanthi C. R. Pelawa Watagoda and David J. Olive. Comparing six shrinkage estimators with large sample theory and asymptotically optimal prediction intervals. Statistical Papers, 62 (5):2407–2431, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01193-1.

Ou:2021:NFS

[2613] Zujun Ou and Hongyi Li. A new foldover strategy and optimal foldover plans for three-level design. Statistical Papers, 62(5):2433–2451, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01194-0.

Dizicheh:2021:BMJ

[2614] Mozhgan Alirezaei Dizicheh, Nasrollah Iranpanah, and Ehsan Zamanzade. Bootstrap methods for judgment post stratification. Statistical Papers, 62(5): 2453–2471, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01197-x.

Rueda:2021:PEL

[2615] Maria del Mar Rueda, Maria Giovanna Ranalli, Antonio Arcos, and David Molina. Population empirical likelihood estimation in dual frame surveys. Statistical Papers, 62(5): 2473–2490, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01200-5.

Hudecova:2021:TSA

[2616] Sárka Hudecová and Miroslav Siman.

Testing symmetry around a subspace. Statistical Papers, 62(5):2491–2508, October 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01201-4.

Liang:2021:PTT

[2617] Yuli Liang, Dietrich von Rosen, and Tatjana von Rosen. On properties of Toeplitz-type covariance matrices in models with nested random effects. Statistical Papers, 62(6): 2509–2528, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01202-3;   http:
//link.springer.com/content/pdf/
10.1007/s00362-020-01202-3.pdf.

Wang:2021:ECP

[2618] Fangfang Wang, Lu Lin, Lei Liu, and Kangning Wang. Estimation and clustering for partially heterogeneous single index model. Statistical Papers, 62(6): 2529–2556, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01203-2.

Kawka:2021:LTL

[2619] Rafael Kawka. Limit theorems for locally stationary processes. Statistical Papers, 62(6):2557–2571, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01204-1;   http://link.springer.
com/content/pdf/10.1007/s00362-
020-01204-1.pdf.

Cai:2021:GSI

[2620] Li Cai and Suojin Wang. Global statistical inference for the difference between two regression mean curves with covariates possibly partially missing. Statistical Papers, 62(6): 2573–2602, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01208-x.

Ahmadi:2021:CCS

[2621] Jafar Ahmadi. Characterization of continuous symmetric distributions using information measures of records. Statistical Papers, 62(6):2603–2626, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01206-z.

Liu:2021:MPV

[2622] Li Liu, Hao Wang, Yanyan Liu, and Jian Huang. Model pursuit and variable selection in the additive accelerated failure time model. Statistical Papers, 62(6):2627–2659, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01205-0.

Tu:2021:MAM

[2623] Jingwen Tu, Hu Yang, Chaohui Guo, and Jing Lv. Model averaging marginal regression for high dimensional conditional quantile prediction. Statistical Papers, 62(6):2661–2689, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01212-1.

Li:2021:SPT

[2624] Wenlong Li, Bing Guo, Hengzhen Huang, and Min-Qian Liu. Semifoldover plans for three-level orthogonal arrays with quantitative factors. Statistical Papers, 62(6):2691– 2709, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01211-2.

Zhang:2021:MFF

[2625] Jing Zhang, Yanyan Liu, and Hengjian Cui. Model-free feature screening via distance correlation for ultrahigh dimensional survival data. Statistical Papers, 62(6):2711–2738, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01210-3.

Berger:2021:BMU

[2626] Yves G. Berger, Paola M. Chiodini, and Mariangela Zenga. Bounds for monetary-unit sampling in auditing: an adjusted empirical likelihood approach. Statistical Papers, 62(6): 2739–2761, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01209-w;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-020-01209-w.pdf.

Kotb:2021:ERM

[2627] M. S. Kotb and M. Z. Raqab. Estimation of reliability for multi-component stress-strength model based on modified Weibull distribution. Statistical Papers, 62(6):2763–2797, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01213-0.

Qingguo:2021:EFL

[2628] Tang Qingguo and Bian Minjie. Estimation for functional linear semiparametric model. Statistical Papers, 62(6): 2799–2823, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01215-y.

Chen:2021:TEP

[2629] Mingjing Chen. Tests for the explanatory power of latent factors. Statistical Papers, 62(6):2825–2856, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01216-x.

Burgard:2021:GER

[2630] Jan Pablo Burgard, Joscha Krause, Dennis Kreber, and Domingo Morales. The generalized equivalence of regularization and min-max robustification in linear mixed models. Statistical Papers, 62(6):2857–2883, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-020-
01214-z;   http://link.springer.
com/content/pdf/10.1007/s00362-
020-01214-z.pdf.

Xia:2021:MAP

[2631] Xiaochao Xia. Model averaging prediction for nonparametric varyingcoefficient models with B-spline smoothing. Statistical Papers, 62(6): 2885–2905, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01218-9.

Li:2021:NMC

[2632] Hongyi Li, Xingyou Huang, Huili Xue, and Hong Qin. A novel method for constructing mixed two- and threelevel uniform factorials with large run sizes. Statistical Papers, 62(6): 2907–2921, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-020-01219-8.

Elsawah:2021:MDS

[2633] A. M. Elsawah. Multiple doubling: a simple effective construction technique for optimal two-level experimental designs. Statistical Papers, 62(6): 2923–2967, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-020-01221-0.

Cahoy:2021:FMO

[2634] Dexter Cahoy, Elvira Di Nardo, and Federico Polito. Flexible models for overdispersed and underdispersed count data. Statistical Papers, 62(6): 2969–2990, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01222-7;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-021-01222-7.pdf.

Ferreira:2021:EBP

[2635] Helena Ferreira, Ana Paula Martins, and Maria da Graça Temido. Extremal behaviour of a periodically controlled sequence with imputed values. Statistical Papers, 62(6):2991–3013, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-020-
                                        01217-w.

Khan:2021:EAT

[2636] Ruhul Ali Khan, Dhrubasish Bhattacharyya, and Murari Mitra. Exact and asymptotic tests of exponentiality against nonmonotonic mean time to failure type alternatives. Statistical Papers, 62(6):3015–3045, December 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-021-
01226-3.

Zhao:2022:RES

[2637] Jun Zhao, Guan’ao Yan, and Yi Zhang.

Robust estimation and shrinkage in ultrahigh dimensional expectile regression with heavy tails and variance heterogeneity. Statistical Papers, 63 (1):1–28, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-021-01227-2.

Beran:2022:NRB

[2638] Jan Beran, Britta Steffens, and Sucharita Ghosh. On nonparametric regression for bivariate circular longmemory time series. Statistical Papers, 63(1):29–52, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-021-01228-1;   http:
//link.springer.com/content/pdf/
10.1007/s00362-021-01228-1.pdf.

Wu:2022:PCL

[2639] Xiaofei Wu, Rongmei Liang, and Hu Yang. Penalized and constrained LAD estimation in fixed and high dimension. Statistical Papers, 63 (1):53–95, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-021-01229-0;   http:
//link.springer.com/content/pdf/
10.1007/s00362-021-01229-0.pdf.

Yang:2022:BEL

[2640] Kai Yang, Xue Ding, and Xiaohui Yuan. Bayesian empirical likelihood inference and order shrinkage for autoregressive models. Statistical Papers, 63 (1):97–121, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01231-6.

Matsuura:2022:OEU

[2641] Shun Matsuura and Hiroshi Kurata. Optimal estimator under risk matrix in a seemingly unrelated regression model and its generalized least squares expression. Statistical Papers, 63(1): 123–141, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01232-5.

Hoga:2022:QDD

[2642] Yannick Hoga. Quantifying the datadredging bias in structural break tests. Statistical Papers, 63(1): 143–155, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01233-4;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-021-01233-4.pdf.

Antognini:2022:SSI

[2643] Alessandro Baldi Antognini, Marco Novelli, and Maroussa Zagoraiou. A simple solution to the inadequacy of asymptotic likelihood-based inference for response-adaptive clinical trials. Statistical Papers, 63(1): 157–180, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-021-01234-3;   http:
//link.springer.com/content/pdf/
10.1007/s00362-021-01234-3.pdf.

Hassler:2022:UNC

[2644] Uwe Hassler and Mehdi Hosseinkouchack. Understanding nonsense correlation between (independent) random walks in finite samples. Statistical Papers, 63(1):181–195, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00362-021-
01237-0;   http://link.springer.
com/content/pdf/10.1007/s00362-
021-01237-0.pdf.

Xu:2022:PGR

[2645] Long-Hao Xu, Kai-Tai Fang, and Ping He. Properties and generation of representative points of the exponential distribution. Statistical Papers, 63 (1):197–223, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-021-01236-1.

Parkinson-Schwarz:2022:TED

[2646] Judith H. Parkinson-Schwarz and Arne C. Bathke. Testing for equality of distributions using the concept of (niche) overlap. Statistical Papers, 63 (1):225–242, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/10.
1007/s00362-021-01239-y;   http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00362-021-01239-y.pdf.

Li:2022:VSH

[2647] Tizheng Li and Xiaojuan Kang. Variable selection of higher-order partially linear spatial autoregressive model with a diverging number of parameters. Statistical Papers, 63(1):243–285, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-021-
                                        01241-4.

Ngatchou-Wandji:2022:CPT

[2648] Joseph Ngatchou-Wandji, Echarif Elharfaoui, and Michel Harel. On changepoints tests based on two-samples Ustatistics for weakly dependent observations. Statistical Papers, 63(1): 287–316, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01242-3.

Liu:2022:PSA

[2649] Zhan Liu and Chun Yip Yau. A propensity score adjustment method for longitudinal time series models under nonignorable nonresponse. Statistical Papers, 63(1):317–342, February 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00362-021-
                                        01261-0.

Silva:2022:KDM

[2650] Ivair R. Silva, Yan Zhuang, and Julio C. A. da Silva Junior. Kronecker delta method for testing independence between two vectors in high-dimension. Statistical Papers, 63(2):343–365, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01238-z.

Luo:2022:AWN

[2651] Jing Luo, Tour Liu, and Qiuping Wang. Affiliation weighted networks with a differentially private degree sequence. Statistical Papers, 63 (2):367–395, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01243-2.

Cheng:2022:CMC

[2652] Nan Cheng, Chao Lu, and Xuejun Wang. Complete moment convergence for randomly weighted sums of extended negatively dependent random variables with application to semiparametric regression models. Statistical Papers, 63 (2):397–419, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01244-1.

Ogundimu:2022:RVS

[2653] Emmanuel O. Ogundimu. Regularization and variable selection in Heckman selection model. Statistical Papers, 63(2):421–439, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01246-z.

Gao:2022:LAU

[2654] Yan-Ping Gao, Si-Yu Yi, and Yong-Dao Zhou. Level-augmented uniform designs. Statistical Papers, 63 (2):441–460, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01247-y.

Arellano-Valle:2022:SPU

[2655] Reinaldo B. Arellano-Valle and Adelchi Azzalini. Some properties of the unified skew-normal distribution. Statistical Papers, 63(2):461–487, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-021-01235-

2. See correction [2847].

Benrabah:2022:LLE

[2656] Ouafae Benrabah, Feriel Bouhadjera, and Elias Ould Saïd. Local linear estimation of the regression function for twice censored data. Statistical Papers, 63(2):489–514, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01240-5.

Li:2022:CCO

[2657] Wenlong Li, Min-Qian Liu, and Jian-Feng Yang. Construction of column-orthogonal strong orthogonal arrays. Statistical Papers, 63 (2):515–530, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01249-w.

Katsaounis:2022:MER

[2658] T. I. Katsaounis. On multistage experiments with restrictions in the randomization of treatments. Statistical Papers, 63(2):531–541, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-021-01254-
z.

Krupskii:2022:ALP

[2659] Pavel Krupskii and Harry Joe. Approximate likelihood with proxy variables for parameter estimation in high-dimensional factor copula models. Statistical Papers, 63(2): 543–569, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01252-1.

Faridi:2022:PGN

[2660] Masoud Faridi and Majid Jafari Khaledi. The polar-generalized normal distribution: properties, Bayesian estimation and applications. Statistical Papers, 63(2): 571–603, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01245-0.

Zou:2022:CLS

[2661] Tingting Zou, Shurong Zheng, and Hongtu Zhu. CLT for linear spectral statistics of large dimensional sample covariance matrices with dependent data. Statistical Papers, 63 (2):605–664, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01250-3.

Ling:2022:MRR

[2662] Nengxiang Ling, Lilei Cheng, and Hui Ding. Missing responses at random in functional single index model for time series data. Statistical Papers, 63(2):665–692, April 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01251-2.

Levy:2022:DSI

[2663] Edmond Levy. On the density for sums of independent exponential, Erlang and gamma variates. Statistical Papers, 63(3):693–721, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01256-x.

Fotopoulos:2022:CPD

[2664] Stergios B. Fotopoulos, Alex Paparas, and Venkata K. Jandhyala. Change point detection and estimation methods under gamma series of observations. Statistical Papers, 63 (3):723–754, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01248-x.

Mainassara:2022:PTA

[2665] Yacouba Boubacar Maïnassara, Othman Kadmiri, and Bruno Saussereau. Portmanteau test for the asymmetric power GARCH model when the power is unknown. Statistical Papers, 63(3):755–793, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01257-w.

Dang:2022:CIP

[2666] Bao-Anh Dang and K. Krishnamoorthy. Confidence intervals, prediction intervals and tolerance intervals for negative binomial distributions. Statistical Papers, 63 (3):795–820, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01255-y.

Bourguignon:2022:SUR

[2667] Marcelo Bourguignon, Diego I. Gallardo, and Rodrigo M. R. Medeiros. A simple and useful regression model for underdispersed count data based on Bernoulli–Poisson convolution. Statistical Papers, 63(3): 821–848, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01253-0.

Laketa:2022:HDG

[2668] Petra Laketa and Stanislav Nagy. Halfspace depth for general measures: the ray basis theorem and its consequences. Statistical Papers, 63 (3):849–883, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01259-8.

Zhang:2022:MNK

[2669] Jun Zhang, Bingqing Lin, and Yiping Yang. Maximum nonparametric kernel likelihood estimation for multiplicative linear regression models. Statistical Papers, 63(3): 885–918, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01258-9.

Figueroa-Zuniga:2022:RBR

[2670] Jorge I. Figueroa-Zúñiga, Cristian L. Bayes, and Shuangzhe Liu. Robust beta regression modeling with errors-in-variables: a Bayesian approach and numerical applications. Statistical Papers, 63(3):919–942, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-021-01260-
                                        1.

Kirkby:2022:IFS

[2671] J. Lars Kirkby, Dang H. Nguyen, and Nhu N. Nguyen. Inversion-free subsampling Newton’s method for large sample logistic regression. Statistical Papers, 63(3):943–963, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01263-y.

Mohan:2022:NPT

[2672] Renjith Mohan, Sreelakshmi N, and Sudheesh K. Kattumannil. Nonparametric test for decreasing renewal dichotomous Markov noise shock model. Statistical Papers, 63 (3):965–982, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01264-x.

Trifonov:2022:BRA

[2673] Mikhail Trifonov. Bessel representation for amplitude distribution of noisy sinusoidal signals. Statistical Papers, 63(3):983–988, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01262-z.

Kreiss:2022:BRS

[2674] Alexander Kreiß. Book review: Skyler J. Cranmer, Bruce A. Desmarais, Jason W. Morgan: Inferential network analysis. Statistical Papers, 63 (3):989–990, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01302-2.

Groll:2022:BRP

[2675] Andreas Groll and Carsten Jentsch.

Book review: Paola Zuccolotto and Marica Manisera (2020): Basketball Data Science: With Applications in R, CRC Press, 243 pp., £80.50 (Hardcover), ISBN: 978-1-138-60079-9. Statistical Papers, 63 (3):991–993, June 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01303-1.

Slynko:2022:AAR

[2676] Alla Slynko. Asymptotic analysis of reliability measures for an imperfect dichotomous test. Statistical Papers, 63 (4):995–1012, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01266-9.

Huskova:2022:THT

[2677] Marie Husková, Simos G. Meintanis, and Charl Pretorius. Tests for heteroskedasticity in transformation models. Statistical Papers, 63(4): 1013–1049, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01267-8.

Yuan:2022:HTS

[2678] Mingao Yuan, Fan Yang, and Zuofeng Shang. Hypothesis testing in sparse weighted stochastic block model. Statistical Papers, 63(4): 1051–1073, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-021-01269-6.

Choi:2022:QCC

[2679] Ji-Eun Choi and Dong Wan Shin. Quantile correlation coefficient: a new tail dependence measure. Statistical Papers, 63(4):1075–1104, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-021-01268-
                                        7.

Zhang:2022:THD

[2680] Jin-Ting Zhang, Bu Zhou, and Jia Guo. Testing high-dimensional mean vector with applications. Statistical Papers, 63 (4):1105–1137, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01270-z.

Shao:2022:OSC

[2681] Yujing Shao and Lei Wang. Optimal subsampling for composite quantile regression model in massive data. Statistical Papers, 63(4): 1139–1161, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01271-y.

Forughi:2022:PTG

[2682] Masoomeh Forughi, Zohreh Shishebor, and Atefeh Zamani. Portmanteau tests for generalized integervalued autoregressive time series models. Statistical Papers, 63(4): 1163–1185, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01274-9.

Nasari:2022:CIH

[2683] Masoud M. Nasari and Mohamedou Ould-Haye. Confidence intervals with higher accuracy for short and long-memory linear processes. Statistical Papers, 63(4): 1187–1220, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01265-w.

Gower:2022:PID

[2684] John C. Gower, Niël J. Le Roux, and Sugnet Gardner-Lubbe. Properties of individual differences scaling and its interpretation. Statistical Papers, 63 (4):1221–1245, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01275-8.

Weissbach:2022:TEU

[2685] Rafael Weißbach and Dominik Wied. Truncating the exponential with a uniform distribution. Statistical Papers, 63 (4):1247–1270, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01272-x.

Lando:2022:TCG

[2686] Tommaso Lando. Testing convexity of the generalised hazard function. Statistical Papers, 63(4): 1271–1289, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01273-w. See cor-

rection [2687].

Lando:2022:CTC

[2687] Tommaso Lando. Correction to: Testing convexity of the generalised hazard function. Statistical Papers, 63 (4):1291–1293, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01311-1.  See

[2686].

Hung:2022:GIC

[2688] Hung Hung, Su-Yun Huang, and Ching-Kang Ing. A generalized information criterion for high-dimensional PCA rank selection. Statistical Papers, 63(4):1295–1321, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-021-01276-
7.

Ciuperca:2022:RTD

[2689] Gabriela Ciuperca. Real-time detection of a change-point in a linear expectile model. Statistical Papers, 63 (4):1323–1367, August 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01278-5.

Chen:2022:GFT

[2690] Yiran Chen and Giray Ökten. A goodness-of-fit test for copulas based on the collision test. Statistical Papers, 63(5):1369–1385, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-021-01277-
6.

Marquis:2022:ICB

[2691] Bastien Marquis and Maarten Jansen.

Information criteria bias correction for group selection. Statistical Papers, 63 (5):1387–1414, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01283-8.

Wei:2022:USJ

[2692] Yang Wei, Zhouping Li, and Yunqiu Dai. Unified smoothed jackknife empirical likelihood tests for comparing income inequality indices. Statistical Papers, 63(5): 1415–1475, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01281-w.

Aoki:2022:LID

[2693] Reiko Aoki, Juan P. M. Bustamante, and Gilberto A. Paula. Local influence diagnostics with forward search in regression analysis. Statistical Papers, 63 (5):1477–1497, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01279-4.

Baksalary:2022:ALL

[2694] Oskar Maria Baksalary and Götz Trenkler. An alternative look at the linear regression model. Statistical Papers, 63 (5):1499–1509, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01280-x.

Gkelsinis:2022:SIB

[2695] Thomas Gkelsinis, Alex Karagrigoriou, and Vlad Stefan Barbu. Statistical inference based on weighted divergence measures with simulations and applications. Statistical Papers, 63(5): 1511–1536, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01286-z.

Fontana:2022:CRD

[2696] Roberto Fontana, Fabio Rapallo, and Henry P. Wynn. Circuits for robust designs. Statistical Papers, 63 (5):1537–1560, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01285-6.

Campos:2022:RMC

[2697] Laíla Luana Campos and Daniel Furtado Ferreira. Robust modified classical spherical tests in the presence of outliers. Statistical Papers, 63(5): 1561–1576, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01289-w.

Kupper:2022:RSD

[2698] Lawrence L. Kupper and Sandra L.

Martin. Replication study design: confidence intervals and commentary. Statistical Papers, 63(5): 1577–1583, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01291-2.

Dong:2022:EBN

[2699] Yuge Dong, Qingtong Xie, and Hu Wang. The evaluation of bivariate normal probabilities for failure of parallel systems. Statistical Papers, 63 (5):1585–1614, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01282-9.

Chesneau:2022:PEH

[2700] Christophe Chesneau, Salima El Kolei, and Fabien Navarro. Parametric estimation of hidden Markov models by least squares type estimation and deconvolution. Statistical Papers, 63(5): 1615–1648, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01288-x.

Yuan:2022:OSC

[2701] Xiaohui Yuan, Yong Li, and Tianqing Liu. Optimal subsampling for composite quantile regression in big data. Statistical Papers, 63(5): 1649–1676, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01292-1.

Mojirsheibani:2022:MDK

[2702] Majid Mojirsheibani. On the maximal deviation of kernel regression estimators with NMAR response variables. Statistical Papers, 63(5): 1677–1705, October 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01293-0.

Hunt:2022:EMS

[2703] Richard Hunt, Shelton Peiris, and Neville Weber. Estimation methods for stationary Gegenbauer processes. Statistical Papers, 63(6): 1707–1741, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01290-3.

Khodsiani:2022:OCE

[2704] Razieh Khodsiani and Saeid Pooladsaz. Optimality of circular equineighbored block designs under correlated observations. Statistical Papers, 63(6): 1743–1755, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01287-y.

Xiong:2022:SSP

[2705] Zikang Xiong, Wenjie Liu, Jianhui Ning, and Hong Qin. Sequential support points. Statistical Papers, 63(6): 1757–1775, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01294-z.

Mahdizadeh:2022:NIG

[2706] M. Mahdizadeh and Ehsan Zamanzade. New insights on goodnessof-fit tests for ranked set samples. Statistical Papers, 63(6):1777– 1799, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-021-01284-7.

Barakat:2022:EPI

[2707] H. M. Barakat and Haidy A. Newer.

Exact prediction intervals for future exponential and Pareto lifetimes based on ordered ranked set sampling of non-random and random size. Statistical Papers, 63(6): 1801–1827, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01295-y.

Ferri-Garcia:2022:VSP

[2708] Ramón Ferri-García and María del Mar Rueda. Variable selection in Propensity Score Adjustment to mitigate selection bias in online surveys. Statistical Papers, 63(6):1829– 1881, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01296-x.

Yu:2022:SSA

[2709] Jun Yu and HaiYing Wang. Subdata selection algorithm for linear model discrimination. Statistical Papers, 63(6): 1883–1906, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01299-8.

Kato:2022:CBM

[2710] Shogo Kato, Toshinao Yoshiba, and Shinto Eguchi. Copula-based measures of asymmetry between the lower and upper tail probabilities. Statistical Papers, 63(6):1907– 1929, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01297-w.

Patriota:2022:MEB

[2711] Alexandre Galvão Patriota. A measure of evidence based on the likelihood-ratio statistics. Statistical Papers, 63(6):1931–1951, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s00362-022-01301-
3.

Cardozo:2022:GLG

[2712] Carlos A. Cardozo, Gilberto A. Paula, and Luiz H. Vanegas. Generalized log-gamma additive partial linear models with P-spline smoothing. Statistical Papers, 63(6):1953– 1978, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01300-4.

Ozkale:2022:IRO

[2713] M. Revan Özkale and Atif Abbasi. Iterative restricted OK estimator in generalized linear models and the selection of tuning parameters via MSE and genetic algorithm. Statistical Papers, 63(6): 1979–2040, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01304-0.

Qingguo:2022:EPL

[2714] Tang Qingguo and Chen Wenyu. Estimation for partially linear additive regression with spatial data. Statistical Papers, 63(6):2041–2063, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01326-
8.

Zielinski:2022:SCI

[2715] Wojciech Zieliński. The shortest confidence interval for Poisson mean. Statistical Papers, 63(6): 2065–2072, December 2022. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01305-z.

Jain:2023:RMR

[2716] Kanchan Jain, Sudheesh K. Kattumannil, and Anjana Rajagopal. Replacement model with random replacement time. Statistical Papers, 64 (1):1–15, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01306-y.

Juodakis:2023:ECD

[2717] Julius Juodakis and Stephen Marsland. Epidemic changepoint detection in the presence of nuisance changes. Statistical Papers, 64(1): 17–39, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01307-x.

Lazecka:2023:SEB

[2718] Małgorzata Lazecka and Jan Mielniczuk. Squared error-based shrinkage estimators of discrete probabilities and their application to variable selection. Statistical Papers, 64 (1):41–72, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01308-w.

Mirezi:2023:ALE

[2719] Buatikan Mirezi and Selahattin Kaçiranlar. Admissible linear estimators in the general Gauss–Markov model under generalized extended balanced loss function. Statistical Papers, 64(1): 73–92, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01298-9.

Weng:2023:DIN

[2720] Lin-Chen Weng, Kai-Tai Fang, and A. M. Elsawah. Degree of isomorphism: a novel criterion for identifying and classifying orthogonal designs. Statistical Papers, 64(1): 93–116, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01310-2.

Sheena:2023:CED

[2721] Yo Sheena. Convergence of estimative density: criterion for model complexity and sample size. Statistical Papers, 64 (1):117–137, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01309-9.

Aly:2023:PID

[2722] Amany E. Aly. Predictive inference of dual generalized order statistics from the inverse Weibull distribution. Statistical Papers, 64(1): 139–160, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01312-0.

Zamanzade:2023:RBE

[2723] Elham Zamanzade, Majid Asadi, Afshin Parvardeh, and Ehsan Zamanzade. A ranked-based estimator of the mean past lifetime with an application. Statistical Papers, 64(1): 161–177, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01314-y.

Curto:2023:IAA

[2724] José Dias Curto. Inference about the arithmetic average of log transformed data. Statistical Papers, 64 (1):179–204, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01315-x.

Yue:2023:COD

[2725] Zongzhi Yue, Xiaoqing Zhang, P. van den Driessche, and Julie Zhou. Constructing K-optimal designs for regression models. Statistical Papers, 64 (1):205–226, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01317-9.

Laria:2023:GLA

[2726] Juan C. Laria, M. Carmen Aguilera-Morillo, and Rosa E. Lillo. Group linear algorithm with sparse principal decomposition: a variable selection and clustering method for generalized linear models. Statistical Papers, 64 (1):227–253, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01313-z.

Dehghan:2023:TDS

[2727] Sakineh Dehghan, Mohammad Reza Faridrohani, and Zahra Barzegar. Testing for diagonal symmetry based on center-outward ranking. Statistical Papers, 64(1):255–283, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01316-
w.

Maddanu:2023:FHP

[2728] Federico Maddanu. Forecasting highly persistent time series with bounded spectrum processes. Statistical Papers, 64(1):285–319, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01321-
z.

Ribeiro:2023:REB

[2729] Terezinha K. A. Ribeiro and Silvia L. P. Ferrari. Robust estimation in beta regression via maximum L_{q}- likelihood. Statistical Papers, 64(1): 321–353, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01320-0.

Aisbett:2023:AGF

[2730] Janet Aisbett, Eric J. Drinkwater, Kenneth L. Quarrie, and Stephen Woodcock. Applying generalized funnel plots to help design statistical analyses. Statistical Papers, 64(1): 355–364, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01322-y.

Nkurunziza:2023:CES

[2731] Sévérien Nkurunziza. Correction to: On efficiency of some restricted estimators in a multivariate regression model. Statistical Papers, 64 (1):365, February 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01339-3.  See

[2742].

Lin:2023:RFA

[2732] Tsung-I Lin, I-An Chen, and Wan-Lun Wang. A robust factor analysis model based on the canonical fundamental skew-t distribution. Statistical Papers, 64(2):367–393, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01318-8.

Clarke:2023:FSC

[2733] Brenton R. Clarke and Andrew Grose. A further study comparing forward search multivariate outlier methods including ATLA with an application to clustering. Statistical Papers, 64 (2):395–420, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01319-7.

Bouhadjera:2023:SUC

[2734] Feriel Bouhadjera, Mohamed Lemdani, and Elias Ould Saïd. Strong uniform consistency of the local linear relative error regression estimator under left truncation. Statistical Papers, 64(2):421–447, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01325-9.

Rai:2023:PCB

[2735] Mritunjay Rai, Tanmoy Maity, Agha Asim Husain, and R. K. Yadav. Pearson’s correlation and background subtraction (BGS) based approach for object’s motion detection in infrared video frame sequences. Statistical Papers, 64(2):449–475, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01323-x.

Wu:2023:EII

[2736] Hong-Jiang Wu, Ying-Ying Zhang, and Han-Yu Li. Expectation identities from integration by parts for univariate continuous random variables with applications to high-order moments. Statistical Papers, 64 (2):477–496, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01329-5.

Alruwaili:2023:MSD

[2737] Bader Alruwaili. The modality of skew t-distribution. Statistical Papers, 64(2):497–507, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01328-6.

Eberl:2023:SOM

[2738] Andreas Eberl and Bernhard Klar.

Stochastic orders and measures of skewness and dispersion based on expectiles. Statistical Papers, 64 (2):509–527, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01331-x.

Li:2023:MTH

[2739] Xu Li, Wenjuan Hu, and Baoxue Zhang. Measuring and testing homogeneity of distributions by characteristic distance. Statistical Papers, 64(2):529–556, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01327-7.

Mesfioui:2023:CBL

[2740] M. Mesfioui, T. Bouezmarni, and M. Belalia. Copula-based link functions in binary regression models. Statistical Papers, 64(2): 557–585, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01330-y.

Duarte:2023:ODE

[2741] Belmiro P. M. Duarte, Anthony C. Atkinson, Satya P. Singh, and Marco S. Reis. Optimal design of experiments for hypothesis testing on ordered treatments via intersectionunion tests. Statistical Papers, 64 (2):587–615, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01334-8.

Nkurunziza:2023:ESR

[2742] Sévérien Nkurunziza. On efficiency of some restricted estimators in a multivariate regression model. Statistical Papers, 64(2): 617–642, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01324-w. See cor-

rection [2731].

Chakraborty:2023:CCC

[2743] Saptarshi Chakraborty and Samuel W. K. Wong. On the circular correlation coefficients for bivariate von Mises distributions on a torus. Statistical Papers, 64 (2):643–675, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01333-9.

Filipiak:2023:TIU

[2744] Katarzyna Filipiak, Mateusz John, and Daniel Klein. Testing independence under a block compound symmetry covariance structure. Statistical Papers, 64(2):677–704, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01335-7.

Kulathinal:2023:WSL

[2745] Sangita Kulathinal and Isha Dewan. Weighted U-statistics for likelihood-ratio ordering of bivariate data. Statistical Papers, 64 (2):705–735, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01332-w.

Zhao:2023:RAE

[2746] Yichuan Zhao and Min Lu. RE-TRACTED ARTICLE: Empirical likelihood inference for the calibration regression model with lifetime medical cost. Statistical Papers, 64(2):737, April 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-010-0346-1.

Ebner:2023:EAL

[2747] Bruno Ebner and Norbert Henze. On the eigenvalues associated with the limit null distribution of the Epps–Pulley test of normality. Statistical Papers, 64(3):739–752, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01336-
6.

Hu:2023:WSB

[2748] Junwei Hu and Lihong Wang. A weighted U-statistic based change point test for multivariate time series. Statistical Papers, 64(3): 753–778, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01341-9.

DAngelo:2023:LIS

[2749] Nicoletta D’Angelo, Giada Adelfio, and Jorge Mateu. Local inhomogeneous second-order characteristics for spatiotemporal point processes occurring on linear networks. Statistical Papers, 64(3):779–805, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01338-4.

Wang:2023:CCF

[2750] Ping Wang and Lu Lin. Conditional characteristic feature screening for massive imbalanced data. Statistical Papers, 64(3):807–834, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01342-
8.

Parsa:2023:AFT

[2751] Motahareh Parsa and Ingrid Van Keilegom. Accelerated failure time vs Cox proportional hazards mixture cure models: David vs Goliath? Statistical Papers, 64 (3):835–855, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01345-5.

Leckey:2023:SPR

[2752] Kevin Leckey, Dennis Malcherczyk, Melanie Horn, and Christine H. Müller. Simple powerful robust tests based on sign depth. Statistical Papers, 64(3):857–882, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01337-5.

Perrone:2023:SUC

[2753] Gabriele Perrone and Gabriele Soffritti. Seemingly unrelated clusterwise linear regression for contaminated data. Statistical Papers, 64 (3):883–921, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01344-6.

Hu:2023:LDL

[2754] Mingyue Hu and Yongcheng Qi. Limiting distributions of the likelihood ratio test statistics for independence of normal random vectors. Statistical Papers, 64(3):923–954, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01348-2.

Facevicova:2023:CCN

[2755] Kamila Facevicová, Peter Filzmoser, and Karel Hron. Compositional cubes: a new concept for multi-factorial compositions. Statistical Papers, 64 (3):955–985, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01350-8.

Guo:2023:COG

[2756] Bing Guo, Xiao-Rong Li, Min-Qian Liu, and Xue Yang. Construction of orthogonal general sliced Latin hypercube designs. Statistical Papers, 64 (3):987–1014, June 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01347-3.

Woods:2023:ESI

[2757] David C. Woods, Stefanie Biedermann, and Chiara Tommasi. Editorial for the special issue for mODa 13: model-oriented data analysis and optimum design. Statistical Papers, 64 (4):1015–1019, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01453-w.

Radloff:2023:ONO

[2758] Martin Radloff and Rainer Schwabe.

D-optimal and nearly D-optimal exact designs for binary response on the ball. Statistical Papers, 64(4): 1021–1040, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01434-z.

Tommasi:2023:ETD

[2759] Chiara Tommasi, Juan M. Rodríguez-Díaz, and Jesús F. López-Fidalgo. An equivalence theorem for design optimality with respect to a multi-objective criterion. Statistical Papers, 64(4): 1041–1056, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01431-2.

Prus:2023:ODP

[2760] Maryna Prus. Optimal designs for prediction in random coefficient regression with one observation per individual. Statistical Papers, 64(4): 1057–1068, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01440-1.

Yu:2023:IBO

[2761] Jun Yu, Jiaqi Liu, and HaiYing Wang. Information-based optimal subdata selection for non-linear models. Statistical Papers, 64(4): 1069–1093, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01430-3.

Reuter:2023:OSD

[2762] Torsten Reuter and Rainer Schwabe. Optimal subsampling design for polynomial regression in one covariate. Statistical Papers, 64(4): 1095–1117, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01425-0.

Deldossi:2023:AOO

[2763] Laura Deldossi, Elena Pesce, and Chiara Tommasi. Accounting for outliers in optimal subsampling methods. Statistical Papers, 64(4): 1119–1135, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01422-3.

Mahendran:2023:MRS

[2764] Amalan Mahendran, Helen Thompson, and James M. McGree. A model robust subsampling approach for generalised linear models in big data settings. Statistical Papers, 64(4): 1137–1157, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01446-9.

Muller:2023:ODS

[2765] Christine H. Müller and Kirsten Schorning. A-optimal designs for state estimation in networks. Statistical Papers, 64(4):1159–1186, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01435-
y.

Pronzato:2023:BAO

[2766] Luc Pronzato and Anatoly Zhigljavsky. BLUE against OLSE in the location model: energy minimization and asymptotic considerations. Statistical Papers, 64(4): 1187–1208, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01423-2.

Strouwen:2023:ARE

[2767] Arno Strouwen, Bart M. Nicolaï, and Peter Goos. Adaptive and robust experimental design for linear dynamical models using Kalman filter. Statistical Papers, 64(4): 1209–1231, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01438-9.

Maruri-Aguilar:2023:SPP

[2768] Hugo Maruri-Aguilar and Henry Wynn. Sparse polynomial prediction. Statistical Papers, 64(4): 1233–1249, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01439-8.

Fontana:2023:DEM

[2769] Roberto Fontana, Alberto Molena, Luca Pegoraro, and Luigi Salmaso. Design of experiments and machine learning with application to industrial experiments. Statistical Papers, 64(4): 1251–1274, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01437-w.

Yousefi:2023:DBG

[2770] Elham Yousefi, Luc Pronzato, Markus Hainy, Werner G. Müller, and Henry P. Wynn. Discrimination between Gaussian process models: active learning and static constructions. Statistical Papers, 64(4): 1275–1304, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01436-x.

Antognini:2023:NIA

[2771] Alessandro Baldi Antognini, Rosamarie Frieri, and Maroussa Zagoraiou. New insights into adaptive enrichment designs. Statistical Papers, 64(4): 1305–1328, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01433-0.

Flournoy:2023:PAI

[2772] Nancy Flournoy and Sergey Tarima. Posterior alternatives with informative early stopping. Statistical Papers, 64 (4):1329–1341, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01429-w.

Chen:2023:LMD

[2773] Younan Chen, Michael Fries, and Sergei Leonov. Longitudinal model for a dose-finding study for a rare disease treatment. Statistical Papers, 64 (4):1343–1360, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01424-1.

Tarima:2023:GST

[2774] Sergey Tarima and Nancy Flournoy.

Group sequential tests: beyond exponential family models. Statistical Papers, 64(4):1361–1372, August 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01432-
1.

Chadjiconstantinidis:2023:CWT

[2775] Stathis Chadjiconstantinidis and Serkan Eryilmaz. Computing waiting time probabilities related to (k_{1}, k_{2}, …, k_{l}) pattern. Statistical Papers, 64(5): 1373–1390, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01351-7.

Crescenzo:2023:SAF

[2776] Antonio Di Crescenzo, Paola Paraggio, Patricia Román-Román, and Francisco Torres-Ruiz. Statistical analysis and first-passage-time applications of a lognormal diffusion process with multi-sigmoidal logistic mean. Statistical Papers, 64(5): 1391–1438, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01349-1.

Nascimento:2023:DBT

[2777] Abraão Nascimento, Jodavid Ferreira, and Alisson Silva. Divergence-based tests for the bivariate gamma distribution applied to polarimetric synthetic aperture radar. Statistical Papers, 64 (5):1439–1463, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01354-4.

Zuo:2023:NAA

[2778] Yijun Zuo. Non-asymptotic analysis and inference for an outlyingness induced winsorized mean. Statistical Papers, 64(5):1465–1481, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01353-
5.

Perri:2023:AEI

[2779] Pier Francesco Perri, Eleni Manoli, and Tasos C. Christofides. Assessing the effectiveness of indirect questioning techniques by detecting liars. Statistical Papers, 64(5): 1483–1506, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01352-6.

Zhao:2023:OET

[2780] Renren Zhao and Robert L. Paige. Optimal equivalence testing in exponential families. Statistical Papers, 64(5): 1507–1525, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01346-4.

Siray:2023:SPU

[2781] Gülesen Üstündag Siray. Simultaneous prediction using target function based on principal components estimator with correlated errors. Statistical Papers, 64(5): 1527–1628, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01340-w.

Hron:2023:BDB

[2782] Karel Hron, Jitka Machalová, and Alessandra Menafoglio. Bivariate densities in Bayes spaces: orthogonal decomposition and spline representation. Statistical Papers, 64(5): 1629–1667, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01359-z.

Barbiero:2023:DAC

[2783] Alessandro Barbiero and Asmerilda Hitaj. Discrete approximations of continuous probability distributions obtained by minimizing Cramér–von Mises-type distances. Statistical Papers, 64(5): 1669–1697, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01356-2.

Pan:2023:PUN

[2784] Siyu Pan, Jie Li, Zujun Ou, and Peng Zhu. Projection uniformity of nearly balanced designs. Statistical Papers, 64 (5):1699–1720, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01358-0.

Otto:2023:GFS

[2785] Philipp Otto and Wolfgang Schmid. A general framework for spatial GARCH models. Statistical Papers, 64(5): 1721–1747, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01357-1.

Livingston:2023:BIM

[2786] G. C. Livingston, Jr. and Darfiana Nur. Bayesian inference of multivariate-GARCH-BEKK models. Statistical Papers, 64(5):1749– 1774, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01360-6.

Jena:2023:PIE

[2787] Pravash Jena, Manas Ranjan Tripathy, and Somesh Kumar. Point and interval estimation of powers of scale parameters for two normal populations with a common mean. Statistical Papers, 64 (5):1775–1804, October 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01361-5.

Wang:2023:PSE

[2788] Yijun Wang, Weiwei Wang, and Xiaobing Zhao. Polynomial spline estimation of panel count data model with an unknown link function. Statistical Papers, 64(6):1805– 1832, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01364-2.

Modarres:2023:NCH

[2789] Reza Modarres. Nonparametric classification of high dimensional observations. Statistical Papers, 64(6): 1833–1859, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01363-3.

Ejsmont:2023:TNI

[2790] Wiktor Ejsmont, Bojana Milosević, and Marko Obradović. A test for normality and independence based on characteristic function. Statistical Papers, 64(6):1861–1889, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01365-
1.

Belaghi:2023:ISE

[2791] Reza Arabi Belaghi, Yasin Asar, and Rolf Larsson. Improved shrinkage estimators in the beta regression model with application in econometric and educational data. Statistical Papers, 64(6):1891–1912, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-022-01355-
                                        3.

Mondal:2023:TOA

[2792] Anjana Mondal, Markus Pauly, and Somesh Kumar. Testing for ordered alternatives in heteroscedastic ANOVA under normality. Statistical Papers, 64(6):1913– 1941, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01366-0.

Yan:2023:OSF

[2793] Qian Yan, Hanyu Li, and Chengmei Niu. Optimal subsampling for functional quantile regression. Statistical Papers, 64(6):1943–1968, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-022-01367-
                                        z.

Biswas:2023:NRA

[2794] Aniket Biswas and Gaurangadeb Chattopadhyay. New results for adaptive false discovery rate control with p-value weighting. Statistical Papers, 64(6): 1969–1996, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01369-x.

He:2023:BMO

[2795] Lei He and Daojiang He. Bayesian and maximin optimal designs for heteroscedastic multi-factor regression models. Statistical Papers, 64(6): 1997–2013, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01368-y.

Klement:2023:MCG

[2796] Erich Peter Klement, Damjana Kokol Bukovšek, Matjaž Omladič, Susanne Saminger-Platz, and Nik Stopar. Multivariate copulas with given values at two arbitrary points. Statistical Papers, 64(6):2015–2055, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01362-
4.

Chen:2023:SDM

[2797] Hui Chen, Linhan Ouyang, Lijun Liu, and Yizhong Ma. Sequential design of multi-fidelity computer experiments with effect sparsity. Statistical Papers, 64(6):2057– 2080, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01370-4.

Yin:2023:BGM

[2798] Xuehua Yin, Narayanaswamy Balakrishnan, and Chuancun Yin. Bounds for Gini’s mean difference based on first four moments, with some applications. Statistical Papers, 64(6): 2081–2100, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01374-0.

Al-Momani:2023:PSE

[2799] Marwan Al-Momani, M. Riaz, and M. F. Saleh. Pretest and shrinkage estimation of the regression parameter vector of the marginal model with multinomial responses. Statistical Papers, 64(6):2101–2117, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-022-01372-
                                        2.

Zhao:2023:NCF

[2800] Jun Zhao and Hyoung-Moon Kim. New closed-form efficient estimators for the negative binomial distribution. Statistical Papers, 64(6):2119– 2135, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01373-1.

Singh:2023:CEB

[2801] Rahul Singh and Neeraj Misra. A class of estimators based on overlapping sample spacings. Statistical Papers, 64(6):2137–2160, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-022-01377-
                                        x.

Maria:2023:SMM

[2802] Chiara Di Maria. Structural multilevel models for longitudinal mediation analysis: a definition variable approach. Statistical Papers, 64(6): 2161–2182, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01378-w.

Cascos:2023:ZRP

[2803] Ignacio Cascos, Giuseppe Pandolfo, and Beatriz Sinova. The zonoid region parameter depth. Statistical Papers, 64(6):2183–2205, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-022-01380-
2.

Lai:2023:VSN

[2804] Peng Lai, Xi Yan, Xin Sun, Haozhe Pang, and Yanqiu Zhou. Variable selection for nonparametric quantile regression via measurement error model. Statistical Papers, 64(6): 2207–2224, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01376-y.

Hamura:2023:RTE

[2805] Yasuyuki Hamura and Tatsuya Kubokawa.

Robustness of a truncated estimator for the smaller of two ordered means. Statistical Papers, 64(6): 2225–2244, December 2023. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01371-3.

Jokiel-Rokita:2024:EPQ

[2806] Alicja Jokiel-Rokita and Sylwester Piatek. Estimation of parameters and quantiles of the Weibull distribution. Statistical Papers, 65 (1):1–18, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01379-9.

Antognini:2024:ECA

[2807] Alessandro Baldi Antognini, Rosamarie Frieri, Maroussa Zagoraiou, and Marco Novelli. The Efficient Covariate– Adaptive Design for high-order balancing of quantitative and qualitative covariates. Statistical Papers, 65 (1):19–44, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01381-1.

Ebner:2024:CCN

[2808] Bruno Ebner, Lena Eid, and Bernhard Klar. Cauchy or not Cauchy? New goodness-of-fit tests for the Cauchy distribution. Statistical Papers, 65 (1):45–78, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01382-0.

Mathew:2024:JEL

[2809] Deemat C. Mathew, Reeba Mary Alex, and Sudheesh K. Kattumannil. Jackknife empirical likelihood ratio test for testing mean time to failure order. Statistical Papers, 65 (1):79–92, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01385-x.

Liu:2024:EVS

[2810] Yanxia Liu, Zhihao Wang, Maozai Tian, and Keming Yu. Estimation and variable selection for generalized functional partially varying coefficient hybrid models. Statistical Papers, 65 (1):93–119, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01383-z.

Mathai:2024:DSS

[2811] A. M. Mathai and Serge B. Provost.

On the distribution of sample scale-free scatter matrices. Statistical Papers, 65 (1):121–138, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01388-8.

Roychowdhury:2024:REA

[2812] Sayoni Roychowdhury, Indrila Ganguly, and Abhik Ghosh. Robust estimation of average treatment effects from panel data. Statistical Papers, 65 (1):139–179, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01389-7.

Mohammad:2024:PLE

[2813] Khandoker Akib Mohammad, Yuichi Hirose, Budhi Surya, and Yuan Yao. Profile likelihood estimation for the Cox proportional hazards (PH) cure model and standard errors. Statistical Papers, 65(1): 181–201, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01387-9.

Davidov:2024:UHE

[2814] Ori Davidov and Tamás Rudas. On the use of historical estimates. Statistical Papers, 65(1):203–236, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-022-01375-
                                        z.

Fang:2024:HDP

[2815] Xiao Fang and Malay Ghosh. Highdimensional properties for empirical priors in linear regression with unknown error variance. Statistical Papers, 65(1):237–262, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-022-01390-
                                        0.

Cruz:2024:CSA

[2816] N. A. Cruz, O. O. Melo, and C. A. Martinez. A correlation structure for the analysis of Gaussian and non-Gaussian responses in crossover experimental designs with repeated measures. Statistical Papers, 65(1): 263–290, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-022-01391-z.

Gupta:2024:SCC

[2817] Nitin Gupta and Santosh Kumar Chaudhary. Some characterizations of continuous symmetric distributions based on extropy of record values. Statistical Papers, 65(1): 291–308, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01392-y.

Evans:2024:SPB

[2818] Michael Evans, Miaoshiqi Liu, Michael Moon, Sabrina Sixta, Siyi Wei, and Siyue Yang. On some problems of Bayesian region construction with guaranteed coverages. Statistical Papers, 65(1):309–334, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01394-
4.

El-Horbaty:2024:MCP

[2819] Yahia S. El-Horbaty and Eman M.

Hanafy. A Monte Carlo permutation procedure for testing variance components using robust estimation methods. Statistical Papers, 65(1): 335–356, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01396-2.

Li:2024:SNC

[2820] Wenlong Li, Min-Qian Liu, and Jian-Feng Yang. Several new classes of space-filling designs. Statistical Papers, 65(1):357–379, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01402-
7.

Cmiel:2024:GSD

[2821] Bogdan Ćmiel, Jakub Nawała, Lucjan Janowski, and Krzysztof Rusek.

Generalised score distribution: underdispersed continuation of the betabinomial distribution. Statistical Papers, 65(1):381–413, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01398-
0.

Eberl:2024:CFK

[2822] Andreas Eberl and Bernhard Klar. Centre-free kurtosis orderings for asymmetric distributions. Statistical Papers, 65(1):415–433, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01403-
6.

Jimenez-Gamero:2024:TNL

[2823] M. D. Jiménez-Gamero. Testing normality of a large number of populations. Statistical Papers, 65(1): 435–465, February 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01384-y.

Yu:2024:RDI

[2824] Jun Yu, Mingyao Ai, and Zhiqiang Ye. A review on design inspired subsampling for big data. Statistical Papers, 65(2):467–510, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-022-01386-w.

Mondal:2024:MMS

[2825] Sagnik Mondal, Reinaldo B. Arellano-Valle, and Marc G. Genton. A multivariate modified skew-normal distribution. Statistical Papers, 65 (2):511–555, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01397-1.

Baringhaus:2024:DMR

[2826] Ludwig Baringhaus and Rudolf Grübel.

Discrete mixture representations of spherical distributions. Statistical Papers, 65(2):557–596, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01393-
5.

Mammadova:2024:DSC

[2827] Ulduz Mammadova and M. Revan Özkale. Detecting shifts in Conway– Maxwell–Poisson profile with deviance residual-based CUSUM and EWMA charts under multicollinearity. Statistical Papers, 65(2): 597–643, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01399-z.

Faouzi:2024:CAB

[2828] Tarik Faouzi, Emilio Porcu, Igor Kondrashuk, and Moreno Bevilacqua. Convergence arguments to bridge Cauchy and Matérn covariance functions. Statistical Papers, 65(2):645–660, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01400-9.

Buono:2024:SPF

[2829] Francesco Buono and Jorge Navarro. Simulations and predictions of future values in the time-homogeneous loadsharing model. Statistical Papers, 65 (2):661–685, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01404-5.

Ahmad:2024:MCH

[2830] Amar S. Ahmad, Munther Al-Hassan, Hamid Y. Hussain, Nirmin F. Juber, Fred N. Kiwanuka, Mohammed Hag-Ali, and Raghib Ali. A method of correction for heaping error in the variables using validation data. Statistical Papers, 65 (2):687–704, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01405-4.

Aghabozorgi:2024:CEF

[2831] Hamid Haji Aghabozorgi and Farzad Eskandari. Clustering and estimation of finite mixture models under bivariate ranked set sampling with application to a breast cancer study. Statistical Papers, 65 (2):705–736, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01411-6.

Choi:2024:IVE

[2832] Byeong Yeob Choi. Instrumental variable estimation of weighted local average treatment effects. Statistical Papers, 65(2):737–770, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01415-
2.

Hennig:2024:PEI

[2833] Christian Hennig. Parameters not empirically identifiable or distinguishable, including correlation between Gaussian observations. Statistical Papers, 65(2):771–794, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01414-3.

Centofanti:2024:SSF

[2834] Fabio Centofanti, Antonio Lepore, and Biagio Palumbo. Sparse and smooth functional data clustering. Statistical Papers, 65(2): 795–825, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01408-1.

Xue:2024:ERE

[2835] Liugen Xue and Junshan Xie. Efficient robust estimation for singleindex mixed effects models with missing observations. Statistical Papers, 65(2):827–864, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01407-2.

Prus:2024:CAE

[2836] Maryna Prus and Lenka Filová.

Computational aspects of experimental designs in multiple-group mixed models. Statistical Papers, 65(2):

865–886, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01416-1.

Liu:2024:NGB

[2837] Yonghui Liu, Yichen Lin, Xin Song, Conan Liu, and Shuangzhe Liu. Nonnegative group bridge and application in financial index tracking. Statistical Papers, 65(2): 887–907, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01406-3.

Adjakossa:2024:KRA

[2838] Eric Adjakossa, Yannig Goude, and Olivier Wintenberger. Kalman recursions aggregated online. Statistical Papers, 65(2):909–944, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01410-
                                        7.

Meilan-Vila:2024:NEF

[2839] Andrea Meilán-Vila, Rosa M. Crujeiras, and Mario Francisco-Fernández. Nonparametric estimation for a functionalcircular regression model. Statistical Papers, 65(2):945–974, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01420-
                                        5.

Guler:2024:FRC

[2840] Nesrin Güler and Melek Eris Büyükkaya. Further remarks on constrained overparameterized linear models. Statistical Papers, 65(2):975–988, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01426-
z.

Yang:2024:CGS

[2841] Fang Yang, Liangliang Zhang, Jingyi Zheng, and Xuan Cao. Consistent group selection using nonlocal priors in regression. Statistical Papers, 65 (2):989–1019, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01441-0.

Cuparic:2024:IAM

[2842] Marija Cuparić and Bojana Milosević. To impute or to adapt? Model specification tests’ perspective. Statistical Papers, 65(2): 1021–1039, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01421-4.

Pumi:2024:NCB

[2843] Guilherme Pumi, Taiane S. Prass, and Sílvia R. C. Lopes. A novel copulabased approach for parametric estimation of univariate time series through its covariance decay. Statistical Papers, 65(2):1041–1063, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01418-z.

Karimi:2024:HMC

[2844] Omid Karimi. A Hamiltonian Monte Carlo EM algorithm for generalized linear mixed models with spatial skew latent variables. Statistical Papers, 65 (2):1065–1084, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01419-y.

Li:2024:SDH

[2845] Tizheng Li, Yuping Wang, and Ke Fang. A semiparametric dynamic higher-order spatial autoregressive model. Statistical Papers, 65 (2):1085–1123, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01489-y.

Dimitriadis:2024:OPI

[2846] Timo Dimitriadis, Tobias Fissler, and Johanna Ziegel. Osband’s principle for identification functions. Statistical Papers, 65(2): 1125–1132, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01428-x.

Arellano-Valle:2024:CSP

[2847] Reinaldo B. Arellano-Valle and Adelchi Azzalini. Correction to: Some properties of the unified skew-normal distribution. Statistical Papers, 65 (2):1133, April 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.

1007/s00362-023-01412-5.  See

[2655].

Sun:2024:NLC

[2848] Yang Sun and Xiangzhong Fang. A new L_{2} calibration procedure of computer models based on the smoothing spline ANOVA. Statistical Papers, 65 (4):1901–1926, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01478-1.

DeSantis:2024:LDP

[2849] Fulvio De Santis and Stefania Gubbiotti. On the limit distribution of the power function induced by a design prior. Statistical Papers, 65 (4):1927–1945, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01462-9.

Garg:2024:USE

[2850] Naresh Garg and Neeraj Misra.

A unified study for estimation of order restricted parameters of a general bivariate model under the generalized Pitman nearness criterion. Statistical Papers, 65(4): 1947–1983, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01464-7.

Dette:2024:PRM

[2851] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. Prediction in regression models with continuous observations. Statistical Papers, 65 (4):1985–2009, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01465-6.

Rankin:2024:BMO

[2852] Isaac Rankin and Julie Zhou. Bayesian and maximin A-optimal designs for spline regression models with unknown knots. Statistical Papers, 65 (4):2011–2032, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01469-2.

Lemyre:2024:KTB

[2853] Félix Camirand Lemyre and Jean-François Quessy. Kendall’s tau-based inference for gradually changing dependence structures. Statistical Papers, 65 (4):2033–2075, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01471-8.

Liu:2024:SEL

[2854] Pangpang Liu and Yichuan Zhao. Smoothed empirical likelihood for the difference of two quantiles with the paired sample. Statistical Papers, 65 (4):2077–2108, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01476-3.

Song:2024:MFC

[2855] Wookyeong Song, Hee-Seok Oh, Ying Kuen Cheung, and Yaeji Lim. Multi-feature clustering of step data using multivariate functional principal component analysis. Statistical Papers, 65(4):2109–2134, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01467-4.

DAlberto:2024:IRT

[2856] Riccardo D’Alberto and Meri Raggi.

Integrating rather than collecting: statistical matching in the data flood era. Statistical Papers, 65 (4):2135–2163, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01468-3.

Ye:2024:PPH

[2857] Xin Ye, Baihua He, Yanyan Liu, and Shuangge Ma. Privacy-preserving and homogeneity-pursuit integrative analysis for high-dimensional censored data. Statistical Papers, 65 (4):2165–2190, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01470-9.

Magalhaes:2024:BCZ

[2858] Tiago M. Magalhães, Gustavo H. A.

Pereira, Denise A. Botter, and Mônica C. Sandoval. Bartlett corrections for zero-adjusted generalized linear models. Statistical Papers, 65 (4):2191–2209, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01477-2. See cor-

rection [2996].

Zhang:2024:UMP

[2859] Jin Zhang. Uniformly most powerful tests under weak restrictions. Statistical Papers, 65(4):2211–2220, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01479-
                                        0.

Ren:2024:ROS

[2860] Min Ren, Shengli Zhao, Mingqiu Wang, and Xinbei Zhu. Robust optimal subsampling based on weighted asymmetric least squares. Statistical Papers, 65(4):2221–2251, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01480-7.

Lu:2024:GEV

[2861] Fang Lu, Guoliang Tian, and Jing Yang. GMM estimation and variable selection of partially linear additive spatial autoregressive model. Statistical Papers, 65(4):2253–2288, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01481-
                                        6.

Bieniek:2024:CFB

[2862] Mariusz Bieniek and Tomasz Rychlik. Conditions for finiteness and bounds on moments of generalized order statistics. Statistical Papers, 65 (4):2289–2312, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01482-5.

Dey:2024:BGF

[2863] Monitirtha Dey and Subir Kumar Bhandari. Bounds on generalized family-wise error rates for normal distributions. Statistical Papers, 65 (4):2313–2326, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01487-0.

Duker:2024:DMC

[2864] Marie-Christine Düker, Seok-Oh Jeong, Taewook Lee, and Changryong Baek. Detection of multiple change-points in high-dimensional panel data with cross-sectional and temporal dependence. Statistical Papers, 65(4): 2327–2359, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01484-3.

Khademnoe:2024:VBH

[2865] Omid Khademnoe and S. Mohammad E. Hosseini-Nasab. On the validity of the bootstrap hypothesis testing in functional linear regression. Statistical Papers, 65(4): 2361–2396, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01488-z.

Xi:2024:CPV

[2866] Daiqing Xi and Tianxiao Pang. Change point in variance of fractionally integrated noise. Statistical Papers, 65 (4):2397–2439, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01490-5.

Wei:2024:LSE

[2867] Chao Wei. Least squares estimation for a class of uncertain Vasicek model and its application to interest rates. Statistical Papers, 65 (4):2441–2459, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01494-1.

Zhao:2024:VSP

[2868] Bo Zhao, Shuying Wang, and Chunjie Wang. Variable selection in proportional odds model with informatively interval-censored data. Statistical Papers, 65(4): 2461–2488, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01486-1.

Farne:2024:RCO

[2869] Matteo Farnè and Angelos Vouldis. ROBOUT: a conditional outlier detection methodology for high-dimensional data. Statistical Papers, 65(4): 2489–2525, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01492-3.

Beran:2024:SDZ

[2870] Jan Beran and Frieder Droullier. On strongly dependent zero-inflated INAR(1) processes. Statistical Papers, 65(4):2527–2553, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01496-z.

Lanconelli:2024:MLT

[2871] Alberto Lanconelli and Christopher S. A. Lauria. Maximum likelihood with a time varying parameter. Statistical Papers, 65(4): 2555–2566, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01497-y.

Xu:2024:QRV

[2872] Hong-Xia Xu, Guo-Liang Fan, and Han-Ying Liang. Quantile regression for varying-coefficient partially nonlinear models with randomly truncated data. Statistical Papers, 65 (4):2567–2604, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01498-x.

Liu:2024:PHN

[2873] Shuangzhe Liu, Götz Trenkler, Tõnu Kollo, Dietrich von Rosen, and Oskar Maria Baksalary. Professor Heinz Neudecker and matrix differential calculus. Statistical Papers, 65 (4):2605–2639, June 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01499-w.

Qasim:2024:WAL

[2874] Muhammad Qasim. A weighted average limited information maximum likelihood estimator. Statistical Papers, 65(5):2641–2666, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01485-2.

Jourdan:2024:SFD

[2875] Astrid Jourdan. Space-filling designs with a Dirichlet distribution for mixture experiments. Statistical Papers, 65(5):2667–2686, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01493-2.

Chen:2024:TCB

[2876] Yongshuai Chen, Wenwen Guo, and Hengjian Cui. On the test of covariance between two high-dimensional random vectors. Statistical Papers, 65 (5):2687–2717, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01500-6.

Yuan:2024:FCP

[2877] Panxu Yuan, Yinfei Kong, and Gaorong Li. FDR control and power analysis for high-dimensional logistic regression via StabKoff. Statistical Papers, 65(5):2719–2749, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01501-
                                        5.

Liu:2024:PLI

[2878] Yang Liu, Rong Kuang, and Guanfu Liu. Penalized likelihood inference for the finite mixture of Poisson distributions from capture-recapture data. Statistical Papers, 65(5): 2751–2773, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01503-3.

Ewnetu:2024:TPD

[2879] Worku Biyadgie Ewnetu, Irène Gijbels, and Anneleen Verhasselt. Two-piece distribution based semi-parametric quantile regression for right censored data. Statistical Papers, 65 (5):2775–2810, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01475-4.

Freise:2024:SAS

[2880] Fritjof Freise, Norbert Gaffke, and Rainer Schwabe. A p-step-ahead sequential adaptive algorithm for D-optimal nonlinear regression design. Statistical Papers, 65(5): 2811–2834, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01502-4.

Liu:2024:ACI

[2881] Xu-Qing Liu, Xiao-Cai Wang, Li Tao, Feng-Xian An, and Gui-Ren Jiang. Alleviating conditional independence assumption of naive Bayes. Statistical Papers, 65(5):2835–2863, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01474-

Reisen:2024:DRF

[2882] Valdério Anselmo Reisen, Céline Lévy-Leduc, Edson Zambon Monte, and Pascal Bondon. A dimension reduction factor approach for multivariate time series with long-memory: a robust alternative method. Statistical Papers, 65(5):2865–2886, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01504-2.

Fotopoulos:2024:APC

[2883] Stergios B. Fotopoulos, Abhishek Kaul, Vasileios Pavlopoulos, and Venkata K. Jandhyala. Adaptive parametric change point inference under covariance structure changes. Statistical Papers, 65(5): 2887–2913, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01495-0.

Wang:2024:RAC

[2884] Weiwei Wang, Zhiyang Cui, Ruijie Chen, Yijun Wang, and Xiaobing Zhao. Regression analysis of clustered panel count data with additive mean models. Statistical Papers, 65 (5):2915–2936, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01511-3.

Coupek:2024:FAG

[2885] Petr Čoupek, Viktor Dolník, Zdeněk Hlávka, and Daniel Hlubinka. Fourier approach to goodness-of-fit tests for Gaussian random processes. Statistical Papers, 65(5):2937–2972, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01510-
4.

Nasir:2024:ASB

[2886] Muhammad Jaffri Mohd Nasir, Ramzan Nazim Khan, Gopalan Nair, and Darfiana Nur. Active-set based block coordinate descent algorithm in group LASSO for self-exciting threshold autoregressive model. Statistical Papers, 65 (5):2973–3006, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01472-7.

Virta:2024:RSD

[2887] Joni Virta, Niko Lietzén, and Henri Nyberg. Robust signal dimension estimation via SURE. Statistical Papers, 65(5):3007–3038, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01512-2.

Satter:2024:ELI

[2888] Faysal Satter, Yichuan Zhao, and Ni Li.

Empirical likelihood inference for the panel count data with informative observation process. Statistical Papers, 65(5):3039–3061, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01506-0.

Liu:2024:APR

[2889] Huilan Liu, Xiawei Zhang, Huaiqing Hu, and Junjie Ma. Analysis of the positive response data with the varying coefficient partially nonlinear multiplicative model. Statistical Papers, 65(5):3063–3092, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01516-y.

Beccarini:2024:TOV

[2890] Andrea Beccarini. Testing omitted variables in VARs. Statistical Papers, 65(5):3093–3109, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01513-1.

Haye:2024:ICT

[2891] Mohamedou Ould Haye, Anne Philippe, and Caroline Robet. Inference for continuous-time long memory randomly sampled processes. Statistical Papers, 65(5):3111–3134, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01515-
                                        z.

Kotb:2024:EER

[2892] Mohammed S. Kotb and Huda M. Alomari. Estimating the entropy of a Rayleigh model under progressive firstfailure censoring. Statistical Papers, 65(5):3135–3154, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01508-y.

Wiemann:2024:USF

[2893] Paul F. V. Wiemann, Thomas Kneib, and Julien Hambuckers. Using the softplus function to construct alternative link functions in generalized linear models and beyond. Statistical Papers, 65(5):3155–3180, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01509-x.

Liu:2024:LOD

[2894] Chang-Yu Liu, Xin Liu, and Rong-Xian Yue. Locally optimal designs for comparing curves in generalized linear models. Statistical Papers, 65 (5):3181–3201, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01514-0.

Lin:2024:IPE

[2895] Yucong Lin, Jinhua Su, Yang Liu, Jue Hou, and Feifei Wang. Implicit profiling estimation for semiparametric models with bundled parameters. Statistical Papers, 65(5): 3203–3234, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01519-9.

Du:2024:SEG

[2896] Jierui Du and Xia Cui. Semiparametric estimation in generalized additive partial linear models with nonignorable nonresponse data. Statistical Papers, 65(5):3235–3259, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01522-0.

Zheng:2024:ASF

[2897] Linjuan Zheng, Beiting Liang, and Guochang Wang. Adaptive slicing for functional slice inverse regression. Statistical Papers, 65(5): 3261–3284, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01518-w.

Jhan:2024:ODB

[2898] Yan ni Jhan, Wan cen Li, Shin hui Ruan, Jia jyun Sie, and Iebin Lian. Optimal dichotomization of bimodal Gaussian mixtures. Statistical Papers, 65(5):3285–3301, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01521-1.

Omer:2024:IBR

[2899] Talha Omer, Kristofer Månsson, Pär Sjölander, and B. M. Golam Kibria. Improved Breitung and Roling estimator for mixed-frequency models with application to forecasting inflation rates. Statistical Papers, 65 (5):3303–3325, July 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01520-2.

Li:2024:SAC

[2900] Weirong Li and Wensheng Zhu. Subgroup analysis with concave pairwise fusion penalty for ordinal response. Statistical Papers, 65(6):

3327–3355, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01526-w.

Zhou:2024:SCT

[2901] Jinyu Zhou, Jigao Yan, and Dongya Cheng. Strong consistency of tail value-at-risk estimator and corresponding general results under widely orthant dependent samples. Statistical Papers, 65(6):3357–3394, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-023-01525-
x.

Benhenni:2024:ECE

[2902] Karim Benhenni, Ali Hajj Hassan, and Yingcai Su. The effect of correlated errors on the performance of local linear estimation of regression function based on random functional design. Statistical Papers, 65 (6):3395–3423, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01523-z.

Ahmad:2024:FII

[2903] Rauf Ahmad, Per Johansson, and Mårten Schultzberg. Is Fisher inference inferior to Neyman inference for policy analysis? Statistical Papers, 65 (6):3425–3445, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01528-2.

Ebner:2024:UAG

[2904] Bruno Ebner, Norbert Henze, and Simos Meintanis. A unified approach to goodness-of-fit testing for spherical and hyperspherical data. Statistical Papers, 65(6): 3447–3475, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01529-1.

Cao:2024:SIT

[2905] Mingxiang Cao, Ziyang Cheng, Kai Xu, and Daojiang He. A scale-invariant test for linear hypothesis of means in high dimensions. Statistical Papers, 65 (6):3477–3497, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01530-8.

Liang:2024:HTI

[2906] Shuyi Liang, Kai-Tai Fang, Xin-Wei Huang, Yijing Xin, and Chang-Xing Ma. Homogeneity tests and interval estimations of risk differences for stratified bilateral and unilateral correlated data. Statistical Papers, 65 (6):3499–3543, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01532-6.

Li:2024:MWL

[2907] Fengying Li, Yuqiang Li, and Xianyi Wu. Minimax weight learning for absorbing MDPs. Statistical Papers, 65 (6):3545–3582, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01491-4.

Parchami:2024:SSL

[2908] Abbas Parchami, Przemyslaw Grzegorzewski, and Maciej Romaniuk. Statistical simulations with LR random fuzzy numbers. Statistical Papers, 65 (6):3583–3600, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01533-5.

Li:2024:SON

[2909] Zhaoyang Li and Yuehan Yang. A semi-orthogonal nonnegative matrix tri-factorization algorithm for overlapping community detection. Statistical Papers, 65(6):3601–3619, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01537-
1.

Xu:2024:RMR

[2910] Long-Hao Xu, Yinan Li, and Kai-Tai Fang. The resampling method via representative points. Statistical Papers, 65(6):3621–3649, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01536-
2.

DAngelo:2024:MCF

[2911] Nicoletta D’Angelo and Giada Adelfio.

Minimum contrast for the firstorder intensity estimation of spatial and spatio-temporal point processes. Statistical Papers, 65(6):

3651–3679, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01541-5.

Liebscher:2024:FCC

[2912] Eckhard Liebscher. Fitting copulas in the case of missing data. Statistical Papers, 65(6):3681–3711, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01535-
                                        3.

Chen:2024:EVS

[2913] Yinjun Chen, Hao Ming, and Hu Yang. Efficient variable selection for highdimensional multiplicative models: a novel LPRE-based approach. Statistical Papers, 65(6):3713–3737, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01545-
                                        1.

Ghezal:2024:EST

[2914] Ahmed Ghezal, Maddalena Cavicchioli, and Imane Zemmouri. On the existence of stationary threshold bilinear processes. Statistical Papers, 65 (6):3739–3767, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01539-z.

Burak:2024:BGL

[2915] Katherine Burak and Adam Kashlak. Bootstrapping generalized linear models to accommodate overdispersed count data. Statistical Papers, 65(6):

3769–3788, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01534-4.

Melilli:2024:CDH

[2916] Eugenio Melilli and Piero Veronese.

Confidence distributions and hypothesis testing. Statistical Papers, 65(6): 3789–3820, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01542-4.

Momoki:2024:HTV

[2917] Koki Momoki and Takuma Yoshida.

Hypothesis testing for varying coefficient models in tail index regression. Statistical Papers, 65(6): 3821–3852, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01538-0.

Alevizakos:2024:EEW

[2918] Vasileios Alevizakos, Arpita Chatterjee, Kashinath Chatterjee, and Christos Koukouvinos. The exponentiated exponentially weighted moving average control chart. Statistical Papers, 65 (6):3853–3891, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01544-2.

Ying:2024:SFS

[2919] Baolong Ying, Qijing Yan, Zehua Chen, and Jinchao Du. A sequential feature selection approach to change point detection in mean-shift change point models. Statistical Papers, 65 (6):3893–3915, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01548-y.

Dudek:2024:SPT

[2920] Dagmara Dudek and Anna Kuczmaszewska. Some practical and theoretical issues related to the quantile estimators. Statistical Papers, 65 (6):3917–3933, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01543-3.

Sun:2024:MVG

[2921] Tianqi Sun, Weiyu Li, and Lu Lin. Matrix-variate generalized linear model with measurement error. Statistical Papers, 65(6):3935–3958, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01540-
6.

Ouimet:2024:DBM

[2922] Frédéric Ouimet. Deficiency bounds for the multivariate inverse hypergeometric distribution. Statistical Papers, 65 (6):3959–3969, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01524-y.

Gross:2024:SAR

[2923] Jürgen Groß and Annette Möller. Some additional remarks on statistical properties of Cohen’s d in the presence of covariates. Statistical Papers, 65 (6):3971–3979, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01527-9.

Curtis:2024:WTM

[2924] David Curtis. Welch’s t test is more sensitive to real world violations of distributional assumptions than Student’s t test but logistic regression is more robust than either. Statistical Papers, 65 (6):3981–3989, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01531-7.

denReijer:2024:HSP

[2925] Ard H. J. den Reijer, Pieter W. Otter, and Jan P. A. M. Jacobs. An heuristic scree plot criterion for the number of factors. Statistical Papers, 65 (6):3991–4000, August 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-023-01517-x.

Baz:2024:FDS

[2926] Juan Baz, Diego García-Zamora, Irene Díaz, Susana Montes, and Luis Martínez. Flexible-dimensional L-statistic for mean estimation of symmetric distributions. Statistical Papers, 65(7):4001–4024, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01547-
z.

Park:2024:HDS

[2927] Hyung Park, Thaddeus Tarpey, Eva Petkova, and R. Todd Ogden. A highdimensional single-index regression for interactions between treatment and covariates. Statistical Papers, 65(7): 4025–4056, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01546-0.

Wang:2024:SDR

[2928] Guochang Wang, Zengyao Wen, Shanming Jia, and Shanshan Liang. Supervised dimension reduction for functional time series. Statistical Papers, 65(7):4057–4077, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-023-01505-
                                        1.

Sharghi:2024:SIM

[2929] Sima Sharghi, Kevin Stoll, and Wei Ning. Statistical inferences for missing response problems based on modified empirical likelihood. Statistical Papers, 65(7):4079–4120, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01553-
                                        1.

Ongaro:2024:TPR

[2930] Andrea Ongaro, Sonia Migliorati, Roberto Ascari, and Enrico Ripamonti. Testing practical relevance of treatment effects. Statistical Papers, 65(7): 4121–4145, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01549-x.

BenSaber:2024:SSL

[2931] Asma Ben Saber and Abderrazek Karoui. On some stable linear functional regression estimators based on random projections. Statistical Papers, 65(7):4147–4178, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01554-
0.

Yu:2024:TFA

[2932] Zhou Yu, Niloufar Dousti Mousavi, and Jie Yang. A trigamma-free approach for computing information matrices related to trigamma function. Statistical Papers, 65(7):4179– 4199, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01552-2.

Wang:2024:EPL

[2933] Ke Wang and Dehui Wang. Estimation for partially linear single-index spatial autoregressive model with covariate measurement errors. Statistical Papers, 65(7):4201–4241, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01551-
3.

Parveen:2024:TSN

[2934] Lisa Parveen, Ruhul Ali Khan, and Murari Mitra. A two sample nonparametric test for variability via empirical likelihood methods. Statistical Papers, 65(7):4243– 4265, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01555-z.

Xu:2024:MRC

[2935] Jianbo Xu. Multiple random change points in survival analysis with applications to clinical trials. Statistical Papers, 65(7):4267– 4298, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-023-01507-z.

Wang:2024:NSL

[2936] Xiaodi Wang and Hengzhen Huang. Nested symmetrical Latin hypercube designs. Statistical Papers, 65(7): 4299–4330, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01556-y.

Longla:2024:NCF

[2937] Martial Longla. New copula families and mixing properties. Statistical Papers, 65(7):4331–4363, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01559-
                                        9.

Sahoo:2024:MSC

[2938] Tanmay Sahoo, Nil Kamal Hazra, and Narayanaswamy Balakrishnan. Multivariate stochastic comparisons of sequential order statistics with non-identical components. Statistical Papers, 65(7):4365–4404, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01558-
w.

Sun:2024:VCB

[2939] Ping Sun, Ze-Chun Hu, and Wei Sun. Variation comparison between infinitely divisible distributions and the normal distribution. Statistical Papers, 65(7):4405– 4429, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01561-1.

Akhtar:2024:AQP

[2940] Nadeem Akhtar, Sajjad Ahmad Khan, Emad A. A. Ismail, Fuad A. Awwad, Akbar Ali Khan, Taza Gul, and Haifa Alqahtani. Analyzing quantitative performance: Bayesian estimation of 3-component mixture geometric distributions based on Kumaraswamy prior. Statistical Papers, 65(7):4431– 4451, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01562-0.

Ip:2024:MDM

[2941] Ryan H. L. Ip and K. Y. K. Wu.

A mixture distribution for modelling bivariate ordinal data. Statistical Papers, 65(7):4453–4488, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01560-
                                        2.

Wu:2024:NCP

[2942] Guangyu Wu and Anders Lindquist. A non-classical parameterization for density estimation using sample moments. Statistical Papers, 65(7): 4489–4513, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01563-z.

Dhull:2024:GID

[2943] Monika S. Dhull and Arun Kumar. Geometric infinitely divisible autoregressive models. Statistical Papers, 65(7): 4515–4536, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01564-y.

Filipiak:2024:CST

[2944] Katarzyna Filipiak and Tõnu Kollo. Covariance structure tests for multivariate t-distribution. Statistical Papers, 65(7):4537–4566, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01569-
                                        7.

Peng:2024:SRS

[2945] Ling Peng, Xiaohui Liu, Xiangyong Tan, Yiweng Zhou, and Shihua Luo. The statistical rate for support matrix machines under low rankness and row (column) sparsity. Statistical Papers, 65(7):4567– 4598, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01570-0.

Mohammadi:2024:WVD

[2946] Morteza Mohammadi and Majid Hashempour. On weighted version of dynamic cumulative residual inaccuracy measure based on extropy. Statistical Papers, 65(7):4599– 4629, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01568-8.

Zhu:2024:DRB

[2947] Xuehu Zhu, Rongzhu Zhao, Dan Zeng, Qian Zhao, and Jun Zhang. Dimension reduction-based adaptiveto-model semi-supervised classification. Statistical Papers, 65(7):4631– 4675, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01578-6.

Zhou:2024:IME

[2948] Qian M. Zhou. Information matrix equivalence in the presence of censoring: a goodness-of-fit test for semiparametric copula models with multivariate survival data. Statistical Papers, 65(7): 4677–4713, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01566-w.

Yin:2024:HIH

[2949] Chuancun Yin, Jing Yao, and Yang Yang. Hessian and increasing-Hessian orderings of multivariate skewelliptical random vectors with applications in actuarial science. Statistical Papers, 65(7):4715–4744, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01580-
y.

Araya:2024:LSE

[2950] Héctor Araya, Soledad Torres, and Ciprian A. Tudor. Least squares estimation for the Ornstein–Uhlenbeck process with small Hermite noise. Statistical Papers, 65(7):4745–4766, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https://link.

springer.com/article/10.1007/s00362-
024-01579-5.

Wegner:2024:RCP

[2951] Lea Wegner and Martin Wendler. Robust change-point detection for functional time series based on Ustatistics and dependent wild bootstrap. Statistical Papers, 65(7):4767– 4810, September 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01577-7.

Belalia:2024:GSM

[2952] Mohamed Belalia and Jean-François Quessy. Generalized simulated method-of-moments estimators for multivariate copulas. Statistical Papers, 65(8):4811–4841, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/

article/10.1007/s00362-024-01574-
w.

Sang:2024:ANM

[2953] Yongli Sang and Xin Dang. Asymptotic normality of a modified estimator of Gini distance correlation. Statistical Papers, 65(8): 4843–4860, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01575-9.

Novoselsky:2024:DBT

[2954] Alexander Novoselsky and Eugene Kagan. A distance based two-sample test of means difference for multivariate datasets. Statistical Papers, 65(8): 4861–4874, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01576-8.

Garcin:2024:NET

[2955] Matthieu Garcin and Maxime L. D.

Nicolas. Nonparametric estimator of the tail dependence coefficient: balancing bias and variance. Statistical Papers, 65(8): 4875–4913, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01582-w.

Li:2024:EFE

[2956] Bogui Li, Jianbao Chen, and Hao Chen. Estimation of fixed effects semiparametric single-index panel model with spatio-temporal correlated errors. Statistical Papers, 65(8): 4915–4953, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01584-8.

Monfared:2024:EDC

[2957] Mohammad Esmail Dehghan Monfared. Exact distribution of changepoint MLE for a multivariate normal sequence. Statistical Papers, 65(8): 4955–4970, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01572-y.

Hayashi:2024:NID

[2958] Kenichi Hayashi and Shinto Eguchi. A new integrated discrimination improvement index via odds. Statistical Papers, 65(8):4971–4990, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01585-
                                        7.

Mao:2024:SAH

[2959] Junfan Mao, Zhigen Gao, Bing-Yi Jing, and Jianhua Guo. On the statistical analysis of high-dimensional factor models. Statistical Papers, 65(8): 4991–5019, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01557-x.

Fathi:2024:IWI

[2960] Abdullah Fathi, Al-Wageh A. Farghal, and Ahmed A. Soliman. Inference on Weibull inverted exponential distribution under progressive first-failure censoring with constant-stress partially accelerated life test. Statistical Papers, 65 (8):5021–5053, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01583-9.

Ochieng:2024:MTI

[2961] Daniel Ochieng. Multiple testing of interval composite null hypotheses using randomized p-values. Statistical Papers, 65(8):5055–5076, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01591-
9.

Jana:2024:EMS

[2962] Nabakumar Jana and Samadrita Bera. Estimation of multicomponent system reliability for inverse Weibull distribution using survival signature. Statistical Papers, 65(8): 5077–5108, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01588-4.

Tian:2024:PQM

[2963] Ruiqin Tian, Miaojie Xia, and Dengke Xu. Profile quasi-maximum likelihood estimation for semiparametric varying-coefficient spatial autoregressive panel models with fixed effects. Statistical Papers, 65(8): 5109–5143, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01586-6.

Chou-Chen:2024:APL

[2964] Shu Wei Chou-Chen, Rodrigo A. Oliveira, Irina Raicher, and Gilberto A. Paula. Additive partial linear models with autoregressive symmetric errors and its application to the hospitalizations for respiratory diseases. Statistical Papers, 65(8): 5145–5166, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01590-w.

Berrendero:2024:FRM

[2965] José R. Berrendero, Alejandro Cholaquidis, and Antonio Cuevas. On the functional regression model and its finitedimensional approximations. Statistical Papers, 65(8):5167–5201, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01567-
                                        9.

Modarres:2024:HTT

[2966] Reza Modarres. Hotelling T^{2} test in high dimensions with application to Wilks outlier method. Statistical Papers, 65(8):5203–5218, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01587-
                                        5.

Yoo:2024:GCB

[2967] Na Young Yoo and Ji Hwan Cha. General classes of bivariate distributions for modeling data with common observations. Statistical Papers, 65(8): 5219–5238, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01589-3.

Kozan:2024:ESB

[2968] Agah Kozan, Burak Uyar, and Halil Tanil. Exceedance statistics based on bottom-k-lists. Statistical Papers, 65 (8):5239–5251, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01581-x. See cor-

rection [3069].

Bourguignon:2024:CIS

[2969] Marcelo Bourguignon and Célestin C.

Kokonendji. Coherent indexes for shifted count and semicontinuous models. Statistical Papers, 65(8): 5253–5271, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01598-2.

Alhakim:2024:HMQ

[2970] Abbas Alhakim. Hadamard matrices, quaternions, and the Pearson chisquare statistic. Statistical Papers, 65 (8):5273–5291, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01602-9.

Saleh:2024:RBE

[2971] Asma Saleh. Reduced bias estimation of the log odds ratio. Statistical Papers, 65(8):5293–5331, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01593-
                                        7.

Shan:2024:CIO

[2972] Guogen Shan, Xinlin Lu, Yahui Zhang, and Samuel S. Wu. Confidence intervals for overall response rate difference in the sequential parallel comparison design. Statistical Papers, 65 (8):5333–5349, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01606-5.

Skarupski:2024:CBC

[2973] Marek Skarupski and Qinhao Wu. Confidence bounds for compound Poisson process. Statistical Papers, 65(8): 5351–5377, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01604-7.

Haq:2024:CNE

[2974] Abdul Haq and William H. Woodall. A critical note on the exponentiated EWMA chart. Statistical Papers, 65 (8):5379–5387, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01601-w.

Bickel:2024:BFI

[2975] David R. Bickel. Bayesian and frequentist inference derived from the maximum entropy principle with applications to propagating uncertainty about statistical methods. Statistical Papers, 65(8):5389–5407, October 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01597-
3.

Kreer:2024:MLE

[2976] Markus Kreer, Ayşe Kızılersü, Jake Guscott, Lukas Christopher Schmitz, and Anthony W. Thomas. Maximum likelihood estimation for lefttruncated log-logistic distributions with a given truncation point. Statistical Papers, 65(9):5409–5445, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01603-
8.

Josse:2024:CSL

[2977] Julie Josse, Jacob M. Chen, Nicolas Prost, Gaël Varoquaux, and Erwan Scornet. On the consistency of supervised learning with missing values. Statistical Papers, 65(9): 5447–5479, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01550-4.

Tian:2024:TTV

[2978] Lingling Tian, Yunan Su, and Chuanhua Wei. Tests for timevarying coefficient spatial autoregressive panel data model with fixed effects. Statistical Papers, 65(9): 5481–5501, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01607-4.

Zheng:2024:NSO

[2979] Chunwei Zheng, Wenlong Li, and Jian-Feng Yang. Nested strong orthogonal arrays. Statistical Papers, 65(9): 5503–5528, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01609-2.

Tarima:2024:CSA

[2980] Sergey Tarima and Nancy Flournoy. The cost of sequential adaptation and the lower bound for mean squared error. Statistical Papers, 65(9):5529– 5553, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01565-x.

Mariani:2024:DPR

[2981] Francesco Mariani, Fulvio De Santis, and Stefania Gubbiotti. The distribution of power-related random variables (and their use in clinical trials). Statistical Papers, 65(9): 5555–5574, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01599-1.

Liu:2024:CED

[2982] Xirui Liu, Mixia Wu, and Liwen Xu. Communication-efficient distributed EM algorithm. Statistical Papers, 65(9):5575–5592, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01594-
                                        6.

Chen:2024:ETE

[2983] Baolin Chen, Shanshan Song, and Yong Zhou. Estimation and testing of expectile regression with efficient subsampling for massive data. Statistical Papers, 65(9):5593– 5613, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01571-z.

Mojirsheibani:2024:SOK

[2984] Majid Mojirsheibani. Strong optimality of kernel functional regression in L^{p} norms with partial response variables and applications. Statistical Papers, 65(9): 5615–5648, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01611-8.

Garay:2024:BAL

[2985] Aldo M. Garay, Francyelle L. Medina, Suelem Torres de Freitas, and Víctor H. Lachos. Bayesian analysis of linear regression models with autoregressive symmetrical errors and incomplete data. Statistical Papers, 65(9): 5649–5690, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01612-7.

Dey:2024:LBS

[2986] Monitirtha Dey. On limiting behaviors of stepwise multiple testing procedures. Statistical Papers, 65(9): 5691–5717, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01613-6.

Garg:2024:IEL

[2987] Naresh Garg, Lakshmi Kanta Patra, and Neeraj Misra. On improved estimation of the larger location parameter. Statistical Papers, 65(9): 5719–5752, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01614-5.

Hu:2024:LSE

[2988] Lixia Hu, Baolin Chen, and Jinhong You. Locally sparse estimator for functional linear panel models with fixed effects. Statistical Papers, 65(9): 5753–5773, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01595-5.

Xiao:2024:TAA

[2989] Yao Xiao, Hong Qin, Kashinath Chatterjee, and Na Zou. Theory and application of absolute discrepancy in experimental designs. Statistical Papers, 65(9):5775– 5795, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01617-2.

Urso:2024:MSM

[2990] Furio Urso, Antonino Abbruzzo, Marcello Chiodi, and Maria Francesca Cracolici. Model selection for mixture hidden Markov models: an application to clickstream data. Statistical Papers, 65(9):5797–5834, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01608-
3.

vanWieringen:2024:RTC

[2991] Wessel N. van Wieringen and Gwenaël G. R. Leday. Ridge-type covariance and precision matrix estimators of the multivariate normal distribution. Statistical Papers, 65(9): 5835–5849, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01610-9.

Pareto:2024:CCB

[2992] Adriano Pareto. On a correlation coefficient based on the L_{1}- norm. Statistical Papers, 65(9): 5851–5871, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01616-3.

Fan:2024:NIV

[2993] Yixuan Fan, Jianhua Cheng, and Dehui Wang. A new integer-valued threshold autoregressive process based on modified negative binomial operator driven by explanatory variables. Statistical Papers, 65(9):5873– 5901, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01605-6.

Wu:2024:MFF

[2994] Jiujing Wu and Hengjian Cui. Modelfree feature screening based on Hellinger distance for ultrahigh dimensional data. Statistical Papers, 65(9): 5903–5930, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01615-4.

Scarlatti:2024:ECS

[2995] S. Scarlatti. Enhanced Cauchy– Schwarz inequality and some of its statistical applications. Statistical Papers, 65(9):5931–5940, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00362-024-01600-
                                        x.

Magalhaes:2024:CBC

[2996] Tiago M. Magalhães, Gustavo H. A. Pereira, Denise A. Botter, and Mônica C. Sandoval. Correction to: Bartlett corrections for zeroadjusted generalized linear models. Statistical Papers, 65(9): 5941, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01592-8.  See

[2858].

Veraverbeke:2024:BEC

[2997] Noël Veraverbeke. Bernstein estimator for conditional copulas. Statistical Papers, 65(9):5943–5954, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01573-
x.

Lukic:2024:CPA

[2998] Zikica Lukić and Bojana Milosević.

Change-point analysis for matrix data: the empirical Hankel transform approach. Statistical Papers, 65(9): 5955–5980, December 2024. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01596-4.

Jin:2025:CQR

[2999] Jun Jin, Chenyan Hao, and Yewen Chen. Composite quantile regression for a distributed system with non-randomly distributed data. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01625-
2.

Pal:2025:RDM

[3000] Samyajoy Pal and Christian Heumann.

Revisiting Dirichlet Mixture Model: unraveling deeper insights and practical applications. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01627-0.

Xia:2025:CCI

[3001] Liqi Xia, Ruiyuan Cao, Jiang Du, and Jun Dai. Consistent complete independence test in high dimensions based on Chatterjee correlation coefficient. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01618-1.

Li:2025:KDR

[3002] Facheng Li and Huilan Liu. Kernel density regression in the additive model: a B-spline approach. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01621-6.

Xu:2025:TTG

[3003] Tianming Xu, Dong Jiang, Yuesong Wei, and Chong Wang. A test for trend gradual changes in heavy tailed AR (p) sequences. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01626-1.

Basu:2025:DPD

[3004] Suparna Basu and Hon Keung Tony Ng. A density power divergence measure to discriminate between generalized exponential and Weibull distributions. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01637-y.

Lou:2025:JET

[3005] Dan Lou and Yuehan Yang. Joint estimation of transfer learning on time series data. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01629-y.

Li:2025:THD

[3006] Jianghao Li, Shizhe Hong, Zhenzhen Niu, and Zhidong Bai. Test for high-dimensional linear hypothesis of mean vectors via random integration. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01624-3.

Zhao:2025:EPD

[3007] Yan-Yong Zhao, Ling-Ling Ge, and Yuan Liu. Estimation of panel data partially linear time-varying coefficient models with cross-sectional spatial autoregressive errors. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01620-7.

Kong:2025:CPR

[3008] Jingsen Kong, Yiming Liu, Guangren Yang, and Wang Zhou. Conformal prediction for robust deep nonparametric regression. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01631-4.

Haines:2025:ODS

[3009] Linda M. Haines. Optimal designs for spherical harmonic regression. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01630-5.

Wong:2025:INC

[3010] A. Wong and X. Shi. Inference for the normal coefficient of variation: an approximate marginal likelihood approach. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01622-5.

Mohaoui:2025:DDE

[3011] Nezha Mohaoui, Hamid Mraoui, and Abdelilah Monir. Distribution and density estimation based on variation-diminishing spline approximation. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01619-0.

Harrouche:2025:BPR

[3012] Lyasmine Harrouche, Hocine Fellag, and Lynda Atil. Bayesian prior robustness using general ϕdivergence measure. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01628-z.

Liu:2025:BQR

[3013] Changsheng Liu, Hanying Liang, and Yongmei Li. Bayesian quantile regression for partially linear single-index model with longitudinal data. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01633-2.

Suresh:2025:JEL

[3014] Saparya Suresh and Sudheesh K. Kattumannil. Jackknife empirical likelihood ratio test for testing the equality of semivariance. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01636-z.

Han:2025:BEP

[3015] Ming Han. Bayesian estimation and posterior risk under the generalized weighted squared error loss function and its applications. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01643-0.

Zhang:2025:SSL

[3016] Menghua Zhang, Mengjiao Peng, and Yong Zhou. Semi-supervised learning for various comparison functions across two populations. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01632-3.

Ahanda:2025:CND

[3017] Benoit Ahanda, Leif Ellingson, and Daniel E. Osborne. The Cholesky normal distribution for SPD matrices and inference for the mean. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01640-3.

Brunner:2025:URB

[3018] Edgar Brunner and Frank Konietschke. An unbiased rank-based estimator of the Mann–Whitney variance including the case of ties. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01635-0.

Amato:2025:FTS

[3019] Umberto Amato, Anestis Antoniadis, Italia De Feis, and Irène Gijbels. Functional time series forecasting: a systematic review. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01645-y.

Kangro:2025:PLA

[3020] Raul Kangro, Kristi Kuljus, and Jüri Lember. Pseudo-likelihood approach for parameter estimation in univariate normal mixture models. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01642-1.

Zhang:2025:UEG

[3021] Xianhua Zhang, Lu Lin, and Qihua Wang. Updatable estimation in generalized linear models with missing data. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01623-4.

Fonseca:2025:OPQ

[3022] Giovanni Fonseca, Federica Giummolè, and Paolo Vidoni. Optimal prediction for quantiles and probabilities. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01641-2.

Yang:2025:MRN

[3023] Zhen-Hang Yang and Jing-Feng Tian.

Monotonicity results and new bounds for the Mills ratio. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01646-x.

Mendo:2025:EOL

[3024] Luis Mendo. Estimating odds and log odds with guaranteed accuracy. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01639-w.

Mukasa:2025:EBT

[3025] Hirai Mukasa and Koji Tsukuda. Equality between two general ridge estimators and equivalence of their residual sums of squares. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01644-z.

Mostofi:2025:NBI

[3026] Amin Ghalamfarsa Mostofi and Mahmood Kharrati-Kopaei. Nonparametric Bayesian inferences on the skewed data using a Dirichlet process mixture model. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01658-7.

Sukeda:2025:MIC

[3027] Issey Sukeda and Tomonari Sei. On the minimum information checkerboard copula under fixed Kendall’s τ. Statistical Papers, 66(1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01648-
9.

Dou:2025:EEB

[3028] Xiaoling Dou, Satoshi Kuriki, Gwo Dong Lin, and Donald Richards. EM estimation of the B-spline copula with penalized pseudo-likelihood functions. Statistical Papers, 66 (1):??, January 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01647-w.

Patilea:2025:DMC

[3029] Valentin Patilea and François Portier.

Density model checks via the lackof-fitness. Statistical Papers, 66 (2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01655-w.

Jimenez-Gamero:2025:THL

[3030] M. Dolores Jiménez-Gamero, Marina Valdora, and Daniela Rodríguez. Testing homoscedasticity of a large number of populations. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01650-1.

Neumeyer:2025:TCE

[3031] Natalie Neumeyer and Leonie Selk.

Testing for changes in the error distribution in functional linear models. Statistical Papers, 66 (2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01656-9.

Meintanis:2025:ODP

[3032] Simos G. Meintanis, Joseph Ngatchou-Wandji, and Sárka Hudecová. Omnibus diagnostic procedures for vector multiplicative errors models. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01653-y.

Wang:2025:BID

[3033] Qingrui Wang and Zhao Yao. Bayesian influence diagnostics for a multivariate GARCH model. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01649-8.

Wang:2025:BEL

[3034] Wenshan Wang, Xinyuan Song, Guichen Han, and Kai Yang. Bayesian empirical likelihood inference and order shrinkage for a hysteretic autoregressive model. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-025-01659-0.

Ye:2025:DSC

[3035] Fei Ye, Jingsong Xiao, Weidong Ma, Shiwen Jin, and Ying Yang. Detecting small clusters in the stochastic block model. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-025-01660-7.

Yu:2025:FSF

[3036] Congran Yu and Hengjian Cui. Feature screening via false discovery rate control for linear model with multivariate responses. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01661-6.

Durante:2025:LDB

[3037] F. Durante, C. Ignazzi, and P. Jaworski. The limiting distribution of a bivariate random vector under univariate truncation. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01663-4.

Lopez-Perez:2025:EST

[3038] A. López-Pérez, M. Febrero-Bande, and W. González-Manteiga. Estimation and specification test for diffusion models with stochastic volatility. Statistical Papers, 66 (2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01652-z.

Xia:2025:CEM

[3039] Xiaochao Xia, Sijin He, and Naiwen Pang. Communication-efficient model averaging prediction for massive data with asymptotic optimality. Statistical Papers, 66 (2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01664-3.

Gonzalez-Perez:2025:NTF

[3040] Ignacio González-Pérez, María Isabel Borrajo, and Wenceslao González-Manteiga. Nonparametric testing of first-order structure in point processes on linear networks. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01657-8.

Chen:2025:TLS

[3041] Xuan Chen and Yunquan Song. Transfer learning for semiparametric varying coefficient spatial autoregressive models. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01662-5.

Sobh:2025:ABB

[3042] M. E. Sobh, H. M. Barakat, Magdy E. El-Adll, and Amany E. Aly. Asymptotic behavior of bootstrapped extreme order statistics under unknown power normalizing constants. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01665-2.

Wang:2025:ECC

[3043] Miaomiao Wang, Shunping Zheng, and Xuejun Wang. Equivalent conditions of complete moment convergence for randomly weighted sums of random variables and some applications with random design. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01666-1.

Forzani:2025:PME

[3044] Liliana Forzani, Daniela Rodriguez, and Mariela Sued. A penalization method to estimate the intrinsic dimensionality of data. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01667-0.

Das:2025:CNT

[3045] Koushik Das and Shyamal Ghosh.

A class of nonparametric tests for DMTTF alternatives based on moment inequality. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01670-5.

Sun:2025:EBP

[3046] Ran Sun, Mohamed Belalia, and Sévérien Nkurunziza. The empirical Bernstein process with application to uniformity testing. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01668-z.

Emami:2025:SPL

[3047] Hadi Emami and Omid Khademnoe.

Shrinkage and pretest Liu estimators in semiparametric linear measurement error models. Statistical Papers, 66(2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01671-4.

Lian:2025:MML

[3048] Chengdi Lian, Camila Borelli Zeller, Ke Yang, and Weihu Cheng. Modified maximum likelihood estimator for censored linear regression model with two-piece generalized t distribution. Statistical Papers, 66 (2):??, February 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-024-01634-1.

Sun:2025:FRE

[3049] Qian Sun, Yang Zhou, and Shouyou Huang. Fast rates of exponential cost function. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-025-01672-3.

Li:2025:GSO

[3050] Pengnan Li, Chunjie Wang, and Bochuan Jiang. Group strong orthogonal arrays and their construction methods. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-025-01674-1.

Li:2025:VSN

[3051] Jie Li and Yunquan Song. Variable selection for nonparametric spatial additive autoregressive model via deep learning. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01669-y.

Meintanis:2025:GFT

[3052] Simos G. Meintanis, Joseph Ngatchou-Wandji, Leonard Santana, and Marius Smuts. Goodness-of-fit tests for discrete response models with covariates. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-024-01654-
x.

Cetinkaya:2025:NEM

[3053] Merve Kandemir Çetinkaya. A new estimator for the multicollinear logistic regression model. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01675-0.

Akhtar:2025:BAH

[3054] Nadeem Akhtar and Muteb Faraj Alharthi. Bayesian analysis of heterogeneous data outliers using a censored mixture model. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-024-01638-x.

Mahdavi:2025:BDM

[3055] Abbas Mahdavi and Javier E. Contreras-Reyes. Bounded data modeling using logit-skew-normal mixtures. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01677-y.

Nam:2025:DRM

[3056] Donghwi Nam, Ja-Yong Koo, and Kwan-Young Bak. Dimensionality reduction in multivariate nonparametric regression via nuclear norm penalization. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01676-z.

Shi:2025:THD

[3057] Hongwei Shi, Weichao Yang, Bowen Sun, and Xu Guo. Tests for high-dimensional partially linear regression models. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01679-w.

Sun:2025:DPF

[3058] Tianqi Sun, Weiyu Li, and Lu Lin. Distributed penalizing function criterion for local polynomial estimation in nonparametric regression with massive data. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01678-x.

Tang:2025:ELN

[3059] Wenjing Tang and Yongsong Qin. Empirical likelihood for nonparametric regression functions under ρ-mixing high-frequency data. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01683-0.

Li:2025:SPL

[3060] Tizheng Li, Lin Li, and Yanhui Li.

Semiparametric partially linear varying coefficient higher-order spatial autoregressive model. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01681-2.

Kawamoto:2025:SCA

[3061] Kohei Kawamoto, Yuichi Goto, and Koji Tsukuda. Spectral clustering algorithm for the Allometric Extension Model. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01680-3.

Bernardi:2025:NTO

[3062] Enrico Bernardi, Alberto Lanconelli, Christopher S. A. Lauria, and Berk Tan Perçin. Non trivial optimal sampling rate for estimating a Lipschitz-continuous function in presence of mean-reverting Ornstein– Uhlenbeck noise. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01685-y.

Amado:2025:REH

[3063] Conceição Amado, Ana M. Bianco, Graciela Boente, and Isabel M. Rodrigues. Robust estimation of heteroscedastic regression models: a brief overview and new proposals. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01686-x.

Bickel:2025:SSB

[3064] David R. Bickel. A small-sample Bayesian information criterion that does not overstate the evidence, with an application to calibrating p-values from likelihood-ratio tests. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01682-1.

Chen:2025:GER

[3065] Jianbao Chen, Bogui Li, and Shuangshuang Li. GMM estimation of random effects semiparametric additive SAR panel model with spatiotemporal correlated errors. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01687-w.

Suzuki:2025:TSM

[3066] Akira Suzuki and Hidetoshi Murakami. The two-sample Mood statistic for clustered data. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01692-z.

Ge:2025:UCC

[3067] Meimei Ge, Yan Wang, Xin Deng, Xuejun Wang, and Aiting Shen. The uniformly complete consistency of generalized edge frequency polygon estimator for asymptotically negatively associated samples and an application. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01690-1.

Jureckova:2025:NTS

[3068] Jana Jurečková, Olcay Arslan, Yeşim Güney, Yetkin Tuaç, Jan Picek, and Martin Schindler. Nonparametric tests for serial independence in linear model against a possible autoregression of error terms. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

//link.springer.com/article/10.
1007/s00362-025-01689-8.

Kozan:2025:CES

[3069] Agah Kozan, Burak Uyar, and Halil Tanil. Correction to: Exceedance statistics based on bottom-k-lists. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL https://link.springer.com/
article/10.1007/s00362-025-01688-

9. See [2968].

Pei:2025:FND

[3070] Jian Pei and Yang Lu. Forecasting natural disaster frequencies using nonstationary count time series models. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-025-01691-0.

Torres-Signes:2025:GFR

[3071] A. Torres-Signes, M. P. Frías, and M. D. Ruiz-Medina. Global Fréchet regression from time correlated bivariate curve data in manifolds. Statistical Papers, 66(3):??, April 2025. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00362-025-01684-z.

Miller:2015:BLT

[3072] Steven J. Miller, editor. Benford’s Law: theory and applications. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). xxvi + 438 pp. LCCN QA273.6 .B46 2015.

                                        URL http://press.princeton.edu/
                                        titles/10527.html.

Bethlehem:2017:PPO

[3073] Jelke G. Bethlehem. Polling Public Opinion. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2017. ISBN 1-138-06655-9, 1-4987-6974-8 (paperback), 1-4987-6975-6 (e-book). xi + 286 pp. LCCN HM1236 .B48 2017.

Weihs:2017:MDA

[3074] Claus Weihs, Dietmar Jannach, Igor Vatolkin, and Günter Rudolph, editors. Music Data Analysis: Foundations and Applications. Chapman et Hall/CRC computer science and data analysis series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2017. ISBN 1-4987-1957-0 (e-book), 1-4987-1956-2 (hardcover), 1-315-35383-0 (ePub ebook), 1-315-33477-1 (Mobipocket ebook), 1-315-37099-9 (e-book). xviii + 675 pp. LCCN ML74 .M877 2017. URL http:

//www.crcnetbase.com/isbn/978-1-
4987-1956-8.

Woodward:2017:ATS

[3075] Wayne A. Woodward, Henry L. Gray, and Alan C. Elliott. Applied time series analysis, with R. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, second edition, 2017. ISBN 1-4987-3422-7 (hardcover), 1-4987-3431-6 (ePub e-book), 1-4987-3427-8 (PDF e-book), 1-4987-3429-4 (VitalBook ebook). xv + 618 pp. LCCN QA280 .W68 2017.